A Complete Bibliography of Publications in Future Internet

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

31 December 2024 Version 1.52

Title word cross-reference

3 [DOB21]. d [KK23a]. K [AAJ^{+}22, SDDB24]. κ [TH21]. μ [TH21]. n [REP22]. n × n [TDT^{+}21]. Q [DWHY24]. t [ZH21].

-Anonymity-Based [AAJ^{+}22]. -Aware [EMHF19]. -Based [ELCS20]. -Identification [SDDB24]. -Regular [KK23a]. -th [REP22].

.eu [GPLK22].

1 [SSM^{+}24]. 1-D [SSM^{+}24]. 1.0 [GM23]. 10 [AMZ^{+}20]. 11 [FC19]. 15 [LJR24]. 17-Generations [IHMG22]. 1800 [DMZ^{+}19]. 19 [ABSG21, AAN^{+}21, AA21a, BCT21, CBRS21, CYC^{+}21, CQWC22, CTM22a, Del21, ESB^{+}20, FFVP21, GNGHEM21, GFPD21, HGF21, Kra20, MPG20, OFV21, RRLT20, RFSOHMSB21, SG21b, TC21, WKD22, YWLY22, YHN21, dCMA21]. 1st [RFSOHMSB21].

2 [CR21, GMLS22, LSSD22]. 2.0 [DPR^{+}21, ERBL12, GM23, NPB12, NB12, PR12, RMB^{+}12, Whe09]. 2020 [CR21]. 2023 [TWW24]. 230 [DMZ^{+}19]. 2D [GNGL23]. 2LCCM [SNZ21].

3 [KHC^{+}23, LRD23, LB23]. 3-D [LB23]. 3.0 [BM23, DV23, FMA^{+}23, FHS^{+}10]. 301 [VBB24b]. 32-Bit [MLPH23]. 395 [LJR24]. 3D [GM20, AWUF19, BZMB23, CRMT23, DWY19b, RES23, SSR^{+}21, VRA^{+}18, WLCC24, YCHG18, ZN22]. 3D-CNN-Based [AWUF19]. 3D-Obstacle [GM20]. 3G [KOJP13]. 3GPP [ACM^{+}19, HAL18, PP22b].

4.0 [AMVM22, BPLBM21, CEZC19, DTL^{+}21, DPR^{+}21, dSFdRP24, JNARV^{+}24, Lel19, LH21b, LTP24, LBBPM21, MRRMC21, NNRR^{+}21, Pil18, VBWW22, ZCdRR^{+}23]. 40 [WZG^{+}18]. 4G [AMZ^{+}20, CTP^{+}22, KDJ24].

5.0 [KHC^{+}24, MAP23]. 5G [AZH22, AMG22, AAA^{+}24a, ACD^{+}22, ASM^{+}24, ASMM22, ABH24, AMSM23, BKL^{+}22, BTBK21, BCM^{+}19a, BAS24, CCD^{+}21, CFP19, CGM^{+}19, DJC^{+}22, DGCK24, FN23, GRM23, GFPVLR^{+}23, GMS^{+}20, HAL18, HGL^{+}23, JLGF19, JNARV^{+}24, KLKP19, KDJ24, KLG^{+}24, KIB^{+}22, KSE^{+}20, LQFI17, LPVM24, LTW21, LCV^{+}17, Mac22, MAW^{+}23, MMF^{+}19, MGV17, MOAI17, MS23, MCH^{+}20, MVMHL24, Mou24, NB24, OGB^{+}24, PSM^{+}18, PP22b, ST22, SBRZ24, SHC22, SOS^{+}23, SMN22, SRSDF23, SAM^{+}18, SP22, VLC^{+}19, WGN23, WWB^{+}20, WUH^{+}23, XDH24, YHSW19, ZOAR24, ZHW23]. 5G-Enabled [GMS^{+}20, WUH^{+}23]. 5G-ICN [ZHW23]. 5G-MEC [WGN23]. 5G-V2X [MCH^{+}20]. 5G/6G [OGB^{+}24]. 5G/B5G [BAS24]. 5G/Satellite [CFP19]. 5GCAR [CGM^{+}19]. 5th [HUM^{+}21].

61499 [DSM21]. 6G [CALHG^{+}24, NA23, AALS21, AAAAM^{+}21, ASD^{+}22, BTS^{+}21, MSAD22, NA23, OSdS24, OGB^{+}24, RNST23, ST22, SAL21, SCC^{+}23, SMN22, SPM^{+}22, VM24b]. 6G-Enabled [NA23]. 6G-Intended [CALHG^{+}24]. 6G-IoV [NA23].

7 [JNARV^{+}24]. 7R [KS14].

802.11 [MSK^{+}18, SE23]. 802.11bc [GCEOBRT23]. 802.16e [SIJA10].

95 [SU17a]. 9th [Bod22].

ABC [MMMH19]. Abdominal [ANAAM21]. ABE [WZX20]. Above [JDH21]. Above-the-Fold [JDH21]. Abstract [GDS23]. Abstracting [VDSK23]. Abstraction [DB23]. Abstractive [BTB22, FTAA21]. Abuse [VYN19]. Abysmal [RKY20]. Academic [GKK^{+}19, GPL24, HT14, PCROB21, PFT^{+}24, UFK^{+}10, VCGOMC^{+}21, YLPW21, RCGSL19]. Accelerated [SPS^{+}21]. Accelerating [CBM^{+}20, YHG^{+}21]. Accelerators [AKM22]. Accent [VLC^{+}19]. Acceptance [AMAA^{+}21, ASG24, FKF24, KA21, KLTC17, NPWC24, RYAA22, SRM20, SIZD12, TN20, ZY20a]. Accepting [KDKG22]. Access [AA21b, AAA^{+}20, AKJB20, BZDP19, CCCPV^{+}24, CWF13, CBKL22, CDLY24, DCG12, DL19, GDARM18, GM16, GRRZ18, HLZ^{+}18, JK24, LWY20, LZY^{+}24, MG10, MAW^{+}19, MGV17, MHQ^{+}23, OF23, OSdS24, PP22b, QSF^{+}17, AJT^{+}23, SHHM21, SZL21, SRSDF23, SSX20, TTKZ19, TKH^{+}11, TRST24, VL24, WZG^{+}18, WLCS17, WZX20, YZP22, ZWL^{+}24]. Accessibility [AF22, KBBH21]. Accountability [BDL22, LB24, RL24, RBSMRN21]. Accounting [DCD24, PLA^{+}24, ZS18]. Accounts [AAAR24, AZV16, SZZL23, ZAV14]. Accuracy [HNHH23, KPH^{+}24, ZSF20]. Accurate [AHH^{+}23, CW23, DAMAS^{+}19]. Achieve [GLRM19, SIJA10, ZZL^{+}23]. Achieving [LB24, URH^{+}23]. Acid [RIS^{+}17]. Acknowledgement [Fut20, Off14, Off15, Off16, Off17, Off18, Off19]. Acknowledgment [Fut21, Off22]. Acoustic [CW23, CBD^{+}22, LLBB18]. Acquired [ZMDCEMI22]. Acquisition [LNT^{+}23, PG14, YWZ^{+}24]. Across [BBH^{+}24, BSEL20, HWZP18, HWCH12, LTM^{+}22, SZFC18, Tre22]. Action [AWUF19, LDZ^{+}23, MKKS18, PDMK19, SOSR^{+}16, SS14, YCHG18]. Action-Research [SOSR^{+}16]. Actions [GLRM19, RK19]. Activation [SDL^{+}19]. Activations [GRB^{+}21]. Active [HAA23, LHW18, LNH23, LLW^{+}24, LL20c, ZLZ^{+}24]. Active-RIS-Assisted [LLW^{+}24]. Activities [ADCM24, CFG20, CEZC19, FOJE19]. Activity [AYH^{+}21, CLTP22, CTM19, DDV22, FM20a, GNV21, GBDDVS23, KKL^{+}20, LSM^{+}12, MMA10, MMMH19, PRDK22, PFT^{+}24, PHG^{+}20, PHGZ20, SFW23, SOA^{+}20, Zaf12, vSR24]. Activity-Based [MMMH19]. Actor [BBBHA24]. Actor-Critic [BBBHA24]. Actuating [ZLS19]. Ad [AASAI22, BRP^{+}13, CGT21, DLK10, FR16, IZK^{+}23, IT24, KH19, MZH22, SS15, TT22, YT18]. Ad-Hoc [KH19]. Adaboost [JK24]. ADAPT [GTWN24]. Adaptation [MCR^{+}23]. Adaptative [VOP22]. Adapted [OLCMdA20, SIZD12]. Adapter [KMS^{+}12]. Adapting [KPB23, Lai24, Sac19, WKF^{+}21]. Adaption [GCEOBRT23]. Adaptive [AAR^{+}19, AD20, ABK22, AS15, AMZ^{+}20, BHC23, CD23, CAL23, DRK^{+}22, GTM22, UGJA18, HF20, HXHP17, KALY17, KHAM24, LCC^{+}24, LSZ21, LZL^{+}23, LXG^{+}23, LW21, MAW^{+}23, NNT22, NRS24, SU17a, SCA^{+}19, SLW17, USS19, VFFHF19, WCM19, WJ21, WLG^{+}22, XCSM18, XY21, YW16, YPMM12, ZSZ18]. Adaptive-Greedy [CD23]. Adaptively [WKW18]. ADASs [LLCL23]. Added [HLLT10]. Adders [FG17]. Addiction [DIG20, RNFS20]. Addictions [RRS20]. Addition [RK19]. Additional [ALST20]. Additive [GDN^{+}23, XLW^{+}24]. Addressing [PA13, RES23, VKK^{+}19, XKKL23].

Adjacency [LYD22].

Adjacency-Information-Entropy-Based [LYD22]. Administration [GTM22, Sta23]. Admitted [PLOT23]. Adolescent [ISG20, SRMJ21]. Adopt [Kha24, MSC19]. Adoption [AA20, GFPD21, KNN^{+}24, LPPG22, LFGR20, MDPHB23, NPWC24, PR12, RHHN21, VLMP21, Yan17, ZKV12]. Ads [PDI21, ASAH^{+}23]. Adult [ADCM24, BBM20]. Adults [HS18b, Pri23, VYN19]. Advanced [DIE^{+}21, EKKS24, EARP^{+}24, DMP^{+}23, HLC24, KGKP20, KAEAB24, KM24, LF24, MAW^{+}19, NSV17, PGSL20, TDM^{+}24, WSN24]. Advancements [KBF^{+}22, PKF24]. Advances [CC24, Fer12, HAL18, KTUF19, LARPHA23, Mal20, MTM22, SHC18, Wan23]. Advancing [AA24b, PBP^{+}24, PLL11, XLW^{+}24]. Adversarial [ALZ^{+}23, AR23, CZZZ22, FAB^{+}22, FLC21, FHXL19, GS23a, GHL^{+}23, IAM24, JGY^{+}24, KRSK24, KL21, OAH^{+}24, QDL22, RD18, URH^{+}23, VR21, VOP22, ZZZ^{+}23]. Advert [MT17a]. Advertisements [MT17b]. Advertising [LOL23, TOLG21, ZY20a]. Aerial [LS21, MAM^{+}21]. Aerosol [IHMG22]. AES [QAQA21]. Aesthetic [LPPG22]. Aesthetics [GPL24]. Affairs [KP15]. Affect [BCT21]. Affected [TC21]. Affecting [GN16, MDPHB23, NPWC24, VLMP21, Yan19]. Affections [FDN^{+}24]. Affective [SAF14]. Affiliation [SS22]. Affiliations [VKK^{+}19]. Africa [APAR22]. African [GS11, KNN^{+}24, Smi13]. after [KHAA21, MPG20]. Again [DGGS19]. Against [FHXL19, ALZ^{+}23, AKDL24, AR23, AV23, CYC^{+}21, ESB^{+}20, GS23b, IAM24, KMK24, MS22, PMP^{+}22b, QDL22, SZZL23]. Age [AARTSBAE24, KKP22b, Lai24, PVM22, Rah22]. Age-of-Information-Aware [PVM22]. Ageing [ACK^{+}16]. Agencies [THL21]. Agency [TTO^{+}19, TL24b]. Agenda [HPC12, RGdSK22, vdSFRF23]. Agent [AAE^{+}22, Cap23, CTM19, DWHY24, DPR^{+}21, DSM21, GB23, HD19, JUN^{+}24, JHBW24, KK21, KBHP24, LZLW22, LFM^{+}19, MMS^{+}19, SV24, Suf23, WLLY24, Zar22, ZZY19]. Agent-Based [DSM21, HD19]. Agent-Oriented [DPR^{+}21]. Agents [CCCA22, CB24, TTN24, WMS23]. Agglomerative [BOB^{+}20]. Agglutinative [PLYD20]. Aggregate [IZK^{+}23]. Aggregated [KZ20]. Aggregation [AAAIA24, KES22, KCT^{+}24, LT24a, SJA^{+}23, YSAY23]. Aggression [ER21]. Agile [IPL^{+}18, MUS11, SU17c, UB18, VBWW22, ASL22]. Aging [LNB12]. Agnostic [TXJ23]. Agreement [AMI23, BBSR24, Li23, MMN19, WHW22, ZOAR24]. Agribusiness [dSRRdOT21, dSRRMC^{+}23]. Agricultural [CH16, KPRP24]. Agriculture [BYG22, CGRV20, CVG^{+}22, JGV^{+}21, MRM^{+}24, PKMS23, TL19]. Ahead [BFG^{+}23, KFP10]. AHP [KKTS22]. AI [ZLW21, AJA^{+}24, AAK24, AAN^{+}24, AGB22, AHP23, ABLR23, AHS^{+}23, Bad23, BAK23, CBR^{+}24, CSA^{+}24, CDD^{+}21, EKG24, ELCS20, FCG22, IPP24, KSG^{+}24b, LS23, MPK^{+}23, Orl22, OR24, PBP^{+}24, Pic21, Pri23, RFZ22, RT23, SKKS22, Suf23, TTN24, TZS^{+}21, TZE^{+}21, VM24a, YLX^{+}24, ZSB24, ZGL20]. AI-Based [CDD^{+}21, Orl22, Suf23]. AI-Driven [CBR^{+}24, EKG24, PBP^{+}24]. AI-Empowered [CSA^{+}24]. AI-Enabled [TZS^{+}21, TZE^{+}21]. Aided [DN20, MEH^{+}22, MKPG22, SF22]. Aimed [OT16]. AIoT [XSSA24]. Air [FHZ^{+}24, JMZ^{+}22, LLL21, VLC^{+}19]. Airborne [IGL^{+}23]. Airport [ASG24]. Airspace [BSM22]. al [LJR24, VBB24b]. Alert [ASMM22]. Alerting [HMR^{+}21]. Algebraic [KGKP20]. Algorithm [AMK^{+}21, AIP^{+}24, AAM^{+}24, ADM^{+}19b, AKM^{+}23, AK22, ADS21, ASL24, AMSM23, Bi14, CD23, CZZH15, CXG^{+}22, ELCS20, FSY^{+}22, FKZ22, FHZ^{+}24, GAMMPCRÁ24, Han13, HA23, HF20, HC21, JHL^{+}18, KSG24a, KKEV17, KLTC17, KHAM24, LWHR14, LSC^{+}17, LTY18, LNA17, LW19, LSJ19, LSZ21, LXY17, LZL^{+}21, LLSS24, MTA^{+}20, MKR22, MS22, OLCMdA20, ON23, RRMI22, RGCM21, Rók23, RCL21, SGG^{+}22, SFEK18, SOK24, SDM22, Str22, TAMA22, TSHO24, TTKZ19, Tra18, WYX18, WLL20, WZL^{+}20, XCW^{+}22, XQLZ19, XGN19, YC16, YYW20, YLY^{+}22, YBO^{+}23, YWW^{+}19, ZWWW17, ZCZ17, ZSZ18, ZMZ^{+}22, ZHH24]. Algorithms [ARAAA19, BPLBM21, CRMT23, GFPVLR^{+}23, GX20, KGK^{+}24, KK21, KK23a, KBF23, KHRG19, LBBPM21, LN19, LIM23, MWY19, MWL20, OOM^{+}23, RZCL19, RPHJ11, SLZJ23, SWY^{+}18, TBT^{+}23, TH20, VCJAMN22, WL20, WY19, XCW^{+}22, ZWS20, ZZY19]. Alignment [CDLY24, GSO^{+}10, RFD23, WJBZ22]. Allocation [ASL24, APAR22, CCCPV^{+}24, DNZ22, ELCS20, ENT21, GWF^{+}23, HF20, JWRX17, JHBW24, KSG24a, KALY17, KCY22, LTY18, LJR23, LJR24, LH21c, LWZL22, LNH23, LLSS24, NST23, OFK19, PAK^{+}23, QAM17, SDM21, SIJA10, SXXH19, SHN^{+}19, SNH^{+}19, TRST24, UFK^{+}10, WYS17, WSM20, YPMM12, ZWL^{+}24, ZX24]. Allow [KLKA19]. ALOHA [SVFV23, YLYL22]. Alpha [GBDDVS23]. Alquist [KHC^{+}24]. Altering [YG20]. Altitude [TSHO24, ZMZ^{+}22]. Alzheimer [DCA23]. Amazon [FFM^{+}23]. Ambalytics [KGRP21]. Ambient [AYH^{+}21, FAC22, GRJ^{+}19]. Ambivalence [Dav12]. American [ASFAAV24, GS11, LPA^{+}22]. Americans [Smi13]. Amharic [YAV^{+}21]. among [ASFAV23, CGMM22, HS18b, KVL^{+}22]. Amplify [SJG18]. An-Eye-Tracking [PDI21]. Analogy [DGHH^{+}20]. Analyse [MSZ23]. Analyses [GKFV24, KGRP21, KA23]. Analysing [DBD^{+}14, HDL^{+}14, MKKG19]. Analysis [AZAV22, AWBS23, AAN^{+}21, ACD^{+}22, ASM^{+}24, ASMM22, AMY24, ALDS19, ACAP20, BCSB24, BKR24, BAR^{+}21, BMGI21, BFVM14, BBH^{+}24, BPG19, CGMM22, CR21, CGRV20, CT20, CLC^{+}22, CMW18, CKK21, CWC^{+}22, COL22, CDD^{+}21, CGMVdUC19, CL19, DGMP21, DeD16, DWZN16, EvdMB20, EWGK24, EAKM22, FGSD22, FLLG24, FSG22, FSG^{+}16, FMMS21, GMD^{+}22, GMGF24, GRM23, GMH18, GVY^{+}23, GFPVLR^{+}23, GDS23, HM24, HT14, HOV23, HBM^{+}21, HKH^{+}24, HNK^{+}21, JTA^{+}21, JXA19, JAS21, JLW19, JSAO^{+}23, JSPH21, KDEA^{+}23, KCK^{+}19, KPRP24, KDE24, KBF23, LS15, LZ23, LSF19, LLJ22, LLY24, LLZ21, LLBB18, LFM^{+}19, LS23, MBH22, MIK^{+}23, MSR^{+}14, MEJA23, MK20, MC24, MKK^{+}22, MKPG22, MSJ^{+}23, MF20, OO19, OFV21, OCCB20, OLCMdA20, OMTFFL21, OAH^{+}24, Oza23, PR11, PMA^{+}22, PDT23, PBP^{+}24, PXC18, PVKG23, PBK22, PNdC^{+}23]. Analysis [PR20, RGB^{+}24, RBG^{+}22, RGMFM12, RGD24, SVV^{+}22, SLS20, SHC22, SFD20, SAR24, SHB23, SRSDF23, SL24b, TH21, TPKA20, TCG19, TPP^{+}23, TC21, Tre21, TMS24, TSDG22, THL21, UJ22, VCRCSV^{+}21, VM24b, WX20, WBK^{+}23, YWS22, YPG20, YLYL22, YMSP24, YST24, YSEK20, YZC^{+}20, YWW^{+}23, ZLL18, ZL19a, ZLS19, ZZ19, ZLW21, ZZL^{+}23]. Analytic [HNHH23, TWM20]. Analytical [DG13, EAKM22, GLRM19, KAEAB24, LAG^{+}18, SLY^{+}12]. Analytics [AKYP23, AFB23, AH21, BMS20, CL19, GMMI16, MEH^{+}22, MT20, MTK^{+}24, SPSS17, SJ12, VDSK23, VI16]. Analytics-as-a-Service [GMMI16]. Analyze [CTM19, SG21b]. Analyzing [AMG22, GRG21, JLH^{+}13, LF16, PL22a, SHC24, VVK^{+}21, YA24]. Anchor [VGK21]. Android [CKK21, CYO^{+}17, DDV22, JLH^{+}13, KAS^{+}20, TPKA20]. Anemia [RHV17]. Anemia-Like [RHV17]. Angle [AMSM23, ITH^{+}21]. Animal [MRM^{+}24, VCPT22]. Animation [LWW^{+}23]. ANN [LLYL22]. Annealing [JHL^{+}18, TL24a]. AnnoMI [WBK^{+}23]. Annotated [RDASB22, WBK^{+}23]. Annotating [PD23]. Annotation [AWBS23, LPM10, Sor12]. Annotations [HG12]. Annotopia [HG12]. Annotopia-Enabling [HG12]. Announcing [Giu18a]. Annoyance [BKR24]. Annual [Bod22]. Anomalies [DKH24]. Anomalous [CC22a]. Anomaly [CPRS20, FCR23, GRA^{+}23, GSP15, GTWN24, GKW^{+}10, KDE24, MNP^{+}24, MSN13, RGB^{+}24, SBRZ24, WBS^{+}24]. Anonymity [AH21, AAJ^{+}22, OKH13]. Anonymization [GM19, MBF^{+}12, TRB22]. Anonymized [ACAP20]. Anonymous [BDL22, ZLLW18]. Answering [ZZZL23a, ZZZL23b]. Answers [LS15]. Ant [GW15, LW19]. Antecedents [CFGD20]. Antenna [CLZL18, CLH^{+}23, JKM^{+}23, PBAAN19, USS19, YBO^{+}23]. Antennas [CLW23]. Anti [FLC21, HRKL21, HJA18, Yiu21a, Yiu21b]. Anti-Collision [HJA18]. Anti-Counterfeiting [Yiu21a, Yiu21b]. Anti-Forensics [FLC21]. Anti-Science [HRKL21]. Anticipation [Pin10]. Antisocial [MKSJ22]. Antonyms [SPGS23]. Anxiety [BBM20, SN23c]. Any [Mat20]. Anyone [Mot24]. AODV [SS15]. AOSR [DPR^{+}21]. AP [LTY18]. APBT [ZWWW17]. APCO [HP20]. API [AAY21, CH16, OAH^{+}24]. APM [dVBA23]. App [GM16, LLLD21, PRDK22, PJMM22]. Appliance [HSM19]. Applicable [LW21]. Application [AJ18, ADM^{+}19b, AV23, AAF^{+}22, ALST20, AGA^{+}22, CLC^{+}22, Chu23, DDV22, DCHM16, DCA23, FOJE19, GXWD19, GBKN23, Goe12, Gro17, HP16, ITH^{+}21, KVWC24, KTCI21, KA21, LWW^{+}23, LRdLS19, MNWK21, OP22, dSRRMC^{+}23, SLP^{+}22, SACP24, SOB^{+}19, SKA^{+}23, SFSN^{+}24, SN23c, SB23, VKV^{+}22, WSF23, XZWL24, YA24, ZJX18]. Application-Aware [AJ18]. Application-Partitioning [ADM^{+}19b]. Applications [AAY21, AN23, AWSA23, AHKC23, AN22, AMCI17, ADS21, AAN^{+}24, AA23b, ALLR16, ASN22, ASD^{+}22, ACA^{+}21, AAKS23, AKJB20, BFL20, BFG^{+}23, BM14, BCM^{+}19b, CFV19, CLZL18, CH16, CALHG^{+}24, DNKDN^{+}24, DBD^{+}14, DVED20, DV23, DHEK19, ESBE23, EMMP22, FLRS12, FAC22, FR16, GHFL24, HSI^{+}22, HSC19, HSC^{+}22, HPB^{+}24, HLLT10, JGV^{+}21, KAEAB24, KI20, KKS^{+}19, LTW21, LW21, LA20, LG23, MW24, MMT24, NBT22, NDKBL19, OFV21, OBK23, PKF24, PBAAN19, PBK22, PRS22, PNdC^{+}23, PNM^{+}20, QAQA21, QLR^{+}22, RYY10, RYAA22, Rei23, RNST23, dSRRdOT21, SCB21, ST22, SG19, SA21, SN23a, TFSBSPSC24, TAAK21, TN20, URH^{+}23, VKM^{+}19, VSE21, VLK24, WGN23, ZMKR22, Zar22, dVBA23]. Applied [AWUF19, BdBC23, KT22, MKSK23, RDASB22, SJ12, VL24, XJ10]. Applying [AZS23, FFM^{+}23, HBFJ20, PTK^{+}23, SOSR^{+}16]. Approach [Adi23, ATY20, ASCM22, ARAAA19, AVA^{+}22, ALL24, ATH18, ADK^{+}24, AAC21, AAARZ24, ACK^{+}16, AIPV21, ABS20, BCNS20, BdCRM10, BHC23, BC19, BHS24, BFFZ^{+}12, BRC24, CCC^{+}24, CBR^{+}24, CN24, CZY17, CHW^{+}21, CHMZ21, Cro19, DGADE14, DCA16, DPR^{+}21, DIE^{+}21, DGKL15, FOJE19, FC22, FML^{+}22, FMA^{+}23, GMRRRQ^{+}16, GHFM20, GGD23, GMH18, GRJ^{+}19, GB23, Gro17, HSSPTM^{+}23, HSL^{+}23, JUN^{+}24, JHL^{+}18, KHAA21, KPT^{+}24, KG18, KHC^{+}24, KFK^{+}24, LA23, LLZL19, LYD22, LJR23, LJR24, LBK22, Mah17, MSZ23, MBH22, MP16, MGBG21, MK20, MSC^{+}21, NLPV12, NPMPM23, NAFM^{+}24, OFK19, OAH^{+}24, PCC^{+}18, Pet20, PRS22, Pri10, QQM24, QAM17, RRMI22, RGM^{+}21, SU17c, SYB^{+}18, SE23, SOE^{+}24, SGWK23, SVGGO24, SV24, SSDS23, SWJ20, SDM22, SFSN^{+}24, SF22, SWY^{+}18, SYLL21, SB23, TSZ^{+}18, VGK21, VBWW22, VC23, WLHR13]. Approach [WKW18, WWLZ21, WLWZ23, XY21, YHG^{+}21, YYY^{+}23]. Approaches [AJ24, Ala20, AHS^{+}23, FWLY23, FDN^{+}24, GHA^{+}21, GZC21, GKS10, KGKP20, KST19, KPKM23, LWW22, LGV13, MAW^{+}23, MHSH16, PSTZ23, RGB^{+}24, SVK^{+}22, SE15]. Appropriate [ZL14]. Approximate [HNHH23, MPG20, QSF^{+}17]. Apps [AZS23, AP21, CKK21, MVAHBL^{+}23, RRLT20, TPKA20]. Aquila [TAMA22]. ARAAC [QAM17]. Arabia [AAN^{+}21]. Arabic [ADAq^{+}22, AA19a, ATH18, AAC21, AFAAM19, FGSD22, GMA20]. ARBiLSTM [YZCR24]. ARBiLSTM-Based [YZCR24]. Architecting [AA19b, DSM21]. Architectural [BM14, Dua21, MZS19, PR11, SKFM24, Tas10]. Architecture [ASMM22, AAA^{+}19, ACM^{+}19, BHS24, BCFP21, CTP^{+}22, CGM^{+}19, CAL23, DCL24, EABE19, FWVF19, FMMS21, GANT21, GRA^{+}23, HHLZ23, HEP^{+}11, KGRP21, KPP^{+}20a, KE22, KDJ24, LBG^{+}14, LW11, LZL^{+}17, LW21, LFM^{+}19, MTF^{+}21, NIK^{+}21, PPKS21, PPR^{+}20, Pec18, PORM^{+}23, PLA^{+}24, PCF^{+}14, PP22b, QLR^{+}22, RC18, RDD^{+}14, RAA^{+}18, RF23, SG12, SADP21, Sem11, SGW22, SNKG20, SPT23, TMP22, TKH^{+}11, YHSW19, ZTBD20]. Architectures [AKK^{+}24, ASN22, BML19, CBM^{+}20, CDLY24, CFPP21, ELC^{+}19, ELCS20, FP19, GKS10, GRA21, KKD24, MEH^{+}22, MR23b, SU17a]. Architectures-Overlay-Based [GKS10]. Archival [LPPG22]. Archives [DBD^{+}14, HDL^{+}14]. Archiving [RDD^{+}14]. ARCOMEM [PCF^{+}14, RDD^{+}14]. Arduino [QNJ24]. Area [BL22, DWZN16, OAG21, SYGY21, TSHO24, TFSBSPSC24, YT15]. Areas [AMC^{+}24, BM23, HD19, Hel15, OT16, OFK19, USS19, VLBRMP21]. ARIBC [GK21]. ARIMA [CMP^{+}19, KKBG23, KPKM23]. Arising [AAN^{+}24, BTS^{+}21, dSFdRP24]. Array [DKH24, LWHR14, USS19, WLHR13]. Arrival [CZZH15]. Arrivals [SVFV23]. Art [AZH22, BM23, CEP24, HNK^{+}21, HSTK23, KSA^{+}21, KVL^{+}22, MHQ^{+}23, PKF24, PVKG23, QQX^{+}24, Ray23, RGdSK22, XLW^{+}24]. Articles [DMS^{+}22, KLKA19, TZE^{+}21]. Artificial [BRC24, CCTC23, CL19, DVSBV^{+}24, DN20, DRK^{+}22, DAMAS^{+}19, EWGK24, ELCS20, FLLC24, KS24, Khr20, Lai24, MDS^{+}24, Min20, MGLBMMSC20, OP22, PMP^{+}22b, RF24, RTN^{+}22, SCB21, SG19, SK24, VTPT24, VL24, VMAG23, VCGOMC^{+}21, VCJAMN22]. Artificial-Intelligence-Based [CCTC23]. Artwork [LWBM22]. Aspect [ZL19a, ZGZ^{+}18]. Aspect-Level [ZL19a, ZGZ^{+}18]. Aspects [GPMO20, LZC21, SE15, Tri14]. Aspirations [HT14]. Assembling [LL20a]. Assess [CS11, GDCM19, HBFJ20, RRLT20, TRLR24]. Assessing [AA23a, AZV16, DN20, GBDDVS23, KAL^{+}19, PP22b, RABC21, SSR17, SG21b, VCB14, ZKV12, TN20]. Assessment [CR23a, CCM^{+}17, DHG21, DGF23, FZ15, GL21, JCPV22, KHAA21, KKK^{+}20, KZ20, KID^{+}15, LZ11, MS23, MM23, MBF^{+}12, NJH^{+}24, NE21, NPMPM23, PMGG21, Pet20, RGB^{+}23, SRS23, SWJ20, TWM20, TMTMB20, TA15, ZL14]. Assessments [PKCF21]. Assessor [FAS^{+}23]. Assets [SOSC^{+}16]. Assignment [KSSF19, YLK^{+}19]. Assist [VKM^{+}19]. Assistance [ASMM22, VCJAMN22, YLX^{+}24, ZZZ^{+}22]. Assistant [VCGOMC^{+}21]. Assisted [AIP^{+}24, CSDAM^{+}23, CLH^{+}23, FAC22, GHZY23, IHCSG^{+}24, LLCL23, LLW^{+}24, NBT22, NE21, SHC22, SOK24, SXXW20, ZPX24, ZWL^{+}24]. Assisting [You23]. Assistive [PS17]. Associated [KIG23]. Association [GRB^{+}21, LTY18]. Assurance [AAD^{+}12]. Asthma [AML22]. Astronaut [WBG12]. AT-Text [LL20a]. Atoms [Jur12]. ATT&CK [BMB^{+}23b]. Attachment [Hua15]. Attack [AJ24, AKDL24, AKM^{+}23, AMEF21, AN22, AYRA21, AAA^{+}24b, BSM22, DSMM21, DL19, GRA^{+}23, GHL^{+}23, HSC19, KDD^{+}21, KBF23, LBC18, MAVS20, MPCS^{+}23, MPM24, MS22, QDL22, SBN^{+}23, SA21, SZL21, SS15, SHLH24, VBB24a, VBB24b, WKW18, WLCS17, WQH23, ZHZ^{+}19, ZXWZ21]. Attacks [AAM^{+}24, AAA^{+}24a, Ala20, AR23, AV23, BMB^{+}23a, BB17, FLC21, FHXL19, FMMS21, GS23b, GMGK20, HCXH16, HAA23, HK18, IAM24, KMK24, KAAAA22, KRSK24, LMK18, LNT^{+}23, MIK^{+}23, MEJA23, NPMPM23, OOM^{+}23, OB13, PB23, SGG^{+}22, SE23, SOE^{+}24, SK19, SGDT19, TS22, YC22, YG20, ZHZ^{+}19]. Attention [CZZZ22, CGnCD20, CYC^{+}21, CZL21, CCYL24, DDMA24, KES24, LH21a, LLZL19, LDZ^{+}23, LMPT19, PDI21, PW21, PB23, SBN^{+}23, SJ20, TCH23, VLK24, WX20, WYZ24, WZ18, WCYL20, XDS^{+}19, YWS22, YLPW21, YWW^{+}23, ZLXY19, ZL19a, ZCHG19, ZLW21, ZGZ^{+}18, ZXMB22]. Attention-Based [KES24, TCH23]. Attentional [SW21a, WCZ23]. Attentive [KES22]. Attitude [ASAH^{+}23, LL20b, SG21a]. Attitudes [HRKL21]. Attraction [GLRM19]. Attribute [MKPG22, MLL^{+}22, SHHM21, SYZ19, SSX20]. Attribute-Based [SHHM21, SSX20]. Attributed [LLLD21]. Attributes [KDKG22, PDTM15, VKK^{+}19, VKKG22]. Attribution [FRKS22]. Atypical [HLC24]. Audience [LOL23]. Audio [Hor19, ITH^{+}21, Lal14, LF23]. Audio-Visual [Hor19, ITH^{+}21, Lal14]. Audiovisual [VKV^{+}22]. Auditable [MJG24]. Augmentation [MKSJ22, VR21, WWLZ21]. Augmented [Ant12, BS17, FHL^{+}24, HF22, Jur12, MAM^{+}21, PEN23]. Augmenting [RGP22]. Australia [KYG22]. Australian [EWGK24, ERBL12]. Austrian [ESB^{+}20]. Authenticated [DNKDN^{+}24]. Authentication [ADM19a, AFS^{+}22, Alb21, ADS20, ADS21, AMI23, BBSR24, BAR^{+}21, BKCL23, CCC19, ESBE23, EAHO21, GPPB^{+}21, GMS^{+}20, JUH^{+}23, Li23, LDZP24, MMS22, OPGH23, PM24, PAM^{+}24, RSX18, SBS18, SST23, VKV^{+}22, WD24, ZOAR24]. Authentication-Based [GMS^{+}20]. Author [BHH12, DMS^{+}22]. Authority [WZX20]. Authorizations [SWGL19]. Authorship [CGMM22, FRKS22, KRSF22, RKS^{+}21]. Autism [Pri23]. Auto [AMZ^{+}20]. Auto-Closing [AMZ^{+}20]. Autoencoder [DDD^{+}22, LNT^{+}24, MA23, SLZY21, YMH22]. Autoencoders [ZH21]. Automated [ABCC22, KMS^{+}12, KSE^{+}20, LPPG22, LLY24, PG19, SKM21, TZ22, WML24, YMSP24]. Automatic [AGO^{+}23b, GPA23, GSO^{+}10, MT23, Ngu19, PCM22, RK19, SYLL21, VKKA17]. Automatically [MMT24, WZL23]. Automating [AF22, BGD09, PMGG21]. Automation [CSIS19, CMH10, LK20, MSAD22, MUS11, WMS23]. Automations [KVVD22]. Automaton [PG23]. AutoML [GOM23]. AutoML-Driven [GOM23]. Automotive [HH20, SVGGO24, SACP24].

Autonomic [ALLR16, AS15, AAE^{+}22, LBG^{+}14, SKT^{+}19]. Autonomous [Adi23, BZDP19, Dua21, BBBHA24, FAT19, HS20, IPP24, KUBS22, MCH^{+}20, TTO^{+}19, YKY18]. Autoregressive [CMP^{+}19, TRLR24]. AUV [LLBB18]. Availability [Kab23, SN23b, SHB23, WC23, YT18]. Available [GMB^{+}21]. Avatar [TEE10]. Average [CMP^{+}19, KK23a]. Aviation [KS24]. Avoidance [GM20, HS20, KHAM24]. Avoiding [GB20, IGL^{+}23]. Award [Giu18a]. Aware [AJ18, Alb19, ALLR16, AS24, AKJB20, BHH12, CLT^{+}22, CGT21, CFP17, CMR24, EH21, EVCL21, GM20, HZQ24, KALY17, Lee17b, LWH24, LW15, MEO18a, MEO18b, MKK17a, MMS^{+}19, MDRR24, PDTM15, PVM22, BNK^{+}23, SJ21, SH21, Sul22, SMK^{+}22, TCH18, VL18a, VM24a, WBS^{+}24, XSX18, YSAY23, ZJJS^{+}23, EMHF19]. Awareness [AS19, CSA^{+}24, EKG24, GWK17, GRJ^{+}19, HSC19, HT14, KKTS22, LLB24, LCV^{+}17, PR12, SAM^{+}18]. Azure [HAA23, RHV17]. Azure-Based [RHV17].

B2C [TLC15]. B5G [BAS24]. BA [HJA18, KSO19]. Baby [YST24]. Back [SLW17]. Back-Off [SLW17]. Backbone [SMG13, UFK^{+}10]. Backbones [FPKK22]. Background [DWL21, ESB^{+}20, HT14, JLW19, Ray23]. Backhaul [GAMMPCRÁ24, PP22b, VLC^{+}19]. Backpressure [LRL^{+}24]. Bagging [YYW20]. Baidu [ZLW21]. Balanced [BdCRM10]. Balancing [ARAAA19, CMPT24, GW15, LW19, LZL^{+}21, SHBS19]. Ball [LA23]. Band [ACD^{+}22, DDV22, LZL^{+}23, PBAAN19, SOS^{+}23]. Bands [SS23]. Bandwidth [CR11, ELC^{+}19, LTY18, LWZL22, LNH23, LHL24b, Rók23]. Banging [KYT^{+}23]. Bangla [DMS^{+}22]. Bank [ASFAAV24, CMS21, LLHW20, LDZ^{+}23, MDPHB23, WWYQ19]. Banking [ASWAE23, CPM17, OBK23]. Bankruptcy [LPA^{+}22, PLA^{+}24]. Banks [JAS21]. Banner [PDI21]. BART [LC23]. BART-IT [LC23]. Base [BZDP19, CPKK23, ENT21, WJBZ22, YPXH19]. Based [AIP^{+}24, AAM^{+}24, ADM19a, AAY21, Adi23, AAK24, AMG22, AN23, ATFMOU^{+}24, ATB^{+}24, ABCC22, AAAIA24, AHKC23, AK22, AlQ23, AA23a, AEM^{+}23, Alb21, ASF^{+}23, ALL24, AASAI22, ACF24, AV23, AA19b, ALLR16, AA21c, AF21, AMZ^{+}20, AMSM23, ASL17, AWUF19, ARM20, AAE^{+}22, AAJ^{+}22, AMY24, AMI23, AHP23, AAA^{+}20, AIPV21, ADG^{+}19, AHS^{+}23, ABS20, BYG22, BBSR24, BCSB24, BKL^{+}22, BZP^{+}21, BZDP19, BP24, BKW^{+}24, BAR^{+}21, BKCL23, BL22, BTB22, BOB^{+}20, BHS24, BCFP21, BZNB23, BMGI21, BCMPW21, BDL22, BM23, BRP^{+}13, BMP21, CP10, CMPT24, CBRS21, CDM23, CB24, CLC^{+}22, CVG^{+}22, CN24, CCKH21, CYC^{+}21, CTP^{+}22, CZY17, CGLR19, CZL21, CHMZ21, CQWC22, CWC^{+}22, CXZ23, CSA^{+}24, CDLY24, CZZH15, CCTC23, COL21, COL22, Chu23, CTM19, CTM22b, CT24b, CDD^{+}21, DTR22, DQ18, DXL^{+}22, DH22, DJC^{+}22, DHDAu22]. Based [DWHY24, DDPVP18, DZH^{+}22, DDMA24, DVED20, DDZ18, DHEK19, DLX^{+}19, DWY19b, DYLZ21, DSM21, DCA23, ESBE23, ELCS20, ENT21, FOJE19, FSY^{+}22, FKM24, FHXL19, FC19, FLRS12, FK21, FTAA21, FHZ^{+}24, FBDR23, FMA^{+}20, FSG22, FR15, FP19, FR16, FLSL24, GAGB19, GPPB^{+}21, Gao21, GWF^{+}23, GZ23, GLZL24, GR22a, GFHK^{+}12, GL21, GM20, GDARM18, GMH18, GCEOBRT23, GSL20b, GAMMPCRÁ24, GDS23, GRA^{+}23, GKS10, GK21, GNK^{+}10, GOM23, GHL^{+}23, GRRZ18, GVC13, GMS^{+}20, GAA21, GKW^{+}10, HD19, HCXH16, HK18, Han13, HLZ^{+}18, HHLZ23, HEP^{+}11, HZWJ21, HHM^{+}23, HNHH23, HCHP16, HWZP18, HLZ23, HKZ^{+}22, HN18, HAL^{+}22, HMR^{+}21, HZCZ22, HGL^{+}23, HG12, HBMS20, HP19, IFF21, ISLARP^{+}22, ITH^{+}21, JBS^{+}23, JXA19, JEJ^{+}24, JHBW24, JP21, JLGF19, JII17, KKD24, KKP22a, KSD^{+}23, KES24, KAM^{+}19, KSG24a, KST19, KDEA^{+}23, KKL^{+}20, KHN^{+}22, KPRP24]. Based [KYS21, KE22, KTAH^{+}23, KPC19, KK20, KAL^{+}19, KBP22, KHM20, KSO19, KFP10, KHRG19, KH19, KSJZ23, KKT20, LWW^{+}23, LA23, LBA^{+}24, LWY20, LH21a, LPVM24, LDd^{+}24, LWY^{+}15, LNA17, LWX18, LMK18, Li18, LZS18, LW19, LSJ19, LLZL19, LMH19, LSZ21, LLLD21, LSH21, LYD22, LLJ22, LLYL22, LWW22, LLY23, LJR23, Li23, LLY24, LRL^{+}24, LJR24, LYY^{+}24, LDZP24, LLZ21, LZL^{+}23, LBB23, LLW21, LWH24, LXG^{+}23, LLJ16, LXY17, LHW18, LCW^{+}18, LH21c, LW21, LNH23, LZY^{+}24, LCCC24, LT24b, LT24a, LLSS24, LHL24a, LHL24b, LGCM24, LCV^{+}17, LHSS12, LSM^{+}12, LW15, LBK22, MXT^{+}16, MLPH23, MCS^{+}24, MB22, MDS^{+}24, MAW^{+}23, MDB22, MKKS18, MMD22, MMS22, MBH22, MMM21, MRRMC21, MGV17, MWY19, MMMH19, MRM^{+}24, MK20, MCPLBH^{+}24, MM21, MA23, MM23, MKM^{+}19, MUS11, NLT^{+}23, NJH^{+}24, NLD^{+}24, NMT23, NSJGN21, NRS24, OS23, OMD^{+}22, OFV21, OT16, Orl22, OSdS24]. Based [OP22, OAH^{+}24, PCROB21, PPR^{+}20, PLL20, PMP22a, PPMMS19, PPGC16, PZH^{+}18, PP22a, PCC^{+}18, PORM^{+}23, PZC21, PW21, PC23, PCM22, PBK22, PYC21, PLG23, PNM^{+}20, QLK^{+}24, QMN19, RGM^{+}21, RAU^{+}24, RLB^{+}23, RL24, RIS^{+}17, RRT^{+}23, RRT20, RPB^{+}17, RRdBSS22, RDASB22, RGCM21, RAK^{+}20, RHV17, RSG24, RSL^{+}24, SAJ24, SU17c, SBN^{+}23, SHA23, SBS18, SJG18, SE23, SJA^{+}23, SGMMRG21, SACP24, SMB23, SHHM21, SCA^{+}19, SSM^{+}24, SAR24, SA21, SFEK18, SW21b, SC23, SOB^{+}19, SZL21, SWC^{+}23, SSD24, SSR^{+}21, SOK24, STK21, SDM21, SNKG20, SST23, SMS^{+}24, SM19, SWJ20, SGDT19, Str22, SL24a, Suf23, Sul22, SKHK23, SLY^{+}12, SZFC18, SZC19, SSX20, SYLL21, SZZL23, SMK^{+}22, TSHO24, TTKZ19, TCH23, TCB^{+}19, TPKA20, TMP22, TFKY19, TL17, TYH23, TZ22, Tra18, TWW24, TRST24, TKT^{+}18, TZE^{+}21, TCYZ22, USS19]. Based [UANU20, VC23, VMD^{+}22, VBB^{+}24c, WLH^{+}17, WHXL18, WYX18, WCSZ18, WLL^{+}19, WWYQ19, WLL20, WSM20, WL20, WX20, WYL^{+}21, WLJM21, WL21, WWD21, WHW22, WXL22, WL22, WLG^{+}22, WXL^{+}23, WC23, WLLY24, WLCC24, WUH^{+}23, WS21, WZG^{+}18, WSD22, WD24, WZ18, WY19, WC20, WZX20, WRHH23, XGY^{+}22, XDH24, XDS^{+}19, XGN19, XNC21, XLM^{+}23, XNZ23a, XSSA24, YCS17, YC22, YWS22, YCHG18, YPG20, YYW20, YZZ^{+}20, YHG^{+}21, YLL21, YZP22, YWLY22, YLYL22, YYY^{+}23, YZCR24, YHSW19, YPMM12, YZ23, YSEK20, YLPW21, YXL^{+}21, YWW^{+}19, YPLZ19, YLK^{+}19, YLJ^{+}19, YWW^{+}23, ZAS^{+}24, ZWWW17, ZWZ17, ZLL18, ZL19b, ZCHG19, ZWPL19, ZL20, ZLT22, ZLZ^{+}24, ZX24, Zha19, ZZ19, ZGL20, ZY20a, ZLW21, ZZYC21, ZHW23, ZHH24, ZYLM17, ZLZZ23, ZQCC16, ZZZ^{+}22, ZZZL23b, ZJJS^{+}23, ZSS^{+}22, DH18, TADS20]. Baseline [GSB^{+}23, KDE24]. Baselines [APJ24, MKK^{+}22]. Bases [AMR^{+}21]. Basic [RMB^{+}12]. Basics [LBBPM21]. Basis [BPLBM21, LLHW20, SSOÁ^{+}16]. BASN [WCYL20]. BASN-Learning [WCYL20]. Bat [ASL24, FKZ22, SGMBK24]. Bat-Inspired [ASL24]. Bat-Optimized [SGMBK24]. Batteries [RIS^{+}17]. Battery [BCZ^{+}23, RBG^{+}22]. Batteryless [SOK24]. Battlefield [FPKK22]. Bayes [HKZ^{+}22, SWJ20]. Bayesian [AES21, CR21, SWJ20, WQH23]. BB84 [FPGO24]. BDA [BKCL23]. Be [RBVV22, SE15]. Beam [LSJ19, LLW^{+}24]. Beam-Forming [LLW^{+}24]. Beamforming [AMSM23, SRS^{+}24]. Bearer [ZHW23]. Beat [VBB24a, VBB24b]. Bed [GNNZRCRG18]. Bee [RAU^{+}24]. Before [MPG20]. Beginners [Fra23]. Behavior [AMAE22, AASAI22, ASMM22, BM19, EHJ17, GS23a, GRB^{+}21, JLH^{+}13, LLZ21, MKSJ22, PVKG23, RSX18, SW21a, TN20, XZY^{+}24, YKY18, ZLOW24]. Behavioral [AA23a, CMR24, ESLB20, FNN21, HMR^{+}21, LOL23]. Behaviors [GRG21]. Behaviour [CP10, MK20, TZ22]. Behaviour-Based [TZ22]. Behavioural [MT17a]. Behaviours [AAC21, SAF14]. Behind [MNH^{+}24]. Beidou [DLX^{+}19]. Being [DIG20, HCW20, LL20b, PPDC22]. Belief [ZL20]. Beliefs [GCA^{+}22]. Beliefs-Predicting [GCA^{+}22]. Benchmark [AAAT21, NLD^{+}24, SO24, TMS24]. Benchmarking [HBMS20, SLZJ23]. Benchmarks [LPA^{+}22]. Benefit [PCM24]. Benefits [ASL22, HSI^{+}22, HSTK23, KAL^{+}19, Kra22, MW24]. Bergen [DIG20]. BERT [BRM23, LLJ22, SPC22, SPGS23]. BERT- [LLJ22]. BERT4Loc [BRM23]. Best [KTUF19]. Better [AGB22, AF22, GLRM19, KID^{+}15, LOL21]. Between [BCM^{+}19b, RBSMRN21, VM24b, vSR24, AHMI22, GSD^{+}17, GRB^{+}21, HT14, HKZ^{+}22, JJ21, LL23b, MS20, Nil12, PL22a, PNS23, QWR18, RFS22, SMG13, SMR24, SRM20, THT18, TTN24, TEE10, ZLY^{+}20]. Beyond [ABH24, HL23, JNARV^{+}24, KHC^{+}23, ND24, YHN21, ZSB24, Del23, Ves18]. BGP [SAR24, TRC^{+}21]. Bi [ZGZ^{+}18]. Bi-Directional [ZGZ^{+}18]. Bias [AAAT21, AYRA21, GR22b, HP20, PP22a, RCL21]. Biases [RS22]. Bibliographic [AZAV22]. Bibliometric [KGRP21, PNdC^{+}23, YMJ20].

Bidirectional [ATH18, RAA^{+}22, WYW^{+}22, WZ18]. Big [AKYP23, AFB23, AEM^{+}23, AAA^{+}24b, BMS20, BJL^{+}22, CGLR19, CL19, DVED20, GRRZ18, KGK^{+}24, KBF^{+}22, LG20, MSFM16, Pec18, Pil18, PH16, QMN19, SAZ18, TG13, TZE^{+}21, XSM22]. BiGRU [ZLXY19]. Billing [EAHO21]. BiLSTM [LLJ22, YWW^{+}23]. BiLSTM-Based [LLJ22]. BIM [CTM22b, CT24a, WS21]. BIM-Based [WS21]. Binarized [NB24]. Binary [CD23, EAKM22, FKZ22, GB23, LBBPM21, OMTFFL21, RKY20, SDDB24, WCYL20]. Binding [LBC18]. Biogas [MSJ^{+}23]. Biologists [KH10, NPB12]. Biomedical [MSL^{+}23, XSM22]. Biometric [AFS^{+}22, KDKG22, PPR^{+}20]. Biometrics [RSG24]. Biotec [OLMBRCRC23]. Bipartite [KK23a, PZC21, TAHO23]. Bit [MLPH23]. Bitcoin [CTM22a, CMS21, WKD22, CTM19, LF16, OKH13, Tre22]. Bits [Jur12]. Black [BN14, GHL^{+}23, ZHZ^{+}19]. Black-Box [GHL^{+}23]. Blackhole [SS15]. Blacklist [SZL21]. BlackWatch [HSC19]. BLE [DDV22, QLK^{+}24]. BLE-Connected [DDV22]. BlenderBot [KHC^{+}23]. Blind [CZK^{+}17]. Block [LCCC24, SBN^{+}23]. Block-Wise [LCCC24]. Blockage [KIB^{+}22]. Blockchain [ASWAE23, ATB^{+}24, ACS24, AMEF21, AM23, AA20, AYH23, AAAKJ22, ABB23, AHP23, BARL24, BBGP19, BKCL23, BOH23, BFL20, BL22, BMS24, BDL22, Cal22, CMPT24, CVG^{+}22, CGLR19, CTNS21, CPM17, CTM21, CTM22b, CT24a, CT24b, CCA^{+}20, CAS^{+}20, DCD24, DT24, DWL21, ESLB20, EKG24, EWGK24, FWLY23, FKM24, FWL22, FAC22, Fra24, GPPB^{+}21, GLD^{+}18, GdSdC23, HSI^{+}22, HM20, HWCL17, HW16, HP19, IPP24, IDR23, JJFC22, JJ21, KTCI21, KT22, KPP^{+}20b, KDG24, KAS^{+}22, KK20, KXZ^{+}24, KHM24, KKTK20, KL23, LB24, LLHW20, LT24b, LT24a, LHL24a, LHL24b, MCS^{+}24, MCA23, MCT^{+}23, MIK^{+}23, MDPHB23, MZ19, NB24, NPWC24, OO19, OHSN24, PNMK22, PPR^{+}20, PKMS23, PM24, PBK22, PYC21, PCM24, PLOT23, PRS22, PSGM22, RL24, RKT19, dSRRdOT21, dSRRMC^{+}23, RH24, SVGGO24, SLS20, SMB23, SG19, SW21b, SN23b, SNKG20]. Blockchain [SST23, SPT23, SA24, SGDT19, SNS22, TADS20, Tre21, VSE21, VFS^{+}19, WL21, WHW22, WADL^{+}24, WRHH23, XKKL23, XCW^{+}22, YHG^{+}21, Yiu21a, Yiu21b, ZMKR22, ZPX24, ZDP^{+}22]. Blockchain- [LT24b]. Blockchain-Assisted [ZPX24]. Blockchain-Based [ATB^{+}24, AHP23, BKCL23, BL22, BDL22, CMPT24, CVG^{+}22, CGLR19, CT24b, FKM24, LT24a, LHL24b, MCS^{+}24, PBK22, RL24, SMB23, SW21b, SGDT19, WRHH23]. Blockchain-Driven [ACS24, RH24]. Blockchain-Empowered [HWCL17]. Blockchain-Enabled [DWL21, MDPHB23, PM24, SST23, SPT23, WADL^{+}24, Yiu21b]. Blockchain-IoT [CTNS21]. Blockchained [CKFH22]. Blockchains [GTQ^{+}22, QXC^{+}21, Sta22, SG23]. Blocks [DWY19b, YXL^{+}21]. Blog [YYW20]. Blogger [ZVV12]. Blogosphere [ZVV12]. Blogs [Zaf12]. Blueprinting [NLPV12]. Bluetooth [ZSB19]. Boards [VMD^{+}22]. Body [GR22a, HRAA19, SG21a]. Body-to-Body [HRAA19]. Book [OF23]. Boomerang [BCG10]. Boost [ESB^{+}20]. Boosting [KHC^{+}23]. Bootstrap [SKM21]. Border [KSE^{+}20]. Borders [MAW^{+}19]. Botnet [HN18]. Botnets [GKM24, SGDT19, TZ22, VSEH^{+}21]. Bottleneck [LWZL22]. Boundaries [Del23, ND24]. Bounded [LYP^{+}19]. Box [BN14, GHL^{+}23, VGK21, ZXJ22]. Box [FMA^{+}23]. BPET [FKM24]. BPMNE4IoT [KGWR23]. Brain [BPLBM21, JPVC21]. Branchial [IKS^{+}21].

Brand [HP20, Zha19]. Brazil [CBB17]. Brazilian [Uto13]. Breaches [RSV^{+}24]. Breakdowns [BLW^{+}17]. Breaking [NSD^{+}22]. Breast [OMD^{+}22]. Bridge [SPM^{+}22, WL20]. Bridging [BCL11, HL19]. Brief [FK24, GRA21]. Bringing [GNK^{+}10]. Broadband [IUIL23]. Broadcast [KH19, NSV17, YK21]. Broadening [SSS^{+}19]. BROV [TRC^{+}21]. Browser [IFF21, KSO19, PV20, PV22]. Browsing [JDH21]. BS [BZDP19]. Bubble [BdBCCG21]. Budget [ZY15]. Buffer [LLCL23, LHW24, LW15, WJ21, YY17]. Buffered [XDH24]. Build [NLD^{+}24]. Building [AAAT21, CTM22b, GHFL24, HBM^{+}21, LLJ16, MCA23, WLJM21]. Buildings [JK24, ZGL20]. Built [ZLS19]. Built-in [ZLS19]. Bursty [LDM22]. Bus [Ant24, DLX^{+}19, KAM^{+}19, YZ23, YKY18]. Business [CNC^{+}21, FFP^{+}22, GBKN23, HL19, KA23, KV23, LSN21, MUS11, RGdSK22, RR24b, SA24, SKM21]. Button [PDMK19]. Buzzwords [LLJ22]. Byproduct [BdBCCG21]. Byzantine [CAS^{+}20].

C [BHC23, FMA^{+}23]. C-Box [FMA^{+}23]. C-V2X [BHC23]. CA [TDW24]. CABEI [ASFAAV24]. Cache [DWHY24, DNZ22, LNA17, LZS18, LXG^{+}23, LH21c, LZY^{+}24, YC22, YSEK20, ZWZ19]. Cache-Enabled [LXG^{+}23]. Cache-Sharing [LZY^{+}24]. CacheHawkeye [YC22]. Caches [PVM22]. Caching [HNHH23, HF22, LSH21, RF21, SB22, SSF^{+}18, TTKZ19, ZHW23]. Cadastral [DGADE14]. Cafés [PBB20]. Calculated [GRA^{+}23]. Calculation [DCHM16]. Call [KE22, OAH^{+}24]. Calls [YGQ^{+}22]. Camera [KES24, KKL^{+}20, TCH18]. Camera-Radar [KES24]. Cameras [XGY^{+}22]. Campania [TC21]. Campus [JK24, VCPPRC20]. Can [BCT21, DGGS19, EM09, LPM10, MKHR21, SJ12, Tre22, ZZL^{+}23]. Canada [Roy14]. Cancellation [LZL^{+}23]. Cancer [OMD^{+}22]. Canny [Vem22]. Capabilities [AlQ23, KKBD10, RSL^{+}24, UFK^{+}10]. Capacity [ADAB21, KDJ24, KLG^{+}24, WYL18]. Capital [AP21, PKCF21, Smi13]. Capsicum [AMR^{+}21]. Capsule [KYS21, NB24]. Capsule-Based [KYS21]. Captioning [LH21a]. Capture [HJA18]. Car [CKK21, Chu23, HZW^{+}20, HSC^{+}22, ZYLM17]. Car-Following [HZW^{+}20, HSC^{+}22]. Carbon [VM24a]. Carbon-Aware [VM24a]. Card [BARL24]. Cardiac [ABLR23]. CARDIAN [SACP24]. Cards [ASWAE23]. Care [ACBP19, DVSBV^{+}24, HMR^{+}21, MDS^{+}24, MGFF14, Mot24, PTK^{+}23]. Career [VNJ18]. Cargo [RF23]. Carlo [PPM21]. Carrier [Lee17a, LWZ^{+}21]. Carry [FOJE19, KFP10]. Carrying [HGL^{+}23]. Cars [GAGB19, Gog12]. Cascade [AD20, IAG^{+}21]. Cascaded [TH21]. Case [AP21, ADAq^{+}22, CR21, DMS^{+}22, ERBL12, FCR23, FZ15, FMNS11, FFM^{+}23, Gar12, dPCGMC20, GMMI16, HDK^{+}12, Hua15, JHC^{+}20, JP21, KPP^{+}20b, KG18, KDKG22, Kra22, KNN^{+}24, LWBM22, MMF^{+}19, MMH16, MS20, MK20, MTHN22, NZ14, NSJGN21, OLMBRCRC23, PHG^{+}20, AJT^{+}23, RHHN21, SU17b, Sem11, TSD24, THL21, Uto13, VLMP21, WLJM21, ZAV14, ZXJ22]. Case-PrepaTec [JP21]. Cases [AAAAM^{+}21, ASG24, BTS^{+}21, GGT^{+}23, KRSF22, SG19, VM24b]. CatBoost [AAM^{+}24]. Categories [vSR24]. Category [LCC^{+}24]. Causal [ZX24]. Causal-Structure-Based [ZX24]. CBDCs [JJ21]. CCAM [NMH^{+}21]. CCFD [BARL24]. CCrFS [MSAA22]. CCTA [DOB21]. CDN [GZ23, IFF21]. CDNs [JLGF19]. CEEMDAN [YZCR24]. Cell [KDJ24, YBO^{+}23]. Cells [QQM24]. Cellular [AZ19, BCM^{+}19a, CXZ23, KSG24a, KP19, KTP17, KFP10, OB13, OAI^{+}22, OFK19, SSF^{+}18, SHC18, SXXW20, WSM20, YPXH19, YPG20, YK21]. Cellular-V2X [BCM^{+}19a]. CEMR [GXWD19]. Census [RL24]. Center [ACG22, AVA^{+}22, FML^{+}22, GDN^{+}23, LRL^{+}24, LZL^{+}21, QAM17, WLLY24]. Centered [GPRMGP21, SIZD12, SMK^{+}22, VKV^{+}22, YB22]. Centerline [DOB21]. Centers [AJ18, AGC18, SZFC18]. Central [CMS21, KNN^{+}24, MDPHB23, ASFAAV24]. Centralized [KSD^{+}23, ZCZ17, ZPX24]. Centre [GHGL22]. Centric [ACM^{+}19, DNZ22, DNZW22, FP19, Fot20, KC19, LSH21, LH21c, LHL24b, LARPHA23, MZLP22, RYY10, SYB^{+}18, SM19, SNZ21, Tri14, XNZ23a, XNZ23b, ZWZ19]. Certificateless [IZK^{+}23, WHW22, YJSL18]. CF [MKK17a]. CFO [SLS20]. Chain [ACS24, CTM21, FHZ^{+}24, MCS^{+}24, MSC^{+}21, RKT19, YOI19, Yiu21a, Yiu21b]. Chaining [AP22, GGK18, MP16]. Chains [GDS23, GKG19, WADL^{+}24]. Challenge [IP16, LMK18]. Challenge-Based [LMK18]. Challenges [AMAJ12, AJ24, AP22, AAAAM^{+}21, AM23, AYH23, AAAKJ22, ASD^{+}22, Bad23, BAK23, BCG^{+}23, BFL20, BM23, BOKC19, CBRS21, CCA^{+}20, CBB17, DDZ18, ESBE23, GZC21, GTQ^{+}22, GGT^{+}23, HSI^{+}22, HLG21, HQC^{+}24, HSTK23, IZK^{+}20, KTCI21, KSA^{+}21, KIG23, KK23b, KPC^{+}23, KSE^{+}20, KTUF19, Lal14, LAG^{+}18, LMF23, LFGR20, MZS19, MW24, MSIP20, MNWK21, MTM22, MZLP22, PEN23, PSGM22, RAC24, Ray23, ST22, SVK^{+}22, SCC^{+}23, SK23, SAZ18, SEB^{+}19, TPJ21, TMTMB20, VL24, Ves18, ZR15, ZMKR22]. Championship [CR21]. Change [CCCA22, CS11, CFGD20, EM09, MMH16, SLZJ23]. Changed [Den12]. Changes [AZAV22, dSRRMC^{+}23]. Changing [JP13, Pet20, SYGY21]. Channel [AAR^{+}19, CBKL22, DNPC^{+}22, DSMM21, HF20, KOJP13, LWY20, LH21a, LNT^{+}23, LNT^{+}24, ON23, SJG18, SDDB24, VBB24a, VBB24b, YC22, YWW^{+}23]. Channels [KKBD10, MR23a, SOS^{+}23, TH21, WK24, ZSZ18]. Chaos [FKZ22]. Character [YWS22]. Character-Level [YWS22]. Characterising [KDE24]. Characteristics [AAAAM^{+}21, BM19, BOH23, CGMVdUC19, MC12, VKKA17, WKD22, ZVV12]. Characterization [DNPC^{+}22, HOPGRV21, KIB^{+}22, OGB^{+}24, RMLAZOPC21]. Characters [RGM^{+}21]. Charger [CCTC23]. Charging [EAHO21, WHXL18, ZX24]. Charting [RKM^{+}22]. Chatbot [KVVD22, LS23, MV20]. Chatbots [ASG24, Lai24]. ChatGPT [LWW^{+}23, LS23, RT23]. ChatGPT-Based [LWW^{+}23]. Chebyshev [SST23]. Check [KSJZ23]. Checking [BFFZ^{+}12, MSZ23]. Checkpoint [MA23]. Checks [JLW19]. Chemists [KH10]. Chest [DDV22, VR21]. Chicago [DKH24]. Child [SWP^{+}22]. Childhood [HP20, NB12]. Children [LYMG17, VCJAMN22, ZWS20]. Chilean [HOPGRV21]. Chimp [SOK24]. China [HSM19, JHC^{+}20, LLJ16]. Chinese [AFAAM19, Gao21, JLW19, LWX18, LLJ22, LMPT19, WCZ23, WZ18, WJBZ22, Yan17, YWW^{+}23]. Choi [FC19]. Choices [BM14]. Choosing [BM18]. Choreography [ADG^{+}19]. Choreography-Based [ADG^{+}19]. Chronic [Bad23]. Chronological [ESB^{+}20]. Chunks [YL21]. Churn [AAC21, CN24]. CIAA [TPKA20]. CIAA-RepDroid [TPKA20]. CIC [TWW24]. CIC-IoT-2023 [TWW24]. CIDOC [FM21]. CIDOC-CRM [FM21]. Cinema [LL20c]. Ciphertext [SHHM21]. Circuit [VMD^{+}22]. Circuits [AARTSBAE24]. Circular [WZ18]. Citation [RCGSL19]. Citie [BMS20]. Cities [AEM^{+}23, AEA^{+}23, BG24, BZDP19, BFVM14, DT24, GHA^{+}21, MW24, Mal20, McK20, MLL^{+}22, OGR^{+}20, PSC19, PNdC^{+}23, SNKG20, TWM20, WZL^{+}20, XSSA24, YBMK21]. Citizen [DCA16, JP13, MT20, SMK^{+}22]. Citizen-Centered [SMK^{+}22]. Citizens [Gar12, OGR^{+}20]. Citizenship [Mar17]. City [AKYP23, BSM22, DTR22, GKFV24, KSD^{+}23, KAM^{+}19, KSG^{+}24b, LZ11, LMY21, LG23, MMF^{+}19, SSO^{+}19, URH^{+}23, ADG^{+}19, Rot12]. Civic [AGN21]. CKMI [CSIS19]. Class [GSL20a, KM24, KBP22, SDL^{+}19, TWW24, YSEK20]. Classification [ALZ^{+}23, ADAq^{+}22, AAE^{+}22, AF22, BPLBM21, BHSAY24, BG09, BCMPW21, CGnCD20, CT20, GDS23, GB23, GTWN24, HKZ^{+}22, ITH^{+}21, IT24, JJFC22, KVST19, KAS^{+}20, LWX18, LLZL19, LGCM24, LBBPM21, MLWS24, MMD22, MTKS24, MDM21, NAFM^{+}24, PHG^{+}20, SYB^{+}18, SZT23, SWD^{+}18, STS23, TTS24, THL21, YYW20, YZZ^{+}20, YGQ^{+}22, YLPW21, ZLXY19, ZZZ^{+}23, ZGZ^{+}18, ZQL^{+}24]. Classification-Based [ITH^{+}21]. Classifications [WLJM21]. Classified [TMP22]. Classifier [AAM^{+}24, ASLG22, KDD^{+}21]. Classifiers [KAAAA22, MAVS20, PCM22, RPB^{+}17]. Classifiers-Based [RPB^{+}17]. Classifying [HP21, ZJJS^{+}23]. Classmates [ISG20]. Classroom [AARTSBAE24, JPVC21, LF23, PLSF14, hSgLH17, VBWW22]. Clean [Fri13]. Client [YSAY23]. Climate [CS11, CH16, SLY^{+}12]. Clinical [GVY^{+}23, YGQ^{+}22]. Cliques [FAB^{+}22]. Clock [DDZ18, HNHH23]. Clock-Based [HNHH23]. Closed [vSR24]. Closed-Source [vSR24]. Closest [WLCC24]. Closing [AMZ^{+}20]. Cloud [AJ18, AH22, ADM^{+}19b, AMAJ12, AAY21, AAMD21, AWSA23, AK22, AlQ23, ADK^{+}24, AAA^{+}19, AGC18, AD18, BRF^{+}23, BSEL20, BFG^{+}23, BB17, BCM^{+}19b, CD23, CBRS21, CVG^{+}22, CC24, CALHG^{+}24, DGD^{+}19, DCHM16, DVED20, Dua21, DGCK24, ELC^{+}19, ELCS20, FLRS12, FYWS16, FFM^{+}23, FBOC15, GW15, GL21, GDARM18, GGT^{+}23, GMH18, GVV^{+}23, GHGL22, GMMI16, HSI^{+}22, HPB^{+}24, HWZP18, HC21, HWCH12, IIOW16, JII17, Kab23, KVWC24, KCY22, KLZ16, Kra22, LKIL19, LZS18, LFGR20, MSC19, MMN19, MSIP20, NLPV12, OS23, OPGH23, PPKS21, PdRB19, PSTZ23, PNM^{+}20, QAM17, RGMFM12, SAJ24, SHHM21, SOR16, Sul22, TDW24, TXJ23, VM24a, VFFHF19, WCSZ18, WXL^{+}23, YCS17, YJSL18, YA24, ZOAR24, ZQL^{+}24, ZSS^{+}22]. Cloud-API-Based [AAY21]. Cloud-Based [AK22, AlQ23, CBRS21, GL21, PNM^{+}20, ZSS^{+}22]. Cloud-Edge [DVED20]. Cloud-Edge-End [WXL^{+}23]. Cloud-Enabled [CVG^{+}22]. Cloud-Native [Dua21, DGCK24, GVV^{+}23, Kra22]. Cloud-to-Thing [BSEL20]. Clouds [DGD^{+}19, SMK^{+}22, XCSM18]. Clubs [TCG19]. Cluster [DGCK24, JOL22, Jeo20, SJA^{+}23]. Cluster-Based [SJA^{+}23]. Clustered [KKEV17]. Clustering [ARAAA19, ASBC22, AMY24, BG24, BOB^{+}20, BZNB23, CLC^{+}22, CCKH21, CR23b, CGT21, DKH24, FAT19, GOM23, LOL21, TZ22, YW16, YWW^{+}19, ZZZ^{+}23, ZQCC16]. Clusters [PL22b, TXJ23, VLPR20]. CM [PLA17]. CMOs [KPP^{+}20b]. CMRS [PLA17]. CMS [LG20]. CMSTW’2021 [Bod22]. CNN [AWUF19, CT20, GMA20, KM24, LBK22, PRDK22, PW21, SJ21, YLJ^{+}19, ZLXY19, ZWPL19, ZGZ^{+}18]. CNN-LSTM [GMA20]. CNNs [DSMM21]. Co [BMB23c, CGMM22, FHS^{+}10, GLH^{+}22, KKG20, KOJP13, KA23, PMT^{+}23, SMS^{+}24, TLV23]. Co-Authorship [CGMM22]. Co-Channel [KOJP13]. Co-Creation [TLV23]. Co-Design [BMB23c, PMT^{+}23]. Co-Engineering [KKG20]. Co-Existence [SMS^{+}24]. Co-Operation [FHS^{+}10]. Co-Simulation [GLH^{+}22]. Co-Word [KA23]. Coalition [AKT24]. Coarse [WJBZ22]. Coarse-to-Fine [WJBZ22]. Cobot [SU17b, SU17c]. Code [AKAS22, GDS23, KRSF22, MT23, MNH^{+}24, SKHK23, SHLH24, SB23, YFY24, YLJ^{+}19, dCdPC^{+}21]. Code-Based [SKHK23]. Codec [WLYL20]. Coded [AKYP23, SVFV23]. Codes [PTM20]. Coding [BKL^{+}19, CCKH21, EMHF19, GVC13, LW15, MTA^{+}20, NBT22, ZQCC16]. Coding-Aware [LW15]. Coefficients [BZMB23]. Coexistence [SRMJ21]. Cognition [FHS^{+}10, KI20]. Cognitive [AMF^{+}10, AKJB20, Bi14, CZK^{+}17, CS22, FAB^{+}22, FM16, KLTC17, Ler16, Min20, MPCS^{+}23, MPM24, NIK^{+}21, Pri10, RRS20, SAL21, SS23, SN23a, WWYQ19, ZCZ17, ZSZ18]. Coherent [AAR^{+}19, AD20, DDZ18, GLW^{+}19]. Coin [IPL^{+}18]. Coined [YKC19]. CoKnowEMe [AALS21]. Cold [CDM23]. Collaborates [OY19]. Collaboration [CGMM22, Del23, MSR^{+}14, Nil12, RMB^{+}12, YLX^{+}24]. Collaborations [LSN21]. Collaborative [AJA^{+}24, AK19, ADG^{+}19, CBRS21, ERBL12, FGBMG19, GDLG18, GGCY20, HK18, KDEA^{+}23, LMK18, LL23b, LHL24a, MNWK21, ON23, PLG23, PFL^{+}12, Rot12, RGP22, SYGY21, SGDT19, Tag20, VP22, Whe09, YLL21, YPLZ19, ZLOW24, ZQL^{+}24]. Collect [MMQ^{+}24]. Collecting [HP21]. Collection [APB^{+}16, IMR^{+}18, LA23]. Collections [CFV19, Cop14, LWBM22]. Collective [ISG20, KPP^{+}20b, MS20, OC20, SAL21]. College [Gao21, hSgLH17, ZZZ^{+}22]. Collision [HJA18, KHAM24]. Collisions [YK21]. Collocation [Bea10]. Colluding [TH21]. Colombia [FSG^{+}16, MFO^{+}20, OLMBRCRC23]. Colombo [AKYP23]. Colony [GW15, LW19, RAU^{+}24]. Colony-Based [RAU^{+}24]. Color [GB20, MCM^{+}10]. Color-Avoiding [GB20]. Color-Favored [GB20]. Colored [DCHM16]. Coloring [MCM^{+}10]. Colors [LWBM22]. Combat [CBRS21, THV^{+}22]. Combating [SHC24, VMGT22, ZXXB22]. Combination [LDd^{+}24, LWW^{+}24]. Combinational [YY17]. Combine [MSAA22]. Combined [ANHMS24, HSL^{+}23, LMPT19, LBK22]. Combining [ASL22, CMP^{+}19, HLC24, HYLP19, JGY^{+}24, LS15, MGBG21, SU17a, TZS^{+}21, Whe09, ZBK^{+}24, ZLXY19, ZLT22]. Come [DGD^{+}19, JDH21]. Comfort [ZGL20]. Coming [Tre22]. Commands [AM21]. Comment [BTB22, KLKA19, FC19]. Comments [LY19]. Commerce [AZS23, CMP^{+}19, GTM22, Khr20, TLC15, WYX18, ZSRC23]. Common [NDKBL19, SMS^{+}24]. Commonly [JGV^{+}21]. Communication [AK18, ADM19a, ASBC22, Alb21, ASM^{+}24, AF21, AMZ^{+}20, ASD^{+}22, ACA^{+}23, AS24, BCNS20, BF18, BCM^{+}19b, BBK^{+}20, CISM22, FM16, FLM23, FHS^{+}10, GPMO20, HAL18, HGL^{+}23, HS20, IHCSG^{+}24, IPP24, Jal19, KZRH22, KALY17, KL23, LH19, LSJ19, LWW22, LLCL23, LA20, MCCOCE24, NMH^{+}21, OFK19, PLA^{+}21, QQM24, RRS20, SDDB24, SHC22, SGR^{+}13, SOS^{+}23, SSS^{+}19, SST23, SM19, SXXW20, SF22, THT18, TL24a, TTN24, dSVJdALRfS21, WUH^{+}23, WY19, YT19, YT18, YK21, ZZZ^{+}22, ZZY19, dCdZCC24]. Communication-Efficient [ASBC22]. Communication-Historical [GPMO20]. Communication-Traffic-Assisted [LLCL23]. Communications [ASL17, AP19, CRMT23, CLH^{+}23, CFP17, CFP19, HDBD23, HYL13, ISLARP^{+}22, KFP10, Lee18, MCH^{+}20, PSC19, PAM^{+}24, PSM^{+}18, QSML24, SHC18, SHN^{+}19, TSHO24, WSM20, ZAS^{+}24]. Communities [BBA21, FAB^{+}22]. Community [BdBCCG21, BLW^{+}17, Cop14, GSD^{+}17, Hua15, KYG22, LN19, NZZ13, PZC21, SL24a, SHA^{+}21, WZC19, WXL22, ZVV12, Zha19, ZGH13]. Community-Driven [ZGH13]. CoMP [ENT21]. CoMP-Enabled [ENT21]. Compact [CLZL18]. Companies [RSV^{+}24, RH24]. Company [SLS20, ZCdRR^{+}23]. Comparative [AGO23a, AMAE22, Bod22, COL22, GRA21, Jal19, KK21, KK23a, LS23, MKPG22, MF20, NLD^{+}24, RCGSL19, SAR24, SHB23, SFSN^{+}24, TBT^{+}23, ZLOW24]. Comparing [KKTK20, PSTZ23]. Comparison [CTM19, GM16, GFPVLR^{+}23, HF14, HDT22, LMY21, MS23, NZZ13, Sta23, SNS22, TLK^{+}20, VM24b, ZSS^{+}22, vSR24]. Compensated [LMH19]. Compensation [AD20, MMN19]. Competence [GDRCDBLÁ23, dPCGMC20, Pet20]. Competences [VCPT22]. Competencies [DHG21, FMA^{+}23, SSOÁ^{+}16]. Competitive [MATP19]. Complete [CC22b, Mou24]. Completion [ZC21]. Complex [BB23, CPI^{+}18, DLX^{+}19, KRSF22, Min20]. Complexity [BR21, FSG^{+}16, MTA^{+}20, SP22, WZL^{+}20, ZLT22]. Compliance [BFFZ^{+}12, GSB^{+}23, PYC21, RSV^{+}24]. Compliant [GdSdC23, vSR24]. Component [GFPVLR^{+}23]. Components [LL20a, TGS^{+}19]. Composite [ZZWP19]. Composition [ALB24, BBH^{+}24, DCHM16, KKT20, MHS23, MKSV^{+}19, RRMI22, SWJ20, SKFM24]. Composting [NPM^{+}18]. Comprehensive [AP22, AAAT21, ACG22, AMCI17, AEA^{+}23, AYH23, ASM^{+}24, CSIS19, CGRV20, DJC^{+}22, KKG20, MSZ23, MDT20, NNT22, PNS23, RGB^{+}24, SSDS23, TPP^{+}23, VLK24, YJSL18, YHN21]. Compressed [CCKH21]. Compression [KL21, PTZM22, SWC^{+}23]. Compressive [GTC19, LXY17]. COMPRIME COMPRI MOv [dCMA21]. Computation [DeD16, LLL21, LZY^{+}24, SZW^{+}23, TLC15, ZWL^{+}24]. Computational [AK19, ATY20, BOKC19, DVED20, HWCL17, Let16, LAG^{+}18, PSTZ23, RPHJ11]. Computations [PH16]. Compute [GVC13]. Computer [BPLBM21, CB16, Del21, KPRP24, MMQ^{+}24, MSP21, Rei23, TA15, Wan23]. Computer-Mediated [TA15]. Computers [Rah22]. Computing [AH22, AAMD21, AAAAM^{+}21, AAA^{+}19, AGC18, ARC^{+}19, And21, AD18, AVE^{+}22, ALDS19, BMS20, BRF^{+}23, BFY22, BZP^{+}21, BMJ^{+}22, BB17, CD23, CPM^{+}19, CC24, DGD^{+}19, GW15, GHZY23, GMH18, HSI^{+}22, HC21, IBBA20, IIOW16, JMJ^{+}23, JHBW24, KLZ16, LW21, LZY^{+}24, LFGR20, MSC19, MDT20, MC24, MATP19, NBT22, NLPV12, PNS23, PAK^{+}23, Pec18, PdRB19, Poz24, RBVV22, RAC24, RAR^{+}23, RRdBSS22, RR24b, BNK^{+}23, SC23, SK23, SFSN^{+}24, Sul22, SEB^{+}19, TDW24, TTKZ19, TL19, TRST24, TCYZ22, VdHH^{+}19, VKM^{+}19, WSF23, XQLZ19, YPZ18, ZOAR24, ZWL^{+}24, ZQL^{+}24]. Concentration [GBDDVS23]. Concept [BMB23c, ESB^{+}20, GPCP12, GKG19, KGRP21, SLZJ23, SGR^{+}13, TPJ21, VLC^{+}19]. Concepts [LA20, RMB^{+}12]. Conceptual [AAGB^{+}20, BRBC14, FAB^{+}22, For14, PRS22, RRMI22]. Concerning [FZ15]. Concerns [GM16, MKKG19]. Concomitant [LQFI17]. Concrete [TPJ21]. Condition [ZLS19]. Conditions [HGF21, MLK22]. Conference [Bod22, MNWK21]. Conferences [Sem11]. Conferencing [MN21, TTO^{+}19]. Confidence [DHEK19, dSMdO22]. Confidential [AJT^{+}23]. Configurable [MPL^{+}21]. Configuration [BN14, YBO^{+}23]. Configurations [AWSA23]. Confirmation [PLL11, UANU20, VM24a]. Conflict [AM24, DeD16, MCM^{+}10]. Conflicting [YLYL22]. Conflicts [Kos16]. Conformal [WLHR13]. Conformity [HC19]. Congestion [ALST20, DNZW22, HA23, LRL^{+}24, LNH23, LIM23, MMMH19, PKB^{+}24, PKS^{+}24, SPK24, SNZ21, SZFC18, WWD21, ZLL^{+}21]. Conjecture [PLL11]. Conjectures [Min20]. Conjunction [DGGS19]. Connect [AYH22]. Connected [ACG^{+}24, DDV22, KSE^{+}20, MCH^{+}20, RZQS18, You23, ZHZ^{+}19, ZAS^{+}24]. Connecting [ZSRC23]. Connection [KMK24]. Connections [Tsi14]. Connectivity [EARP^{+}24, GHFM20, NMH^{+}21, VC19, Zaf12]. Conscious [CDD^{+}21, RSL^{+}24]. Consensus [CCA^{+}20, CAS^{+}20, KK21, KK23a, LLW^{+}17, QXC^{+}21, SVGGO24, WLL20, XCW^{+}22, XNC21]. Consensuses [AMEF21]. Consent [Pic21]. Consequences [BM19, Ler16]. Conservation [NPB12, NMT23]. Considerations [LBB23, RRAGMMGG20]. Considered [SE15]. Considering [CZK^{+}17, HZW^{+}20, HOPGRV21, MKY24, TGS^{+}19, UDF^{+}19, WZLW19, YLK^{+}19, ZY15]. Consistency [DCB19, FLLG24, LWZ^{+}21, MLWS24]. Consistent [ZLY^{+}20]. Consolidation [AJ18, XCSM18]. Consortium [KPP^{+}20b, PYC21]. Constellations [MR23a]. Constrained [FRR^{+}24, GMD^{+}22, SA21, dOB19]. Constraint [XQLZ19, ZLZ^{+}24, ZY15]. Constraints [AD18, FML^{+}22, PP22b, SWGL19, SSX20, TCH18]. Construction [BFFZ^{+}12, CR23b, CT24a, DG18, DWY^{+}19a, LLJ16, PLA17, WADL^{+}24, PLA17]. Consumer [ASAH^{+}23, EHJ17, Khr20, RGD24, TN20]. Consumption [FGBMG19, HLC24, HAL18, NMT23, SMG13, TBT^{+}23, WLHR13, ZWZ19]. Contact [GFPD21, RRLT20]. Container [SHB23]. Containerization [SFSN^{+}24]. Containerized [YA24]. Containers [YCS17]. Contemporary [HNK^{+}21]. Content [AA19a, BN14, Cap12, CSL^{+}20, CDD^{+}21, DWHY24, GZ23, HSSPTM^{+}23, IP16, KHM24, KC19, KA23, LNA17, LHD^{+}24, LH21c, MAW^{+}19, MOAI17, MSC^{+}21, PVM22, PDI21, SB22, SYB^{+}18, SSS21, SS22, SHLH24, VLPR20, VKV^{+}22]. Content-Caching [SB22]. Content-Centric [KC19]. Content-Oriented [SHLH24]. Contents [BCG10]. Contests [Yan19]. Context [AAGB^{+}20, AAARZ24, ALLR16, AYH^{+}21, AF22, CZL21, CMR24, EKG24, FZ15, FDN^{+}24, GMH18, GRJ^{+}19, GNGHEM21, HBFJ20, HY22, LCMV17, LHSS12, MEO18a, MMS^{+}19, MXL^{+}21, MSJ^{+}23, MKHR21, PCROB21, PDTM15, PPGC16, PSC19, PW21, BNK^{+}23, SMK^{+}22, ZZ18, vSR24]. Context-Aware [ALLR16, CMR24, MEO18a, MMS^{+}19, PDTM15, BNK^{+}23, SMK^{+}22]. Context-Based [LHSS12, PPGC16]. Context-Induced [AYH^{+}21]. Contexts [HL19, PL22a, WL22]. Contextual [DB23, KHN^{+}22, LZY22, SH21]. Continuity [LPVM24]. Continuous [DHDAu22, GTWN24, LG20, TL24a]. Continuum [BSEL20, GGT^{+}23, PSTZ23, SFSN^{+}24]. Contract [LT24b, PTM20]. Contract-Based [LT24b]. Contracts [BHSAY24, Fra24, GLD^{+}18, KPP^{+}20b, KPP^{+}20a, LHL24a, ZL22]. Contrastive [ZZZL23a]. Contribution [TL19]. Contributions [GS12, KTCI21, KST24, NZZ13]. Control [ATFMOU^{+}24, APB^{+}16, Aln22, AAA^{+}20, AKJB20, BM20, BCFP21, CSIS19, CFPP21, DCG12, DL19, DNZW22, EW19, GS23b, GAMMPCRÁ24, GGK18, GRRZ18, HA23, UGJA18, JGY^{+}24, KL23, LRL^{+}24, LB23, LNH23, LXQ24, LHSS12, LIM23, MSN13, OOM^{+}23, OF23, Oza23, PKB^{+}24, PRS22, PKS^{+}24, QNJ24, RRP17, RGP22, SU17a, SZL21, SPK24, SNZ21, SZFC18, SSX20, Tra18, UFK^{+}10, WWD21, WLCS17, WZX20, YWLY22, YBO^{+}23, ZBK^{+}24, ZJJS^{+}23]. Controllable [KLG^{+}24]. Controlled [MTA^{+}20]. Controller [DNKDN^{+}24, RAJJ23, SHBS19, WKW18, YLK^{+}19]. Controllers [SDM21]. Controlling [HKZ^{+}22, OTR^{+}20, PKCF21, ZGL20]. Controls [AA21b]. Convergecast [WC17]. Converged [FMMS21, MKvD11]. Convergence [MZLP22]. Converging [PMGG21]. Conversational [MMS^{+}19]. Conversations [KHC^{+}24, MT20]. Conversion [LMLW18, LMH19]. ConvNets [DWY19b]. Convolution [TDT^{+}21, WX22]. Convolutional [ALH24, BFY22, CBM^{+}20, CYC^{+}21, CH23, CKSG21, DDD^{+}22, EVCL21, GTC19, GR22b, JTA^{+}21, LNT^{+}24, LQY^{+}19, SSM^{+}24, SYZ19, VGK21, VMGT22, VR21, WYL18, YCHG18, YST24, dOB19]. Convoy [SLB24]. Cooking [ZLOW24]. Cooling [PC23]. Cooperation [GDLG18, GFL21, HRAA19, SU17b, TEE10]. Cooperative [CFP17, DH22, FTHY21, KALY17, LYD22, MSK^{+}18, MKPG22, MPCS^{+}23, MPM24, SU17a, SU17b, SSM^{+}24, ZZY19]. Coordinate [ZQCC16]. Coordinated [SWY^{+}18]. Coordination [Pet11, TGS^{+}19, TFKY19]. Coping [BM19, BBM20, HJA18]. Coprocessor [QMN19]. Coprocessor-Based [QMN19]. Coproduction [DCA16]. Copy [LSC^{+}17]. Copyright [DN20, Fra24, ZWWW17]. Core [DGCK24, FWVF19, KZRH22, MPL^{+}21, SDS^{+}22, YPG20]. Cores [MLPH23]. Coronary [DOB21]. Correct [SOSC^{+}16]. Correction [LRD23, LJR24, VBB24b]. Correlated [AWSA23, SVFV23]. Correlates [HRKL21]. Correlation [CLC^{+}22, GKK^{+}19, MSAA22, MSJ^{+}23, PL22a, XNZ23a]. Correlations [WKD22]. Cosine [OY19]. Cost [AHMI22, BP24, BFG^{+}23, CTP^{+}22, CZK^{+}17, DVSBV^{+}24, EVCL21, GDLG18, HC21, LLJ16, MDT20, MKK17a, PR11, Sul22, TPD^{+}20, VL18a, dVBA23]. Cost-Aware [EVCL21, MKK17a, Sul22, VL18a]. Cost-Profiling [dVBA23]. Costing [OS23]. Costs [CPM17, MGV17, RPHJ11]. Council [RBSMRN21]. Counselling [WBK^{+}23]. Counter [SZL21]. Counterfeiting [Yiu21a, Yiu21b]. Countering [AAAR24]. Countermeasures [ADS20, LL23a, OB13, RSV^{+}24]. Counting [DIE^{+}21, SC23]. Countrie [MDPHB23]. Countries [ASFAAV24, AT14, BZDP19, CR21, GMRRRQ^{+}16, GPK^{+}20, JII17]. Country [DTL^{+}21, GPLK22]. Counts [RCGSL19]. Coupled [ZWPL19]. Course [Del21, HRGQHOA21]. Courses [AGO^{+}23b, PLA^{+}21]. Courts [CBB17]. Cousins [BPBF12]. COV [CR21]. Cover [Mah17, RFZ22]. Coverage [ADAB21, Cro19, GM20, GHFM20, KDJ24, SSS^{+}19, WX24, WLWZ23, YPG20]. COVID [ABSG21, AAN^{+}21, AA21a, BCT21, CBRS21, CYC^{+}21, CQWC22, CTM22a, Del21, ESB^{+}20, FFVP21, GNGHEM21, GFPD21, HGF21, Kra20, MPG20, OFV21, PMP^{+}22b, RRLT20, RFSOHMSB21, SG21b, TC21, WKD22, YWLY22, YHN21, dCMA21]. COVID-19 [ABSG21, AAN^{+}21, AA21a, BCT21, CBRS21, CYC^{+}21, CQWC22, CTM22a, Del21, ESB^{+}20, FFVP21, GNGHEM21, GFPD21, HGF21, Kra20, MPG20, OFV21, RRLT20, RFSOHMSB21, SG21b, TC21, YWLY22, YHN21, dCMA21]. COVID-19-Robocov [PMP^{+}22b]. COVIDNet [MTF^{+}21]. CP [WZX20]. CP-ABE [WZX20]. CPA [SZL21]. CPS [PGSL20]. Craft [MBAS20]. Cram [LL20b]. Crashes [Chu23]. Crater [WBG12]. Crawling [PCF^{+}14]. Create [SRS23, Whe09]. Created [Kra20]. Creating [HDL^{+}14, HPC12, LOL21, SC20, SWP^{+}22].

Creation [CYO^{+}17, Goe12, LWW^{+}23, MCT^{+}23, PLA^{+}21, TLV23, WBK^{+}23]. Creativity [ATY20]. Credentials [FSXP22]. Credibility [CLM^{+}21]. Credit [BARL24]. CRF [XGN19]. Crisis [MV20]. CRISOLA [FMNS11]. Criteria [ANHMS24, DDPVP18]. Criterion [Rók23]. Critic [BBBHA24]. Critical [AOI^{+}23, CCC^{+}24, DPC^{+}24, DTL^{+}21, HBMS20, JSAO^{+}23, Roy14]. CRM [FM21]. Croatia [OBK23]. Crop [KVWC24]. Cross [AWSA23, ALL24, BM19, BRP^{+}13, CZZZ22, CQWC22, DGCK24, IT24, KSE^{+}20, LJWX23, MKKS18, MSR^{+}14, QDL22, SKT^{+}19, WZC19, WKD22, ZZYC21, ZSRC23]. Cross-Border [KSE^{+}20]. Cross-Cluster [DGCK24]. Cross-Correlated [AWSA23]. Cross-Correlations [WKD22]. Cross-Disciplinary [MSR^{+}14]. Cross-Domain [ALL24, CQWC22, QDL22]. Cross-Layer [IT24, SKT^{+}19]. Cross-Modal [CZZZ22]. Cross-Network [BRP^{+}13]. Cross-Platform [ZSRC23]. Cross-Project [ZZYC21]. Cross-Sectional [BM19]. Cross-View [LJWX23, MKKS18]. Crossing [MAW^{+}19, PCC^{+}18, VCPT22]. Crowd [Lal14, NPB12, SVGGO24, YPZ18, ASN22]. Crowd-IoT [ASN22]. Crowd-Sourced [SVGGO24]. CROWDMATCH [AKT24]. Crowdsensing [KK20, Pil18, SSO^{+}19]. Crowdsensing-Based [KK20]. Crowdsourced [Goe12, NZZ12]. Crowdsourcing [AKT24, ASN22, DQ18, EH21, LGCM24, PG14, PLSF14, Pil18, VKV^{+}22, Yan19]. Crowns [LBK22]. Cruise [CTP^{+}22]. Cry [YST24]. Crypto [DMO^{+}19, SDS^{+}22]. Crypto-Currency [DMO^{+}19]. Cryptocurrency [FCG22, RM23]. Cryptograph [Li23]. Cryptographic [MLPH23, MKK17b, RPHJ11]. Cryptography [AOI^{+}23, AZS23, CISM22, FLM23, GK21, HEP^{+}11, LBA^{+}24, TSDG22, WUH^{+}23]. Cryptography-Based [WUH^{+}23]. CSI [LLYL22]. Cues [PB23]. Cultural [AIPV21, FM21]. Culture [GNGHEM21, GSD^{+}17, PVKG23, SOSC^{+}16]. Cumulative [DTTK19, YB22]. Curating [TG13]. Curation [WKPC18, ZGH13]. Currencies [CMS21, MDPHB23]. Currency [DMO^{+}19, GANT21]. Current [HE20, KTCI21, KXZ^{+}24, KPC^{+}23, PSGM22, SRMJ21, VTPT24, VSEH^{+}21, YBMK21]. Curricula [SSOÁ^{+}16]. Curriculum [VC23]. Curriculum-Based [VC23]. Curve [AKDL24, CISM22]. Custodianship [GANT21]. Customer [AAC21, ASG24, CCD^{+}21, CN24, FRE^{+}22, HC19, KPP^{+}20a, KA21, LYP^{+}19, ZS18, ZY15, ASWAE23]. Customers [FN23]. Customized [AA24a]. Customizing [SSS^{+}19]. Cutoff [AM21]. Cutting [ZYWS24]. CvAMoS [DB23]. CWM [KH10]. Cyber [AS19, BSM22, BCM^{+}23, CCC^{+}24, CPP^{+}20, DGD^{+}19, DSM21, FSG^{+}16, GW13, GTQ^{+}22, Gro22, HBFJ20, JWD^{+}24, JSAO^{+}23, Kos16, LH19, Lee20, LGV13, MAW^{+}23, MIK^{+}23, MKHR21, MHL^{+}23, NPMPM23, OOM^{+}23, OCCB20, PMP22a, PB23, PGSL20, QLR^{+}22, QQX^{+}24, SKI^{+}19, SKFM24, Suf23, SYLL21, TKT^{+}18, VYN19, WLH^{+}17, YOI19, dCdG20]. Cyber-Attacks [OOM^{+}23]. Cyber-Physical [BCM^{+}23, CCC^{+}24, CPP^{+}20, DSM21, GW13, GTQ^{+}22, JSAO^{+}23, LH19, MKHR21, MHL^{+}23, OCCB20, PMP22a, PGSL20, TKT^{+}18]. Cyber-Sextortion [PB23]. Cyber-Storms [DGD^{+}19]. Cyberattack [KHAA21]. Cyberbullying [ER21, HHM^{+}23, MF20, OMTFFL21]. Cybercrime [AAAR24]. Cyberphysical [KKG20]. Cybersecurity [CR23a, FKF24, FSI^{+}24, JWD^{+}24, KPT^{+}24, KKG20, Lee20, Rah22, SACP24, Tag20, TSZ^{+}18, UW21].

Cyberstalking [BM19, BBM20]. Cyberworld [VDKL21]. Cycling [MMQ^{+}24]. Cyst [IKS^{+}21]. Czech [MKK^{+}22].

D [DOB21, LB23, MKPG22, SSM^{+}24]. D-GNSS [MKPG22]. D-UNet [DOB21]. D2D [ASL17, HWCL17, HAL18, KSG24a, MOAI17, PSM^{+}18, SSF^{+}18, SHC22, SHC18, SHN^{+}19, Wan17, WSM20, ZWL^{+}24]. D2D-Assisted [ZWL^{+}24]. D2D-Enabled [KSG24a]. D2D-Enhanced [MOAI17]. DA-GAN [CZZZ22]. DAG [GPPB^{+}21]. DAG-Based [GPPB^{+}21]. Daily [ADCM24, CFG20]. Dairy [CNC^{+}21, NSJGN21, VLMP21, VLBRMP21]. Damage [BZP^{+}21, KHAA21]. DASH [NBT22]. Dashboard [VI16, dCMA21]. Dashboards [Smi15]. Data [AJ18, AKYP23, AAAHA22, AAAIA24, AFB23, AEM^{+}23, AWBS23, ACG22, AVA^{+}22, AAAAK24, APB^{+}16, ACD^{+}22, Aln22, AGC18, AAA^{+}24b, ABH24, ARC^{+}19, AH21, ACA^{+}23, ACAP20, BM20, BMS20, BCSB24, BMB23c, BAR^{+}21, BNJ24, BBGP19, BKCL23, BC19, BMS24, BHM^{+}19, BB23, BJL^{+}22, CCTV24, CMO^{+}21, CBRS21, CMP^{+}19, CFV19, CCKH21, CGLR19, CWC^{+}22, CSZ^{+}24, CDLY24, CH16, CGMVdUC19, CAL23, CL19, CC22b, DGADE14, DH22, DGHH^{+}20, DVED20, DCL24, DGF23, EHJ17, EvdMB20, EKKS24, EKG24, EAAA^{+}23, FSI^{+}24, Fio10, FMNS11, FML^{+}22, GSL20a, GSL22, GPPB^{+}21, GPMO20, GDN^{+}23, GL21, GKFV24, GM16, GVY^{+}23, GRA^{+}23, GHGL22, GOM23, GMB^{+}21, GM19, GRRZ18, GMS^{+}20, HCXH16, HS18a, HWZP18, HKZ^{+}22, HBM^{+}21, HN18, HWCH12, ISLARP^{+}22, JXA19, Jeo20, JP20, JT23, KST19, KGK^{+}24, KS14, KCT^{+}24, KSW21, KBF23]. Data [KPKM23, KPCS13, KBF^{+}22, KVL^{+}22, KH19, KSJZ23, KHAM24, KV23, LPPG22, LB24, LDM22, LG20, LLY23, LLY24, LRL^{+}24, LTTS21, LNT^{+}24, LW24, LZL^{+}21, LT24a, LFM^{+}19, LG23, MXT^{+}16, MGM^{+}23, MSC^{+}24, MEH^{+}22, MMH16, MKSJ22, MS16, MT20, MSR^{+}14, MATP19, MCPLBH^{+}24, MSFM16, MTK^{+}24, MKHR21, MP13, NZZ13, NJH^{+}24, OF23, PG23, PRDK22, Pec18, PLA^{+}24, PCM22, PST^{+}21, PYC21, Pil18, PHG^{+}20, PHGZ20, PPM21, PR20, Por21, PH16, PM17, PNM^{+}20, QMN19, QA24, QAM17, RBG^{+}22, RSV^{+}24, RAA^{+}18, Roy14, SB22, SJA^{+}23, SVGGO24, SHHM21, SSR17, SFW23, SN23b, SMR24, SZT23, SLSK21, SH21, Sta23, SOR16, SAZ18, SZFC18, SZZL23, TAD23, TG13, TRB22, TC21, Tos23, TZE^{+}21, UDF^{+}19, UJ22, VPCE20, VR21, Ves18, VDSK23, VCB14, WLL^{+}19, WWLZ21, WL21, WLLY24, WLCC24, WKF^{+}21, WC20, XDH24]. Data [XSM22, XNZ23a, YJSL18, YL21, YSEK20, YWW^{+}19, ZCG^{+}24, ZZ21, ZQL^{+}24, vSR24]. Data-Driven [ACD^{+}22, AAA^{+}24b, EAAA^{+}23, JT23, Ves18, WKF^{+}21]. Data-Enabled [MMH16]. Data-Intensive [CMO^{+}21]. Data-Link-Layer [CAL23]. Data-Scheduling [XNZ23a]. Database [ARM20, EvdMB20, GM19, JT23, KHAA21, WJ21]. Databases [DGADE14, DCB19, FK21, GRG21, SO24]. Datacenters [CZY17, IBBA20]. Dataset [BGD09, GFPVLR^{+}23, HP21, LNT^{+}24, LPA^{+}22, MAVS20, NLD^{+}24, RDASB22, TBT^{+}23, VKKA17, WBK^{+}23, TWW24]. Datasets [AAAT21, BMB^{+}23a, BS17, LCC^{+}24, TMP22, YAV^{+}21]. Dataspaces [GHFL24]. DataStream [MGM^{+}23]. Dating [VYN19]. Day [CS22, HCXH16, MNH^{+}24]. DB [FBDR23]. DB-Based [FBDR23]. DBA [Rók23]. dBFT [CCA^{+}20]. DCAM [DNZ22]. DCT [SS17]. DD [AVE^{+}22]. DD-FoG [AVE^{+}22]. DDoS [AKM^{+}23, AAA^{+}24b, FMMS21, HK18, HP16, KAAAA22, MA23, WKW18, YMH22].

DDPG [PKS^{+}24]. DDPG-MPCC [PKS^{+}24]. DDS [DGF23]. De-Clustering [CCKH21]. De-Duplication [KCT^{+}24]. De-Identification [LTTS21]. Deadline [AD18, XQLZ19]. Deadlines [oRL19]. Deadlocks [Cap23]. Dealing [Cap23]. Debate [VLPR20, OC20]. Debt [PLOT23]. Decentralized [AJA^{+}24, AAF^{+}22, ACA^{+}21, BM20, BKCL23, CCCPV^{+}24, Cal24, CZK^{+}17, CLT^{+}22, FWLY23, FSXP22, FK24, GHZY23, KPP^{+}20a, LWW22, MT17a, OF23, SCA^{+}19]. Decentralizing [CTNS21, Yiu21a]. Deception [ITZ20, PC24, RFSOHMSB21]. Decision [ANHMS24, ARM20, AFS^{+}23, BM18, CGRV20, CHMZ21, DCPG21, FMA^{+}20, LLL21, MHSH16, MKPG22, MLL^{+}22, Orl22, PR12, Rók23, SWJ20, SZC19, VP22, YKC19]. Decision-Making [MHSH16]. Decisions [AK19]. Declarative [OTR^{+}20]. Decoding [BG24, BHS24, MTA^{+}20]. Decoding-Complexity [MTA^{+}20]. Decorrelation [PP22a]. Decorrelation-Based [PP22a]. Decoupled [MEO18a]. Decoupling [DMO^{+}19, LWY20]. Decoupling-Based [LWY20]. Deducing [EHJ17]. Deep [AKM^{+}23, AN22, Alb19, AAAR24, AAG^{+}22, AAA^{+}24b, AKM22, BHSAY24, BHS24, BCMPW21, CBM^{+}20, CLC^{+}22, CHMZ21, CWC^{+}22, CASN21, Chu23, CT24b, CCYL24, DXL^{+}22, DWHY24, DKV24, DIE^{+}21, FWLY23, FHZ^{+}24, FFP^{+}22, FM20a, GWF^{+}23, GRA21, HHM^{+}23, HBM^{+}21, JXA19, JOL22, JMJ^{+}23, KM24, KPH^{+}24, KTD24, KRSF22, LWX18, LLZL19, LWH24, MCS^{+}24, MAW^{+}23, MMD22, MKSJ22, MCR^{+}23, MBLC22, MSC^{+}21, MA23, MHL^{+}23, NB24, NLD^{+}24, NAFM^{+}24, PLL20, PBP^{+}24, PP22a, PHG^{+}20, Pri23, QSML24, RRT20, RPB^{+}17, RKS^{+}21, SJ21, SGG^{+}22, SESO24, SSR^{+}21, SOK24, SPK24, SWD^{+}18, SF22, TRHU23, TLK^{+}20, TCYZ22, VGK21, VR21, VMD^{+}22, WWD21, WL22, WLLY24, WYL18, YBO^{+}23, YMH22, YLJ^{+}19, ZL20, ZZZ^{+}23, ZZSX19, ZHH24, ZXWZ21, ZLZZ23]. Deep- [DWHY24]. Deep-Learning [CHMZ21]. Deep-Learning-Based [Chu23, HHM^{+}23]. DeepDiver [RKY20]. Deepest [GCA^{+}22]. Deepfake [FLC21, KYS21, QDL22]. Deepfake-Image [FLC21]. DeepLabCut [FSG22]. Deeply [RKY20]. Defeating [HVS17]. Defect [Alb19, ZZYC21]. Defects [VMD^{+}22]. Defence [LGV13]. Defend [FHXL19]. Defending [AV23, MSC^{+}24]. Defense [AR23, LLB24, ZLZ^{+}24]. Defenses [BB17]. Defensive [PH20]. Deferred [KH19]. Deficit [SIJA10]. Defined [AA21c, BCFP21, BM14, EW19, FML^{+}22, Gal24, GLZL24, GTQ^{+}22, GGW18, IZK^{+}20, LZL^{+}23, LXL^{+}19, MZS19, MTKS24, MP16, NAFM^{+}24, Pap20, RAJJ23, SN23a, SMN22, STS23, SKT^{+}19, SAM^{+}18, TRC^{+}21, VL20]. Defining [BCS24]. Definitions [BFL20, VDKL21]. Degradation [SMVP21]. Degraded [QQM24]. Degree [CEZC19, DH18, DTTK19, GRM23, LHW18]. Delay [AMK^{+}21, BKL^{+}19, CL16, KKP22b, KST24, MKK17b, QWR18, USS19]. Delay-Tolerant [KKP22b, KST24, USS19]. Delegation [WYW^{+}22]. Deliberation [SK24]. Delineation [LBK22]. Deliveries [EAAA^{+}23]. Delivering [Nil12]. Delivery [BNJ24, DWHY24, GZ23, LNA17, dMLMS22, LCW^{+}18, MKY24, MSC^{+}21, PP22b, SH20]. Delta [SSS^{+}19, PPP^{+}23]. Demand [AYH^{+}21, DTTK19, JRP21, KT22, Khr20, KAL^{+}19, SWZ18, SWY^{+}18, dSVJdALRfS21]. Demand-Responsive [SWZ18, SWY^{+}18]. Demands [Pin10]. Democracy [Rot12]. Demographic [KDKG22]. Demonstration [GKG19, LWBM22]. Denial [AAM^{+}24, BB17, SGDT19, STS23]. Denial-of-Service [BB17]. Denied [FKK20]. Denoising [TDM^{+}24]. Dense [HSSPTM^{+}23, IUIL23, KES22, LTY18, LWY20, LL20a, MGV17, ZZZL23a]. Dense/Ultra [MGV17]. Dense/Ultra-Dense [MGV17]. Densely [OFK19, YCHG18]. Densely-Populated [OFK19]. Density [BOK^{+}21, FPGO24, HBMS20, PZC21, YPXH19, YW16, ZQCC16]. Density-Based [HBMS20]. Dependability [BBGP19, Tri14]. Dependencies [GGD23]. Dependency [LZY22]. Dependent [JLGF19, LJR23, LJR24, ZGH13]. Deploy [GMD^{+}22]. Deployed [KOJP13, RS17]. Deploying [YA24]. Deployment [CPP^{+}20, CCTC23, FHZ^{+}24, GMGF24, GRM23, MKSV^{+}19, MS23, MSC^{+}21, NPB^{+}19, SGLZ20, SFSN^{+}24, WX24]. Deployments [ADAB21, BOK^{+}21, DSB23, KIB^{+}22, SKKS22]. Depressive [BBM20]. Depth [DHEK19, GZS^{+}24, PMT^{+}23, RKY20, TCH18, ZDP^{+}22]. Derived [ALZ^{+}23]. Description [GL12, LWBM22, SZT23]. Descriptive [SRS23, SSS21]. Descriptors [NLD^{+}24]. Design [AAAKJ22, AAKS23, ALDS19, BMB23c, BM14, CSIS19, CFPP21, Ciu19, CFX20, DDV22, DL10, DCA23, FSC^{+}23, GHZY23, GPRMGP21, HDK^{+}12, Hua15, JLZ18, KYT^{+}23, KG18, KPCS13, LLHW20, Let16, LBB23, LL20c, LN19, LYMG17, MLPH23, MZS19, MPT23, MV20, MMH16, MBAS20, NPB^{+}19, Orl22, OR24, PR11, PMT^{+}23, PB23, QAQA21, RIS^{+}17, RSF23, Sem11, STG^{+}20, SH20, SIZD12, SWZ18, SWY^{+}18, Tag20, THV^{+}22, Tri14, WK24, WYL^{+}21, YB22, YQ24, YXL^{+}21, ZN22, ZYWS24]. Design-Based [DCA23]. Design-Two [HDK^{+}12]. Designing [AA21c, CHH^{+}16, KID^{+}15, MNWK21, PMP^{+}22b, RS17, RSF23, ZZZ^{+}22]. Designs [SRS23]. Destination [OT16]. Destroy [KMK24]. Detect [AAM^{+}24, ASF^{+}23, CLTP22, KDD^{+}21, LMLW18, MRM^{+}24, WKW18, XLJ^{+}22].

Detectable [YLYL22]. Detecting [AAA^{+}24a, AA21a, AASAI22, Chu23, DSB23, KAAAA22, MKR22, MSAA22, RD18, SE23, SOE^{+}24, SFEK18, TS22, YC22]. Detection [AdMTJ20, AKM^{+}23, AZS23, AGO23a, AN22, AKAS22, ASF^{+}23, ALL24, AAAR24, Alh22, AAG^{+}22, AR23, AA23b, AAA^{+}24b, AAARZ24, AA21c, ASLG22, AOF24, APJ24, AGA^{+}22, AHS^{+}23, BARL24, Bad23, BFY22, BSM22, BHSAY24, BZP^{+}21, BCNS20, BCM^{+}23, BBA21, BOB^{+}20, BMP21, CDM23, CPRS20, Chu23, CCYL24, DKV24, DYLZ21, FHXL19, FRR^{+}24, FMA^{+}20, FCR23, GNV21, GS23b, GMA20, GRA^{+}23, GMMI16, GKW^{+}10, HCXH16, HHM^{+}23, HP16, HN18, HY22, HBMS20, IGL^{+}23, IAM24, ITZ20, JEJ^{+}24, JP21, KES22, KES24, KHP^{+}22, KM24, KVST19, KHN^{+}22, KDE24, KYS21, KPH^{+}24, KSG^{+}24b, LDd^{+}24, LMK18, LL20a, LZY22, LNT^{+}23, LNT^{+}24, LQY^{+}19, LRdLS19, LN19, LBK22, MNP^{+}24, MDS^{+}24, MCM^{+}10, MAVS20, MSN13, MMS22, MKSJ22, MFO^{+}20, MPL^{+}21, Mou24, MSP21, dSMdO22, MA23, MF20, MPCS^{+}23, MPM24, MTHN22, NB24, NLT^{+}23]. Detection [NSD^{+}22, ND24, OOM^{+}23, OAH^{+}24, PMP22a, PBP^{+}24, PRDK22, PZC21, PW21, PCM22, QQM24, RGB^{+}24, RM23, RLB^{+}23, RES23, RRT^{+}23, RPB^{+}17, RAA^{+}22, SBN^{+}23, SGG^{+}22, SSF^{+}18, SPC22, SFW23, SQK^{+}17, SA21, SC23, SWC^{+}23, SVFV23, SGDT19, SKHK23, SHLH24, TMP22, TLK^{+}20, TVAT24, TLM^{+}20, TTS24, TWW24, VGK21, VBB24a, VBB24b, VPCE20, VKKA17, VOP22, VMD^{+}22, VSEH^{+}21, WCF^{+}16, WLL^{+}19, WWLZ21, WBS^{+}24, WQH23, XGY^{+}22, XSX18, XY21, YLY^{+}22, YMH22, YZC^{+}20, ZMDCEMI22, ZHZ^{+}19, ZZ18, ZCHG19, ZLT22, ZMZ^{+}22, ZL22, ZZSX19, ZHH24, MGFF14]. Detections [SWGL19]. Detector [RDASB22, SLZJ23, Vem22]. Detectors [GSP15]. Determinants [Kha24, LFGR20, Zha19]. Determination [NE21]. Determining [VKKG22]. Deterministic [SDDB24]. Detonator [LT24b]. Develop [MRRMC21, NNRR^{+}21, PTK^{+}23, SOSR^{+}16]. Developed [AAG^{+}22]. Developing [AT14, CYO^{+}17, DWY^{+}19a, GMRRRQ^{+}16, JII17, MVAHBL^{+}23, ML20, PLG23, SGWK23]. Development [ASK23, ACA^{+}21, AAKS23, BMB23c, CGBF24, ESB^{+}20, FLRS12, FFP^{+}22, dSFdRP24, GS12, HPC12, LA23, LWY^{+}15, LXQ24, LG23, MBAS20, MS16, MGLBMMSC20, NZZ13, OLMBRCRC23, PLL11, RIS^{+}17, SSOÁ^{+}16, STG^{+}20, SS14, SLSK21, ST20, THT18, YBMK21]. Developments [GNGL23, NZ14, Rei23]. Device [ADM19a, AN23, ADAB21, AMY24, BAS24, CFP17, JWRX17, KTP17, KALY17, KTAH^{+}23, KBF23, LH21b, NSV17, SXXW20, SF22, TL24b, VMAG23, WD24, ZLS19]. Device-to-Device [ADM19a, BAS24, CFP17, JWRX17, KALY17, NSV17, SXXW20, SF22]. Devices [ABK22, ABCC22, AG22, AAG^{+}22, AGA^{+}22, BKSS19, BC23, CLW23, CLT^{+}22, CBD^{+}22, DSMM21, FRR^{+}24, GSL20a, HOV23, JEJ^{+}24, LXL^{+}20, MS22, RC18, RSX18, RPHJ11, WZL23, YZC^{+}20, dOB19]. DevOps [ASL22]. DF [SHC22]. Dfake [KYS21]. DFT [RGCM21]. DFT-Based [RGCM21]. DGA [YLY^{+}22]. Diabetes [AEM^{+}23, NB12, PTK^{+}23]. Diagnosis [DSM21, OMD^{+}22, PC23, VMAG23, WZL^{+}20]. Dialing [KE22]. Dialogue [KHC^{+}24]. Dialogues [WBK^{+}23]. Didactic [ASFAV23]. Diesel [Tra18]. Difference [FLLG24, TTS24]. Differences [GS11]. Different [ANHMS24, RF21, SLZJ23, SRS23, ZSS^{+}22]. Different-Sized [SRS23]. Differential [AMZ^{+}20, AVV^{+}23, ABBV23, DQ18, MS22, PC23]. Differential-Linear [MS22]. Differentiated [JP21]. Differentiation [Söd13]. Difficulty [BR21]. Diffusion [VLPR20]. Digit [NBL^{+}24]. Digital [AAK24, AAGB^{+}20, AARTSBAE24, AT14, AA24b, AGN21, BKR24, BD24, BCL11, CGRV20, CTP^{+}22, CT24a, CT24b, Cop14, CGBF24, CALHG^{+}24, CMS21, FGSD22, dSFdRP24, FR15, FMA^{+}23, Fra24, FLSL24, GDRCDBLÁ23, GLRM19, GWK17, GKFV24, GNGHEM21, GANT21, GS11, Gra13, dPCGMC20, GTWN24, HSM19, HOPGRV21, HOV23, HRGQHOA21, HSTK23, IP16, IL14, JWD^{+}24, KVWC24, KIG23, KFK^{+}24, LWW^{+}23, LNB12, Lee17a, LWBM22, LT24b, MEH^{+}22, MSIP20, MK20, MDPHB23, MCCOCE24, MKHR21, ND24, NPWC24, OLMBRCRC23, OGR^{+}20, Orl22, PH20, PC23, PLA^{+}21, PKCF21, QLR^{+}22, RNFS20, RABC21, RSG24, SACG23, SPSS17, SK24, Söd13, SPM^{+}22, TOLG21, TCG19, TPJ21, Tom21, Yan17, ZYWS24]. Digital-Twin-Aided [MEH^{+}22]. Digital-Twin-Based [FLSL24]. Digitizing [APB^{+}16]. Dilemmas [GFL21]. Diluted [GB20]. Dimensional [CXLB21, Goe12, KES24, SOK24, YLJ^{+}19]. Dimensionalities [McK20]. Dimensioning [ADAB21, KDJ24]. Dingo [CD23]. Direction [ASF^{+}23, CZZH15, DZH^{+}22, HE20]. Directional [CLW23, LLZ21, WX24, ZGZ^{+}18, ZZY19]. Directions [AZH22, AAMD21, AEA^{+}23, AYH23, AA23b, Bad23, BSEL20, BPG19, GNGL23, HNK^{+}21, JSJM23, KK23b, MZLP22, MHQ^{+}23, QLR^{+}22, QQX^{+}24, RGB^{+}24, Ray23, XZWL24]. Directory [BC19, HAA23]. Dirichlet [ZH21]. Dis [RFZ22]. Dis-Cover [RFZ22]. Disabilities [CWF13]. Disability [Söd13]. Disadvantage [Smi13]. Disaggregation [PPGC16, YWW^{+}19]. Disambiguation [Bea10, OY19]. Disassociation [ACAP20]. Disaster [CSA^{+}24, KYG22, PST^{+}21, SH21, TZE^{+}21, UUT^{+}19, UANU20, YT15].

Disasters [BOKC19, QSML24, SLSK21]. Disastrous [PFL^{+}12]. Disciplinary [MSR^{+}14]. Discourse [DL10, TPP^{+}23]. Discover [CGMM22, KKBD10]. Discovery [BP24, BGL^{+}22, HPB^{+}24, JMJ^{+}23, KS19, Por21, WZC19, YZC^{+}20]. Discriminant [DWZN16]. Discriminative [SQK^{+}17]. Discussion [Hel15]. Discussions [BBA21]. Disease [CTM22a, OHSN24, ZC21]. Diseases [Bad23]. Disinformation [PMA^{+}22]. Disorders [JTA^{+}21]. Displacement [WL20]. Display [ASAH^{+}23, BRC24, HSM19]. Disruptive [BJL^{+}22, KMV^{+}23]. DISSECT [MKK17a]. DISSECT-CF [MKK17a]. Dissemination [ASMM22, BRP^{+}13, CSL^{+}20, GMS^{+}20, HLZ23, LH21c, MV20, UDF^{+}19]. Distance [Alh22, ERBL12, FTHY21, GPM21]. Distillation [RRT^{+}23]. Distillation-Based [RRT^{+}23]. Distinguishing [RS22]. Distorted [JMZ^{+}22]. Distortion [Lee17a]. Distortions [MR23a]. Distracted [WWLZ21]. Distributed [AMEF21, ACA^{+}21, AVE^{+}22, BNJ24, BD24, BMS24, BB17, CALHG^{+}24, EPC24, FWLY23, HEP^{+}11, HDT22, JOL22, KGRP21, KK21, KK23a, LSF19, LWZL22, LZY^{+}24, LHL24a, MXT^{+}16, MMT24, MBF^{+}12, Oza23, RR24b, SA21, SW21b, SGDT19, STS23, SPS^{+}21, Tas10, Tre21, VdHH^{+}19, VDSK23, VKM^{+}19, WYS17, WSF23, WXL^{+}23, XNC21, XSM22, ZTBD20]. Distribution [AZH22, AMZ^{+}20, BTBK21, CGT21, DGF23, EW19, MT17a, MT17b, MCPLBH^{+}24, PL22b, YG20]. Distributions [GNK^{+}10, PPM21]. Disturbances [TGELGH13]. Dive [LWH24]. Diverse [Bi14, MLK22]. Diversity [VL20]. Divide [AT14, BCL11, SPM^{+}22]. Diving [RKY20]. Division [LWW^{+}24, NAV^{+}23]. Divisive [VLPR20]. DL [MS22, Mou24, TMS24]. DLT [AKK^{+}24]. DNN [MPCS^{+}23]. DNS [HN18, MMM21]. Do [Mar13]. DOA [LWHR14, WLHR13]. Document [KHP^{+}22, MKK^{+}22, PPR^{+}20]. Document-Level [MKK^{+}22]. Documents [ADSAKAD22, DMS^{+}22, AJT^{+}23]. Does [Mot24]. Dog [FSG22]. Domain [ALL24, BGL^{+}22, CQWC22, DCG12, GPCP12, KS19, MMS^{+}19, MCR^{+}23, PPKS21, QDL22, SAM^{+}18, YLY^{+}22, ZZZL23a, ZWZ17, ZZZL23b]. Domain-Specific [KS19]. Domains [HWCH12]. Domestic [MCPLBH^{+}24]. Dominance [GPK^{+}20]. Dominant [ARAAA19]. Dominating [ZHZ^{+}19]. Don’t [Tsi14]. Doocing [FC19]. DOORS [Tas10]. Doping [PRS22]. Double [AMEF21, BCG10, GFL21, RD18]. Double-Boomerang [BCG10]. Double-Layer [GFL21]. Double-Spending [AMEF21]. Down [ALH24, Kra20]. Downlink [MEO18a, SHN^{+}19, YZP22]. Downlinks [KLKP19]. Downloading [BC19]. Downward [WCM19]. Downward/Upward [WCM19]. DPDK [LSZ21]. DPDK-Based [LSZ21]. DPIA [GL21, HBFJ20]. DPoS [WLL20]. DQN [XLM^{+}23]. DQN-Based [XLM^{+}23]. DQSM [CSDAM^{+}23]. Draft [Mot24]. Dragon Pi [LNT^{+}24]. Dragonfly [MKR22]. Dramatic [BCT21]. Drift [SLZJ23]. Drifting [LGY23]. Drinking [ATE24]. Drinking-Water [ATE24]. Driven [AIP^{+}24, ACS24, AMR^{+}21, ACD^{+}22, AAA^{+}24b, AS15, CBR^{+}24, EKG24, EAAA^{+}23, GOM23, GTWN24, JBS^{+}23, JT23, KHM24, KGWR23, KPKM23, PBP^{+}24, PLL11, PKS^{+}24, RRT20, RDD^{+}14, RH24, Sof19, TAAK21, Ves18, WKF^{+}21, YPMM12, ZGH13]. Driver [ASMM22, LQY^{+}19, YZ23]. Driverless [ZYLM17]. Drivers [AAAAM^{+}21, You23]. Driving [BBBHA24, Gog12, HZCZ22, LLR19, WWLZ21, YZ23, YKY18]. DRL [BCSB24]. DRL-Based [BCSB24]. Drone [AJ24, BSM22, KL23, NSD^{+}22, QSML24, TAAK21, ZMDCEMI22]. Drone-Acquired [ZMDCEMI22]. Drones [RS17, dCdZCC24]. Dropout [dSPR^{+}22]. Drug [CYC^{+}21, JMJ^{+}23, ZC21]. DSP [DDZ18, GDARM18, WZG^{+}18]. DSP-Based [DDZ18, GDARM18, WZG^{+}18]. DT [ABLR23]. DTN [KST24, OT16, THT18, YT15]. Dual [CZZZ22, DYS15, JUH^{+}23, ON23, PBAAN19, ZHZ^{+}19, ZXMB22, YWW^{+}23]. Dual-Band [PBAAN19]. Dual-Channel [ON23, YWW^{+}23]. Dualism [Ant12]. Duality [YB24]. Dump [RM23]. Duplex [AZ19, LZL^{+}23, PP22b, SNH^{+}19, SXXW20]. Duplication [KCT^{+}24]. Duqu [BPBF12]. During [Del21, FDN^{+}24, TMS24, ZLOW24, CTM22a, GBDDVS23, GFPD21, MV20, MPG20, PS17, PLA^{+}21, RFSOHMSB21, YHN21]. Duty [SWGL19, SSX20]. Dynamic [AIP^{+}24, ADM^{+}19b, AYH22, AVE^{+}22, BCG10, BCFP21, CR23a, CZY17, DH18, DYLZ21, DNZ22, EAHO21, FSG^{+}16, GW15, Gro17, HVS17, IUK^{+}23, IMLL24, KP19, KIB^{+}22, KKT20, Lee17b, LLW21, LGCM24, MP16, MATP19, NRS24, OS23, OGB^{+}24, PDTM15, POC20, QLL^{+}21, RC18, RRMI22, RAA^{+}18, SAJ24, SV24, SDM21, SHBS19, SZFC18, TSHO24, TOLG21, TTN24, UFK^{+}10, VL18a, WC23, WYZ24, WZLL19, XCSM18, XZY^{+}24, XDS^{+}19, XQLZ19, YWS22, YBO^{+}23, YQ24, YLK^{+}19, ZL19b, ZXXB22, ZX24]. Dynamical [TLC15]. Dynamics [MGFF14, SAL21]. Dynamis [PDTM15].

e-Book [OF23]. E-Commerce [AZS23, CMP^{+}19, GTM22, Khr20, TLC15, WYX18, ZSRC23]. E-Customer [KA21]. E-Democracy [Rot12]. e-Governance [LSN21]. E-Government [JII17, Yan17, AZV16, DGKL15, GS12, VLMP21, ZKV12, ZAV14]. E-Health [Nil12, RHHN21, AMAJ12, SIZD12].

E-Learning [BOH23, ESB^{+}20, MFH22, PFT^{+}24, PLA^{+}21, JGV^{+}21, ZSS^{+}22, ESB^{+}20]. E-Mail [SG21a, SKM21]. E-Marketplace [MBAS20]. E-Training [ACBP19]. E-Vote [CKFH22]. Eagle [RRMI22, RR24a]. EAOA [GM20]. Early [Bad23, BCSB24, DKV24, GS23b, GMMI16, HP16]. Earth [BC19, Kra20]. Earthquake [PFL^{+}12]. Ease [BFFZ^{+}12, KMS^{+}12]. East [JAS21]. EASY [LSM^{+}12, DSMM21]. Eavesdropper [THV^{+}22]. Eavesdroppers [TH21]. Eavesdropping [HS20]. EBBA [FKZ22]. eBooks [SC20]. ECG [Ciu19]. ECHO [KP15]. ECIES [KAS^{+}22]. Ecological [ACK^{+}16]. EconLedger [XNC21]. Economic [ASFAAV24, KAL^{+}19, SVV^{+}22]. Economical [HC21]. Economy [KHM24, LSN21, LGMZ19]. Ecosystem [DGHH^{+}20, FSG^{+}16, Fri13, GPL24, Li18, RF24, SCC^{+}23]. Ecosystemic [Giu18b]. Ecosystems [FYWS16, GVV^{+}23, HPC12, KNN^{+}24]. EDA [LF24, Mou24]. EDAS [AS15]. Edge [AJA^{+}24, AAM^{+}24, AALS21, AAAAM^{+}21, AAG^{+}22, ALDS19, AGA^{+}22, Bad23, BFY22, BMJ^{+}22, BCM^{+}19b, BC23, CD23, CCCPV^{+}24, CPM^{+}19, CC22a, CALHG^{+}24, DSMM21, DVED20, DHEK19, DWL21, DMP^{+}23, GHZY23, GGT^{+}23, GBKN23, IG10, UMS24, JMJ^{+}23, JHBW24, Kab23, KVWC24, LKIL19, LPVM24, LJR23, LJR24, LW24, LW21, LZY^{+}24, LHW24, MRM^{+}24, MATP19, MVMHL24, MDRR24, MZLP22, PPKS21, PVM22, PNS23, PAK^{+}23, Poz24, QHX^{+}24, RAC24, BNK^{+}23, SSF^{+}18, SBRZ24, SC23, SGLZ20, SK23, SFSN^{+}24, SKHK23, SEB^{+}19, TTKZ19, TCH23, TRST24, TCYZ22, Vem22, VKM^{+}19, VLC^{+}19, WXL^{+}23, YPZ18, ZZWP19, ZOAR24, ZWL^{+}24, ZQL^{+}24]. Edge-Computing-Assisted [GHZY23]. Edge-Computing-Based [SC23].

Edge-Enhanced [SBRZ24]. Edge-Scale [QHX^{+}24]. Edited [MC12]. Editor [DeD16, MHS23]. Editorial [Bod22, Mac22, Tos23, ZMKR22]. Educate [SAF14]. Education [Alh22, BCSB24, BBK^{+}20, Del21, Del23, EWGK24, ERBL12, GPM21, KPB23, KYG22, LLB24, LL20c, Mar17, MRRMC21, PJMM22, PEN23, RMLAZOPC21, TPJ21, UW21, XZWL24, ZLW21, BPLBM21, LBBPM21, MRRMC21, NNRR^{+}21, VBWW22]. Education-Case [ERBL12]. Educational [BRF^{+}23, DRK^{+}22, GNGHEM21, LHD^{+}24, MGLBMMSC20, OMTFFL21, SRMJ21, TA15]. EEG [BCMPW21, GBDDVS23, PBP^{+}24, ZBK^{+}24]. Effect [DTTK19, HCW20, HC19, HJA18, LLW21, OCCB20, WZLW19, ZS18]. Effective [AK19, AKM^{+}23, ALH24, Chu23, HA23, KM24, KMK24, PDTM15, Rók23, SOE^{+}24, SESO24]. Effectively [ASK23]. Effectiveness [BKR24, ELC^{+}19, KYG22, KID^{+}15, PKMS23, VM24a]. Effects [BHH12, DPBG18, ESB^{+}20, GL19, LLZ21, LTW21, PDI21, PDMK19]. Efficacy [MTF^{+}21]. Efficiency [BARL24, BCZ^{+}23, BAS24, CXZ23, HRAA19, HKZ^{+}22, KZ24, KPC^{+}23, OGB^{+}24, PGFH24, SKI^{+}19, SNH^{+}19, TAMA22, TDW24, Wan17, YPG20]. Efficient [AK19, AKM^{+}23, ASBC22, AFB23, ADS21, AS24, BMS24, BZNB23, CBM^{+}20, CFP19, CGBF24, CAL23, ENT21, FWL22, ISLARP^{+}22, IZK^{+}23, JWRX17, KES22, KZ24, KCT^{+}24, KFP10, KSSF19, LC23, LL20a, LCW^{+}18, LXL^{+}19, LHSS12, MMA10, Mah17, MLL^{+}22, NST23, PZH^{+}18, RRP17, SGLZ20, SMR24, SOR16, Sul22, TSHO24, TAHO23, TH20, TCYZ22, VGK21, WYS17, XZ19, XLM^{+}23, YPXH19, YLH^{+}17, ZBN21, ZWL^{+}24]. Effort [Alb19]. Effort-Aware [Alb19]. Effortless [GOM23]. eHealth [HP19, IKP^{+}13]. eIDAS [RSG24]. EIGRP [SAR24]. Elastic [AD18, AFS^{+}23, FSY^{+}22, KLG^{+}24, VdHH^{+}19, YCS17]. ElasticSearch [BSE^{+}21]. Elderly [DVSBV^{+}24, FR16, HMR^{+}21, LA20, PS17, TVAT24]. eLearning [SC20, Smi15]. Electric [EAHO21, SVV^{+}22, YKY18, ZX24]. Electrical [AARTSBAE24]. Electricity [HSL^{+}23, TBT^{+}23]. Electrocardiogram [Ciu19]. Electroencephalography [HYLP19]. Electromyography [DAMAS^{+}19]. Electronic [CCTV24, CCCA22, GRG21, LDM22, LTTS21, LT24b, PPR^{+}20, PTK^{+}23, VMAG23]. Electronics [GPM21, KKL^{+}20]. Element [BBK^{+}20, WX20, XGN19]. Elements [PDI21]. Elevators [SC23]. Elgamal [ADM19a]. Eliminating [DCG12, Söd13]. Elixir [LF24]. Elixir-EDA-MQTT [LF24]. Elliptic [ADM19a, AKDL24, CISM22]. Elliptic-Elgamal-Based [ADM19a]. Elusive [VI16]. Email [AJT^{+}23]. Embedded [BKSS19, PPMMS19, SS17, ZZSX19]. Embedding [CASN21, FR15, LLLD21, MGM^{+}23, SPC22]. Embeddings [BOB^{+}20, CN24, LHD^{+}24, LTM^{+}22]. Emergence [Cal24]. Emergencies [Han13, Pic21]. Emergency [CTP^{+}22, DG13, DCA16, GW13, LGV13, MDRR24, QSML24, SMS^{+}24, SLB24]. Emerging [BOH23, DTL^{+}21, DWL21, GDRCDBLÁ23, IG10, KDG24, KIG23, KPC^{+}23, MR23b, PNdC^{+}23, WSF23]. Emilia [TC21]. Emissions [SSO^{+}19]. Emitting [HLG21]. Emojis [AF22]. Emote [DG18]. Emoticon [YKC19]. Emotion [AWBS23, CKSG21, GCA^{+}22, HYLP19, LY19, PL22a, RSG21, SGMMRG21, VCG24]. Emotional [AF22, CCM^{+}17, GRB^{+}21, YKC19]. Emotionality [BCT21]. Emotions [CDM23, FSG22, FDN^{+}24]. Empire [Ant12]. Empirical [AMAA^{+}21, AAC21, DTL^{+}21, EAKM22, HSM19, HUM^{+}21, ISG20, OBK23, RAK^{+}20, TMS24]. Empirically [HP20]. Employing [HHL24, JK24, MV20, NAFM^{+}24]. Empowered [CSA^{+}24, HWCL17]. Empowering [MEH^{+}22, YWZ^{+}24]. Empowerment [FBOC15]. EmuCD [CSL^{+}20]. Emulated [PC24]. Emulating [Mat20]. Emulator [CSL^{+}20]. Enable [KMS^{+}12, PPP^{+}23, QXC^{+}21, SU17c]. Enabled [Alb21, CRMT23, CVG^{+}22, DWL21, ENT21, FMMS21, GWF^{+}23, GMS^{+}20, GTWN24, Lee17a, Lee17b, LXG^{+}23, MMH16, MDPHB23, NA23, PM24, SJA^{+}23, SG12, SST23, SPT23, TZS^{+}21, TH20, TZE^{+}21, WUH^{+}23, WADL^{+}24, Yiu21b, ZZWP19, KSG24a]. Enabling [AN23, AEA^{+}23, AMC^{+}24, CGBF24, FM16, GDARM18, HG12, JNARV^{+}24, KGKP20, Mac22, MDRR24, NA23, PPKS21, PSM^{+}18, POC20]. Enactments [LSN21]. Enclaves [WYL^{+}21]. Enclosed [YWLY22]. Encrypted [GKG19, HWZP18, PRDK22, SOR16, Uto13, WLCC24]. Encryption [ABBV23, FQ21, KHRG19, NDS^{+}23, SHHM21, TDW24]. Encyclopedia [WJBZ22]. End [BdCRM10, CGBF24, DTR22, HS18a, Kab23, KKP22a, MS16, MSAD22, NPB^{+}19, WXL^{+}23]. End-To-End [NPB^{+}19, BdCRM10, DTR22, Kab23, KKP22a, MSAD22]. End-User [CGBF24, MS16]. Energy [AGC18, BAS24, CMPT24, CLW23, CXZ23, CFP19, CGBF24, CBD^{+}22, DXL^{+}22, EHJ17, ENT21, FKM24, FML^{+}22, Gal24, GL13, GM20, HRAA19, HZQ24, HLC24, HAL18, JWRX17, KCT^{+}24, KAL^{+}19, KPC^{+}23, KFP10, Lee17b, LZ23, LXL^{+}20, LYP^{+}19, LXL^{+}19, MMA10, Mah17, MPT23, MCA23, MSJ^{+}23, MLL^{+}22, NMT23, OO19, Oza23, PGFH24, PVM22, PPGC16, QSML24, RBG^{+}22, RPHJ11, SOK24, SXXH19, SNH^{+}19, Sul22, TAMA22, TH20, WLHR13, WYS17, Wan17, WZL^{+}20, YPXH19, YPG20, YG20, YLH^{+}17, YXL^{+}21, ZWZ19]. Energy-Aware [GM20, HZQ24]. Energy-Based [KAL^{+}19]. Energy-Consumption [HLC24]. Energy-Defined [Gal24]. Energy-Efficient [ENT21, JWRX17, KCT^{+}24, LXL^{+}19, MLL^{+}22, Sul22, YPXH19]. Energy-Neutral [CBD^{+}22]. Energy-Optimized [PVM22]. Energy-QoS [GL13]. Energy-Saving [LYP^{+}19]. ENF [XNC21]. Engagement [ACBP19, Gar12]. Engaging [BR21]. Engine [BMB^{+}23b, DSM21, KH10, MDM21, Tra18, VG20]. Engineering [AS19, CPRS20, CPP^{+}20, DL10, EMMP22, GSB^{+}23, GPM21, JJFC22, KKG20, Orl22, PB23, RRP17, SK19, SSX20, Tos23, UB18, VP22, ZLL18]. Engines [MVAHBL^{+}23]. English [AdMTJ20, Gao21, GMA20, GPK^{+}20, LL20b, hSgLH17]. EngraveChain [SW21b]. Enhance [AAAAK24, BDL22, HDBD23, Hua15, HYLP19, KMS^{+}12, PLSF14, SK24, VBWW22, ZN22]. Enhanced [ABH24, AVV^{+}23, AS24, CH23, DWZN16, FKZ22, FSI^{+}24, FPT20, GZ23, GS23a, JWD^{+}24, KPT^{+}24, KYG22, KZ24, KAS^{+}22, KTD24, KFK^{+}24, LSJ19, LGY23, MEO18a, MOAI17, MZH22, ON23, Pec18, QNJ24, RL24, Rec21, SBRZ24, SESO24, SE15, SS15, SZT23, TVAT24, TRC^{+}21, WX22, ZOAR24, ZX24]. Enhancement [AIPV21, CHMZ21, HHL24, WX24]. Enhancements [CCA^{+}20, KHC^{+}23, WC17]. Enhancing [AKK^{+}24, AAAIA24, ACM^{+}19, BARL24, BBGP19, BOH23, CSA^{+}24, DVSBV^{+}24, EARP^{+}24, BBBHA24, GKFV24, GZS^{+}24, IPP24, KKP22b, KDEA^{+}23, KHC^{+}24, KBF23, KSG^{+}24b, OHSN24, PCROB21, PGFH24, PNM^{+}20, SE23, Smi13, TAMA22, TDW24, TDM^{+}24, TT22, TA15, ZXJ22, dCdZCC24]. Enough [APJ24, GLD^{+}18, LMY21]. Enriched [LG23, PLA^{+}21]. Enriching [DBD^{+}14, RTN^{+}22]. Ensemble [ASCM22, Alb19, BKW^{+}24, FRR^{+}24, JJFC22, MAVS20, SOE^{+}24, TS22, WLL^{+}19]. Ensemble-Based [BKW^{+}24]. Ensembler [AAA^{+}24b]. Ensure [CNC^{+}21, PKMS23]. Entailment [WZL21]. Entering [BM23]. Enterprise [Pet11]. Enterprises [Kha24]. Entitie [Zha19]. Entity [ATH18, BKW^{+}24, GXWD19, LMPT19, MSL^{+}23, WX22, WJBZ22]. Entropy [KDE24, LYD22, MPCS^{+}23]. Enumerated [Tsi14]. Environment [AK18, AYH^{+}21, CKFH22, GMD^{+}22, Gra13, HZW^{+}20, HSC^{+}22, IUIL23, Jal19, JP13, LWY^{+}15, McK20, MSIP20, MSJ^{+}23, NIK^{+}21, NPB^{+}19, PPKS21, PG23, PAM^{+}24, RYAA22, RZQS18, RK19, SAJ24, SGG^{+}22, SPSS17, VFFHF19, WHB^{+}21, You23, Zar22, ZJ18, ZYLM17, ZQL^{+}24]. Environmental [Cro19, FZ15, GAGB19, GPPB^{+}21, GN16, KPB23, KKTS22, LZ11, PL22a, PPP^{+}23]. Environmental-Based [GAGB19]. Environmentally [XSX18]. Environments [ACG^{+}24, ACG22, ADAB21, AMC^{+}24, ALLR16, ARC^{+}19, AD18, AMY24, BMJ^{+}22, CD23, CPI^{+}18, CFGD20, DIE^{+}21, DGKL15, DPBG18, EARP^{+}24, FMCT09, FM20a, GSL22, GSD^{+}17, GGCY20, HXHP17, HOPGRV21, HS20, UMS24, NST23, NNRR^{+}21, OMTFFL21, PCROB21, PM24, PFT^{+}24, RGMFM12, SBRZ24, SKT^{+}19, VL24, VFS^{+}19, YOI19, ZCG^{+}24, ZDP^{+}22]. Envisioned [ASMM22]. Envisioning [SACG23]. EO [TZE^{+}21]. eParticipation [VTPT24]. Epidemic [ESB^{+}20, HD19]. Epidemics [LZC21]. Episodes [ML20]. Equalization [CXG^{+}22]. Equation [FLM23, PC23]. Equipment [KWI^{+}19, NZU20]. Equivalence [CB24]. Equivalent [LSSD22]. Era [BCM^{+}19a, DGD^{+}19, Yan17, YLX^{+}24, ZCdRR^{+}23]. ERGCN [WX22]. ERMOCTAVE [MSC19]. Error [WX20]. Errors [QLK^{+}24]. Escalation [HAA23]. Escape [WLCS17, FAT19]. Escaping [SUF^{+}22]. ESP32 [LL23a]. ESTA [SS15]. Establishing [FKF24]. Establishment [DNKDN^{+}24]. Estimates [ALST20]. Estimating [FPGO24, MN21]. Estimation [AMSM23, BHS24, COL22, DWY19b, GZS^{+}24, KKP22a, KKL^{+}20, KWI^{+}19, KK21, LWHR14, LLJ16, NRS24, QLK^{+}24, VPCE20, WLHR13, WLJM21]. Estimations [Cop14]. Ether [WKD22]. Ethereum [BHSAY24, PTM20, ZL22]. Ethernet [MM21]. Ethernet-Based [MM21]. Ethical [AGB22, VI16]. EU [GPK^{+}20, RSG24]. Europe [TCG19]. European [CR21, FMNS11, GPLK22, MCR^{+}23]. EV [EAHO21, WHXL18]. Evacuation [CTP^{+}22, DG13, FAT19, WS21]. Evacuees [Bi14]. Evaluate [GMRRRQ^{+}16, dCdG20]. Evaluating [BSE^{+}21, CEZC19, IAM24, KTUF19, LHD^{+}24, LIM23, OBK23, TAD23, TMS24, YST24]. Evaluation [AAY21, ABSG21, AALS21, AMC^{+}24, ACF24, AA24a, AS15, ABB23, BMB23c, BAK23, BTBK21, BS17, BCG10, CCCPV^{+}24, DDMA24, DHEK19, GPL24, GVY^{+}23, HCXH16, HAL^{+}22, KP19, KHN^{+}22, KUBS22, KVVD22, KAS^{+}20, LRD23, LLY23, LBB23, LXG^{+}23, MDS^{+}24, MV20, MLK22, MKDG20, MGBG21, MMMH19, MCCOCE24, dSMdO22, NNT22, OAG21, OSdS24, QAQA21, RR24a, RNST23, RRdBSS22, SSOÁ^{+}16, SS22, hSgLH17, SP22, TRB22, Tom21, VR21, WK24, WBK^{+}23, YAV^{+}21, ZCG^{+}24, ZLZZ23]. Event [AS15, CLTP22, DB23, MGM^{+}23, MFO^{+}20, PCROB21, RAA^{+}18, SGWK23, SLB24, SJ20, VdHH^{+}19, WZ18, ZL21]. Event-Driven [AS15]. Events [BCT21, GBKN23, PFL^{+}12, TADS20, YC22]. Ever [RBVV22]. Everything [ASM^{+}24, LSN21, MAEP18, WUH^{+}23, NSJGN21, SNKG20]. Evidence [HCHP16, ISG20, Lal14, PDI21]. Evolution [DYS15, Giu18b, GM23, HD16, Let16, MP20, MKvD11, NZZ12, PFdRGG21, STS23, TTK24, VSEH^{+}21, WXL22, YBO^{+}23, YCD^{+}23, ZZ19]. Evolutionary [SAL21, YCD^{+}23]. Evolving [LGY23, MUS11]. Ex [LAG^{+}18]. Examination [Roy14]. Examine [HT14]. Examining [GS11, GR22b, HCW20]. Example [GL19, TPJ21, ZLW21]. Examples [ALZ^{+}23, FLC21]. Excellence [GKK^{+}19]. Exchange [CGLR19, HWCH12, Jeo20, PAM^{+}24, SMR24, TRC^{+}21]. Excision [IKS^{+}21]. Exclusion [PE13, WT13]. Executing [KGWR23]. Execution [CMO^{+}21, IUK^{+}23, SL24b, SHLH24]. Executions [SFW23]. Exhibitors [VNJ18]. Existence [SMS^{+}24]. Existing [KDJ24]. Exoskeletons [TPP^{+}23]. Expanding [KHC^{+}23]. Expansion [ZZZL23b]. Expectation [LNA17, ZSB24]. Expectations [PGSL20]. Experience [AIPV21, EMHF19, FC19, FDN^{+}24, Gao21, GPL24, GBDDVS23, HE20, HCW20, KTS18, KPC19, LOL21, MGBG21, NKSF24, PBB20, PMT^{+}23, PDMK19, PKS^{+}24, PLA^{+}21, SS14, TKT^{+}18, ZLOW24]. Experiences [GHFL24, Øie12]. Experiential [Dav12]. Experiment [HZWJ21, TL19]. Experimental [AAY21, BRP^{+}13, CCCPV^{+}24, FGBMG19, MM21, OAG21, PHGZ20, QAQA21, RGB^{+}24, RS22, SDM22, TRB22]. Experimentation [ACG22, GSL22, MLK22, YAV^{+}21]. Experimenting [AVA^{+}22, LGMZ19]. Experiments [CBD^{+}22, MLK22, NIK^{+}21, SUF^{+}22]. Expert [KSG^{+}24b, Pet20, PKCF21, SSOÁ^{+}16, WBK^{+}23]. Expert-Annotated [WBK^{+}23]. Expertise [ZGH13]. Explainability [GVY^{+}23]. Explainable [BRC24, FLLC24, FCG22, MPK^{+}23, SBN^{+}23, SESO24, SK24]. exPlanation [GDN^{+}23]. Explanations [RTN^{+}22]. Exploitation [LL23a, LLCL23, SWP^{+}22]. Exploiting [CMO^{+}21, GKW^{+}10, HAA23, HDL^{+}14, KKBD10, NSV17, SDS^{+}22, SS23, VL18b]. Exploits [MNH^{+}24]. Exploration [BC23, VRA^{+}18]. Explorative [GDCM19]. Exploratory [Sac19, SRM20, ZN22]. Explore [RRLT20]. Exploring [AMR^{+}21, ASL22, ASG24, ABBV23, BKR24, BHH12, GPK^{+}20, HGF21, JOL22, KBBH21, KVL^{+}22, McK20, MDPHB23, PNS23, PM17, SKFM24, TOLG21, Tre21, ZN22, ZCG^{+}24]. Exposed [FKK20]. Exposing [TG13]. Expression [HAL^{+}22, ZXMB22]. Expressions [HYLP19, JPVC21]. Extended [BLW^{+}17, HZW^{+}20, KAEAB24, MVAHBL^{+}23]. Extending [Del23]. Extensible [MGM^{+}23, PG19]. Extension [MGM^{+}23, MKK17a, SSR17, LSM^{+}12]. Extensive [TEE10]. External [GPLK22, JGV^{+}21]. Extracted [NLD^{+}24]. Extracting [CDM23]. Extraction [DOB21, ITH^{+}21, LCC^{+}24, LPPG22, LTM^{+}22, QQM24, SKM21, TTS24, VMD^{+}22, WZ18]. Extraneous [HP20]. Extreme [KTD24]. Exuberance [PFdRGG21]. Eye [BKR24, NKSF24, PDI21]. Eye-Tracking [BKR24].

FaaS [RFS22]. Fabric

[LLHW20, LLW21, PORM^{+}23].

Fabric-Based [LLW21, PORM^{+}23]. Face [AYRA21, Gar12, GCA^{+}22, GRA21, LWZ^{+}21, RSX18, ZZ18]. Face-to-Face [Gar12]. Facebook [GM16]. FaceMashup [MS16]. Faceted [AMR^{+}21, Kra22]. Facial [GCA^{+}22, GR22b, HAL^{+}22, HYLP19, JPVC21, LQY^{+}19, RSG21, SWD^{+}18, ZXMB22]. Facilitate [KPC^{+}23, TLV23]. Facilitating [LSM^{+}12]. Facilitation [GGCY20]. Facility [BD24]. Factor [ASCM22, ADS20, ADS21, NKSF24, OPGH23]. Factorization [DDD^{+}22]. Factors [AAKS23, DTL^{+}21, GN16, MDPHB23, NPWC24, PL22a, RHHN21, SS22, VLMP21, YB22, Yan19, ZVKK19]. Factual [LS23, SPGS23, SSS21]. Faculty [ASFAV23]. FaDe [CGLR19]. Faded [MR23a]. Fading [TH21]. Failover [SAR24]. Failure [AHMI22, GMMI16, OCCB20, PDT23, PG19]. Failures [SW21a]. Fair [CGLR19, HHLZ23, HWCL17, KSSF19, MSK^{+}18, Pri23, QXC^{+}21, SVGGO24]. Fairness [CLT^{+}22, CFGD20, DPBG18, LWY20, LWZL22, SPK24, SIJA10, YZP22]. Fairness-Aware [CLT^{+}22]. Fairs [VNJ18]. Fairway [Ngu19]. Fake [AdMTJ20, ALL24, AAAR24, AAARZ24, AGN21, DGMP21, HP21, LZC21, PCVH24, SSS21, VMGT22]. FakeNewsLab [RS22]. Fall [MDS^{+}24, TVAT24]. Fallacy [DGGS19]. False [APJ24, RS22]. Family [ZLOW24]. Fan [PC23]. FANET [SHA23]. Farm [CNC^{+}21, LGY23, LSM^{+}12, PKMS23]. Farmers [VLMP21]. Farming [GBKN23, KVWC24, VLBRMP21]. Farms [GMRRRQ^{+}16]. Fashion [IMLL24, RGD24]. Fast [AMZ^{+}20, EMHF19, NSV17, TDT^{+}21, WC23, ZZ18]. Fatigue [LQY^{+}19, RSG21, vdSFRF23]. Fault [BP24, BRC24, CBR^{+}24, CAS^{+}20, DSM21, RK19, WZL^{+}20]. Fault-Tolerance [RK19]. Favored [GB20]. FC [GSL20a]. FDM [DNPC^{+}22]. FDMA [YXL^{+}21]. Feasibility [MKY24, PS17, SRSDF23]. Feature [AWUF19, BG24, BAR^{+}21, CPRS20, CR23b, FRR^{+}24, FRKS22, FKZ22, GDN^{+}23, HLC24, ITH^{+}21, JJFC22, KES22, LCC^{+}24, LQY^{+}19, NLD^{+}24, ON23, QQM24, RGM^{+}21, SQK^{+}17, STK21, TCB^{+}19, TTS24, VMD^{+}22, WXL22, ZWWW17, ZZ18, ZLXY19, ZCHG19, ZZYC21, ZLZZ23, ZXMB22]. Feature-Based [TCB^{+}19, ZWWW17]. Feature-Selection [HLC24]. Features [AdMTJ20, AFAAM19, CXLB21, DMS^{+}22, LMF23, LPM10, MKKS18, MSN13, MXL^{+}21, MSAA22, NLD^{+}24, PNMK22, PMP^{+}22b, RRT20, VDKL21, WZ18, XY21, XSM22, YWS22, YA24, YPZ18]. Featuring [XSM22]. Fed [Giu18b]. FedCO [ASBC22]. Federated [ASBC22, AM23, AG22, AYH22, AA23b, AVV^{+}23, AAJ^{+}22, ABBV23, BARL24, CD23, DDPVP18, DCL24, EKKS24, FQ21, FK24, HHLZ23, LCC^{+}24, LLLD21, LW24, LT24a, MSC^{+}24, NBL^{+}24, NST23, OHSN24, PKF24, RNST23, URH^{+}23, WRHH23, XLM^{+}23, XSSA24, YSAY23, ZPX24, ZXWZ21]. Federations [TXJ23]. Fee [WC23]. Feed [ZY20a]. Feedback [GCEOBRT23, LRL^{+}24, SAF14, TA15, WYX18, YK21]. Feeder [SWZ18, SWY^{+}18]. Feedforward [YMH22]. Felix [PG19]. Femtocells [UGJA18, KOJP13]. Few [LDZ^{+}23]. Few-Shot [LDZ^{+}23]. FHIR [SSH20]. Fi [BCM^{+}19a, Lee17a, Lee17b, AHH^{+}23, CBKL22, LOL23, MN21, MTM22]. Fiber [CCD^{+}21, LBA^{+}24, MG10, GLW^{+}19]. Fiber-Based [LBA^{+}24]. fiber-induced [GLW^{+}19]. Fiber-Wireless [MG10]. Field [FRE^{+}22, LRD23, PD23, XJ10]. Fields [PKMS23, WKF^{+}21]. FIFO [HNHH23]. Fifth [WK24]. Fifth-Generation [WK24]. File [BMS24, LB24]. Files [Uto13]. Filesystems [CC22b, CC23]. Filling [MCPLBH^{+}24]. Filter [LZL^{+}23, QLK^{+}24, WWYQ19]. Filter-Based [LZL^{+}23]. Filtering [DH18, KDEA^{+}23, ON23, YLL21]. Filtering-Based [KDEA^{+}23]. Filters [AD20, FSY^{+}22]. Financial [CLC^{+}22, DCD24, LLY23, LLY24, PS16, TZS^{+}21, ZQL^{+}24]. Finding [HP16]. Findings [KP15]. Fine [BCMPW21, TPKA20, WLJM21, WCZ23, WLLY24, WJBZ22]. Fine-Grained [TPKA20, WLLY24]. Fine-Scale [WLJM21]. Fine-Tuning [WCZ23]. Fingerprint [JK24, LLYL22].

Fingerprinting [FNN21, KTAH^{+}23, dSMdO22, QLK^{+}24, SOA^{+}20, Str22]. Fingerprinting-Based [KTAH^{+}23]. Finite [DeD16, PG23]. Finite-State [DeD16]. Finnish [JSPH21]. Fire [AHS^{+}23, MSP21, NB24, WS21, XSX18]. Firefighter [SGR^{+}13]. Firefighting [BCM^{+}23]. Firewall [AV23, MMM21]. Firewalls [AA21b]. Firework [MLM14]. Firms [DWY^{+}19a, PLA17]. First [AML22, GSL20a, LSZ21, LF16, VLC^{+}19]. First-Class [GSL20a]. Fisher [DWZN16]. FIVADMI [MNP^{+}24]. FiWi [MG10]. Fixed [OT16, SRSDF23]. Fixed-Route [OT16]. Fixing [FKK20]. FL [AN23]. FL-LoRaMAC [AN23]. Flagship [FRE^{+}22]. FLAME [MM23, BPBF12]. FLAME-VQA [MM23]. Flash [BZMB23]. Flat [EW19, LHW18]. Fleet [RS17]. Flexibility [MCA23]. Flexible [CALHG^{+}24, Gro17, KAL^{+}19, NMH^{+}21, SMVP21, Sor12]. Flight [FKZ22]. Flipped [JPVC21]. Floating [PDMK19]. Flock [LLSS24]. FlockAI [TAAK21]. Flooding [SLSK21]. Floor [JK24]. Florence [PBB20]. Flow [FML^{+}22, IZK^{+}20, MB22, MPT23, SU17b, TCH18, TDT^{+}21, WHB^{+}21, WYZ24, ZJJS^{+}23]. Flow-Based [MB22]. Flows [KDE24]. Fluent [LS23]. Fly [PMGG21]. Flying [CPKK23, KSJZ23, SJA^{+}23]. FM [CDLY24]. FM-Based [CDLY24]. Focus [Gar12, GPRMGP21, GSO^{+}10, KHM24, SRM20, ZQCC16]. Focused [DBD^{+}14, SFSN^{+}24]. Fog [ADM^{+}19b, ASMM22, ARC^{+}19, BMS20, BRF^{+}23, BP24, IFF21, Kab23, KVWC24, MDT20, MBH22, MC24, NBT22, Pec18, PdRB19, Poz24, RRdBSS22, SV24, SNKG20, SK23, Sul22, SEB^{+}19, VdHH^{+}19, ZOAR24, AVE^{+}22]. Fog-Based [IFF21]. Fog-Computing-Assisted [NBT22]. Fog-IoCV [ASMM22]. Fold [JDH21]. Followers [ZAV14]. Following [HZW^{+}20, HSC^{+}22]. FollowMe [YPLZ19]. Food [BYS^{+}15, CTM21, DCA23, PPMMS19, BYS^{+}15]. Foot [SSR^{+}21]. Football [CR21, MPK^{+}23, TCG19]. Force [SHA23]. Force-Based [SHA23]. Forcing [GVC13]. Forecast [HLZ^{+}18, LGY23, MSIP20, WX20]. Forecasting [ASCM22, ABH24, CMP^{+}19, DZH^{+}22, FFP^{+}22, HSL^{+}23, JRP21, KKBG23, KPKM23, RSMC21, TBT^{+}23, TOLG21, YZCR24, dSPR^{+}22]. Forecasts [GKFV24]. Forensic [AJ24, PV20, PV22, SKI^{+}19, SK23]. Forensics [AAK24, FLC21, HCHP16, IL14]. Forest [ASF^{+}23, AHS^{+}23, OAG21, OAI^{+}22, XSX18]. Forex [DMO^{+}19]. Forged [LDd^{+}24]. Forgery [LMLW18]. Form [AK22, SOE^{+}24]. Formal [DGF23, MBH22]. Format [KMS^{+}12]. Formation [BdBCCG21, KL23]. Formative [NE21]. Formats [BAK23]. Forming [LLW^{+}24]. Forms [Cal24]. Forward [GVC13, KFP10, LCW^{+}18, SJG18, XGN19]. Forward-Looking [XGN19]. Forwarding [AS24, ISLARP^{+}22, KTUF19]. Foster [ACK^{+}16]. Fostering [CL19]. Foundations [FHS^{+}10]. Four [AIPV21, LF16, TL24b]. Four-Pillars [AIPV21]. FPGA [MLPH23]. Fractal [KKK^{+}20]. Fractional [CL16]. Fracture [GAA21]. Fragmentation [SIJA10]. Fragments [RTN^{+}22]. Frame [Lee18, LMLW18, LMH19, MM23]. Frameless [YLYL22]. Framework [AJA^{+}24, AKK^{+}24, Adi23, AMG22, AN23, ACS24, ABCC22, AlQ23, ACBP19, AVE^{+}22, AAJ^{+}22, AHP23, BMB^{+}23b, BRBC14, BG09, BKSS19, BLW^{+}17, BM18, BDL22, CGRV20, CLT^{+}22, CLTP22, CCM^{+}17, DQ18, DGCK24, EH21, FAB^{+}22, FKM24, FHL^{+}24, FSI^{+}24, Fio10, GSL20b, GMS^{+}20, HHLZ23, HBM^{+}21, IFF21, IIOW16, JLH^{+}13, JII17, KAEAB24, KVST19, KP19, KS14, KPH^{+}24, KGWR23, KVVD22, KKT20, LDM22, LF24, LNT^{+}23, LZL^{+}17, LNG19, LLB24, LBK22, MSC19, MNP^{+}24, MKDG20, MSJ^{+}23, NMH^{+}21, NLT^{+}23, NSD^{+}22, NE21, NRS24, OR24, PC24, PPN18, PG19, PTZM22, PW22, QDL22, RRLT20, RMLAZOPC21, RNST23, RAK^{+}20, RD18, BNK^{+}23, SBN^{+}23, SBRZ24, SRS23, SMN22, Sor12, Sul22, TZE^{+}21, VdHH^{+}19, VCRCSV^{+}21, VCB14, YQ24, ZPX24]. Frameworks [AP22, DWY^{+}19a, GSB^{+}23]. Franchise [SOB^{+}19]. Fraud [BARL24, LRdLS19, MT17a, MTHN22]. Frauds [VKKA17]. FREDY [AVV^{+}23]. Free [BKW^{+}24, CCKH21, GPMO20]. Free-Space [GPMO20]. Freeway [SYGY21]. French [FTAA21]. Frequencies [ACA^{+}23]. Frequency [BCMPW21, DZH^{+}22, Lee17a, OFK19, VPCE20, ZZL^{+}23, dCdZCC24]. Friendly [RRP17]. Friends [MSX^{+}21]. Friendship [SJ21]. Fronthaul [MGV17, MM21]. Frontier [ASCM22, AA24b]. Frontiers [Del23]. Frontline [SGR^{+}13]. Frugal [EPC24]. Fruit [AGA^{+}22]. Fruition [AIPV21]. FttC [MGV17]. FttC-Based [MGV17]. Full [AZ19, LZL^{+}23, SNH^{+}19, SXXW20]. Full-Duplex [LZL^{+}23, SNH^{+}19, SXXW20]. Fully [BHS24, Del21, HPF^{+}24, JKM^{+}23, Pri10]. Fun [CGGK22]. Function [AP22, Aln22, CASN21, ELC^{+}19, ELCS20, FR24, FHZ^{+}24, GS23a, GVV^{+}23, GGK18, GKG19, JK20, KAS^{+}22, MSC^{+}21, Pap20, STCP21, SN23a, SZW^{+}23, VL20, WYL^{+}21]. Functional [HZCZ22]. Functionality [GPL24]. Functions [MP16, MR23a, SZW^{+}23]. Fundamentals [PKF24]. Fungible [GM23]. Furniture [FRE^{+}22]. Fused [AWUF19]. Fusepool [KS14]. Fusion [ABH24, CZL21, CSZ^{+}24, FSY^{+}22, KES24, KTAH^{+}23, LQY^{+}19, RGM^{+}21, SYLL21, WX20, WXL22, XY21, YWW^{+}19, YWW^{+}23, ZZ18, ZLXY19, ZCHG19, ZC21, ZLZZ23, ZQCC16, ZXMB22]. Fusion-Based [RGM^{+}21, WXL22]. Future [AZH22, AAMD21, AYH23, AA23b, AAE^{+}22, Bad23, BSEL20, BLU^{+}15, BPG19, CEP24, CFV19, DPH17, DPH19, DDPVP18, Del23, Dua21, EWGK24, FC19, FBDR23, FDN^{+}24, FGBMG19, Fra18, Fri13, GAGB19, Gal24, GPCP12, GW13, GNGL23, HE20, HNK^{+}21, HS09, IBBA20, IKP^{+}13, JMJ^{+}23, JSJM23, KGKP20, KHN^{+}11, KST24, LJR24, LZL^{+}17, LCMV17, LL20c, LFGR20, MG10, MMF^{+}19, MGJ22, MSIP20, MKvD11, MAEP18, MR23b, MAP23, MZLP22, MHQ^{+}23, NAV^{+}23, NZ14, NSD^{+}22, PKF24, Pri10, PSGM22, QLR^{+}22, QQX^{+}24, Rah22, RGB^{+}24, Ray23, RKM^{+}22, SVV^{+}22, SDDB24, SG19, SUF^{+}22, STG^{+}20, SMN22, VBB24b, XZWL24, YBMK21, Fut20, Fut21, Giu18a, Off15, Off16, Off17, Off18, Off19, Off22]. Futures [DZH^{+}22]. Fuzzy [DDPVP18, DCHM16, HWZP18, JP21, KKEV17, LA23, MM23, PCC^{+}18, RFD23, Tra18, YPMM12]. Fuzzy-Based [PCC^{+}18]. Fuzzy-NN [KKEV17]. Fuzzy-Rule-based [YPMM12]. FWA [GRM23].

G [FG17]. G-Networks [FG17]. G.723.1 [WLYL20]. Gaia [GHFL24]. Gaia-X-Med [GHFL24]. Game [AZ19, FC22, GFL21, HLZ23, KHN^{+}22, LSF19, LJR23, LJR24, MEO18a, MHSH16, PW22, RDASB22, SAJ24, TKT^{+}18, VCPT22]. Game-Based [KHN^{+}22, LJR23, LJR24, SAJ24]. Game-Theoretic [FC22, LSF19, PW22]. Games [AKT24, BB20, GBDDVS23, MVAHBL^{+}23, PNMK22]. Gamification [MKKG19, PJMM22]. Gamified [MVAHBL^{+}23]. Gamifying [KYG22]. Gaming [AWSA23, BHH12, NKSF24]. GAN [CZZZ22, dMLMS22]. Gap [GAMMPCRÁ24, SPM^{+}22]. Gap-Based [GAMMPCRÁ24]. Gaps [RES23]. Garden [PPDC22]. Gas [OOM^{+}23]. Gated [OAH^{+}24, SJ20]. Gateway [GHZY23, LHW24]. Gauss [BPBF12]. Gaussian [TGELGH13, YLL21]. Gazetteers [Lau15]. GB [WZG^{+}18]. GB/s [WZG^{+}18]. GDPR [GdSdC23, PYC21, Ves18, vSR24]. GDPR-Compliance [PYC21]. GDPR-Compliant [GdSdC23, vSR24]. Gender [AYRA21, AT14, dPCGMC20, GSD^{+}17, GR22b, Mar13, SRM20, VYN19]. General [AA23b, LNG19]. Generalization [RH24]. Generalized [HZW^{+}20, JKM^{+}23, SIJA10, WHB^{+}21]. Generate [DDD^{+}22, TMP22]. Generated [FR16, ND24]. Generating [BJL^{+}22, GGD23, GMLS22, JLH^{+}13, LTTS21, SZT23, UJ22, vSR24]. Generation [AlQ23, AEA^{+}23, ASM^{+}24, APP22, AMI23, CC24, FWVF19, HYMT24, HEP^{+}11, HPF^{+}24, HLLT10, HUM^{+}21, IG10, KP19, LGY23, MKSV^{+}19, MSJ^{+}23, RYY10, SAZ18, SYLL21, TZE^{+}21, WK24, WML24, WZG^{+}18, YKY18, ZZZL23b]. Generation-Based [ZZZL23b]. Generational [KHC^{+}23]. Generations [IHMG22]. Generative [BAK23, CZZZ22, GS23a, JGY^{+}24, KHC^{+}24, Lai24, OAH^{+}24, VR21, ZSB24]. Generator [AOF24]. Generator-Powered [AOF24]. Generators [CHH^{+}16]. Generic [MWL20]. Genetic [JHL^{+}18, MTKS24, SWY^{+}18, TTKZ19]. Genres [Hor19]. Geo [AKYP23, KLKP19, PLSF14]. Geo-Coded [AKYP23]. Geo-Wiki [PLSF14]. Geocache [GN16]. Geocentric [Mat20]. Geodata [Goe12]. Geographic [CR11, FZ15, GLZL24, Lau15, MZH22, NZZ13, NZ14, PA13]. Geographical [PLSF14]. Geography [PLSF14]. Geolocalization [MCCHO19]. Geomatics [HDK^{+}12]. Geometric [TCH18]. Geospatial [BGD09, KKK^{+}20, LW11, PFL^{+}12, RMB^{+}12, WSN24]. Geostationary [Mat20]. GeoSurf [MR23a].

Geovisual [SJ12, SLY^{+}12]. German [BKW^{+}24]. Germany [NZZ12]. Gesture [DMP^{+}23, RSX18, ZL19b, ZZZ^{+}22]. Getting [MRRMC21]. GIS [CDM23]. GIS-Based [CDM23]. GitHub [WKPC18]. GLI-Split [MHK13]. Global [CGnCD20, HY22, KH10, Kos16, KFK^{+}24, LZY22, LZLW22, LPM10, Mal20, MHK13, PH20]. Globally [And21]. Globes [SPC^{+}11]. Glove [DMP^{+}23]. GNSS [MKPG22, WL22]. GNSS-Based [WL22]. Go [PdRB19]. Goal [SKM21]. Gold [CTM19]. Good [LMY21, Pir22]. Google [ASAH^{+}23, CMP^{+}19, RCGSL19, RCL21, ZVKK19]. Gossip [KK21, LBB23]. Governance [AHP23, Cal24, DGHH^{+}20, DCL24, KPP^{+}20b, KHC^{+}23, LSN21, LMY21, Roy14, Tag20]. Government [AZV16, DGKL15, GS12, HPC12, JII17, PR20, QA24, Sca14, VLMP21, Yan17, ZKV12, ZAV14]. GPS [RRT^{+}23]. GPSR [GLZL24]. GPT [SB23]. GPU [RPB^{+}17]. GPU-Powered [RPB^{+}17]. GPUs [XZ19, ZZ21]. Gradient [SDL^{+}19]. Gradient-Weighted [SDL^{+}19]. Grained [TPKA20, WJ21, WLLY24]. Grammatical [STS23, TTK24]. Graph [AFS^{+}23, BMB^{+}23b, CYC^{+}21, CQWC22, CH23, CSA^{+}24, DG13, DDMA24, EvdMB20, FAS^{+}23, FBDR23, HCHP16, JTA^{+}21, KBP22, KVL^{+}22, OKH13, OFK19, SLSK21, SYZ19, SLZY21, VMGT22, VLK24, WX22, WLG^{+}22, WQH23, XZY^{+}24, YLPW21, ZC21, ZL22, ZH21, ZZ21]. Graph-Based [CSA^{+}24, KBP22]. Graph-Enhanced [CH23]. Graph-Labeling [OFK19]. Graph-Matching [ZL22]. GRAPH4 [GRA^{+}23]. Graphical [KHP^{+}22, LSSD22]. Graphol [LSSD22]. Graphs [FAB^{+}22, FM21, GGD23, GRA^{+}23, KK23a, LHD^{+}24, TAHO23, VRA^{+}18, WX22, YCD^{+}23]. GRAPHYP [AFS^{+}23, FAS^{+}23]. Gray [HLZ^{+}18, ZHZ^{+}19]. Greece [SS14, ZKV12, ZAV14]. Greedy [CD23, GLZL24]. Greek [KDKG22, ZVV12, Zaf12]. Green [MDB22, MPT23, NSJGN21, PG23]. Greenhouse [VBB^{+}24c]. GreenLab [VBB^{+}24c]. Grid [ANHMS24, BBSR24, ENT21, GM20, HGL^{+}23, JBS^{+}23, LLW^{+}17, Mal20, WHW22, MIK^{+}23]. Grid-Based [GM20, JBS^{+}23]. Gridded [CH16]. Gridless [AAR^{+}19]. Grids [CMPT24, CBR^{+}24, GPM21]. Grinding [PGSL20]. Groceries [PKMS23]. Ground [FHZ^{+}24, LA23, LLL21]. Groundwater [AZAV22]. Group [Alb21, AF21, ASL17, BBK^{+}20, DGGS19, FM20a, FSXP22, GPRMGP21, GFL21, SST23, THT18, WHW22, WLCC24, XDS^{+}19, YJSL18]. Grouping [KAEAB24, ZLLW18, ZZY19]. Grouping-Proof [ZLLW18]. Groups [ISG20]. Growth [BdBC23, MAEP18]. GRU [LWH24, RAA^{+}22]. GRU-Based [LWH24]. GSM [SS23, ZLT22]. Guarantee [Pin10]. Guarantees [AFB23, BDL22]. Guard [KSJZ23]. Guide [Fra23, GRG21]. Guided [CH23]. Guidelines [ACA^{+}21, HF14, KP15, MNWK21, Sof19]. Guiding [PS17]. Guizhou [ZLW21].

H [FC19]. H2O [FNN21]. Habit [LMSC19]. Hadoop [VCRCSV^{+}21]. Half [DDD^{+}22, PP22b]. Half-Duplex [PP22b]. Hand [BCMPW21]. Handheld [RPHJ11]. Handle [Tre22]. Handoff [CZK^{+}17]. Handover [DYS15, MEO18a, MEO18b]. Handshake [YT15]. Handwriting [BAR^{+}21]. Handwritten [RGM^{+}21]. HAPS [YBO^{+}23]. HAPT [NNRR^{+}21]. Haptic [DMP^{+}23, NNRR^{+}21]. Hard [YPG20]. Hard-Core [YPG20]. Hardware [AKDL24, CPM^{+}19, CBM^{+}20, DDV22, GMD^{+}22, MPL^{+}21, MSP21, NLT^{+}23]. Hardware-Based [NLT^{+}23]. Harnessing [ADK^{+}24, BRF^{+}23, GLW^{+}19]. Harvesting [CLW23, LZ23, SOK24, SXXH19, YXL^{+}21]. Hash [KAS^{+}22]. Hashing [Alb21].

Hashing-Based [Alb21]. Hashtag [AHD21]. Hate [VVK^{+}21]. haul [GLW^{+}19]. Haven [CTM22a]. Hawk [NB24]. Hazard [NJH^{+}24]. Hazards [Cro19]. HBase [MXT^{+}16]. HDMM [Tas10]. Head [BRC24, HY22, PLA^{+}24, YWS22]. Head-Mounted [BRC24]. Heads [SPGS23]. Healing [ADSAKAD22, JSAO^{+}23]. Health [AMAJ12, AEM^{+}23, AH21, BMB23c, BZP^{+}21, BGW16, BN14, CCTV24, CCCA22, GL21, GVY^{+}23, KS24, KP15, LTTS21, Nil12, PMT^{+}23, PORM^{+}23, Pic21, RHHN21, SIZD12, SCM12, SMK^{+}22, TBT^{+}23, TDM^{+}24]. Healthcare [AFS^{+}22, AAJ^{+}22, AMI23, AA24b, BHM^{+}19, BGW16, CT24b, GHA^{+}21, HBMS20, KHAA21, KCT^{+}24, LQFI17, RAC24, RAR^{+}23, RF24, SLP^{+}22, Tri14, ZSB24]. Healthcare-Data [BHM^{+}19]. Healthchain [WL21]. HealthFetch [SMK^{+}22]. Healthiness [UMS24]. Hearing [ZZZ^{+}22]. Hearing-Impaired [ZZZ^{+}22]. Heart [ABLR23, GMMI16, KDD^{+}21, OHSN24]. Heat [VBB24a, VBB24b]. Heaven [LCS12]. Heavily [MC12]. Heavy [SVFV23, VPCE20]. Heavy-Tailed [SVFV23]. HeFUN [NDS^{+}23]. Help [FAS^{+}23, FOJE19]. Henhouse [LWY^{+}15]. Heritage [AIPV21, FM21]. Heterogeneous [AFB23, AA20, CLT^{+}22, CXZ23, CDLY24, DDPVP18, Kab23, LLLD21, LW24, LLSS24, MZS19, MEO18b, NLT^{+}23, SFD20, WJBZ22, XQLZ19, YPXH19, YPG20]. HetNet [ENT21]. Heuristics [AA24a, JK20]. HEVC [EMHF19, MTA^{+}20]. HH [NLT^{+}23]. HH-NIDS [NLT^{+}23]. Hidden [BBA21, RKY20, WY19]. Hiding [CCKH21, Fio10, GM19, LCCC24, SHHM21, WLCC24]. Hierarchical [AES21, AAJ^{+}22, CGnCD20, CSA^{+}24, Chu23, DGKL15, KAEAB24, LL23b, LHW18, LH21c, SJ20]. Hierarchy [TWM20]. High [CC24, CR11, DNPC^{+}22, DZH^{+}22, IBBA20, JKM^{+}23, JP20, JP21, KKS^{+}19, KKM^{+}22, KLTC17, LTY18, LFGR20, MM23, MTF^{+}21, SOS^{+}23, SHB23, SZC19, TSHO24, VL20, XSM22, YCS17, ZZ21]. High-Altitude [TSHO24]. High-Band [SOS^{+}23]. High-Dense [LTY18]. High-Frequency [DZH^{+}22]. High-Level [SZC19]. High-Performance [IBBA20, XSM22, YCS17]. High-Rate [JKM^{+}23]. High-Speed [KKS^{+}19]. High-Voltage [DNPC^{+}22, KKM^{+}22]. Higher [Del21, EWGK24, GDS23, PJMM22, SIZD12, UW21, VBWW22, VDSK23]. Higher-Order [GDS23]. Highly [CYO^{+}17, YYY^{+}23]. Highways [NJH^{+}24]. HiiMap [HEP^{+}11]. Hijacking [TRC^{+}21]. Hiking [MMQ^{+}24]. Hindi [RGM^{+}21]. Hip [GAA21]. Historical [GPMO20, HSC^{+}22, PPDC22, PPM21]. Hitters [VPCE20]. HMI [MAP23]. HMM [WKW18]. HMM-R [WKW18]. Hoaxes [RFSOHMSB21]. Hoc [AASAI22, CGT21, DLK10, FR16, IT24, KH19, MZH22, SS15, TT22, YT18, BRP^{+}13, IZK^{+}23]. Hole [CXZ23, ZHZ^{+}19]. Holistic [GGD23, GHGL22, HKH^{+}24, KRSK24]. Holonic [SU17a]. Holt [KKBG23]. Home [ABCC22, AMY24, COL21, GRJ^{+}19, HSM19, HXHP17, KHN^{+}11, LYP^{+}19, LK20, MDS^{+}24, Oza23, PPN18, ZLOW24]. Home-Oriented [ZLOW24]. Homepages [PFdRGG21]. Homes [CGBF24, JEJ^{+}24, KCK^{+}19, LNB12, PPGC16, RGB^{+}24]. Homogeneous [PHG^{+}20]. Homomorphic [ABBV23, FQ21, KHRG19, NDS^{+}23, TDW24]. Honeypot [YYY^{+}23]. Honeypot-Based [YYY^{+}23]. Honeypots [FC22, LSF19, MEJA23, SWP^{+}22]. Hop [CGT21, LXL^{+}19, MZH22]. Hopping [DL19, ZLZ^{+}24, dCdZCC24]. Horizon [Tre22]. Horizons [VCPT22]. Horizontal [AKDL24, LBG^{+}14]. Horses [CKSG21]. Hospital [TBT^{+}23, TRHU23]. Hospitals [NZU20, SRS23]. Host [HCXH16]. Hostile [IGL^{+}23]. Hot [CDM23, HP16, WZC19]. Hot-IP [HP16]. HOTS [VBWW22]. House [SGWK23]. HPC [CC24]. HRI [KG18]. HTTP [IFF21, KMS^{+}12]. Hub [LH19]. Huber [HHL24]. Human [CB16, CYO^{+}17, COL22, FHL^{+}24, DMP^{+}23, GR22a, HDK^{+}12, IHMG22, KKL^{+}20, LWW^{+}23, LMSC19, Ler16, MKKS18, McK20, MAP23, MKHR21, NKSF24, PL22a, Pet20, PHG^{+}20, PHGZ20, PPDC22, RFZ22, SSR^{+}21, SOA^{+}20, TL24b, Tsi14, VKV^{+}22, YB22, YCHG18, YLX^{+}24]. Human-Centered [VKV^{+}22, YB22]. Human-Computer [CB16]. Human-Like [CYO^{+}17]. Human-Smart [McK20]. Humanities [LS23, WKF^{+}21]. Humanoid [PS17, Rec21]. Hunting [RKY20]. Hybrid [AOI^{+}23, AMEF21, AZS23, ABS20, BYG22, BD24, CBD^{+}22, ENT21, FLM23, GMA20, JUN^{+}24, JLZ18, JHL^{+}18, KPT^{+}24, KM24, KPCS13, KTD24, KRSF22, LSC^{+}17, LSH21, LHW24, LLBB18, LLB24, MA23, NMH^{+}21, PC23, RGM^{+}21, SGG^{+}22, SVGGO24, SCA^{+}19, SSDS23, SRS^{+}24, SGW22, SKHK23, TL24a, TRHU23, dSVJdALRfS21, WLL20, WC20, XY18, YW16, YMSP24, YL21, YLK^{+}19, ZOAR24, ZY20a, ZHW23, ZZZL23b]. Hybrid-Copy [LSC^{+}17]. Hybrid-Feature [RGM^{+}21]. Hybrid-Powered [ENT21]. Hybridizing [RFD23]. Hydrakon [PC24]. Hydraulic [FCR23]. HyDSMaaS [dSVJdALRfS21]. Hyper [CALHG^{+}24]. Hyper-Distributed [CALHG^{+}24]. Hyperledger [LLHW20, LLW21, PORM^{+}23]. Hyperparameter [OMD^{+}22]. Hyperparameter-Optimized [OMD^{+}22]. Hyperthermia [OP22]. Hypothesis [RBVV22].

I/O [JLH^{+}13]. I4.0 [dSFdRP24]. I4.0I [ZCdRR^{+}23]. IaaS [FFM^{+}23, XCSM18]. IBM [GMMI16]. iCaps [KYS21]. iCaps-Dfake [KYS21]. ICN [Fot20, GZ23, JLGF19, LNA17, LYD22, LL23b, LHW24, XDH24, ZHW23]. ICN-Based [GZ23, JLGF19]. ICNs [LARPHA23]. ICS [MEJA23]. ICT [RRS20, Hua15, TMTMB20, VLBRMP21]. ICT-Mediated [TMTMB20]. ID [MMS22]. Ideas [NSJGN21]. Ideation [JSJM23, OR24]. Identifiability [DCPG21]. Identification [AMY24, CZL21, CXLB21, DMS^{+}22, IAG^{+}21, KRSF22, LLY24, LTTS21, MXL^{+}21, PPR^{+}20, PMP22a, RKS^{+}21, SDDB24, Tag20, WSD22, ZCZ17]. Identified [LPPG22]. Identifier [LHW18, MHK13]. Identifiers [AAF^{+}22, FSXP22]. Identify [BMB^{+}23a, CS22, JLW19, MPK^{+}23, ZHZ^{+}19]. Identifying [AAK24, FMCT09, FDN^{+}24, HSSPTM^{+}23, MKKG19, MNH^{+}24, PD23, WLH^{+}17, WZL23]. Identities [KHN^{+}11, LBC18]. Identity [ASWAE23, AYH22, CTM21, CT24b, FR15, Gla11, GK21, HOPGRV21, ITZ20, Li23, LWZ^{+}21, POC20, RSG24, SG12, SZZL23, VL24, VKKG22, VFFHF19, WD24]. Identity-as-a-Service [VFFHF19]. Identity-Based [FR15, GK21, Li23]. Identity-Enabled [SG12]. Idling [Kra20]. IDS [HCXH16, AAAT21, GSP15]. IEC [DSM21]. IEEE [GCEOBRT23, MSK^{+}18, SE23, SIJA10]. iForest [ASF^{+}23]. iForest-Based [ASF^{+}23]. IgA [ZWS20]. II [KKTS22, SG23, Tos23]. IID [EKKS24, LCC^{+}24]. IIoU [RES23]. Ill [VI16]. Ill-Structured [VI16]. Illicit [KS19]. Im2Graph [GGD23]. Image [AKYP23, Alb21, AGA^{+}22, BS17, CGnCD20, CCC19, Chu23, FLC21, GRRZ18, KTS18, KYS21, LJWX23, LGCM24, LPM10, MMD22, MTF^{+}21, Rei23, SGW22, SWC^{+}23, SS17, SKHK23, TCB^{+}19, USS19, WYL18, ZWWW17, ZWZ17, ZQCC16]. Image-Processing [Chu23]. Imagery [BPLBM21, CL19, GMLS22, IMLL24, KKM^{+}22, MMD22]. Images [CCKH21, CDD^{+}21, Cop14, DYLZ21, GR22a, JMZ^{+}22, KHP^{+}22, LDd^{+}24, LBK22, PPP^{+}23, VR21, ZMDCEMI22]. Imaging [XZ19]. IMBA [ASL24]. Imbalance [KM24]. Imbalanced [BMB^{+}23a, JXA19, QQM24]. Imitation [JGY^{+}24]. Immersive [DHEK19, FOJE19, KTS18, ML20]. Immigration [VLPR20]. Impact [AP21, BCT21, FC19, Gao21, GL21, Hel15, IUK^{+}23, KDJ24, KPC19, KKS^{+}19, LL20b, MPG20, MGBG21, NKSF24, PMGG21, PKB^{+}24, RSV^{+}24, SLS20, SS22, TRLR24, VMD^{+}22, YB22, YLK^{+}19]. Impacting [APB^{+}16]. Impacts [HSM19, WHB^{+}21]. Impaired [FBOC15, ZZZ^{+}22]. Imperfect [ALB24]. Imperviousness [GMLS22]. Implementation [AGO^{+}23b, BTBK21, DTL^{+}21, dSFdRP24, GBKN23, HBM^{+}21, IAG^{+}21, JEJ^{+}24, JLZ18, JJ21, KPP^{+}20a, KPCS13, LZL^{+}23, LXQ24, MV20, MMF^{+}19, MFH22, MMMH19, NPM^{+}18, RL24, RNST23, TPD^{+}20, Vem22, VCGOMC^{+}21, WYL^{+}21, YFY24, ZYWS24]. Implementations [AP22, SG19]. Implementing [Ant24, BD24, BCS24, CHH^{+}16, DCL24, FR24, GdSdC23, HLLT10, MTF^{+}21, SE15]. Implications [DN20, LIM23, MZ19, RRAGMMGG20, Sta23, ST20]. Implicit [HL23, WYX18]. Importance [BG24, SOSC^{+}16, YPLZ19]. Importance-Based [YPLZ19]. Important [LZLW22, ZVKK19]. Improve [CP10, CPRS20, CMR24, HHL24, KKM^{+}22, LPM10, PPN18, TMP22, YT19, ZBN21]. Improved [ADM^{+}19b, AK22, ADAq^{+}22, AFAAM19, BAS24, GXWD19, JJFC22, JHL^{+}18, KES24, KC19, LLZL19, dMLMS22, LLSS24, MKR22, MA23, QLK^{+}24, RRMI22, RR24a, RFD23, SGG^{+}22, TL17, YMH22, ZSZ18, ZL20].

Improvement [DVSBV^{+}24, DGF23, HRAA19, JHC^{+}20]. Improvements [LK20]. Improving [AK22, ACBP19, BCZ^{+}23, BRC24, CMO^{+}21, CCD^{+}21, DHDAu22, GPRMGP21, GKW^{+}10, HM20, KPH^{+}24, KID^{+}15, KL21, LWY20, MLM14, PHGZ20, PW22, RRP17, SKI^{+}19, THL21, VCPPRC20, ZSF20, ZVKK19]. Imputation [AAAHA22, AAAIA24, PHGZ20, ZZ21]. IMTIBOT [GKM24]. In-App [PRDK22]. In-Band [LZL^{+}23]. In-Depth [PMT^{+}23, ZDP^{+}22]. In-Home [COL21, MDS^{+}24]. In-Memory [CMO^{+}21]. In-Network [DH22, KKBD10]. In-Situ [GRRZ18]. In-Store [FRE^{+}22]. In-Transit [MDRR24]. In-Vehicle [MNP^{+}24, RYAA22, BCNS20]. Incentive [LHL24b]. Incentives [VSEH^{+}21]. Incidence [LCV^{+}17]. Incident [XGY^{+}22]. Inclusion [PE13]. Income [LRdLS19, MTHN22, PPM21, SS22]. Incorporating [JLW19, WQH23]. Increased [KPC^{+}23]. Increasing [HSC19, RSX18, Rók23]. Incremental [WZL23, WQH23]. Independent [AdMTJ20, TPD^{+}20, WY19]. Index [LNA17, LLCC24, TTS24, WLYL20, YZP22, ZLW21]. Indexed [YMJ20]. IndexedDB [PV20, PV22]. Indexing [ACG^{+}24, KBF^{+}22]. Indian [STV^{+}22]. Indicator [CCKH21]. Indicator-Free [CCKH21]. Indicators [AK22, BJL^{+}22, GPRMGP21, PC24, PPDC22, SSR17]. Indices [KZ20, LGCM24]. Individual [GDCM19, MZ19]. Individuals [PS17]. Indonesia [RHHN21]. Indoor [ACD^{+}22, CW23, DWZN16, EUDW23, FSY^{+}22, FHL^{+}24, Goe12, GZS^{+}24, HUM^{+}21, JK24, KTAH^{+}23, LLYL22, MCCHO19, dSMdO22, QLK^{+}24, RSMC21, Str22]. Induced [AYH^{+}21, GNV21, GLW^{+}19]. Industrial [BML19, BD24, CSIS19, Cap23, CMW18, CMH10, DSM21, DGF23, GS23b, Jal19, Kha24, KIB^{+}22, LG20, LXQ24, MSN13, PC23, Pil18, RSF23, RGP22, SU17a, SLS20, TFSBSPSC24, UB18, VL24, WXL^{+}23, YQ24, LMF23]. Industrie [PKCF21]. Industry [BFFZ^{+}12, CTM19, CALHG^{+}24, KXZ^{+}24, MBAS20, NSJGN21, PGSL20, AMVM22, CEZC19, DTL^{+}21, DPR^{+}21, dSFdRP24, JNARV^{+}24, Lel19, LH21b, LTP24, Pil18, ZCdRR^{+}23]. Inefficiency [GSP15]. Inequalities [CWF13]. Inequality [Gra13]. Inertial [DDV22]. Infeasible [AZ19]. Infection [VMAG23]. Infectious [Bad23]. Inference [BZMB23, JOL22, KKP22a, NDS^{+}23, PPKS21, TMS24, ZZZL23a, dOB19]. Inferring [GMB^{+}21]. Infinite [Tsi14]. Inflated [UJ22, ZSB24]. Influence [CLM^{+}21, DGGS19, ESLB20, LMK18, NKSF24, RZCL19, SMK^{+}22, WZLW19, YPZ18, ZVV12]. Influence-Based [SMK^{+}22]. Influences [HP20]. Influencing [PB23]. Influential [AAKS23, SSF^{+}18]. Infodemic [AAN^{+}21]. Information [AOI^{+}23, AZS23, ACM^{+}19, BCG^{+}23, BRP^{+}13, CP10, CZL21, CFG20, CTM22b, DGMP21, DGADE14, Del23, DNZ22, DNZW22, FLLG24, DB23, FZ15, FP19, Fot20, GLZL24, GL19, HZWJ21, HLZ23, HOV23, ICF^{+}11, KKP22b, LZ23, Ler16, LW11, LSH21, LYD22, LH21c, LCCC24, LHL24b, LARPHA23, MWY19, MGJ22, MHSH16, MMQ^{+}24, MZH22, MZLP22, NZZ13, NZ14, OT16, PS16, PVM22, PW21, PA13, PPDC22, PNdC^{+}23, RRS20, RBSMRN21, SYB^{+}18, SOSR^{+}16, Sof19, SM19, SNZ21, SYLL21, Tri14, Uto13, UUT^{+}19, WHB^{+}21, XNZ23a, XNZ23b, YHG^{+}21, Zha19, ZY20a, ZWZ19]. Information-Centric [ACM^{+}19, DNZ22, DNZW22, FP19, Fot20, LSH21, LH21c, LHL24b, LARPHA23, MZLP22, SYB^{+}18, SM19, SNZ21, XNZ23a, XNZ23b, ZWZ19]. Information-Driven [Sof19]. Informational [KDD^{+}21]. Informed [Pic21]. Infrared [GR22a]. Infrastructure [AMC^{+}24, CSIS19, CB22, FHL^{+}24, FN23, GMD^{+}22, GMGF24, HEP^{+}11, KVWC24, KMS^{+}12, KDJ24, MDB22, NPB^{+}19, OPGH23, SSS^{+}19, Tas10, dSVJdALRfS21, VFFHF19, XSM22]. Infrastructure-Limited [AMC^{+}24]. Infrastructures [CCC^{+}24, DPC^{+}24, HDT22, HBMS20, KGKP20, MVMHL24, SCB21, SHB23]. Infringing [HSSPTM^{+}23]. Inhibition [GGCY20]. Initial [IPL^{+}18]. Initialization [DDD^{+}22]. Initiatives [OO19, VCB14, VM24b]. Injecting [MMT24]. Injection [AV23, GHL^{+}23, SCA^{+}19]. Innovating [DCD24]. Innovation [ACK^{+}16, BJL^{+}22, CMK^{+}16, CGMVdUC19, KA23, OGR^{+}20, PKCF21]. Innovations [GMRRRQ^{+}16, RAC24]. Innovative [GX20, LLCC24, LCCC24, MGBG21, SIZD12]. Inpainting [SSD24]. Input [AFAAM19, BAK23, CFX20, oRL19, Smi15, SZW^{+}23, ZCG^{+}24]. Insider [SZZL23]. Insights [BBH^{+}24, GSL22, PFT^{+}24, TLV23, YBMK21, VDE^{+}20]. Inspection [SB23]. Inspired [ASL24, CCA^{+}20, Min20]. Instagram [AAAR24, BGW16, HS18b, IP16, PV22]. Installation [SOSR^{+}16]. Instance [YLJ^{+}19]. Instances [VM24a]. Instantaneous [JMZ^{+}22]. Institution [APAR22]. Institutional [GPRMGP21]. Institutions [BRF^{+}23]. Instrument [RL24]. Instruments [CCCA22, PS16]. Insulators [KKM^{+}22]. Insurance [GLD^{+}18, MMN19, RH24]. Insurer [CS22]. Integer [GVC13, PTZM22]. Integrated [CMP^{+}19, Chu23, FKZ22, FHZ^{+}24, FMNS11, IAG^{+}21, KYS21, Li18, LLL21, MCS^{+}24, MKM^{+}19, PP22b, QLL^{+}21, RLB^{+}23, SP22, VLC^{+}19, ZLL18]. Integrates [YQ24]. Integrating [Al 20, AHP23, BGL^{+}22, CISM22, IP16, JP21, KZRH22, KVVD22, MLWS24, PEN23, QA24, ZC21]. Integration [AKT24, AM23, AFB23, AAAKJ22, ASFAAV24, AGO^{+}23b, AA24b, DMZ^{+}19, EWGK24, HSI^{+}22, KAAAA22, LWH24, LG23, OHSN24, SOSK12, Tri14, VLC^{+}19, ZBK^{+}24]. Integrative [RHHN21]. Integrity [HP19, LB24, XGN19]. Intellectual [CWF13, PLA^{+}21, PKCF21]. Intellectually [FBOC15]. Intelligence [AK19, AAAAM^{+}21, AEM^{+}23, BG24, BRC24, CCTC23, CL19, DVSBV^{+}24, DN20, DRK^{+}22, DWL21, EWGK24, ELCS20, FLLC24, GRJ^{+}19, GB23, ISG20, IG10, UMS24, KS24, Khr20, KV23, Lai24, MDS^{+}24, Min20, MGLBMMSC20, MZLP22, OC20, OP22, PSTZ23, PMP^{+}22b, RF24, RTN^{+}22, SCB21, SKI^{+}19, SG19, SK24, Suf23, SYLL21, VTPT24, VL24, VMAG23, VCGOMC^{+}21, VCJAMN22, YWZ^{+}24, dCdG20]. Intelligent [AFS^{+}22, AYH^{+}21, APP22, AVE^{+}22, BF18, CNC^{+}21, DPH17, DPH19, DRK^{+}22, Dua21, FFVP21, GKM24, IHCSG^{+}24, IL14, JKM^{+}23, KE22, MATP19, MVMHL24, MSAD22, MKM^{+}19, PSC19, SYB^{+}18, SBRZ24, SRS^{+}24, SA21, SOK24, SF22, TCH23, VP22, VFS^{+}19, WMS23, WCF^{+}16, WWD21, WUH^{+}23, ZJ18, ZZZ^{+}23, ZX24, ZGL20, ZLL^{+}21, ZZY19]. Intended [CALHG^{+}24]. Intensity [CMW18]. Intensive [CMO^{+}21]. Intention [HSM19, HC19, WLL^{+}19]. Inter [BKL^{+}19, CXG^{+}22, HHL24, Lee18, YT19]. Inter-Frame [Lee18]. Inter-Packet [BKL^{+}19]. Inter-Satellite [CXG^{+}22]. Inter-SubNet [HHL24]. Inter-Vehicle [YT19]. Interaction [CB16, DMP^{+}23, FR16, GNNZRCRG18, HC19, KYT^{+}23, LWH24, LF23, RGP22, SU17a, SFD20, SN23c, ZCG^{+}24, ZLOW24]. Interactions [BM20, BMGI21, DeD16, FNN21, MS20, McK20]. Interactive [AA23a, OOM^{+}18, SC20, SFW23, SPB^{+}11, SKA^{+}23, VRA^{+}18, YYY^{+}23, ZZZ^{+}22].

Interactivity [Zha19]. Intercell [TFKY19]. Interception [FPGO24]. Interchange [LDM22]. Intercultural [OMTFFL21]. Interest [AAK24, AYH^{+}21, DTTK19, HS18b, RGD24, Sor12]. Interestingness [MGFF14]. Interface [BS17, BHH12, FHL^{+}24, JMZ^{+}22, MAP23, VVK^{+}21, ZLOW24]. Interfaces [BPLBM21, LB23]. Interference [AZ19, UGJA18, HJA18, KSG24a, KTP17, KOJP13, Lee17b, LZL^{+}23, TFKY19]. Interference-Aware [Lee17b]. Internal [KSW21]. International [GSB^{+}23, LMY21]. Internet [FC19, LMF23, LJR24, VBB24b, WYW^{+}22, AMK^{+}21, AOI^{+}23, AAM^{+}24, AAAHA22, ALB24, AAGB^{+}20, AM23, AlQ23, AA20, ACF24, AA19b, AA23b, AAAKJ22, AS15, ASN22, ABB23, AMY24, BML19, BRF^{+}23, BBMR19, BTS^{+}21, BZDP19, BFL20, BCFP21, BYS^{+}15, BMGI21, BLU^{+}15, BCM^{+}19b, CEP24, Cal24, CPM^{+}19, CMK^{+}16, CFV19, CWF13, CVG^{+}22, CMW18, CLW23, CZY17, CFPP21, CC24, DKH24, DDPVP18, Del21, Del23, DWY^{+}19a, DCPG21, EAHO21, Far12, FR24, FBDR23, FMA^{+}20, FDN^{+}24, FPKK22, Fri13, Gal24, Gao21, GZC21, Gog12, GS11, Gro22, HDBD23, HRAA19, HEP^{+}11, HT14, HZQ24, Hel15, HF14, HKH^{+}24, HS09, IG10, UMS24, Jal19, JUH^{+}23, JMJ^{+}23, Jia19, KTCI21, KDG24, KYT^{+}23, Kha24, KPH^{+}24, KHRG19, KH10, KST24, KNN^{+}24, KA23, Lee20, Lel19, Li18, dMLMS22, LZL^{+}17, LLW^{+}24, LIM23, MKSK23, MGJ22]. Internet [MAEP18, MR23b, MTK^{+}24, NA23, NDKBL19, NJH^{+}24, NBL^{+}24, NZU20, NSJGN21, Off14, OMD^{+}22, ORK^{+}19, OC20, PAM^{+}24, Pet11, POC20, Poz24, Pri10, QSF^{+}17, RGB^{+}23, RR24a, RAU^{+}24, RIS^{+}17, RGdSK22, RKT19, RRdBSS22, RRAGMMGG20, RR24b, RSL^{+}24, BNK^{+}23, SCB21, SJA^{+}23, SVK^{+}22, SS23, SPK24, SDM21, SNKG20, SMN22, SST23, SPT23, Söd13, SGDT19, TSD24, TN20, UB18, VC19, VCPPRC20, WZL^{+}20, WXL^{+}23, WSD22, WY19, ZR15, ZZZ^{+}23, ZX24, ZJX18, ZSF20, ZZY19, Fut20, Fut21, Giu18a, Off15, Off16, Off17, Off18, Off19, Off22]. Internet-Based [Gao21]. Internet-of-Things [AMY24, CLW23, FMA^{+}20, HDBD23, LLW^{+}24, OMD^{+}22]. Internet-Supported [Del21]. Internets [Gog12]. Internetworking [Jia19]. Interoperability [AFS^{+}23, CNC^{+}21, HG12, JJ21, KMS^{+}12, Lel19, RKM^{+}22, RFS22]. Interoperable [MEH^{+}22, PM24, SSH20]. Interplanetary [BMS24, LB24]. Interplay [PNS23]. Interpretability [CT20]. Interpretable [FAS^{+}23, SLZY21]. Intersection [AAN^{+}24, YT19]. Intervals [AM21, NRS24]. Intervention [PTK^{+}23]. Interventions [TRLR24]. Intra [EMHF19]. Intra-Prediction [EMHF19]. Intransitiveness [BB20]. Intraoperative [IKS^{+}21]. Intrinsic [BdCRM10, SSR17]. Introducing [Ant24, YAV^{+}21]. Introduction [CB16, Far12, Fer12, FK24, Gra13, Sac19]. Intruders [WCF^{+}16]. Intrusion [AGO23a, AR23, AA23b, AA21c, BMB^{+}23a, BCNS20, CCYL24, FRR^{+}24, FMA^{+}20, HCXH16, JEJ^{+}24, KM24, KPH^{+}24, LMK18, LNT^{+}23, LNT^{+}24, MAVS20, MRM^{+}24, Mou24, NB24, NLT^{+}23, RLB^{+}23, SQK^{+}17, TMP22, TLK^{+}20, TTS24, TWW24, VOP22, WQH23, ZHZ^{+}19]. Intuitively [LWBM22]. iNUIT [CMK^{+}16]. Invariant [MKKS18]. Inventor [PJMM22]. Inventory [MCS^{+}24]. Inversion [CR21]. Invertible [TDT^{+}21]. Investigating [AHH^{+}23, FTHY21, GPLK22, HP20, LMK18, LCY^{+}23, SMG13, SRM20]. Investigation [BHM^{+}19, ELCS20, EVCL21, FLLC24, HBMS20, PLOT23, RNFS20, SMVP21, SK23, TPP^{+}23]. Investigations [PV20, PV22]. Investor [CS22]. Investors [FCG22]. Involved [ZCdRR^{+}23]. Involvement [SS14]. IoCV [ASMM22]. IoE [SCC^{+}23]. IoET [AAN^{+}24]. IoH [GRJ^{+}19]. IoMT [RGB^{+}23, Adi23, AALS21, AAN^{+}24, PORM^{+}23, SESO24]. IoMT-Based [PORM^{+}23]. IoRT [MKSK23]. IoT [ABB23, HKH^{+}24, Lee20, NZU20, TWW24, TN20, VL18b, AMK^{+}21, AN23, ATE24, ABCC22, AHKC23, AK22, AAA^{+}24a, AG22, Alb21, ADAB21, ASL24, AAN^{+}24, AF21, ARC^{+}19, ASN22, Ant24, ALDS19, AMI23, AKJB20, ABLR23, AHS^{+}23, BM20, BYG22, BP24, BBGP19, BHS24, CTP^{+}22, CTNS21, CR23b, CLH^{+}23, CFP19, CDBF19, CTM22b, CALHG^{+}24, CCYL24, DTR22, DNKDN^{+}24, DHDAu22, DGD^{+}19, DT24, DWL21, EKG24, EAAA^{+}23, FOJE19, FWLY23, FRR^{+}24, FNN21, FSXP22, GSL20a, GPPB^{+}21, GKM24, GMD^{+}22, GHA^{+}21, GGT^{+}23, GBKN23, GSL20b, GAA21, GTWN24, HA23, HBM^{+}21, JRP21, JEJ^{+}24, Jeo20, KZRH22, KAEAB24, KPT^{+}24, KAM^{+}19, KGK^{+}24, KCT^{+}24, KAS^{+}22, KRSK24, KGWR23, KBF23, KBF^{+}22, Lee20, LWY^{+}15, LNT^{+}23, LNT^{+}24, LW21, LXQ24, LA20, LG23, MJG24, MGM^{+}23, MMS22, MKK17a, MKR22, MOAI17, MPL^{+}21, MCPLBH^{+}24]. IoT [MKM^{+}19, MSJ^{+}23, MDRR24, MHQ^{+}23, MS22, NLT^{+}23, NPM^{+}18, NPB^{+}19, PKF24, PGFH24, PVM22, PPMMS19, PPP^{+}23, PLG23, PPP21, QAQA21, QNJ24, QXC^{+}21, RC18, RRMI22, RLB^{+}23, Ray23, RNST23, SBN^{+}23, SGG^{+}22, SBRZ24, SKKS22, SFW23, SA21, SOK24, SH20, Sta22, SG23, SNS22, TAD23, TFSBSPSC24, TS22, TZ22, THV^{+}22, TWW24, TPD^{+}20, TN20, TCYZ22, VDE^{+}20, VBB^{+}24c, WZL23, WX24, YWLY22, YHN21, YZC^{+}20, ZKKK19, ZGL20, ZDP^{+}22, dOB19]. IoT-Based [AMI23, BYG22, DHDAu22, GSL20b, GAA21, LW21, MCPLBH^{+}24, PLG23, VBB^{+}24c, YWLY22]. IoT-Blockchain [SNS22]. IoT-Driven [GTWN24, KGWR23]. IoT-Edge-Cloud [GGT^{+}23]. IoT-Enabled [Alb21]. IoT-Enhanced [QNJ24]. IoT-Fog [BP24]. IoT-Mist [ASL24]. IoT-NDN [PGFH24]. IoT-NOMA [THV^{+}22]. IoT-Oriented [CDBF19]. IoT-Portrait [WZL23]. IoT-Scenarios [Ant24]. IOTA [GPPB^{+}21]. IoTs [Lee17a, Lee17b]. IoTSRM2 [PPP21]. IoTwins [BD24]. IoU [RES23]. IoV [NA23, KT22, MZLP22, WZS^{+}22]. IoVT [XNC21]. IP [Al 20, HP16, LHW24, Pin10]. IP-ICN [LHW24]. IP/Optical [Pin10]. IPsec [LBA^{+}24]. IPTV [KKBD10, LCY^{+}23]. IPv6 [HF14, IKP^{+}13, SAR24]. IPv6-Based [SAR24]. IR [MXL^{+}21, QNJ24]. Irradiance [KLKP19]. Irrigated [CGRV20]. Irrigation [KKEV17, QQX^{+}24]. ISA [BGL^{+}22, SU17a]. ISA-95 [SU17a]. Isolated [GKG19, NIK^{+}21]. Isolation [GGK18, MNP^{+}24]. Issue [Bod22, CB16, EMMP22, Far12, Fer12, Gra13, JP13, Mac22, PSC19, Sta22, SG23, Tos23, ZMKR22]. Issues [AS19, AIR^{+}18, AYH23, AAD^{+}12, DWL21, GPCP12, GZC21, GGT^{+}23, KBF^{+}22, MSFM16, VKK^{+}19, XKKL23]. IT/OT [FMMS21]. Italian [AGO^{+}23b, BM19, BBM20, CGMM22, GSD^{+}17, LC23, MPG20, VLPR20, ZSS^{+}22]. Italy [CEP24, FZ15, GFPD21, TC21, ZL14]. Item [DDD^{+}22]. Iterative [AKDL24, LXY17]. ITU [VM24b].

J [FC19]. Jammer [CLH^{+}23].

Jammer-Assisted [CLH^{+}23]. Japan [CS11]. Japanese [LWBM22, UFK^{+}10]. Java [MT23]. JavaScript [AKAS22]. Jitter [LW15]. Job [SZT23]. Join [CFV19]. Joining [TYH23]. Joint [AFB23, BZNB23, ENT21, FK24, JLGF19, LTY18, LLW^{+}24, MMS22, NBT22, OAI^{+}22, SXXH19, SHN^{+}19, YPXH19, YXL^{+}21]. Jointly [LXL^{+}20]. JoSDW [ZXXB22]. Journalism [DN20, KVVD22, NSD^{+}22, RBSMRN21]. Journey [FRE^{+}22]. JPEG [PTZM22]. JTAG [VL18b]. Judgment [MBLC22]. Jump [ZYLM17]. Junctions [PCC^{+}18]. Jurimetrics [CBB17]. Just [Alb19]. Just-In-Time [Alb19]. JVM [BM23]. JVM-Based [BM23].

kClusterHub [GOM23]. Keeping [ZLY^{+}20]. Key [AKK^{+}24, AZH22, AMI23, BBSR24, CSIS19, DNKDN^{+}24, FLM23, GR22a, HEP^{+}11, HDT22, Li23, LCMV17, MKK17b, MPK^{+}23, PNMK22, PL22b, PAM^{+}24, SPM^{+}22, VCPT22, WHW22, ZOAR24]. Keyword [BC23, HWZP18, ZZL^{+}23]. Kim [FC19]. Kits [DG18]. KNN [LBBPM21, SGMBK24]. Know [KPP^{+}20a]. Knowledge [AMR^{+}21, AFS^{+}23, BdBC23, BGL^{+}22, CYC^{+}21, CQWC22, CH23, FAB^{+}22, FM21, For14, ICF^{+}11, KS19, KVL^{+}22, LHD^{+}24, LTM^{+}22, MSL^{+}23, OTR^{+}20, PG14, PLL11, PLA^{+}21, PKCF21, Por21, RFZ22, RRT^{+}23, RRT20, RTN^{+}22, SU17c, SPGS23, SLSK21, SCM12, TLV23, WX22, WJBZ22, YCD^{+}23, ZL20, ZC21, ZY20b, ZGH13]. Knowledge-Driven [RRT20]. Knowledge-Graph-Based [CYC^{+}21]. Known [KAAAA22]. Kohonen [BCNS20]. KPIs [SPSS17]. Kr00k [SE23]. Krack [SE23]. KubeHound [DSB23]. Kubernetes [DSB23, PPKS21, PL22b, TXJ23]. kV [AMZ^{+}20]. KYC [KPP^{+}20a].

L [WKW18]. L-DDoS [WKW18]. L2S [GVV^{+}23]. L2S-M [GVV^{+}23]. Lab [ACK^{+}16]. Labeling [KBP22, LBK22, MLWS24, OFK19, SSS21]. Labels [CLTP22, SGW22, SHA^{+}21, ZXXB22]. Laboratories [ZJ18]. Laboratory [GPM21, HZWJ21, RHHN21, RSF23]. Labs [GTWN24]. Lagrange [HKZ^{+}22]. Lake [DGHH^{+}20]. Land [MGUD12]. Landscape [BR23, KXZ^{+}24, PR11, SPC^{+}11, SPB^{+}11]. Landscapes [CR23b]. Lane [LRD23, SYGY21, WZG^{+}18]. Language [AdMTJ20, AM24, ADAq^{+}22, CS22, CMR24, EMMP22, FRKS22, GPK^{+}20, HYMT24, HOV23, HZCZ22, JSPH21, KI20, KKT20, LSSD22, LHD^{+}24, PLYD20, PCVH24, PD23, RKS^{+}21, RCL21, STV^{+}22, SZT23, SG21a, SB23, VCG24, YMSP24, ZSS^{+}22, vSR24]. Language-Independent [AdMTJ20]. Languages [BTB22, MKK^{+}22]. LANs [Lee18]. L’Aquila [ADG^{+}19]. Large [ACG^{+}24, AM24, CMR24, HYMT24, HP21, JXA19, KGK^{+}24, PCVH24, SADP21, SKKS22, SZT23, VCG24, XSSA24, YMSP24, vSR24]. Large-Scale [JXA19, KGK^{+}24, XSSA24]. Largest [Kra20]. Latches [SDS^{+}22]. Latency [BNJ24, BBH^{+}24, IFF21, MLPH23, MM21, PBK22, PP22b, VKM^{+}19, YSAY23, YLK^{+}19]. Latency-Aware [YSAY23]. Latency-Based [YLK^{+}19]. Latent [LWH24, TRLR24, ZY20b]. Latest [HAL18]. Latitude [Mat20]. Lattices [GB20]. Law [LAG^{+}18, PPM21]. Layer [AKM^{+}23, AAF^{+}22, BOK^{+}21, CCKH21, CAL23, GVV^{+}23, GTC19, GFL21, GVC13, HNK^{+}21, IT24, MMS22, ORK^{+}19, PKB^{+}24, RF21, SSD24, SKT^{+}19, TH21, TSDG22, WXL22, WX24, AKK^{+}24]. Layer-Wise [GTC19]. Layered [ASMM22, QWR18]. Layers [LL23b]. LBCS [MGUD12]. LBP [SKHK23]. LDA [LBBPM21]. Lead [RIS^{+}17]. Lead-Acid [RIS^{+}17]. Leak [AOF24]. Leakage [AZS23, SZZL23]. Leaks [FKK20]. Learner [CWC^{+}22, VI16]. Learner-Managed [VI16]. Learning [AMK^{+}21, AIP^{+}24, AZAV22, AKDL24, AN23, ABSG21, AKM^{+}23, ASBC22, AM23, AAA^{+}24a, AGO23a, AG22, AN22, Alb19, AA21a, AAAR24, Alh22, AAG^{+}22, AR23, AA23b, AAA^{+}24b, AA21c, AAJ^{+}22, APJ24, AGN21, ABBV23, BARL24, BSM22, BHSAY24, BHC23, BOH23, BHS24, BCMPW21, CD23, CMO^{+}21, CGnCD20, CXLB21, CHMZ21, CQWC22, CLT^{+}22, CWC^{+}22, CSA^{+}24, CASN21, Chu23, CMR24, CGBF24, DXL^{+}22, DKV24, Del23, DZH^{+}22, DIE^{+}21, DT24, ESB^{+}20, EKKS24, EAAA^{+}23, ERBL12, FOJE19, FWLY23, FHXL19, FQ21, FLLG24, DMP^{+}23, FRKS22, FHZ^{+}24, FFP^{+}22, FMCT09, FM20a, FDN^{+}24, FFM^{+}23, FK24, Gao21, GWF^{+}23, GDN^{+}23, GHA^{+}21, GVY^{+}23, GFPVLR^{+}23, GHGL22, GHL^{+}23, GMGK20, GTWN24, GPM21, HHLZ23, HHM^{+}23, HXHP17, HOPGRV21, HBM^{+}21, HN18, Hua15, HRGQHOA21, IHCSG^{+}24, IAM24, IHMG22, IDR23, JXA19, JUN^{+}24, JOL22, JMJ^{+}23, JJFC22, JEJ^{+}24]. Learning [JHC^{+}20, JHBW24, JGY^{+}24, JGV^{+}21, JSAO^{+}23, KPT^{+}24, KM24, KSG24a, KPRP24, KALY17, KRSK24, KK23b, KPH^{+}24, KBF23, KPKM23, KID^{+}15, KTD24, KV23, Lai24, LCC^{+}24, LS15, LWX18, LG20, LLLD21, LZLW22, LJR23, LJR24, LLZ21, LT24a, LHL24a, LL20b, LPA^{+}22, LL20c, LYMG17, MCS^{+}24, MB22, MAW^{+}23, MMD22, MAW^{+}19, MSC^{+}24, MMS22, MMM21, MKSJ22, MDM21, MRRMC21, MFH22, MIK^{+}23, MCR^{+}23, MSAA22, MSC^{+}21, MN21, MA23, MKM^{+}19, MP13, MF20, NLD^{+}24, NBL^{+}24, NST23, NAFM^{+}24, OOM^{+}23, OLMBRCRC23, OOM^{+}18, OHSN24, PLL20, PKF24, PBP^{+}24, PP22a, PFT^{+}24, Pec18, PHG^{+}20, PLA^{+}21, PSTZ23, Pri23, QLK^{+}24, QSML24, QWR18, RGB^{+}24, RM23, RGB^{+}23, RSMC21, RLB^{+}23, Rec21, RNST23, RDASB22, RFD23, SGG^{+}22, SE23, SLP^{+}22, SESO24, SPC22, SE15, SGW22, SOK24, SPK24, SJ12, SF22]. Learning [TAD23, TCH23, TDM^{+}24, TSZ^{+}18, TMP22, TFKY19, TCH18, TLK^{+}20, TS22, TMTMB20, TKT^{+}18, TCYZ22, VCPT22, VCRCSV^{+}21, VOP22, VC23, VMD^{+}22, VI16, WLL^{+}19, WLG^{+}22, WZL23, WLLY24, WSD22, Whe09, WCYL20, WZL21, WLWZ23, WQH23, WRHH23, XLW^{+}24, XZY^{+}24, XLJ^{+}22, XLM^{+}23, XSSA24, YBO^{+}23, YQ24, YSAY23, YLJ^{+}19, ZKKK19, ZWS20, ZPX24, ZX24, ZZSX19, ZHH24, ZY20b, ZXWZ21, ZJJS^{+}23, ZTBD20, ZSS^{+}22, GLW^{+}19, ESB^{+}20]. Learning-Assisted [AIP^{+}24]. Learning-Based [BHS24, BCMPW21, GWF^{+}23, JXA19, JEJ^{+}24, KSG24a, KPRP24, MAW^{+}23, MKM^{+}19, QLK^{+}24, RLB^{+}23, RDASB22, SE23, SOK24, WLLY24, ZJJS^{+}23]. Learning-Enhanced [SESO24]. Learning-Powered [CGBF24]. Learning-Using [VI16]. Learnt [KG18]. Ledger [ACA^{+}21, FK21, SPS^{+}21, Tre21, VDSK23, XNC21]. Legacy [YLK^{+}19]. Legal [DN20, Fra18, HWCH12, LAG^{+}18]. Legged [MKSK23]. Legislative [SS14]. Legislators [Ves18]. Legitimacy [WD24]. Length [TRHU23]. Length-of-Stay [TRHU23]. Lens [AKYP23]. LEO [ISLARP^{+}22]. Lessons [KG18]. Level [AZAV22, CGM^{+}19, GKFV24, LMLW18, MMN19, MKK^{+}22, MCPLBH^{+}24, MPK^{+}23, SNZ21, SZC19, VDSK23, XY21, XSSA24, YWS22, ZL19a, ZLZZ23, ZGZ^{+}18]. Levels [GBDDVS23]. Leveraging [ABH24, FCG22, GSB^{+}23, JWD^{+}24, MPK^{+}23, NJH^{+}24, RKT19, SSF^{+}18]. Lévy [FKZ22]. Lexical [ND24]. Li [LJR24]. Librarie [TG13]. Libraries [BM23, IP16, ND24]. Library [Cop14, SPB^{+}11]. LibreSocial [MKDG20]. License [RPB^{+}17]. Lie [RDASB22]. Life [MK20, NRS24]. Lifelong [MRRMC21, VCPT22]. Lifetime [TAMA22, ZBN21]. Light [HLG21, KC19, PCC^{+}18, QNJ24]. Light-Weight [KC19]. Lightweight [ADM19a, AMI23, BL22, BMP21, FRR^{+}24, JEJ^{+}24, MLPH23, NBL^{+}24, SBN^{+}23, SBS18, SFEK18, SMN22, WHW22, XNC21, ZZ18, ZMZ^{+}22]. Like [CYO^{+}17, RHV17]. Likelihood [SJG18]. Limit [Chu23]. Limitations [ABB23, CR11, Del23, KTCI21, PGSL20]. Limited [AMC^{+}24, BZDP19, BM18]. Limited-Resource [BM18]. Limits [Ler16]. Linda [DGGS19]. Linear [CC23, LLY23, LBB23, MR23a, MS22]. Lines [DNPC^{+}22, KKM^{+}22]. Linguistic [AFAAM19]. Link [CAL23, FSC^{+}23, GDRCDBLÁ23, GVV^{+}23, KKP22a, SAR24, SRSDF23]. Link-Failover [SAR24]. Link-Layer [GVV^{+}23]. Linked [BB23, QA24, RAA^{+}18, SB22]. Linking [SW21a]. Links [LWZL22, Mat20]. Linux [SHB23]. Lipreading [ITH^{+}21]. Liquidity [DCD24]. Listed [LS15]. Listless [SS17]. Lists [DCG12]. Literary [DMS^{+}22, FRKS22]. Literature [AN22, AA21b, AAA^{+}20, BG09, CR23a, CGMVdUC19, ESBE23, GGW18, KT22, KIG23, KMV^{+}23, Lee20, LH21b, PAK^{+}23, RNFS20, SMVP21, SE15, YMSP24, ZC21]. Live [AGC18, LSC^{+}17, MSC^{+}21]. Live-Streaming [MSC^{+}21]. Lived [PMT^{+}23]. Livestock [GMRRRQ^{+}16]. Living [ACK^{+}16, ADCM24, CFG20, FAC22]. LLM [CN24, ND24, YLX^{+}24]. LLM-Generated [ND24]. LLMs [PCVH24, YWZ^{+}24]. Load [ANHMS24, ARAAA19, BdCRM10, GW15, HSL^{+}23, LW19, LZL^{+}21, MKM^{+}19, OT16, SAR24, SHBS19, YZCR24, YG20, YWW^{+}19]. Load-Balanced [BdCRM10]. Load-Sharing [SAR24]. Loading [LTP24]. Loans [APAR22]. Local [CPRS20, DQ18, DWZN16, DWY19b, GS12, HGF21, HY22, KKM^{+}22, MSR^{+}14, MHK13, RFSOHMSB21, USS19, Yan17, ZYLM17, ZQCC16]. Localization [AAG^{+}22, GZC21, JK24, KTAH^{+}23, LJWX23, dSMdO22, PG19, RPB^{+}17]. Location [BRM23, BHH12, DQ18, DLX^{+}19, GLZL24, GR22a, ISLARP^{+}22, JLGF19, KE22, KSA^{+}21, KH19, OFV21, QLK^{+}24, RSX18, RAK^{+}20, VKK^{+}19, WL22]. Location-Aware [BHH12]. Location-Based [ISLARP^{+}22, KH19, OFV21, RAK^{+}20]. Location-Dependent [JLGF19]. Location-POI [BRM23]. Locative [Øie12]. Locator [LHW18, MHK13]. Lock [CZZH15]. Lockdown [HGF21, MPG20, PBB20]. LOFAR [CHMZ21]. Log [SJG18, SW21b, ZYWS24]. Logging [Kra22]. Logic [CPP^{+}20, GL12, JP21, LA23, MM23, Tra18]. Logic-Based [MM23]. Logically [KSD^{+}23]. Logically-Centralized [KSD^{+}23]. Logistic [CN24, OMTFFL21, SQK^{+}17]. Logistics [CALHG^{+}24, SZZL23]. Lognormal [KP19]. Logo [PMP22a]. Logo-Based [PMP22a]. Logs [CLTP22, HPB^{+}24, MGM^{+}23]. Lombardy [TC21]. Long [AOF24, BKL^{+}22, DYS15, FM20b, KTD24, SSD24, SRSDF23, SLB24, TFSBSPSC24, GLW^{+}19]. long-haul [GLW^{+}19]. Long-Range [AOF24, FM20b, KTD24, TFSBSPSC24]. Long-Term [DYS15, SRSDF23]. Longevity [Jeo20]. Look [BFG^{+}23, KFP10]. Look-Ahead [BFG^{+}23, KFP10]. Looking [XGN19]. Loop [CZZH15, OTR^{+}20]. LoRa [AN23, ACA^{+}23, CAL23, GHZY23, LBB23, LA20, OAG21]. LoRa-Based [AN23, LBB23]. LoRaCommunication [KKEV17]. LoRaMAC [AN23]. LoraMesh [dSVJdALRfS21]. LoRaWAN [ADAB21, BZDP19, BPG19, EABE19, KHAM24, MCCHO19, VC19, dSVJdALRfS21]. Lose [Tsi14]. Loss [GS23a, RES23, SOS^{+}23]. Losses [HHL24]. Lot [TCB^{+}19]. Lots [WC23]. Low [ABK22, ABB23, BNJ24, BKL^{+}19, BCMPW21, CKFH22, CTP^{+}22, CXLB21, CXG^{+}22, DVSBV^{+}24, KKD24, KMK24, LW15, MLPH23, MCR^{+}23, MM21, PR11, SKKS22, TPD^{+}20, VKM^{+}19, WLHR13, WZL^{+}20, YT19, YWW^{+}19, ZLY^{+}20, ZLT22, ZMZ^{+}22, dCdPC^{+}21]. Low-Altitude [ZMZ^{+}22]. Low-Code [dCdPC^{+}21]. Low-Complexity [WZL^{+}20, ZLT22]. Low-Cost [CTP^{+}22, DVSBV^{+}24, PR11, TPD^{+}20]. Low-Dimensional [CXLB21]. Low-Jitter [LW15]. Low-Latency [BNJ24, MLPH23, MM21]. Low-Orbit [CXG^{+}22]. Low-Overhead [KKD24]. Low-Powered [ABB23]. Low-Priority [YT19]. Low-Rate [KMK24]. Low-Resource [MCR^{+}23, ZLY^{+}20]. Loyalty [GLRM19, SMB23]. LPG [MCPLBH^{+}24]. LRU [HNHH23]. LSSDNF [SMN22]. LSTM [AWUF19, EVCL21, GMA20, HSL^{+}23, PLA^{+}24, RAA^{+}22, TCYZ22, XGN19, ZWPL19, ZGZ^{+}18]. LSTM-CNN [ZGZ^{+}18]. LSTM-CRF [XGN19]. LTC [KE22]. LTE [ABK22, AMZ^{+}20, CTP^{+}22, DMZ^{+}19, NSV17, QQM24, TFKY19]. LTE-Advanced [NSV17]. LTE-R [QQM24]. LTE/4G [CTP^{+}22]. Lyrebird [AAM^{+}24].

M [FC19, GVV^{+}23]. M2M [ABK22, BL22, LBG^{+}14]. M2X [LSN21]. MAC [TPD^{+}20, HNK^{+}21, MM21, MSK^{+}18, ORK^{+}19, TSDG22, YLH^{+}17]. MAC-PHY [MM21]. Machina [LAG^{+}18]. Machine [AMK^{+}21, AIP^{+}24, AZAV22, AKDL24, AAA^{+}24a, AGO23a, AA21a, Alh22, AR23, AA21c, AFAAM19, APJ24, BSM22, CMO^{+}21, CMR24, CGBF24, DeD16, DZH^{+}22, DL19, DT24, EAAA^{+}23, FQ21, DMP^{+}23, FRKS22, FFM^{+}23, GDN^{+}23, GHA^{+}21, GVY^{+}23, GMGK20, HN18, IHCSG^{+}24, IAM24, IHMG22, IDR23, JEJ^{+}24, JSAO^{+}23, KSG24a, KPRP24, KRSK24, KK23b, KBF23, KPKM23, KTD24, KV23, LSC^{+}17, LSN21, LG20, LPA^{+}22, MB22, MMS22, MMM21, MDM21, MIK^{+}23, MSAA22, MN21, MAP23, MKM^{+}19, MF20, OOM^{+}23, PLYD20, PDT23, QLK^{+}24, QWR18, RGB^{+}24, RM23, RGB^{+}23, RSMC21, RLB^{+}23, RDASB22, RKS^{+}21, RFD23, SE23, SLP^{+}22, TAD23, TSZ^{+}18, TLK^{+}20, TS22, WLCS17, XLW^{+}24, XLJ^{+}22, YQ24, ZKKK19, ZWS20, ZLY^{+}20, ZTBD20, GLW^{+}19]. Machine-Learning [LG20]. Machine-Learning-Based [DZH^{+}22]. Machine-to-Everything [LSN21]. Machines [CMO^{+}21]. Macro [OGB^{+}24]. Macro- [OGB^{+}24]. Madness [ATY20]. Magnetic [OP22]. Mail [SG21a, SKM21]. Main [MSFM16, TCG19]. Mainstream [SDM22]. Maintaining [TTO^{+}19]. Maintenance [CSZ^{+}24, FHL^{+}24, NZU20, NRS24, SOSR^{+}16]. Major [CBRS21, MKKG19]. Makers [Ves18]. Makes [BR21]. Making [ANHMS24, ARM20, AFS^{+}23, CGGK22, CMS21, DSMM21, DCPG21, Let16, MHSH16, MKPG22, MLL^{+}22, Orl22, PR12, SZC19]. Male [VYN19]. Malicious [AKAS22, AASAI22, ASLG22, SZL21, ZCZ17]. Malta [FMNS11, For14]. Malware [IAM24, KAS^{+}20, LCY^{+}23, OAH^{+}24, Uto13, YZZ^{+}20, ZJJS^{+}23]. Man [OOM^{+}23]. Man-in-the-Middle [OOM^{+}23]. Managed [NST23, VI16]. Management [AKK^{+}24, AM24, ATB^{+}24, ATE24, AML22, AZ19, Al 20, AlQ23, AAAAK24, APB^{+}16, AYH22, APP22, AAE^{+}22, BMS20, BKL^{+}22, BHC23, BNJ24, BD24, BSEL20, BN14, CSIS19, CKK21, CTM22b, CGBF24, DCD24, DDPVP18, DCA16, Dua21, DCA23, EKG24, FC19, FLRS12, FSXP22, FMA^{+}23, GW13, Gla11, GSL20b, UGJA18, IP16, ICF^{+}11, IZK^{+}20, KT22, KPP^{+}20b, KGK^{+}24, KTP17, KUBS22, KOJP13, KLZ16, KID^{+}15, Lee20, LYP^{+}19, LHW24, LCMV17, LCV^{+}17, LGV13, LSM^{+}12, LW15, MSC19, MKK17b, MMMH19, NMH^{+}21, PPMMS19, PBP^{+}24, PLA17, PCC^{+}18, PPP^{+}23, PST^{+}21, PSTZ23, PPP21, QMN19, RC18, RKT19, RZQS18, RH24, SOSR^{+}16, SOSC^{+}16, SMB23, Sof19, SLSK21, SKA^{+}23, SKT^{+}19, SP22, TADS20, TH20, VP22, VL24, dSVJdALRfS21, VCPPRC20, VCGOMC^{+}21, WJ21, WZS^{+}22, WLZ18, YCS17, YYY^{+}23, YPMM12, YY17]. Management-Complexity [SP22]. Managers [ATR20]. Managing [Han13, Lal14, AJT^{+}23, SSS^{+}19]. MANET [Al 20, MLK22, SHA23]. MANET/FANET [SHA23]. MANETs [Alb21, MLM14, PZH^{+}18, ZHZ^{+}19]. Manifold [KL21, ZZYC21]. Manufacturing [BD24, JNARV^{+}24, LH21b, LHSS12, MEH^{+}22, PG23, PGSL20, SU17c, TPJ21, WCSZ18, XLW^{+}24, ZLL18]. Many [CC23, Kra22, REP22]. Many-Faceted [Kra22]. Map [DHEK19]. Mapper [AJ18]. Mapping [CRMT23, GBDDVS23, JLGF19, KMS^{+}12, KVL^{+}22, LA23, LYY^{+}24, LHW18, PR11, PPDC22, UUT^{+}19, VTPT24, VDE^{+}20]. Maps [AWUF19, GMLS22, NZZ12, PFL^{+}12, SDL^{+}19]. Maria [WBG12]. Marine [Jia19, LLBB18, Tra18, ZJX18]. Maritime [KXZ^{+}24, ML20, WLWZ23]. Marked [SDL^{+}19]. Market [AAAAM^{+}21, CLC^{+}22, DZH^{+}22, DMO^{+}19, LPA^{+}22, MPT23, MATP19, OS23, RM23, SVV^{+}22]. Marketing [BKR24, GLRM19, NSJGN21, SPSS17, WLL^{+}19]. Marketplace [MBAS20, PNM^{+}20]. Markets [ASCM22, KSW21]. Markov [BOB^{+}20, GDS23, SJG18]. Markup [DL10]. Mars [BC19, WBG12]. Masculine [VYN19]. Mashup [FMCT09, HS18a, MHS23]. Mashups [BG09, Whe09]. Mask [LBK22]. Masked [GPPB^{+}21]. Masketeer [BKW^{+}24]. Masses [GNK^{+}10]. Massive [HRGQHOA21, WYS17, Wan17]. Master [ZL14]. Masterplans [FZ15, LZ11]. Match [PMGG21]. Matching [AKT24, LLW21, MEO18a, REP22, RFD23, ZL22, ZY20b]. Material [WADL^{+}24]. Materialist [LMY21]. Matérn [YPG20]. Maternal [SESO24]. Mathematical [AZAV22, KSSF19]. Matrix [BFVM14, DDD^{+}22]. Matter [Mar13]. Mature [GLD^{+}18]. Maximization [GHFM20, RZCL19, SNH^{+}19, WZLW19]. Maximize [dOB19]. Maximizing [WLWZ23]. Maximum [ZLT22]. MCCM [GHFM20]. MCSDN [RAJJ23]. Me [AF22, TRST24]. Mean [ARAAA19]. Meaning [CGGK22]. Meaning-Making [CGGK22]. Meaningful [DGADE14, GBKN23]. Means [Rók23]. Measure [LL20c]. Measurement [AA23a, CW23, DHEK19, LB23, MBF^{+}12, PXC18, YPMM12]. Measurement-Based [YPMM12]. Measurements [KKP22a]. Measures [PH20]. Measuring [AGB22, PC24, VM24a, VMAG23]. MEC [DXL^{+}22, LXL^{+}20, SDM21, WGN23, YXL^{+}21]. MEC-WPT [LXL^{+}20]. Mechanical [NZU20]. Mechanism [ADM^{+}19b, AMEF21, BKCL23, CCCPV^{+}24, CYC^{+}21, CZL21, DDMA24, DNZ22, DNZW22, KCT^{+}24, Lee17b, LWY20, LLZL19, LRL^{+}24, LMPT19, LLW21, LZY^{+}24, LHL24b, RAU^{+}24, SB22, SNZ21, SLW17, WHXL18, WCYL20, XDS^{+}19, ZLXY19, ZCHG19]. Mechanisms [AYH23, ARC^{+}19, CCC19, CCYL24, HZQ24, TFSBSPSC24]. Med [GHFL24]. Media [AKYP23, AWBS23, AAAR24, BdBCCG21, Bod22, Dav12, DV23, DIG20, FRKS22, FZ15, GPA23, HC19, JP13, Jur12, KVVD22, LPPG22, MGFF14, NWG20, PST^{+}21, SW21a, STCP21, SKA^{+}23, ST20, SG21b, SHA^{+}21, TKH^{+}11, TLV23, VKKG22, WLL^{+}19, WSN24]. Media-Related [FZ15]. Median [IKS^{+}21]. Mediated [DCA16, RRS20, TMTMB20, TA15]. Mediation [HL19]. Medical [AOI^{+}23, ALB24, ATB^{+}24, AMAA^{+}21, ACBP19, BKW^{+}24, BFL20, GHFL24, HRAA19, JUH^{+}23, JMJ^{+}23, LH21b, MTK^{+}24, PTK^{+}23, RGB^{+}23, SGW22, TSD24, WL21, ZC21, CCCA22, OMD^{+}22]. Medicine [ZSB24, AA24b]. Medium [AKJB20, Kha24, SDM22]. Medium-Sized [Kha24, SDM22]. Meet [HYMT24, Jur12, KHC^{+}24]. Meeting [HPF^{+}24]. Meets [BCG^{+}23, MW24, PFLT12]. Mega [WYL18]. Megalos [SADP21]. Mekong [PPP^{+}23]. Member [GPLK22, SHA^{+}21]. Members [THT18]. Membership [FSXP22]. Meme [RFSOHMSB21]. Memetics [RFSOHMSB21]. Memory [AKM22, BR21, BKL^{+}22, CMO^{+}21, CDLY24, DTTK19, RC18, SSD24, WZLW19, YC22]. MEMP [ZL19b]. Mental [PMT^{+}23, ZS18]. Merging [CCTV24, SYGY21]. Merkle [CCC19]. Mermaids [MP20]. MES [WCSZ18]. Mesh [CDBF19, DCL24, EW19, HF20, KSD^{+}23, KLTC17, WLCC24]. Meshing [GRM23]. Message [ASMM22, GKW^{+}10, HC19, LB24, THT18, YCS17]. Messages [CGT21, GMA20, LDM22, UDF^{+}19]. Messaging [GPPB^{+}21]. Meta [GVY^{+}23, JSJM23, JGY^{+}24, TDM^{+}24, ZJJS^{+}23]. Meta-Evaluation [GVY^{+}23]. Meta-Learning [TDM^{+}24]. Meta-Metaverse [JSJM23]. Meta-Reinforcement [JGY^{+}24]. Metadata [BGD09, Bea10, Gla11, Tas10]. Metaheuristic [MKSK23]. Metaheuristics [RAJJ23]. Metaphors [PD23]. MetaSSI [FSI^{+}24]. Metastability [SDS^{+}22]. Metaverse [BM23, FSI^{+}24, HQC^{+}24, HSTK23, JSJM23, MW24, Mot24, PM24, QHX^{+}24, RSF23, RGD24, SACG23, SGWK23, AAKS23]. Meteorological [MXT^{+}16]. Meter [EHJ17]. Metering [NBL^{+}24]. Method [Alb19, Ant24, AZV16, AHS^{+}23, BGD09, CDM23, CT20, DCHM16, DH18, DZH^{+}22, DLX^{+}19, GXWD19, GLZL24, GHL^{+}23, HCHP16, HXHP17, HKZ^{+}22, HGL^{+}23, JGY^{+}24, JJ21, KKK^{+}20, LLYL22, LL23b, LLY23, LYY^{+}24, LLZ21, LNH23, LXQ24, LHW24, PPM21, QLK^{+}24, QLL^{+}21, SOSR^{+}16, SFW23, SLB24, SSB21, SKHK23, hSgLH17, TCB^{+}19, TYH23, WLH^{+}17, WWD21, WHW22, WX24, WZLL19, WZX20, WQH23, XDH24, YYW20, YKY18, YLPW21, YWW^{+}19, YWZ^{+}24, ZZ18, ZZZ^{+}23, ZLZ^{+}24, ZY20a, ZZYC21, ZYLM17, ZQCC16, ZZZL23b, ZQL^{+}24]. Methodological [CCC^{+}24, GMRRRQ^{+}16, HSSPTM^{+}23, SACG23]. Methodologies [ADAB21, AIPV21, Rei23]. Methodology [DVED20, GNV21, NNRR^{+}21, RLB^{+}23, SOSR^{+}16, Uto13, dCdG20]. Methods [AHD21, CFX20, EMMP22, FRKS22, GM16, HH20, HLC24, HBM^{+}21, HP16, HDT22, IT24, JSPH21, KKS^{+}19, LDd^{+}24, MMD22, MKSJ22, MIK^{+}23, MC24, OC20, PR11, PAK^{+}23, PD23, RSMC21, SK24, VKKA17, VLK24]. Metric [CXLB21, LGCM24, YT18]. Metrics [ALDS19, BAK23, GRA^{+}23, HC19, JP21, MC24, MPK^{+}23, PTM20]. MeVer [PMA^{+}22]. Mexico [ASFAV23, JP21]. MFCNet [MXL^{+}21]. MHBase [MXT^{+}16]. Micro [CXG^{+}22, DNKDN^{+}24, LYY^{+}24, OGB^{+}24, YYW20]. Micro-Blog [YYW20]. Micro-Controller [DNKDN^{+}24]. Micro-Mobilities [OGB^{+}24]. Micro-Satellites [CXG^{+}22]. Micro-Segmentation [LYY^{+}24]. Microblog [HL19, YWS22]. Microblogs [EM09, JLW19, ZL21, ZXJ22]. Microcontroller [SA21]. Microcontroller-Based [SA21]. Microfinance [APAR22]. Microgrid [KAL^{+}19]. Microgrids [GPM21, TGS^{+}19]. Microservice [DSB23, KKD24, MRM^{+}24, SGLZ20, ZAS^{+}24, dVBA23]. Microservice-Based [MRM^{+}24, ZAS^{+}24]. Microservices [ATFMOU^{+}24, AALS21]. Microservices-Based [ATFMOU^{+}24]. Microsoft [HAA23, RCGSL19].

Microverse [QHX^{+}24]. Microwave [FSC^{+}23]. Mid [ACD^{+}22, SOS^{+}23]. Mid-Band [ACD^{+}22, SOS^{+}23]. Middle [ER21, OOM^{+}23, JAS21]. Middleware [AMVM22, EKG24, HVS17, PLG23]. Middlewares [FMCT09]. Migration [AGC18, KMS^{+}12, LSC^{+}17, RGMFM12, TTKZ19, WZLL19]. Military [GPCP12, GGW18, LMPT19]. Millimeter [KIB^{+}22, LSJ19, MS23, SRS^{+}24]. Millimeter-Wave [LSJ19, SRS^{+}24]. MILP [CAS^{+}20]. MIMO [CSDAM^{+}23, WYS17, Wan17]. MIMO-DQSM [CSDAM^{+}23]. Minds [RFZ22]. Minecraft [For14]. Miner [QXC^{+}21]. MinHash [HWZP18]. MinHash-Based [HWZP18]. Minimap [ZN22]. Minimax [GS23a]. Minimization [REP22]. Minimizing [IFF21, Mah17]. Minimum [BP24, NDKBL19, SWJ20]. Minimum-Cost-Based [BP24]. Mining [AAC21, CLTP22, CTM19, LMSC19, LLCL23, LY19, MSZ23, MXL^{+}21, OOM^{+}18, RZCL19, SSB21, SWGL19]. MinT [QXC^{+}21]. Minutes [HPF^{+}24]. Mirai [YFY24]. Mirror [BBGP19]. Mirrors [GPMO20, GCA^{+}22]. Misbehavior [AMF^{+}10]. Misconfiguration [AA21b, HAA23]. Misinformation [AA21a, JSPH21]. Misinformation-Related [JSPH21]. Misleading [Sta23]. Missing [AAAHA22, AAAIA24, PHGZ20]. Mist [ASL24]. Misuse [Aln22]. Mitigate [KSG24a, SS15]. Mitigating [GMGK20, SE23, SN23c]. Mitigation [AYRA21, CS11, GKM24, HK18, ITZ20, KM24, PP22a, VL18b, YZC^{+}20, GLW^{+}19]. MitM [OOM^{+}23]. MITRE [BMB^{+}23b]. Mixed [IKS^{+}21, KDJ24, SN23c, ZN22, ZCG^{+}24]. Mixed-Cell [KDJ24]. Mixed-Reality [ZN22]. Mixture [YLL21]. ML [DJC^{+}22, TAAK21]. ML-Based [DJC^{+}22].

ML-Driven [TAAK21]. MLLMs

[XZWL24]. mmWave

[BOK^{+}21, GRM23, OGB^{+}24, SHC22].

mmWave/sub [OGB^{+}24].

mmWave/sub-THz [OGB^{+}24]. Mobile [ADS20, ADS21, AAA^{+}19, ALLR16, ALST20, AS24, BTBK21, BKSS19, BCL11, BHH12, CPKK23, DYS15, DSMM21, DCA23, Far12, FWVF19, FR16, FBOC15, GSL22, GL13, Gog12, GKS10, HF22, IUIL23, JAS21, JHBW24, Jur12, KMW18, KA21, KKBG23, LPVM24, LXL^{+}20, LLLD21, LBC18, MT17b, MATP19, MDRR24, MPCS^{+}23, MPM24, NKSF24, NE21, OB13, OFV21, OBK23, PNS23, PRDK22, PSC19, PH16, RSX18, RPB^{+}17, SFEK18, SS15, SPT23, SIJA10, SNS22, TSHO24, TAHO23, TDW24, TEE10, WCM19, YA24, YT18, YY17, YPZ18]. Mobile-Assisted [NE21]. Mobile-Sensor [WCM19]. MobileNet [SC23]. Mobilities [OGB^{+}24]. Mobility [AFB23, Al 20, ADG^{+}19, BNJ24, CGT21, EAAA^{+}23, HGF21, IMR^{+}18, KC19, KSE^{+}20, LTM^{+}22, MDB22, MLM14, OGR^{+}20, Sof19, WC17, XDH24, YB24, Zar22, ZZWP19]. Mobility-Aware [CGT21]. Mobility-Enabled [ZZWP19]. Mobility-Pollution [AFB23]. Modal [CZZZ22, ZBK^{+}24, ZHH24]. Modality [FLLG24]. Modbus [CISM22]. Mode [DDZ18, OCCB20]. Model [AFS^{+}22, AML22, AAGB^{+}20, AAAIA24, ANAAM21, AAA^{+}24a, ACF24, AAA^{+}20, BdBCCG21, BdBC23, BAR^{+}21, BBGP19, BCS24, BGL^{+}22, BZNB23, CMP^{+}19, CT20, CHW^{+}21, Chu23, CTM19, CT24b, CAS^{+}20, CNC^{+}21, DDMA24, DHEK19, DL19, ESLB20, FKF24, FHXL19, dSFdRP24, GNV21, GFHK^{+}12, GMA20, GDCM19, GGK18, GRRZ18, GFL21, GAA21, HZW^{+}20, HSC^{+}22, HP20, HLZ23, HL19, HP19, IHMG22, JMZ^{+}22, JPVC21, KM24, KWI^{+}19, KYS21, KCY22, KP15, KHC^{+}23, KSO19, KSSF19, KTD24, KA23, KL21, LC23, LRD23, LDM22, LWX18, LOL23, LY19, MKKS18, MSZ23, MFO^{+}20, MA23, MM23, MSJ^{+}23, Orl22, OAH^{+}24, PPKS21, PLYD20, PPM21, PPP21, PM17, QDL22, RR24a, RRdBSS22, SAJ24, SGG^{+}22, SESO24, SC20, SS23, SSR^{+}21, SA24, SH21, SL24a, SHLH24, TRHU23, TLC15, TS22, TMS24, TTS24, TN20, VCPPRC20, WWB^{+}20, WLG^{+}22, WYW^{+}22, WSD22, WLCS17]. Model [WY19, XGN19, XLJ^{+}22, YLL21, YZCR24, YMSP24, YSAY23, ZLXY19, ZL20, ZC21, ZXWZ21, ZLZZ23, ZGZ^{+}18]. Model-Based [BZNB23, GFHK^{+}12]. Modeling [AMG22, ATE24, AES21, AMAE22, Alh22, BZMB23, DTTK19, EAKM22, FYWS16, FTHY21, GL19, GNK^{+}10, HH20, HMR^{+}21, KGWR23, LSSD22, LMSC19, MDT20, OAI^{+}22, PC23, RYAA22, RSX18, SDM22, VSE21, YOI19, YCD^{+}23]. Modelling [GWK17, MHSH16, YLYL22]. Models [ALZ^{+}23, AZAV22, AM24, AAAT21, ADS20, AAG^{+}22, BAK23, BMGI21, CFV19, CMR24, DKV24, DG13, DCB19, FFVP21, GVY^{+}23, GLH^{+}22, HD19, HYMT24, HAL18, IAM24, JT23, KK23b, KHC^{+}24, KKBG23, KKM^{+}22, LHD^{+}24, LZC21, MSC^{+}24, NNT22, OC20, OSdS24, PCVH24, RGdSK22, RR24b, RT23, RAA^{+}22, SMVP21, SSM^{+}24, SOS^{+}23, SZT23, Smi15, SM19, SB23, TOLG21, TLK^{+}20, TRLR24, VCG24, VDKL21, WLCC24, YAV^{+}21, ZLY^{+}20, ZGH13, vSR24]. Moderated [HL19]. Moderating [HCW20]. Moderation [HC19]. Modern [BCZ^{+}23, KBHP24, KKS^{+}19, Sta23]. Modification [Aln22]. Modified [HJA18, WLL20]. MODIS [PPP^{+}23]. Modulation [AM21, DNPC^{+}22, JKM^{+}23, MLWS24, WLYL20]. Module [LRD23, MKPG22, SBN^{+}23, WZL21]. Modules [Oza23, SMR24]. Moment [BOB^{+}20]. Monetization [SVGGO24]. Money [ADS20, ADS21, CMS21, JAS21]. MongoDB [CFV19]. Monitoring [Adi23, ANAAM21, AEM^{+}23, ADK^{+}24, AASAI22, AYH^{+}21, ABLR23, BMB23c, BZP^{+}21, CVG^{+}22, CC22a, DHDAu22, DDV22, DVSBV^{+}24, FLSL24, GRA^{+}23, GAA21, JCPV22, KS24, KGWR23, KTD24, LWY^{+}15, LS21, LLBB18, LA20, MNP^{+}24, MDRR24, NMT23, OAG21, PC24, PBP^{+}24, PMT^{+}23, PORM^{+}23, PHGZ20, RS17, RIS^{+}17, RHV17, SESO24, SSO^{+}19, TDM^{+}24, TVAT24, VGS^{+}19, WL20, XSX18, ZJ18, dCMA21]. Monoaural [ALH24]. Monocular [GZS^{+}24, TCH18]. Monopole [CLZL18, PBAAN19]. MONs [YPLZ19]. Monte [PPM21]. MOOCs [CWC^{+}22, OLMBRCRC23]. Mood [WLH^{+}17]. Morals [GCA^{+}22]. Moroccan [FOJE19]. Mortality [SESO24]. Most [MCCOCE24]. Most-Valued [MCCOCE24]. Motion [LMH19, OOM^{+}18, ZWPL19]. Motion-Compensated [LMH19]. Motivation [LL20b]. Motives [HS18b]. Motor [BPLBM21]. Mountain [KNN^{+}24]. Mounted [BRC24]. Move [KC19]. Movement [ASF^{+}23, BTBK21, GAA21, JTA^{+}21, MZH22]. Movements [BCMPW21]. Movie [CH23, ZXJ22]. Moving [CMP^{+}19, ZSB24]. MPCC [PKS^{+}24]. mPOC [Adi23]. MPTCP [LIM23]. MQTT [DTR22, LF24]. MQTT-Based [DTR22]. Mscf [GR22a]. Mscf-ResNet [GR22a]. MSEN [YCD^{+}23]. MU [YLJ^{+}19]. MUI [ZLOW24]. Multi [ABK22, AKM^{+}23, ANHMS24, AlQ23, ADS21, ABH24, AAE^{+}22, ALH24, BOK^{+}21, Cap23, CSDAM^{+}23, CKK21, CHW^{+}21, CHMZ21, CLTP22, CLH^{+}23, CSZ^{+}24, CGT21, DWHY24, DDPVP18, DDZ18, DIE^{+}21, FSY^{+}22, FRE^{+}22, FLM23, GMLS22, GPL24, GB23, HF20, HF22, ITH^{+}21, JUN^{+}24, JT23, JHBW24, JK24, Kab23, KK21, KBHP24, KK20, Kra22, LWY20, LZLW22, LQY^{+}19, LXL^{+}19, LF23, LZY^{+}24, LLSS24, LFM^{+}19, MCA23, MP16, MWY19, MKPG22, MLL^{+}22, OPGH23, PLYD20, PLA^{+}24, RAJJ23, RZCL19, oRL19, SV24, SSD24, SRSDF23, SVFV23, SOR16, SYLL21, SAM^{+}18, TBT^{+}23, TTKZ19, TWW24, TRST24, VGK21, WL20, WX20, WJ21, WXL22, WLLY24, WLCC24, WD24, WZX20, WZL21, XCSM18, XY21, XLM^{+}23, XSSA24, YWS22, YBO^{+}23, YCD^{+}23, Zar22, ZBK^{+}24, ZZ18, ZLXY19, ZL19a, ZZWP19, ZLZ^{+}24, ZWL^{+}24, ZHH24, ZLZZ23, ZLOW24, ZQCC16]. Multi-Access [LZY^{+}24, TTKZ19, TRST24, ZWL^{+}24]. Multi-Agent [DWHY24, GB23, JUN^{+}24, JHBW24, KK21, KBHP24, LZLW22, LFM^{+}19, SV24, WLLY24, Zar22]. Multi-Agent-Based [AAE^{+}22]. Multi-Angle [ITH^{+}21]. Multi-Antenna [CLH^{+}23]. Multi-Attention [ZLXY19, ZL19a]. Multi-Attribute [MLL^{+}22]. Multi-Authority [WZX20]. Multi-Blockchain [KK20]. Multi-Case [Kra22]. Multi-Cell [YBO^{+}23]. Multi-Channel [HF20]. Multi-Class [TWW24]. Multi-Constraint [ZLZ^{+}24]. Multi-Context [ZZ18]. Multi-Controller [RAJJ23]. Multi-Criteria [ANHMS24, DDPVP18]. Multi-Database [JT23]. Multi-Decision [CHMZ21]. Multi-Domain [SAM^{+}18]. Multi-Element [WX20]. Multi-Energy [MCA23]. Multi-Equation [FLM23]. Multi-Facial [LQY^{+}19]. Multi-Factor [ADS21, OPGH23]. Multi-Floor [JK24]. Multi-Focus [ZQCC16]. Multi-Grained [WJ21]. Multi-Group [WLCC24]. Multi-Head [PLA^{+}24, YWS22]. Multi-Hop [CGT21, LXL^{+}19]. Multi-Input [oRL19]. Multi-Key [FLM23]. Multi-Layer [AKM^{+}23, BOK^{+}21, SSD24, WXL22]. Multi-Level [XY21, XSSA24, ZLZZ23]. Multi-Modal [ZBK^{+}24, ZHH24]. Multi-Mode [DDZ18]. Multi-Model [CHW^{+}21]. Multi-Object [DIE^{+}21]. Multi-Objective [XLM^{+}23]. Multi-Output [oRL19]. Multi-Policy [ZLZ^{+}24]. Multi-Premise [WZL21]. Multi-Provider [MP16]. Multi-Radio [HF20]. Multi-Rate [LWY20]. Multi-Scale [LF23, VGK21, XY21, YCD^{+}23]. Multi-Sensory [FRE^{+}22]. Multi-Service [ABK22]. Multi-Site [GPL24]. Multi-Slice [AlQ23]. Multi-Social [RZCL19]. Multi-Source [ABH24, CSZ^{+}24, FSY^{+}22, PLYD20, SYLL21]. Multi-Spectral [GMLS22]. Multi-Stage [TBT^{+}23]. Multi-Targets [WL20]. Multi-Task [ABH24]. Multi-Thresholds [XCSM18]. Multi-Tier [CKK21, Kab23, SRSDF23]. Multi-Token [MCA23]. Multi-Topology [MWY19]. Multi-Unmanned [LLSS24]. Multi-User [ANHMS24, CSDAM^{+}23, LXL^{+}19, oRL19, SVFV23, SOR16, WD24, ZZWP19, ZLOW24]. Multi-View [ALH24, CLTP22, HF22]. Multi-WiIR [WD24]. Multiagent [TCH23, WLWZ23]. Multiband [NAV^{+}23]. Multicast [CFP17, CFP19, DLK10, DNZ22, DNZW22, JWRX17, KKBD10, PSM^{+}18, TYH23, YC16]. Multicell [KTP17]. Multiclass [SBRZ24]. Multidimensional [OFV21]. Multidisciplinary [RSF23]. Multidiscipline [ZLL18]. Multifaceted [EWGK24, PNS23]. Multifractal [WKD22]. Multihoming [YB24]. Multihop [NSV17]. Multilane [SYGY21]. Multilayer [FPKK22]. Multilevel [HT14]. Multilingualism [Lau15]. Multimedia [BCG10, HDL^{+}14, HE20, RRT20, SCB21, Wan23, WY19]. Multimodal [AFS^{+}22, CSA^{+}24, CCM^{+}17, FLLG24, FM20a, RGP22, VCG24]. Multimodel [RAA^{+}22]. Multiparty [AKK^{+}24]. Multipath [KSSF19, KLTC17, LHW24, PKB^{+}24, PW22, PKS^{+}24, WY19, XNZ23a, XNZ23b]. Multiperspective [MHS23]. Multiple [BZNB23, GLH^{+}22, HLZ^{+}18, HWZP18, LXL^{+}20, LLY23, LWZL22, MPM24, RIS^{+}17, SWZ18, WYL^{+}21, YZP22, YAV^{+}21, YXL^{+}21, ZZ21]. Multiplex [SAL21]. Multiplexing [NAV^{+}23]. Multiplication [AKDL24]. Multiproduct [RH24]. Multiscalar [dCMA21]. Multitask [FLLG24, SPC22]. Multitasking [SW21a]. Multitier [DGKL15]. Multiverse [FAS^{+}23]. Municipal [FZ15, ZL14]. Municipality [Yan17]. Musculoskeletal [JTA^{+}21]. Museum [FDN^{+}24]. Musical [KPP^{+}20b, KYT^{+}23]. Mutual [AdMTJ20, CLM^{+}21]. My [MSR^{+}14, UDF^{+}19]. myDIG [KS19]. MySQL [SO24].

N [YLY^{+}22]. N-Trans [YLY^{+}22]. Naïve [HKZ^{+}22]. Name [AAAAK24, Bea10, FP19, GSL20a, KHM20, LYD22, LL23b]. Name-Based [FP19]. Named [AAAAK24, ATH18, ARC^{+}19, BNJ24, BKCL23, GSL22, GXWD19, ISLARP^{+}22, KST19, KH19, LMPT19, MSL^{+}23, SB22, YSEK20]. Names [YLY^{+}22]. Naming [WC20]. NAND [BZMB23]. Nano [MAEP18]. Nano-Things [MAEP18]. Nanoscopic [MDB22]. Narrative [ATY20]. Narratives [DGGS19]. Narrowband [HDBD23, NBL^{+}24]. Nation [Cal24]. Nation-Statehood [Cal24]. National [GVY^{+}23, RL24, TBT^{+}23]. Nationality [Mar13]. Native [Dua21, DGCK24, GVV^{+}23, Kra22]. Natives [BKR24, GWK17]. Natural [CS22, DGHH^{+}20, EMMP22, HOV23, ICF^{+}11, NPB^{+}19, PD23, STV^{+}22, SN23c, VCJAMN22, ZCG^{+}24, ZSS^{+}22]. Navigation [AIP^{+}24, BBBHA24, FSY^{+}22, HS20, IAG^{+}21, Ngu19, ZN22, ZYLM17]. NB [ADAB21, KZRH22, MOAI17]. NB-IoT [ADAB21, KZRH22, MOAI17]. NDN [BKCL23, KSD^{+}23, KHM20, PGFH24, WZX20, YHG^{+}21]. NDN-BDA [BKCL23]. ndnIoT [GSL20a]. ndnIoT-FC [GSL20a]. NDNs [HKZ^{+}22]. Near [BZP^{+}21, CC23]. Near-Sensor [BZP^{+}21]. Neck [IKS^{+}21]. Need [LNG19, SSS21]. Needs [PEN23, Tag20]. Negative [FC19, KPC19]. Negotiation [WHXL18]. Neighbor [DH22]. Neighborhood [GKFV24]. Neighbour [BP24]. Neither [FKK20]. Neo [CCA^{+}20, MDS^{+}24]. NeoGeography [For14]. Nephropathy [ZWS20]. Nerve [SWD^{+}18]. Nested [PKB^{+}24]. Net [GRA21, SS23, ZMDCEMI22]. NetKAT [PG19]. Nets [DCHM16]. Network [ALZ^{+}23, AP22, ASCM22, AGO23a, ATH18, AA21b, Aln22, AA21c, ABB23, AAE^{+}22, AKM22, ALH24, BMB^{+}23a, BKL^{+}22, BTBK21, BZP^{+}21, BZDP19, BKSS19, Bi14, BBH^{+}24, BRP^{+}13, BC23, CZZZ22, CCCPV^{+}24, CPRS20, CYC^{+}21, CTNS21, CH23, CASN21, CGT21, CMH10, DH22, DJC^{+}22, DWHY24, DWZN16, DDMA24, DLX^{+}19, DAMAS^{+}19, EW19, ELC^{+}19, ELCS20, EVCL21, FC19, FR24, FHZ^{+}24, FC22, FWVF19, FML^{+}22, GS23a, GRM23, GVV^{+}23, GHGL22, GRG21, GVC13, Han13, HLZ^{+}18, UGJA18, HCHP16, HD16, HWCL17, HY22, HGL^{+}23, IG10, ITZ20, JK20, JTA^{+}21, KGKP20, KES24, KGRP21, KKK^{+}20, KWI^{+}19, KK21, KDE24, KZ24, KPC19, KKBG23, KLG^{+}24, KSO19, KBF23, KLTC17, KTD24, KKBD10, KRSF22, LLW^{+}17, LLZL19, LLLD21, LZY22, LTW21, LF16, LQY^{+}19, LXL^{+}19, LLL21, LNH23, LZY^{+}24, MMA10, MAVS20, MSN13, MP16, MS16, MGBG21, MGV17, MXL^{+}21]. Network [Mou24, MA23, MDRR24, MBF^{+}12, MLL^{+}22, NB24, NZZ12, NLT^{+}23, NDS^{+}23, NMT23, OAG21, OAH^{+}24, PMA^{+}22, Pap20, PPMMS19, PXC18, PLG23, RS17, RF21, RD18, SBN^{+}23, SMG13, SE23, SHC22, SAL21, SADP21, SSM^{+}24, SQK^{+}17, SAR24, SN23a, SYGY21, SSS^{+}19, SMN22, SST23, Smi13, SXXH19, SYZ19, SK23, SKT^{+}19, SL24a, SG21a, SP22, TSHO24, TCH23, TLK^{+}20, UFK^{+}10, VGK21, VL20, Vem22, WLH^{+}17, WZLW19, WSM20, WWD21, WX22, WXL^{+}23, WYZ24, WZ18, WYL18, WQH23, YPXH19, YZZ^{+}20, YYY^{+}23, YT18, YSEK20, YLPW21, YCD^{+}23, YY17, YLJ^{+}19, ZBN21, ZAS^{+}24, ZL19b, ZL19a, ZCHG19, ZL20, ZL22, ZZZ^{+}23, Zha19, ZSRC23, ZHW23, ZJJS^{+}23, ZSF20, PPN18]. Network-Based [BKL^{+}22, BZP^{+}21, DWHY24, DDMA24, Han13, NMT23, SBN^{+}23, SSM^{+}24, Zha19]. Networked [EUDW23, FHL^{+}24, LS21, Tag20, TEE10]. Networking [ATFMOU^{+}24, AAAAK24, AP21, ACM^{+}19, ARC^{+}19, BNJ24, BKCL23, BCFP21, BM14, CPKK23, CDBF19, DNZ22, DNZW22, FP19, Fot20, Gal24, GJ15, GPCP12, GTQ^{+}22, GVV^{+}23, HYMT24, ISLARP^{+}22, IZK^{+}20, JNARV^{+}24, KST19, KST24, KH19, LHL24b, Mac22, MZS19, MTKS24, MCH^{+}20, NAFM^{+}24, Pap20, PKCF21, QSF^{+}17, RAJJ23, SYB^{+}18, SUF^{+}22, SN23a, Smi13, SM19, SNZ21, STS23, VL20, XNZ23a, XNZ23b, ZWZ19]. Networks [AAR^{+}19, AAM^{+}24, ADM^{+}19b, AZH22, AMG22, AAAHA22, ALB24, AK22, AAA^{+}24a, AIR^{+}18, ACD^{+}22, AASAI22, ASL24, ABH24, AA21c, AMZ^{+}20, AMSM23, APP22, AHH^{+}23, ABB23, AMF^{+}10, AAE^{+}22, AS24, BFY22, BCNS20, BdCRM10, BP24, BF18, BCG10, BL22, BKL^{+}19, BRP^{+}13, CBM^{+}20, CLM^{+}21, CGMM22, CCD^{+}21, CSL^{+}20, CBKL22, CZK^{+}17, CXZ23, CLH^{+}23, CCTC23, CS22, CKSG21, CAL23, CCYL24, DPC^{+}24, DWHY24, DPH17, DPH19, DVSBV^{+}24, DMZ^{+}19, DLK10, Dua21, EVCL21, FMA^{+}20, FWVF19, FMMS21, FG17, FML^{+}22, GSL20a, GSL22, GWF^{+}23, GNV21, GMGF24, GDARM18, GGW18, GAMMPCRÁ24, GX20, GTC19, Gro17, GMB^{+}21, GSP15, GFL21, GR22b, HF20, HLZ23, HSSPTM^{+}23, HLLT10, HNK^{+}21, IGL^{+}23, IKP^{+}13, IUIL23, IZK^{+}23, IDR23, IT24, JBS^{+}23, JUN^{+}24, JCPV22, JP20, Kab23, KSD^{+}23, KMW18, KSG24a, KLKP19, KTP17, KRSK24]. Networks [KHN^{+}11, KPKM23, KC19, KTUF19, LBG^{+}14, LZ23, LPVM24, LDd^{+}24, LMK18, LSH21, LWW22, LRL^{+}24, LBB23, LCW^{+}18, LH21c, LZL^{+}21, LGY23, LLW^{+}24, LARPHA23, LCMV17, LRdLS19, LZC21, LN19, MCM^{+}10, MG10, MEO18b, MP20, MP16, MKDG20, MT20, MKK17b, MWY19, MSX^{+}21, MKvD11, MBLC22, MZH22, MSC^{+}21, MN21, MSAD22, MPCS^{+}23, MPM24, MTHN22, MZLP22, MHL^{+}23, MHQ^{+}23, NAV^{+}23, NB24, NPMPM23, OOM^{+}23, OB13, OMD^{+}22, OLCMdA20, OAI^{+}22, OTR^{+}20, PNS23, PZC21, Pet20, Pin10, PSM^{+}18, QLL^{+}21, QAM17, RLB^{+}23, RZCL19, oRL19, RSX18, Rók23, RKS^{+}21, REP22, SB22, SSF^{+}18, SJA^{+}23, SDDB24, SRMJ21, SCC^{+}23, SS23, SSDS23, SOS^{+}23, SSD24, SSR^{+}21, SOK24, SSS^{+}19, SS15, SIJA10, SWD^{+}18, SXXW20, SNS22, SAZ18, SLW17, SAM^{+}18, TAMA22, TAHO23, TDW24, TFKY19, TL17, TFSBSPSC24, THV^{+}22, TSDG22, TWW24, USS19, VP22, VMGT22]. Networks [VR21, VLK24, WCF^{+}16, WCM19, WZC19, WWB^{+}20, WZS^{+}22, WLLY24, WZG^{+}18, WLWZ23, XSX18, XNC21, YCHG18, YKC19, YPG20, YGQ^{+}22, YHSW19, YMH22, YW16, YST24, YLH^{+}17, ZOAR24, dOB19, ZX24]. NetworkX [PMA^{+}22]. Neural [ASCM22, ATH18, AKM22, BFY22, BZP^{+}21, BC23, CBM^{+}20, CKSG21, DDMA24, DAMAS^{+}19, EVCL21, GX20, GTC19, GR22b, Han13, HSSPTM^{+}23, HY22, HGL^{+}23, JTA^{+}21, KWI^{+}19, KRSF22, LDd^{+}24, LLZL19, LWZ^{+}21, LGY23, LRdLS19, MGBG21, MBLC22, MTHN22, MHL^{+}23, NB24, NDS^{+}23, OMD^{+}22, OTR^{+}20, PLYD20, PC23, RSX18, RKS^{+}21, SSM^{+}24, SSR^{+}21, SWD^{+}18, TSHO24, VGK21, Vem22, WLG^{+}22, WZ18, YZZ^{+}20, YGQ^{+}22, YMH22, YST24, ZLY^{+}20, ZJJS^{+}23, dOB19]. Neural-Network-Based [TSHO24]. Neuro [PBP^{+}24]. Neuro-Monitoring [PBP^{+}24]. Neurological [JTA^{+}21]. Neurologist [SWD^{+}18]. Neutral [CBD^{+}22]. Neutrality [Cap12]. Newly [YKC19]. News [AdMTJ20, ALL24, AAARZ24, APJ24, AGN21, CLM^{+}21, DN20, HP21, KLKA19, LZC21, MV20, NJH^{+}24, NSD^{+}22, Øie12, PCVH24, PFdRGG21, RS22, SSS21, SDM22, TZS^{+}21, TZE^{+}21, VMGT22]. News-Based [NJH^{+}24]. Newspaper [PFdRGG21]. Next [AlQ23, AEA^{+}23, APP22, AMI23, AA24b, CC24, FWVF19, HYMT24, HEP^{+}11, HLLT10, IG10, KP19, MKSV^{+}19, RYY10, SAZ18, Tre21, WZG^{+}18]. Next-Generation [AlQ23, AEA^{+}23, APP22, AMI23, FWVF19, HYMT24, IG10, RYY10, WZG^{+}18]. NextDet [KES22]. NeXtFusion [KES24]. NextGen [DGCK24, SPT23]. NFC [ESBE23]. NFT [GM23, KFK^{+}24]. NFV [Pap20, SDM21, SP22, TH20]. NFV-Based [SDM21]. NFV-Enabled [TH20]. NG [Rók23]. NG-PON2 [Rók23]. NGN [MKvD11]. NGOs [MCCOCE24]. NHS [TBT^{+}23]. NIDS [NLT^{+}23]. NLP [AAK24]. NLP-Based [AAK24]. NN [KKEV17]. No [DHEK19, KS19, SLP^{+}22, WL22]. No-Reference [DHEK19]. No-Show [SLP^{+}22]. Nodal [BTBK21, CASN21]. Node [AASAI22, BP24, CCCPV^{+}24, LCW^{+}18, LNH23, MMA10, MCM^{+}10, TYH23, VLC^{+}19, WLZ18, WY19, HSL^{+}23]. Node.js [KKD24]. Node.js-Based [KKD24]. Nodes [AASAI22, DYS15, EUDW23, GZC21, IGL^{+}23, MMS22, OT16, RZCL19, SAL21, SA21, SFEK18]. Noise [LMLW18, TGELGH13]. Noisy [CR23b, FPGO24, SGW22, ZXXB22]. NOMA [ASL17, GWF^{+}23, THV^{+}22, WSM20]. NOMA-Based [WSM20]. NOMA-Enabled [GWF^{+}23]. Non [AKDL24, Aln22, CDLY24, DWY19b, EKKS24, FOJE19, GM23, KDD^{+}21, LCC^{+}24, LLZ21, TH21, TH20, YZP22]. Non-Colluding [TH21]. Non-Control [Aln22]. Non-Directional [LLZ21]. Non-Fungible [GM23]. Non-IID [EKKS24, LCC^{+}24]. Non-Immersive [FOJE19]. Non-Informational [KDD^{+}21]. Non-Local [DWY19b]. Non-Orthogonal [YZP22]. Non-Profiled [AKDL24]. Non-Renewable [TH20]. Non-Uniform [CDLY24]. Noninformative [SLZY21]. Nonlinear [ZLS19]. nonlinearity [GLW^{+}19]. Nonparametric [BZMB23]. nor [FKK20]. Norm [HD16]. Normalization [PHG^{+}20]. Norms [VYN19]. North [VLMP21]. North-West [VLMP21]. Norway [NWG20]. Norwegian [VNJ18]. NoSQL [DCB19]. Nostalgia [HP20]. Notification [AHS^{+}23]. Notifications [LBC18, SLB24]. Notion [AAD^{+}12]. Novel [AN23, ASCM22, AHKC23, ARAAA19, ADAq^{+}22, ADK^{+}24, BMJ^{+}22, BHS24, DPR^{+}21, DGCK24, EKG24, HSL^{+}23, IHMG22, KPT^{+}24, KHC^{+}24, KFK^{+}24, LWHR14, LSC^{+}17, LXY17, LXQ24, MAVS20, MKKS18, NAFM^{+}24, OLCMdA20, PMP22a, PZC21, QWR18, RDASB22, Rók23, SOK24, SKHK23, TTKZ19, WZL^{+}20, WSM20, XCSM18, XNZ23b, YC16, YLYL22, ZQCC16]. Novelty [dSMdO22]. NR [AMSM23, BOK^{+}21]. ns [AVA^{+}22]. ns-3 [AVA^{+}22]. Nudge [LGMZ19]. Number [MSX^{+}21, Tsi14, VMD^{+}22, YLYL22]. Numbers [VCJAMN22].

O [JLH^{+}13]. Obfuscated [AKAS22, ZJJS^{+}23]. Object [BFY22, BCM^{+}19b, Chu23, DIE^{+}21, DMP^{+}23, KES22, KES24, KHP^{+}22, LZY22, MPL^{+}21, RES23, SPB^{+}11, SC23, Sor12, VGK21, ZCHG19]. Object-of-Interest [Sor12]. Objective [AWSA23, XLM^{+}23]. Objectives [AP22, ANHMS24, LS15, Tag20]. Objects [DYLZ21, GNNZRCRG18, KVL^{+}22, MAW^{+}19, MC12, Pir22, RSL^{+}24, DG18]. OBS [BdCRM10]. Observability [Kra22]. Observation [CH16, SOSK12]. Observing [Yan17]. Obstacle [GM20]. OCARI [MMA10]. Occlusion [TCH18]. Occlusion-Aware [TCH18]. Occupancy [EUDW23, HBM^{+}21, HZCZ22, TCB^{+}19]. Oceania [SRSDF23]. Oceans [Jia19, ML20]. ODK [APB^{+}16]. Odometry [NLD^{+}24]. OFDM [GLW^{+}19]. Off [HKZ^{+}22, LMK18, SLW17]. Offering [FGBMG19]. Offerings [IPL^{+}18]. Office [FM20a, ZXJ22]. Offices [PPM21, RBSMRN21]. Official [CKK21]. Offline [MFO^{+}20, RGM^{+}21, ZLLW18]. Offload [SPT23]. Offloading [ADM^{+}19b, DH22, DVED20, GWF^{+}23, JHBW24, LLL21, LZY^{+}24, MATP19, QWR18, TCYZ22, YXL^{+}21, ZWL^{+}24]. Offs [GL13, TMS24, AHMI22]. Offset [Lee17a]. Oil [KTD24, OOM^{+}23, Sta23]. Old [Roy14]. Older [ADCM24, LNB12]. OLP [dCdPC^{+}21]. OLSR [CTNS21]. On-Demand [AYH^{+}21, DTTK19]. On-Device [AN23]. On-Manifold [KL21]. On-Off [LMK18]. On-Ramp [SYGY21]. On-the-Fly [PMGG21]. Onboard [ML20]. One [DNKDN^{+}24, JK24, MZH22, REP22, SU17b, SZZL23, YPLZ19]. One- [MZH22]. One-Pass [DNKDN^{+}24]. One-Time-Use [SZZL23]. Ongoing [AS19]. Online [ABSG21, AIR^{+}18, AGN21, BLW^{+}17, BBK^{+}20, CLM^{+}21, CS11, DWY^{+}19a, DCPG21, Gar12, GK21, GTWN24, HZWJ21, HLZ23, HAL^{+}22, HCW20, HRGQHOA21, ISG20, ICF^{+}11, ITZ20, LS15, LLJ22, LLZ21, LOL21, LY19, MBAS20, MKDG20, MKY24, MSX^{+}21, MSC^{+}21, PDI21, PFdRGG21, PC23, PVKG23, RGD24, SWP^{+}22, SPB^{+}11, Sor12, SSS21, SS22, SG21b, SHA^{+}21, VKKA17, WSN24, WT13, Yan19, YQ24, ZS18, vdSFRF23]. Ontological [KPCS13]. Ontological-Relational [KPCS13]. Ontologies [AAMD21, BYS^{+}15, CCTV24, Lau15]. Ontology [ASK23, AML22, AMR^{+}21, CP10, FM21, GSO^{+}10, LZ11, LSSD22, LLJ16, MGUD12, PCROB21, RFD23, RSL^{+}24, SU17c, SOB^{+}19, STK21, SKA^{+}23, SZC19, ZL14]. Ontology-Based [CP10, LLJ16, PCROB21, RSL^{+}24, SU17c, SOB^{+}19, STK21]. Ontology-Driven [AMR^{+}21]. Open [AP22, ASWAE23, AYH23, BBK^{+}20, CTP^{+}22, Ciu19, CGMVdUC19, CSC21, DTR22, DWL21, GMD^{+}22, GMGF24, GZC21, GGT^{+}23, HPC12, HPF^{+}24, HRGQHOA21, KS14, KBF^{+}22, Li18, LXQ24, LIM23, MZS19, MHQ^{+}23, NB12, PLOT23, PKCF21, PFL^{+}12, PR20, QA24, RT23, Roy14, SGWK23, Sca14, SLSK21, TRB22, VCB14, VDE^{+}20, YWZ^{+}24, ZZZL23a, ZZZL23b, dCdPC^{+}21, vSR24, OLMBRCRC23]. Open- [vSR24]. Open-AI [RT23]. Open-Domain [ZZZL23a, ZZZL23b]. Open-House [SGWK23]. Open-Source [CTP^{+}22, DTR22, HPF^{+}24, PLOT23, TRB22, YWZ^{+}24]. OpenFlow [BM14, IZK^{+}20]. OpenGov [SS14]. OpenID [AYH22]. Opening [BN14]. OpenStreetMap [MC12, NZZ12, NZ14, PM17, SSR17]. Operating [BR23, KBBH21]. Operation [BCG^{+}23, FAT19, FHS^{+}10, KSE^{+}20, SBS18, TGS^{+}19]. Operational [ACD^{+}22, AHMI22, LCMV17]. Operations [CFV19, GL19, KP15]. Operative [GAA21]. Operators [FFP^{+}22]. Opinion [MGFF14, Mot24, YKC19]. Opportunistic [AS24, KZ24, KTUF19, Lee17b, LCW^{+}18, MZH22, TL19, UDF^{+}19, YY17]. Opportunities [AMAJ12, ASD^{+}22, Bad23, BTS^{+}21, BCG^{+}23, BOKC19, CWF13, GS23b, HQC^{+}24, KG18, LMF23, MRRMC21, PEN23, RAC24, Rot12, Roy14, TA15, VL24, ZR15]. Opportunity [SL24a]. Optic [CCD^{+}21]. Optical [AAR^{+}19, AD20, DDZ18, GPMO20, Gro17, KLKP19, LLBB18, NAV^{+}23, Pin10, TCH18, GLW^{+}19]. Optical-Acoustic [LLBB18]. Optimal [ANHMS24, AlQ23, BAR^{+}21, FN23, GRM23, HF22, JUH^{+}23, LYP^{+}19, SXXH19, SWZ18, YC16, ZY15]. Optimally [Han13]. Optimisation [Gro17, LTP24]. Optimising [ASL24]. Optimization [AIP^{+}24, AK22, ASBC22, ADAq^{+}22, AGC18, APAR22, CD23, CBR^{+}24, DYS15, DDV22, DLX^{+}19, EKKS24, GW15, LTY18, LKIL19, LZS18, LW19, LW24, LLW^{+}24, LLSS24, LGCM24, MCS^{+}24, MKSK23, MDM21, MKY24, MAM^{+}21, MKR22, OAI^{+}22, PGFH24, PAK^{+}23, Rók23, SYB^{+}18, SGG^{+}22, SCA^{+}19, SOK24, SKT^{+}19, SWGL19, SSX20, TSHO24, TGS^{+}19, TTN24, TGELGH13, Tra18, TTK24, VG20, WX22, YFY24, YPXH19, YQ24, YL21, YWW^{+}19, ZBK^{+}24, ZWL^{+}24, ZYWS24, ZWZ19]. Optimizations [GDN^{+}23]. Optimize [LXL^{+}20, MCPLBH^{+}24, WX24]. Optimized [AAM^{+}24, ATB^{+}24, BCM^{+}19b, FSC^{+}23, OMD^{+}22, PVM22, QNJ24, RAJJ23, RAU^{+}24, SGMBK24, TRHU23]. Optimizer [TAMA22]. Optimizing [AKT24, CDLY24, FHXL19, IUIL23, IUK^{+}23, KAEAB24, LWH24, QSML24, SAR24]. Optimum [AAR^{+}19, BAS24]. Option [QA24]. Options [EAKM22, KBBH21, MS23]. Oracle [Cal22]. Orange [LBK22]. Orbit [CXG^{+}22, Mat20]. Orchestration [ARC^{+}19, LHSS12, MVMHL24, OS23]. Order [CCKH21, GDS23, REP22, VBWW22]. Ordinary [PC23]. Organisation [Tag20]. Organization [MFH22, SDM22]. Organizational [YOI19]. Organizations [GL21, KPP^{+}20b, MCCOCE24]. Organize [SU17a]. Organized [PTM20]. Organizing [CPP^{+}20, SHA23]. Oriented [AWBS23, CDBF19, DPR^{+}21, KPP^{+}20a, LW11, LZL^{+}21, LLL21, MP16, PKS^{+}24, QHX^{+}24, SHLH24, TKH^{+}11, ZLOW24].

Origin [GPLK22]. Originality [LS23]. Orthogonal [YZP22]. OSPF [SAR24]. OT [FMMS21]. Other [Ler16, LS23, VM24b]. Outage [CASN21]. Outbreak [Adi23]. Outbreaks [SG21b]. Outcome [FC19, FFVP21, KPC19, PLA^{+}21, PKCF21]. Outcome-Experience [PLA^{+}21]. Outdoor [ACD^{+}22, FSY^{+}22]. Outdoor-to-Indoor [ACD^{+}22]. Outdoors [WL22]. Outlets [LPPG22]. Outlier [ASF^{+}23, MKR22]. Output [BAK23, oRL19, Smi15, Uto13]. Outs [PG23]. Outsourcing [SZW^{+}23]. Over-The-Air [VLC^{+}19]. Overdue [APAR22]. Overflow [LL23a, LLCL23]. Overhead [KKD24]. Overlapping [PZC21]. Overlay [GMGF24, GKS10, LWW22, LZY^{+}24]. Overlooked [Tre21]. Overtaking [PG23]. Overview [AP19, BGW16, Cal22, DPC^{+}24, KXZ^{+}24, KPC^{+}23, Ray23, SM19, SEB^{+}19, YMJ20]. OWA [AAAIA24]. OWA-Based [AAAIA24]. OWL [LSSD22]. Owner [GANT21]. Owner-Custodianship [GANT21]. Owners [ZSF20].

P4 [MMMH19, MMMH19]. P4-ABC [MMMH19]. PACER [KAS^{+}20]. Packet [Bi14, BKL^{+}19, CLH^{+}23, KDE24, LSZ21, PP22b, SHLH24, YK21]. Page [MMD22, MGBG21, MDM21]. Pages [DH18]. Pair [WLCC24]. Pairing [WHW22]. Pairings [REP22]. Pakistan [APB^{+}16]. Pallor [RHV17]. PANAS [GRB^{+}21]. Pandemic [CBRS21, Del21, GNGHEM21, GFPD21, PLA^{+}21, TC21, dCMA21]. Paper [BTS^{+}21]. Papers [Bod22]. Paradigm [ABBV23, KS24, PPMMS19, Rot12]. Parallel [CMO^{+}21, EPC24, HM20, IAG^{+}21, MT23, MTF^{+}21, YLY^{+}22]. Parallelism [CDLY24]. Parallelized [KHAA21]. Paralysis [SWD^{+}18]. Parameter [KKP22a, PGFH24, YFY24]. Parameters [DYS15, MDB22, MSJ^{+}23, YQ24]. Paranoid [ATY20]. Parcel [SH20]. Park [FC19]. Parking [KZRH22, KK20, Str22, TCB^{+}19, WC23]. Parkinson [ZC21]. Parliament [DBD^{+}14]. Parliaments [KMV^{+}23]. Part [BGL^{+}22, WQH23]. Part-of [BGL^{+}22]. Participant [XLM^{+}23]. Participating [Mot24]. Participation [DCA16, Gar12, GSD^{+}17, Smi13, Yan19, YPZ18, ZS18, Zha19]. Participatory [Hua15]. Particle [ADAq^{+}22, FSY^{+}22, OAI^{+}22, QLK^{+}24, Tra18]. Partition [CXG^{+}22, GOM23, WLCC24]. Partition-Based [GOM23]. Partitioning [ADM^{+}19b, BFG^{+}23, KMK24, dOB19]. Partnership [ASM^{+}24]. Pass [DNKDN^{+}24]. Passenger [SWZ18, VGS^{+}19, XZY^{+}24]. Passive [HMR^{+}21, LDd^{+}24, YLH^{+}17]. Password [ATR20]. Past [RKM^{+}22]. Pat [OTR^{+}20]. Pat-in-the-Loop [OTR^{+}20]. Patent [UJ22]. Path [CRMT23, EARP^{+}24, GM20, LSH21, MAM^{+}21, SOS^{+}23, TYH23, XNZ23a, ZLZ^{+}24, ZYLM17, ZXMB22]. Pathloss [OAI^{+}22]. Pathologies [ABLR23]. Patient [GAA21, HBMS20, MKM^{+}19, NB12, PTK^{+}23, Tri14, SLP^{+}22]. Patient-Centric [Tri14]. Patients [ANAAM21, CCCA22, DCA23, FFVP21, MDRR24, PTK^{+}23]. Patrol [KSJZ23]. Pattern [ADCM24, BCSB24, FLRS12, GMH18, MJG24, MPT23, MK20, SSB21]. Pattern-Based [FLRS12, GMH18]. Patterns [Aln22, BHM^{+}19, CGMM22, DVED20, FAS^{+}23, FAB^{+}12, HLC24, HGF21, KVST19, PFT^{+}24, QNJ24, SG21a, VOP22]. Payload [KDE24, SHLH24]. Payment [NPWC24]. Payments [KIG23]. PBFT [CCA^{+}20]. PBFT-Inspired [CCA^{+}20]. PCA [GFPVLR^{+}23]. PdM [NZU20]. PE [IAM24]. Peach [AGA^{+}22]. Peak [ZSB24]. Peaks [LGY23, ZQCC16]. Peaks-Over-Threshold-Enhanced [LGY23]. PECSA [LW21]. Pedagogical [LS15]. Pedestrian [PCC^{+}18, SYZ19, VC23, XY21, ZN22, ZMZ^{+}22, ZHH24]. Peer [BBK^{+}20, CLT^{+}22, FKM24, GKS10, HWCH12, LWW22, LGCM24, MKDG20, PTK^{+}23]. Peer-Based [LGCM24]. Peer-Communication [BBK^{+}20]. Peer-to-Peer [CLT^{+}22, FKM24, GKS10, LWW22, MKDG20, PTK^{+}23]. Peers [LWW22]. PEKS [KHM20]. PEKS-Based [KHM20]. Pen [BAR^{+}21]. Pen-Tablet [BAR^{+}21]. Penetration [ABCC22, CMW18]. People [CWF13, FR16, GN16, LNB12, LA20, Mot24, PMT^{+}23, RRdBSS22, SC23, Söd13]. People-Counting [SC23]. Perceived [AP21, OBK23, RRAGMMGG20]. Perception [ASFAV23, CWC^{+}22, CCM^{+}17, GM16, GSD^{+}17, HL19, JDH21, SRMJ21, SS22]. Perceptions [LL20c, NWG20, PNMK22, SHA^{+}21]. Perceptron [AKM^{+}23, LBBPM21]. Percolation [BBMR19, GB20]. Perfect [PMGG21, WL22]. Perform [DSMM21]. Performance [AAY21, AZV16, AHH^{+}23, BARL24, BCSB24, BCS24, BMGI21, BBH^{+}24, CC24, DDMA24, GRM23, GKK^{+}19, HDBD23, HNHH23, HUM^{+}21, IBBA20, IZK^{+}20, KP19, KKS^{+}19, KHC^{+}23, KHRG19, KTD24, KAS^{+}20, LRD23, LZ23, LLZ21, LBB23, LLBB18, LFGR20, MLK22, MGV17, MS23, MPK^{+}23, MBF^{+}12, OLCMdA20, PFT^{+}24, PXC18, PKB^{+}24, PW22, PKS^{+}24, QQM24, RRP17, RRdBSS22, SO24, SHC22, SAR24, SOS^{+}23, SP22, TMP22, TSDG22, WWB^{+}20, XSM22, YCS17, YA24, YSEK20, ZZ21, ZSS^{+}22]. Performances [GMD^{+}22, GMGF24]. Perimeter [GLZL24]. Period [MPG20, WSD22, ZSZ18]. Permissioned [HP19]. Perpetration [ER21]. Persecution [PH20]. Persistence [CR23b, OF23]. Persistent [ABK22, PV20, PV22, YK21].

Persisting [AJ24]. Person [CZL21, CXLB21, MXL^{+}21]. Personal [Cop14, FOJE19, FMCT09, FSI^{+}24, GM19, HOPGRV21, LRdLS19]. Personalised [AFS^{+}22, SWY^{+}18]. Personality [CPI^{+}18, GCA^{+}22, SRM20, YYW20]. Personalized [ASK23, AML22, AMR^{+}21, ASMM22, AA24b, CH23, KS19, LCC^{+}24, TKH^{+}11, WYX18, ZPX24, ZSRC23]. Personalizing [KKTS22]. Persons [AAK24]. Perspective [AIPV21, CDBF19, Del23, Dua21, JII17, MDT20, QQX^{+}24, SVV^{+}22, SACG23, SSO^{+}19, SRM20, SAZ18, VLBRMP21, WZS^{+}22, ZS18]. Perspectives [BKR24, BCM^{+}19a, CMH10, HH20, HKH^{+}24, Mar13, PSC19, dSRRMC^{+}23, SLZJ23, VNJ18, YLX^{+}24]. Perturbation [VOP22]. Perturbations [RD18]. Pervasive [ALLR16, FFVP21]. Petri [DCHM16, SS23]. PETs [AKK^{+}24]. pFedBASC [ZPX24]. Phases [IAG^{+}21]. PhD [Giu18a]. Phishing [Ala20, BMP21, FLLC24, KHN^{+}22, LLB24, MSAA22, PMP22a, PB23, RAA^{+}22, SOE^{+}24]. Phone [KE22, KE22]. Phones [BCL11, MT17b]. Photogrammetry [SSR^{+}21]. Phubber [GRB^{+}21]. Phubbing [GDCM19, GRB^{+}21, SW21a]. PHY [MM21]. Physical [BCM^{+}23, CCC^{+}24, CPP^{+}20, DSM21, GW13, GTQ^{+}22, Gro22, GVC13, HBFJ20, JWD^{+}24, JSAO^{+}23, LH19, MMS22, MKHR21, MHL^{+}23, NPMPM23, OCCB20, PMP22a, PGSL20, QLR^{+}22, QQX^{+}24, QWR18, TH21, TKT^{+}18]. Physical-Layer [TH21]. Physically [OOM^{+}18]. Physics [Del21]. Pi [JOL22, RNST23]. Pickup [MKY24]. Pico [YPXH19]. Pico-Base-Station [YPXH19]. Piezoelectric [DMP^{+}23]. Piezoresistive [DDV22]. Pillars [AIPV21, Pil18]. Pipeline [KTD24, Mou24, NMT23]. Piracy [Tom21]. Pitfalls [AS19, NDKBL19, RS22]. Pixel [LCCC24]. PKI [TADS20]. Placement [BCZ^{+}23, BMJ^{+}22, BZNB23, JK20, LKIL19, RAJJ23, SV24, TL24a, YL21, ZWZ19]. Plan [AHMI22, DCA23]. Plan4all [FMNS11]. Plane [ATFMOU^{+}24, CFPP21, DH22, EW19, GRA^{+}23, GGK18]. Planes [VRA^{+}18]. Planner [DPR^{+}21]. Planning [CRMT23, CSC21, DGADE14, EARP^{+}24, FSC^{+}23, GM20, HC21, KAL^{+}19, KLG^{+}24, MGUD12, PG14, PFL^{+}12, Rot12, SPC^{+}11, SPB^{+}11, YKY18, ZYLM17, ZLL^{+}21]. Plans [ZL14]. Plant [ADK^{+}24, MCPLBH^{+}24]. Plate [RPB^{+}17]. Platform [Adi23, ACBP19, ALLR16, AHH^{+}23, BYG22, CKFH22, CALHG^{+}24, DCHM16, GRJ^{+}19, GMMI16, KZRH22, KHM24, KAS^{+}20, LGMZ19, Li18, LG20, LGCM24, LG23, MLK22, MBAS20, MKSV^{+}19, NMH^{+}21, RPB^{+}17, SSO^{+}19, SS14, TSHO24, TZS^{+}21, Tri14, VdHH^{+}19, VDE^{+}20, WSF23, ZLL18, ZSRC23, ZSS^{+}22, dCdPC^{+}21]. Platforms [AOI^{+}23, DVED20, FR24, FSI^{+}24, FGBMG19, GPA23, KDG24, MKvD11, MKSV^{+}19, NIK^{+}21, RLB^{+}23, RNST23, RFS22, SA24, XQLZ19, YA24, ZGH13, dCdG20, LAG^{+}18]. Plausible [GL12]. Play [LXY17]. Player [BHH12, GBDDVS23, OOM^{+}18]. Playful [CGGK22, LYMG17]. PLC [DNPC^{+}22]. Plug [LXY17]. Plug-and-Play [LXY17]. POI [BRM23, SJ21]. Point [GR22a, JK24, KSJZ23, Mar17, SH20, Tre22, ZYLM17]. Poisoning [ZXWZ21]. Poisson [CXZ23]. Polarity [SPC22, YKC19]. Policies [CS11, CDD^{+}21, DGKL15, HGF21, HWCH12, SRS23]. Policy [DWHY24, JAS21, Let16, LWW^{+}24, RH24, SHHM21, SSX20, TPD^{+}20, Ves18, YPMM12, ZLZ^{+}24, ZWL^{+}24]. Policy-Driven [YPMM12]. Policy-Engineering [SSX20]. Policy-Makers [Ves18]. Polish [OC20]. Politeness [LLR19]. Political [RBSMRN21, SS22, ZVV12, Zaf12]. Polling [TFSBSPSC24]. Pollution [AFB23, KKM^{+}22, XLJ^{+}22]. Polymediation [Hor19]. Polynomial [SST23, SZW^{+}23]. Polynomial-Based [SST23]. PON2 [Rók23]. Pooling [YWS22, YCHG18]. Pools [BTB22]. Popular [CQWC22, GRA21]. Popularity [UDF^{+}19, YSEK20, ZL21]. Populated [OFK19]. Population [Gla11, MPG20, WLJM21]. Port [FFP^{+}22, HVS17]. Portability [RFS22]. Portal [NB12, Tas10]. PORTAL-DOORS [Tas10]. Portals [Sor12]. Portfolio [PLL20]. Portrait [CHW^{+}21, WZL23]. Ports [RF23]. Portugal [dCMA21]. Portuguese [AdMTJ20, GVY^{+}23, MBAS20, MCR^{+}23]. Pose [COL22, KTAH^{+}23, TCH18, TVAT24]. PoseNet [COL21]. PoseNet-Based [COL21]. Position [SOR16]. Positioning [CZZH15, CW23, CBD^{+}22, DWZN16, FSY^{+}22, LLYL22]. Positive [LNB12]. Positivity [HP20]. Possession [YJSL18]. Possibilities [TMTMB20]. Possible [OB13, RBVV22]. Post [GAA21, HRGQHOA21, LBA^{+}24, LMY21, OLMBRCRC23, WKD22, ZMDCEMI22]. Post-Compensation [AD20]. Post-COVID-19 [WKD22]. Post-Digital [HRGQHOA21, OLMBRCRC23]. Post-Materialist [LMY21]. Post-Operative [GAA21]. Post-Processing [ZMDCEMI22]. Post-Quantum-Cryptography-Secured [LBA^{+}24]. Postgraduate [AP21]. PostgreSQL [SO24]. Posting [BCT21, SRS23]. Posts [BTB22]. Posture [FSG22]. PoT [KE22]. Potential [AAAAM^{+}21, GNK^{+}10, KKP22b, KBF^{+}22, PCM24, Tre21]. Potholes [Chu23]. PoW [QXC^{+}21]. Power [ABK22, BRF^{+}23, BAK23, CMPT24, CKFH22, DNPC^{+}22, DMZ^{+}19, ENT21, FCR23, GPM21, UGJA18, JWRX17, LNT^{+}23, LNT^{+}24, PPM21, SMG13, SKKS22, SXXH19, SNH^{+}19, SG21b, WYS17, WLZ18, YPXH19, YG20]. Power-Law [PPM21]. Powered [ABB23, AOF24, CGBF24, ENT21, HS20, RPB^{+}17, SNH^{+}19, WLZ18, YXL^{+}21]. Powerful [MKKS18, NPB12]. PQoS [MN21]. Practical [FOJE19, Gro17, LW21, OB13, TL19, VSE21]. Practice [AM24, GS11, KS14, MRRMC21, NS12]. Practice-Based [MRRMC21]. Practices [DN20, GNGHEM21, HKH^{+}24, IPL^{+}18, KLKA19, KTUF19, NDKBL19, SACG23, SSB21, Zaf12]. Practitioners [ACBP19]. Pragmatic [HL23]. Pre [AD20, AMC^{+}24, Lee17a, LHD^{+}24, LL20c]. Pre-Distortion [Lee17a]. Pre-School [LL20c]. Pre-Signature [AMC^{+}24]. Pre-Trained [LHD^{+}24]. Pre/Post [AD20]. Pre/Post-Compensation [AD20]. Preceding [HZW^{+}20, WHB^{+}21]. Precision [BYG22, KKS^{+}19]. Precoded [GVC13]. Precoding [CSDAM^{+}23]. Predict [DGMP21, FFVP21, GNK^{+}10, LCW^{+}18, SLP^{+}22, SG21a, WC23]. Predicting [AMAE22, AA19a, BS17, CCD^{+}21, CFG20, FSG22, GCA^{+}22, JDH21, MPK^{+}23]. Prediction [AMK^{+}21, Adi23, Alb19, AES21, AAC21, AAE^{+}22, BCSB24, CBR^{+}24, CN24, CQWC22, CSZ^{+}24, DAMAS^{+}19, EMHF19, EVCL21, FFM^{+}23, GS23a, GS23b, GVY^{+}23, HBM^{+}21, IHMG22, KTS18, LCC^{+}24, LMH19, LLLD21, LPA^{+}22, MKM^{+}19, OHSN24, PDT23, PLA^{+}24, RGCM21, SH21, SJ20, TBT^{+}23, TRHU23, TZS^{+}21, TCYZ22, WX20, WX22, WXL22, WLG^{+}22, WCZ23, WYZ24, WLCC24, XZY^{+}24, YZ23, ZWPL19, ZL20, ZWS20, ZZ19, ZLW21, ZZYC21, ZL21, ZXJ22, ZLZZ23]. Prediction-Based [LMH19]. Predictive [KPRP24, LZL^{+}21, NZU20, OAI^{+}22, Smi15, SG21b, WLZ18, XLJ^{+}22, YCS17]. Predictors [GFPD21]. Preemption [AMG22]. Preemption-Based [AMG22]. Preference [MSX^{+}21]. Prefetch [SMK^{+}22]. Prefix [SYB^{+}18]. Prehabilitation [ANAAM21].

Preliminary [CBKL22, KP15, RT23]. Premise [WZL21]. Preoperative [ANAAM21]. Preparedness [For14]. PrepaTec [JP21]. Preprocessing [NSD^{+}22]. Preschool [VCJAMN22]. Prescription [ATB^{+}24]. Prescription-Management [ATB^{+}24]. Presence [RK19]. Present [MSIP20, RKM^{+}22]. Presentation [AYRA21]. Preservation [AG22, Cop14, YJSL18]. Preserve [RFZ22]. Preserving [AMI23, BARL24, BFY22, CFV19, DQ18, FQ21, LDZP24, MTK^{+}24, SZZL23, VFFHF19, ZZL^{+}23]. Press [RBSMRN21]. Pretraining [ZLY^{+}20]. Prevent [SWP^{+}22]. Preventing [ABLR23, DL19, MT17a, RS22, SOE^{+}24, WLCS17]. Prevention [AMEF21, ACBP19, YWLY22]. Preventive [NRS24]. Price [FFM^{+}23, WHXL18]. Prices [CMP^{+}19]. Pricing [FN23, JLGF19, SAJ24, TOLG21]. Primary [LL20c, PTK^{+}23]. Principal [GFPVLR^{+}23]. Principles [DCG12, HKH^{+}24, PKF24]. Print [LWBM22]. Printed [CLZL18, VMD^{+}22]. Prior [SLZY21, TRLR24, ZH21]. Priorities [IUK^{+}23, oRL19]. Prioritisation [Tag20]. Prioritization [AMG22, KAEAB24]. Priority [CL16, YT19]. Priors [LXY17]. Privacy [AH22, AM23, AG22, AIR^{+}18, AAAKJ22, AVV^{+}23, AH21, AMI23, AS24, ABBV23, BARL24, BFY22, CDD^{+}21, DQ18, DGKL15, FKK20, FQ21, FC19, FSI^{+}24, Fri13, GJ15, GM16, GANT21, GSD^{+}17, HXHP17, HLZ23, HBFJ20, HBMS20, JAS21, KMW18, KPP^{+}20a, KSA^{+}21, KAS^{+}22, KPC19, KHM20, LDZP24, MSC^{+}24, MKKG19, MGJ22, MTK^{+}24, PMGG21, PPN18, PH16, RL24, SG12, Sca14, SRS23, SHC24, SZW^{+}23, Sta22, SG23, SZZL23, URH^{+}23, Ves18, VKK^{+}19, VKKG22, VFFHF19, WL21, YJSL18, ZR15, ZZL^{+}23, vdSFRF23]. Privacy-Aware [AS24].

Privacy-Conscious [CDD^{+}21]. Privacy-Oriented [KPP^{+}20a]. Privacy-Preserving [AMI23, BFY22, LDZP24, MTK^{+}24, SZZL23, VFFHF19, ZZL^{+}23]. Private [CTNS21, DMZ^{+}19, GMGF24, MT17b, SCC^{+}23]. Privilege [HAA23]. Privileged [SB22]. Proactive [HF22, LW11, SSF^{+}18]. Probabilistic [ABS20, GCEOBRT23, TPKA20, ZLL18]. Probabilistic-Feedback [GCEOBRT23]. Probability [JK24, KIB^{+}22, WLL20]. Probability-One [JK24]. Problem [GDLG18, GLH^{+}22, MKY24, SU17b]. Problem-Solving [GDLG18]. Problems [Alh22, Lal14, LBBPM21, SRMJ21, VI16, ZCG^{+}24]. Process [BGL^{+}22, CLTP22, CXZ23, ER10, FMNS11, FLSL24, KAEAB24, MSZ23, MBAS20, OOM^{+}23, SFW23, TWM20, TRLR24, Tra18, TLV23, VCPPRC20, WCSZ18, YPG20, ZBN21, ZH21]. Processes [APB^{+}16, CGRV20, KGWR23, LHSS12, MSZ23, PE13, Pil18, PLA^{+}21, PR12, TMTMB20, YMSP24, ZLL18]. Processing [Chu23, CS22, HOV23, LSZ21, PSGM22, PD23, Rei23, RAA^{+}18, SSR17, STV^{+}22, VdHH^{+}19, WZL21, YQ24, ZMDCEMI22, vSR24]. Processor [MLPH23]. Procurement [DWY^{+}19a]. Producer [KC19]. Product [LTP24, YWW^{+}23, ZLZZ23]. Production [FLSL24, JWD^{+}24, MKHR21, VP22]. Products [CMP^{+}19, HSM19, NDKBL19, TN20]. Products-Empirical [HSM19]. PROFEE [GCEOBRT23]. Professional [ABSG21, KVVD22]. Professionals [CCCA22, Nil12]. Professor [DHG21]. Professors [RRAGMMGG20]. Proficiency [LS23]. Profile [LLZ21, LCW^{+}18, RMLAZOPC21]. Profiled [AKDL24]. Profiles [OOM^{+}18, TL24b]. Profiling [GTM22, SAL21, SL24b, WL22, ZGH13, dVBA23]. Prognosis [SCM12]. Program [ANAAM21, APB^{+}16, GDS23, RSF23]. Programmable [DH22]. Programming [CZY17, DH18, KS19, MB22, MTKS24]. Programs [AS19, GL12, KPB23]. Programs-Pitfalls [AS19]. Progress [XCW^{+}22]. Progressive [JHC^{+}20, WCZ23]. Project [CSZ^{+}24, CGM^{+}19, FMNS11, ZZYC21, ASM^{+}24]. Projects [BOH23, OC20, PR11, PLOT23]. PROMETHEE [KKTS22]. Promising [AAN^{+}24]. Promote [ALST20, NE21]. Promoting [AGN21, RSG24]. Prompt [ALL24]. Prompt-Based [ALL24]. Pronunciation [KI20]. Proof [AMEF21, BMB23c, GKG19, HM20, HDT22, MPT23, QXC^{+}21, SW21b, WLL20, XNC21, ZLLW18]. Proof-of-Concept [BMB23c, GKG19]. Proof-of-ENF [XNC21]. Proof-of-Probability [WLL20]. Proof-of-Stake [AMEF21]. Proof-of-Work [AMEF21, QXC^{+}21]. Propaganda [RBSMRN21]. Propagation [ACD^{+}22, GZS^{+}24, SLB24]. Prophet [YZCR24]. Proportion [MBLC22]. Proposal [CTM22b, CT24b, ELCS20, EVCL21, FMMS21, MFO^{+}20, Mot24, NNRR^{+}21, ZCG^{+}24]. Proposed [KKEV17]. Prosocial [CGGK22]. Prospective [ST20]. Prospects [GTQ^{+}22]. Prosumerism [PE13]. Protecting [SSDS23, SZW^{+}23]. Protection [AM23, AMZ^{+}20, FSI^{+}24, Fra24, GL21, HXHP17, HWCH12, JT23, SSS21, TRC^{+}21, VKKG22, WL21, ZWWW17]. Protocol [AAAAK24, AMI23, AKJB20, BBSR24, BAS24, CFPP21, DGF23, EABE19, FPGO24, GPPB^{+}21, GMS^{+}20, HLZ^{+}18, HS18a, HF14, HJA18, KMS^{+}12, KKEV17, KC19, MSK^{+}18, OLCMdA20, POC20, RRP17, SCA^{+}19, SS15, WCM19, YT19, YT15, YLYL22, YW16, YLH^{+}17, ZOAR24, ZLLW18, CISM22]. Protocols [AK18, Al 20, ACG22, AVA^{+}22, BM14, CSL^{+}20, DPC^{+}24, Fra23, GKW^{+}10, HNK^{+}21, Jal19, KKP22b, KTUF19, MLK22, ORK^{+}19, OSdS24, SBS18, TDW24, TT22]. Prototype [AS15, LSM^{+}12, Tri14, VKV^{+}22, WBG12]. Prototyping [PORM^{+}23]. Provable [YJSL18]. Provenance [WADL^{+}24]. Provide [ALST20]. Provided [LS15]. Provider [MP16]. Providing [CH16, RFS22]. Province [ZLW21]. Provision [PCM24]. Provisioning [BdCRM10, BOK^{+}21, ELC^{+}19, KOJP13, MG10, YC16]. Proximal [ZWL^{+}24]. Proximate [KKBD10]. Proximity [AIP^{+}24]. Proximity-Driven [AIP^{+}24]. Pseudo [MLWS24]. Pseudo-Labeling [MLWS24]. Pseudolite [LLYL22]. Pseudonymization [BKW^{+}24]. PSI [TDW24]. PSI-CA [TDW24]. Psychological [CFGD20, MPG20, NIK^{+}21, VGS^{+}19]. Psychometrics [WLH^{+}17]. Psychophysiological [KKL^{+}20]. Public [CB22, CS11, Cop14, FAT19, Gar12, GTM22, HEP^{+}11, HDT22, LMY21, MLL^{+}22, PCM24, Pic21, QA24, SS14, TPP^{+}23, VKKA17]. Publication [GM19]. Publicly [GMB^{+}21, RSV^{+}24, SZW^{+}23]. Publishing [ACAP20, Fra18, LCS12, PR11]. PUE [MPCS^{+}23, MPM24]. PUF [BBSR24, LDZP24, SDS^{+}22]. PUF-Based [BBSR24]. Pulverization [CPP^{+}20]. Pump [RM23]. Punishment [Mar13]. Purchase [HSM19]. Purposes [AMAA^{+}21]. Push [LBC18, SM19]. Push-Based [SM19]. Pyramid [YCHG18].

Q [CASN21, WWD21]. Q-Learning

[CASN21]. QCA [ZY20a]. QGIS [SSR17]. QoE [EMHF19, HE20, AWSA23, BS17, DTTK19, EH21, LXG^{+}23, LWZL22, NNT22]. QoE-Aware [EH21]. QoE-Based [LXG^{+}23]. QoS [AWSA23, AALS21, AZ19, AFB23, BdCRM10, BS17, BCG10, CL16, DLK10, GL13, KCY22, KOJP13, KC19, MG10, MDRR24, Pin10, Sul22, UFK^{+}10, YC16, YPMM12]. QoS-Aware [MDRR24]. QoS-Based [Sul22]. QoS-Satisfied [KCY22]. QRNG [dCdZCC24]. Qualitative [FGSD22, FMMS21, NB12, Oza23]. Qualities [ALDS19]. Quality [AK22, ASL17, BAS24, CMH10, CTM21, CR11, CMR24, DHEK19, EMHF19, Gao21, GKK^{+}19, HE20, HAL^{+}22, IUIL23, JCPV22, KTS18, KZ20, LLY23, MM23, NKSF24, OBK23, PZH^{+}18, PDMK19, SB22, SYB^{+}18, SMG13, SSR17, hSgLH17, TKT^{+}18, XLJ^{+}22, YT18, Zha19]. Quality-of-Service-Linked [SB22]. Quantifying [BOK^{+}21]. Quantile [EPC24]. Quantitative [FGSD22, LS15, SL24b]. Quantization [WLYL20]. Quantum [AZH22, CFPP21, Del23, FPGO24, IUK^{+}23, LBA^{+}24, PL22b, Rah22, RGB^{+}23, RAR^{+}23]. Quarantine [MPG20]. Queries [ARM20, BB23, RGCM21]. Query [HN18, JBS^{+}23, MXT^{+}16, PSGM22, XY18, ZZL^{+}23, ZZZL23b]. Query-Driven [JBS^{+}23]. Querying [ACG^{+}24, BB23]. Question [ZZZL23a, ZZZL23b]. Questionnaire [BBM20, LL20c]. Questions [LS15]. Queue [RZQS18]. Queueing [BMGI21, LDM22, RRdBSS22]. Queueing-Based [RRdBSS22]. Queues [CL16]. Queuing [KLG^{+}24, LB24]. QuickFaaS [RFS22]. Quickly [Ant24]. Quo [dSRRMC^{+}23, ST20].

R [LBK22, QQM24, WKW18, YLJ^{+}19]. R-CNN [LBK22, YLJ^{+}19]. Racial [WT13]. Rack [CZY17]. Radar [KES24]. Radio [ADAB21, ASM^{+}24, AMF^{+}10, BKL^{+}22, CZK^{+}17, FM20b, HF20, KZRH22, LZL^{+}23, MPCS^{+}23, MPM24, MHQ^{+}23, OSdS24, SS23, SKT^{+}19, WK24, WWYQ19, ZSZ18]. Railway [KPC^{+}23, SMS^{+}24]. Rain [MR23a]. Raising [PR12]. Ramp [SYGY21]. Random [ASF^{+}23, BB20, Han13, OAI^{+}22]. Range [AOF24, ACA^{+}23, FM20b, KTD24, TFSBSPSC24]. Ranges [SKFM24]. Rank [ZVKK19, ZCdRR^{+}23]. Ranking [RCGSL19, RCL21]. Ransomware [ADSAKAD22, DKV24, GS23a, GS23b]. Ransomware-Resilient [ADSAKAD22]. Rapid [GPCP12, SLB24]. Rare [BMB^{+}23a, LWBM22, TADS20]. Raspberry [JOL22, RNST23]. Rate [ACA^{+}23, CR21, DNZW22, GCEOBRT23, JKM^{+}23, KMK24, KHAM24, LWY20, LMLW18, LMH19, LNH23, MTA^{+}20, MM23, OGB^{+}24, QWR18, RRP17, REP22, WZG^{+}18, YZP22, dOB19]. Rate-Controlled [MTA^{+}20]. Rating [FPT20]. Ratio [KLTC17, SJG18, ZLT22]. Rational [QAM17]. Rationality [LYP^{+}19]. Raw [DGADE14]. ray [VR21]. RDF [VRA^{+}18]. RDTIDS [FMA^{+}20]. Re [CZL21, CXLB21, HCW20, MAW^{+}19, MXL^{+}21, THL21]. Re-examining [HCW20]. Re-Identification [CZL21, CXLB21, MXL^{+}21]. RE-SWOT [THL21]. Re-Use [MAW^{+}19]. Reachability [MLM14]. Reactive [TT22]. Read [CDLY24]. Readability [MG12]. Reader [AM21, JLZ18]. Reader-Tag [AM21]. Reader/Router [JLZ18]. Readiness [ESB^{+}20]. Reading [ATY20, Øie12]. Readmission [TRHU23]. Ready [MRRMC21]. Real [AMCI17, AAG^{+}22, AOF24, AGO^{+}23b, AGA^{+}22, AGN21, BSE^{+}21, CMPT24, CBKL22, CALHG^{+}24, DCA23, For14, GSP15, GPM21, HQC^{+}24, IMR^{+}18, ISG20, LCS12, LBC18, MXT^{+}16, MDB22, MMF^{+}19, NPM^{+}18, OO19, PG23, PPDC22, RR24a, RIS^{+}17, SH21, TVAT24, TZE^{+}21, VGK21, VGS^{+}19]. Real-Time [AMCI17, AAG^{+}22, AOF24, AGA^{+}22, CMPT24, CBKL22, CALHG^{+}24, DCA23, GPM21, HQC^{+}24, IMR^{+}18, MXT^{+}16, MDB22, PPDC22, RR24a, RIS^{+}17, SH21, TVAT24, TZE^{+}21, VGK21, VGS^{+}19]. Real-to-Virtual [For14]. Real-World [AGO^{+}23b, BSE^{+}21, GSP15, NPM^{+}18, OO19, PG23]. Realistic [IAM24, LZC21, MGV17, VOP22]. Reality [AARTSBAE24, Ant12, ASFAV23, ASFAAV24, FHL^{+}24, FSI^{+}24, FRE^{+}22, GBDDVS23, HF22, IKS^{+}21, JGV^{+}21, KTS18, MVAHBL^{+}23, PEN23, SN23c, ZN22, ZCG^{+}24]. Realization [DNKDN^{+}24]. Reasoning [AGN21, CCCA22, DCHM16, DGGS19, FAS^{+}23, GL12, HCHP16, LLJ16, MKHR21, PCROB21, SU17c]. RECAP [BSE^{+}21]. Receiver [AAR^{+}19, BHS24, YT15]. Receiver-Triggered [YT15]. Rechargeable [CCTC23]. Recipe [GDLG18]. Recognition [ATH18, AWUF19, AF22, ADCM24, BKW^{+}24, CHMZ21, CC22a, CKSG21, DYLZ21, DMP^{+}23, FM20a, GXWD19, GCA^{+}22, GRA21, GR22b, HAL^{+}22, HYLP19, ITH^{+}21, LJWX23, LDZ^{+}23, LMPT19, LF23, MKKS18, MSL^{+}23, NBL^{+}24, RGM^{+}21, RPB^{+}17, RSG21, SBRZ24, SGMMRG21, SOA^{+}20, SYZ19, SDL^{+}19, USS19, VCG24, VGK21, XGN19, YCHG18, ZL19b, ZZSX19, ZZZ^{+}22, ZXMB22]. Recognition-Based [USS19]. Recommendation [ASK23, ACF24, AHD21, AYH^{+}21, BdBCCG21, CH23, GAGB19, LHD^{+}24, LPM10, ON23, RAK^{+}20, SJ21, WYX18, XY18, XDS^{+}19, ZSF20]. Recommendation-Based [ACF24]. Recommendations [ITZ20, KPB23, KKTK20, PEN23, TL17]. Recommender [BRM23, KDEA^{+}23, SGMMRG21, SOB^{+}19, SDM22, ZSRC23]. Recommenders [LWH24]. Recompression [LLCC24]. Reconfigurable [IHCSG^{+}24, IBBA20, SOK24, WWYQ19, ZX24].

Reconfigurable-Intelligent-Surface-Enhanced [ZX24]. Reconfiguration [CBR^{+}24, QLL^{+}21, TXJ23]. Reconnaissance [BMB^{+}23b]. Reconstruction [SSR^{+}21, TDT^{+}21]. Reconstructor [LH21a]. Record [CCCA22]. Records [CCTV24, DKH24, GRG21, LTTS21, PTK^{+}23, SYLL21, vSR24]. Recovering [BLW^{+}17]. Recovery [AHMI22, DDZ18, SNS22, ZS18]. RecPOID [SJ21]. Recruiters [VNJ18]. Recruitment [VKKA17]. Recurrent [ATH18, OAH^{+}24, SJ20]. Recycling [ZTBD20]. Redactable [AYH23]. Reddit [BCT21, BTB22]. Reduce [QLK^{+}24, SOSR^{+}16]. Reduced [GFPVLR^{+}23, QAQA21]. Reducing [AAAT21, ELC^{+}19, OT16, RSG21, SAF14]. Reduction [LNH23, NMT23, SESO24]. Redundancy [Fio10, OS23]. Redundant [CLTP22]. Reference [BML19, DHEK19, MEH^{+}22, PPP21, RF23]. Referencing [WSN24]. Referral [ZY15]. Refined [MLWS24]. Refinement [HY22, ZZZL23a]. Reflecting [SF22]. Reflective [VI16, Whe09]. Reframing [KG18]. Refreshing [PVM22]. ReFuse [GMLS22]. Regard [PG23]. Regards [SOSC^{+}16]. Region [KNN^{+}24, TC21]. Regional [GGD23]. Regions [NZZ13]. Register [LL23a, SZL21]. Registers [Gla11, KFK^{+}24]. Registry [MJG24]. Regression [BZMB23, CN24, LLY23, MBLC22, SGMBK24, SQK^{+}17, YWW^{+}19]. Regressions [OMTFFL21]. Regular [KK23a]. Regularization [MLWS24]. Regulated [Lai24]. Regulation [BFFZ^{+}12, LGMZ19, PKCF21]. Regulatory [GSB^{+}23, LMF23]. Rehabilitation [COL21, GAA21, MMS^{+}19, SCM12]. Reinforcement [BHC23, DXL^{+}22, FHXL19, FHZ^{+}24, GWF^{+}23, GHGL22, GHL^{+}23, JUN^{+}24, JHBW24, JGY^{+}24, KALY17, LZLW22, MCS^{+}24, MSC^{+}21, PSTZ23, QSML24, QWR18, SOK24, SPK24, SF22, TCH23, TFKY19, TKT^{+}18, TCYZ22, VC23, WZLW19, WLLY24, WLWZ23, YBO^{+}23, ZX24, ZY20b]. Related [AWBS23, Dua21, FZ15, FGBMG19, JSPH21, Min20, SLSK21, YMJ20, ZL14]. Relation [DGADE14, KA23, RSV^{+}24]. Relational [KPCS13, LDZ^{+}23, WX22]. Relations [KVL^{+}22, PM17]. Relationship [CPI^{+}18, HT14, Mar13, SRM20, WZL21]. Relationships [TL17, TL24b]. Relay [CFP17, DYS15, EARP^{+}24, JP20, SJG18, SHC22, SXXH19, SNH^{+}19, THV^{+}22]. Relay-Assisted [SHC22]. Relaying [KFP10, SJG18, SXXH19]. Released [OFV21]. Relevance [RCGSL19, XY18]. Relevant [CDBF19, Tre22]. Reliability [DCHM16, FAB^{+}12, SWJ20]. Reliable [AAF^{+}22, BZMB23, CFP19, DNZ22, Lee17a, MKK^{+}22, MM21, NSV17, ZAS^{+}24]. Remaining [NRS24]. Remote [ANAAM21, AMI23, Ciu19, DDV22, KV23, LWY^{+}15, MJG24, POC20, SHLH24, YLX^{+}24]. Removal [Fio10]. Renaissance [KVL^{+}22]. Renewable [CMPT24, KAL^{+}19, TH20]. Rent [PW22]. Rent-Seeking [PW22]. Rented [FN23]. Reopening [CCD^{+}21]. Reordering [LCCC24, LHW24]. RepDroid [TPKA20]. Repetition [LMLW18]. Replacement [DWHY24, LNA17]. Replica [QLR^{+}22]. Replicas [LZS18]. Report [MT17a]. Reported [SW21a]. Reporting [DCD24, GK21, KAS^{+}20, TZE^{+}21]. Repositioning [CYC^{+}21]. Repositories [GPRMGP21, SKM21, WSN24, WKPC18]. Repository [KPCS13, PTM20]. Representation [LLJ16, PCROB21, SU17c, SCM12, SSX20, XZY^{+}24, ZZZ^{+}22]. Representational [DGADE14]. Representations [GDS23, HL23]. Representative [CLTP22]. Representing [FM21]. Reproducing [Söd13]. Reproduction [CCCA22]. Reptile [JGY^{+}24]. Repurposing [ZC21]. Reputation [CPI^{+}18, CFGD20, DPBG18, DCPG21, KST19, KUBS22, LGMZ19, TPKA20, TLC15]. Reputation-Based [KST19]. Request [JLGF19, LNA17]. Requests [VL18a]. Requirement [STG^{+}20, ZZ19]. Requirements [AP22, AZ19, BAK23, CPM^{+}19, KAEAB24, LCMV17, MNWK21, RF21, SCC^{+}23, TH20]. Requiring [ANAAM21]. Resampling [BMB^{+}23a, KM24]. Rescue [KVST19, SZC19]. Research [AP22, AEA^{+}23, AYH23, BM23, Cal22, CXG^{+}22, CEZC19, DHG21, Del23, DCA23, GHZY23, GPCP12, GW13, HPC12, ITZ20, JSPH21, KK23b, KA23, LLL21, LTM^{+}22, NZ14, Ngu19, OC20, OP22, PNdC^{+}23, QLR^{+}22, RNFS20, RGdSK22, RKM^{+}22, SOSR^{+}16, SWC^{+}23, SN23b, SL24a, SEB^{+}19, TSZ^{+}18, Tre22, VTPT24, WHB^{+}21, WKF^{+}21, WZLL19, XGY^{+}22, XCW^{+}22, XSM22, Yan19, ZHH24, ZLL^{+}21, ZZY19, vdSFRF23]. Researchers [CGMM22]. Reshape [RF24]. Reshaping [Cal24]. Resident [CCCA22]. Residual [CCYL24, LXL^{+}20]. Resilience [AVV^{+}23, GKFV24, KYG22, SRSDF23, TSD24, VDKL21, WZS^{+}22]. Resilient [ADSAKAD22]. ResNet [GR22a]. Resolution [AM24, LYD22, LL23b, dMLMS22, LHW18]. Resolving [AAAAK24, YK21]. Resource [AJ18, ACG^{+}24, AlQ23, ASL24, BKL^{+}22, BHC23, BSEL20, BM18, ELCS20, EVCL21, FRR^{+}24, GWF^{+}23, HKZ^{+}22, HWCL17, HZCZ22, ICF^{+}11, JWRX17, JHBW24, KSG24a, KALY17, KCY22, KLG^{+}24, KSSF19, KLZ16, LJR23, LJR24, LHSS12, MCR^{+}23, MVMHL24, NST23, SA21, SGLZ20, SIJA10, SHN^{+}19, TRST24, UFK^{+}10, WSM20, YPMM12, ZLY^{+}20, ZWL^{+}24, ZX24]. Resource-Constrained [FRR^{+}24, SA21]. Resource-Efficient [LHSS12]. Resources [ELC^{+}19, Gao21, ICF^{+}11, KAL^{+}19, LHL24b, Pet20, SFD20, TCG19, YLPW21]. Respiratory [DDV22, IHMG22, VMAG23]. Response [ABSG21, Cro19, KT22, PFL^{+}12]. Responses [CS11, CSA^{+}24, HC19]. Responsible [AHP23]. Responsive [SWZ18, SWY^{+}18]. ReSQoV [KCY22]. REST [HS18a, LNG19]. REST-Security [LNG19]. RESTful [KMS^{+}12, dCdPC^{+}21]. Restoration [JMZ^{+}22]. Results [BRP^{+}13, HNHH23, KA23, SDDB24, SRM20]. Resume [SZT23]. Retailers [TLV23]. Retailing [FRE^{+}22]. Retention [AH21]. Rethinking [KXZ^{+}24]. Retrieval [BMS24, CZZZ22, DPR^{+}21, FWL22, LWBM22, RRT20, ZZZL23a]. Retrieving [FAB^{+}22, Uto13]. Retrospective [AAN^{+}21]. Reuse [ASK23, WADL^{+}24]. Reverse [TYH23]. Reversible [CCKH21, WLCC24]. Review [AK18, AH22, AZH22, AP22, AZAV22, AAMD21, AN22, ACG22, ADS20, AEA^{+}23, AYH23, AA21b, AHD21, ACA^{+}21, AAA^{+}20, BMS20, BAK23, BTS^{+}21, BG09, BFL20, CPKK23, CR23a, CWC^{+}22, DRK^{+}22, DIG20, ER10, ESBE23, EWGK24, EUDW23, FAC22, GHA^{+}21, GDARM18, GGW18, GM23, HSC^{+}22, HYMT24, HHM^{+}23, HP16, HAL18, IHCSG^{+}24, IDR23, IT24, KGKP20, KT22, KIG23, KRSK24, KPKM23, KMV^{+}23, Lee20, LH21b, LOL21, LTP24, LGV13, MVAHBL^{+}23, MCT^{+}23, MR23b, Mot24, NPWC24, PAK^{+}23, PJMM22, RNFS20, RGB^{+}24, RAR^{+}23, dSRRdOT21, RT23, SMVP21, SGMMRG21, SKKS22, SE15, SRS^{+}24, SOS^{+}23, SKFM24, TSD24, UW21, VLBRMP21, VLK24, WMS23, XLW^{+}24, XCW^{+}22, YMSP24, YHN21, ZKKK19, ZDP^{+}22, vdSFRF23]. Reviewers [Fut20, Fut21, Off14, Off15, Off16, Off17, Off18, Off19, Off22]. Reviewing [AS19, DCPG21]. Reviews [TPKA20, YWW^{+}23]. Revisiting [IBBA20]. Revitalization [MBAS20]. Revocable [WZX20]. Revolution [CMS21, Jur12]. Revolutionize [ACS24, LQFI17]. Reward [ZY15]. RF [CBD^{+}22, LZ23, XZ19]. RF-Acoustic [CBD^{+}22]. RFI [LOL23]. RFID [ASF^{+}23, AM21, ESBE23, HJA18, JLZ18, PPMMS19, PBAAN19, SBS18, SOA^{+}20, WCSZ18, ZLLW18]. RFID-WSN [JLZ18]. RGB [DYLZ21, LBK22, MXL^{+}21]. RGB-IR [MXL^{+}21]. Rich [XSM22]. Rich-Features [XSM22]. Rider [MDB22]. Rights [KPP^{+}20b, Tsi14]. Rigidity [PFdRGG21]. Rigorous [MKPG22, Tre22]. Rikki [Tsi14]. Ring [HHLZ23]. Ring-Architecture-Based [HHLZ23]. RingFFL [HHLZ23]. RIS [CSDAM^{+}23, LLW^{+}24]. RIS-Assisted [CSDAM^{+}23]. RISC [DNKDN^{+}24, MLPH23]. RISC-V [DNKDN^{+}24, MLPH23]. Risk [ASCM22, AHMI22, AAA^{+}20, BPG19, CR23a, FC19, GDCM19, Lee20, MSC19, NWG20, PLA17, PPP21, PR12, SWJ20, SG21a]. Risk-Based [AAA^{+}20]. Risk-Taking [SG21a]. Risks [dSFdRP24, HBFJ20, PS16, Rot12, SESO24, UW21]. Risky [JLW19, SAF14]. RLXSS [FHXL19]. Road [ADG^{+}19, KKK^{+}20, LLW^{+}17, RYAA22, SMS^{+}24, USS19, YT19]. Roadshow [LOL21]. Robin [IUK^{+}23, SIJA10]. Robocov [PMP^{+}22b]. Robogames [OOM^{+}18]. Robot [AAG^{+}22, MKSK23, PS17, PMP^{+}22b, SU17a, SZC19, TTO^{+}19]. Robotic [AMVM22, MKSK23, SVK^{+}22, TEE10]. Robotics [CCM^{+}17, DG18, KG18, SU17b]. Robotics-Case-Study [SU17b]. Robots [CYO^{+}17, CSC21, FHL^{+}24, KG18, LYMG17, Rec21]. Robust [ABSG21, AAA^{+}24a, BCSB24, SGW22, SS17, SKHK23, TDT^{+}21, VOP22, YHSW19, ZWZ17]. Robust-Learning [SGW22]. Robustness [HCXH16, KAAAA22, KL21, MMT24]. Rockfall [NJH^{+}24]. Rockfalls [Chu23]. Rogue [AA19a]. Role [AG22, DTTK19, DCPG21, ER21, FC19, GTQ^{+}22, GPLK22, GSD^{+}17, HCW20, JNARV^{+}24, Khr20, KA21, KPC19, KDKG22, LZLW22, PB23, SLS20, SWGL19, TG13, VKKG22]. Role-Mining [SWGL19]. Roles [HGF21]. Roll [ZWPL19]. Romanian [ND24]. Rome [PBB20]. Rooms [RGP22]. Rosenbrock [LSJ19]. Rotary [ZYWS24]. Rotation [LWHR14, SBS18]. Rotational [Tra18]. Round [IUK^{+}23, SIJA10]. Route [OT16, TRC^{+}21, YT18]. RouteNet [DDMA24]. Router [JLZ18, XDH24]. Routers [YLK^{+}19]. Routes [MCPLBH^{+}24, OT16]. Routines [NS12]. Routing [ACG22, AVA^{+}22, AAF^{+}22, ABS20, BdCRM10, Bi14, BAS24, CXG^{+}22, ELC^{+}19, GLZL24, Goe12, GLH^{+}22, JBS^{+}23, KKP22b, KZ24, KLTC17, LBB23, LCW^{+}18, LXL^{+}19, LZL^{+}21, LTP24, LW15, MMA10, MLK22, MWY19, MZH22, MLM14, OT16, PZH^{+}18, TAHO23, TT22, VL18a, WCM19, WY19, YC16, YW16]. Routing-Delivery [LCW^{+}18]. RPL [WC17]. RSS [SOA^{+}20]. RSSI [KTAH^{+}23]. RT [LLYL22]. RT-ANN [LLYL22]. Rule [ARM20, KKT20, Let16, YPMM12]. Rule-Based [ARM20, KKT20]. Rules [FMA^{+}20, HZCZ22]. Rumor [PW21]. Running [NIK^{+}21, RGCM21]. Rural [AMC^{+}24, EARP^{+}24, Hel15, Hua15, OLMBRCRC23, RS17, VLBRMP21]. Russia [YBMK21]. Russian [Ant12, FRKS22, RKS^{+}21]. Russian-Language [FRKS22, RKS^{+}21]. Rwanda [NZU20].

S [FC19, WZG^{+}18]. S2NetM [PLG23]. S4 [PMP^{+}22b]. SA [RAJJ23]. SA-MCSDN [RAJJ23]. SAES [Sac19]. Safe [CTM22a, EAAA^{+}23, ML20]. Safer [RYAA22]. Safety [AGO^{+}23b, CGT21, DHDAu22, DGF23, FTHY21, FAT19, HT14, HZCZ22, HYL13, IPP24, UMS24, KKG20, KPC^{+}23, KSG^{+}24b, LT24b, MDB22, ML20, UANU20].

Saharan [APAR22]. Sales [ZXJ22]. Salp [CD23]. Same [LLW^{+}17]. Sample [BM19, BBM20, LRdLS19, VYN19]. Samples [UJ22, VMD^{+}22]. Sampling [JHL^{+}18, LXY17, YWW^{+}19]. Sandbox [NIK^{+}21]. Santa [WBG12]. Sardinia [FZ15, ZL14]. SARS [CR21]. SARS-COV-2 [CR21]. SASLedger [SPS^{+}21]. Satellite [CXG^{+}22, CFP19, CFPP21, FSC^{+}23, GMLS22, HLZ^{+}18, ISLARP^{+}22, KLKP19, KKM^{+}22, KV23, MR23a, QLL^{+}21, REP22, VL20, VLC^{+}19]. Satellite-Terrestrial [QLL^{+}21]. Satellites [CXG^{+}22]. Satisfaction [GHGL22, HKZ^{+}22, STG^{+}20, SWZ18, ZS18, ZN22]. Satisfied [KCY22]. Saturation [KAAAA22]. Saudi [AAN^{+}21, AP21]. Saving [DXL^{+}22, Lee17b, LYP^{+}19]. Savings [CPM17]. Sayonara [CYO^{+}17]. SBNNR [SGMBK24]. SBT [Ant24]. SC-FDM [DNPC^{+}22]. SCADA [CISM22, MEJA23, TSZ^{+}18]. Scaffolding [BdBC23]. Scalability [ABB23, HM20, SN23b]. Scalable [AJA^{+}24, AGN21, KGRP21, KCY22, NBT22, PNM^{+}20, SADP21, SPS^{+}21]. Scalar [AKDL24]. Scale [AFB23, AA24a, GMRRRQ^{+}16, HP21, JXA19, JHC^{+}20, KGK^{+}24, LF23, MR23b, QHX^{+}24, SKKS22, VGK21, VBB^{+}24c, WLJM21, XY21, XSSA24, YCD^{+}23]. Scales [DIG20]. Scaling [DGCK24, TMS24]. Scan [APB^{+}16]. Scanning [HVS17]. Scenario [AALS21, FN23, OSdS24, SZC19, Tre21, WML24]. Scenario-Based [OSdS24]. Scenarios [Ant24, AMF^{+}10, BBH^{+}24, DTR22, KVWC24, MGV17, MHSH16, NA23, SADP21, SMS^{+}24, SYZ19, SEB^{+}19, Tre22, VM24b, WC17, ZMZ^{+}22]. Scene [GGD23, LJWX23, LL20a]. Scheduler [ABK22, AGC18]. Scheduling [ANHMS24, ARAAA19, And21, AD18, CD23, FML^{+}22, IUK^{+}23, KAL^{+}19, LW19, LSZ21, MMA10, MC24, NRS24, oRL19, SU17b, STCP21, SGLZ20, SIJA10, Sul22, SZFC18, TGS^{+}19, WHXL18, WLLY24, XQLZ19, XNZ23a, YCS17]. Schema [DH22]. Scheme [AJ18, AOI^{+}23, ADM19a, AALS21, AZ19, Alb21, ADS20, AASAI22, AMC^{+}24, ABS20, BP24, BL22, CGLR19, CCC19, HK18, IZK^{+}23, JUH^{+}23, LWW22, Li23, LLCC24, LCW^{+}18, LH21c, LCCC24, LN19, MXT^{+}16, Mah17, PZH^{+}18, PM24, PAM^{+}24, PYC21, QMN19, SJG18, SZL21, SOK24, SXXH19, SZZL23, SMK^{+}22, TPKA20, WSM20, WC20, XSX18, XNZ23b, YY17, YPLZ19, ZWZ17]. Schemes [JKM^{+}23]. Schizophrenia [PBP^{+}24]. Scholar [LTM^{+}22, RCGSL19, RCL21]. School [ER21, GNGHEM21, JP21, LL20b, LL20c]. Schools [MS20, TC21]. Science [BBMR19, BOKC19, Del21, HRKL21, Let16, LAG^{+}18, MT20, PBB20, Por21, Tos23, MGLBMMSC20, YMJ20]. Sciences [WKF^{+}21]. Scientific [AD18, EM09, ER10, FAB^{+}22, FAS^{+}23, HC21, LS23, MGLBMMSC20, WBG12]. Scientists [Ves18]. Scientometric [SGMMRG21]. SciFlow [ER10]. SCMC [Mah17]. Scope [HNHH23, LMF23]. Scoping [MVAHBL^{+}23, vdSFRF23]. Scopus [RCGSL19]. Score [AA24a, MGBG21]. Scores [FPT20]. Scoring [AGO^{+}23b]. Scratch [PJMM22]. Screen [CCCA22]. Screening [GVY^{+}23, Pri23]. SD [EH21, GLZL24, TRC^{+}21]. SD-BROV [TRC^{+}21]. SD-GPSR [GLZL24]. SD-WAN [EH21]. SDCG [ZQL^{+}24]. SDMw [HVS17]. SDN [Pap20, AV23, AF21, AHMI22, FMMS21, HK18, KSD^{+}23, KAAAA22, MWY19, NB24, PG19, SDM21, SMS^{+}24, SHBS19, SZFC18, SP22, WKW18, YHSW19, ZLZ^{+}24]. SDN-Based [AV23, AF21, KSD^{+}23, MWY19, SMS^{+}24, YHSW19]. SDN-Enabled [FMMS21]. SDN-NFV [SP22]. SDNs [YLK^{+}19]. SDR [MM21]. SDR-Based [MM21]. SDWSN [AHKC23]. SDWSN-Based [AHKC23]. Sea [WLG^{+}22]. Seagull [AK22]. Seamless [XDH24, ZAS^{+}24]. Search [AMR^{+}21, CP10, CCKH21, FPT20, HWZP18, KVST19, KH10, MDM21, SOR16, SZC19, VG20, XJ10, ZBN21, ZYLM17, ZVKK19, KH10]. Search-Order [CCKH21]. Searching [LWBM22, PVKG23, RAU^{+}24, XJ10]. Seasonal [KKBG23]. Secondary [HT14]. Secret [BLU^{+}15]. Sectional [BM19]. Sector [CTM22b, CT24a, CT24b, DV23, SLP^{+}22, SS14]. Secure [ADM19a, ATB^{+}24, ACS24, AZS23, Alb21, ADS21, AAJ^{+}22, ABBV23, BFY22, BL22, BMS24, BFG^{+}23, CLH^{+}23, EKG24, FNN21, FLM23, GPCP12, GMS^{+}20, HVS17, HS20, IZK^{+}23, KZ24, KAS^{+}22, KSW21, KHN^{+}11, KL23, Lee18, LT24a, MT17b, MJG24, NDS^{+}23, PZH^{+}18, PM24, PAM^{+}24, POC20, RC18, SS15, SMN22, SST23, SMS^{+}24, SXXW20, SL24a, WUH^{+}23, ZZZ^{+}23, POC20]. Secured [ASWAE23, LBA^{+}24, SPS^{+}21]. Securing [AOI^{+}23, AAM^{+}24, ALZ^{+}23, AF21, BSM22, CCC^{+}24, CPKK23, GPPB^{+}21, IPP24, IDR23, KDG24, MS22, PAM^{+}24, SA21]. Securities [FWLY23]. Security [AH22, AJ24, AS19, AIR^{+}18, AAN^{+}24, AAAKJ22, AS15, ASN22, BBSR24, BPG19, CVG^{+}22, CKK21, CMH10, DJC^{+}22, DGD^{+}19, DSB23, ESLB20, FAB^{+}12, FYWS16, FC22, FSG^{+}16, FP19, GMGF24, GWK17, GRA^{+}23, GKS10, GSB^{+}23, HLG21, HH20, HZQ24, HF14, HKH^{+}24, IIOW16, IZK^{+}20, KMW18, KSA^{+}21, KBF23, LZL^{+}17, LNG19, LLB24, MCS^{+}24, MAW^{+}23, MSC^{+}24, MMS22, MIK^{+}23, MGJ22, MSFM16, NLT^{+}23, PS16, PMGG21, PPN18, PORM^{+}23, PNdC^{+}23, PPP21, AJT^{+}23, RGB^{+}23, RYY10, RL24, RGMFM12, SBS18, SGG^{+}22, SE23, SOSR^{+}16, SOSC^{+}16, SSDS23, SAF14, Sta22, SG23, SWGL19, SAM^{+}18, SB23, TDW24, TH21, TSDG22, URH^{+}23, VFFHF19, WYL^{+}21, XKKL23, YHSW19, YOI19, YZC^{+}20, ZR15, dCdZCC24]. SEDIA [LG23]. Seed [ACS24, RZCL19]. SeedChain [ACS24]. Seeing [JMZ^{+}22]. Seek [KMK24]. Seek-and-Destroy [KMK24]. Seeking [PW22]. Segment [ELC^{+}19]. Segmentation [GNGL23, LSH21, LYY^{+}24, LOL23, MMD22, PTK^{+}23, SGW22, YLJ^{+}19, ZHW23]. Segmentation-Based [LSH21, ZHW23]. Seismic [MFO^{+}20]. Select [Cop14]. Selected [Bod22, NZZ13]. Selection [BAR^{+}21, CLC^{+}22, CZK^{+}17, CFP17, CFX20, EMHF19, FRR^{+}24, FRKS22, FKZ22, GDN^{+}23, GL13, HLC24, JKM^{+}23, JK24, MSAA22, MLL^{+}22, PDTM15, SJG18, SQK^{+}17, STK21, SXXH19, TYH23, TTS24, TRST24, XLM^{+}23, YSAY23, ZZWP19]. Selective [Fio10]. Self [ASWAE23, ADSAKAD22, CPP^{+}20, CTM21, CT24b, DYS15, DHG21, GZS^{+}24, JSAO^{+}23, KOJP13, Lai24, LMPT19, LZL^{+}23, LGY23, NE21, PLOT23, Sac19, SU17a, SHA23, SGG^{+}22, SW21a, WLL^{+}19, WYZ24, WSD22, XCSM18, YWS22, YW16, YWW^{+}23]. Self-Adapting [Sac19]. Self-Adaptive [XCSM18, YW16]. Self-Admitted [PLOT23]. Self-Assessment [DHG21]. Self-Attention [LMPT19, WYZ24, YWS22, YWW^{+}23]. Self-Determination [NE21]. Self-Evolving [LGY23]. Self-Healing [ADSAKAD22, JSAO^{+}23]. Self-Improved [SGG^{+}22]. Self-Interference [LZL^{+}23]. Self-Media [WLL^{+}19]. Self-Optimization [DYS15]. Self-Organize [SU17a]. Self-Organizing [CPP^{+}20, SHA23]. Self-Propagation [GZS^{+}24]. Self-Provisioning [KOJP13]. Self-Regulated [Lai24]. Self-Reported [SW21a]. Self-Sovereign [ASWAE23, CT24b]. Self-Supervised [GZS^{+}24, WSD22]. SEM [ZY20a]. Semantic [ALLR16, AFS^{+}23, CSC21, EKG24, FBDR23, FPT20, GNNZRCRG18, GNGL23, HL23, HWCH12, HG12, KBP22, LPPG22, LH21a, LDZ^{+}23, LCS12, MMD22, NSD^{+}22, PFLT12, PA13, PLG23, PNM^{+}20, RDD^{+}14, SOSK12, SJ20, Suf23, VRA^{+}18, WZ18, YAV^{+}21, ZGH13, BFFZ^{+}12, DBD^{+}14, Fra18, PFLT12, RKM^{+}22]. Semantic-Based [ALLR16]. Semantic-Driven [RDD^{+}14]. Semantically [KMS^{+}12, LG23]. Semantics [RRT20]. Semi [ABK22, GSO^{+}10, KHP^{+}22, KAAAA22, MLWS24, TTO^{+}19, YMSP24, YSAY23, ZPX24, ZJX18]. Semi-Automated [YMSP24]. Semi-Automatic [GSO^{+}10]. Semi-Autonomous [TTO^{+}19]. Semi-Centralized [ZPX24]. Semi-Persistent [ABK22]. Semi-Supervised [KHP^{+}22, KAAAA22, MLWS24]. Semi-Synchronous [YSAY23]. Semi-TCP [ZJX18]. Sending [HYL13]. Senior [ACK^{+}16]. Sense [CMS21, GSD^{+}17, OY19, TTO^{+}19]. Sensemaking [GPCP12, VI16]. Sensing [BYG22, EUDW23, UGJA18, Hel15, HMR^{+}21, KV23, Øie12, SSM^{+}24, TZS^{+}21, WWYQ19, WX24, XZ19, ZCZ17]. Sensitive [ATFMOU^{+}24, JNARV^{+}24, LLY24, PCM22, ZQL^{+}24]. Sensor [AAM^{+}24, AAAHA22, AAAIA24, ABB23, BZP^{+}21, BAR^{+}21, BF18, CH16, CCTC23, CMH10, CFX20, DPC^{+}24, DVSBV^{+}24, IDR23, JBS^{+}23, JCPV22, LZ23, MMA10, MCM^{+}10, MGM^{+}23, MCPLBH^{+}24, MP13, NMT23, SJA^{+}23, SLW17, TAMA22, TAHO23, TSDG22, WCF^{+}16, WCM19, WXL^{+}23, WLZ18, XSX18, YW16, YLH^{+}17, ZLS19]. Sensorial [BKSS19]. Sensors [ABLR23, FHL^{+}24, DMP^{+}23, MKR22, QNJ24, TDM^{+}24, TEE10, WX24]. Sensory [DMP^{+}23, FRE^{+}22, PHGZ20]. Sentence [BOB^{+}20]. Sentiment [AWBS23, CWC^{+}22, FLLG24, JLW19, KDEA^{+}23, LLZL19, LLJ22, MKK^{+}22, PPN18, SH21, TPKA20, THL21, WCZ23, YWS22, YYW20, YWW^{+}23, ZL19a, ZZ19]. Sentiment-Aware [SH21]. Sentimental [JXA19]. Sentinel [GMLS22]. Sentinel-2 [GMLS22]. SEO [GKK^{+}19, KLKA19]. Separate [AGN21]. Separated [CPP^{+}20]. Separation [ALH24, SWGL19, SSX20]. Separation-of-Duty [SWGL19, SSX20]. Sequence [ARAAA19, LC23, OAH^{+}24, QQM24]. Sequence-to-Sequence [LC23]. Sequential [DYLZ21, MT23]. Series [CR23b, KLKP19, KPKM23, PLA^{+}24, PW21, TOLG21, TRLR24, WX20, WLG^{+}22, WBS^{+}24]. Serious [MVAHBL^{+}23]. Serpent [MS22]. Server [PXC18, TRST24, ZZWP19]. Serverless [FR24, YA24]. Servers [HWZP18, SGLZ20]. Service [AAM^{+}24, AP22, AMG22, ABK22, ARC^{+}19, AAD^{+}12, ASL17, ASG24, BB17, BAS24, CMPT24, CCD^{+}21, CL16, CC22a, CH16, CMH10, Ciu19, CSC21, DCHM16, DGF23, ESLB20, FR24, FHZ^{+}24, FFM^{+}23, GL13, GMMI16, GGK18, GKG19, IKP^{+}13, IUIL23, Kab23, KTS18, KMS^{+}12, KG18, KA21, KKT20, LBG^{+}14, LPVM24, LW11, LW21, MHS23, MP16, MMN19, MKvD11, MKSV^{+}19, MSC^{+}21, NSV17, NPM^{+}18, PDTM15, PZH^{+}18, PSTZ23, PNM^{+}20, QLL^{+}21, RRMI22, RF21, SB22, SYB^{+}18, SMG13, SSH20, SV24, SPT23, SMS^{+}24, SWJ20, SGDT19, STS23, SWY^{+}18, TBT^{+}23, TKH^{+}11, UB18, dSVJdALRfS21, VFFHF19, WWB^{+}20, WYW^{+}22, GVY^{+}23]. Service-Oriented [LW11, MP16]. Services [AFS^{+}22, BOK^{+}21, BDL22, CHH^{+}16, CFP19, CR11, DTTK19, FC19, Fer12, FAB^{+}12, FGBMG19, Fra18, GTM22, HGL^{+}23, IFF21, JRP21, JAS21, KMS^{+}12, KCY22, KPC19, LW11, MKSV^{+}19, NMH^{+}21, NSD^{+}22, PCM24, PSM^{+}18, RAK^{+}20, SIZD12, SWZ18, Tri14, VP22, WZLL19, Yan17, YKC19, ZKV12, ZZWP19, ZSS^{+}22, RGdSK22]. Serving [LG20]. Session [LWH24]. Session-Aware [LWH24]. Set [ACAP20, GM19, Mah17, ZHZ^{+}19]. Set-Valued [ACAP20, GM19]. Sets [HCXH16]. Settings [AAR^{+}19, BD24, BLW^{+}17, BM18, SK24]. Setup [KP15, SOR16]. Sextortion [PB23]. Shadow [ZMDCEMI22]. Shadowing [KP19]. Shallow [YZZ^{+}20]. SHAP [GDN^{+}23, MPK^{+}23]. Shape [SSR^{+}21]. Shapelet [KVST19]. Shaping [EWGK24, HGF21, Khr20, KA21]. Shapley [GDN^{+}23]. Shared [EAAA^{+}23, MAW^{+}23, WC23, YJSL18]. Shares [BLU^{+}15]. Sharing [AGC18, CCCA22, CGLR19, CDD^{+}21, FM16, FMNS11, HWCL17, JUN^{+}24, LZY^{+}24, LT24a, MCPLBH^{+}24, PYC21, QA24, SAR24, ZAV14]. Sheep [LLSS24]. SHFL [AAJ^{+}22]. Shift [ARAAA19]. Shifts [FAS^{+}23]. Ship [Ngu19, ZZSX19]. Ships [CTP^{+}22]. Shop [SU17b]. Shops [PBB20]. Short [BKL^{+}22, CLH^{+}23, CDLY24, FR15, Fra23, GPA23, HSL^{+}23, KKBG23, LGY23, PCM24, PLA^{+}21, SSD24, WCZ23, YZCR24, ZXJ22]. Short-Read [CDLY24]. Short-Term [BKL^{+}22, HSL^{+}23, KKBG23, LGY23, PLA^{+}21, SSD24, YZCR24, ZXJ22]. Short-Text [WCZ23]. Shortest [TYH23]. Shot [HY22, LDZ^{+}23]. Should [MGFF14, PdRB19]. Show [SLP^{+}22]. Siamese [ZJJS^{+}23]. Side [DSMM21, LNT^{+}23, LNT^{+}24, PXC18, VBB24a, VBB24b, dSVJdALRfS21, YC22]. Side-Channel [DSMM21, LNT^{+}23, LNT^{+}24]. Sidecar [BCM^{+}19b]. Sidelink [BCM^{+}19a]. SIEM [KPCS13]. Sighted [YLX^{+}24]. Sign [DVSBV^{+}24, JDH21, STV^{+}22]. Signage [CTP^{+}22, TOLG21]. Signal [LLW^{+}17]. Signals [DAMAS^{+}19, JPVC21, KLG^{+}24, OAI^{+}22].

Signature [AMC^{+}24, FR15, IZK^{+}23, JUH^{+}23]. Signatures [DMO^{+}19]. Significance [AWSA23]. Signs [Chu23, You23]. Silence [ZSZ18]. Silicon [KNN^{+}24]. Similarity [DH18, OY19, YZ23]. Simple [AK19, BdBC23, BR21, Gro17, JK20, KWI^{+}19, LBBPM21]. Simplifying [ER10]. Simulated [JHL^{+}18, TL24a]. Simulating [AMAE22, BSEL20, BSE^{+}21, GDLG18, SEB^{+}19]. Simulation [AIP^{+}24, AVA^{+}22, BdBCCG21, BMGI21, GLH^{+}22, KHRG19, LZC21, MDT20, VC23]. Simulation-Based [AIP^{+}24, BMGI21, KHRG19]. Simulations [SUF^{+}22]. Simulator [BSE^{+}21, FPGO24]. Simulators [ALDS19, ZDP^{+}22]. Sina [AHD21]. Single [ARM20, BTB22, DDZ18, DNZW22, HY22, SC23, TBT^{+}23]. Single-Mode [DDZ18]. Single-Rate [DNZW22]. Single-Shot [HY22]. Single-Stage [SC23]. Singular [KHC^{+}23]. Sink [BZNB23, TAHO23]. SIoT [ZBN21]. Site [AIPV21, GPL24, VL20]. Sites [AP21, Smi13]. Situ [GRRZ18]. Situation [CSA^{+}24, YBMK21]. Situational [LCV^{+}17]. Situations [LLR19, UANU20, ZZL^{+}23]. Six [SOK24]. Six-Dimensional [SOK24]. Size [EMHF19, KK21, LH21c, SGMBK24]. Sized [Kha24, SRS23, SDM22]. Skeleton [COL22]. Skeleton-Based [COL22]. Skies [AJ24]. Skills [VBWW22]. Slab [FLSL24]. SLAM [NLD^{+}24]. Sleep [CMR24, DVSBV^{+}24]. Slice [AlQ23, MAW^{+}23]. Sliceable [NAV^{+}23]. Slicing [AMG22, DJC^{+}22, HGL^{+}23, KLG^{+}24, LTW21]. Slotted [YLYL22]. Slotted-ALOHA [YLYL22]. Slovak [MKK^{+}22]. Slum [PR11]. Slum-Upgrading [PR11]. Small [GMRRRQ^{+}16, JHC^{+}20, Kha24, RRT^{+}23, SGMBK24, VBB^{+}24c]. Small-Scale [GMRRRQ^{+}16, VBB^{+}24c]. Small-Size [SGMBK24]. Smart [ABCC22, AHKC23, AMAA^{+}21, ANHMS24, AEM^{+}23, AEA^{+}23, ADK^{+}24, ARC^{+}19, AH21, AAJ^{+}22, AMY24, AMI23, BMS20, BBSR24, BCSB24, BSM22, BHSAY24, BCZ^{+}23, BG24, BZDP19, BFVM14, CBR^{+}24, CVG^{+}22, CBKL22, CC22a, CGBF24, DTR22, DAMAS^{+}19, DT24, EHJ17, ENT21, FM16, FM20a, Fra24, GAGB19, GLD^{+}18, GHA^{+}21, Giu18b, GKFV24, GBKN23, GSL20b, GNNZRCRG18, GTWN24, GPM21, HXHP17, HGL^{+}23, IMR^{+}18, UMS24, JEJ^{+}24, JNARV^{+}24, KZRH22, KSD^{+}23, KVWC24, KAM^{+}19, KPP^{+}20b, KPP^{+}20a, KCK^{+}19, KK20, KAL^{+}19, KL23, KSG^{+}24b, LNB12, LMSC19, Let16, LF24, LMY21, LT24b, LHL24a, LG23, MW24, Mal20, MDB22, MAW^{+}19, Mar17, MIK^{+}23, McK20, MRM^{+}24, MLL^{+}22, MHQ^{+}23, NBL^{+}24, OGR^{+}20, Oza23, PG23, PPGC16, PCC^{+}18, PSC19, PPN18, PPP^{+}23, PTM20, Pir22, PLG23, PNdC^{+}23, QQX^{+}24, RGB^{+}24, RHV17, RSL^{+}24, BNK^{+}23, SA21, SSO^{+}19]. Smart [SNKG20, Str22, SZC19, TDM^{+}24, TL24b, TL19, URH^{+}23, VL20, VBWW22, VBB^{+}24c, WZL^{+}20, WHW22, WS21, XSSA24, YBMK21, ZKKK19, ZL22, ZLS19, ZTBD20]. Smart-City [KSD^{+}23]. Smart-Home-Based [ABCC22]. SmartDED [LT24b]. Smarter [Mal20]. SmartLab [TPJ21]. Smartness [McK20]. Smartphone [CW23, KKL^{+}20, KTAH^{+}23, SRM20, BRP^{+}13, UDF^{+}19]. Smartphone-Based [BRP^{+}13]. Smartphones [KKTS22]. Smartwatch [BMB23c]. Smartwatches [AA24a]. Smells [DSB23]. SMEs [DPR^{+}21, SOSC^{+}16]. SMOTE [ASF^{+}23]. SMS [FTAA21, GMA20]. SMYOLO [ZMZ^{+}22]. Snack [KWI^{+}19]. SNR [OGB^{+}24, SZL21]. SNS [THT18]. SNS24 [YGQ^{+}22]. SOAP [KMS^{+}12]. SoC [MLPH23]. Social [AKYP23, AWBS23, AS19, AP21, AAAR24, AIR^{+}18, ACK^{+}16, AS24, BdBCCG21, BCT21, BOKC19, CLM^{+}21, CDM23, CFP17, CCM^{+}17, Dav12, DIG20, FC19, Far12, FRKS22, FZ15, FBOC15, GJ15, GPCP12, GNV21, GPA23, Gog12, GMB^{+}21, GFL21, HL23, HT14, HLZ23, HCW20, HRGQHOA21, ITZ20, Jur12, KG18, KPC19, KDKG22, Let16, LZC21, MP20, MMH16, MKDG20, MS16, MGFF14, MSX^{+}21, MOAI17, MCCOCE24, Mot24, NWG20, PB23, PST^{+}21, PFLT12, Pir22, PLG23, RZCL19, RDD^{+}14, RSL^{+}24, SK19, SRMJ21, SAL21, Smi13, ST20, SG21b, THT18, TZS^{+}21, TL17, TLV23, TKT^{+}18, Ves18, VKK^{+}19, VKKG22, VCB14, WLH^{+}17, WZC19, WZLW19, WKF^{+}21, YKC19, YPLZ19, ZX24, Zha19, ZSRC23, ZSF20, CB16, WYW^{+}22]. Social- [RDD^{+}14]. Social-Aware [CFP17]. SocialBot [KHC^{+}24]. Socialism [HW16]. Socially [PS17]. Societal [MZ19]. Societies [Let16, LMY21]. Society [ACK^{+}16, Tsi14, MAP23]. Socioeconomic [HGF21, HRKL21]. Soft [CFX20, BBBHA24, LH21a]. Software [AA19b, AAA^{+}19, AA21c, Ant24, BCFP21, BM14, DDV22, EW19, FML^{+}22, GLZL24, GTQ^{+}22, GGW18, IZK^{+}20, LZL^{+}23, LXL^{+}19, MZS19, MTKS24, MP16, MSP21, NAFM^{+}24, Pap20, QAQA21, RAJJ23, RKY20, SN23a, SMN22, STS23, SKT^{+}19, SAM^{+}18, Tos23, TRC^{+}21, VL20]. Software-Bus-Toolchain [Ant24]. Software-Defined [BM14, FML^{+}22, GLZL24, GTQ^{+}22, LXL^{+}19, MZS19, MTKS24, MP16, NAFM^{+}24, RAJJ23, STS23, SAM^{+}18, TRC^{+}21, VL20]. Softwarized [EVCL21, Mou24, RS17]. SoK [HM24]. Solar [ADK^{+}24, HS20]. Solar-Powered [HS20]. Solid [BZMB23]. Solid-State [BZMB23]. SOLIOT [BM20]. Solution [CC22b, FBDR23, GVV^{+}23, KZRH22, MCA23, SU17b, SPS^{+}21, TCH23, VI16]. Solutions [AM23, AAN^{+}24, GPCP12, KSA^{+}21, KBF^{+}22, KSE^{+}20, MIK^{+}23, NBT22, PLG23, RSG24, ZMKR22]. Solve [CBRS21]. Solver [Yan19]. Solving [GDLG18]. SOM [BCNS20]. Some [BZDP19]. Sort [Sta23]. Sorting [AIP^{+}24]. Sound [MDS^{+}24, MTF^{+}21, YST24]. Sound-Based [MDS^{+}24]. Source [ABH24, CTP^{+}22, CSZ^{+}24, DTR22, FSY^{+}22, GMD^{+}22, GMGF24, GDS23, HPF^{+}24, KRSF22, Li18, PLYD20, PTM20, PLOT23, PFL^{+}12, SYLL21, SB23, TRB22, YFY24, YWZ^{+}24, ZZY19, vSR24]. Sourced [Lal14, SVGGO24]. Sources [RBSMRN21]. Sourcing [NPB12]. South [PH20]. Sovereign [ASWAE23, CTM21, CT24b]. Space [FHZ^{+}24, GPMO20, Lee18, LLL21, LGCM24, MWY19, TL24a, Whe09]. Spaceborne [SWC^{+}23]. Spaces [LMSC19, MS20, Whe09, YWLY22]. Spacing [AAR^{+}19]. Spain [VLMP21]. Spam [GMA20]. Spammers [AA19a]. Spanish [AdMTJ20, Hor19, LL20c, RRAGMMGG20]. Spark [JOL22]. SPARQL [RAA^{+}18]. Sparse [KES22, SQK^{+}17]. Sparse-to-Dense [KES22]. Spatial [BR21, DQ18, FMNS11, JKM^{+}23, LMH19, MHL^{+}23, NAV^{+}23, PG14, SS14, WX20, WYZ24, YCHG18, ZWZ17, PL22a]. Spatial-Temporal [MHL^{+}23, WX20]. Spatio [CFG20, EvdMB20, SKA^{+}23, dCMA21]. Spatio-Temporal [CFG20, EvdMB20, SKA^{+}23, dCMA21]. Spatiotemporal [AES21]. SPDepth [GZS^{+}24]. Special [Bod22, CB16, EMMP22, Far12, Fer12, Gra13, LMK18, Mac22, PG23, PSC19, Sta22, SG23, Tos23, ZMKR22]. Specialists [JP13]. Species [GNK^{+}10]. Specific [GPCP12, KS19, LY19, MSN13]. Specifications [MUS11]. Spectral [GMLS22, GAMMPCRÁ24, OGB^{+}24]. Spectrogram [LF23]. Spectrum [CZK^{+}17, CHMZ21, FM16, JUN^{+}24, JRP21, SSM^{+}24, Wan17, WWYQ19, ZCZ17, dCdZCC24].

Speech [ALH24, HHL24, ITH^{+}21, MCR^{+}23, STV^{+}22, SSD24, VVK^{+}21]. Speech-to-Text [MCR^{+}23]. Speed [Chu23, DWY19b, GAGB19, GCEOBRT23, HM20, KKS^{+}19, Tra18]. Spending [AMEF21]. Sphere [PE13]. Spherical [Mat20]. SPIHT [SS17]. Spiking [BZP^{+}21, NB24, Vem22]. Spillover [RZQS18]. Split [KMK24, MHK13, MM21, XSSA24]. Split-and-Destroy [KMK24]. Spoofing [MMS22, RRT^{+}23]. Sports [PRS22]. Spot [OS23]. Spots [ALST20, CDM23]. Spotting [BC23, DKH24]. Spread [Bad23, DGMP21, HGF21, dCdZCC24]. Spreading [CLM^{+}21, HD19, RFSOHMSB21]. SQL [AV23, GHL^{+}23]. SRMCC [DNZW22]. SRUP [POC20]. SRv6 [LPVM24, LWW^{+}24]. SRv6-Based [LPVM24]. SSI [CTM22b, CT24a, PBK22]. SSQLi [GHL^{+}23]. SSS [AMSM23]. Stack [AFS^{+}23, JLH^{+}13, dVBA23]. Stacked [MA23, ZXMB22]. Stacking [ASCM22, WLL^{+}19]. Stacking-Based [WLL^{+}19]. Stage [AMY24, SC23, TBT^{+}23]. Staging [CYO^{+}17]. Stake [AMEF21]. Stalking [GJ15]. Stance [AAARZ24]. Standalone [KDJ24]. Standard [SWD^{+}18]. Standardization [HAL18]. Standardized [JP21]. Standards [DPC^{+}24, Dua21, FAB^{+}12, KKTK20, MCA23, dCdG20]. Star [SOB^{+}19]. Start [LLLD21]. Start-Up [LLLD21]. Starting [Tre22]. State [AZH22, BZMB23, BM23, CEP24, CCM^{+}17, DeD16, HNK^{+}21, KSA^{+}21, LZLW22, MHQ^{+}23, PG23, PKF24, PSGM22, QQX^{+}24, Ray23, RGdSK22, SJG18, VCRCSV^{+}21, XLW^{+}24]. State-of-the-Art [CEP24, HNK^{+}21, KSA^{+}21, Ray23, XLW^{+}24]. Statehood [Cal24]. Stateless [GLZL24]. Statements [DCD24, MMT24, TZS^{+}21]. States [GPLK22, SG21b, WLH^{+}17]. Static [BdCRM10, GSP15, IMLL24, SNS22].

Station

[CPKK23, ENT21, TSHO24, YPXH19]. Stationary [MCPLBH^{+}24]. Stations [BZDP19]. Statistical [CMW18, MSZ23, Smi15]. Statistics [SJ12]. Status [CBD^{+}22, GGT^{+}23, HE20, SCM12]. Stay [PdRB19, TRHU23]. Stealth [HCXH16]. Steam [AOF24]. Steganalysis [WLYL20]. Steganographic [WK24]. Steganography [PTZM22, SKHK23, WYL18, WCYL20, WLYL20]. StegNet [WYL18]. Steps [AML22, For14, LCS12]. Stereotypes [Söd13]. Stickiness [HL19]. Sticky [FN23]. Stiffness [DGMP21]. Still [QA24]. Stimulus [CB24]. Stochastic [MSZ23, ZQCC16]. Stock [CLC^{+}22, LPA^{+}22, TZS^{+}21]. Stop [You23]. Stopovers [ALST20]. Stopping [FTHY21]. Stops [KAM^{+}19]. Storage [BMS24, BZMB23, DH22, DPR^{+}21, HOV23, HDT22, LZS18, LL23b, MMN19, OF23, PV20, PV22, SHHM21, SN23b, SPS^{+}21, SMK^{+}22, TTN24, XSM22, YJSL18, YL21]. Store [FRE^{+}22, KFP10]. Store-Carry [KFP10]. Stores [OF23]. Storm [SPM^{+}22]. Storms [DGD^{+}19]. Storytelling [BHH12]. Stranger [DCPG21]. Strategic [FZ15, LZ11]. Strategies [AK19, AR23, BdCRM10, BM19, BBM20, EAKM22, IHCSG^{+}24, KSD^{+}23, KFP10, MSC^{+}24, MP20, MCCOCE24, RRP17, URH^{+}23, VM24a, XJ10]. Strategy [BMJ^{+}22, CPRS20, DXL^{+}22, FAT19, GW15, ISLARP^{+}22, JHBW24, KYT^{+}23, KHM20, LZS18, LYP^{+}19, LWZL22, MBH22, PPP21, RRMI22, RR24a, SGLZ20, SYGY21, SIZD12, TTKZ19, VBWW22, XCSM18, XNZ23a, ZBN21, ZWZ19, ZZY19]. Strategy-Based [MBH22]. Stream [GL12, LQY^{+}19, MGM^{+}23, SBRZ24, SSR^{+}21, VdHH^{+}19]. Streaming [HF22, LNA17, LXG^{+}23, LGY23, LIM23, MSC^{+}21, NNT22, SA24, TCH23, VdHH^{+}19]. Streams [CDM23, LWZL22, Pec18, VPCE20]. Street [MSR^{+}14, NZZ12]. Strengthen [GPM21]. Strengthening [WZS^{+}22]. Stress [FC19, KPC19, LRD23, VGS^{+}19]. Stress-Outcome [FC19, KPC19]. Stressor [FC19, KPC19]. String [RFD23]. Stroke [NS12]. Strong [APJ24]. Structural [BZP^{+}21]. Structure [AAAAK24, Jeo20, KK20, NZU20, OKH13, TCG19, WC20, ZX24, ZY20b]. Structured [HS18a, Kra22, VI16]. Structures [GKW^{+}10, Sac19]. Structuring [BML19]. Student [Alh22, AGN21, BCSB24, BBK^{+}20, LLZ21, NE21, PFT^{+}24, SSOÁ^{+}16, TRLR24, ZH21]. Student- [TRLR24]. Students [AP21, BM19, Del21, ER21, FOJE19, Gao21, Giu18a, HT14, HOPGRV21, LL20b, VCRCSV^{+}21, ZZZ^{+}22, dSPR^{+}22]. Studies [Bod22, ERBL12, HDK^{+}12, MMH16, SLY^{+}12, Uto13, YMJ20]. Study [AJ24, AMAA^{+}21, AGO23a, AP21, AMAE22, ADAq^{+}22, ASM^{+}24, AAC21, AAN^{+}24, AYRA21, ARC^{+}19, BMB23c, BM19, CRMT23, CFV19, DVSBV^{+}24, DMS^{+}22, DTL^{+}21, FCR23, FRE^{+}22, FPGO24, FZ15, FGBMG19, Gar12, GLRM19, dPCGMC20, GMMI16, GRA21, HSM19, HUM^{+}21, Hua15, Jal19, JHC^{+}20, KYG22, KLKA19, KK21, KK23a, KG18, KDKG22, Kra22, LXG^{+}23, MPG20, MS20, MFH22, MK20, MM21, MTHN22, NLD^{+}24, NB12, OLMBRCRC23, OBK23, PDI21, PS17, PHG^{+}20, PHGZ20, AJT^{+}23, RCGSL19, RS22, SU17b, SRS23, Sem11, STG^{+}20, SRM20, SVFV23, SJ12, hSgLH17, TBT^{+}23, TSD24, THL21, VLMP21, VM24a, VDE^{+}20, WC17, WL20, WLJM21, WCZ23, ZXJ22, ZLOW24, dSPR^{+}22]. Studying [ZJX18]. Stuxnet [BPBF12, Den12]. Style [FOJE19, Tas10]. Styles [YZ23]. Sub [RD18, APAR22]. Sub-Network [RD18]. Sub-Saharan [APAR22]. sub-THz [OGB^{+}24]. Subjected [NPMPM23].

Subjective [AWSA23, BS17, HCW20]. Subjectivity [SPC22]. SubNet [HHL24]. Substitute [RZCL19]. Substrate [SVGGO24]. Substring [SOR16]. Subsystems [AKM22]. Success [MP20]. Successful [Hor19]. SUDC [LWW^{+}24]. Suitability [EvdMB20]. Suite [TAAK21]. Summarization [BTB22, GPA23, LC23]. Summarizer [FTAA21]. Super [HWCH12, dMLMS22]. Super-Peer [HWCH12]. Super-Resolution [dMLMS22]. Supercomputer [Kra20]. Supercomputing [DCHM16]. Supermarket [AIP^{+}24]. Supervised [GGD23, GFPVLR^{+}23, GZS^{+}24, HXHP17, HOV23, KHP^{+}22, KAAAA22, LTTS21, MLWS24, TMP22, WSD22]. Supervision [HP21, LT24b]. Supplementary [RL24]. Supply [ACS24, CTM21, GSL20b, MCS^{+}24, PR20, RKT19, WADL^{+}24, YOI19, Yiu21a, Yiu21b]. Support [ASK23, BNJ24, CBKL22, CTM21, Del23, FCG22, HCW20, IKP^{+}13, LA20, LSM^{+}12, MKHR21, NLPV12, Ngu19, NSD^{+}22, RKS^{+}21, RAK^{+}20, SGR^{+}13, SCC^{+}23, SPB^{+}11, VI16, VKV^{+}22, XDH24, YWW^{+}19]. Supported [Del21, GSO^{+}10, RRdBSS22]. Supporting [FFP^{+}22, FR16, MP13, PFL^{+}12, PH16, SG12, SOR16, VP22]. Supports [WKPC18]. Surface [DAMAS^{+}19, IHCSG^{+}24, JMZ^{+}22, SOK24, SF22, WLG^{+}22, ZWPL19, ZX24]. Surface-Aided [SF22]. Surface-Assisted [IHCSG^{+}24, SOK24]. Surfaces [Mat20]. Surge [WSD22]. Surgery [ANAAM21]. Surveillance [LZS18, LS21, SYZ19, USS19]. Survey [AAAAM^{+}21, Al 20, Ala20, AMCI17, AR23, AKM22, AIPV21, BCM^{+}19a, BBM20, BSEL20, BB17, CPM^{+}19, CBM^{+}20, CMH10, CDBF19, CBB17, CFX20, DJC^{+}22, DLK10, DT24, Dua21, EABE19, FWLY23, FK21, FAB^{+}12, FM20b, FK24, GZC21, GGT^{+}23, GKS10, GSO^{+}10, HH20, HQC^{+}24, HZQ24, IZK^{+}20, JCPV22, KDG24, KKG20, KSA^{+}21, KBF^{+}22, LS21, MSC^{+}24, MDT20, MKK17b, MHSH16, MTM22, MHL^{+}23, ORK^{+}19, OPGH23, PH20, PCVH24, PNS23, PST^{+}21, PCM24, RM23, SK19, STK21, SRM20, TLM^{+}20, TEE10, VMGT22, VL18b, VSEH^{+}21, WML24, WRHH23, XZWL24, YZC^{+}20]. Surveying [LMSC19, Mot24]. Surveys [GdSdC23]. Survival [PDT23]. Susceptibility [FLLC24, NJH^{+}24]. Sustainability [KZ24, KPC^{+}23, MMN19]. Sustainable [AAKS23, CT24a, Gal24, NSJGN21, OFK19, PPMMS19, URH^{+}23, WADL^{+}24]. SVD [ZWWW17, ZWZ17]. Swapping [LWZ^{+}21]. Swarm [ADAq^{+}22, ABS20, CD23, FK24, GB23, OAI^{+}22, SHA23, Tra18]. Swarm-Based [ABS20]. Swarms [KL23, LLBB18]. SWIM [LW19, WZLL19, WC20]. SWIPT [LLW^{+}24]. Switches [IZK^{+}20]. Switching [ENT21, GRG21, THV^{+}22]. SWOT [THL21]. Sybil [SFEK18]. Symbolic [ASLG22]. Symmetric [SDDB24]. Symptom [RRLT20]. Symptomizing [MAW^{+}23]. Symptoms [BBM20]. Synchronization [BC19, CC22b]. Synchronizing [CC23]. Synchronous [LWW^{+}24, YSAY23]. Synergistic [LF24]. Synergy [EAAA^{+}23, JMJ^{+}23]. Synonyms [SPGS23]. Syntax [GDS23]. Synthesis [IMLL24, LJWX23]. Synthesizer [KLKP19]. Synthetic [LTTS21, SZT23, UJ22]. Syscall [VBB24a, VBB24b]. System [ASK23, ATB^{+}24, AEM^{+}23, AA24a, AA21c, AMZ^{+}20, APAR22, AGO^{+}23b, ALDS19, ADG^{+}19, BdBCCG21, BCSB24, BRM23, BCM^{+}23, BSE^{+}21, BR23, BMS24, CTP^{+}22, COL21, CTM22b, CT24a, CGM^{+}19, CNC^{+}21, DTR22, DHDAu22, DRK^{+}22, DAMAS^{+}19, EARP^{+}24, FWL22, FMA^{+}20, FBOC15, FLSL24, GNK^{+}10, GMMI16, HZWJ21, HWCL17, HMR^{+}21, IAG^{+}21, JEJ^{+}24, JLZ18, KGRP21, KDEA^{+}23, KWI^{+}19, KE22, KK20, KP15, KLG^{+}24, KTD24, LB24, LLHW20, LZ23, LWY^{+}15, LXL^{+}20, LOL23, LLW21, LHW18, LXQ24, LHW24, LT24b, LK20, LSM^{+}12, MDS^{+}24, MDB22, MSN13, MRM^{+}24, MFO^{+}20, dSMdO22, NB24, Ngu19, NBL^{+}24, PPR^{+}20, PPMMS19, PLA17, PPP^{+}23, QNJ24, RIS^{+}17, RRT20, RHV17, SSOÁ^{+}16, SMB23, SCA^{+}19, SW21b, SC23, SOB^{+}19, SZL21, SOA^{+}20, SHN^{+}19, SH20, SLY^{+}12, TADS20, Tas10, TSZ^{+}18, TPD^{+}20, UUT^{+}19, UANU20, VL20, VCJAMN22, VSE21, Wan17, WL21, WXL^{+}23, WC23]. System [WS21, XSSA24, YZP22, YWLY22, YBO^{+}23, ZHZ^{+}19, ZJ18, ZCZ17, ZSZ18, ZGL20, ZSRC23, ZYWS24, ZZZ^{+}22]. System-Level [CGM^{+}19]. Systematic [AZAV22, AN22, AAA^{+}20, CR23a, DIG20, ESBE23, EWGK24, GGW18, KT22, KIG23, LH21b, MKDG20, MR23b, NPWC24, OO19, OPGH23, PAK^{+}23, PJMM22, RNFS20, dSRRdOT21, SMVP21, SKKS22, SKFM24, UW21, VLBRMP21, VDE^{+}20, YMSP24, vSR24]. Systemic [DGHH^{+}20]. Systems [ATE24, AZ19, AA23a, AM21, AR23, AA19b, AAA^{+}19, AA23b, ASD^{+}22, ARM20, AAE^{+}22, AAJ^{+}22, AMVM22, BARL24, BRC24, CSIS19, CCC^{+}24, Cap12, Cap23, CRMT23, CPP^{+}20, CSDAM^{+}23, CH23, CFGD20, DPH17, DPH19, DN20, DDZ18, DSM21, DT24, FM16, FAT19, GPCP12, GS23b, GW13, GL19, GTQ^{+}22, Giu18b, GGT^{+}23, GMH18, GSL20b, Gro22, HCXH16, HLG21, HM20, HBFJ20, HAL18, HJA18, IHCSG^{+}24, IPP24, JWD^{+}24, JSAO^{+}23, KS24, KHAA21, KM24, KP19, KGK^{+}24, KKG20, KPRP24, KK21, KBHP24, KTAH^{+}23, KDKG22, KKEV17, KPCS13, KBBH21, KID^{+}15, KSG^{+}24b, LH19, LDM22, LZS18, LSZ21, LL23b, LF24, LYP^{+}19, LLSS24, MSC19, MT17a, MJG24, MAW^{+}19, MBH22, MEJA23, MOAI17, Min20, MS23, MKM^{+}19, MKHR21, MHL^{+}23, MHQ^{+}23, NMT23, NPWC24, OCCB20, OFK19, OGB^{+}24, Oza23]. Systems [PMP22a, PC24, PVM22, PSC19, PORM^{+}23, PA13, PGSL20, PH16, PSGM22, QQX^{+}24, RYY10, RR24b, SBS18, SMVP21, SGMMRG21, SOSR^{+}16, SAR24, SRS^{+}24, STG^{+}20, SNH^{+}19, SPS^{+}21, SN23c, TMP22, TLK^{+}20, TKT^{+}18, VP22, VDSK23, WYS17, WJ21, WSF23, WUH^{+}23, YG20, Yiu21a, YL21]. Systems-Aided [DN20].

t [TRLR24, KC19]. T-Move [KC19]. T5 [FTAA21]. T5-Based [FTAA21]. Table [IZK^{+}20, LSJ19]. Tables [LLCC24]. Tablet [BAR^{+}21]. Tabu [LSJ19]. Tacit [PKCF21]. Tackling [TPJ21]. Tactic [BMB^{+}23b]. Tactical [IGL^{+}23]. Tactile [AMK^{+}21]. Tag [AM21, LPM10]. Tags [ASF^{+}23]. Tailed [SVFV23]. Taiwan [CMW18, Hua15, LL20b]. Take [SPM^{+}22]. Taking [SS14, SG21a, ZLW21]. Tales [XJ10]. TalkRoBots [AMVM22]. TAM [JGV^{+}21]. Tamper [HDT22, SW21b]. Tamper-Proof [HDT22, SW21b]. Tangle [GPPB^{+}21]. Tanks [MCPLBH^{+}24]. TAP [HHL24]. Target [AWBS23, CHMZ21, HP16, SWC^{+}23, ZMZ^{+}22]. Target-Oriented [AWBS23]. Targeted [MT17b]. Targets [WL20]. Task [ARAAA19, ABH24, CD23, CSC21, GDS23, IUK^{+}23, JHBW24, LW19, LJR23, LJR24, LLL21, LLSS24, MC24, OS23, PAK^{+}23, QHX^{+}24, QWR18, TTKZ19, TCYZ22, WMS23, XQLZ19, ZZZL23a, ZJJS^{+}23]. Task-Aware [ZJJS^{+}23]. Task-Based [OS23]. Task-Dependent [LJR23, LJR24]. Task-Oriented [LLL21, QHX^{+}24]. Tasks [BR21, EMMP22, MWL20, WX22, YAV^{+}21]. tattles [UDF^{+}19]. Tax [LRdLS19, MTHN22]. Taxonomic [KBP22]. Taxonomical [GSB^{+}23]. Taxonomy [AA23b, IP16, MHS23]. Taxpayers [LRdLS19]. TCN [CLC^{+}22]. TCN-Deep [CLC^{+}22]. TCP [CISM22, AHH^{+}23, OLCMdA20, PW22, RRP17, WWD21, ZJX18]. TDLearning [LHL24a]. TDOA [CW23]. TE [BdCRM10]. Teacher [KID^{+}15, LL20c, MRRMC21, PEN23]. Teachers [dPCGMC20, ZKV12]. Teaching [AARTSBAE24, BPLBM21, Del21, GDRCDBLÁ23, HAL^{+}22, JHC^{+}20, JP21, LF23, LBBPM21, PLSF14, RMLAZOPC21, SSB21, hSgLH17, TMTMB20, VBWW22, VCJAMN22]. Team [MPK^{+}23]. Team-Level [MPK^{+}23]. Teamwork [AM24, AGB22]. Technical [AH22, Ala20, AAAKJ22, Oza23, PLOT23]. Technique [ADAq^{+}22, GKM24, GS23a, UGJA18, KUBS22, LWHR14, MKKS18, PMP22a, ZHZ^{+}19]. Techniques [AA23a, AMAE22, ABBV23, Chu23, EKKS24, GPA23, GKFV24, GSO^{+}10, GMGK20, HM24, HOV23, HN18, KM24, LMSC19, LOL21, LTP24, MLWS24, MG10, MKSK23, MKSJ22, MSL^{+}23, MF20, NMT23, PAK^{+}23, Pri23, RGP22, SLP^{+}22, TDM^{+}24, TTK24, Wan23, WCZ23, WSN24, XLW^{+}24, ZCG^{+}24]. Techno [KAL^{+}19, LGMZ19]. Techno-Economic [KAL^{+}19]. Techno-Regulation-Experimenting [LGMZ19]. Technological [CBD^{+}22, GL19, LL20c, SN23c]. Technologies [AEA^{+}23, AIPV21, BM18, CTP^{+}22, CMH10, CDBF19, EABE19, ERBL12, FR24, FBDR23, FM20b, GDRCDBLÁ23, GDARM18, GNGHEM21, HYMT24, HUM^{+}21, UMS24, KIG23, KXZ^{+}24, KPC^{+}23, KMV^{+}23, LQFI17, LH21b, LMF23, LCMV17, LA20, Mac22, MTM22, NA23, PMT^{+}23, Pic21, Pir22, PPDC22, PR12, RC18, RRS20, ST22, SHC24, SPM^{+}22, YHN21]. Technology [ABH24, ACA^{+}21, AA24b, ASG24, BCG^{+}23, BN14, CEP24, CBRS21, CPM17, DCA16, EvdMB20, ELC^{+}19, FWLY23, FM16, FK21, FAC22, GLD^{+}18, HSI^{+}22, IPP24, IKS^{+}21, JWD^{+}24, KZRH22, KT22, KDG24, KL23, KNN^{+}24, KFK^{+}24, LWY^{+}15, LYY^{+}24, ML20, NB24, OGR^{+}20, Orl22, PCM24, PPM21, Rec21, RKT19, dSRRMC^{+}23, SLS20, SE15, SRM20, SA24, SH20, Tre21, VFS^{+}19, XKKL23, Yiu21a, ZSB19]. Technology-Based [FK21]. Technology-Enhanced [SE15]. Technology-Mediated [DCA16]. Teens [BCL11]. Telco [JLGF19]. Telecommunication [BM18]. Telehealth [COL21]. Telemedicine [MDRR24, VMAG23]. Telemetry [LB24]. Telerobotic [BBH^{+}24]. Television [SOB^{+}19]. Tell [AF22]. Temperature [DSMM21, RSMC21, WLG^{+}22]. TempoFuseNet [SBRZ24]. Temporal [CH23, CFG20, EvdMB20, MP20, MHL^{+}23, PL22a, SKA^{+}23, WX20, WX22, WYZ24, YCD^{+}23, dCMA21]. Temporal-Guided [CH23]. Tendencies [MRRMC21]. Tendency [GM19, TTS24]. Tennis [LA23]. Tensions [Roy14]. Tensor [XZ19]. Term [BKL^{+}22, DYS15, HSL^{+}23, KKBG23, LGY23, PLA^{+}21, SSD24, SRSDF23, YZCR24, ZXJ22]. Terminal [FFP^{+}22]. Terminology [SSH20]. Terrestrial [FSC^{+}23, QLL^{+}21]. Territorial [KKK^{+}20]. Tesla [LRD23]. Test [GNNZRCRG18, LRD23, MUS11, PLL11]. Testbed [AHKC23, BRP^{+}13, TSZ^{+}18]. Testbeds [CALHG^{+}24, DDPVP18, WGN23]. Testing [ABCC22, CB24, GFHK^{+}12, KYG22, TAAK21, WML24, YB22, RBVV22]. Tests [BCS24, VLC^{+}19, YB22]. Text [ADAq^{+}22, BKW^{+}24, CT20, FTAA21, GPA23, GKW^{+}10, HSSPTM^{+}23, JSPH21, LC23, LWX18, LLZL19, LL20a, LWBM22, MCR^{+}23, ND24, RKS^{+}21, WCZ23, ZLXY19, ZGZ^{+}18, ZZZL23b, LL20a]. Text-Based [GKW^{+}10]. Textbook [SFD20]. Textbooks [FGSD22]. Texts [FRKS22].

Textual [Ant12, SOE^{+}24]. Texture

[KWI^{+}19]. th [REP22]. Thanks [CPM17]. THBase [QMN19]. Theatre [CYO^{+}17]. Their [GDRCDBLÁ23, HOPGRV21, KLKA19, Min20, SOSC^{+}16, TGS^{+}19]. Thematic [FMNS11]. Theoretic [AZ19, FC22, LSF19, MHSH16, PW22]. Theoretical [CC22b, dSFdRP24, FHS^{+}10]. Theories [JSAO^{+}23]. Theory [AKT24, DV23, FKZ22, GRG21, HLZ23, JP21, TKT^{+}18]. Theory-Based [HLZ23]. there [GKK^{+}19]. Thermal [VBB24a, VBB24b, ZGL20]. ThermalAttackNet [DSMM21]. Thermoelectric [AOF24]. These [TPJ21]. Thesis [JHL^{+}18]. Thickness [LLW21]. Thickness-Based [LLW21]. Thing [AA19b, BSEL20]. Things [AYH^{+}21, CVG^{+}22, DKH24, GNNZRCRG18, HRAA19, JDH21, KS24, KE22, Lee20, MKSK23, MAEP18, PLG23, QLR^{+}22, RGB^{+}23, TL19, TN20, AOI^{+}23, AAM^{+}24, AAAHA22, ALB24, AAGB^{+}20, AM23, AA20, ACF24, AAAKJ22, AS15, ASN22, ABB23, AMY24, BML19, BRF^{+}23, BTS^{+}21, BFL20, BCFP21, BYS^{+}15, BMGI21, BCM^{+}19b, CPM^{+}19, CMK^{+}16, CLW23, DKH24, FR24, FMA^{+}20, FPKK22, GZC21, Gro22, HDBD23, HZQ24, HKH^{+}24, UMS24, Jal19, JUH^{+}23, JMJ^{+}23, KDG24, Kha24, KPH^{+}24, KHRG19, KNN^{+}24, KA23, Lel19, Li18, LMF23, LZL^{+}17, LLW^{+}24, MAEP18, MTK^{+}24, NDKBL19, NBL^{+}24, NZU20, OMD^{+}22, ORK^{+}19, PAM^{+}24, POC20, Poz24, RR24a, RAU^{+}24, RIS^{+}17, RKT19, RRAGMMGG20, RR24b, RSL^{+}24, BNK^{+}23, SJA^{+}23, SVK^{+}22, SS23, SST23, SGDT19, TSD24, UB18, VC19, VCPPRC20, WXL^{+}23, WYW^{+}22, WY19, ZZZ^{+}23, ZZY19]. Things-Based [AAM^{+}24]. Things-of-Interest [AYH^{+}21]. Think [DG18]. Thinking [VBWW22]. Third [ASM^{+}24]. Thorough [DPR^{+}21]. Threat [ADS20, Aln22, ESBE23, KCK^{+}19, Kos16, KAS^{+}20, SKI^{+}19, SYLL21, YYY^{+}23, YOI19, dCdG20]. Threats [AAN^{+}24, LGV13, MAW^{+}23]. Three [ASMM22, BTB22, Goe12, KES24, LCS12, SJG18]. Three-Dimensional [Goe12, KES24]. Threshold [AKK^{+}24, HEP^{+}11, LGY23, WUH^{+}23]. Thresholding [LXY17]. Thresholds [XCSM18]. Throughput [JP20, KALY17, KLTC17, LWY20, LSZ21, VL20]. Throughput-Aware [KALY17]. Throughput-First [LSZ21]. Thwart [YG20]. THz [OGB^{+}24]. Ticket [CCD^{+}21]. Tier [CKK21, Kab23, MEO18b, SRSDF23, YPG20]. Tiered [CL16, GGK18]. TiMAC [TPD^{+}20]. Time [ATFMOU^{+}24, Alb19, AMCI17, AAG^{+}22, AOF24, AGA^{+}22, CMPT24, CBKL22, CR23b, CALHG^{+}24, CC23, DCA23, FML^{+}22, GPM21, HQC^{+}24, HC21, IMR^{+}18, IUK^{+}23, JDH21, JNARV^{+}24, KLKP19, KPKM23, LLHW20, Lel19, MXT^{+}16, MDB22, MAM^{+}21, PG23, PLA^{+}24, PW21, PPDC22, RR24a, RIS^{+}17, RGCM21, SNH^{+}19, SH21, SS22, SNS22, SL24b, SWZ18, SZZL23, TSHO24, TOLG21, TVAT24, TRLR24, TZE^{+}21, VGK21, VGS^{+}19, WX20, WLG^{+}22, WKD22, WBS^{+}24, Yan19, YZ23, YXL^{+}21, ZSZ18, ZHW23, ZGH13]. Time-Dependent [ZGH13]. Time-Efficient [TSHO24]. Time-Outs [PG23]. Time-Sensitive [ATFMOU^{+}24, JNARV^{+}24]. Time-Series [TOLG21]. Time-Variant [ZSZ18]. TimeBank [LLW21]. Times [ELC^{+}19, SOSR^{+}16]. Timing [KKS^{+}19]. TinyML [FK24, KGK^{+}24, SKKS22]. Tips [CMR24]. TLC [WC20]. TLD [GPLK22]. TMR [dSFdRP24]. TMR-I4.0 [dSFdRP24]. Today [Bod22]. Toddlers [Pri23]. Token [KHM24, MCA23]. Tokenization [KFK^{+}24, MZ19]. Tokens [GM23, MCT^{+}23]. Tolerance [BP24, RK19]. Tolerant [CAS^{+}20, KKP22b, KST24, MKK17b, MLM14, USS19]. Tomographic [XZ19]. Tomography [KGKP20]. Tongue [SDL^{+}19]. Tool [BMB23c, BKW^{+}24, ESLB20, FSC^{+}23, GOM23, HPF^{+}24, KMK24, LS15, MBAS20, MS16, Mot24, NPB12, OR24, PG14, AJT^{+}23, SPC^{+}11, VMAG23]. Toolbox [SSR17]. Toolchain [Ant24, RLB^{+}23]. Toolkits [YWZ^{+}24]. Tools [ASFAV23, BM23, CS11, JSAO^{+}23, KID^{+}15, MDT20, MTKS24, NS12, PLA^{+}21, PFL^{+}12, Ray23, TRB22, Whe09]. Tooth [SDL^{+}19]. Tooth-Marked [SDL^{+}19]. Top [ALH24]. Top-Down [ALH24]. Topic [Alh22, BOB^{+}20, LLY24, LY19, VLPR20, WZC19]. Topic-Specific [LY19]. Topics [CQWC22, HS18b, LTM^{+}22]. Topologies [GX20]. Topology [GAMMPCRÁ24, KKP22a, MWY19, TPD^{+}20]. Tor [MBF^{+}12]. Tourism [AWBS23, ALST20, FGBMG19]. Tourist [MK20, SGMMRG21, SS14]. Tower [TL24a]. Town [UUT^{+}19]. Town-Watching [UUT^{+}19]. Toy [SPC^{+}11]. TPM [KHN^{+}11]. TPU [AAG^{+}22]. TQU [NLD^{+}24]. TQU-SLAM [NLD^{+}24]. Trace [JLH^{+}13, SL24b]. Traceability [GFHK^{+}12, PKMS23, Yiu21a, Yiu21b]. Traceroute [PG19]. Traces [IMR^{+}18]. Tracing [GFPD21, KKD24, PMA^{+}22, PFT^{+}24, RRLT20]. Tracking [BKR24, CZZH15, EPC24, KES24, NKSF24, PDI21, RRLT20, SHC24, WL20, WADL^{+}24, ZHH24]. Tract [IHMG22]. Trade [AHMI22, GL13, GPLK22, HKZ^{+}22, TMS24]. Trade-Off [HKZ^{+}22]. Trade-Offs [GL13, TMS24, AHMI22]. Traded [RSV^{+}24]. Tradeoff [QWR18]. Tradeoffs [SMG13]. Trading [FKM24, MCA23, WKD22]. Traditional [EM09, NS12]. Traffic [ALZ^{+}23, AES21, AMAE22, ABH24, APP22, ALST20, AAE^{+}22, AMY24, CBKL22, CC22a, DHDAu22, ESLB20, GSL20a, KKBG23, KLG^{+}24, LLCL23, MSN13, MTKS24, NAFM^{+}24, PCC^{+}18, Pin10, QNJ24, RRP17, SZFC18, WHB^{+}21, WYZ24, WLLY24, WSD22, XGY^{+}22, YT19, ZZZ^{+}23, ZLL^{+}21]. Trails [MMQ^{+}24]. Trained [LHD^{+}24]. Training [AM24, AS19, ACBP19, FKF24, dPCGMC20, GTC19, KI20, KHN^{+}22, KL21, LSJ19, LTTS21, ML20, RRS20, URH^{+}23, VM24a, ZZZ^{+}23]. Trajectories [LTM^{+}22, TGELGH13]. Trajectory [QMN19, YKY18]. Trans [YLY^{+}22]. Transaction [AZS23, FWL22, HM20, OKH13, SCA^{+}19]. Transactions [JJFC22]. Transactive [YG20]. Transceiver [NAV^{+}23]. Transcoding [NBT22]. Transcripts [CS22]. Transfer [CQWC22, For14, HRGQHOA21, LWZ^{+}21, MR23a, MSL^{+}23, MKM^{+}19, PPKS21, PS16, PPM21, WZL21, ZL20, ZY20b]. Transferable [KPH^{+}24]. Transform [PTZM22, Pil18]. Transformation [CT24a, CT24b, MKHR21, RABC21, ZZYC21]. Transformations [Far12]. Transformer [ATY20, ABH24, BTB22, CB24, KM24, LF23, MT23, PCM22, TWW24, WZL23]. Transformer-Based [BTB22, CB24, PCM22]. Transformer-CNN [KM24]. Transformers [HSSPTM^{+}23, YST24]. Transforming [AARTSBAE24, BRF^{+}23, GBKN23]. Transit [MDRR24, SWZ18, SWY^{+}18]. Translating [STV^{+}22]. Translation [AFAAM19, PLYD20, ZLY^{+}20]. Transmission [CSDAM^{+}23, GPMO20, JWRX17, KKM^{+}22, LZ23, LNH23, LHW24, LW15, XNZ23b, YHG^{+}21]. Transmissions [NSV17]. Transmit [JKM^{+}23, YPXH19]. Transparency [ATR20, IIOW16, KE22, RL24, SK24, VdHH^{+}19]. Transparent [ACS24, CT24a, KKD24]. Transponders [AD20]. Transport [IHMG22, KPC^{+}23, KSSF19, LB24, LTW21, MLL^{+}22, PKB^{+}24, SRS^{+}24]. Transport-Layer [PKB^{+}24]. Transportation [HD19, LF24, NPMPM23, PSC19, WUH^{+}23, ZKKK19]. Traps [AOF24]. Travel [Giu18a, THL21, YZ23]. Traveling [OT16]. Treatment [NS12, RRS20]. Tree [CCC19, FMA^{+}20, LBK22, TYH23]. Tree-Based [FMA^{+}20, TYH23]. Trees [GDS23, KHC^{+}24]. Trek [SOB^{+}19]. Trend [RNFS20, ZL21]. Trends [BOH23, CS22, DWL21, KBHP24, KA23, KTUF19, LPPG22, Mal20, MAEP18, NZ14, PKF24, SPM^{+}22, VSEH^{+}21, CMP^{+}19]. Triage [SKI^{+}19, VMAG23]. Trial [YGQ^{+}22]. Trigger [MEO18b]. Triggered [YT15]. Troll [TLM^{+}20]. True [RS22]. Trunk [AAG^{+}22]. Trust [ATE24, AMC^{+}24, ACF24, ATR20, BLW^{+}17, DDPVP18, DPBG18, KST19, MCS^{+}24, MGJ22, NPWC24, PKMS23, PZH^{+}18, Pic21, SG12, SK24, SL24a, TAD23, TLC15, TL17, TT22, TRST24, WZS^{+}22, WYW^{+}22, TRST24]. Trust-Based [PZH^{+}18, TRST24]. TRUST-ME [TRST24]. Trusted [ALB24, CB22, LHL24a, SS15]. Trustless [CMPT24]. Trustworthiness [MOAI17, RSX18, SS22]. Trustworthy [LH19, PSM^{+}18]. TSN [JNARV^{+}24]. TSP [LLW^{+}17]. Tuberculosis [APB^{+}16, ACBP19]. Tuning [WCZ23, YFY24]. Tunnel [LBA^{+}24]. Tunneling [PKB^{+}24]. Turkish [GSD^{+}17]. Turning [KLZ16]. Tutorial [Jia19]. Tutorials [AGN21]. TVWS [ACA^{+}23]. Tweet [MSR^{+}14]. Tweeting [UUT^{+}19]. Tweets [HP21, KDD^{+}21]. Twin [AAGB^{+}20, CGBF24, CALHG^{+}24, FLSL24, GKFV24, GTWN24, JWD^{+}24, MEH^{+}22, Orl22, PC23, QLR^{+}22, ZYWS24, GTWN24]. TWIN-ADAPT [GTWN24]. Twin-Enabled [GTWN24]. TwinNet [RD18]. Twins [AA24b, BD24, KVWC24, QXC^{+}21].

Twitter [AA21a, AA19a, AAC21, AHD21, AZV16, BBA21, BTB22, JXA19, JSPH21, MT20, MSR^{+}14, MF20, SH21, TPP^{+}23, UANU20, VLPR20, YMJ20, ZAV14]. Twitter-Based [UANU20]. Twitter-Related [YMJ20]. Two [ADS20, AMY24, CCKH21, HDK^{+}12, LQY^{+}19, MEO18b, MMH16, MZH22, QWR18, RF21, SBRZ24, SXXH19, SNZ21, YPG20, YLJ^{+}19]. Two-Dimensional [YLJ^{+}19]. Two-Factor [ADS20]. Two-Hop [MZH22]. Two-Layer [CCKH21, RF21]. Two-Layered [QWR18]. Two-Level [SNZ21]. Two-Stage [AMY24]. Two-Stream [LQY^{+}19, SBRZ24]. Two-Tier [MEO18b, YPG20]. Two-Way [SXXH19]. Type [HC19, JP21, PTK^{+}23]. Types [Ala20, GOM23, ZS18].

U [Rot12, ZMDCEMI22]. U-City [Rot12]. U-NET [ZMDCEMI22]. U.S. [RSV^{+}24, SG21b]. UAV [CRMT23, CPKK23, DXL^{+}22, EARP^{+}24, HS20, JP20, JGY^{+}24, LBK22, RRT^{+}23, SHA23, WLWZ23]. UAV-Based [LBK22]. UAV-Enabled [CRMT23]. UAVs [BNJ24, GM20, MSP21]. Ubimus [KYT^{+}23]. Ubimus-Design [KYT^{+}23]. Ubiquitous [BKSS19, TWM20]. UI [BR21]. UK [CS22, TBT^{+}23]. UK-National [TBT^{+}23]. Ukiyo [LWBM22]. Ukiyo-e [LWBM22]. Ukraine [MFH22]. Ukrainian [OC20]. Ultra [MR23a, SBS18, SKKS22, VKM^{+}19, WLHR13, ZAS^{+}24]. Ultra-Dense [MGV17]. Ultra-Lightweight [SBS18]. Ultra-Low [SKKS22]. Ultra-Low-Latency [VKM^{+}19]. Ultra-Reliable [ZAS^{+}24]. Ultra-Wideband [MR23a, WLHR13]. UMLS [SSH20]. Unbalanced [TDW24]. Uncertainty [CGRV20, CEZC19, WBS^{+}24]. Uncertainty-Aware [WBS^{+}24]. Unchangeable [RK19]. Unconscious [Min20]. Unconstrained [NDS^{+}23, RGM^{+}21]. Uncover [HLC24].

Underlay [GKS10, JWRX17, KTP17, Wan17]. Underlay-Based [GKS10]. Understand [MT20]. Understanding [CWC^{+}22, FHS^{+}10, LFGR20, SPSS17, SSS^{+}19, SKM21, WKPC18, ZSS^{+}22, Zha19]. Underwater [ACF24, CHMZ21]. UNet [DOB21]. Unified [FKM24, Kra22, LBK22, SDS^{+}22]. Uniform [CDLY24, VdHH^{+}19]. Union [GPLK22]. Unit [EMHF19, OAH^{+}24, SMVP21, SJ20]. Units [FCR23]. Universal [QSF^{+}17, RD18, TKH^{+}11]. Universities [GNGHEM21, RABC21, RRAGMMGG20]. University [BM19, CKFH22, ESB^{+}20, HOPGRV21, PPM21, SHA^{+}21, VCPPRC20, VCGOMC^{+}21, ZJ18, dSPR^{+}22, ERBL12, VCRCSV^{+}21]. Unknown [KAAAA22, WSD22]. Unmanned [LA23, LS21, LLSS24, MAM^{+}21, ZWPL19]. Unreachable [LWW22]. Unsignalized [YT19]. Unsteady [WX20]. Unstructured [BKW^{+}24]. Unsupervised [AKDL24, BCNS20, HOV23, HMR^{+}21, KPT^{+}24, LNT^{+}24, LTM^{+}22, TCH18, VMD^{+}22, ZZZ^{+}23]. Unveiling [BR23]. Up-Conversion [LMLW18, LMH19]. UP-SDCG [ZQL^{+}24]. Update [LWW^{+}24, POC20]. Updated [CBM^{+}20]. Upgradation [SMVP21]. Upgrading [PR11]. Uplink [ABK22, MEO18a, SHN^{+}19, SXXW20, WYS17, Wan17]. Uploading [MOAI17]. upon [MDB22]. Upper [HT14, IHMG22]. Upward [WCM19]. Urban [AK18, ADAB21, BG24, CMK^{+}16, CC22a, DGADE14, DIE^{+}21, GKFV24, GMB^{+}21, HD19, HDK^{+}12, HS20, IUIL23, KPRP24, PPDC22, RLB^{+}23, XGY^{+}22, ZN22, ZYLM17, ZLL^{+}21]. URL [BMP21, RAA^{+}22]. URL-Based [BMP21]. Usability [AA24a, MBF^{+}12, YB22, YA24]. Usage [AFS^{+}23, FAS^{+}23, GS11, LS15, SS22]. Use [AAAAM^{+}21, ASFAV23, ASFAAV24, ASG24, BTS^{+}21, Cop14, FM16, GGT^{+}23, HS18b, IKS^{+}21, KPP^{+}20b, LBK22, MAW^{+}19, MMS22, MKSJ22, MSL^{+}23, MGUD12, Nil12, NWG20, PCVH24, PST^{+}21, POC20, QSML24, RNFS20, RSMC21, Söd13, SZZL23, TCG19, VMGT22, VTPT24, VCRCSV^{+}21, VCGOMC^{+}21, VCJAMN22, VM24b, Zar22, SG19]. Used [JGV^{+}21, MS22, NZU20]. Useful [NRS24]. Usefulness [RRAGMMGG20]. User [AOI^{+}23, AMAA^{+}21, ANHMS24, AMI23, BAR^{+}21, CSDAM^{+}23, CHW^{+}21, CGBF24, DDD^{+}22, FKF24, GTM22, GPL24, GWK17, GM16, GPRMGP21, GRG21, HKZ^{+}22, HL19, KDKG22, LLZ21, LB23, LOL21, LXL^{+}19, LHL24b, MS16, MGBG21, Øie12, PNMK22, PRDK22, PVKG23, RYY10, oRL19, RYAA22, SSF^{+}18, SZL21, SHC24, SXXW20, SVFV23, SIZD12, SOR16, VKKG22, VFFHF19, WYX18, WD24, ZCZ17, ZZWP19, ZY20a, ZLOW24]. User-Centered [GPRMGP21, SIZD12]. User-Centric [RYY10]. Users [CLM^{+}21, CCM^{+}17, GLRM19, KLKA19, MMQ^{+}24, SAF14, WLH^{+}17, YLYL22]. Uses [YT19]. Using [AOI^{+}23, ALZ^{+}23, AAK24, AKDL24, AKM^{+}23, AK22, ARAAA19, AA23a, AES21, ALL24, ADAq^{+}22, AAAR24, Alh22, AAG^{+}22, AAC21, AYH22, AA24a, AAF^{+}22, ASLG22, ACA^{+}23, ABLR23, BMB^{+}23b, BMB23c, BHSAY24, BCZ^{+}23, BdCRM10, BAR^{+}21, BSE^{+}21, BMS24, BJL^{+}22, BM14, CD23, CVG^{+}22, CLW23, CS11, CCC19, CWC^{+}22, CR23b, CASN21, CGT21, COL21, CBB17, CCYL24, DCHM16, DZH^{+}22, DAMAS^{+}19, DOB21, BBBHA24, ERBL12, FWLY23, FHL^{+}24, DB23, FRKS22, FSG22, FC22, FSXP22, FLM23, GPPB^{+}21, GDN^{+}23, GMH18, Goe12, GdSdC23, GHGL22, GB23, GRA21, HD19, Han13, HT14, HXHP17, HP21, HOV23, HBM^{+}21, HN18, HBMS20, IFF21, IDR23, JTA^{+}21, JOL22, JJFC22, JPVC21, JSAO^{+}23, KAEAB24, KM24, KI20, KVST19, KDEA^{+}23, KWI^{+}19, KG18, KDJ24, KHN^{+}11, KKBG23, KBF23, KVL^{+}22, KKM^{+}22, KKTS22, KTD24, KRSF22, LWHR14]. Using [LB24, LDd^{+}24, LMLW18, LZL^{+}23, LQY^{+}19, LRdLS19, LBC18, LA20, MAVS20, Mah17, MAW^{+}23, MCCHO19, MTKS24, MCA23, MKSJ22, MDM21, MIK^{+}23, MT20, MAM^{+}21, MCR^{+}23, MMMH19, MEJA23, MMQ^{+}24, MSAA22, MN21, dSMdO22, MSJ^{+}23, MPCS^{+}23, MTHN22, NBL^{+}24, NZU20, OS23, OOM^{+}23, OMD^{+}22, OY19, OMTFFL21, PTK^{+}23, PDTM15, PS16, PDT23, PRDK22, PG19, PLSF14, PB23, PCM22, PTZM22, PPM21, POC20, Pri23, RAJJ23, RNFS20, RM23, Rec21, RSX18, RKS^{+}21, RAA^{+}22, SVGGO24, SSH20, SHHM21, SSR17, SQK^{+}17, STV^{+}22, SC23, SSR^{+}21, SS17, SSS^{+}19, SA24, SLSK21, SH21, STS23, SDL^{+}19, TAMA22, TBT^{+}23, TAHO23, TWM20, TRHU23, TDM^{+}24, TSZ^{+}18, TVAT24, TRB22, TS22, TWW24, TTK24, UJ22, UUT^{+}19, VCG24, VGK21, Vem22, VI16, XCSM18, XY18, YFY24, YCHG18, Yiu21a, ZMDCEMI22, ZHZ^{+}19, ZWWW17, ZL22]. Using [ZZZ^{+}23, ZXMB22, ZSF20, ZTBD20, dCdZCC24, dVBA23]. UTAD [dSPR^{+}22]. Utilising [SAF14]. Utility [FGBMG19, GHGL22]. Utilization [CZY17, QWR18, Rók23]. Utilize [TLV23]. Utilizing [ASF^{+}23, DDD^{+}22, KAS^{+}22, KP15, KSG^{+}24b, LHL24b, NB24, PV20, PV22, SB23, YMH22]. UWB [CLZL18]. UWF [BMB^{+}23b]. UWF-ZeekData22 [BMB^{+}23b]. UX [KG18].

V [DNKDN^{+}24, MLPH23, AJ18]. v-Mapper [AJ18]. V2V [Li23]. V2X [WUH^{+}23, ACM^{+}19, BHC23, BCM^{+}19a, CGM^{+}19, HZW^{+}20, HSC^{+}22, MCH^{+}20, OSdS24, SGLZ20, SYGY21, WGN23, WHB^{+}21, YK21]. Vadis [dSRRMC^{+}23, ST20]. Validating [AWSA23, AA20, FM21, SDM22]. Validation [AIP^{+}24, DPR^{+}21, GWK17, KHM24, KHRG19, LL20c, PG23, TADS20, TRC^{+}21, WML24]. Valuation [MCT^{+}23]. Value [Cop14, HLLT10, JK20, KS14, KHM24, TLV23, VCB14]. Value-Driven [KHM24]. Valued [ACAP20, GM19, MCCOCE24]. Values [AAAIA24, AGB22, LMY21]. VANET [GMS^{+}20, VKM^{+}19]. VANETs [BRP^{+}13, AK18, AKK^{+}24, KSA^{+}21, LDZP24]. Variability [KBBH21]. Variables [ALZ^{+}23, JGV^{+}21]. Variance [YZP22]. Variant [ZSZ18]. Variational [SLZY21, YLL21, ZH21]. Varied [GOM23, MDB22]. Varieties [TL24b]. Various [BBH^{+}24, IAG^{+}21, LN19, YAV^{+}21]. Varying [BBA21]. Vasilas [VBB24b]. VDTN [ABS20]. Vector [RKS^{+}21, YWW^{+}19]. Vectors [AJ24, Ala20, DDD^{+}22]. Vegetation [OAG21]. Vehicle [AK18, ASM^{+}24, AMC^{+}24, BCNS20, CC22a, CSZ^{+}24, DLX^{+}19, DWY19b, FTHY21, GLH^{+}22, HYL13, IPP24, LA23, LLR19, MNP^{+}24, RYAA22, RZQS18, RBG^{+}22, SLB24, WHB^{+}21, WML24, WUH^{+}23, XGY^{+}22, YT19, You23, ZWPL19, ZX24]. Vehicle-to-Everything [ASM^{+}24, WUH^{+}23]. Vehicle-to-Vehicle [AMC^{+}24, FTHY21]. Vehicles [AA23b, HZW^{+}20, KUBS22, LS21, MCCHO19, MAM^{+}21, MCH^{+}20, SVV^{+}22, YT19, ZAS^{+}24, EAHO21, KTCI21, SDM21, ZSF20]. Vehicular [AASAI22, AP19, BFY22, BCM^{+}19a, BKCL23, BKL^{+}19, BRP^{+}13, CSL^{+}20, CGT21, DHDAu22, DIE^{+}21, HDBD23, HNK^{+}21, IMR^{+}18, IKP^{+}13, IZK^{+}23, JUN^{+}24, MZS19, NMH^{+}21, PNS23, SSS^{+}19, SSO^{+}19, TGELGH13, USS19, ZR15, ZAS^{+}24]. Velocity [MEO18b]. Velocity-Aware [MEO18b]. Vendor [TXJ23].

Vendor-Agnostic [TXJ23]. Verbal [ER21]. Verifiable [FSXP22, SZW^{+}23]. Verification [HZCZ22, KSJZ23, KKT20, MT17a, SCA^{+}19, SL24b]. Verification-Based [SCA^{+}19]. Verify [Pic21]. Versatile [Ant24, MLK22]. versus [AAAT21, BCM^{+}19a]. Verticals [NA23]. Veterans [KP15]. via [AM24, ASBC22, AM21, ARC^{+}19, ACAP20, BZNB23, BMGI21, CMR24, EUDW23, FNN21, GFPVLR^{+}23, GZS^{+}24, GR22b, HM20, HSSPTM^{+}23, ICF^{+}11, KKP22b, KAAAA22, LLZL19, NMT23, PPKS21, PGFH24, QSML24, TAD23, TMS24, TDT^{+}21, VBB24a, VBB24b, WZL23, ZBK^{+}24]. Viable [NDKBL19, QA24]. Victimization [ER21]. Victims [BBM20]. Victorian [ICF^{+}11]. Video [DHEK19, DTTK19, FLM23, GWF^{+}23, GBDDVS23, HSM19, HYL13, IMLL24, KYS21, KLZ16, LH21a, LZS18, LMLW18, dMLMS22, LXG^{+}23, LWZL22, MTA^{+}20, MN21, MM23, NBT22, PTZM22, SBRZ24, SSR^{+}21, Sor12, TCH23, VCPT22, WWB^{+}20, YCHG18]. Video-Based [YCHG18]. Video-Coding [MTA^{+}20]. Videoconferencing [Nil12, RSG21]. Videos [JPVC21, LS15, VCG24]. Vietnamese [TLV23]. View [ALH24, CLTP22, HF22, LJWX23, MKKS18, Mar17, PGSL20, Sta23]. View-Invariant [MKKS18]. Views [NB12, PEN23]. Vine [AAG^{+}22]. Violation [DHDAu22, HBMS20]. Violations [FC19, KPC19, MMN19]. Violence [JLW19, KSG^{+}24b, Mar13]. Virtual [AARTSBAE24, ASFAV23, ASFAAV24, CASN21, CPI^{+}18, CFGD20, DL19, DPBG18, FSI^{+}24, FRE^{+}22, For14, GMGF24, GBDDVS23, GVV^{+}23, GSD^{+}17, GGCY20, HZWJ21, JBS^{+}23, JGV^{+}21, JPVC21, KTS18, KSO19, LSC^{+}17, LBC18, Mar13, ML20, MS20, MP16, MNWK21, PTK^{+}23, Pec18, SGWK23, Sem11, STG^{+}20, TMTMB20, VL20, VNJ18, VMAG23, VCGOMC^{+}21, WBG12, WYL^{+}21, WLCS17, SPC^{+}11]. Virtualisation [GVV^{+}23]. Virtualization [AP22, Aln22, ELC^{+}19, ELCS20, FR24, KKS^{+}19, Pap20, RC18, SADP21, STCP21, SN23a, SFSN^{+}24]. Virtualized [JK20, MSC^{+}21, NST23]. Virtuous [SHHM21]. Virus [Ler16]. VIS [NNRR^{+}21]. VIS-HAPT [NNRR^{+}21]. Vision [BS17, BFL20, KPRP24, LA23, MMQ^{+}24, MSP21, NA23, Rei23, SCC^{+}23, Wan23, YST24]. Vision-Based [LA23]. Visual [BCSB24, BR21, BS17, DMS^{+}22, Hor19, ITH^{+}21, JCPV22, KBBH21, KVL^{+}22, Lal14, LDZ^{+}23, LPM10, NLD^{+}24, PDI21, SSX20]. Visualisation [OR24]. Visualization [PMA^{+}22, RBG^{+}22, SFD20, WBG12]. Visualizations [VDSK23]. Visualize [BMB^{+}23b]. Visually [GSO^{+}10]. Visuo [NNRR^{+}21]. Visuo-Haptic [NNRR^{+}21]. Vital [DVSBV^{+}24]. VLAN [LYY^{+}24]. VLAN-VxLAN [LYY^{+}24]. VM [XCSM18]. VNF [BMJ^{+}22, LKIL19]. Vnode [KKD24]. Vocabulary [OR24]. Vocal [FR16]. Voice [ZLS19]. Voltage [DNPC^{+}22, KKM^{+}22]. Volume [SG23, Tos23]. Volumes [BBA21]. Volumetric [LTP24]. Volunteer [And21, Kra20, XQLZ19]. Volunteered [NZZ13, NZ14]. Vote [CKFH22, TMP22]. Vote-Based [TMP22]. VPN [GMD^{+}22]. VQ [CCKH21, LLCC24]. VQ-Compressed [CCKH21]. VQA [MM23]. VR [BOK^{+}21, FOJE19, LB23]. vs [BTB22, GKS10, KDD^{+}21, KPKM23, MKKG19, NS12, PdRB19, PSTZ23, SFSN^{+}24]. Vulnerabilities [BR23, HAA23, ITZ20, LLCL23, RKY20]. Vulnerability [AN22, BHSAY24, HCHP16, NPMPM23, RBVV22, TWM20, YZC^{+}20, ZL22]. VxLAN [LYY^{+}24].

W6 [TVAT24]. Walking [PS17]. Walks [BB20]. WAN [EH21]. Warehouses [BCZ^{+}23]. Warfare [MHSH16]. Warning [BCSB24]. Wartime [MFH22]. Watch [AJ24, AMAA^{+}21]. Watchdog [KSJZ23, SFEK18]. Watchdog-Based [KSJZ23, SFEK18]. Watching [UUT^{+}19]. Water [ATE24, GSL20b, JMZ^{+}22, NMT23, XLJ^{+}22]. Watermark [FR15, GRRZ18]. Watermark-Based [GRRZ18]. Watermarking [BCG10, Fra23, SS17, ZWWW17, ZWZ17]. WATS [FTAA21]. WATS-SMS [FTAA21]. Watson [GMMI16]. Wave [KIB^{+}22, LSJ19, MS23, SRS^{+}24, Tre21]. Wavelet [PTZM22]. Waves [JMZ^{+}22]. Wavy [JMZ^{+}22]. Way [KLKA19, SXXH19, WC23, ZCdRR^{+}23]. Weak [HP21]. Weakly [GGD23]. Wearable [Adi23, AA24a, KKL^{+}20, SOA^{+}20, TDM^{+}24]. Wearables [BS17]. Weather [CHH^{+}16]. Web [AA23a, AN22, AV23, AAD^{+}12, AFS^{+}23, BFFZ^{+}12, BM23, CP10, CB16, CHH^{+}16, DBD^{+}14, DCHM16, DV23, ERBL12, FAS^{+}23, Fer12, FAB^{+}12, Fra18, FMA^{+}23, FHS^{+}10, GPL24, Goe12, GNK^{+}10, GNNZRCRG18, HSC19, HL23, HDL^{+}14, HG12, IFF21, JDH21, Jur12, KMS^{+}12, KSO19, LPPG22, LSM^{+}12, MMD22, MG12, MGBG21, MDM21, MGLBMMSC20, MUS11, NPB12, NIK^{+}21, NB12, PDTM15, PV20, PFLT12, PR12, RRMI22, RKM^{+}22, RDD^{+}14, RMB^{+}12, RGCM21, SPSS17, SHC24, SWJ20, SLY^{+}12, SB23, TL19, VL18a, VVK^{+}21, VKV^{+}22, Whe09, XJ10, YB22, YA24, YMJ20]. Web-Based [AA23a, HG12, KSO19, LSM^{+}12, MUS11, SLY^{+}12]. Web3 [AAKS23, Cal24]. Web3/Metaverse [AAKS23]. WebAssembly [HM24, Ray23, SMR24]. Webcam [ZBK^{+}24]. Weblogs [EM09]. Webpage [SRS23]. WebRTC [FKK20]. Webshell [GMGK20]. Website [BN14, GKK^{+}19, ICF^{+}11, KZ20]. Websites [ASLG22, GPK^{+}20, HSSPTM^{+}23, KLKA19, MSAA22, TCG19]. Weibo [AHD21]. Weight [KC19, ZXXB22]. Weighted [SIJA10, SDL^{+}19, XY21]. Well [DIG20, HCW20, LL20b, PPDC22]. Well-Being [DIG20, HCW20, LL20b, PPDC22]. Went [GFPD21]. West [VLMP21]. WhatsApp [BGW16, PV20]. Wheel [MKSK23]. Wheel-Legged [MKSK23]. Wheelchair [ZBK^{+}24]. Where [GN16]. Which [MSC19]. White [YFY24]. White-Hat [YFY24]. Who [GJ15]. Wi [BCM^{+}19a, Lee17a, Lee17b, AHH^{+}23, CBKL22, LOL23, MN21, MTM22]. Wi-Fi [BCM^{+}19a, Lee17a, Lee17b, AHH^{+}23, CBKL22, LOL23, MN21, MTM22]. Wide [FM16, TFSBSPSC24]. Wide-Area [TFSBSPSC24]. Wideband [MR23a, WLHR13, WWYQ19]. WiFi [JK24, JNARV^{+}24, SPK24, WD24]. WiIR [WD24]. Wiki [Cap12, PLSF14]. WikiGIS [BRBC14, RMB^{+}12]. Wikipedia [BdBC23, DeD16, FTAA21, HD16]. Wild [LCY^{+}23]. Wildfire [BCM^{+}23]. Wildlife [OAG21]. Will [LQFI17, Pil18, RBVV22]. WiMAX [CLZL18, HYL13, SIJA10]. WiMAX/WLAN [CLZL18]. Win [AFS^{+}23]. Wind [LGY23, WLZ18]. Window [LL23a]. Windows [SWZ18, HCXH16, IAM24]. Winter [USS19]. Winters [KKBG23]. Wireless [AAM^{+}24, AMZ^{+}20, ASD^{+}22, AOF24, BF18, BCG10, BL22, CB22, CLW23, CCTC23, CMH10, CDBF19, DPC^{+}24, DYS15, DVSBV^{+}24, DWZN16, DMZ^{+}19, EAHO21, EW19, FM16, FM20b, GSL22, GAMMPCRÁ24, HF20, HUM^{+}21, IDR23, JBS^{+}23, JCPV22, JP20, KSD^{+}23, KMW18, KLTC17, LQFI17, LBA^{+}24, Lee18, LXL^{+}19, LW15, Mac22, MMA10, MCM^{+}10, MG10, MP13, NMT23, OLCMdA20, OAI^{+}22, Oza23, PSM^{+}18, RYY10, oRL19, SCB21, SRS^{+}24, SUF^{+}22, SRSDF23, SNH^{+}19, SLW17, TAMA22, TAHO23, TL24a, THV^{+}22, TSDG22, TTS24, WCF^{+}16, WLZ18, WLWZ23, XSX18, YW16, YLH^{+}17, YXL^{+}21, YSAY23]. Wireless-Powered [SNH^{+}19]. Wise [GTC19, LCCC24]. Within [KNN^{+}24, AAGB^{+}20, DCG12, GNGHEM21, HSC19, LH21b, RMLAZOPC21, TEE10, ZVV12, dCdG20]. without [WHW22]. WLAN [AKJB20, CLZL18, MSK^{+}18]. WLANs [LTY18, LWY20]. WMNs [YC16]. WMS [DPR^{+}21]. Word [KA23, LTM^{+}22, OY19, YWS22]. Word-Level [YWS22]. Word2vec [OY19]. WordNet [OY19]. WordPress [MNH^{+}24]. Words [YKC19]. Work [AMEF21, HM20, QXC^{+}21]. Workcell [SU17a]. Worker [DQ18, SU17a, SU17b, SU17c]. Worker-Cobot [SU17c]. Worker-Industrial [SU17a]. Workflow [AFS^{+}23, LCS12, SL24b, YMSP24]. Workflows [AD18, CMO^{+}21, HC21]. Working [BR21]. Workload [JLH^{+}13]. Workplace [AGO^{+}23b]. Workshop [UUT^{+}19]. World [AGO^{+}23b, BSE^{+}21, Bod22, BLU^{+}15, CPM^{+}19, CB22, For14, GSP15, KSO19, LNB12, LCS12, NZZ13, NPM^{+}18, OO19, PG23, SGWK23, STG^{+}20, WT13]. Worlds [CSC21]. Worldwide [MCCOCE24, VC19]. Worms [YFY24]. WoS [RCGSL19]. WPT [LXL^{+}20]. Writing [EM09, ER10, LS23]. Wrong [GFPD21]. WSN [AMCI17, JLZ18]. WSNs [BZNB23, Mah17, SFEK18]. Wuhan [WLJM21].

X [GHFL24, HC19, VR21]. X-ray [VR21]. xAAL [LK20]. XES [MGM^{+}23]. XML [ADSAKAD22]. XR [MVAHBL^{+}23]. XSS [AAA^{+}24a, FHXL19].

Year [RFSOHMSB21]. Years

[LF16, ZSB19]. Yielding [YT19]. Yolov5 [SWC^{+}23]. YOLOv7 [TVAT24].

YOLOv7-W6-Pose [TVAT24]. Young [BKR24, BBM20, HS18b, NB12, Söd13]. Youths [Tom21]. YouTube [JPVC21]. YouTuber’s [HC19, Hor19].

ZeekData22 [BMB^{+}23b]. Zenith [Mat20]. Zero [HCXH16, MCS^{+}24, MNH^{+}24, RBVV22, UJ22]. Zero-Day [HCXH16, MNH^{+}24]. Zero-Inflated [UJ22]. Zero-Trust [MCS^{+}24]. Zone [RS17]. ZSM [XKKL23]. ZVC [RBVV22].

References

Alharbi:2019:PRC

[AA19a] Adel R. Alharbi and Amer Aljaedi. Predicting rogue content and Arabic spammers on Twitter. Future Internet, 11(11):229, October 30, 2019. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
11/229.

Alreshidi:2019:ASI

[AA19b] Abdulrahman Alreshidi and Aakash Ahmad. Architecting software for the Internet of thing based systems. Future Internet, 11(7):153, July 10, 2019. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
7/153.

AlSuwaidan:2020:VAH

[AA20] Lulwah AlSuwaidan and Nuha Almegren. Validating the adoption of heterogeneous Internet of Things with blockchain. Future Internet, 12(6):107, June 21, 2020. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/12/6/107.

Alenezi:2021:MLD

[AA21a] Mohammed N. Alenezi and Zainab M. Alqenaei. Machine learning in detecting COVID-19 misinformation on Twitter. Future Internet, 13(10):244, September 23, 2021. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
10/244.

Alicea:2021:MFN

[AA21b] Michael Alicea and Izzat Alsmadi. Misconfiguration in firewalls and network access controls: Literature review. Future Internet, 13(11):283, November 08, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/11/283.

Alzahrani:2021:DNI

[AA21c] Abdulsalam O. Alzahrani and Mohammed J. F. Alenazi. Designing a network intrusion detection system based on machine learning for software defined networks. Future Internet, 13(5):111, April 28, 2021. CODEN ???? ISSN 1999-

5903. URL https://www.

mdpi.com/1999-5903/13/
5/111.

AlSalem:2023:AIW

[AA23a] Thanaa Saad AlSalem and Majed Aadi Al-Shamari. Assessing interactive Web-based systems using behavioral measurement techniques. Future Internet, 15(11):365, November 11, 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
11/365.

Alsamiri:2023:FLI

[AA23b] Jadil Alsamiri and Khalid Alsubhi. Federated learning for intrusion detection systems in Internet of vehicles: a general taxonomy, applications, and future directions. Future Internet, 15(12):403, December 14, 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
12/403.

Alshamari:2024:UEW

[AA24a] Majed A. Alshamari and Maha M. Althobaiti. Usability evaluation of wearable smartwatches using customized heuristics and system usability scale score. Future Internet, 16 (6):204, June 6, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.

                                        mdpi.com/1999-5903/16/
                                        6/204.

Attaran:2024:AHT

[AA24b] Sharmin Attaran and Mohsen Attaran. Advancing healthcare through the integration of digital twins technology: Personalized Medicine’s next frontier. Future Internet, 16(12): 477, December 20, 2024. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/16/
                                        12/477.

Alreshidi:2019:SAM

[AAA^{+}19] Abdulrahman Alreshidi, Aakash Ahmad, Ahmed B. Altamimi, Khalid Sultan, and Rashid Mehmood. Software architecture for mobile cloud computing systems. Future Internet, 11(11):238, November 13, 2019. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/11/11/238.

Atlam:2020:RBA

[AAA^{+}20] Hany F. Atlam, Muhammad Ajmal Azad, Madini O. Alassafi, Abdulrahman A. Alshdadi, and Ahmed Alenezi. Riskbased access control model: a systematic literature review. Future Internet, 12 (6):103, June 11, 2020. CODEN ???? ISSN 1999-

                                        5903. URL https://www.

mdpi.com/1999-5903/12/
6/103.

AlJamal:2024:RML

[AAA^{+}24a] Mahmoud AlJamal, Rabee Alquran, Ayoub Alsarhan, Mohammad Aljaidi, Mohammad Alhmmad, Wafa’ Q. Al-Jamal, and Nasser Albalawi. A robust machine learning model for detecting XSS attacks on IoT over 5G networks. Future Internet, 16(12):482, December 23, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
12/482.

Alshdadi:2024:BDD

[AAA^{+}24b] Abdulrahman A. Alshdadi, Abdulwahab Ali Almazroi, Nasir Ayub, Miltiadis D. Lytras, Eesa Alsolami, and Faisal S. Alsubaei. Big data-driven deep learning ensembler for DDoS attack detection. Future Internet, 16(12): 458, December 5, 2024. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
12/458.

Aldaoud:2024:DSM

[AAAAK24] Manar Aldaoud, Dawood Al-Abri, Medhat Awadalla, and Firdous Kausar. Data structure and management protocol to enhance name resolving in named data networking. Future Internet, 16 (4):118, March 30, 2024. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
4/118.

Al-Ansi:2021:SIE

[AAAAM^{+}21] Ahmed Al-Ansi, Abdullah M. Al-Ansi, Ammar Muthanna, Ibrahim A. Elgendy, and Andrey Koucheryavy. Survey on intelligence edge computing in 6G: Characteristics, challenges, potential use cases, and market drivers. Future Internet, 13(5):118, April 30, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/5/118.

Agbo:2022:MDI

[AAAHA22] Benjamin Agbo, Hussain Al-Aqrabi, Richard Hill, and Tariq Alsboui. Missing data imputation in the Internet of Things sensor networks. Future Internet, 14(5):143, May 06, 2022. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
5/143.

Al-Amidie:2024:ESD

[AAAIA24] Muthana Al-Amidie, Laith Alzubaidi, Muhammad Aminul Islam, and Derek T. Anderson. Enhancing sensor data imputation: OWAbased model aggregation for missing values. Future Internet, 16(6):193, May 31, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/6/193.

Alzoubi:2022:ITB

[AAAKJ22] Yehia Ibrahim Alzoubi, Ahmad Al-Ahmad, Hasan Kahtan, and Ashraf Jaradat. Internet of Things and blockchain integration: Security, privacy, technical, and design challenges. Future Internet, 14(7):216, July 21, 2022. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
7/216.

Alharbi:2024:CSM

[AAAR24] Najla Alharbi, Bashayer Alkalifah, Ghaida Alqarawi, and Murad A. Rassam. Countering social media cybercrime using deep learning: Instagram fake accounts detection. Future Internet, 16(10), October 11, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/10/367.

Alsmadi:2024:SDC

[AAARZ24] Izzat Alsmadi, Iyad Alazzam, Mohammad Al-Ramahi, and Mohammad Zarour. Stance detection in the context of fake news — a new approach. Future Internet, 16(10), October 6, 2024. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/16/10/364.

Ahmad:2021:MVD

[AAAT21] Rasheed Ahmad, Izzat Alsmadi, Wasim Alhamdani, and Lo’ai Tawalbeh. Models versus datasets: Reducing bias through building a comprehensive IDS benchmark. Future Internet, 13(12):318, December 17, 2021. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/13/
                                        12/318.

Almuqren:2021:ESC

[AAC21] Latifah Almuqren, Fatma S. Alrayes, and Alexandra I. Cristea. An empirical study on customer churn behaviours prediction using Arabic Twitter mining approach. Future Internet, 13(7):175, July 05, 2021. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/13/
                                        7/175.

Anisetti:2012:WSA

[AAD^{+}12] Marco Anisetti, Claudio A. Ardagna, Ernesto Damiani, Fulvio Frati, Hausi A. Müller, and Atousa Pahlevan. Web service assurance: The notion and the issues. Future Internet, 4(1):92–109, February 14, 2012. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/4/1/92.

Arzo:2022:MAB

[AAE^{+}22] Sisay Tadesse Arzo, Zeinab Akhavan, Mona Esmaeili, Michael Devetsikiotis, and Fabrizio Granelli. Multiagent-based traffic prediction and traffic classification for autonomic network management systems for future networks. Future Internet, 14(8):230, July 28, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/8/230.

Alsubhi:2022:RAL

[AAF^{+}22] Khalid Alsubhi, Bander Alzahrani, Nikos Fotiou, Aiiad Albeshri, and Mohammed Alreshoodi. Reliable application layer routing using decentralized identifiers. Future Internet, 14(11):322, November 06, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/11/322.

Alibabaei:2022:RTD

[AAG^{+}22] Khadijeh Alibabaei, Eduardo Assunção, Pedro D. Gaspar, Vasco N. G. J. Soares, and João M. L. P. Caldeira. Real-time detection of vine trunk for robot localization using deep learning models developed for edge TPU devices. Future Internet, 14 (7):199, June 29, 2022. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/14/
                                        7/199.

Al-Ali:2020:DTC

[AAGB^{+}20] A. R. Al-Ali, Ragini Gupta, Tasneem Zaman Batool, Taha Landolsi, Fadi Aloul, and Ahmad Al Nabulsi. Digital twin conceptual model within the context of Internet of Things. Future Internet, 12(10):163, September 26, 2020. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/12/10/163.

Asad:2022:SKA

[AAJ^{+}22] Muhammad Asad, Muhammad Aslam, Syeda Fizzah Jilani, Saima Shaukat, and Manabu Tsukada. SHFL: K-anonymity-based secure hierarchical federated learning framework for smart healthcare systems. Future Internet, 14(11): 338, November 18, 2022. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/14/
                                        11/338.

Adkins:2024:IPI

[AAK24] Jonathan Adkins, Ali Al Bataineh, and Majd Khalaf. Identifying persons of interest in digital forensics using NLP-based AI. Future Internet, 16(11):426, November 18, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
11/426.

Aria:2023:IFD

[AAKS23] Reza Aria, Norm Archer, Moein Khanlari, and Bharat Shah. Influential factors in the design and development of a sustainable Web3/Metaverse and its applications. Future Internet, 15(4):131, March 30, 2023. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/15/4/131.

Aiosa:2021:CEE

[AALS21] Grazia Veronica Aiosa, Barbara Attanasio, Aurelio La Corte, and Marialisa Scatá. CoKnowEMe: an edge evaluation scheme for QoS of IoMT microservices in 6G scenario. Future Internet, 13(7):177, July 07, 2021. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
7/177.

Abinayaa:2024:SEC

[AAM^{+}24] Sennanur Srinivasan Abinayaa, Prakash Arumugam, Divya Bhavani Mohan, Anand Rajendran, Abderezak Lashab, Baoze Wei, and Josep M. Guerrero. Securing the edge: CatBoost classifier optimized by the lyrebird algorithm to detect denial of service attacks in Internet of Things-based wireless sensor networks. Future Internet, 16(10), October 19, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
10/381.

Agbaegbu:2021:OCC

[AAMD21] JohnBosco Agbaegbu, Oluwasefunmi Tale Arogundade, Sanjay Misra, and Robertas Damasevicius. Ontologies in cloud computing — review and future directions. Future Internet, 13(12):302, November 26, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/12/302.

Alasmari:2021:RAC

[AAN^{+}21] Ashwag Alasmari, Aseel Addawood, Mariam Nouh, Wajanat Rayes, and Areej Al-Wabil. A retrospective analysis of the COVID-19 infodemic in Saudi Arabia. Future Internet, 13(10): 254, September 30, 2021. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
10/254.

Alrubayyi:2024:STP

[AAN^{+}24] Hadeel Alrubayyi, Moudy Sharaf Alshareef, Zunaira Nadeem, Ahmed M. Abdelmoniem, and Mona Jaber. Security threats and promising solutions arising from the intersection of AI and IoT: a study of IoMT and IoET applications. Future Internet, 16(3):85, February 29, 2024. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/16/3/85.

Abdo:2019:ACR

[AAR^{+}19] Ahmad Abdo, Sadok Aouini, Bilal Riaz, Naim Ben-Hamida, and Claude D’Amours. Adaptive coherent receiver settings for optimum channel spacing in gridless optical networks. Future Internet, 11(10):206, September 25, 2019. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/11/10/206.

Albarracin-Acero:2024:VRC

[AARTSBAE24] Diego Alejandro Albarracin-Acero, Fidel Alfonso Romero-Toledo, Claudia Esperanza Saavedra-Bautista, and Edwan Anderson Ariza-Echeverri. Virtual reality in the classroom: Transforming the teaching of electrical circuits in the digital age. Future Internet, 16(8):??, August 5, 2024. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/16/
                                        8/279.

Alkhalidy:2022:NSD

[AASAI22] Muhsen Alkhalidy, Atalla Fahed Al-Serhan, Ayoub Alsarhan, and Bashar Igried. A new scheme for detecting malicious nodes in vehicular ad hoc networks based on monitoring node behavior. Future Internet, 14(8):223, July 26, 2022. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/14/
                                        8/223.

Abuzrieq:2021:EPE

[AAY21] Yara Abuzrieq, Amro Al-Said Ahmad, and Maram Bani Younes. An experimental performance evaluation of cloud-API-based applications. Future Internet, 13(12):314, December 13, 2021. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/13/
                                        12/314.

Arachchige:2023:EBN

[ABB23] Kithmini Godewatte Arachchige, Philip Branch, , and Jason But. Evaluation of blockchain networks’ scalability limitations in low-powered Internet of Things (IoT) sensor networks. Future Internet, 15(9):317, September 21, 2023. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/15/9/317.

Aziz:2023:EHE

[ABBV23] Rezak Aziz, Soumya Banerjee, Samia Bouzefrane, and Thinh Le Vinh. Exploring homomorphic encryption and differential privacy techniques towards secure federated learning paradigm. Future Internet, 15(9):310, September 13, 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
9/310.

Akhilesh:2022:APT

[ABCC22] Rohit Akhilesh, Oliver Bills, Naveen Chilamkurti, and Mohammad Jabed Morshed Chowdhury. Automated penetration testing framework for smarthome-based IoT devices. Future Internet, 14(10): 276, September 27, 2022. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
10/276.

Althamary:2024:EMT

[ABH24] Ibrahim Althamary, Rubbens Boisguene, and Chih-Wei Huang. Enhanced multi-task traffic forecasting in beyond 5G networks: Leveraging transformer technology and multi-source data fusion. Future Internet, 16(5):159, May 5, 2024. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/16/5/159.

Afrin:2022:MSA

[ABK22] Nusrat Afrin, Jason Brown, and Jamil Y. Khan. A multi-service adaptive semi-persistent LTE uplink scheduler for low power M2M devices. Future Internet, 14(4):107– ??, March 27, 2022. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/14/
                                        4/107.

Avanzato:2023:HDM

[ABLR23] Roberta Avanzato, Francesco Beritelli, Alfio Lombardo, and Carmelo Ricci. Heart DT: Monitoring and preventing cardiac pathologies using AI and IoT sensors. Future Internet, 15 (7):223–??, July 2023. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/15/
                                        7/223.

Azzoug:2020:PVR

[ABS20] Youcef Azzoug, Abdelmadjid Boukra, and Vasco N. G. J. Soares. A probabilistic VDTN routing scheme based on hybrid swarm-based approach. Future Internet, 12(11):

192, November 07, 2020. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/12/
11/192.

Ahadi:2021:OPL

[ABSG21] Alireza Ahadi, Matt Bower, Abhay Singh, and Michael Garrett. Online professional learning in response to COVID-19 — towards robust evaluation. Future Internet, 13(3):56, February 24, 2021. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/13/3/56.

Antal:2021:DLT

[ACA^{+}21] Claudia Antal, Tudor Cioara, Ionut Anghel, Marcel Antal, and Ioan Salomie. Distributed ledger technology review and decentralized applications development guidelines. Future Internet, 13(3):62, February 27, 2021. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
3/62.

Askhedkar:2023:LCU

[ACA^{+}23] Anjali R. Askhedkar, Bharat S. Chaudhari, Maha Abdelhaq, Raed Alsaqour, Rashid Saeed, and Marco Zennaro. LoRa communication using TVWS frequencies: Range and data rate. Future Internet, 15(8):270, August 14, 2023. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/15/8/270.

Awad:2020:PAS

[ACAP20] Nancy Awad, Jean-François Couchot, Bechara Al Bouna, and Laurent Philippe. Publishing anonymized set-valued data via disassociation towards analysis. Future Internet, 12(4):71, April 17, 2020. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/12/4/71.

Ali:2019:FIE

[ACBP19] Syed Mustafa Ali, Ana Filomena Curralo, Maged N. Kamel Boulos, and Sara Paiva. A framework for improving the engagement of medical practitioners in an etraining platform for tuberculosis care and prevention. Future Internet, 11 (1):6, December 28, 2019. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/11/
                                        1/6.

Ali:2022:DDA

[ACD^{+}22] Usman Ali, Giuseppe Caso, Luca De Nardis, Konstantinos Kousias, Mohammad Rajiullah, Özgü Alay, Marco Neri, Anna Brunstrom, and Maria-Gabriella Di Benedetto. Data-driven analysis of outdoor-to-indoor propagation for 5G mid-band operational networks. Future Internet, 14(8):239, August 11, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
8/239.

Almutairi:2024:RBT

[ACF24] Abeer Almutairi, Xavier Carpent, and Steven Furnell. Recommendationbased trust evaluation model for the Internet of Underwater Things. Future Internet, 16(9), September 23, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
9/346.

Alberro:2022:EED

[ACG22] Leonardo Alberro, Alberto Castro, and Eduardo Grampin. Experimentation environments for data center routing protocols: a comprehensive review. Future Internet, 14(1):29, January 17, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
1/29.

Achkouty:2024:RIQ

[ACG^{+}24] Fouad Achkouty, Richard Chbeir, Laurent Gallon, Elio Mansour, and Antonio Corral. Resource indexing and querying in large connected environments. Future Internet, 16 (1):15, December 30, 2024. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/16/
                                        1/15.

Angelini:2016:SLL

[ACK^{+}16] Leonardo Angelini, Stefano Carrino, Omar Abou Khaled, Susie Riva-Mossman, and Elena Mugellini. Senior living lab: An ecological approach to Foster social innovation in an ageing society. Future Internet, 8(4):50, October 21, 2016. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/8/4/50.

Amadeo:2019:EVA

[ACM^{+}19] Marica Amadeo, Claudia Campolo, Antonella Molinaro, Jerome Harri, Christian Esteve Rothenberg, and Alexey Vinel. Enhancing the 3GPP V2X architecture with informationcentric networking. Future Internet, 11(9):199, September 18, 2019. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/11/
                                        9/199.

Ahuja:2024:SST

[ACS24] Rohit Ahuja, Sahil Chugh, and Raman Singh. Seed-Chain: A secure and transparent blockchain-driven framework to revolutionize the seed supply chain. Future Internet, 16(4):132, April 15, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
4/132.

Anwar:2018:ESS

[AD18] Nazia Anwar and Huifang Deng. Elastic scheduling of scientific workflows under deadline constraints in cloud computing environments. Future Internet, 10 (1):5, January 07, 2018. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/10/
1/5.

Abdo:2020:APP

[AD20] Ahmad Abdo and Claude D’Amours. Adaptive pre/ post-compensation of cascade filters in coherent optical transponders. Future Internet, 12(2):21, January 24, 2020. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/12/2/21.

Aldabbagh:2021:RCD

[ADAB21] Ghadah Aldabbagh, Nikos Dimitriou, Samar Alkhuraiji, and Omaimah Bamasag. Radio coverage and device capacity dimensioning methodologies for IoT LoRaWAN and NB-IoT deployments in urban environments. Future Internet, 13(6):144, May 30, 2021. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
6/144.

Alhaj:2022:NTC

[ADAq^{+}22] Yousif A. Alhaj, Abdelghani Dahou, Mohammed A. A. Al-qaness, Laith Abualigah, Aaqif Afzaal Abbasi, Nasser Ahmed Obad Almaweri, Mohamed Abd Elaziz, and Robertas Damasevicius. A novel text classification technique using improved particle swarm optimization: a case study of Arabic language. Future Internet, 14(7):194, June 27, 2022. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
7/194.

Augusto:2024:PRO

[ADCM24] Gonçalo Augusto, Rui Duarte, Carlos Cunha, and Ana Matos. Pattern recognition in older adults’ activities of daily living. Future Internet, 16(12): 476, December 20, 2024. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
12/476.

Autili:2019:CBC

[ADG^{+}19] Marco Autili, Amleto Di Salle, Francesco Gallo, Claudio Pompilio, and Massimo Tivoli. A choreography-based and collaborative road mobility system for L’Aquila City. Future Internet, 11 (6):132, June 14, 2019. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
6/132.

Adibi:2023:MFA

[Adi23] Sasan Adibi. The mPOC framework: An autonomous outbreak prediction and monitoring platform based on wearable IoMT approach. Future Internet, 15(8):257, July 30, 2023. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
8/257.

Ali:2024:HCN

[ADK^{+}24] Mohammad Imran Ali, Shahi Dost, Khurram Shehzad Khattak, Muhammad Imran Khan, and Riaz Muhammad. Harnessing the cloud: a novel approach to smart solar plant monitoring. Future Internet, 16(6):191, May 29, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/6/191.

Abro:2019:LEE

[ADM19a] Adeel Abro, Zhongliang Deng, and Kamran Ali Memon. A lightweight elliptic-Elgamal-based authentication scheme for secure device-to-device communication. Future Internet, 11(5):108, May 07, 2019. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/11/5/108.

Abro:2019:DAP

[ADM^{+}19b] Adeel Abro, Zhongliang Deng, Kamran Ali Memon, Asif Ali Laghari, Khalid Hussain Mohammadani, and Noor ul Ain. A dynamic application-partitioning algorithm with improved offloading mechanism for fog cloud networks. Future Internet, 11(7):141, June 28, 2019. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/11/7/141.

Abonizio:2020:LIF

[AdMTJ20] Hugo Queiroz Abonizio, Janaina Ignacio de Morais, Gabriel Marques Tavares, and Sylvio Barbon Junior. Language-independent fake news detection: English, Portuguese, and Spanish mutual features. Future Internet, 12(5):87, May 11, 2020. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/12/5/87.

Ali:2020:TFA

[ADS20] Guma Ali, Mussa Ally Dida, and Anael Elikana Sam. Two-factor authentication scheme for mobile money: a review of threat models and countermeasures. Future Internet, 12(10):160, September 24, 2020. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/12/10/160.

Ali:2021:SEM

[ADS21] Guma Ali, Mussa Ally Dida, and Anael Elikana Sam. A secure and efficient multi-factor authentication algorithm for mobile money applications. Future Internet, 13(12): 299, November 25, 2021. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
12/299.

Al-Dwairi:2022:RRS

[ADSAKAD22] Mahmoud Al-Dwairi, Ahmed S.

Shatnawi, Osama Al-Khaleel, and Basheer Al-Duwairi. Ransomwareresilient self-healing XML documents. Future Internet, 14(4):115–??, April 07, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/4/115.

Ali:2023:ETN

[AEA^{+}23] Shrouk A. Ali, Shaimaa Ahmed Elsaid, Abdelhamied A. Ateya, Mohammed ElAffendi, and Ahmed A. Abd El-Latif. Enabling technologies for next-generation smart cities: a comprehensive review and research directions. Future Internet, 15(12):398, December 09, 2023. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/15/12/398.

AlZubi:2023:DMS

[AEM^{+}23] Shadi AlZu’bi, Mohammad Elbes, Ala Mughaid, Noor Bdair, Laith Abualigah, Agostino Forestiero, and Raed Abu Zitar. Diabetes monitoring system in smart health cities based on big data intelligence. Future Internet, 15(2):85, February 20, 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
2/85.

Alghamdi:2021:STP

[AES21] Taghreed Alghamdi, Khalid Elgazzar, and Taysseer Sharaf. Spatiotemporal traffic prediction using hierarchical Bayesian modeling. Future Internet, 13 (9):225, August 30, 2021. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
9/225.

Alzahrani:2021:SSB

[AF21] Bander Alzahrani and Nikos Fotiou. Securing SDN-based IoT group communication. Future Internet, 13(8):207, August 09, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/8/207.

Atif:2022:TMM

[AF22] Muhammad Atif and Valentina Franzoni. Tell me more: Automating emojis classification for better accessibility and emotional context recognition. Future Internet, 14(5):142, May 05, 2022. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
5/142.

Aqlan:2019:IAC

[AFAAM19] Fares Aqlan, Xiaoping Fan, Abdullah Alqwbani, and Akram Al-Mansoub. Improved Arabic–Chinese machine translation with linguistic input features. Future Internet, 11(1):22, January 19, 2019. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
1/22.

AlJawarneh:2023:EIH

[AFB23] Isam Mashhour Al Jawarneh, Luca Foschini, and Paolo Bellavista. Efficient integration of heterogeneous mobility-pollution big data for joint analytics at scale with QoS guarantees. Future Internet, 15 (8):263, August 07, 2023. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/15/
                                        8/263.

Ahamed:2022:IMB

[AFS^{+}22] Farhad Ahamed, Farnaz Farid, Basem Suleiman, Zohaib Jan, Luay A. Wahsheh, and Seyed Shahrestani. An intelligent multimodal biometric authentication model for personalised healthcare services. Future Internet, 14(8):222, July 26, 2022. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/14/
                                        8/222.

Azeroual:2023:ESG

[AFS^{+}23] Otmane Azeroual, Renaud Fabre, Uta Störl, , and Ruidong Qi. Elastic stack and GRAPHYP knowledge graph of Web usage: a win–win workflow for semantic interoperability in decision making. Future Internet, 15(6): 190–??, June 2023. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/15/
                                        6/190.

Alam:2022:FLR

[AG22] Tanweer Alam and Ruchi Gupta. Federated learning and its role in the privacy preservation of IoT devices. Future Internet, 14 (9):246, August 23, 2022. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
9/246.

Assuncao:2022:RTI

[AGA^{+}22] Eduardo Assunção, Pedro D. Gaspar, Khadijeh Alibabaei, Maria P. Simões, Hugo Proença, Vasco N. G. J. Soares, and João M. L. P. Caldeira. Real-time image detection for edge devices: a peach fruit detection application. Future Internet, 14(11): 323, November 08, 2022. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
11/323.

Altuntas:2022:MEV

[AGB22] Erkin Altuntas, Peter A.

Gloor, and Pascal Budner. Measuring ethical values with AI for better teamwork. Future Internet, 14 (5):133, April 27, 2022. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
5/133.

Alshathri:2018:SLM

[AGC18] Samah Alshathri, Bogdan Ghita, and Nathan Clarke. Sharing with live migration energy optimization scheduler for cloud computing data centers. Future Internet, 10(9):86, September 06, 2018. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/10/
                                        9/86.

Axelsson:2021:LHS

[AGN21] Carl-Anton Werner Axelsson, Mona Guath, and Thomas Nygren. Learning how to separate fake from real news: Scalable digital tutorials promoting students’ civic online reasoning. Future Internet, 13(3):60, February 27, 2021. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/13/3/60.

AlLail:2023:MLN

[AGO23a] Mustafa Al Lail, Alejandro Garcia, and Saul Olivo. Machine learning for network intrusion detection — a comparative study. Future Internet, 15(7): 243–??, July 2023. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/15/
                                        7/243.

Arici:2023:RWI

[AGO^{+}23b] Nicola Arici, Alfonso Emilio Gerevini, Matteo Olivato, Luca Putelli, Luca Sigalini, and Ivan Serina. Real-world implementation and integration of an automatic scoring system for workplace safety courses in Italian. Future Internet, 15(8):268, August 12, 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
8/268.

Arca:2021:AAP

[AH21] Sevgi Arca and Rattikorn Hewett. Analytics on anonymity for privacy retention in smart health data. Future Internet, 13(11):274, October 28, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/11/274.

Abdulsalam:2022:SPC

[AH22] Yunusa Simpa Abdulsalam and Mustapha Hedabou. Security and privacy in cloud computing: Technical review. Future Internet, 14(1):11, December 27, 2022. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/14/1/11.

Alsini:2021:HRM

[AHD21] Areej Alsini, Du Q.

Huynh, and Amitava Datta. Hashtag recommendation methods for Twitter and Sina Weibo: a review. Future Internet, 13(5):129, May 14, 2021. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
5/129.

Aoyagi:2023:API

[AHH^{+}23] Shunji Aoyagi, Yuki Horie, Do Thi Thu Hien, Thanh Duc Ngo, Duy-Dinh Le, Kien Nguyen, and Hiroo Sekiya. An accurate platform for investigating TCP performance in Wi-Fi networks. Future Internet, 15 (7):246–??, July 2023. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/15/
                                        7/246.

Al-Hamid:2023:NSB

[AHKC23] Duaa Zuhair Al-Hamid, Pejman A. Karegar, and Peter Han Joo Chong. A novel SDWSN-based testbed for IoT smart applications. Future Internet, 15(9):291, August 28, 2023. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/15/9/291.

Astaneh:2022:TOB

[AHMI22] Saeed A. Astaneh, Shahram Shah Heydari, Sara Taghavi Motlagh, and Alireza Izaddoost. Trade-offs between risk and operational cost in SDN failure recovery plan. Future Internet, 14(9):263, September 13, 2022. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/14/9/263.

Asif:2023:IBB

[AHP23] Rameez Asif, Syed Raheel Hassan, and Gerard Parr. Integrating a blockchain-based governance framework for responsible AI. Future Internet, 15(3):97, February 28, 2023. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/15/3/97.

Avazov:2023:FFD

[AHS^{+}23] Kuldoshbay Avazov, An Eui Hyun, Alabdulwahab Abrar Sami S, Azizbek Khaitov, Akmalbek Bobomirzaevich Abdusalomov, and Young Im Cho. Forest fire detection and notification method based on AI and IoT approaches. Future Internet, 15(2):61, January 31, 2023. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/15/2/61.

Abella:2024:MLA

[AIP^{+}24] Vincent Abella, Johnfil Initan, Jake Mark Perez, Philip Virgil Astillo, Jr. Luis Gerardo Cañete, and Gaurav Choudhary. Machine learning-assisted dynamic proximity-driven sorting algorithm for supermarket navigation optimization: a simulationbased validation. Future Internet, 16(8):??, August 2, 2024. CODEN ????

ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/16/8/277.

Augello:2021:SEE

[AIPV21] Agnese Augello, Ignazio Infantino, Giovanni Pilato, and Gianpaolo Vitale. Site experience enhancement and perspective in cultural heritage fruition — a survey on new technologies and methodologies based on a “Four-Pillars” approach. Future Internet, 13(4):92, April 04, 2021. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/13/
                                        4/92.

Ali:2018:PSI

[AIR^{+}18] Shaukat Ali, Naveed Islam, Azhar Rauf, Ikram Ud Din, Mohsen Guizani, and Joel J. P. C. Rodrigues. Privacy and security issues in online social networks. Future Internet, 10(12): 114, November 22, 2018. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/10/
                                        12/114.

Abbasi:2018:VMA

[AJ18] Aaqif Afzaal Abbasi and Hai Jin. v-Mapper: An application-aware resource consolidation scheme for cloud data centers. Future Internet, 10(9):90, September 15, 2018. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/10/
9/90.

Adel:2024:WSS

[AJ24] Amr Adel and Tony Jan.

Watch the skies: a study on drone attack vectors, forensic approaches, and persisting security challenges. Future Internet, 16 (7):??, July 13, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
7/250.

Abdelmoniem:2024:TDC

[AJA^{+}24] Ahmed M. Abdelmoniem, Mona Jaber, Ali Anwar, Yuchao Zhang, and Mingliang Gao. Towards a decentralized collaborative framework for scalable edge AI. Future Internet, 16(11):421, November 14, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/11/421.

Qahtani:2023:MAC

[AJT^{+}23] Elham Al Qahtani, Yousra Javed, Sarah Tabassum, Lipsarani Sahoo, and Mohamed Shehab. Managing access to confidential documents: a case study of an email security tool. Future Internet, 15(11):356, October 28, 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.

                                        mdpi.com/1999-5903/15/
                                        11/356.

Abbasi:2018:RVV

[AK18] Irshad Ahmed Abbasi and Adnan Shahid Khan. A review of vehicle to vehicle communication protocols for VANETs in the urban environment. Future Internet, 10(2):14, January 31, 2018. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/10/2/14.

Aboagye:2019:SEC

[AK19] Emelia Opoku Aboagye and Rajesh Kumar. Simple and efficient computational intelligence strategies for effective collaborative decisions. Future Internet, 11(1):24, January 21, 2019. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/11/1/24.

Al-Khafaji:2022:IQI

[AK22] Hamza Mohammed Ridha Al-Khafaji. Improving quality indicators of the cloud-based IoT networks using an improved form of seagull optimization algorithm. Future Internet, 14(10):281, September 29, 2022. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/14/10/281.

Alazab:2022:DOM

[AKAS22] Ammar Alazab, Ansam Khraisat, Moutaz Alazab, and Sarabjot Singh. Detection of obfuscated malicious JavaScript code. Future Internet, 14(8): 217, July 22, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
8/217.

Aftowicz:2024:NPU

[AKDL24] Marcin Aftowicz, Ievgen Kabin, Zoya Dyka, and Peter Langendörfer. Nonprofiled unsupervised horizontal iterative attack against hardware elliptic curve scalar multiplication using machine learning. Future Internet, 16(2):45, January 29, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
2/45.

Ausaf:2020:WAC

[AKJB20] Asfund Ausaf, Mohammad Zubair Khan, Muhammad Awais Javed, and Ali Kashif Bashir. WLAN aware cognitive medium access control protocol for IoT applications. Future Internet, 12(1):11, January 11, 2020. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/12/1/11.

Adarbah:2024:NFE

[AKK^{+}24] Haitham Y. Adarbah, Mehmet Sabir Kiraz, Suleyman Kardas, Ali H. Al-Bayatti, and Hilal M. Y. Al-Bayatti. A new framework for enhancing VANETs through Layer 2 DLT architectures with multiparty threshold key management and PETs. Future Internet, 16(9), September 9, 2024. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
9/328.

Asad:2022:SMS

[AKM22] Arghavan Asad, Rupinder Kaur, and Farah Mohammadi. A survey on memory subsystems for deep neural network accelerators. Future Internet, 14 (5):146, May 10, 2022. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
5/146.

Ahmed:2023:EED

[AKM^{+}23] Sheeraz Ahmed, Zahoor Ali Khan, Syed Muhammad Mohsin, Shahid Latif, Sheraz Aslam, Hana Mujlid, Muhammad Adil, and Zeeshan Najam. Effective and efficient DDoS attack detection using deep learning algorithm, multilayer perceptron. Future Internet, 15(2):76, February 15, 2023. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/15/2/76.

Adesokan:2024:COC

[AKT24] Adedamola Adesokan, Rowan Kinney, and Eirini Eleni Tsiropoulou. CROWD-MATCH: Optimizing crowdsourcing matching through the integration of matching theory and coalition games. Future Internet, 16 (2):58, February 11, 2024. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
2/58.

Abesinghe:2023:ICT

[AKYP23] Sandulika Abesinghe, Nayomi Kankanamge, Tan Yigitcanlar, and Surabhi Pancholi. Image of a city through big data analytics: Colombo from the lens of geo-coded social media data. Future Internet, 15 (1):32, January 09, 2023. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
1/32.

AlMojamed:2020:IIM

[Al 20] Mohammad Al Mojamed.

Integrating IP mobility management protocols and MANET: A survey. Future Internet, 12(9):150, September 03, 2020. CO-DEN ???? ISSN 1999-

5903. URL https://www.

                                        mdpi.com/1999-5903/12/
                                        9/150.

Alabdan:2020:PAS

[Ala20] Rana Alabdan. Phishing attacks survey: Types, vectors, and technical approaches. Future Internet, 12(10):168, September 30, 2020. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/12/10/168.

Albahli:2019:DEL

[Alb19] Saleh Albahli. A deep ensemble learning method for effort-aware just-in-time defect prediction. Future Internet, 11(12):246, November 20, 2019. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/11/
                                        12/246.

Albeshri:2021:IHB

[Alb21] Aiiad Albeshri. An image hashing-based authentication and secure group communication scheme for IoT-enabled MANETs. Future Internet, 13(7):166, June 27, 2021. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/13/
                                        7/166.

Ahmad:2024:TCI

[ALB24] Ehsan Ahmad, Brian Larson, and Abdulbasid Banga. Trusted composition of Internet of Medical Things over imperfect networks. Future Internet, 16 (7):??, June 28, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
7/230.

Ashouri:2019:ECS

[ALDS19] Majid Ashouri, Fabian Lorig, Paul Davidsson, and Romina Spalazzese. Edge computing simulators for IoT system design: An analysis of qualities and metrics. Future Internet, 11(11):235, November 08, 2019. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
11/235.

Alhazmi:2022:DSP

[Alh22] Huda Alhazmi. Detection of students’ problems in distance education using topic modeling and machine learning. Future Internet, 14(6):170, May 31, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/6/170.

Aung:2024:EMS

[ALH24] Aye Nyein Aung, Che-Wei Liao, and Jeih-Weih Hung. Effective monoaural speech separation through convolutional top-down multi-view network. Future Internet, 16(5):151, April 28, 2024. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/16/
                                        5/151.

Alghamdi:2024:CDF

[ALL24] Jawaher Alghamdi, Yuqing Lin, and Suhuai Luo. Cross-domain fake news detection using a promptbased approach. Future Internet, 16(8):??, August 8, 2024. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/16/8/286.

Alti:2016:ASB

[ALLR16] Adel Alti, Abderrahim Lakehal, Sébastien Laborie, and Philippe Roose. Autonomic semantic-based context-aware platform for mobile applications in pervasive environments. Future Internet, 8(4):48, September 29, 2016. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/8/
                                        4/48.

Alnaim:2022:MPT

[Aln22] Abdulrahman K. Alnaim. Misuse patterns from the threat of modification of non-control data in network function virtualization. Future Internet, 14 (7):201, June 30, 2022. CODEN ???? ISSN 1999-

                                        5903. URL https://www.

mdpi.com/1999-5903/14/
7/201.

AlQahtani:2023:TOC

[AlQ23] Salman Ali AlQahtani. Towards an optimal cloudbased resource management framework for nextgeneration Internet with multi-slice capabilities. Future Internet, 15(10): 343, October 19, 2023. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
10/343.

Aoyagi:2020:MAP

[ALST20] Saizo Aoyagi, Yiping Le, Tetsuo Shimizu, and Kazuki Takahashi. Mobile application to provide traffic congestion estimates and tourism spots to promote additional stopovers. Future Internet, 12(5):83, April 29, 2020. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/12/5/83.

Adeke:2023:SNT

[ALZ^{+}23] James Msughter Adeke, Guangjie Liu, Junjie Zhao, Nannan Wu, and Hafsat Muhammad Bashir. Securing network traffic classification models against adversarial examples using derived variables. Future Internet, 15(12):405, December 16, 2023. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/15/12/405.

Almaaitah:2021:RTC

[AM21] Abdallah Y. Alma’aitah and Mohammad A. Massad. Reader-tag commands via modulation cutoff intervals in RFID systems. Future Internet, 13(9):235, September 16, 2021. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/13/9/235.

AlAsqah:2023:FLB

[AM23] Muneerah Al Asqah and Tarek Moulahi. Federated learning and blockchain integration for privacy protection in the Internet of Things: Challenges and solutions. Future Internet, 15(6):203–??, June 2023. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/15/
                                        6/203.

Aggrawal:2024:TCM

[AM24] Sakhi Aggrawal and Alejandra J. Magana. Teamwork conflict management training and conflict resolution practice via large language models. Future Internet, 16(5):177, May 19, 2024. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/16/5/177.

Al-Maroof:2021:UAS

[AMAA^{+}21] Rana Saeed Al-Maroof, Khadija Alhumaid, Ahmad Qasim Alhamad, Ahmad Aburayya, and Said Salloum. User acceptance of smart watch for medical purposes: an empirical study. Future Internet, 13(5):127, May 12, 2021. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
5/127.

Alghamdi:2022:CST

[AMAE22] Taghreed Alghamdi, Sifatul Mostafi, Ghadeer Abdelkader, and Khalid Elgazzar. A comparative study on traffic modeling techniques for predicting and simulating traffic behavior. Future Internet, 14(10):294, October 15, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/10/294.

AbuKhousa:2012:HCO

[AMAJ12] Eman AbuKhousa, Nader Mohamed, and Jameela Al-Jaroodi. e-health cloud: Opportunities and challenges. Future Internet, 4 (3):621–645, July 04, 2012. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/4/
3/621.

Almani:2024:EVV

[AMC^{+}24] Dimah Almani, Tim Muller, Xavier Carpent, Takahito Yoshizawa, and Steven Furnell. Enabling vehicleto-vehicle trust in rural areas: an evaluation of a pre-signature scheme for infrastructure-limited environments. Future Internet, 16(3):77, February 26, 2024. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/16/3/77.

Ali:2017:CSR

[AMCI17] Ahmad Ali, Yu Ming, Sagnik Chakraborty, and Saima Iram. A comprehensive survey on realtime applications of WSN. Future Internet, 9(4):77, November 07, 2017. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/9/
4/77.

Akbar:2021:DHD

[AMEF21] Nur Arifin Akbar, Amgad Muneer, Narmine El-Hakim, and Suliman Mohamed Fati. Distributed hybrid double-spending attack prevention mechanism for proof-of-work and proof-of-stake blockchain consensuses. Future Internet, 13(11):285, November 12, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/11/285.

Arkoulis:2010:MSC

[AMF^{+}10] Stamatios Arkoulis, Giannis F. Marias, Pantelis A. Frangoudis, Jens Oberender, Alexandru Popescu, Markus Fiedler, Hermann de Meer, and George C. Polyzos. Misbehavior scenarios in cognitive radio networks. Future Internet, 2(3):212–237, July 29, 2010. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/2/3/212.

Adou:2022:MAP

[AMG22] Yves Adou, Ekaterina Markova, and Yuliya Gaidamaka. Modeling and analyzing preemptionbased service prioritization in 5G networks slicing framework. Future Internet, 14(10):299, October 18, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/10/299.

Ashraf:2023:LPP

[AMI23] Zeeshan Ashraf, Zahid Mahmood, and Muddesar Iqbal. Lightweight privacy-preserving remote user authentication and key agreement protocol for next-generation IoT-based smart healthcare. Future Internet, 15(12):386, November 29, 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
12/386.

Abdellah:2021:MLA

[AMK^{+}21] Ali R. Abdellah, Omar Abdulkareem Mahmood, Ruslan Kirichek, Alexander Paramonov, and Andrey Koucheryavy. Machine learning algorithm for delay prediction in IoT and tactile Internet. Future Internet, 13(12):304, November 26, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/12/304.

Ajami:2022:FSA

[AML22] Hicham Ajami, Hamid Mcheick, and Catherine Laprise. First steps of asthma management with a personalized ontology model. Future Internet, 14(7):190, June 22, 2022. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
7/190.

Akbar:2021:ODP

[AMR^{+}21] Zaenal Akbar, Hani Febri Mustika, Dwi Setyo Rini, Lindung Parningotan Manik, Ariani Indrawati, Agusdin Dharma Fefirenta, and Tutie Djarwaningsih. An ontology-driven personalized faceted search for exploring knowledge bases of Capsicum. Future Internet, 13(7):172, June 30, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/7/172.

Andrade:2023:BBS

[AMSM23] Daniel Andrade, Roberto Magueta, Adão Silva, and Paulo Marques. Beamforming based on a SSS angle estimation algorithm for 5G NR networks. Future Internet, 15(3):105, March 09, 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
3/105.

Ayaida:2022:TMR

[AMVM22] Marwane Ayaida, Nadhir Messai, Frederic Valentin, and Dimitri Marcheras. TalkRoBots: a middleware for robotic systems in Industry 4.0. Future Internet, 14(4):109–??, March 29, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/4/109.

Asano:2024:ITT

[AMY24] Mizuki Asano, Takumi Miyoshi, and Taku Yamazaki. Internet-of-Things traffic analysis and device identification based on two-stage clustering in smart home environments. Future Internet, 16(1):17, December 31, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
1/17.

An:2020:ADP

[AMZ^{+}20] Wen An, Jun Jie Ma, Hong Yang Zhou, Hong Shan Chen, Xu Jun, and Xu Jian. An adaptive differential protection and fast auto-closing system for 10 kV distribution networks based on 4G LTE wireless communication. Future Internet, 12(1):2, December 20, 2020. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/12/
1/2.

Alaoui:2022:DLV

[AN22] Rokia Lamrani Alaoui and El Habib Nfaoui. Deep learning for vulnerability and attack detection on Web applications: a systematic literature review. Future Internet, 14 (4):118–??, April 13, 2022. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
4/118.

Aggarwal:2023:FLN

[AN23] Shobhit Aggarwal and Asis Nasipuri. FL-LoRaMAC: a novel framework for enabling ondevice learning for LoRabased IoT applications.

Future Internet, 15(9):307, September 10, 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
9/307.

Al-Naime:2021:RMM

[ANAAM21] Khalid Al-Naime, Adnan Al-Anbuky, and Grant Mawston. Remote monitoring model for the preoperative prehabilitation program of patients requiring abdominal surgery. Future Internet, 13(5):104, April 22, 2021. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
5/104.

Anderson:2021:GSV

[And21] David P. Anderson. Globally scheduling volunteer computing. Future Internet, 13(9):229, August 31, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/9/229.

Al-Nidawi:2024:MUO

[ANHMS24] Yaarob Al-Nidawi, Haider Tarish Haider, Dhiaa Halboot Muhsen, and Ghadeer Ghazi Shayea. Multi-user optimal load scheduling of different objectives combined with multi-criteria decision making for smart grid. Future Internet, 16 (10), September 29, 2024.

CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/16/
                                        10/355.

Antley:2012:TDA

[Ant12] Jeremy Antley. Textual dualism and augmented reality in the Russian empire. Future Internet, 4 (4):1037–1048, December 10, 2012. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/4/4/1037.

Anton:2024:SBT

[Ant24] Simon D. Duque Anton. Software-bus-toolchain (SBT): Introducing a versatile method for quickly implementing (I)IoT-scenarios. Future Internet, 16(7):??, July 3, 2024. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/16/7/237.

Aragones:2024:TGP

[AOF24] Raúl Aragonés, Joan Oliver, and Carles Ferrer. Thermoelectric generatorpowered long-range wireless real-time steam leak detection in steam traps. Future Internet, 16(12): 474, December 19, 2024. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/16/
                                        12/474.

Abikoye:2023:SCU

[AOI^{+}23] Oluwakemi Christiana Abikoye, Esau Taiwo Oladipupo, Agbotiname Lucky Imoize, Joseph Bamidele Awotunde, Cheng-Chi Lee, and Chun-Ta Li. Securing critical user information over the Internet of Medical Things platforms using a hybrid cryptography scheme. Future Internet, 15(3):99, February 28, 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
3/99.

Arena:2019:OVC

[AP19] Fabio Arena and Giovanni Pau. An overview of vehicular communications. Future Internet, 11(2):27, January 24, 2019. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/11/2/27.

Alghamdi:2021:PIS

[AP21] Abdulelah A. Alghamdi and Margaret Plunkett. The perceived impact of social networking sites and apps on the social capital of Saudi postgraduate students: a case study. Future Internet, 13(1):20, January 16, 2021. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/13/1/20.

Adoga:2022:NFV

[AP22] Haruna Umar Adoga and Dimitrios P. Pezaros. Network function virtualization and service function chaining frameworks: a comprehensive review of requirements, objectives, implementations, and open research challenges. Future Internet, 14 (2):59, February 15, 2022. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
2/59.

Araujo:2022:OSA

[APAR22] Andreia Araújo, Filipe Portela, Filipe Alvelos, and Saulo Ruiz. Optimization of the system of allocation of overdue loans in a Sub-Saharan Africa microfinance institution. Future Internet, 14(6):163, May 27, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/6/163.

Ali:2016:OSD

[APB^{+}16] Syed Mustafa Ali, Rachel Powers, Jeffrey Beorse, Arif Noor, Farah Naureen, Naveed Anjum, Muhammad Ishaq, Javariya Aamir, and Richard Anderson. ODK scan: Digitizing data collection and impacting data management processes in Pakistan’s tuberculosis control program.

Future Internet, 8(4):51, October 24, 2016. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/8/4/51.

Aslan:2024:SBE

[APJ24] Lara Aslan, Michal Ptaszynski, and Jukka Jauhiainen. Are strong baselines enough? false news detection with machine learning. Future Internet, 16(9), September 5, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
9/322.

Aouedi:2022:ITM

[APP22] Ons Aouedi, Kandaraj Piamrat, and Benoît Parrein. Intelligent traffic management in next-generation networks. Future Internet, 14(2):44, January 28, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/2/44.

Alotaibi:2023:AML

[AR23] Afnan Alotaibi and Murad A. Rassam. Adversarial machine learning attacks against intrusion detection systems: a survey on strategies and defense. Future Internet, 15(2):62, January 31, 2023. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/15/2/62.

Al-Rahayfeh:2019:NAT

[ARAAA19] Amer Al-Rahayfeh, Saleh Atiewi, Abdullah Abuhussein, and Muder Almiani. Novel approach to task scheduling and load balancing using the dominant sequence clustering and mean shift clustering algorithms. Future Internet, 11(5):109, May 08, 2019. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
5/109.

Amadeo:2019:FCI

[ARC^{+}19] Marica Amadeo, Giuseppe Ruggeri, Claudia Campolo, Antonella Molinaro, Valeria Loscrí, and Carlos T. Calafate. Fog computing in IoT smart environments via named data networking: a study on service orchestration mechanisms. Future Internet, 11(11):222, October 24, 2019. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/11/11/222.

Artiemjew:2020:ARB

[ARM20] Piotr Artiemjew, Lada Rudikova, and Oleg Myslivets. About rule-based systems: Single database queries for decision making. Future Internet, 12 (12):212, November 27, 2020. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/12/12/212.

Aman:2015:EEP

[AS15] Waqas Aman and Einar Snekkenes. EDAS: An evaluation prototype for autonomic event-driven adaptive security in the Internet of Things. Future Internet, 7(3):225– 256, July 08, 2015. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/7/
3/225.

Aldawood:2019:RCS

[AS19] Hussain Aldawood and Geoffrey Skinner. Reviewing cyber security social engineering training and awareness programspitfalls and ongoing issues. Future Internet, 11(3):73, March 18, 2019. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/11/3/73.

Assiri:2024:EPA

[AS24] Azizah Assiri and Hassen Sallay. Efficient privacy-aware forwarding for enhanced communication privacy in opportunistic mobile social networks. Future Internet, 16(2):48, January 31, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
2/48.

AlKhasawneh:2023:CAT

[ASAH^{+}23] Mohammad Al Khasawneh, Abdel-Aziz Ahmad Sharabati, Shafig Al-Haddad, Rania Al-Daher, Sarah Hammouri, and Sima Shaqman. Consumer’s attitude towards display Google Ads. Future Internet, 15(4):145, April 07, 2023. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/15/4/145.

Al-Saedi:2022:FCE

[ASBC22] Ahmed A. Al-Saedi, Veselka Boeva, and Emiliano Casalicchio. FedCO: Communication-efficient federated learning via clustering optimization. Future Internet, 14(12):377, December 13, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
12/377.

Akter:2022:FRF

[ASCM22] Mst. Shapna Akter, Hossain Shahriar, Reaz Chowdhury, and M. R. C. Mahdy. Forecasting the risk factor of frontier markets: a novel stacking ensemble of neural network approach. Future Internet, 14(9):252, August 25, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
9/252.

Anoh:2022:WCS

[ASD^{+}22] Kelvin Anoh, Chan Hwang See, Yousef Dama, Raed A. Abd-Alhameed, and Simeon Keates. 6G wireless communication systems: Applications, opportunities and challenges. Future Internet, 14(12):379, December 15, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
12/379.

Alfian:2023:URF

[ASF^{+}23] Ganjar Alfian, Muhammad Syafrudin, Norma Latif Fitriyani, Sahirul Alam, Dinar Nugroho Pratomo, Lukman Subekti, Muhammad Qois Huzyan Octava, Ninis Dyah Yulianingsih, Fransiskus Tatas Dwi Atmaji, and Filip Benes. Utilizing random forest with iForest-based outlier detection and SMOTE to detect movement and direction of RFID tags. Future Internet, 15(3):103, March 08, 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
3/103.

Anton-Sancho:2024:UVR

[ASFAAV24] Álvaro Antón-Sancho, Pablo Fernández-Arias, Edwan Anderson Ariza, and Diego Vergara. The use of virtual reality in the countries of the Central American Bank for Economic Integration (CABEI). Future Internet, 16(7):??, July 13, 2024. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/16/7/249.

Anton-Sancho:2023:PUV

[ASFAV23] Álvaro Antón-Sancho, Pablo Fernández-Arias, and Diego Vergara. Perception of the use of virtual reality didactic tools among faculty in Mexico. Future Internet, 15(2):72, February 12, 2023. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/15/2/72.

Auer:2024:CAC

[ASG24] Isabel Auer, Stephan Schlögl, and Gundula Glowka. Chatbots in airport customer service — exploring use cases and technology acceptance. Future Internet, 16(5):175, May 17, 2024. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/16/
                                        5/175.

Abdelreheim:2023:POR

[ASK23] Marwa Abdelreheim, Taysir Hassan A. Soliman, and Friederike Klan. A personalized ontology recommendation system to effectively support ontology development by reuse. Future Internet, 15(10):331, October 07, 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
10/331.

Anwar:2017:QSB

[ASL17] Asim Anwar, Boon-Chong Seet, and Xue Jun Li. Quality of service based NOMA group D2D communications. Future Internet, 9(4):73, November 01, 2017. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/9/4/73.

Almeida:2022:EBC

[ASL22] Fernando Almeida, Jorge Simões, and Sérgio Lopes. Exploring the benefits of combining DevOps and Agile. Future Internet, 14 (2):63, February 19, 2022. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
2/63.

Almudayni:2024:IIM

[ASL24] Ziyad Almudayni, Ben Soh, and Alice Li. IMBA: IoT-mist bat-inspired algorithm for optimising resource allocation in IoT networks. Future Internet, 16(3):93, March 08, 2024. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
3/93.

Andelic:2022:DMW

[ASLG22] Nikola Andelić, Sandi Baressi Segota, Ivan Lorencin, and Matko Glucina. Detection of malicious Websites using symbolic classifier. Future Internet, 14(12):358, November 29, 2022. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/14/12/358.

Ali:2024:CSA

[ASM^{+}24] G. G. Md. Nawaz Ali, Mohammad Nazmus Sadat, Md Suruz Miah, Sameer Ahmed Sharief, and Yun Wang. A comprehensive study and analysis of the third generation partnership Project’s 5G new radio for vehicleto-everything communication. Future Internet, 16 (1):21, January 06, 2024. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/16/
                                        1/21.

Alowish:2022:TLA

[ASMM22] Mazen Alowish, Yoshiaki Shiraishi, Masami Mohri, and Masakatu Morii. Three layered architecture for driver behavior analysis and personalized assistance with alert message dissemination in 5G envisioned fog-IoCV. Future Internet, 14(1):12, December 27, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
1/12.

Ang:2022:TCI

[ASN22] Kenneth Li Minn Ang, Jasmine Kah Phooi Seng, and Ericmoore Ngharamike. Towards crowdsourcing Internet of Things (Crowd-IoT): Architectures, security and applications. Future Internet, 14(2):49, January 31, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
2/49.

Ahmed:2023:BSS

[ASWAE23] Khaled A. M. Ahmed, Sabry F. Saraya, John F. Wanis, and Amr M. T. Ali-Eldin. A blockchain self-sovereign identity for open banking secured by the Customer’s banking cards. Future Internet, 15(6):208–??, June 2023. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
6/208.

Antonio:2014:GDD

[AT14] Amy Antonio and David Tuffley. The gender digital divide in developing countries. Future Internet, 6(4):673–687, October 31, 2014. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/6/4/673.

Ahmed:2024:TOB

[ATB^{+}24] Imen Ahmed, Mariem Turki, Mouna Baklouti, Bouthaina Dammak, and Amnah Alshahrani. Towards an optimized blockchainbased secure medical prescription-management system. Future Internet, 16(7):??, July 9, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
7/243.

Aiche:2024:MTI

[ATE24] Aicha Aiche, Pierre-Martin Tardif, and Mohammed Erritali. Modeling trust in IoT systems for drinkingwater management. Future Internet, 16(8):??, July 30, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/8/273.

Agusti-Torra:2024:MBC

[ATFMOU^{+}24] Anna Agustí-Torra, Marc Ferré-Mancebo, Gabriel David Orozco-Urrutia, David Rincón-Rivera, and David Remondo. A microservicesbased control plane for time-sensitive networking. Future Internet, 16(4):120, April 01, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
4/120.

Ali:2018:BRN

[ATH18] Mohammed N. A. Ali, Guanzheng Tan, and Aamir Hussain. Bidirectional recurrent neural network approach for Arabic named entity recognition. Future Internet, 10(12):123, December 13, 2018. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/10/12/123.

Alodhyani:2020:PMA

[ATR20] Fahad Alodhyani, George Theodorakopoulos, and Philipp Reinecke. Password managers — it’s all about trust and transparency. Future Internet, 12(11):189, October 30, 2020. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/12/11/189.

Agafonova:2020:PTR

[ATY20] Yana Agafonova, Alexey Tikhonov, and Ivan P. Yamshchikov. Paranoid transformer: Reading narrative of madness as computational approach to creativity. Future Internet, 12(11):182, October 27, 2020. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/12/11/182.

Alotaibi:2023:TSB

[AV23] Fahad M. Alotaibi and Vassilios G. Vassilakis. Toward an SDN-based Web application firewall: Defending against SQL injection attacks. Future Internet, 15(5):170, April 29, 2023. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/15/5/170.

Alberro:2022:ERP

[AVA^{+}22] Leonardo Alberro, Felipe Velázquez, Sara Azpiroz, Eduardo Grampin, and Matías Richart. Experimenting with routing protocols in the data center: an ns-3 simulation approach. Future Internet, 14(10):292, October 14, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/10/292.

Artem:2022:DFI

[AVE^{+}22] Volkov Artem, Kovalenko Vadim, Ibrahim A. Elgendy, Ammar Muthanna, and Andrey Koucheryavy. DD-FoG: Intelligent distributed dynamic FoG computing framework. Future Internet, 14(1):13, December 27, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
1/13.

Anastasakis:2023:FFR

[AVV^{+}23] Zacharias Anastasakis, Terpsichori-Helen Velivassaki, Artemis Voulkidis, Stavroula Bourou, Konstantinos Psychogyios, Dimitrios Skias, , and Theodore Zahariadis. FREDY: Federated resilience enhanced with differential privacy. Future Internet, 15(9):296, September 01, 2023. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/15/9/296.

Alaei:2023:TOD

[AWBS23] Alireza Alaei, Ying Wang, Vinh Bui, and Bela Stantic. Target-oriented data annotation for emotion and sentiment analysis in tourism related social media data. Future Internet, 15(4):150, April 19, 2023. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
4/150.

Ahmad:2023:SCC

[AWSA23] Nafi Ahmad, Abdul Wahab, John Schormans, and Ali Adib Arnab. Significance of cross-correlated QoS configurations for validating the subjective and objective QoE of cloud gaming applications. Future Internet, 15(2):64, February 02, 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
2/64.

Arif:2019:CBF

[AWUF19] Sheeraz Arif, Jing Wang, Tehseen Ul Hassan, and Zesong Fei. 3D-CNN-based fused feature maps with LSTM applied to action recognition. Future Internet, 11(2):42, February 13, 2019. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/11/2/42.

Altulyan:2021:CIA

[AYH^{+}21] May Altulyan, Lina Yao, Chaoran Huang, Xianzhi Wang, and Salil S. Kanhere. Context-induced activity monitoring for ondemand things-of-interest recommendation in an ambient intelligent environment. Future Internet, 13(12):305, November 28, 2021. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/13/12/305.

Alsadeh:2022:DFI

[AYH22] Ahmad Alsadeh, Nasri Yatim, and Yousef Hassouneh. A dynamic federated identity management using OpenID connect. Future Internet, 14(11):339, November 21, 2022. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/14/
                                        11/339.

Ali:2023:RBC

[AYH23] Shams Mhmood Abd Ali, Mohd Najwadi Yusoff, and Hasan Falah Hasan. Redactable blockchain: Comprehensive review, mechanisms, challenges, open issues and future research directions. Future Internet, 15(1):35, January 12, 2023. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/15/1/35.

Alshareef:2021:SGB

[AYRA21] Norah Alshareef, Xiaohong Yuan, Kaushik Roy, and Mustafa Atay. A study of gender bias in face presentation attack and its mitigation. Future Internet, 13(9):234, September 14, 2021. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
9/234.

Al-Zahrani:2019:GTI

[AZ19] Ali Y. Al-Zahrani. A game theoretic interference management scheme in full duplex cellular systems under infeasible QoS requirements. Future Internet, 11(7):156, July 16, 2019. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/11/7/156.

Afrifa:2022:MML

[AZAV22] Stephen Afrifa, Tao Zhang, Peter Appiahene, and Vijayakumar Varadarajan. Mathematical and machine learning models for groundwater level changes: a systematic review and bibliographic analysis. Future Internet, 14(9):259, August 30, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
9/259.

Adnan:2022:QKD

[AZH22] Mohd Hirzi Adnan, Zuriati Ahmad Zukarnain, and Nur Ziadah Harun. Quantum key distribution for 5G networks: A review, state of art and future directions. Future Internet, 14(3):73, February 25, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
3/73.

Al-Zubaidie:2023:ADL

[AZS23] Mishall Al-Zubaidie and Ghanima Sabr Shyaa. Applying detection leakage on hybrid cryptography to secure transaction information in e-commerce apps. Future Internet, 15(8):262, August 01, 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
8/262.

Antoniadis:2016:MAP

[AZV16] Konstantinos Antoniadis, Kostas Zafiropoulos, and Vasiliki Vrana. A method for assessing the performance of e-government Twitter accounts. Future Internet, 8(2):12, April 18, 2016. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/8/2/12.

Badidi:2023:EAE

[Bad23] Elarbi Badidi. Edge AI for early detection of chronic diseases and the spread of infectious diseases: Opportunities, challenges, and future directions. Future Internet, 15(11):370, November 18, 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
11/370.

Bandi:2023:PGA

[BAK23] Ajay Bandi, Pydi Venkata Satya Ramesh Adapa, and Yudu Eswar Vinay Pratap Kumar Kuchi. The power of generative AI: A review of requirements, models, input–output formats, evaluation metrics, and challenges. Future Internet, 15(8):260, July 31, 2023. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/15/8/260.

Begum:2021:UAB

[BAR^{+}21] Nasima Begum, Md Azim Hossain Akash, Sayma Rahman, Jungpil Shin, Md Rashedul Islam, and Md Ezharul Islam. User authentication based on handwriting analysis of pen-tablet sensor data using optimal feature selection model. Future Internet, 13(9):231, September 06, 2021. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
9/231.

Baabdullah:2024:EFL

[BARL24] Tahani Baabdullah, Amani Alzahrani, Danda B. Rawat, and Chunmei Liu. Efficiency of federated learning and blockchain in preserving privacy and enhancing the performance of credit card fraud detection (CCFD) systems. Future Internet, 16(6): 196, June 2, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
6/196.

Bunu:2024:IRP

[BAS24] Sanusi Mohammad Bunu, Omar Younis Alani, and Mohammad Saraee. An improved routing protocol for optimum quality of service in device-to-device and energy efficiency in 5G/B5G. Future Internet, 16(9), September 23, 2024. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
9/347.

Bonguet:2017:SDS

[BB17] Adrien Bonguet and Martine Bellaiche. A survey of denial-of-service and distributed denial of service attacks and defenses in cloud computing. Future Internet, 9(3):43, August 05, 2017. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/9/3/43.

Baldi:2020:IGR

[BB20] Alberto Baldi and Franco Bagnoli. Intransitiveness: From games to random walks. Future Internet, 12(9):151, September 03, 2020. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/12/9/151.

Boumechaal:2023:CQQ

[BB23] Hasna Boumechaal and Zizette Boufaida. Complex queries for querying linked data. Future Internet, 15 (3):106, March 09, 2023. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
3/106.

Blekanov:2021:DHC

[BBA21] Ivan Blekanov, Svetlana S.

Bodrunova, and Askar Akhmetov. Detection of hidden communities in Twitter discussions of varying volumes. Future Internet, 13(11):295, November 20, 2021. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
11/295.

Elallid:2024:EAD

[BBBHA24] Badr Ben Elallid, Nabil Benamar, Miloud Bagaa, and Yassine Hadjadj-Aoul. Enhancing autonomous driving navigation using soft actor-critic. Future Internet, 16(7):??, July 4, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/7/238.

Bellini:2019:EID

[BBGP19] Alessandro Bellini, Emanuele Bellini, Monica Gherardelli, and Franco Pirri. Enhancing IoT data dependability through a blockchain mirror model. Future Internet, 11(5): 117, May 21, 2019. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
5/117.

Bray:2024:LCA

[BBH^{+}24] Nick Bray, Matthew Boeding, Michael Hempel, Hamid Sharif, Tapio Heikkilä, Markku Suomalainen, and Tuomas Seppälä. A latency composition analysis for telerobotic performance insights across various network scenarios. Future Internet, 16(12):457, December 4, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/12/457.

Bylieva:2020:OGS

[BBK^{+}20] Daria Bylieva, Zafer Bekirogullari, Dmitry Kuznetsov, Nadezhda Almazova, Victoria Lobatyuk, and Anna Rubtsova. Online group student peercommunication as an element of open education. Future Internet, 12(9):143, August 26, 2020. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/12/
9/143.

Begotti:2020:CSA

[BBM20] Tatiana Begotti, Martina Bollo, and Daniela Acquadro Maran. Coping strategies and anxiety and depressive symptoms in young adult victims of cyberstalking: A questionnaire survey in an Italian sample. Future Internet, 12(8):136, August 12, 2020. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/12/8/136.

Bagnoli:2019:PIS

[BBMR19] Franco Bagnoli, Emanuele Bellini, Emanuele Massaro, and Raúl Rechtman. Percolation and Internet science. Future Internet, 11(2):35, February 02, 2019. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/11/2/35.

Bagheri:2024:SGS

[BBSR24] Nasour Bagheri, Ygal Bendavid, Masoumeh Safkhani, and Samad Rostampour. Smart grid security: a PUF-based authentication and key agreement protocol. Future Internet, 16 (1):9, December 28, 2024. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/16/
                                        1/9.

Bertolazzi:2019:MED

[BC19] Marco Bertolazzi and Carlo Caini. Mars to Earth data downloading: a directory synchronization approach. Future Internet, 11(8):173, August 08, 2019. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/11/8/173.

Bushur:2023:NNE

[BC23] Jacob Bushur and Chao Chen. Neural network exploration for keyword spotting on edge devices. Future Internet, 15(6): 219–??, June 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
6/219.

Bonanni:2021:DCA

[BCFP21] Michele Bonanni, Francesco Chiti, Romano Fantacci, and Laura Pierucci. Dynamic control architecture based on software defined networking for the Internet of Things. Future Internet, 13(5):113, April 28, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/5/113.

Benedetto:2010:DQE

[BCG10] Francesco Benedetto, Alberto Curcio, and Gaetano Giunta. Dynamic QoS evaluation of multimedia contents in wireless networks by “Double-Boomerang” watermarking. Future Internet, 2 (1):60–73, March 08, 2010. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/2/
1/60.

Berardi:2023:WOT

[BCG^{+}23] Davide Berardi, Franco Callegati, Andrea Giovine, Andrea Melis, Marco Prandini, and Lorenzo Rinieri. When operation technology meets information technology: Challenges and opportunities. Future Internet, 15(3):95, February 27, 2023. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/15/
                                        3/95.

Brown:2011:MPB

[BCL11] Katie Brown, Scott W. Campbell, and Rich Ling. Mobile phones bridging the digital divide for teens in the US? Future Internet, 3(2):144–158, May 13, 2011. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/3/2/144.

Bazzi:2019:SPV

[BCM^{+}19a] Alessandro Bazzi, Giammarco Cecchini, Michele Menarini, Barbara M. Masini, and Alberto Zanella. Survey and perspectives of vehicular wi-fi versus sidelink cellular-V2X in the 5G era. Future Internet, 11(6):122, May 29, 2019. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/11/6/122.

Busanelli:2019:SOO

[BCM^{+}19b] Stefano Busanelli, Simone Cirani, Lorenzo Melegari, Marco Picone, Mirco Rosa, and Luca Veltri. A sidecar object for the optimized communication between edge and cloud in Internet of Things applications. Future Internet, 11(7):145, July 05, 2019. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
7/145.

Battistoni:2023:CPS

[BCM^{+}23] Pietro Battistoni, Andrea Antonio Cantone, Gerardo Martino, Valerio Passamano, Marco Romano, Monica Sebillo, and Giuliana Vitiello. A cyberphysical system for wildfire detection and firefighting. Future Internet, 15(7): 237–??, July 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
7/237.

Bressan:2021:DLB

[BCMPW21] Giulia Bressan, Giulia Cisotto, Gernot R. Müller-Putz, and Selina Christin Wriessnegger. Deep learning-based classification of fine hand movements from low frequency EEG. Future Internet, 13 (5):103, April 21, 2021. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
5/103.

Barletta:2020:IDV

[BCNS20] Vita Santa Barletta, Danilo Caivano, Antonella Nannavecchia, and Michele Scalera. Intrusion detection for in-vehicle communication networks: An unsupervised Kohonen SOM approach. Future Internet, 12(7):119, July 14, 2020. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/12/
                                        7/119.

Bolanowski:2024:NMD

[BCS24] Marek Bolanowski, Michał ’Cmil, and Adrian Starzec. New model for defining and implementing performance tests. Future Internet, 16(10), October 10, 2024. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/16/10/366.

Bagunaid:2024:VDP

[BCSB24] Wala Bagunaid, Naveen Chilamkurti, Ahmad Salehi Shahraki, and Saeed Bamashmos. Visual data and pattern analysis for smart education: a robust DRLbased early warning system for student performance prediction. Future Internet, 16(6):206, June 11, 2024. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/16/6/206.

Basile:2021:HDE

[BCT21] Valerio Basile, Francesco Cauteruccio, and Giorgio Terracina. How dramatic events can affect emotionality in social posting: The impact of COVID-19 on Reddit. Future Internet, 13(2):29, January 27, 2021. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/13/2/29.

Baras:2023:IEM

[BCZ^{+}23] Nikolaos Baras, Antonios Chatzisavvas, Dimitris Ziouzios, Ioannis Vanidis, and Minas Dasygenis. Improving the efficiency of modern warehouses using smart battery placement. Future Internet, 15(11): 353, October 26, 2023. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
11/353.

Bellavista:2024:IID

[BD24] Paolo Bellavista and Giuseppe Di Modica. IoTwins: Implementing distributed and hybrid digital twins in industrial manufacturing and facility management settings. Future Internet, 16(2):65, February 17, 2024. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/16/2/65.

Bagnoli:2023:SMK

[BdBC23] Franco Bagnoli and Guido de Bonfioli Cavalcabo’. A simple model of knowledge scaffolding applied to Wikipedia growth. Future Internet, 15(2):67, February 06, 2023. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/15/2/67.

Bagnoli:2021:CFB

[BdBCCG21] Franco Bagnoli, Guido de Bonfioli Cavalcabo’, Banedetto Casu, and Andrea Guazzini. Community formation as a byproduct of a recommendation system: a simulation model for bubble formation in social media. Future Internet, 13(11): 296, November 22, 2021. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/13/
                                        11/296.

Barradas:2010:ITA

[BdCRM10] Alvaro L. Barradas and Maria do Carmo R. Medeiros. An intrinsic TE approach for end-to-end QoS provisioning in OBS networks using static load-balanced routing strategies. Future Internet, 2(4):559– 586, October 22, 2010. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/2/
                                        4/559.

Buccafurri:2022:BBF

[BDL22] Francesco Buccafurri, Vincenzo De Angelis, and Sara Lazzaro. A blockchainbased framework to enhance anonymous services with accountability guarantees. Future Internet, 14 (8):243, August 21, 2022.

CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
8/243.

Beall:2010:MND

[Bea10] Jeffrey Beall. Metadata for name disambiguation and collocation. Future Internet, 2(1):1–15, January 05, 2010. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/2/1/1.

Bendjima:2018:ICW

[BF18] Mostefa Bendjima and Mohammed Feham. Intelligent communication in wireless sensor networks. Future Internet, 10(9):91, September 15, 2018. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/10/
9/91.

Bouzidi:2012:SWA

[BFFZ^{+}12] Khalil Riad Bouzidi, Bruno Fies, Catherine Faron-Zucker, Alain Zarli, and Nhan Le Thanh. Semantic Web approach to ease regulation compliance checking in construction industry. Future Internet, 4(3):830–851, September 11, 2012. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/4/3/830.

Bocci:2023:SPC

[BFG^{+}23] Alessandro Bocci, Stefano Forti, Roberto Guanciale, Gian-Luigi Ferrari, and Antonio Brogi. Secure partitioning of cloud applications, with cost lookahead. Future Internet, 15 (7):224–??, July 2023. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/15/
                                        7/224.

Bigini:2020:RBI

[BFL20] Gioele Bigini, Valerio Freschi, and Emanuele Lattanzi. A review on blockchain for the Internet of Medical Things: Definitions, challenges, applications, and vision. Future Internet, 12(12):208, November 25, 2020. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/12/
                                        12/208.

Branchi:2014:AMS

[BFVM14] Pablo E. Branchi, Carlos Fernández-Valdivielso, and Ignacio R. Matias. Analysis matrix for smart cities. Future Internet, 6(1):61–75, January 22, 2014. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/6/1/61.

Bai:2022:PPO

[BFY22] Tianyu Bai, Song Fu, and Qing Yang. Privacypreserving object detection with secure convolutional neural networks for vehicular edge computing. Future Internet, 14(11): 316, October 31, 2022. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
11/316.

Beemer:2009:MLR

[BG09] Brandon Beemer and Dawn Gregg. Mashups: a literature review and classification framework. Future Internet, 1(1):59–87, December 22, 2009. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/1/
1/59.

Barbierato:2024:DUI

[BG24] Enrico Barbierato and Alice Gatti. Decoding urban intelligence: Clustering and feature importance in smart cities. Future Internet, 16(10), October 5, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/10/362.

Batcheller:2009:MAG

[BGD09] James K. Batcheller, Bruce M. Gittings, and Robert I. Dunfey. A method for automating geospatial dataset metadata. Future Internet, 1 (1):28–46, November 10, 2009. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/1/1/28.

Bottrighi:2022:IIP

[BGL^{+}22] Alessio Bottrighi, Marco Guazzone, Giorgio Leonardi, Stefania Montani, Manuel Striani, and Paolo Terenziani. Integrating ISA and part-of domain knowledge into process model discovery. Future Internet, 14(12):357, November 28, 2022. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/14/12/357.

Boulos:2016:IWH

[BGW16] Maged N. Kamel Boulos, Dean M. Giustini, and Steve Wheeler. Instagram and WhatsApp in health and healthcare: an overview. Future Internet, 8(3):37, July 26, 2016. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/8/
                                        3/37.

Bayu:2023:RLA

[BHC23] Teguh Indra Bayu, Yung-Fa Huang, and Jeang-Kuo Chen. Reinforcement learning approach for adaptive C-V2X resource management. Future Internet, 15(10):339, October 15, 2023. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/15/10/339.

Bunting:2012:PAE

[BHH12] Ben S. Bunting, Jr., Jacob Hughes, and Tim Hetland. The player as author: Exploring the effects of mobile gaming and the location-aware interface on storytelling. Future Internet, 4(1):142–160, February 17, 2012. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/4/1/142.

Boddy:2019:IHD

[BHM^{+}19] Aaron Boddy, William Hurst, Michael Mackay, Abdennour El Rhalibi, Thar Baker, and Casimiro A. Curbelo Montañez. An investigation into healthcare-data patterns. Future Internet, 11(2):30, January 30, 2019. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
2/30.

Boeding:2024:NAT

[BHS24] Matthew Boeding, Michael Hempel, and Hamid Sharif. Novel approach towards a fully deep learning-based IoT receiver architecture: From estimation to decoding. Future Internet, 16(5): 155, April 30, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.

                                        mdpi.com/1999-5903/16/
                                        5/155.

Bani-Hani:2024:VDC

[BHSAY24] Raed M. Bani-Hani, Ahmed S. Shatnawi, and Lana Al-Yahya. Vulnerability detection and classification of Ethereum smart contracts using deep learning. Future Internet, 16(9), September 4, 2024. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/16/
                                        9/321.

Bi:2014:RDE

[Bi14] Huibo Bi. Routing diverse evacuees with the cognitive packet network algorithm. Future Internet, 6(2):203–222, April 17, 2014. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/6/2/203.

Brackin:2022:GID

[BJL^{+}22] Roger C. Brackin, Michael J. Jackson, Andrew Leyshon, Jeremy G. Morley, and Sarah Jewitt. Generating indicators of disruptive innovation using big data. Future Internet, 14(11): 327, November 11, 2022. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/14/
                                        11/327.

Benmoussa:2023:NBB

[BKCL23] Ahmed Benmoussa, Chaker Abdelaziz Kerrache, Carlos T. Calafate, and Nasreddine Lagraa. NDN-BDA: a blockchain-based decentralized data authentication mechanism for vehicular named data networking. Future Internet, 15(5):167, April 29, 2023. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
5/167.

Bocharova:2019:LDI

[BKL^{+}19] Irina Bocharova, Boris Kudryashov, Nikita Lyamin, Erik Frick, Maben Rabi, and Alexey Vinel. Low delay inter-packet coding in vehicular networks. Future Internet, 11(10):212, October 11, 2019. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
10/212.

Balmuri:2022:LST

[BKL^{+}22] Kavitha Rani Balmuri, Srinivas Konda, Wen-Cheng Lai, Parameshachari Bidare Divakarachari, Kavitha Malali Vishveshwarappa Gowda, and Hemalatha Kivudujogappa Lingappa. A long short-term memory network-based radio resource management for 5G network. Future Internet, 14(6):184, June 14, 2022.

CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/14/
                                        6/184.

Balaskas:2024:PYD

[BKR24] Stefanos Balaskas, Georgia Kotsari, and Maria Rigou. Perspectives of young digital natives on digital marketing: Exploring annoyance and effectiveness with eye-tracking analysis. Future Internet, 16(4):125, April 08, 2024. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/16/
                                        4/125.

Behan:2019:SNF

[BKSS19] Miroslav Behan, Ondrej Krejcar, Thabit Sabbah, and Ali Selamat. Sensorial network framework embedded in ubiquitous mobile devices. Future Internet, 11(10):215, October 14, 2019. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/11/10/215.

Baumgartner:2024:MEB

[BKW^{+}24] Martin Baumgartner, Karl Kreiner, Fabian Wiesmüller, Dieter Hayn, Christian Puelacher, and Günter Schreier. Masketeer: an ensemble-based pseudonymization tool with entity recognition for German unstructured medical free text. Future Internet, 16(8):??, August 6, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
8/281.

Bilami:2022:LBB

[BL22] Karam Eddine Bilami and Pascal Lorenz. Lightweight blockchain-based scheme to secure wireless M2M area networks. Future Internet, 14(5):158, May 23, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/5/158.

Buchanan:2015:FIW

[BLU^{+}15] William J. Buchanan, David Lanc, Elochukwu Ukwandu, Lu Fan, Gordon Russell, and Owen Lo. The future Internet: a world of secret shares. Future Internet, 7(4):445–464, November 24, 2015. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/7/4/445.

Binmad:2017:EFR

[BLW^{+}17] Ruchdee Binmad, Mingchu Li, Zhen Wang, Nakema Deonauth, and Chettupally Anil Carie. An extended framework for recovering from trust breakdowns in online community settings. Future Internet, 9(3):36, July 17, 2017. CODEN ???? ISSN 1999-

5903. URL https://www.

                                        mdpi.com/1999-5903/9/
                                        3/36.

Braun:2014:SDN

[BM14] Wolfgang Braun and Michael Menth. Software-defined networking using Open-Flow: Protocols, applications and architectural design choices. Future Internet, 6(2):302–336, May 12, 2014. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/6/2/302.

Brown:2018:DFC

[BM18] Suzana Brown and Alan Mickelson. A decision framework for choosing telecommunication technologies in limitedresource settings. Future Internet, 10(1):8, January 12, 2018. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/10/1/8.

Begotti:2019:CCB

[BM19] Tatiana Begotti and Daniela Acquadro Maran. Characteristics of cyberstalking behavior, consequences, and coping strategies: a crosssectional study in a sample of Italian university students. Future Internet, 11(5):120, May 22, 2019. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/11/
                                        5/120.

Bader:2020:SDD

[BM20] Sebastian R. Bader and Maria Maleshkova. SO-LIOT — decentralized data control and interactions for IoT. Future Internet, 12(6):105, June 16, 2020. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/12/6/105.

Bucur:2023:EMJ

[BM23] Vlad Bucur and Liviu-Cristian Miclea. Entering the metaverse from the JVM: The state of the art, challenges, and research areas of JVM-based Web 3.0 tools and libraries. Future Internet, 15(9):305, September 07, 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
9/305.

Bagui:2023:RIN

[BMB^{+}23a] Sikha Bagui, Dustin Mink, Subhash Bagui, Sakthivel Subramaniam, and Daniel Wallace. Resampling imbalanced network intrusion datasets to identify rare attacks. Future Internet, 15(4):130, March 29, 2023. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/15/4/130.

Bagui:2023:UGE

[BMB^{+}23b] Sikha S. Bagui, Dustin Mink, Subhash C. Bagui, Michael Plain, Jadarius Hill, and Marshall Elam. Using a graph engine to visualize the reconnaissance tactic of the MITRE ATT&CK framework from UWF-ZeekData22. Future Internet, 15(7):236– ??, July 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
7/236.

Bajaj:2023:CDD

[BMB23c] Ruhi Kiran Bajaj, Rebecca Mary Meiring, and Fernando Beltran. Codesign, development, and evaluation of a health monitoring tool using smartwatch data: A proofof-concept study. Future Internet, 15(3):111, March 17, 2023. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/15/3/111.

Bouloukakis:2021:PAI

[BMGI21] Georgios Bouloukakis, Ioannis Moscholios, Nikolaos Georgantas, and Valérie Issarny. Performance analysis of Internet of Things interactions via simulation-based queueing models. Future Internet, 13(4):87, March 29, 2021. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
4/87.

Battisti:2022:NSV

[BMJ^{+}22] Anselmo Luiz Éden Battisti, Evandro Luiz Cardoso Macedo, Marina Ivanov Pereira Josué, Hugo Barbalho, Flávia C. Delicato, Débora Christina Muchaluat-Saade, Paulo F. Pires, Douglas Paulo de Mattos, and Ana Cristina Bernardo de Oliveira. A novel strategy for VNF placement in edge computing environments. Future Internet, 14(12):361, November 30, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
12/361.

Bader:2019:SRA

[BML19] Sebastian R. Bader, Maria Maleshkova, and Steffen Lohmann. Structuring reference architectures for the industrial Internet of Things. Future Internet, 11(7):151, July 08, 2019. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
7/151.

Butnaru:2021:TLU

[BMP21] Andrei Butnaru, Alexios Mylonas, and Nikolaos Pitropakis. Towards lightweight URLbased phishing detection. Future Internet, 13(6):154, June 13, 2021. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/13/
                                        6/154.

Badidi:2020:FCS

[BMS20] Elarbi Badidi, Zineb Mahrez, and Essaid Sabir. Fog computing for smart cities’ big data management and analytics: a review. Future Internet, 12(11):190, October 31, 2020. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/12/11/190.

BinSaif:2024:ESD

[BMS24] Muhammad Bin Saif, Sara Migliorini, and Fausto Spoto. Efficient and secure distributed data storage and retrieval using interplanetary file system and blockchain. Future Internet, 16(3):98, March 15, 2024. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/16/3/98.

Brainin:2014:ITO

[BN14] Esther Brainin and Efrat Neter. Inside technology: Opening the black box of health — Website configuration and content management. Future Internet, 6(4):773–799, December 10, 2014. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/6/4/773.

Bellaj:2024:DMM

[BNJ24] Mohammed Bellaj, Najib Naja, and Abdellah Jamali. Distributed mobility management support for low-latency data delivery in named data networking for UAVs. Future Internet, 16(2):57, February 10, 2024. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/16/2/57.

Sada:2023:CAE

[BNK^{+}23] Abdelkarim Ben Sada, Abdenacer Naouri, Amar Khelloufi, Sahraoui Dhelim, and Huansheng Ning. A context-aware edge computing framework for smart Internet of Things. Future Internet, 15(5):154, April 22, 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
5/154.

Bodrunova:2020:TDB

[BOB^{+}20] Svetlana S. Bodrunova, Andrey V. Orekhov, Ivan S. Blekanov, Nikolay S. Lyudkevich, and Nikita A. Tarasov. Topic detection based on sentence embeddings and agglomerative clustering with Markov moment. Future Internet, 12(9):144, August 26, 2020. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/12/9/144.

Bodrunova:2022:ESI

[Bod22] Svetlana S. Bodrunova. Editorial for the special issue “Selected Papers from the 9th Annual Conference ‘Comparative Media Studies in Today’s World’ (CMSTW’2021)”. Future Internet, 14(11):334, November 16, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
11/334.

Bidry:2023:ELB

[BOH23] Mahmoud Bidry, Abdellah Ouaguid, and Mohamed Hanine. Enhancing elearning with blockchain: Characteristics, projects, and emerging trends. Future Internet, 15(9):293, August 28, 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
9/293.

Beschastnyi:2021:QDM

[BOK^{+}21] Vitalii Beschastnyi, Daria Ostrikova, Roman Konyukhov, Elizaveta Golos, Alexander Chursin, Dmitri Moltchanov, and Yuliya Gaidamaka. Quantifying the density of mmWave NR deployments for provisioning multilayer VR services. Future Internet, 13(7):185, July 20, 2021. CODEN ????

ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/7/185.

Burger:2019:CSS

[BOKC19] Annetta Burger, Talha Oz, William G. Kennedy, and Andrew T. Crooks. Computational social science of disasters: Opportunities and challenges. Future Internet, 11(5):103, April 26, 2019. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/11/5/103.

Baskar:2024:MCB

[BP24] Premalatha Baskar and Prakasam Periasamy. Minimumcost-based neighbour node discovery scheme for fault tolerance under IoT-fog networks. Future Internet, 16(4):123, April 03, 2024. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
4/123.

Bencsath:2012:CSD

[BPBF12] Boldizsár Bencsáth, Gábor Pék, Levente Buttyán, and Márk Félegyházi. The cousins of Stuxnet: Duqu, Flame, and Gauss. Future Internet, 4(4):971– 1003, November 06, 2012. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/4/
4/971.

Butun:2019:SRA

[BPG19] Ismail Butun, Nuno Pereira, and Mikael Gidlund. Security risk analysis of Lo-RaWAN and future directions. Future Internet, 11 (1):3, December 21, 2019. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/11/
                                        1/3.

Balderas:2021:ETB

[BPLBM21] David Balderas, Pedro Ponce, Diego Lopez-Bernal, and Arturo Molina. Education 4.0: Teaching the basis of motor imagery classification algorithms for brain–computer interfaces. Future Internet, 13 (8):202, August 03, 2021. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/13/
                                        8/202.

Bakaev:2021:WMU

[BR21] Maxim Bakaev and Olga Razumnikova. What makes a UI simple? Difficulty and complexity in tasks engaging visual– spatial working memory. Future Internet, 13(1):21, January 19, 2021. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/13/
                                        1/21.

Bhurtel:2023:ULO

[BR23] Manish Bhurtel and Danda B.

Rawat. Unveiling the landscape of operating system vulnerabilities. Future Internet, 15(7):248– ??, July 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
7/248.

Batita:2014:TCF

[BRBC14] Wided Batita, Stéphane Roche, Yvan Bédard, and Claude Caron. Towards a conceptual framework for WikiGIS. Future Internet, 6(4):640–672, October 29, 2014. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/6/4/640.

Bouzidi:2024:EAI

[BRC24] Abdelaziz Bouzidi, Lala Rajaoarisoa, and Luka Claeys. Explainable artificial intelligence approach for improving head-mounted fault display systems. Future Internet, 16(8):??, August 6, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/8/282.

Badshah:2023:TEI

[BRF^{+}23] Afzal Badshah, Ghani Ur Rehman, Haleem Farman, Anwar Ghani, Shahid Sultan, Muhammad Zubair, and Moustafa M. Nasralla. Transforming educational institutions: Harnessing the power of Internet of Things, cloud, and fog computing. Future Internet, 15(11):367, November 13, 2023. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/15/
                                        11/367.

Bashir:2023:BBL

[BRM23] Syed Raza Bashir, Shaina Raza, and Vojislav B. Misic. BERT4Loc: BERT for location-POI recommender system. Future Internet, 15(6):213–??, June 2023. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/15/6/213.

Busanelli:2013:CNI

[BRP^{+}13] Stefano Busanelli, Filippo Rebecchi, Marco Picone, Nicola Iotti, and Gianluigi Ferrari. Cross-network information dissemination in vehicular ad hoc networks (VANETs): Experimental results from a Smartphone-based testbed. Future Internet, 5(3):398– 428, August 05, 2013. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/5/
                                        3/398.

Bauman:2017:VIE

[BS17] Brian Bauman and Patrick Seeling. Visual interface evaluation for wearables datasets: Predicting the subjective augmented vision image QoE and QoS. Future Internet, 9(3):40, July 28, 2017. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/9/3/40.

Bendechache:2021:SER

[BSE^{+}21] Malika Bendechache, Sergej Svorobej, Patricia Takako Endo, Adrian Mihai, and Theo Lynn. Simulating and evaluating a realworld ElasticSearch system using the RECAP DES simulator. Future Internet, 13(4):83, March 24, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/4/83.

Bendechache:2020:SRM

[BSEL20] Malika Bendechache, Sergej Svorobej, Patricia Takako Endo, and Theo Lynn. Simulating resource management across the cloudto-thing continuum: a survey and future directions. Future Internet, 12(6):95, May 29, 2020. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/12/6/95.

Baig:2022:SSC

[BSM22] Zubair Baig, Naeem Syed, and Nazeeruddin Mohammad. Securing the smart city airspace: Drone cyber attack detection through machine learning. Future Internet, 14(7):205, June 30, 2022. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/14/7/205.

Blekanov:2022:TBA

[BTB22] Ivan S. Blekanov, Nikita Tarasov, and Svetlana S. Bodrunova. Transformerbased abstractive summarization for Reddit and Twitter: Single posts vs. comment pools in three languages. Future Internet, 14(3):69, February 23, 2022. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/14/3/69.

Baranov:2021:IEN

[BTBK21] Dmitry Baranov, Alexandr Terekhin, Dmitry Bragin, and Anton Konev. Implementation and evaluation of nodal distribution and movement in a 5G mobile network. Future Internet, 13(12):321, December 20, 2021. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/13/
                                        12/321.

Barakat:2021:OAI

[BTS^{+}21] Basel Barakat, Ahmad Taha, Ryan Samson, Aiste Steponenaite, Shuja Ansari, Patrick M. Langdon, Ian J. Wassell, Qammer H. Abbasi, Muhammad Ali Imran, and Simeon Keates. 6G opportunities arising from Internet of Things use cases: a review paper. Future Internet, 13(6):159, June 18, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/6/159.

Bagha:2022:HSP

[BYG22] Hamid Bagha, Ali Yavari, and Dimitrios Georgakopoulos. Hybrid sensing platform for IoT-based precision agriculture. Future Internet, 14(8):233, July 28, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/8/233.

Boulos:2015:TIF

[BYS^{+}15] Maged N. Kamel Boulos, Abdulslam Yassine, Shervin Shirmohammadi, Chakkrit Snae Namahoot, and Michael Brückner. Towards an “Internet of Food”: Food ontologies for the Internet of Things. Future Internet, 7(4):372– 392, October 01, 2015. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/7/
4/372.

Barro:2019:SCL

[BZDP19] Pape Abdoulaye Barro, Marco Zennaro, Jules Degila, and Ermanno Pietrosemoli. A smart cities LoRaWAN network based on autonomous base stations (BS) for some countries with limited Internet access. Future Internet, 11 (4):93, April 08, 2019. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/11/
                                        4/93.

Borghesi:2023:MNF

[BZMB23] Michela Borghesi, Cristian Zambelli, Rino Micheloni, and Stefano Bonnini. Modeling 3D NAND flash with nonparametric inference on regression coefficients for reliable solid-state storage. Future Internet, 15(10):319, September 26, 2023. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/15/
                                        10/319.

Bouarourou:2023:EMB

[BZNB23] Soukaina Bouarourou, Abderrahim Zannou, El Habib Nfaoui, and Abdelhak Boulaalam. An efficient model-based clustering via joint multiple sink placement for WSNs. Future Internet, 15(2):75, February 15, 2023. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/15/2/75.

Barchi:2021:SNN

[BZP^{+}21] Francesco Barchi, Luca Zanatta, Emanuele Parisi, Alessio Burrello, Davide Brunelli, Andrea Bartolini, and Andrea Acquaviva. Spiking neural network-based near-sensor computing for damage detection in structural health monitoring. Future Internet, 13(8):219, August 23, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/8/219.

Caldarelli:2022:OBO

[Cal22] Giulio Caldarelli. Overview of blockchain oracle research. Future Internet, 14(6):175, June 08, 2022. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
6/175.

Coutinho:2023:EAD

[CAL23] Micael Coutinho, Jose A.

Afonso, and Sérgio F. Lopes. An efficient adaptive data-link-layer architecture for LoRa networks. Future Internet, 15(8):273, August 17, 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
8/273.

Calzada:2024:DWR

[Cal24] Igor Calzada. Decentralized Web3 reshaping Internet governance: Towards the emergence of new forms of nation-statehood? Future Internet, 16(10), October 4, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
10/361.

Crespo-Aguado:2024:FHD

[CALHG^{+}24] Maria Crespo-Aguado, Raul Lozano, Fernando Hernandez-Gobertti, Nuria Molner, and David Gomez-Barquero. Flexible hyperdistributed IoT–edge–cloud platform for real-time digital twin applications on 6g-intended testbeds for logistics and industry. Future Internet, 16(11):431, November 20, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
11/431.

Cap:2012:TCN

[Cap12] Clemens H. Cap. Towards content neutrality in Wiki systems. Future Internet, 4(4):1086–1104, December 19, 2012. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/4/4/1086.

Capkovic:2023:DDI

[Cap23] Frantisek Capkovic. Dealing with deadlocks in industrial multi agent systems. Future Internet, 15 (3):107, March 09, 2023. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
3/107.

Coelho:2020:MMB

[CAS^{+}20] Vitor Nazário Coelho, Rodolfo Pereira Araújo, Haroldo Gambini Santos, Wang Yong Qiang, and Igor Machado Coelho. A MILP model for a Byzantine fault tolerant blockchain consensus. Future Internet, 12(11):185, October 29, 2020. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/12/
11/185.

Chetty:2021:VNF

[CASN21] Swarna Bindu Chetty, Hamed Ahmadi, Sachin Sharma, and Avishek Nag. Virtual network function embedding under nodal outage using deep Qlearning. Future Internet, 13(3):82, March 23, 2021. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
3/82.

Carta:2016:ISI

[CB16] Salvatore Carta and Ludovico Boratto. Introduction to the special issue on human-computer interaction and the Social Web. Future Internet, 8(3):43, September 01, 2016. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/8/3/43.

Carnley:2022:PIT

[CB22] Renee Carnley and Sikha Bagui. A public infrastructure for a trusted wireless world. Future Internet, 14(7):200, June 30, 2022. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/14/
                                        7/200.

Carrillo:2024:TSE

[CB24] Alexis Carrillo and Moisés Betancort. Testing stimulus equivalence in transformerbased agents. Future Internet, 16(8):??, August 9, 2024. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/16/8/289.

Colombo:2017:CWU

[CBB17] Bruna Armonas Colombo, Pedro Buck, and Vinicius Miana Bezerra. Challenges when using jurimetrics in Brazil— a survey of courts. Future Internet, 9(4):68, October 25, 2017. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/9/4/68.

Cox:2022:PEN

[CBD^{+}22] Bert Cox, Chesney Buyle, Daan Delabie, Lieven De Strycker, and Liesbet Van der Perre. Positioning energy-neutral devices: Technological status and hybrid RF-acoustic experiments. Future Internet, 14(5):156, May 20, 2022. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
5/156.

Chemrov:2022:SPC

[CBKL22] Kirill Chemrov, Dmitry Bankov, Evgeny Khorov, and Andrey Lyakhov. Smart preliminary channel access to support real-time traffic in Wi-Fi networks. Future Internet, 14(10): 296, October 16, 2022. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
10/296.

Capra:2020:USE

[CBM^{+}20] Maurizio Capra, Beatrice Bussolino, Alberto Marchisio, Muhammad Shafique, Guido Masera, and Maurizio Martina. An updated survey of efficient hardware architectures for accelerating deep convolutional neural networks. Future Internet, 12(7):113, July 07, 2020. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/12/
7/113.

Carrascal:2024:FPR

[CBR^{+}24] David Carrascal, Paula Bartolomé, Elisa Rojas, Diego Lopez-Pajares, Nicolas Manso, and Javier Diaz-Fuentes. Fault prediction and reconfiguration optimization in smart grids: AI-driven approach. Future Internet, 16(11): 428, November 20, 2024. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
11/428.

Cappellari:2021:CBD

[CBRS21] Max Cappellari, John Belstner, Bryan Rodriguez, and Jeff Sedayao. A cloudbased data collaborative to combat the COVID-19 pandemic and to solve major technology challenges. Future Internet, 13(3):61, February 27, 2021. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
3/61.

Chen:2022:AVR

[CC22a] Ning Chen and Yu Chen. Anomalous vehicle recognition in smart urban traffic monitoring as an edge service. Future Internet, 14(2):54, February 10, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/2/54.

Csirmaz:2022:DSC

[CC22b] Elod P. Csirmaz and Laszlo Csirmaz. Data synchronization: a complete theoretical solution for filesystems. Future Internet, 14(11):344, November 21, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/11/344.

Csirmaz:2023:SMF

[CC23] Elod P. Csirmaz and Laszlo Csirmaz. Synchronizing many filesystems in near linear time. Future Internet, 15(6):198– ??, June 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
6/198.

Chou:2024:CCH

[CC24] Jerry Chou and Wu-Chun Chung. Cloud computing and high performance computing (HPC) advances for next generation Internet. Future Internet, 16(12):465, December 12, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/12/465.

Coelho:2020:CPI

[CCA^{+}20] Igor M. Coelho, Vitor N.

Coelho, Rodolfo P. Araujo, Wang Yong Qiang, and Brett D. Rhodes. Challenges of PBFT-inspired consensus for blockchain and enhancements over neo dBFT. Future Internet, 12(8):129, July 30, 2020. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/12/8/129.

Chen:2019:IAS

[CCC19] Yi-Cheng Chen, Yueh-Peng Chou, and Yung-Chen Chou. An image authentication scheme using Merkle tree mechanisms. Future Internet, 11(7):149, July 06, 2019. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/11/
                                        7/149.

Calabro:2024:MAS

[CCC^{+}24] Antonello Calabrò, Enrico Cambiaso, Manuel Cheminod, Ivan Cibrario Bertolotti, Luca Durante, Agostino Forestiero, Flavio Lombardi, Giuseppe Manco, Eda Marchetti, Albina Orlando, and Giuseppe Papuzzo. A methodological approach to securing cyber-physical systems for critical infrastructures. Future Internet, 16(11):418, November 12, 2024. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/16/11/418.

Campos-Castillo:2022:PCA

[CCCA22] Celeste Campos-Castillo, Noelle Chesley, and Onur Asan. Professionals as change agents or instruments of reproduction? Medical residents’ reasoning for not sharing the electronic health record screen with patients. Future Internet, 14(12):367, December 07, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
12/367.

Calle-Cancho:2024:DME

[CCCPV^{+}24] Jesus Calle-Cancho, Carlos Cañada, Rafael Pastor-Vargas, Mercedes E. Paoletti, and Juan M. Haut. Decentralized mechanism for edge node allocation in access network: an experimental evaluation. Future Internet, 16(9), September 20, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/9/342.

Celsi:2021:PTR

[CCD^{+}21] Lorenzo Ricciardi Celsi, Andrea Caliciotti, Matteo D’Onorio, Eugenio Scocchi, Nour Alhuda Sulieman, and Massimo Villari. On predicting ticket reopening for improving customer service in 5G fiber optic networks. Future Internet, 13(10):259, October 09, 2021. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
10/259.

Chang:2021:TLR

[CCKH21] Chin-Chen Chang, Jui-Feng Chang, Wei-Jiun Kao, and Ji-Hwei Horng. Two-layer reversible data hiding for VQ-compressed images based on declustering and indicatorfree search-order coding. Future Internet, 13(8):215, August 20, 2021. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/13/
                                        8/215.

Cominelli:2017:MPF

[CCM^{+}17] Lorenzo Cominelli, Nicola Carbonaro, Daniele Mazzei, Roberto Garofalo, Alessandro Tognetti, and Danilo De Rossi. A multimodal perception framework for users emotional state assessment in social robotics. Future Internet, 9(3):42, August 01, 2017. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/9/3/42.

Cho:2023:AIB

[CCTC23] Hsin-Hung Cho, Wei-Che Chien, Fan-Hsun Tseng, and Han-Chieh Chao. Artificial-intelligence-based charger deployment in wireless rechargeable sensor networks. Future Internet, 15(3):117, March 22, 2023. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/15/3/117.

Calcagno:2024:MOD

[CCTV24] Salvatore Calcagno, Andrea Calvagna, Emiliano Tramontana, and Gabriella Verga. Merging ontologies and data from electronic health records. Future Internet, 16(2):62, February 17, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
2/62.

Cui:2024:IDI

[CCYL24] Bo Cui, Yachao Chai, Zhen Yang, and Keqin Li. Intrusion detection in IoT using deep residual networks with attention mechanisms. Future Internet, 16(7):??, July 18, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/7/255.

Cai:2023:TSF

[CD23] Weihong Cai and Fengxi Duan. Task scheduling for federated learning in edge cloud computing environments by using adaptivegreedy dingo optimization algorithm and binary salp swarm algorithm. Future Internet, 15(11):357, October 30, 2023. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/15/11/357.

Cilfone:2019:WMN

[CDBF19] Antonio Cilfone, Luca Davoli, Laura Belli, and Gianluigi Ferrari. Wireless mesh networking: An IoToriented perspective survey on relevant technologies. Future Internet, 11 (4):99, April 17, 2019. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/11/
                                        4/99.

Contu:2021:ABA

[CDD^{+}21] Francesco Contu, Andrea Demontis, Stefano Dessì, Marco Muscas, and Daniele Riboni. AI-based analysis of policies and images for privacy-conscious content sharing. Future Internet, 13(6):139, May 21, 2021. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/13/6/139.

Chen:2024:ODP

[CDLY24] Shaolong Chen, Yunzi Dai, Liwei Liu, and Xinting Yu. Optimizing data parallelism for FM-based shortread alignment on the heterogeneous non-uniform memory access architectures. Future Internet, 16(6):217, June 19, 2024. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/16/
                                        6/217.

Cardone:2023:GBH

[CDM23] Barbara Cardone, Ferdinando Di Martino, and Vittorio Miraglia. A GISbased hot and cold spots detection method by extracting emotions from social streams. Future Internet, 15(1):23, December 30, 2023. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/15/1/23.

Cafaro:2024:SAF

[CEP24] Massimo Cafaro, Italo Epicoco, and Marco Pulimeno. State-of-the-art future Internet technology in Italy 2022–2023. Future Internet, 16(2):53, February 06, 2024. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/16/2/53.

Cividino:2019:EDU

[CEZC19] Sirio Cividino, Gianluca Egidi, Ilaria Zambon, and Andrea Colantoni. Evaluating the degree of uncertainty of research activities in Industry 4.0. Future Internet, 11(9):196, September 11, 2019. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
9/196.

Chua:2020:PAD

[CFG20] Sook-Ling Chua, Lee Kien Foo, and Hans W. Guesgen. Predicting activities of daily living with spatiotemporal information. Future Internet, 12(12):214, November 27, 2020. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/12/
                                        12/214.

Collodi:2020:HRS

[CFGD20] Stefania Collodi, Maria Fiorenza, Andrea Guazzini, and Mirko Duradoni. How reputation systems change the psychological antecedents of fairness in virtual environments. Future Internet, 12(8):132, August 09, 2020. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/12/
                                        8/132.

Chiti:2017:SAR

[CFP17] Francesco Chiti, Romano Fantacci, and Laura Pierucci. Social-aware relay selection for cooperative multicast deviceto-device communications. Future Internet, 9(4):92, December 04, 2017. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/9/
                                        4/92.

Chiti:2019:EEC

[CFP19] Francesco Chiti, Romano Fantacci, and Laura Pierucci. Energy efficient communications for reliable IoT multicast 5G/ satellite services. Future Internet, 11(8):164, July 25, 2019. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/11/8/164.

Chiti:2021:TQI

[CFPP21] Francesco Chiti, Romano Fantacci, Roberto Picchi, and Laura Pierucci. Towards the quantum Internet: Satellite control plane architectures and protocol design. Future Internet, 13(8):196, July 30, 2021. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
8/196.

Celesti:2019:SJO

[CFV19] Antonio Celesti, Maria Fazio, and Massimo Villari. A study on join operations in MongoDB preserving collections data models for future Internet applications. Future Internet, 11(4):83, March 27, 2019. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
4/83.

Curreri:2020:ISM

[CFX20] Francesco Curreri, Giacomo Fiumara, and Maria Gabriella Xibilia. Input selection methods for soft sensor design: a survey. Future Internet, 12 (6):97, June 04, 2020. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/12/
6/97.

Cotti:2024:EEU

[CGBF24] Luca Cotti, Davide Guizzardi, Barbara Rita Barricelli, and Daniela Fogli. Enabling end-user development in smart homes: a machine learning-powered digital twin for energy efficient management. Future Internet, 16(6):208, June 14, 2024. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/16/6/208.

Carroll:2022:PMM

[CGGK22] John M. Carroll, Fanlu Gui, Srishti Gupta, and Tiffany Knearem. Playful meaning-making as prosocial fun. Future Internet, 14(10):288, September 30, 2022. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/14/10/288.

Chen:2019:FBB

[CGLR19] Yuling Chen, Jinyi Guo, Changlou Li, and Wei Ren. FaDe: a blockchain-based fair data exchange scheme for big data sharing. Future Internet, 11(11):225, October 24, 2019. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/11/
                                        11/225.

Condoluci:2019:VSL

[CGM^{+}19] Massimo Condoluci, Laurent Gallo, Laurent Mussot, Apostolos Kousaridas, Panagiotis Spapis, Maliheh Mahlouji, and Toktam Mahmoodi. 5G V2X system-level architecture of 5GCAR project. Future Internet, 11(10):217, October 19, 2019. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
10/217.

Carchiolo:2022:CAN

[CGMM22] Vincenza Carchiolo, Marco Grassia, Michele Malgeri, and Giuseppe Mangioni. Co-authorship networks analysis to discover collaboration patterns among Italian researchers. Future Internet, 14(6):187, June 16, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/6/187.

Corrales-Garay:2019:ODO

[CGMVdUC19] Diego Corrales-Garay, Eva-María Mora-Valentín, and Marta Ortiz de Urbina-Criado. Open data for open innovation: An analysis of literature characteristics. Future Internet, 11(3):77, March 24, 2019. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
3/77.

Cao:2020:LHG

[CGnCD20] Kerang Cao, Jingyu Gao, Kwang nam Choi, and Lini Duan. Learning a hierarchical global attention for image classification. Future Internet, 12(11):178, October 22, 2020. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/12/
                                        11/178.

Cavazza:2020:DIA

[CGRV20] Francesco Cavazza, Francesco Galioto, Meri Raggi, and Davide Viaggi. Digital irrigated agriculture: Towards a framework for comprehensive analysis of decision processes under uncertainty. Future Internet, 12(11):181, October 26, 2020. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/12/11/181.

Chiluveru:2021:DSM

[CGT21] Rajeshwari Chiluveru, Nishu Gupta, and Ariel Soares Teles. Distribution of safety messages using mobility-aware multi-hop clustering in vehicular ad hoc network. Future Internet, 13(7):169, June 29, 2021. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/13/7/169.

Chinnachodteeranun:2016:SOS

[CH16] Rassarin Chinnachodteeranun and Kiyoshi Honda. Sensor observation service API for providing gridded climate data to agricultural applications. Future Internet, 8(3):40, August 09, 2016. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/8/3/40.

Chen:2023:TGK

[CH23] Chin-Yi Chen and Jih-Jeng Huang. Temporalguided knowledge graphenhanced graph convolutional network for personalized movie recommendation systems. Future Internet, 15(10):323, September 28, 2023. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/15/10/323.

Chinnachodteeranun:2016:DIW

[CHH^{+}16] Rassarin Chinnachodteeranun, Nguyen Duy Hung, Kiyoshi Honda, Amor V. M. Ines, and Eunjin Han. Designing and implementing weather generators as Web services. Future Internet, 8(4):55, December 15, 2016. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/8/4/55.

Chen:2021:UTR

[CHMZ21] Jie Chen, Bing Han, Xufeng Ma, and Jian Zhang. Underwater target recognition based on multi-decision LOFAR spectrum enhancement: a deep-learning approach. Future Internet, 13(10): 265, October 13, 2021. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/13/
                                        10/265.

Chung:2023:AEH

[Chu23] Yao-Liang Chung. Application of an effective hierarchical deep-learning-based object detection model integrated with image-processing techniques for detecting speed limit signs, rockfalls, potholes, and car crashes. Future Internet, 15(10):322, September 28, 2023. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/15/
                                        10/322.

Chen:2021:MMA

[CHW^{+}21] Yanbo Chen, Jingsha He, Wei Wei, Nafei Zhu, and Cong Yu. A multi-model approach for user portrait. Future Internet, 13 (6):147, May 31, 2021. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/13/
                                        6/147.

Chochtoula:2022:IEC

[CISM22] Despoina Chochtoula, Aristidis Ilias, Yannis C. Stamatiou, and Christos Makris. Integrating elliptic curve cryptography with the Modbus TCP SCADA Communication Protocol. Future Internet, 14(8):232, July 28, 2022. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
8/232.

Ciuffoletti:2019:DOR

[Ciu19] Augusto Ciuffoletti. Design of an open remote electrocardiogram (ECG) service. Future Internet, 11(4):101, April 24, 2019. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
4/101.

Chaabane:2022:LPB

[CKFH22] Faten Chaabane, Jalel Ktari, Tarek Frikha, and Habib Hamam. Low power blockchained E-vote platform for university environment. Future Internet, 14(9):269, September 19, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/9/269.

Chatzoglou:2021:MTS

[CKK21]  Efstratios  Chatzoglou,

Georgios Kambourakis, and Vasileios Kouliaridis. A multi-tier security analysis of official car management apps for Android. Future Internet, 13(3):58, February 25, 2021. CO-DEN ???? ISSN 1999-

5903. URL https://www.

                                        mdpi.com/1999-5903/13/
                                        3/58.

Corujo:2021:ERH

[CKSG21] Luis A. Corujo, Emily Kieson, Timo Schloesser, and Peter A. Gloor. Emotion recognition in horses with convolutional neural networks. Future Internet, 13(10):250, September 28, 2021. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/13/10/250.

Chang:2016:PQF

[CL16] Gary Chang and Chung-Chieh Lee. Priority queues with fractional service for tiered delay QoS. Future Internet, 8(1):1, December 29, 2016. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/8/1/1.

Cremer:2019:AII

[CL19] Stefan Cremer and Claudia Loebbecke. Artificial intelligence imagery analysis fostering big data analytics. Future Internet, 11 (8):178, August 15, 2019. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/11/
                                        8/178.

Cen:2022:FMC

[CLC^{+}22] Yuefeng Cen, Mingxing Luo, Gang Cen, Cheng Zhao, and Zhigang Cheng.

Financial market correlation analysis and stock selection application based on TCN-deep clustering. Future Internet, 14(11): 331, November 14, 2022. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
11/331.

Chen:2023:MAJ

[CLH^{+}23] Dechuan Chen, Jin Li, Jianwei Hu, Xingang Zhang, and Shuai Zhang. Multi-antenna jammerassisted secure short packet communications in IoT networks. Future Internet, 15(10):320, September 26, 2023. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/15/10/320.

Carchiolo:2021:MIU

[CLM^{+}21] Vincenza Carchiolo, Alessandro Longheu, Michele Malgeri, Giuseppe Mangioni, and Marialaura Previti. Mutual influence of users credibility and news spreading in online social networks. Future Internet, 13(5):107, April 25, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/5/107.

Chen:2022:FAP

[CLT^{+}22] Zheyi Chen, Weixian Liao, Pu Tian, Qianlong Wang, and Wei Yu. A fairnessaware peer-to-peer decentralized learning framework with heterogeneous devices. Future Internet, 14(5):138, April 30, 2022. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
5/138.

Chen:2022:MVF

[CLTP22] Qifan Chen, Yang Lu, Charmaine S. Tam, and Simon K. Poon. A multiview framework to detect redundant activity labels for more representative event logs in process mining. Future Internet, 14(6):181, June 09, 2022. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
6/181.

Chang:2023:WEH

[CLW23] Hsiao-Ching Chang, Hsing-Tsung Lin, and Pi-Chung Wang. Wireless energy harvesting for Internet-of-Things devices using directional antennas. Future Internet, 15(9):301, September 03, 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
9/301.

Chen:2018:CPM

[CLZL18] Zubin Chen, Baijun Lu, Yanzhou Zhu, and Hao Lv. A compact printed monopole antenna for WiMAX/WLAN and UWB applications. Future Internet, 10(12):122, December 13, 2018. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/10/12/122.

Christin:2010:SWS

[CMH10] Delphine Christin, Parag S.

Mogre, and Matthias Hollick. Survey on wireless sensor network technologies for industrial automation: The security and quality of service perspectives. Future Internet, 2 (2):96–125, April 08, 2010. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/2/
2/96.

Carrino:2016:IIT

[CMK^{+}16] Francesco Carrino, Elena Mugellini, Omar Abou Khaled, Nabil Ouerhani, and Juergen Ehrensberger. iNUIT: Internet of Things for urban innovation. Future Internet, 8(2):18, May 11, 2016. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/8/2/18.

Cantini:2021:EML

[CMO^{+}21] Riccardo Cantini, Fabrizio Marozzo, Alessio Orsino, Domenico Talia, and Paolo Trunfio. Exploiting machine learning for improving in-memory execution of data-intensive workflows on parallel machines. Future Internet, 13(5):121, May 05, 2021. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
5/121.

Carta:2019:FCP

[CMP^{+}19] Salvatore Carta, Andrea Medda, Alessio Pili, Diego Reforgiato Recupero, and Roberto Saia. Forecasting e-commerce products prices by combining an autoregressive integrated moving average (ARIMA) model and Google Trends data. Future Internet, 11(1):5, December 24, 2019. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/11/1/5.

Calvagna:2024:BBR

[CMPT24] Andrea Calvagna, Giovanni Marotta, Giuseppe Pappalardo, and Emiliano Tramontana. A blockchain-based real-time power balancing service for trustless renewable energy grids. Future Internet, 16 (5):149, April 26, 2024. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
5/149.

Corda:2024:CAB

[CMR24] Erica Corda, Silvia M.

Massa, and Daniele Riboni. Context-aware behavioral tips to improve sleep quality via machine learning and large language models. Future Internet, 16(2):46, January 30, 2024. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/16/2/46.

Cunha:2021:BCB

[CMS21] Paulo Rupino Cunha, Paulo Melo, and Helder Sebastião. From bitcoin to central bank digital currencies: Making sense of the digital money revolution. Future Internet, 13(7):165, June 27, 2021. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
7/165.

Chang:2018:SAI

[CMW18] Chia-Lin Chang, Michael McAleer, and Yu-Chieh Wu. A statistical analysis of industrial penetration and Internet intensity in Taiwan. Future Internet, 10(3):31, March 16, 2018. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/10/
3/31.

Chajia:2024:CCP

[CN24] Meryem Chajia and El Habib Nfaoui. Customer churn prediction approach based on LLM embeddings and logistic regression. Future Internet, 16(12):453, December 3, 2024. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/16/
                                        12/453.

Cretan:2021:ISE

[CNC^{+}21] Adina Cretan, Cristina Nica, Carlos Coutinho, Ricardo Jardim-Goncalves, and Ben Bratu. An intelligent system to ensure interoperability for the dairy farm business model. Future Internet, 13(6):153, June 12, 2021. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/13/
                                        6/153.

Chua:2021:TUP

[COL21] Jiaming Chua, Lee-Yeng Ong, and Meng-Chew Leow. Telehealth using PoseNet-based system for in-home rehabilitation. Future Internet, 13(7):173, July 02, 2021. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/13/
                                        7/173.

Chung:2022:CAS

[COL22] Jen-Li Chung, Lee-Yeng Ong, and Meng-Chew Leow. Comparative analysis of skeleton-based human pose estimation. Future Internet, 14(12):380, December 15, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
12/380.

Copeland:2014:UPV

[Cop14] Andrea Copeland. The use of personal value estimations to select images for preservation in public library digital community collections. Future Internet, 6(2):359–377, May 27, 2014. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/6/2/359.

Calegari:2010:OBI

[CP10] Silvia Calegari and Gabriella Pasi. Ontology-based information behaviour to improve Web search. Future Internet, 2(4):533– 558, October 18, 2010. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/2/
4/533.

Collodi:2018:PRC

[CPI^{+}18] Stefania Collodi, Sara Panerati, Enrico Imbimbo, Federica Stefanelli, Mirko Duradoni, and Andrea Guazzini. Personality and reputation: a complex relationship in virtual environments. Future Internet, 10(12):120, December 01, 2018. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/10/12/120.

Chang:2023:SUF

[CPKK23] Sang-Yoon Chang, Kyungmin Park, Jonghyun Kim, and Jinoh Kim. Securing UAV flying base station for mobile networking: a review. Future Internet, 15(5):176, May 09, 2023. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/15/
                                        5/176.

Cocco:2017:BBC

[CPM17] Luisanna Cocco, Andrea Pinna, and Michele Marchesi. Banking on blockchain: Costs savings thanks to the blockchain technology. Future Internet, 9(3):25, June 27, 2017. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/9/
                                        3/25.

Capra:2019:ECS

[CPM^{+}19] Maurizio Capra, Riccardo Peloso, Guido Masera, Massimo Ruo Roch, and Maurizio Martina. Edge computing: a survey on the hardware requirements in the Internet of Things world. Future Internet, 11(4):100, April 23, 2019. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/11/
                                        4/100.

Casadei:2020:PCP

[CPP^{+}20] Roberto Casadei, Danilo Pianini, Andrea Placuzzi, Mirko Viroli, and Danny Weyns. Pulverization in cyber-physical systems: Engineering the selforganizing logic separated from deployment. Future Internet, 12(11):203, November 19, 2020. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/12/
11/203.

Carta:2020:LFE

[CPRS20] Salvatore Carta, Alessandro Sebastian Podda, Diego Reforgiato Recupero, and Roberto Saia. A local feature engineering strategy to improve network anomaly detection. Future Internet, 12(10): 177, October 21, 2020. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/12/
10/177.

Chen:2022:CDT

[CQWC22] Xiaolin Chen, Qixing Qu, Chengxi Wei, and Shudong Chen. Crossdomain transfer learning prediction of COVID-19 popular topics based on knowledge graph. Future Internet, 14(4):103– ??, March 24, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.

                                        mdpi.com/1999-5903/14/
                                        4/103.

Coltekin:2011:HQG

[CR11] Arzu Coltekin and Tumasch Reichenbacher. High quality geographic services and bandwidth limitations. Future Internet, 3(4):379–396, December 20, 2011. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/3/4/379.

Casini:2021:BAI

[CR21] Luca Casini and Marco Roccetti. A Bayesian analysis of the inversion of the SARS-COV-2 case rate in the countries of the 2020 European Football Championship. Future Internet, 13(8):212, August 17, 2021. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/13/8/212.

Cheimonidis:2023:DRA

[CR23a] Pavlos Cheimonidis and Konstantinos Rantos. Dynamic risk assessment in cybersecurity: a systematic literature review. Future Internet, 15(10):324, September 28, 2023. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/15/
                                        10/324.

Chen:2023:FCU

[CR23b] Renjie Chen and Nalini Ravishanker. Feature construction using persistence landscapes for clustering noisy IoT time series. Future Internet, 15(6):195– ??, June 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
6/195.

Carvajal-Rodriguez:2023:PPA

[CRMT23] Jorge Carvajal-Rodriguez, Marco Morales, and Christian Tipantuña. 3D path planning algorithms in UAV-enabled communications systems: a mapping study. Future Internet, 15 (9):289, August 27, 2023. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
9/289.

Croft:2019:EHC

[Cro19] Paul J. Croft. Environmental hazards: a coverage response approach. Future Internet, 11(3):72, March 14, 2019. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/11/3/72.

Chay:2011:UOT

[CS11] Sengtha Chay and Nophea Sasaki. Using online tools to assess public responses to climate change mitigation policies in Japan.

Future Internet, 3(2):117– 129, April 01, 2011. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/3/
2/117.

Claus:2022:NLP

[CS22] Stefan Claus and Massimo Stella. Natural language processing and cognitive networks identify UK insurers’ trends in investor day transcripts. Future Internet, 14(10):291, October 12, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/10/291.

Chen:2024:AEM

[CSA^{+}24] Jieli Chen, Kah Phooi Seng, Li Minn Ang, Jeremy Smith, and Hanyue Xu. AI-empowered multimodal hierarchical graphbased learning for situation awareness on enhancing disaster responses. Future Internet, 16(5):161, May 7, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/5/161.

Cui:2021:STP

[CSC21] Guowei Cui, Wei Shuai, and Xiaoping Chen. Semantic task planning for service robots in open worlds. Future Internet, 13 (2):49, February 17, 2021. CODEN ???? ISSN 1999-

5903. URL https://www.

mdpi.com/1999-5903/13/
2/49.

Castillo-Soria:2023:PRA

[CSDAM^{+}23] Francisco R. Castillo-Soria, J. Alberto Del Puerto-Flores, Cesar A. Azurdia-Meza, Vinoth Babu Kumaravelu, Jorge Simón, and Carlos A. Gutierrez. Precoding for RIS-assisted multi-user MIMO-DQSM transmission systems. Future Internet, 15(9):299, September 02, 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
9/299.

C:2019:CCK

[CSIS19] Pramod T. C., Thejas G.

S., S. S. Iyengar, and N. R. Sunitha. CKMI: Comprehensive key management infrastructure design for industrial automation and control systems. Future Internet, 11(6):126, June 04, 2019. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/11/6/126.

Chaves:2020:EEC

[CSL^{+}20] Ricardo Chaves, Carlos Senna, Miguel Luís, Susana Sargento, André Moreira, Diogo Recharte, and Ricardo Matos. EmuCD: an emulator for content dissemination protocols in vehicular networks. Future Internet, 12(12):234, December 21, 2020. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/12/
                                        12/234.

Chen:2024:MSD

[CSZ^{+}24] Fanghua Chen, Deguang Shang, Gang Zhou, Ke Ye, and Guofang Wu. Multisource data fusion for vehicle maintenance project prediction. Future Internet, 16(10), October 14, 2024. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/16/10/371.

Ce:2020:AMI

[CT20] Peng Ce and Bao Tie. An analysis method for interpretability of CNN text classification model. Future Internet, 12(12):228, December 13, 2020. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/12/
                                        12/228.

Cocco:2024:DTC

[CT24a] Luisanna Cocco and Roberto Tonelli. Digital transformation in the construction sector: Blockchain, BIM and SSI for a more sustainable and transparent system. Future Internet, 16 (7):??, June 28, 2024. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/16/
                                        7/232.

Cocco:2024:SSI

[CT24b] Luisanna Cocco and Roberto Tonelli. A self-sovereign identity–blockchain-based model proposal for deep digital transformation in the healthcare sector. Future Internet, 16(12):473, December 19, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
12/473.

Cocco:2019:ABM

[CTM19] Luisanna Cocco, Roberto Tonelli, and Michele Marchesi. An agent based model to analyze the Bitcoin mining activity and a comparison with the gold mining industry. Future Internet, 11(1):8, January 02, 2019. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/11/1/8.

Cocco:2021:BSS

[CTM21] Luisanna Cocco, Roberto Tonelli, and Michele Marchesi. Blockchain and self sovereign identity to support quality in the food supply chain. Future Internet, 13(12):301, November 26, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/12/301.

Cocco:2022:BSH

[CTM22a] Luisanna Cocco, Roberto Tonelli, and Michele Marchesi. Bitcoin as a safe haven during COVID-19 disease. Future Internet, 14(4):98– ??, March 22, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
4/98.

Cocco:2022:SPI

[CTM22b] Luisanna Cocco, Roberto Tonelli, and Michele Marchesi. A system proposal for information management in building sector based on BIM, SSI, IoT and blockchain. Future Internet, 14(5):140, April 30, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/5/140.

Chen:2021:DPB

[CTNS21] Xuan Chen, Shujuan Tian, Kien Nguyen, and Hiroo Sekiya. Decentralizing private blockchain-IoT network with OLSR. Future Internet, 13(7):168, June 28, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/7/168.

Cheimaras:2022:LCO

[CTP^{+}22] Vasileios Cheimaras, Athanasios Trigkas, Panagiotis Papageorgas, Dimitrios Piromalis, and Emmanouil Sofianopoulos. A lowcost open-source architecture for a digital signage emergency evacuation system for cruise ships, based on IoT and LTE/4G technologies. Future Internet, 14(12):366, December 07, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/12/366.

Chaganti:2022:BBC

[CVG^{+}22] Rajasekhar Chaganti, Vijayakumar Varadarajan, Venkata Subbarao Gorantla, Thippa Reddy Gadekallu, and Vinayakumar Ravi. Blockchain-based cloudenabled security monitoring using Internet of things in smart agriculture. Future Internet, 14(9):250, August 24, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
9/250.

Cheng:2023:ATM

[CW23] Bingbing Cheng and Jiao Wu. Acoustic TDOA measurement and accurate indoor positioning for smartphone. Future Internet, 15 (7):240–??, July 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
7/240.

Chen:2022:ULP

[CWC^{+}22] Xieling Chen, Fu Lee Wang, Gary Cheng, Man-Kong Chow, and Haoran Xie. Understanding learners’ perception of MOOCs based on review data analysis using deep learning and sentiment analysis. Future Internet, 14(8): 218, July 25, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
8/218.

Chadwick:2013:IAP

[CWF13] Darren Chadwick, Caroline Wesson, and Chris Fullwood. Internet access by people with intellectual disabilities: Inequalities and opportunities. Future Internet, 5 (3):376–397, July 17, 2013. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/5/
3/376.

Cheng:2022:RRE

[CXG^{+}22] Hengfei Cheng, Zhaobin Xu, Xiaoxu Guo, Jia Yang, Kedi Xu, Shuqin Liu, Zhonghe Jin, and Xiaojun Jin. Research on routing equalization algorithm of inter-satellite partition for low-orbit microsatellites. Future Internet, 14(7):207, July 04, 2022. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
7/207.

Chen:2021:PRIb

[CXLB21] Xingyuan Chen, Huahu Xu, Yang Li, and Minjie Bian. Person reidentification by lowdimensional features and metric learning. Future Internet, 13(11):289, November 18, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/11/289.

Chen:2023:EEH

[CXZ23] Yonghong Chen, Lei Xun, and Shibing Zhang. The energy efficiency of heterogeneous cellular networks based on the Poisson hole process. Future Internet, 15(2):56, January 30, 2023. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/15/2/56.

Che:2021:KGB

[CYC^{+}21] Mingxuan Che, Kui Yao, Chao Che, Zhangwei Cao, and Fanchen Kong. Knowledge-graph-based drug repositioning against COVID-19 by graph convolutional network with attention mechanism. Future Internet, 13(1):13, January 07, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/1/13.

Chikaraishi:2017:CSA

[CYO^{+}17]  Takenobu  Chikaraishi,

Yuichiro Yoshikawa, Kohei Ogawa, Oriza Hirata, and Hiroshi Ishiguro. Creation and staging of Android theatre “Sayonara” towards developing highly human-like robots. Future Internet, 9(4):75, November 02, 2017. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/9/4/75.

Chen:2017:DBS

[CZK^{+}17] Yongqun Chen, Huaibei Zhou, Ruoshan Kong, Li Zhu, and Huaqing Mao. Decentralized blind spectrum selection in cognitive radio networks considering handoff cost. Future Internet, 9(2):10, March 31, 2017. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/9/2/10.

Chen:2021:PRIa

[CZL21] Shengbo Chen, Hongchang Zhang, and Zhou Lei. Person re-identification based on attention mechanism and context information fusion. Future Internet, 13(3):72, March 13, 2021. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/13/
                                        3/72.

Chen:2017:TRU

[CZY17] Haibao Chen, Yuyan Zhao, and Chuxiong Yan. Towards rack utilization in Internet datacenters: An approach based on dynamic programming. Future Internet, 9(2):17, May 06, 2017. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/9/2/17.

Cheng:2015:TPA

[CZZH15] Xiu-Zhi Cheng, Da-Rong Zhu, Shen Zhang, and Ping He. Tracking positioning algorithm for direction of arrival based on direction lock loop. Future Internet, 7(3):214– 224, June 26, 2015. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/7/
3/214.

Cai:2022:GDA

[CZZZ22] Liewu Cai, Lei Zhu, Hongyan Zhang, and Xinghui Zhu. DA-GAN: Dual attention generative adversarial network for cross-modal retrieval. Future Internet, 14(2):43, January 27, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
2/43.

Dorgham:2019:SSP

[DAMAS^{+}19] Osama Dorgham, Ibrahim Al-Mherat, Jawdat Al-Shaer, Sulieman Bani-Ahmad, and Stephen Laycock. Smart system for prediction of accurate surface electromyography signals using an artificial neural network. Future Internet, 11(1):25, January 21, 2019. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/11/1/25.

Davis:2012:SME

[Dav12] Jenny L. Davis. Social media and experiential ambivalence. Future Internet, 4(4):955–970, October 26, 2012. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/4/4/955.

Federico:2023:CEA

[DB23] Gemma Di Federico and Andrea Burattin. CvA-MoS — event abstraction using contextual information. Future Internet, 15 (3):113, March 18, 2023. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/15/
                                        3/113.

Demidova:2014:AEF

[DBD^{+}14] Elena Demidova, Nicola Barbieri, Stefan Dietze, Adam Funk, Helge Holzmann, Diana Maynard, Nikolaos Papailiou, Wim Peters, Thomas Risse, and Dimitris Spiliotopoulos. Analysing and enriching focused Semantic Web archives for Parliament applications. Future Internet, 6(3):433– 456, July 30, 2014. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/6/
3/433.

Diaz:2016:CAT

[DCA16] Paloma Díaz, John M.

Carroll, and Ignacio Aedo. Coproduction as an approach to technologymediated citizen participation in emergency management. Future Internet, 8 (3):41, August 10, 2016. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/8/
3/41.

Duarte:2023:MAR

[DCA23] Rui P. Duarte, Carlos A. S. Cunha, and Valter N. N. Alves. Mobile application for real-time food plan management for Alzheimer patients through design-based research. Future Internet, 15(5):168, April 29, 2023. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
5/168.

Diogo:2019:CMN

[DCB19] Miguel Diogo, Bruno Cabral, and Jorge Bernardino. Consistency models of NoSQL databases. Future Internet, 11(2):43, February 14, 2019. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/11/2/43.

Dashkevich:2024:BFS

[DCD24] Natalia Dashkevich, Steve Counsell, and Giuseppe Destefanis. Blockchain financial statements: Innovating financial reporting, accounting, and liquidity management. Future Internet, 16(7):??, July 9, 2024. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/16/7/244.

deMeloeSilva:2020:MES

[dCdG20] Alessandra de Melo e Silva, João José Costa Gondim, Robson de Oliveira Albuquerque, and Luis Javier García Villalba. A methodology to evaluate standards and platforms within cyber threat intelligence. Future Internet, 12(6):108, June 23, 2020. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/12/
                                        6/108.

daCruz:2021:ORO

[dCdPC^{+}21] Mauro A. A. da Cruz, Heitor T. L. de Paula, Bruno P. G. Caputo, Samuel B. Mafra, Pascal Lorenz, and Joel J. P. C. Rodrigues. OLP — a RESTful open low-code platform. Future Internet, 13(10):249, September 25, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/10/249.

deCurto:2024:ECS

[dCdZCC24] J. de Curtò, I. de Zarzà, Juan-Carlos Cano, and Carlos T. Calafate. Enhancing communication security in drones using QRNG in frequency hopping spread spectrum. Future Internet, 16(11):412, November 8, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
11/412.

Davies:2012:PEA

[DCG12] John N. Davies, Paul Comerford, and Vic Grout. Principles of eliminating access control lists within a domain. Future Internet, 4(2):413–429, April 19, 2012. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/4/2/413.

Deng:2016:RCM

[DCHM16] Ziyun Deng, Lei Chen, Tingqing He, and Tao Meng. A reliability calculation method for Web service composition using fuzzy reasoning colored Petri nets and its application on supercomputing cloud platform. Future Internet, 8(4):47, September 27, 2016. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/8/4/47.

Dolhopolov:2024:IFG

[DCL24] Anton Dolhopolov, Arnaud Castelltort, and Anne Laurent. Implementing federated governance in data mesh architecture. Future Internet, 16(4):115, March 29, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
4/115.

daCosta:2021:DCM

[dCMA21] Nuno Marques da Costa, Nelson Mileu, and André Alves. Dashboard COM-PRIME COMPRI MOv: Multiscalar spatio-temporal monitoring of the COVID-19 pandemic in Portugal. Future Internet, 13(2):45, February 12, 2021. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
2/45.

Duradoni:2021:RSI

[DCPG21] Mirko Duradoni, Stefania Collodi, Serena Coppolino Perfumi, and Andrea Guazzini. Reviewing stranger on the Internet: The role of identifiability through “Reputation” in online decision making. Future Internet, 13(5):110, April 27, 2021. CODEN ???? ISSN 1999-

5903. URL https://www.

mdpi.com/1999-5903/13/
5/110.

Duong:2022:UHC

[DDD^{+}22] Tan Nghia Duong, Nguyen Nam Doan, Truong Giang Do, Manh Hoang Tran, Duc Minh Nguyen, and Quang Hieu Dang. Utilizing half convolutional autoencoder to generate user and item vectors for initialization in matrix factorization. Future Internet, 14(1):20, January 04, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/1/20.

Dhamala:2024:PEG

[DDMA24] Binita Kusum Dhamala, Babu R. Dawadi, Pietro Manzoni, and Baikuntha Kumar Acharya. Performance evaluation of graph neural networkbased RouteNet model with attention mechanism. Future Internet, 16(4):116, March 29, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
4/116.

Dechouniotis:2018:FMC

[DDPVP18] Dimitrios Dechouniotis, Ioannis Dimolitsas, Konstantinos Papadakis-Vlachopapadopoulos, and Symeon Papavassiliou. Fuzzy multi-criteria based trust management in heterogeneous federated future Internet testbeds. Future Internet, 10(7):58, June 25, 2018. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/10/7/58.

DeFazio:2022:BCP

[DDV22] Roberto De Fazio, Massimo De Vittorio, and Paolo Visconti. A BLEconnected piezoresistive and inertial chest band for remote monitoring of the respiratory activity by an Android application: Hardware design and software optimization. Future Internet, 14(6):183, June 11, 2022. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/14/6/183.

Diniz:2018:CRC

[DDZ18] Júlio César Medeiros Diniz, Francesco Da Ros, and Darko Zibar. Clock recovery challenges in DSPbased coherent singlemode and multi-mode optical systems. Future Internet, 10(7):59, June 26, 2018. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/10/7/59.

DeDeo:2016:CCW

[DeD16]  Simon DeDeo.  Con-

flict and computation on Wikipedia: a finite-state machine analysis of Editor interactions. Future Internet, 8(3):31, July 08, 2016. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/8/
3/31.

Delgado:2021:TPC

[Del21] Francisco Delgado. Teaching physics for computer science students in higher education during the COVID-19 pandemic: a fully Internet-supported course. Future Internet, 13 (2):35, January 29, 2021. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
2/35.

Delgado:2023:ELC

[Del23] Francisco Delgado. Extending learning and collaboration in quantum information with Internet support: a future perspective on research education beyond boundaries, limitations, and frontiers. Future Internet, 15(5):160, April 26, 2023. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/15/5/160.

Denning:2012:SWC

[Den12] Dorothy E. Denning.

Stuxnet: What has changed? Future Internet, 4(3):672– 687, July 16, 2012. CO-DEN ???? ISSN 1999-

5903. URL https://www.

mdpi.com/1999-5903/4/
3/672.

Desmet:2013:GAM

[DG13] Antoine Desmet and Erol Gelenbe. Graph and analytical models for emergency evacuation. Future Internet, 5(1):46–55, February 21, 2013. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/5/
1/46.

DAmico:2018:RCK

[DG18] Antonella D’Amico and Domenico Guastella. Robotics construction kits: From “Objects to think with” to “Objects to think and to emote with”. Future Internet, 10(2):21, February 24, 2018. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/10/2/21.

Dalmau:2014:RDM

[DGADE14] Francesc Valls Dalmau, Pilar Garcia-Almirall, Ernest Redondo Domínguez, and David Fonseca Escudero. From raw data to meaningful information: a representational approach to cadastral databases in relation to urban planning. Future Internet, 6(4):612– 639, October 24, 2014. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/6/
4/612.

Dumitru-Guzu:2024:NFC

[DGCK24] Oana-Mihaela Dumitru-Guzu, Vladeanu Calin, and Robert Kooij. A novel framework for crosscluster scaling in cloudnative 5G NextGen core. Future Internet, 16(9), September 6, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
9/325.

DeDonno:2019:CSC

[DGD^{+}19] Michele De Donno, Alberto Giaretta, Nicola Dragoni, Antonio Bucchiarone, and Manuel Mazzara. Cyber-storms come from clouds: Security of cloud computing in the IoT era. Future Internet, 11(6):127, June 04, 2019. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/11/6/127.

Du:2023:FSA

[DGF23] Jinze Du, Chengtai Gao, and Tao Feng. Formal safety assessment and improvement of DDS protocol for industrial data distribution service. Future Internet, 15(1):24, December 31, 2023. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/15/1/24.

Donati:2019:ALA

[DGGS19] Camillo Donati, Andrea Guazzini, Giorgio Gronchi, and Andrea Smorti. About Linda again: How narratives and group reasoning can influence conjunction fallacy. Future Internet, 11(10): 210, October 08, 2019. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
10/210.

Derakhshannia:2020:DLG

[DGHH^{+}20] Marzieh Derakhshannia, Carmen Gervet, Hicham Hajj-Hassan, Anne Laurent, and Arnaud Martin. Data lake governance: Towards a systemic and natural ecosystem analogy. Future Internet, 12 (8):126, July 27, 2020. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/12/
8/126.

Drogkaris:2015:HMA

[DGKL15] Prokopios Drogkaris, Stefanos Gritzalis, Christos Kalloniatis, and Costas Lambrinoudakis. A hierarchical multitier approach for privacy policies in egovernment environments. Future Internet, 7(4):500– 515, December 21, 2015. CODEN ???? ISSN 1999-

5903. URL https://www.

mdpi.com/1999-5903/7/
4/500.

DAmbrosio:2021:SAP

[DGMP21] Raffaele D’Ambrosio, Giuseppe Giordano, Serena Mottola, and Beatrice Paternoster. Stiffness analysis to predict the spread out of fake information. Future Internet, 13(9):222, August 28, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/9/222.

Deng:2018:MFP

[DH18] Ziyun Deng and Tingqin He. A method for filtering pages by similarity degree based on dynamic programming. Future Internet, 10(12):124, December 13, 2018. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/10/
12/124.

Dang:2022:NCS

[DH22] Shoujiang Dang and Rui Han. An in-network cooperative storage schema based on neighbor offloading in a programmable data plane. Future Internet, 14(1):18, December 30, 2022. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/14/1/18.

Daraghmi:2022:IBS

[DHDAu22] Yousef-Awwad Daraghmi, Mamoun Abu Helou, Eman-Yasser Daraghmi, and Waheeb Abu-ulbeh. IoT-based system for improving vehicular safety by continuous traffic violation monitoring. Future Internet, 14(11):319, November 02, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
11/319.

Dogan:2019:NRD

[DHEK19] Safak Dogan, Nasser Haddad, Erhan Ekmekcioglu, and Ahmet M. Kondoz. No-reference depth map quality evaluation model based on depth map edge confidence measurement in immersive video applications. Future Internet, 11(10):204, September 20, 2019. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/11/10/204.

Delgado:2021:RPS

[DHG21] Gabriela Torres Delgado and Neil Hernández-Gress. Research professors’ selfassessment of competencies. Future Internet, 13 (2):41, February 04, 2021. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
2/41.

Dirir:2021:ADL

[DIE^{+}21] Ahmed Dirir, Henry Ignatious, Hesham Elsayed, Manzoor Khan, Mohammed Adib, Anas Mahmoud, and Moatasem Al-Gunaid. An advanced deep learning approach for multi-object counting in urban vehicular environments. Future Internet, 13(12):306, November 30, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/12/306.

Duradoni:2020:WBS

[DIG20] Mirko Duradoni, Federico Innocenti, and Andrea Guazzini. Well-being and social media: a systematic review of Bergen addiction scales. Future Internet, 12 (2):24, January 29, 2020. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/12/
2/24.

Dangi:2022:MBN

[DJC^{+}22] Ramraj Dangi, Akshay Jadhav, Gaurav Choudhary, Nicola Dragoni, Manas Kumar Mishra, and Praveen Lalwani. MLbased 5G network slicing security: a comprehensive survey. Future Internet, 14 (4):116–??, April 08, 2022. CODEN ???? ISSN 1999-

5903. URL https://www.

                                        mdpi.com/1999-5903/14/
                                        4/116.

DeMedeiros:2024:CCA

[DKH24] Kyle DeMedeiros, Chan Young Koh, and Abdeltawab Hendawi. Clustering on the Chicago Array of things: Spotting anomalies in the Internet of Things records. Future Internet, 16(1):28, January 16, 2024. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/16/1/28.

Davidian:2024:ERD

[DKV24] Matan Davidian, Michael Kiperberg, and Natalia Vanetik. Early ransomware detection with deep learning models. Future Internet, 16(8):??, August 11, 2024. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/16/
                                        8/291.

Ding:2010:MED

[DL10] Lian Ding and Shaofeng Liu. Markup in engineering design: a discourse. Future Internet, 2(1):74– 95, March 11, 2010. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/2/
                                        1/74.

Dong:2019:ACM

[DL19] Ying Dong and Zhou Lei. An access control model for preventing virtual machine hopping attack. Future Internet, 11(3):82, March 26, 2019. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/11/3/82.

Do:2010:SQM

[DLK10] Viet Thi Minh Do, Lars Landmark, and Øivind Kure. A survey of QoS multicast in ad hoc networks. Future Internet, 2(3):388–416, September 14, 2010. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/2/3/388.

Dong:2019:MBN

[DLX^{+}19] Peixin Dong, Dongyuan Li, Jianping Xing, Haohui Duan, and Yong Wu. A method of bus network optimization based on complex network and beidou vehicle location. Future Internet, 11(4):97, April 15, 2019. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/11/4/97.

Liborio:2022:IVD

[dMLMS22] Joao da Mata Liborio, Cesar Melo, and Marcos Silva. Internet video delivery improved by superresolution with GAN. Future Internet, 14(12):364, December 06, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.

                                        mdpi.com/1999-5903/14/
                                        12/364.

Drozdz:2019:SCC

[DMO^{+}19] Stanisław Drozdz, Ludovico Minati, Paweł Oświȩcimka, Marek Stanuszek, and Marcin Wa̧torek. Signatures of the cryptocurrency market decoupling from the Forex. Future Internet, 11(7):154, July 10, 2019. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/11/
                                        7/154.

Fazio:2023:HMI

[DMP^{+}23] Roberto De Fazio, Vincenzo Mariano Mastronardi, Matteo Petruzzi, Massimo De Vittorio, and Paolo Visconti. Human–machine interaction through advanced haptic sensors: a piezoelectric sensory glove with edge machine learning for gesture and object recognition. Future Internet, 15(1):14, December 27, 2023. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/15/1/14.

Dhar:2022:AIL

[DMS^{+}22] Ankita Dhar, Himadri Mukherjee, Shibaprasad Sen, Md Obaidullah Sk, Amitabha Biswas, Teresa Gonçalves, and Kaushik Roy. Author identification from literary articles with visual features: a case study with Bangla documents. Future Internet, 14(10):272, September 23, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/10/272.

Ding:2019:ILL

[DMZ^{+}19] Zhengyang Ding, Weiwei Miao, Mingxuan Zhang, Wei Li, Rui Liu, Jun Zou, and Chen Xu. Integration of LTE 230 and LTE 1800 in power wireless private networks. Future Internet, 11(11):221, October 23, 2019. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/11/11/221.

Diaz-Noci:2020:AIS

[DN20] Javier Díaz-Noci. Artificial intelligence systemsaided news and copyright: Assessing legal implications for journalism practices. Future Internet, 12 (5):85, May 08, 2020. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/12/
5/85.

Dang:2024:RAO

[DNKDN^{+}24] Tuan-Kiet Dang, Khai-Duy Nguyen, Binh Kieu-Do-Nguyen, Trong-Thuc Hoang, and Cong-Kha Pham. Realization of authenticated one-pass key establishment on RISC-V micro-controller for IoT applications. Future Internet, 16(5):157, May 3, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/5/157.

DelPuerto-Flores:2022:CCS

[DNPC^{+}22] Jose Alberto Del Puerto-Flores, José Luis Naredo, Fernando Peña-Campos, Carolina Del-Valle-Soto, Leonardo J. Valdivia, and Ramón Parra-Michel. Channel characterization and SC-FDM modulation for PLC in high-voltage power lines. Future Internet, 14(5):139, April 30, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/5/139.

Duan:2022:DCA

[DNZ22] Yingjie Duan, Hong Ni, and Xiaoyong Zhu. A dynamic cache allocation mechanism (DCAM) for reliable multicast in information-centric networking. Future Internet, 14(4):105–??, March 25, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/4/105.

Duan:2022:SRM

[DNZW22] Yingjie Duan, Hong Ni, Xiaoyong Zhu, and Xu Wang. A single-rate multicast congestion control (SRMCC) mechanism in informationcentric networking. Future Internet, 14(2):38, January 25, 2022. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/14/2/38.

deOliveira:2019:PCN

[dOB19] Fabíola Martins Campos de Oliveira and Edson Borin. Partitioning convolutional neural networks to maximize the inference rate on constrained IoT devices. Future Internet, 11(10):209, September 29, 2019. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/11/10/209.

Dorobantiu:2021:CCE

[DOB21] Alexandru Dorobantiu, Valentin Ogrean, and Remus Brad. Coronary centerline extraction from CCTA using 3D-UNet. Future Internet, 13(4):101, April 19, 2021. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
4/101.

Duradoni:2018:FTV

[DPBG18] Mirko Duradoni, Mario Paolucci, Franco Bagnoli, and Andrea Guazzini. Fairness and trust in virtual environments: The effects of reputation. Future Internet, 10(6):50, June 09, 2018. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/10/6/50.

Daousis:2024:OPS

[DPC^{+}24] Spyridon Daousis, Nikolaos Peladarinos, Vasileios Cheimaras, Panagiotis Papageorgas, Dimitrios D. Piromalis, and Radu Adrian Munteanu. Overview of protocols and standards for wireless sensor networks in critical infrastructures. Future Internet, 16 (1):33, January 21, 2024. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/16/
                                        1/33.

Grande-de-Prado:2020:DCG

[dPCGMC20] Mario Grande de Prado, Ruth Cañón, Sheila García-Martín, and Isabel Cantón. Digital competence and gender: Teachers in training. A case study. Future Internet, 12(11):204, November 20, 2020. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/12/
                                        11/204.

De-Pablos-Heredero:2017:FIS

[DPH17] Carmen De-Pablos-Heredero. Future intelligent systems and networks. Future Internet, 9(3):49, September 02, 2017. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/9/3/49.

De-Pablos-Heredero:2019:FIS

[DPH19] Carmen De-Pablos-Heredero.

Future intelligent systems and networks. Future Internet, 11(6):140, June 25, 2019. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/11/6/140.

Din:2021:ANA

[DPR^{+}21] Fareed Ud Din, David Paul, Joe Ryan, Frans Henskens, and Mark Wallis. AOSR 2.0: a novel approach and thorough validation of an agent-oriented storage and retrieval WMS planner for SMEs, under Industry 4.0. Future Internet, 13(6):155, June 15, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/6/155.

Dai:2018:PPF

[DQ18] Jiazhu Dai and Keke Qiao. A privacy preserving framework for worker’s location in spatial crowdsourcing based on local differential privacy. Future Internet, 10(6):53, June 14, 2018. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/10/6/53.

Dong:2022:AIA

[DRK^{+}22] Jingwen Dong, Siti Nurulain Mohd Rum, Khairul Azhar Kasmiran, Teh Noranis Mohd Aris, and Raihani Mohamed. Artificial intelligence in adaptive and intelligent educational system: a review. Future Internet, 14(9):245, August 23, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
9/245.

DellImmagine:2023:KDM

[DSB23] Giorgio Dell’Immagine, Jacopo Soldani, and Antonio Brogi. KubeHound: Detecting microservices’ security smells in Kubernetes deployments. Future Internet, 15(7):228– ??, July 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
7/228.

Filho:2024:DTM

[dSFdRP24] Vitor Hugo dos Santos Filho, Luis Maurício Martins de Resende, and Joseane Pontes. Development of a theoretical model for digital risks arising from the implementation of Industry 4.0 (TMR-I4.0). Future Internet, 16 (6):215, June 17, 2024. CODEN ???? ISSN 1999-

5903. URL https://www.

mdpi.com/1999-5903/16/
6/215.

Dowdeswell:2021:AAB

[DSM21] Barry Dowdeswell, Roopak Sinha, and Stephen G. MacDonell. Architecting an agent-based fault diagnosis engine for IEC 61499 industrial cyber-physical systems. Future Internet, 13(8):190, July 23, 2021. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
8/190.

Mourao:2022:ILS

[dSMdO22] Helmer Augusto de Souza Mourão and Horácio Antonio Braga Fernandes de Oliveira. Indoor localization system using fingerprinting and novelty detection for evaluation of confidence. Future Internet, 14(2):51, February 07, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
2/51.

Dey:2021:TCM

[DSMM21] Somdip Dey, Amit Kumar Singh, and Klaus McDonald-Maier. ThermalAttackNet: Are CNNs making it easy to perform temperature side-channel attack in mobile edge devices? Future Internet, 13(6):146, May 31, 2021. CODEN ???? ISSN 1999-

5903. URL https://www.

                                        mdpi.com/1999-5903/13/
                                        6/146.

daSilva:2022:FSD

[dSPR^{+}22] Diogo E. Moreira da Silva, Eduardo J. Solteiro Pires, Arsénio Reis, Paulo B. de Moura Oliveira, and João Barroso. Forecasting students dropout: A UTAD university study. Future Internet, 14(3):76, February 28, 2022. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/14/
                                        3/76.

Rocha:2021:BAA

[dSRRdOT21] Geneci da Silva Ribeiro Rocha, Letícia de Oliveira, and Edson Talamini. Blockchain applications in agribusiness: a systematic review. Future Internet, 13(4):95, April 08, 2021. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/13/4/95.

Rocha:2023:BQV

[dSRRMC^{+}23] Geneci da Silva Ribeiro Rocha, Diego Durante Mühl, Hermenegildo Almeida Chingamba, Letícia de Oliveira, and Edson Talamini. Blockchain, quo vadis? Recent changes in perspectives on the application of technology in agribusiness. Future Internet, 15(1):38, January 16, 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
1/38.

Veloso:2021:HHC

[dSVJdALRfS21] Artur Felipe da Silva Veloso, José Valdemir Reis Júnior, Ricardo de Andrade Lira Rabelo, and Jocines Dela flora Silveira. HyDSMaaS: a hybrid communication infrastructure with Lo-RaWAN and LoraMesh for the demand side management as a service. Future Internet, 13(11):271, October 26, 2021. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
11/271.

Dritsas:2024:MLB

[DT24] Elias Dritsas and Maria Trigka. Machine learning for blockchain and IoT systems in smart cities: a survey. Future Internet, 16 (9), September 6, 2024. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
9/324.

Dikhanbayeva:2021:CFI

[DTL^{+}21] Dinara Dikhanbayeva, Akmaral Tokbergenova, Yevgeniy Lukhmanov, Essam Shehab, Zbigniew Pastuszak, and Ali Turkyilmaz. Critical factors of Industry 4.0 implementation in an emerging country: Empirical study. Future Internet, 13(6):137, May 21, 2021. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/13/6/137.

DOrtona:2022:OSM

[DTR22] Cristian D’Ortona, Daniele Tarchi, and Carla Raffaelli. Open-source MQTTbased end-to-end IoT system for smart city scenarios. Future Internet, 14 (2):57, February 15, 2022. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/14/
                                        2/57.

Duc:2019:MCQ

[DTTK19] Tho Nguyen Duc, Chanh Minh Tran, Phan Xuan Tan, and Eiji Kamioka. Modeling of cumulative QoE in on-demand video services: Role of memory effect and degree of interest. Future Internet, 11 (8):171, August 04, 2019. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/11/
                                        8/171.

Duan:2021:IAM

[Dua21] Qiang Duan. Intelligent and autonomous management in cloud-native future networks — a survey on related standards from an architectural perspective. Future Internet, 13 (2):42, February 05, 2021.

CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
2/42.

Dimoulas:2023:TAW

[DV23] Charalampos A. Dimoulas and Andreas Veglis. Theory and applications of Web 3.0 in the media sector. Future Internet, 15(5): 165, April 28, 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
5/165.

deVries:2023:CPM

[dVBA23] Sjouke de Vries, Frank Blaauw, and Vasilios Andrikopoulos. Cost-profiling microservice applications using an APM stack. Future Internet, 15(1):37, January 13, 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
1/37.

DiMartino:2020:MBC

[DVED20] Beniamino Di Martino, Salvatore Venticinque, Antonio Esposito, and Salvatore D’Angelo. A methodology based on computational patterns for offloading of big data applications on cloud-edge platforms. Future Internet, 12(2):28, February 07, 2020. CO-DEN ???? ISSN 1999-

5903. URL https://www.

                                        mdpi.com/1999-5903/12/
                                        2/28.

Del-Valle-Soto:2024:EEC

[DVSBV^{+}24] Carolina Del-Valle-Soto, Ramon A. Briseño, Ramiro Velázquez, Gabriel Guerra-Rosales, Santiago Perez-Ochoa, Isaac H. Preciado-Bazavilvazo, Paolo Visconti, and José Varela-Aldás. Enhancing elderly care through lowcost wireless sensor networks and artificial intelligence: a study on vital sign monitoring and sleep improvement. Future Internet, 16(9), September 6, 2024. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/16/9/323.

Dassanayake:2024:MAD

[DWHY24] Janith K. Dassanayake, Minxiao Wang, Muhammad Z. Hameed, and Ning Yang. Multi-agent deep-Q network-based cache replacement policy for content delivery networks. Future Internet, 16(8):??, August 14, 2024. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/16/
                                        8/292.

Du:2021:BEE

[DWL21] Yao Du, Zehua Wang, and Victor C. M. Leung. Blockchain-enabled edge intelligence for IoT: Background, emerging trends and open issues. Future Internet, 13(2):48, February 17, 2021. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/13/2/48.

Deraman:2019:DIO

[DWY^{+}19a] Rafikullah Deraman, Chen Wang, Jeffrey Boon Hui Yap, Heng Li, and Faizul Azli Mohd-Rahim. Developing Internet online procurement frameworks for construction firms. Future Internet, 11(6):136, June 20, 2019. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/11/6/136.

Dong:2019:VSE

[DWY19b] Huanan Dong, Ming Wen, and Zhouwang Yang. Vehicle speed estimation based on 3D ConvNets and nonlocal blocks. Future Internet, 11(6):123, May 30, 2019. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/11/6/123.

Deng:2016:ELF

[DWZN16] Zhi-An Deng, Di Wu, Yiran Zhou, and Zhenyu Na. Enhanced local Fisher discriminant analysis for indoor positioning in wireless local area network. Future Internet, 8(2):8, March 25, 2016. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/8/2/8.

Dai:2022:ESS

[DXL^{+}22] Zhiqiang Dai, Gaochao Xu, Ziqi Liu, Jiaqi Ge, and Wei Wang. Energy saving strategy of UAV in MEC based on deep reinforcement learning. Future Internet, 14(8):226, July 26, 2022. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/14/8/226.

Dong:2021:DDR

[DYLZ21] Shuai Dong, Zhihua Yang, Wensheng Li, and Kun Zou. Dynamic detection and recognition of objects based on sequential RGB images. Future Internet, 13(7):176, July 07, 2021. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/13/
                                        7/176.

Davaasambuu:2015:SOH

[DYS15] Battulga Davaasambuu, Keping Yu, and Takuro Sato. Self-optimization of handover parameters for long-term evolution with dual wireless mobile relay nodes. Future Internet, 7(2):196–213, June 11, 2015. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/7/2/196.

Deng:2022:HFD

[DZH^{+}22] Shangkun Deng, Yingke Zhu, Xiaoru Huang, Shuangyang Duan, and Zhe Fu. Highfrequency direction forecasting of the futures market using a machinelearning-based method. Future Internet, 14(6):180, June 09, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
6/180.

Elwy:2023:DDS

[EAAA^{+}23] Fatema Elwy, Raafat Aburukba, A. R. Al-Ali, Ahmad Al Nabulsi, Alaa Tarek, Ameen Ayub, , and Mariam Elsayeh. Datadriven safe deliveries: The synergy of IoT and machine learning in shared mobility. Future Internet, 15(10):333, October 10, 2023. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/15/10/333.

Erturk:2019:SLA

[EABE19] Mehmet Ali Ertürk, Muhammed Ali Aydin, Muhammet Talha Büyükakkaslar, and Hayrettin Evirgen. A survey on LoRaWAN architecture, protocol and technologies. Future Internet, 11(10): 216, October 17, 2019. CODEN ???? ISSN 1999-

5903. URL https://www.

mdpi.com/1999-5903/11/
10/216.

ElGhanam:2021:ABD

[EAHO21] Eiman ElGhanam, Ibtihal Ahmed, Mohamed Hassan, and Ahmed Osman. Authentication and billing for dynamic wireless EV charging in an Internet of Electric Vehicles. Future Internet, 13(10):257, October 08, 2021. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
10/257.

Ertek:2022:AME

[EAKM22] Gurdal Ertek, Aysha Al-Kaabi, and Aktham Issa Maghyereh. Analytical modeling and empirical analysis of binary options strategies. Future Internet, 14(7):208, July 06, 2022. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
7/208.

ElDebeiki:2024:APP

[EARP^{+}24] Mostafa El Debeiki, Saba Al-Rubaye, Adolfo Perrusquía, Christopher Conrad, and Juan Alejandro Flores-Campos. An advanced path planning and UAV relay system: Enhancing connectivity in rural environments. Future Internet, 16(3):89, March 06, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/3/89.

Ellawindy:2021:CFQ

[EH21] Ibtihal Ellawindy and Shahram Shah Heydari. Crowdsourcing framework for QoE-aware SD-WAN. Future Internet, 13(8):209, August 15, 2021. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
8/209.

Ebeid:2017:DEC

[EHJ17] Emad Ebeid, Rune Heick, and Rune Hylsberg Jacobsen. Deducing energy consumer behavior from smart meter data. Future Internet, 9(3):29, July 06, 2017. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/9/3/29.

Elkhodr:2024:NSI

[EKG24] Mahmoud Elkhodr, Samiya Khan, and Ergun Gide. A novel semantic IoT middleware for secure data management: Blockchain and AI-driven context awareness. Future Internet, 16(1):22, January 07, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/1/22.

Efthymiadis:2024:AOT

[EKKS24] Filippos Efthymiadis, Aristeidis Karras, Christos Karras, and Spyros Sioutas. Advanced optimization techniques for federated learning on non-IID data. Future Internet, 16(10), October 13, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
10/370.

Eramo:2019:ESR

[ELC^{+}19] Vincenzo Eramo, Francesco G. Lavacca, Tiziana Catena, Marco Polverini, and Antonio Cianfrani. Effectiveness of segment routing technology in reducing the bandwidth and cloud resources provisioning times in network function virtualization architectures. Future Internet, 11(3):71, March 12, 2019. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/11/3/71.

Eramo:2020:PIA

[ELCS20] Vincenzo Eramo, Francesco Giacinto Lavacca, Tiziana Catena, and Paul Jaime Perez Salazar. Proposal and investigation of an artificial intelligence (AI)-based cloud resource allocation algorithm in network function virtualization architectures. Future Internet, 12(11):196, November 13, 2020. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/12/11/196.

Ebner:2009:CWM

[EM09] Martin Ebner and Hermann Maurer. Can weblogs and microblogs change traditional scientific writing? Future Internet, 1(1):47–58, November 18, 2009. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/1/1/47.

Erabadda:2019:QEQ

[EMHF19] Buddhiprabha Erabadda, Thanuja Mallikarachchi, Chaminda Hewage, and Anil Fernando. Quality of experience (QoE)-Aware fast coding unit size selection for HEVC intraprediction. Future Internet, 11(8):175, August 11, 2019. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/11/8/175.

Esposito:2022:SIN

[EMMP22] Massimo Esposito, Giovanni Luca Masala, Aniello Minutolo, and Marco Pota. Special issue “Natural Language Engineering: Methods, Tasks and Applications”. Future Internet, 14(4):106–??, March 26, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/4/106.

Euttamarajah:2021:EEJ

[ENT21] Shornalatha Euttamarajah, Yin Hoe Ng, and Chee Keong Tan. Energyefficient joint base station switching and power allocation for smart grid based hybrid-powered CoMPenabled HetNet. Future Internet, 13(8):213, August 17, 2021. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
8/213.

Epicoco:2024:PDF

[EPC24] Italo Epicoco, Marco Pulimeno, and Massimo Cafaro. Parallel and distributed frugal tracking of a quantile. Future Internet, 16(9), September 13, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/9/335.

Eichler:2010:SSW

[ER10] Frederik Eichler and Wolfgang Reinhardt. Simplifying the scientific writing and review process with SciFlow. Future Internet, 2(4):645–661, December 06, 2010. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/2/4/645.

Eden:2021:RVA

[ER21] Jen Eden and Anthony J. Roberto. The role of verbal aggression in cyberbullying perpetration and victimization by middle school students. Future Internet, 13(9):223, August 30, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/9/223.

Exter:2012:UWT

[ERBL12] Kristin Den Exter, Stephen Rowe, William Boyd, and David Lloyd. Using Web 2.0 technologies for collaborative learning in distance education-case studies from an Australian University. Future Internet, 4(1):216–237, March 07, 2012. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/4/1/216.

Ebner:2020:CEL

[ESB^{+}20] Martin Ebner, Sandra Schön, Clarissa Braun, Markus Ebner, Ypatios Grigoriadis, Maria Haas, Philipp Leitner, and Behnam Taraghi. COVID-19 epidemic as e-learning boost? Chronological development and effects at an Austrian university against the background of the concept of “E-Learning readiness”. Future Internet, 12(6):94, May 26, 2020. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/12/6/94.

ElGaabouri:2023:SLR

[ESBE23] Ismail El Gaabouri, Mohamed Senhadji, Mostafa Belkasmi, and Brahim El Bhiri. A systematic literature review on authentication and threat challenges on RFID based NFC applications. Future Internet, 15(11):354, October 27, 2023. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/15/11/354.

Elagin:2020:BBT

[ESLB20] Vasiliy Elagin, Anastasia Spirkina, Andrei Levakov, and Ilya Belozertsev. Blockchain behavioral traffic model as a tool to influence service IT security. Future Internet, 12 (4):68, April 15, 2020. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/12/
                                        4/68.

Emad-Ud-Din:2023:IOS

[EUDW23] Muhammad Emad-Ud-Din and Ya Wang. Indoor occupancy sensing via networked nodes (2012– 2022): a review. Future Internet, 15(3):116, March 22, 2023. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/15/3/116.

Eramo:2021:PIC

[EVCL21] Vincenzo Eramo, Francesco Valente, Tiziana Catena, and Francesco Giacinto Lavacca. Proposal and investigation of a convolutional and LSTM neural network for the cost-aware resource prediction in softwarized networks. Future Internet, 13(12):316, December 16, 2021. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
12/316.

Effendi:2020:SGD

[EvdMB20] Sedick Baker Effendi, Brink van der Merwe, and Wolf-Tilo Balke. Suitability of graph database technology for the analysis of spatio-temporal data. Future Internet, 12(5):78, April 26, 2020. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/12/5/78.

Elzain:2019:SDW

[EW19] Hisham Elzain and Yang Wu. Software defined wireless mesh network flat distribution control plane. Future Internet, 11(8): 166, July 25, 2019. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
8/166.

Elkhodr:2024:SRM

[EWGK24] Mahmoud Elkhodr, Ketmanto Wangsa, Ergun Gide, and Shakir Karim. A systematic review and multifaceted analysis of the integration of artificial intelligence and blockchain: Shaping the future of Australian higher education. Future Internet, 16(10), October 18, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/10/378.

Fernandez:2012:SPW

[FAB^{+}12] Eduardo B. Fernandez, Ola Ajaj, Ingrid Buckley, Nelly Delessy-Gassant, Keiko Hashizume, and Maria M. Larrondo-Petrie. A survey of patterns for Web services security and reliability standards. Future Internet, 4(2):430– 450, April 20, 2012. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/4/
2/430.

Fabre:2022:RAC

[FAB^{+}22] Renaud Fabre, Otmane Azeroual, Patrice Bellot, Joachim Schöpfel, and Daniel Egret. Retrieving adversarial cliques in cognitive communities: a new conceptual framework for scientific knowledge graphs. Future Internet, 14(9):262, September 07, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/9/262.

Florea:2022:RBT

[FAC22] Alexandru-Ioan Florea, Ionut Anghel, and Tudor Cioara. A review of blockchain technology applications in ambient assisted living. Future Internet, 14(5):150, May 12, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/5/150.

Farman:2012:IST

[Far12] Jason Farman. Introduction to the social transformations from the mobile Internet special issue. Future Internet, 4(2):545– 550, May 23, 2012. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/4/
2/545.

Fabre:2023:MGH

[FAS^{+}23] Renaud Fabre, Otmane Azeroual, Joachim Schöpfel, Patrice Bellot, and Daniel Egret. A multiverse graph to help scientific reasoning from Web usage: Interpretable patterns of assessor shifts in GRAPHYP. Future Internet, 15(4):147, April 10, 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
4/147.

Fragkos:2019:EES

[FAT19] Georgios Fragkos, Pavlos Athanasios Apostolopoulos, and Eirini Eleni Tsiropoulou. ESCAPE: Evacuation strategy through clustering and autonomous operation in public safety systems. Future Internet, 11(1):20, January 17, 2019. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/11/1/20.

Ferilli:2023:GDB

[FBDR23] Stefano Ferilli, Eleonora Bernasconi, Davide Di Pierro, and Domenico Redavid. A graph DBbased solution for semantic technologies in the future Internet. Future Internet, 15(10):345, October 20, 2023. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/15/10/345.

Freina:2015:SEI

[FBOC15] Laura Freina, Rosa Bottino, Michela Ott, and Filippo Costa. Social empowerment of intellectually impaired through a cloud mobile system. Future Internet, 7(4):429–444, November 17, 2015. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/7/4/429.

Fantinelli:2019:SNS

[FC19] Stefania Fantinelli and Michela Cortini. Social network services management and risk of doocing. Comment on Kim, S.; Park, H.; Choi, M. J. “Negative Impact of Social Network Services Based on Stressor–Stress-Outcome: The Role of Experience of Privacy Violations. Future Internet 2019, 11, 137”. Future Internet, 11(9):191, September 04, 2019. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/

9/191. See [KPC19].

Florea:2022:GTA

[FC22] Razvan Florea and Mitica Craus. A game-theoretic approach for network security using honeypots. Future Internet, 14(12):362, November 30, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
12/362.

Fior:2022:LEA

[FCG22] Jacopo Fior, Luca Cagliero, and Paolo Garza. Leveraging explainable AI to support cryptocurrency investors. Future Internet, 14(9):251, August 24, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/9/251.

Fic:2023:ADH

[FCR23] Paweł Fic, Adam Czornik, and Piotr Rosikowski.

Anomaly detection for hydraulic power units — a case study. Future Internet, 15(6):206–??, June 2023. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/15/6/206.

Fominska:2024:AIE

[FDN^{+}24] Iana Fominska, Stefano Di Tore, Michele Nappi, Gerardo Iovane, Maurizio Sibilio, and Angela Gelo. Approaches to identifying emotions and affections during the museum learning experience in the context of the future Internet. Future Internet, 16(11): 417, November 10, 2024. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/16/
                                        11/417.

Fernandez:2012:ISI

[Fer12] Eduardo B. Fernandez. Introduction to the special issue on recent advances in Web services. Future Internet, 4(3):618–620, June 27, 2012. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/4/3/618.

Fragiadakis:2023:AML

[FFM^{+}23] George Fragiadakis, Evangelia Filiopoulou, Christos Michalakelis, Thomas Kamalakis, and Mara Nikolaidou. Applying machine learning in cloud service price prediction: The case of Amazon IaaS. Future Internet, 15(8):277, August 19, 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
8/277.

Ferretti:2022:DLF

[FFP^{+}22] Marco Ferretti, Ugo Fiore, Francesca Perla, Marcello Risitano, and Salvatore Scognamiglio. Deep learning forecasting for supporting terminal operators in port business development. Future Internet, 14(8):221, July 25, 2022. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
8/221.

Ferreira:2021:PIM

[FFVP21] Ana Teresa Ferreira, Carlos Fernandes, José Vieira, and Filipe Portela. Pervasive intelligent models to predict the outcome of COVID-19 patients. Future Internet, 13(4):102, April 20, 2021. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
4/102.

Fourneau:2017:GNA

[FG17] Jean-Michel Fourneau and Erol Gelenbe. G-networks with adders. Future Internet, 9(3):34, July 10, 2017.

CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/9/
                                        3/34.

Fondevila-Gascon:2019:ESU

[FGBMG19] Joan-Francesc Fondevila-Gascón, Gaspar Berbel, and Mònica Muñoz-González. Experimental study on the utility and future of collaborative consumption platforms offering tourism related services. Future Internet, 11(3):80, March 25, 2019. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/11/3/80.

Fallucchi:2022:DQQ

[FGSD22] Francesca Fallucchi, Bouchra Ghattas, Riem Spielhaus, and Ernesto William De Luca. Digital qualitative and quantitative analysis of Arabic textbooks. Future Internet, 14(8): 237, July 29, 2022. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/14/
                                        8/237.

Fath:2024:IIM

[FHL^{+}24] Alireza Fath, Nicholas Hanna, Yi Liu, Scott Tanch, Tian Xia, and Dryver Huston. Indoor infrastructure maintenance framework using networked sensors, robots, and augmented reality human interface. Future Internet, 16(5):170, May 15, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/5/170.

Fuchs:2010:TFW

[FHS^{+}10] Christian Fuchs, Wolfgang Hofkirchner, Matthias Schafranek, Celina Raffl, Marisol Sandoval, and Robert Bichler. Theoretical foundations of the Web: Cognition, communication, and co-operation. towards an understanding of Web 1.0, 2.0, 3.0. Future Internet, 2(1):41–59, February 19, 2010. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/2/
1/41.

Fang:2019:ROX

[FHXL19] Yong Fang, Cheng Huang, Yijia Xu, and Yang Li. RLXSS: Optimizing XSS detection model to defend against adversarial attacks based on reinforcement learning. Future Internet, 11(8):177, August 14, 2019. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/11/8/177.

Feng:2024:SFC

[FHZ^{+}24] Xu Feng, Mengyang He, Lei Zhuang, Yanrui Song, and Rumeng Peng. Service function chain deployment algorithm based on deep reinforcement learning in space–air–ground integrated network. Future Internet, 16(1):27, January 16, 2024. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/16/1/27.

Fiore:2010:SRR

[Fio10]  Ugo Fiore.  Selective

redundancy removal: a framework for data hiding. Future Internet, 2(1): 30–40, February 15, 2010. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/2/
                                        1/30.

Fekete:2021:SLT

[FK21] Dénes László Fekete and Attila Kiss. A survey of ledger technology-based databases. Future Internet, 13(8):197, July 31, 2021. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/13/8/197.

Fragkou:2024:JSD

[FK24] Evangelia Fragkou and Dimitrios Katsaros. A joint survey in decentralized federated learning and TinyML: a brief introduction to swarm learning. Future Internet, 16(11): 413, November 8, 2024. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/16/
                                        11/413.

Fallatah:2024:EMU

[FKF24] Wesam Fallatah, Joakim Kävrestad, and Steven Furnell. Establishing a model for the user acceptance of cybersecurity training. Future Internet, 16(8):??, August 15, 2024. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
8/294.

Fakis:2020:NDN

[FKK20] Alexandros Fakis, Georgios Karopoulos, and Georgios Kambourakis. Neither denied nor exposed: Fixing WebRTC privacy leaks. Future Internet, 12(5):92, May 22, 2020. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/12/5/92.

Fan:2024:BUB

[FKM24] Caixiang Fan, Hamzeh Khazaei, and Petr Musilek. BPET: a unified blockchainbased framework for peerto-peer energy trading. Future Internet, 16(5):162, May 7, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/5/162.

Feng:2022:EEB

[FKZ22] Jinghui Feng, Haopeng Kuang, and Lihua Zhang. EBBA: an enhanced binary bat algorithm integrated with chaos theory and Lévy flight for feature selection. Future Internet, 14(6):178, June 09, 2022. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
6/178.

Fan:2021:DIA

[FLC21] Li Fan, Wei Li, and Xiaohui Cui. Deepfake-image anti-forensics with adversarial examples attacks. Future Internet, 13(11): 288, November 17, 2021. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
11/288.

Fan:2024:IPS

[FLLC24] Zhengyang Fan, Wanru Li, Kathryn Blackmond Laskey, and Kuo-Chu Chang. Investigation of phishing susceptibility with explainable artificial intelligence. Future Internet, 16(1):31, January 17, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/1/31.

Fang:2024:LMC

[FLLG24] Cheng Fang, Feifei Liang, Tianchi Li, and Fangheng Guan. Learning modality consistency and difference information with multitask learning for multimodal sentiment analysis. Future Internet, 16(6):

213, June 17, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
6/213.

Fouzar:2023:SVC

[FLM23] Youcef Fouzar, Ahmed Lakhssassi, and Ramakrishna Mundugar. Secure video communication using multi-equation multikey hybrid cryptography. Future Internet, 15(12): 387, November 29, 2023. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
12/387.

Fehling:2012:PBD

[FLRS12] Christoph Fehling, Frank Leymann, Jochen Rütschlin, and David Schumm. Patternbased development and management of cloud applications. Future Internet, 4(1):110–141, February 15, 2012. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/4/1/110.

Fu:2024:DTB

[FLSL24] Tianjie Fu, Peiyu Li, Chenke Shi, and Youzhu Liu. Digital-twin-based monitoring system for slab production process. Future Internet, 16(2):59, February 13, 2024. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/16/2/59.

Fantacci:2016:CSS

[FM16] Romano Fantacci and Dania Marabissi. Cognitive spectrum sharing: An enabling wireless communication technology for a wide use of smart systems. Future Internet, 8(2):23, May 20, 2016. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/8/2/23.

Florea:2020:MDL

[FM20a] George Albert Florea and Radu-Casian Mihailescu. Multimodal deep learning for group activity recognition in smart office environments. Future Internet, 12(8):133, August 09, 2020. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/12/8/133.

Foubert:2020:LRW

[FM20b] Brandon Foubert and Nathalie Mitton. Longrange wireless radio technologies: a survey. Future Internet, 12(1):13, January 14, 2020. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/12/1/13.

Faraj:2021:RVC

[FM21] Ghazal Faraj and András Micsik. Representing and validating cultural heritage knowledge graphs in CIDOC-CRM ontology.

Future Internet, 13(11): 277, October 29, 2021. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
11/277.

Ferrag:2020:RRD

[FMA^{+}20] Mohamed Amine Ferrag, Leandros Maglaras, Ahmed Ahmim, Makhlouf Derdour, and Helge Janicke. RDTIDS: Rules and decision tree-based intrusion detection system for Internet-of-Things networks. Future Internet, 12(3):44, March 02, 2020. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/12/
3/44.

Francia:2023:DMC

[FMA^{+}23] Alberto Francia, Stefano Mariani, Giuseppe Adduce, Sandro Vecchiarelli, and Franco Zambonelli. Digital management of competencies in Web 3.0: The C-Box approach. Future Internet, 15(11):350, October 26, 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
11/350.

Fiaidhi:2009:IMM

[FMCT09] Jinan Fiaidhi, Sabah Mohammed, Lyle Chamarette, and David Thomas. Identifying middlewares for mashup personal learning environments. Future Internet, 1(1):14–27, August 05, 2009. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/1/1/14.

Fraga:2022:FSD

[FML^{+}22] Martin Fraga, Matías Micheletto, Andrés Llinás, Rodrigo Santos, and Paula Zabala. Flow scheduling in data center networks with time and energy constraints: a softwaredefined network approach. Future Internet, 14(2):65, February 21, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
2/65.

Foschini:2021:SEA

[FMMS21] Luca Foschini, Valentina Mignardi, Rebecca Montanari, and Domenico Scotece. An SDN-enabled architecture for IT/OT converged networks: a proposal and qualitative analysis under DDoS attacks. Future Internet, 13(10):258, October 08, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/10/258.

Formosa:2011:SIS

[FMNS11] Saviour Formosa, Vincent Magri, Julia Neuschmid, and Manfred Schrenk.

Sharing integrated spatial and thematic data: The CRISOLA case for Malta and the European Project Plan4all process. Future Internet, 3(4):344– 361, December 20, 2011. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/3/
4/344.

Flamini:2023:OPR

[FN23] Marta Flamini and Maurizio Naldi. Optimal pricing in a rented 5G infrastructure scenario with sticky customers. Future Internet, 15(2):82, February 19, 2023. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/15/2/82.

Ferretti:2021:HSI

[FNN21] Marco Ferretti, Serena Nicolazzo, and Antonino Nocera. H2O: Secure interactions in IoT via behavioral fingerprinting. Future Internet, 13(5):117, April 30, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/5/117.

Fahim:2019:ANI

[FOJE19] Mohamed Fahim, Brahim Ouchao, Abdeslam Jakimi, and Lahcen El Bermi. Application of a nonimmersive VR, IoT based approach to help Moroccan students carry out practical activities in a personal learning style. Future Internet, 11(1):11, January 04, 2019. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/11/
                                        1/11.

Formosa:2014:NPR

[For14] Saviour Formosa. Neo-Geography and preparedness for real-to-virtual world knowledge transfer: Conceptual steps to Minecraft Malta. Future Internet, 6(3):542– 555, August 28, 2014. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/6/
                                        3/542.

Fotiou:2020:ICN

[Fot20] Nikos Fotiou. Informationcentric networking (ICN). Future Internet, 12(2):35, February 13, 2020. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/12/
                                        2/35.

Fotiou:2019:NBS

[FP19] Nikos Fotiou and George C. Polyzos. Name-based security for information-centric networking architectures. Future Internet, 11(11): 232, November 01, 2019. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/11/
                                        11/232.

Fiorini:2024:EID

[FPGO24] Francesco Fiorini, Michele Pagano, Rosario Giuseppe Garroppo, and Antonio Osele. Estimating interception density in the BB84 protocol: a study with a noisy quantum simulator. Future Internet, 16(8):??, August 2, 2024. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
8/275.

Fragkou:2022:MBI

[FPKK22] Evangelia Fragkou, Dimitrios Papakostas, Theodoros Kasidakis, and Dimitrios Katsaros. Multilayer backbones for Internet of Battlefield Things. Future Internet, 14(6):186, June 15, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/6/186.

Formica:2020:SSE

[FPT20] Anna Formica, Elaheh Pourabbas, and Francesco Taglino. Semantic search enhanced with rating scores. Future Internet, 12 (4):67, April 15, 2020. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/12/
4/67.

Fang:2021:PPM

[FQ21] Haokun Fang and Quan Qian. Privacy preserving machine learning with homomorphic encryption and federated learning. Future Internet, 13(4):94, April 08, 2021. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/13/4/94.

Fiore:2015:EIB

[FR15] Ugo Fiore and Francesco Rossi. Embedding an identity-based short signature as a digital watermark. Future Internet, 7(4):393–404, October 23, 2015. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/7/4/393.

Francese:2016:SEP

[FR16] Rita Francese and Michele Risi. Supporting elderly people by ad hoc generated mobile applications based on vocal interaction. Future Internet, 8(3):42, August 25, 2016. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/8/3/42.

Femminella:2024:IIT

[FR24] Mauro Femminella and Gianluca Reali. Implementing Internet of Things service platforms with network function virtualization serverless technologies. Future Internet, 16 (3):91, March 08, 2024. CODEN ???? ISSN 1999-

5903. URL https://www.

mdpi.com/1999-5903/16/
3/91.

Francesconi:2018:FLP

[Fra18] Enrico Francesconi. On the future of legal publishing services in the Semantic Web. Future Internet, 10(6):48, June 05, 2018. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/10/
6/48.

Frattolillo:2023:WPS

[Fra23] Franco Frattolillo. Watermarking protocols: a short guide for beginners. Future Internet, 15(5):163, April 28, 2023. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/15/5/163.

Frattolillo:2024:BSC

[Fra24] Franco Frattolillo. Blockchain and smart contracts for digital copyright protection. Future Internet, 16 (5):169, May 14, 2024. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
5/169.

Fiorentino:2022:MSS

[FRE^{+}22] Michele Fiorentino, Marina Ricci, Alessandro Evangelista, Vito Modesto Manghisi, and Antonio Emmanuele Uva. A multi-sensory in-store virtual reality customer journey for retailing: a field study in a furniture flagship store. Future Internet, 14(12):381, December 16, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/12/381.

Fritsch:2013:CPE

[Fri13] Lothar Fritsch. The clean privacy ecosystem of the future Internet. Future Internet, 5(1):34–45, January 14, 2013. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/5/1/34.

Fedotova:2022:AAS

[FRKS22] Anastasia Fedotova, Aleksandr Romanov, Anna Kurtukova, and Alexander Shelupanov. Authorship attribution of social media and literary Russianlanguage texts using machine learning methods and feature selection. Future Internet, 14(1):4, December 22, 2022. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/14/1/4.

Fatima:2024:TEF

[FRR^{+}24] Mahawish Fatima, Osama Rehman, Ibrahim M. H. Rahman, Aisha Ajmal, and Simon Jigwan Park. Towards ensemble feature selection for lightweight intrusion detection in resource-constrained IoT devices. Future Internet, 16(10), October 12, 2024. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
10/368.

Ferreira:2023:OPT

[FSC^{+}23] Eduardo Ferreira, Pedro Sebastião, Francisco Cercas, Carlos Sá Costa, and Américo Correia. An optimized planning tool for microwave terrestrial and satellite link design. Future Internet, 15(2):58, January 31, 2023. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/15/2/58.

Florez:2016:ADC

[FSG^{+}16] Angélica Flórez, Lenin Serrano, Urbano Gómez, Luis Suárez, Alejandro Villarraga, and Hugo Rodríguez. Analysis of dynamic complexity of the cyber security ecosystem of Colombia. Future Internet, 8(3):33, July 19, 2016. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/8/
3/33.

Ferres:2022:PDE

[FSG22] Kim Ferres, Timo Schloesser, and Peter A. Gloor. Predicting dog emotions based on posture analysis using DeepLabCut. Future Internet, 14(4):97–??, March 22, 2022. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/14/4/97.

Fiaz:2024:MFP

[FSI^{+}24] Faisal Fiaz, Syed Muhammad Sajjad, Zafar Iqbal, Muhammad Yousaf, and Zia Muhammad. MetaSSI: a framework for personal data protection, enhanced cybersecurity and privacy in metaverse virtual reality platforms. Future Internet, 16(5):176, May 18, 2024. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/16/
                                        5/176.

Fotiou:2022:IGM

[FSXP22] Nikos Fotiou, Vasilios A. Siris, George Xylomenos, and George C. Polyzos. IoT group membership management using decentralized identifiers and verifiable credentials. Future Internet, 14(6):173, June 01, 2022. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/14/6/173.

Fan:2022:IOM

[FSY^{+}22] Guangwei Fan, Chuanzhen Sheng, Baoguo Yu, Lu Huang, and Qiang Rong. An indoor and outdoor multisource elastic fusion navigation and positioning algorithm based on particle filters. Future Internet, 14(6):169, May 31, 2022. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
6/169.

Fendji:2021:WST

[FTAA21] Jean Louis Ebongue Kedieng Fendji, Désiré Manuel Taira, Marcellin Atemkeng, and Adam Musa Ali. WATS-SMS: a T5-based French Wikipedia abstractive text summarizer for SMS. Future Internet, 13(9):238, September 18, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/9/238.

Flanagan:2021:IMC

[FTHY21] Steven Knowles Flanagan, Zuoyin Tang, Jianhua He, and Irfan Yusoff. Investigating and modeling of cooperative vehicle-to-vehicle safety stopping distance. Future Internet, 13 (3):68, March 10, 2021. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
3/68.

FIEO:2020:ARF

[Fut20] Future Internet Editorial Office. Acknowledgement to reviewers of Future Internet in 2019. Future Internet, 12(1):18, January 20, 2020. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/12/1/18.

FIEO:2021:ARF

[Fut21] Future Internet Editorial Office. Acknowledgment to reviewers of Future Internet in 2020. Future Internet, 13(2):28, January 24, 2021. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/13/2/28.

Feng:2022:EBT

[FWL22] Hangwei Feng, Jinlin Wang, and Yang Li. An efficient blockchain transaction retrieval system. Future Internet, 14(9):267, September 15, 2022. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/14/
                                        9/267.

Falayi:2023:SDD

[FWLY23] Ayodeji Falayi, Qianlong Wang, Weixian Liao, and Wei Yu. Survey of distributed and decentralized IoT securities: Approaches using deep learning and blockchain technology. Future Internet, 15(5):178, May 11, 2023. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/15/
                                        5/178.

Forte:2019:NGC

[FWVF19] Andrea G. Forte, Wei Wang, Luca Veltri, and Gianluigi Ferrari. A nextgeneration core network architecture for mobile networks. Future Internet, 11(7):152, July 09, 2019. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
7/152.

Fernandez:2016:MSC

[FYWS16] Eduardo B. Fernandez, Nobukazu Yoshioka, Hironori Washizaki, and Madiha H. Syed. Modeling and security in cloud ecosystems. Future Internet, 8(2):13, April 20, 2016. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/8/2/13.

Floris:2015:SMR

[FZ15] Roberta Floris and Corrado Zoppi. Social mediarelated geographic information in the context of strategic environmental assessment of municipal masterplans: a case study concerning Sardinia (Italy). Future Internet, 7(3):276–293, August 07, 2015. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/7/3/276.

Gupta:2021:IBP

[GAA21] Akash Gupta and Adnan Al-Anbuky. IoT-based patient movement monitoring: The post-operative hip fracture rehabilitation model. Future Internet, 13(8):195, July 29, 2021. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/13/
                                        8/195.

Galanis:2019:EBS

[GAGB19] Ioannis Galanis, Iraklis Anagnostopoulos, Priyaa Gurunathan, and Dona Burkard. Environmentalbased speed recommendation for future smart cars. Future Internet, 11(3):78, March 24, 2019. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/11/3/78.

Galis:2024:FSI

[Gal24] Alex Galis. Future sustainable Internet energydefined networking. Future Internet, 16(1):23, January 09, 2024. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/16/1/23.

Gonzalez-Ambriz:2024:SGB

[GAMMPCRÁ24] Sergio Jesús González-Ambriz, Rolando Menchaca-Méndez, Sergio Alejandro Pinacho-Castellanos, and Mario Eduardo Rivero-Ángeles. A spectral gapbased topology control algorithm for wireless backhaul networks. Future Internet, 16(2):43, January 26, 2024. CODEN ????

ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/2/43.

Goodell:2021:DCA

[GANT21] Geoffrey Goodell, Hazem Danny Al-Nakib, and Paolo Tasca. A digital currency architecture for privacy and owner-custodianship. Future Internet, 13(5): 130, May 14, 2021. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
5/130.

Gao:2021:IQE

[Gao21] Hui-Li Gao. The impact of quality of experience of Chinese college students on Internet-based resources English learning. Future Internet, 13(7): 162, June 22, 2021. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
7/162.

Garau:2012:FCP

[Gar12] Chiara Garau. Focus on citizens: Public engagement with online and faceto-face participation — a case study. Future Internet, 4(2):592–606, June 15, 2012. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/4/2/592.

Giusfredi:2020:CAC

[GB20] Michele Giusfredi and Franco Bagnoli. From color-avoiding to colorfavored percolation in diluted lattices. Future Internet, 12(8):139, August 18, 2020. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/12/8/139.

Grimes:2023:MAA

[GB23] Sean Grimes and David E. Breen. A multi-agent approach to binary classification using swarm intelligence. Future Internet, 15 (1):36, January 12, 2023. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/15/
                                        1/36.

GomezRomero-Borquez:2023:MEA

[GBDDVS23] Jesus GomezRomero-Borquez, J. Alberto Del Puerto-Flores, and Carolina Del-Valle-Soto. Mapping EEG alpha activity: Assessing concentration levels during player experience in virtual reality video games. Future Internet, 15(8):264, August 09, 2023. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/15/
                                        8/264.

Gkoulis:2023:TIE

[GBKN23] Dimitris Gkoulis, Cleopatra Bardaki, George Kousiouris, and Mara Nikolaidou. Transforming IoT events to meaningful business events on the edge: Implementation for smart farming application. Future Internet, 15(4):135, March 31, 2023. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/15/4/135.

Gloor:2022:YFM

[GCA^{+}22] Peter A. Gloor, Andrea Fronzetti Colladon, Erkin Altuntas, Cengiz Cetinkaya, Maximilian F. Kaiser, Lukas Ripperger, and Tim Schaefer. Your face mirrors your deepest beliefs-predicting personality and morals through facial emotion recognition. Future Internet, 14(1):5, December 22, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
1/5.

Gomez:2023:PPF

[GCEOBRT23] Javier Gomez, Jose Jaime Camacho-Escoto, Luis Orozco-Barbosa, and Diego Rodriguez-Torres. PRO-FEE: a probabilisticfeedback based speed rate adaption for IEEE 802.11bc. Future Internet, 15(12):396, December 09, 2023. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/15/12/396.

Giddings:2018:RDB

[GDARM18] Roger Giddings, Xiao Duan, Ehab Al-Rawachy, and Mingzhi Mao. A review of DSP-based enabling technologies for cloud access networks. Future Internet, 10(11):109, November 15, 2018. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/10/
                                        11/109.

Guazzini:2019:EMA

[GDCM19] Andrea Guazzini, Mirko Duradoni, Ambra Capelli, and Patrizia Meringolo. An explorative model to assess individuals’ phubbing risk. Future Internet, 11(1):21, January 18, 2019. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/11/1/21.

Guazzini:2018:SCC

[GDLG18] Andrea Guazzini, Mirko Duradoni, Alessandro Lazzeri, and Giorgio Gronchi. Simulating the cost of cooperation: a recipe for collaborative problem-solving. Future Internet, 10(6):55, June 19, 2018. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/10/6/55.

Gebreyesus:2023:MLD

[GDN^{+}23] Yibrah Gebreyesus, Damian Dalton, Sebastian Nixon, Davide De Chiara, and Marta Chinnici. Machine learning for data center optimizations: Feature selection using Shapley Additive exPlanation (SHAP). Future Internet, 15(3):88, February 21, 2023. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/15/3/88.

Garcia-Delgado:2023:ETT

[GDRCDBLÁ23] Miguel Ángel García-Delgado, Sonia Rodríguez-Cano, Vanesa Delgado-Benito, and María Lozano-Álvarez. Emerging technologies and their link to digital competence in teaching. Future Internet, 15(4):140, April 05, 2023. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
4/140.

Gorchakov:2023:APR

[GDS23] Artyom V. Gorchakov, Liliya A. Demidova, and Peter N. Sovietov. Analysis of program representations based on abstract syntax trees and higherorder Markov chains for source code classification task. Future Internet, 15(9):314, September 18, 2023. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/15/9/314.

Goncalves:2023:IGC

[GdSdC23] Ricardo Martins Gonçalves, Miguel Mira da Silva, and Paulo Rupino da Cunha. Implementing GDPR-compliant surveys using blockchain. Future Internet, 15(4):143, April 07, 2023. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/15/
                                        4/143.

George:2012:TMB

[GFHK^{+}12] Mathew George, Klaus-Peter Fischer-Hellmann, Martin Knahl, Udo Bleimann, and Shirley Atkinson. Traceability in modelbased testing. Future Internet, 4(4):1026–1036, November 26, 2012. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/4/
                                        4/1026.

Guo:2021:CSD

[GFL21] Dongwei Guo, Mengmeng Fu, and Hai Li. Cooperation in social dilemmas: a group game model with double-layer networks. Future Internet, 13(2):33, January 27, 2021. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/13/
                                        2/33.

Guazzini:2021:WWW

[GFPD21] Andrea Guazzini, Maria Fiorenza, Gabriele Panerai, and Mirko Duradoni. What went wrong? predictors of contact tracing adoption in Italy during COVID-19 pandemic. Future Internet, 13(11):286, November 15, 2021. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
11/286.

Gonzalez-Franco:2023:CSL

[GFPVLR^{+}23] Joan D. Gonzalez-Franco, Jorge E. Preciado-Velasco, Jose E. Lozano-Rizk, Raul Rivera-Rodriguez, Jorge Torres-Rodriguez, and Miguel A. Alonso-Arevalo. Comparison of supervised learning algorithms on a 5G dataset reduced via principal component analysis (PCA). Future Internet, 15(10):335, October 11, 2023. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/15/10/335.

Guazzini:2020:CFC

[GGCY20] Andrea Guazzini, Elisa Guidi, Cristina Cecchini, and Eiko Yoneki. Collaborative facilitation and collaborative inhibition in virtual environments. Future Internet, 12(7):118, July 13, 2020. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/12/
7/118.

Ghosh:2023:IWS

[GGD23] Swarnendu Ghosh, Teresa Gonçalves, and Nibaran Das. Im2Graph: a weakly supervised approach for generating holistic scene graphs from regional dependencies. Future Internet, 15(2):70, February 10, 2023. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/15/2/70.

Gunleifsen:2018:TCP

[GGK18] Håkon Gunleifsen, Vasileios Gkioulos, and Thomas Kemmerich. A tiered control plane model for service function chaining isolation. Future Internet, 10 (6):46, June 04, 2018. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/10/
                                        6/46.

Gkonis:2023:SIE

[GGT^{+}23] Panagiotis Gkonis, Anastasios Giannopoulos, Panagiotis Trakadas, Xavi Masip-Bruin, and Francesco D’Andria. A survey on IoT-edge-cloud continuum systems: Status, challenges, use cases, and open issues. Future Internet, 15(12):383, November 28, 2023. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/15/12/383.

Gkioulos:2018:SLR

[GGW18] Vasileios Gkioulos, Håkon Gunleifsen, and Goitom Kahsay Weldehawaryat. A systematic literature review on military software defined networks. Future Internet, 10(9):88, September 12, 2018. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/10/9/88.

Ghazal:2021:ISC

[GHA^{+}21] Taher M. Ghazal, Mohammad Kamrul Hasan, Muhammad Turki Alshurideh, Haitham M. Alzoubi, Munir Ahmad, Syed Shehryar Akbar, Barween Al Kurdi, and Iman A. Akour. IoT for smart cities: Machine learning approaches in smart healthcare — a review. Future Internet, 13 (8):218, August 23, 2021. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
8/218.

Gerlach:2024:GXM

[GHFL24] Bennet Gerlach, Hannes Hesse, Stefan Fischer, and Martin Leucker. Gaia-X-Med: Experiences with building dataspaces for medical applications. Future Internet, 16(12):463, December 9, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.

                                        mdpi.com/1999-5903/16/
                                        12/463.

Ghaddar:2020:MAC

[GHFM20] Alia Ghaddar, Monah Bou Hatoum, Ghassan Fadlallah, and Hamid Mcheick. MCCM: An approach for connectivity and coverage maximization. Future Internet, 12(2):19, January 21, 2020. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/12/2/19.

Goudarzi:2022:HUS

[GHGL22] Pejman Goudarzi, Mehdi Hosseinpour, Roham Goudarzi, and Jaime Lloret. Holistic utility satisfaction in cloud data centre network using reinforcement learning. Future Internet, 14(12): 368, December 08, 2022. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/14/
                                        12/368.

Guan:2023:SBB

[GHL^{+}23] Yuting Guan, Junjiang He, Tao Li, Hui Zhao, and Baoqiang Ma. SSQLi: a black-box adversarial attack method for SQL injection based on reinforcement learning. Future Internet, 15(4):133, March 30, 2023. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/15/4/133.

Gao:2023:RDD

[GHZY23] Han Gao, Zhangqin Huang, Xiaobo Zhang, and Huapeng Yang. Research and design of a decentralized edgecomputing-assisted LoRa gateway. Future Internet, 15(6):194–??, June 2023. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
6/194.

Giuli:2018:AFI

[Giu18a] Dino Giuli. Announcing the 2018 Future Internet travel award for PhD students. Future Internet, 10(2):11, January 28, 2018. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/10/2/11.

Giuli:2018:EEF

[Giu18b] Dino Giuli. Ecosystemic evolution Fed by smart systems. Future Internet, 10(3):28, March 10, 2018. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/10/
3/28.

Gan:2015:SNP

[GJ15] Diane Gan and Lily R.

Jenkins. Social networking privacy — who’s stalking you? Future Internet, 7 (1):67–93, March 23, 2015. CODEN ???? ISSN 1999-

5903. URL https://www.

mdpi.com/1999-5903/7/
1/67.

Goudosis:2021:AOR

[GK21] Athanasios Goudosis and Sokratis Katsikas. ARIBC: Online reporting based on identity-based cryptography. Future Internet, 13 (2):53, February 21, 2021. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
2/53.

Gkontzis:2024:EUR

[GKFV24] Andreas F. Gkontzis, Sotiris Kotsiantis, Georgios Feretzakis, and Vassilios S. Verykios. Enhancing urban resilience: Smart city data analyses, forecasts, and digital twin techniques at the neighborhood level. Future Internet, 16(2):47, January 30, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/2/47.

Gunleifsen:2019:PCD

[GKG19] Håkon Gunleifsen, Thomas Kemmerich, and Vasileios Gkioulos. A proof-of-concept demonstration of isolated and encrypted service function chains. Future Internet, 11(9):183, August 24, 2019. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
9/183.

Giannakoulopoulos:2019:AEW

[GKK^{+}19] Andreas Giannakoulopoulos, Nikos Konstantinou, Dimitris Koutsompolis, Minas Pergantis, and Iraklis Varlamis. Academic excellence, Website quality, SEO performance: Is there a correlation? Future Internet, 11(11):242, November 18, 2019. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
11/242.

Garg:2024:IIM

[GKM24] Umang Garg, Santosh Kumar, and Aniket Mahanti. IMTIBOT: an intelligent mitigation technique for IoT botnets. Future Internet, 16(6):212, June 17, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/6/212.

Gottron:2010:SSM

[GKS10] Christian Gottron, André König, and Ralf Steinmetz. A survey on security in mobile peer-to-peer architectures-overlay-based vs. underlay-based approaches. Future Internet, 2(4):505–532, October 13, 2010. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/2/4/505.

Guthle:2010:IAD

[GKW^{+}10] Martin Güthle, Jochen Kögel, Stefan Wahl, Matthias Kaschub, and Christian M. Mueller. Improving anomaly detection for text-based protocols by exploiting message structures. Future Internet, 2(4):662–669, December 21, 2010. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/2/4/662.

Groza:2012:PDL

[GL12] Adrian Groza and Ioan Alfred Letia. Plausible description logic programs for stream reasoning. Future Internet, 4(4):865– 881, October 17, 2012. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/4/
                                        4/865.

Gelenbe:2013:EQT

[GL13] Erol Gelenbe and Ricardo Lent. Energy-QoS tradeoffs in mobile service selection. Future Internet, 5(2): 128–139, April 19, 2013. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/5/
                                        2/128.

Geyda:2019:MIO

[GL19] Alexander Geyda and Igor Lysenko. Modeling of information operations effects: Technological systems example. Future Internet, 11(3):62, March 05, 2019. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/11/3/62.

Georgiou:2021:DPI

[GL21] Dimitra Georgiou and Costas Lambrinoudakis. Data protection impact assessment (DPIA) for cloud-based health organizations. Future Internet, 13(3):66, March 07, 2021. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
3/66.

Glassey:2011:MIM

[Gla11] Olivier Glassey. Metadata for identity management of population registers. Future Internet, 3(2): 130–143, April 18, 2011. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/3/
2/130.

Gatteschi:2018:BSC

[GLD^{+}18] Valentina Gatteschi, Fabrizio Lamberti, Claudio Demartini, Chiara Pranteda, and Víctor Santamaría. Blockchain and smart contracts for insurance: Is the technology mature enough? Future Internet, 10(2):20, February 20, 2018. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/10/2/20.

Guia:2022:CSM

[GLH^{+}22] Sana Sahar Guia, Abdelkader Laouid, Mohammad Hammoudeh, Ahcène Bounceur, Mai Alfawair, and Amna Eleyan. Cosimulation of multiple vehicle routing problem models. Future Internet, 14(5):137, April 29, 2022. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
5/137.

Garcia:2019:DMA

[GLRM19] Juan José López García, David Lizcano, Celia M. Q. Ramos, and Nelson Matos. Digital marketing actions that achieve a better attraction and loyalty of users: An analytical study. Future Internet, 11(6):130, June 08, 2019. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
6/130.

Giacoumidis:2019:HML

[GLW^{+}19] Elias Giacoumidis, Yi Lin, Jinlong Wei, Ivan Aldaya, Athanasios Tsokanos, and Liam P. Barry. Harnessing machine learning for fiberinduced nonlinearity mitigation in long-haul coherent optical OFDM. Future Internet, 11(1):2, December 20, 2019. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/11/1/2.

Gao:2024:SGS

[GLZL24] Shaopei Gao, Qiang Liu, Junjie Zeng, and Li Li. SD-GPSR: a software-defined greedy perimeter stateless routing method based on geographic location information. Future Internet, 16(7):??, July 17, 2024. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
7/251.

Golbeck:2016:UPF

[GM16] Jennifer Golbeck and Matthew Louis Mauriello. User perception of Facebook app data access: a comparison of methods and privacy concerns. Future Internet, 8(2):9, March 25, 2016. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/8/2/9.

Gunawan:2019:DAH

[GM19] Dedi Gunawan and Masahiro Mambo. Data anonymization for hiding personal tendency in set-valued database publication. Future Internet, 11(6):138, June 20, 2019. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
6/138.

Ghaddar:2020:EEA

[GM20] Alia Ghaddar and Ahmad Merei. EAOA: Energy-aware grid-based 3d-obstacle avoidance in coverage path planning for UAVs. Future Internet, 12 (2):29, February 08, 2020. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/12/
                                        2/29.

Guidi:2023:NNR

[GM23] Barbara Guidi and Andrea Michienzi. From NFT 1.0 to NFT 2.0: a review of the evolution of non-fungible tokens. Future Internet, 15(6):189– ??, June 2023. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/15/
                                        6/189.

Ghourabi:2020:HCL

[GMA20] Abdallah Ghourabi, Mahmood A. Mahmood, and Qusay M. Alzubi. A hybrid CNN-LSTM model for SMS spam detection in Arabic and English messages. Future Internet, 12(9):156, September 18, 2020. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/12/9/156.

Guarino:2021:IUS

[GMB^{+}21] Stefano Guarino, Enrico Mastrostefano, Massimo Bernaschi, Alessandro Celestini, Marco Cianfriglia, Davide Torre, and Lena Rebecca Zastrow. Inferring urban social networks from publicly available data. Future Internet, 13(5):108, April 26, 2021. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
5/108.

Gentile:2022:VPA

[GMD^{+}22] Antonio Francesco Gentile, Davide Macrì, Floriano De Rango, Mauro Tropea, and Emilio Greco. A VPN performances analysis of constrained hardware open source infrastructure deploy in IoT environment. Future Internet, 14(9):264, September 13, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/9/264.

Gentile:2024:OVP

[GMGF24] Antonio Francesco Gentile, Davide Macrì, Emilio Greco, and Peppino Fazio. Overlay and virtual private networks security performances analysis with open source infrastructure deployment. Future Internet, 16(8):??, August 7, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/8/283.

Guo:2020:MWA

[GMGK20] You Guo, Hector Marco-Gisbert, and Paul Keir. Mitigating webshell attacks through machine learning techniques. Future Internet, 12(1):12, January 14, 2020. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/12/
                                        1/12.

Goeke:2018:CAC

[GMH18] Ludger Goeke, Nazila Gol Mohammadi, and Maritta Heisel. Context analysis of cloud computing systems using a pattern-based approach. Future Internet, 10(8):72, July 31, 2018. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/10/
                                        8/72.

Giacco:2022:RGI

[GMLS22] Giovanni Giacco, Stefano Marrone, Giuliano Langella, and Carlo Sansone. ReFuse: Generating imperviousness maps from multi-spectral Sentinel-2 satellite imagery. Future Internet, 14(10):278, September 28, 2022. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/14/
                                        10/278.

Guidi:2016:CSI

[GMMI16] Gabriele Guidi, Roberto Miniati, Matteo Mazzola, and Ernesto Iadanza. Case study: IBM Watson analytics cloud platform as analytics-as-a-service system for heart failure early detection. Future Internet, 8(3):32, July 13, 2016. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/8/
3/32.

Garcia-Martinez:2016:MAE

[GMRRRQ^{+}16] Antón García-Martínez, José Rivas-Rangel, Jaime Rangel-Quintos, José Antonio Espinosa, Cecilio Barba, and Carmen De-Pablos-Heredero. A methodological approach to evaluate livestock innovations on small-scale farms in developing countries. Future Internet, 8 (2):25, June 03, 2016. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/8/
2/25.

Gupta:2020:ABS

[GMS^{+}20] Nishu Gupta, Ravikanti Manaswini, Bongaram Saikrishna, Francisco Silva, and Ariel Teles. Authenticationbased secure data dissemination protocol and framework for 5G-enabled VANET. Future Internet, 12(4):63, April 01, 2020. CODEN ???? ISSN 1999-

5903. URL https://www.

                                        mdpi.com/1999-5903/12/
                                        4/63.

Golbeck:2016:EFA

[GN16] Jennifer Golbeck and Carman Neustaedter. Environmental factors affecting where people geocache. Future Internet, 8(2):11, April 12, 2016. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/8/2/11.

Gonzalez-Nieto:2021:SCD

[GNGHEM21] Noé Abraham González-Nieto, Caridad García-Hernández, and Margarita Espinosa-Meneses. School culture and digital technologies: Educational practices at universities within the context of the COVID-19 pandemic. Future Internet, 13(10): 246, September 24, 2021. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/13/
                                        10/246.

Guo:2023:SSR

[GNGL23] Yu Guo, Guigen Nie, Wenliang Gao, and Mi Liao. 2D semantic segmentation: Recent developments and future directions. Future Internet, 15(6):205– ??, June 2023. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/15/
                                        6/205.

Graham:2010:BMM

[GNK^{+}10] Jim Graham, Greg Newman, Sunil Kumar, Catherine Jarnevich, Nick Young, Alycia Crall, Thomas J. Stohlgren, and Paul Evangelista. Bringing modeling to the masses: a Web based system to predict potential species distributions. Future Internet, 2(4):624–634, November 11, 2010. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/2/4/624.

Guerrero-Narvaez:2018:TBS

[GNNZRCRG18] Santiago Guerrero-Narváez, Miguel-Ángel Niño-Zambrano, Dalila-Jhoana Riobamba-Calvache, and Gustavo-Adolfo Ramírez-González. Test bed of semantic interaction of smart objects in the Web of things. Future Internet, 10(5):42, May 09, 2018. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/10/5/42.

Gavra:2021:DIA

[GNV21] Dmitrii Gavra, Ksenia Namyatova, and Lidia Vitkova. Detection of induced activity in social networks: Model and methodology. Future Internet, 13(11):297, November 22, 2021. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/13/11/297.

Goetz:2012:UCI

[Goe12]  Marcus Goetz.  Using

crowdsourced indoor geodata for the creation of a three-dimensional indoor routing Web application. Future Internet, 4(2):575– 591, June 06, 2012. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/4/
                                        2/575.

Goggin:2012:DIM

[Gog12] Gerard Goggin. Driving the Internet: Mobile Internets, cars, and the social. Future Internet, 4(1): 306–321, March 20, 2012. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/4/
                                        1/306.

Gratsos:2023:KAD

[GOM23] Konstantinos Gratsos, Stefanos Ougiaroglou, and Dionisis Margaris. kClusterHub: an AutoMLdriven tool for effortless partition-based clustering over varied data types. Future Internet, 15(10):341, October 18, 2023. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/15/
                                        10/341.

Ghanem:2023:AST

[GPA23] Fahd A. Ghanem, M. C.

Padma, and Ramez Alkhatib. Automatic short text summarization techniques in social media platforms. Future Internet, 15(9):311, September 13, 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
9/311.

Garside:2012:SMS

[GPCP12] Debbie Garside, Arjun Ponnusamy, Steve Chan, and Richard Picking. Secure military social networking and rapid sensemaking in domain specific concept systems: Research issues and future solutions. Future Internet, 4(1):253– 264, March 12, 2012. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/4/
1/253.

Giannakoulopoulos:2020:EDE

[GPK^{+}20] Andreas Giannakoulopoulos, Minas Pergantis, Nikos Konstantinou, Aristeidis Lamprogeorgos, Laida Limniati, and Iraklis Varlamis. Exploring the dominance of the English language on the Websites of EU countries. Future Internet, 12(4):76, April 22, 2020. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/12/4/76.

Giannakoulopoulos:2024:UEF

[GPL24] Andreas Giannakoulopoulos, Minas Pergantis, and Aristeidis Lamprogeorgos. User experience, functionality and aesthetics evaluation in an academic multi-site Web ecosystem. Future Internet, 16(3):92, March 08, 2024. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/16/3/92.

Giannakoulopoulos:2022:ICO

[GPLK22] Andreas Giannakoulopoulos, Minas Pergantis, Laida Limniati, and Alexandros Kouretsis. Investigating the country of origin and the role of the .eu TLD in external trade of European Union member states. Future Internet, 14(6):174, June 04, 2022. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
6/174.

Gutierrez:2021:RTP

[GPM21] Juan Roberto López Gutiérrez, Pedro Ponce, and Arturo Molina. Real-time power electronics laboratory to strengthen distance learning engineering education on smart grids and microgrids. Future Internet, 13(9):237, September 17, 2021. CODEN ????

ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/9/237.

Garlinska:2020:MFS

[GPMO20] Magdalena Garlinska, Agnieszka Pregowska, Karol Masztalerz, and Magdalena Osial. From mirrors to free-space optical communication-historical aspects in data transmission. Future Internet, 12(11):179, October 22, 2020. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/12/11/179.

Gangwani:2021:SEI

[GPPB^{+}21] Pranav Gangwani, Alexander Perez-Pons, Tushar Bhardwaj, Himanshu Upadhyay, Santosh Joshi, and Leonel Lagos. Securing environmental IoT data using masked authentication messaging protocol in a DAG-based blockchain: IOTA tangle. Future Internet, 13(12):312, December 06, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/12/312.

Gonzalez-Perez:2021:IIR

[GPRMGP21] Laura Icela González-Pérez, María Soledad Ramírez-Montoya, and Francisco José García-Peñalvo. Improving institutional repositories through user-centered design: Indicators from a focus group. Future Internet, 13(11):282, November 02, 2021. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/13/11/282.

Ge:2022:KPL

[GR22a] Shengguo Ge and Siti Nurulain Mohd Rum. Key points’ location in infrared images of the human body based on Mscf-ResNet. Future Internet, 14(1):15, December 27, 2022. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/14/
                                        1/15.

Gwyn:2022:EGB

[GR22b] Tony Gwyn and Kaushik Roy. Examining gender bias of convolutional neural networks via facial recognition. Future Internet, 14(12):375, December 13, 2022. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/14/
                                        12/375.

Graham:2013:ISI

[Gra13] Roderick Graham. Introduction to the special issue on inequality in the digital environment. Future Internet, 5(4):580–584, November 26, 2013. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/5/4/580.

Gwyn:2021:FRU

[GRA21] Tony Gwyn, Kaushik Roy, and Mustafa Atay. Face recognition using popular deep net architectures: a brief comparative study. Future Internet, 13(7):164, June 25, 2021. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
7/164.

Gori:2023:GSM

[GRA^{+}23] Giacomo Gori, Lorenzo Rinieri, Amir Al Sadi, Andrea Melis, Franco Callegati, and Marco Prandini. GRAPH4: a security monitoring architecture based on data plane anomaly detection metrics calculated over attack graphs. Future Internet, 15(11):368, November 15, 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
11/368.

Guazzini:2021:PEA

[GRB^{+}21] Andrea Guazzini, Tommaso Raimondi, Benedetta Biagini, Franco Bagnoli, and Mirko Duradoni. Phubber’s emotional activations: The association between PANAS and phubbing behavior. Future Internet, 13(12):311, December 04, 2021. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
12/311.

Gronchi:2021:NTS

[GRG21] Giorgio Gronchi, Marco Raglianti, and Fabio Giovannelli. Network theory and switching behaviors: a user guide for analyzing electronic records databases. Future Internet, 13(9):228, August 31, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/9/228.

Gomes:2019:IPI

[GRJ^{+}19] Luis Gomes, Carlos Ramos, Aria Jozi, Bruno Serra, Lucas Paiva, and Zita Vale. IoH: a platform for the intelligence of home with a context awareness and ambient intelligence approach. Future Internet, 11(3):58, March 02, 2019. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
3/58.

Gheyas:2023:OMD

[GRM23] Iffat Gheyas, Alessandro Raschella, and Michael Mackay. Optimal meshing degree performance analysis in a mmWave FWA 5G network deployment. Future Internet, 15(6):218–??, June 2023.

CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
6/218.

Grout:2017:SAD

[Gro17] Vic Grout. A simple approach to dynamic optimisation of flexible optical networks with practical application. Future Internet, 9(2):18, May 23, 2017. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/9/
2/18.

Grobelna:2022:ITC

[Gro22] Iwona Grobelna. Internet of Things and cyber– physical systems. Future Internet, 14(11):337, November 18, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
11/337.

Guo:2018:WBS

[GRRZ18] Jinyi Guo, Wei Ren, Yi Ren, and Tianqing Zhu. A watermark-based in-situ access control model for image big data. Future Internet, 10(8):69, July 29, 2018. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/10/8/69.

Graham:2011:IDP

[GS11] Roderick Graham and Danielle Taana Smith. Internet as digital practice:

Examining differences in African American Internet usage. Future Internet, 3 (3):185–203, July 20, 2011. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/3/
3/185.

Gomes:2012:CDL

[GS12] Rui Gomes and Lígia Sousa. Contributions to the development of local egovernment 2.0. Future Internet, 4(4):882–899, October 22, 2012. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/4/4/882.

Gazzan:2023:EML

[GS23a] Mazen Gazzan and Frederick T. Sheldon. An enhanced minimax loss function technique in generative adversarial network for ransomware behavior prediction. Future Internet, 15(10):318, September 22, 2023. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/15/10/318.

Gazzan:2023:OED

[GS23b] Mazen Gazzan and Frederick T. Sheldon. Opportunities for early detection and prediction of ransomware attacks against industrial control systems. Future Internet, 15(4):144, April 07, 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
4/144.

Grigaliunas:2023:LTE

[GSB^{+}23] Sarunas Grigaliunas, Michael Schmidt, Rasa Bruzgiene, Panayiota Smyrli, and Vladislav Bidikov. Leveraging taxonomical engineering for security baseline compliance in international regulatory frameworks. Future Internet, 15(10):330, October 07, 2023. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/15/10/330.

Guazzini:2017:PPP

[GSD^{+}17] Andrea Guazzini, Ayça Saraç, Camillo Donati, Annalisa Nardi, Daniele Vilone, and Patrizia Meringolo. Participation and privacy perception in virtual environments: The role of sense of community, culture and gender between Italian and Turkish. Future Internet, 9(2):11, April 07, 2017. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/9/2/11.

Gameiro:2020:NFI

[GSL20a] Luís Gameiro, Carlos Senna, and Miguel Luís. ndnIoT-FC: IoT devices as first-class traffic in name data networks. Future Internet, 12(11):207, November 21, 2020. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/12/11/207.

Goncalves:2020:IBF

[GSL20b] Rosiberto Gonçalves, Jesse J. M. Soares, and Ricardo M. F. Lima. An IoT-based framework for smart water supply systems management. Future Internet, 12(7):114, July 07, 2020. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/12/
7/114.

Gameiro:2022:IEN

[GSL22] Luís Gameiro, Carlos Senna, and Miguel Luís. Insights from the experimentation of named data networks in mobile wireless environments. Future Internet, 14(7):196, June 27, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/7/196.

Granitzer:2010:OAS

[GSO^{+}10] Michael Granitzer, Vedran Sabol, Kow Weng Onn, Dickson Lukose, and Klaus Tochtermann. Ontology alignment — a survey with focus on visually supported semi-automatic techniques. Future Internet, 2(3):238–258, August 04, 2010. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/2/3/238.

Guillen:2015:IIS

[GSP15] Edward Guillen, Jeisson Sánchez, and Rafael Paez. Inefficiency of IDS static anomaly detectors in realworld networks. Future Internet, 7(2):94–109, May 06, 2015. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/7/2/94.

Grimaldi:2019:LWC

[GTC19] Matteo Grimaldi, Valerio Tenace, and Andrea Calimera. Layer-wise compressive training for convolutional neural networks. Future Internet, 11(1):7, December 28, 2019. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
1/7.

Gatziolis:2022:AUP

[GTM22] Kleanthis G. Gatziolis, Nikolaos D. Tselikas, and Ioannis D. Moscholios. Adaptive user profiling in e-commerce and administration of public services. Future Internet, 14(5):144, May 09, 2022. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
5/144.

Ghosh:2022:CPS

[GTQ^{+}22] Uttam Ghosh, Deepak Tosh, Nawab Muhammad Faseeh Qureshi, Ali Kashif Bashir, Al-Sakib Khan Pathan, and Zhaolong Ning. Cyberphysical systems: Prospects, challenges and role in software-defined networking and blockchains. Future Internet, 14(12):382, December 18, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
12/382.

Gupta:2024:TAC

[GTWN24] Ragini Gupta, Beitong Tian, Yaohui Wang, and Klara Nahrstedt. TWIN-ADAPT: Continuous learning for digital twin-enabled online anomaly classification in IoT-driven smart labs. Future Internet, 16 (7):??, July 4, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
7/239.

Gupta:2013:PLN

[GVC13] Smrati Gupta and M. A. Vázquez-Castro. Physical layer network coding based on integer forcing precoded compute and forward. Future Internet, 5(3):439–459, August 28, 2013. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/5/3/439.

Gonzalez:2023:LLV

[GVV^{+}23] Luis F. Gonzalez, Ivan Vidal, Francisco Valera, Raul Martin, and Dulce Artalejo. A link-layer virtual networking solution for cloud-native network function virtualisation ecosystems: L2S-M. Future Internet, 15(8):274, August 17, 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
8/274.

Goncalves:2023:CSP

[GVY^{+}23] Teresa Gonçalves, Rute Veladas, Hua Yang, Renata Vieira, Paulo Quaresma, Paulo Infante, Cátia Sousa Pinto, João Oliveira, Maria Cortes Ferreira, Jéssica Morais, Ana Raquel Pereira, Nuno Fernandes, and Carolina Gonçalves. Clinical screening prediction in the Portuguese National Health Service: Data analysis, machine learning models, explainability and meta-evaluation. Future Internet, 15(1):26, January 03, 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
1/26.

Gelenbe:2013:FRC

[GW13] Erol Gelenbe and Fang-Jing Wu. Future research on cyber-physical emergency management systems. Future Internet, 5(3):336–354, June 27, 2013. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/5/3/336.

Gao:2015:DLB

[GW15] Ren Gao and Juebo Wu. Dynamic load balancing strategy for cloud computing with ant colony optimization. Future Internet, 7(4):465–483, November 26, 2015. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/7/4/465.

Gao:2023:DRL

[GWF^{+}23] Siyu Gao, Yuchen Wang, Nan Feng, Zhongcheng Wei, and Jijun Zhao. Deep reinforcement learningbased video offloading and resource allocation in NOMA-enabled networks. Future Internet, 15(5): 184, May 18, 2023. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/15/
                                        5/184.

Gkioulos:2017:UMV

[GWK17] Vasileios Gkioulos, Gaute Wangen, and Sokratis K. Katsikas. User modelling validation over the security awareness of digital natives. Future Internet, 9 (3):32, July 10, 2017. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/9/
3/32.

Graziani:2020:ITA

[GX20] Salvatore Graziani and Maria Gabriella Xibilia. Innovative topologies and algorithms for neural networks. Future Internet, 12(7):117, July 11, 2020. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/12/
7/117.

Gao:2019:IMN

[GXWD19] Ming Gao, Qifeng Xiao, Shaochun Wu, and Kun Deng. An improved method for named entity recognition and its application to CEMR. Future Internet, 11(9):185, August 26, 2019. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
9/185.

Gao:2023:IBE

[GZ23] Lei Gao and Xiaoyong Zhu. ICN-based enhanced content delivery for CDN. Future Internet, 15(12): 390, November 30, 2023. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
12/390.

Ghorpade:2021:SLI

[GZC21] Sheetal Ghorpade, Marco Zennaro, and Bharat Chaudhari. Survey of localization for Internet of Things nodes: Approaches, challenges and open issues. Future Internet, 13(8):210, August 16, 2021. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/13/8/210.

Guo:2024:SES

[GZS^{+}24] Xiaotong Guo, Huijie Zhao, Shuwei Shao, Xudong Li, Baochang Zhang, and Na Li. SPDepth: Enhancing self-supervised indoor monocular depth estimation via self-propagation. Future Internet, 16(10), October 16, 2024. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/16/
                                        10/375.

Hasan:2023:EIC

[HA23] Husam H. Hasan and Zainab T. Alisa. Effective IoT congestion control algorithm. Future Internet, 15(4):136, March 31, 2023. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/15/
                                        4/136.

Haimed:2023:EMV

[HAA23] Ibrahim Bu Haimed, Marwan Albahar, and Ali Alzubaidi. Exploiting misconfiguration vulnerabilities in Microsoft’s Azure active directory for privilege escalation attacks. Future Internet, 15(7): 226–??, July 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
7/226.

Hoyhtya:2018:RLA

[HAL18] Marko Höyhtyä, Olli Apilo, and Mika Lasanen. Review of latest advances in 3GPP standardization: D2D communication in 5G systems and its energy consumption models. Future Internet, 10(1):3, January 03, 2018. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/10/1/3.

Hou:2022:EOT

[HAL^{+}22] Changbo Hou, Jiajun Ai, Yun Lin, Chenyang Guan, Jiawen Li, and Wenyu Zhu. Evaluation of online teaching quality based on facial expression recognition. Future Internet, 14(6):177, June 08, 2022. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
6/177.

Han:2013:MEO

[Han13]  Qing Han.  Managing

emergencies optimally using a random neural network-based algorithm. Future Internet, 5(4):515– 534, October 16, 2013. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/5/
                                        4/515.

Henriksen-Bulmer:2020:DCA

[HBFJ20] Jane Henriksen-Bulmer, Shamal Faily, and Sheridan Jeary. DPIA in context: Applying DPIA to assess privacy risks of cyber physical systems. Future Internet, 12(5):93, May 24, 2020. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/12/5/93.

Hitimana:2021:IIF

[HBM^{+}21] Eric Hitimana, Gaurav Bajpai, Richard Musabe, Louis Sibomana, and Jayavel Kayalvizhi. Implementation of IoT framework with data analysis using deep learning methods for occupancy prediction in a building. Future Internet, 13(3):67, March 09, 2021. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/13/3/67.

Hurst:2020:PPV

[HBMS20] William Hurst, Aaron Boddy, Madjid Merabti, and Nathan Shone. Patient privacy violation detection in healthcare critical infrastructures: an investigation using densitybased benchmarking. Future Internet, 12(6):100, June 08, 2020. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/12/
6/100.

Huertas:2019:IET

[HC19] Melby Karina Zuniga Huertas and Tarcisio Duarte Coelho. The interaction effect of type of message X YouTuber’s media metrics on customers’ responses and the moderation of conformity intention. Future Internet, 11(6):135, June 20, 2019. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/11/6/135.

Hilda:2021:CTE

[HC21] Jabanjalin Hilda and Srimathi Chandrasekaran. Cost and time economical planning algorithm for scientific workflows in cloud computing. Future Internet, 13(10):263, October 13, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/10/263.

He:2016:NFM

[HCHP16] Jingsha He, Chengyue Chang, Peng He, and Muhammad Salman Pathan. Network forensics method based on evidence graph and vulnerability reasoning. Future Internet, 8 (4):54, November 10, 2016. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/8/
4/54.

Hsu:2020:REE

[HCW20] Meng-Hsiang Hsu, Chun-Ming Chang, and Shing-Ling Wu. Re-examining the effect of online social support on subjective well-being: The moderating role of experience. Future Internet, 12(5):88, May 15, 2020. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/12/5/88.

Haider:2016:WBD

[HCXH16] Waqas Haider, Gideon Creech, Yi Xie, and Jiankun Hu. Windows based data sets for evaluation of robustness of host based intrusion detection systems (IDS) to zero-day and stealth attacks. Future Internet, 8(3):29, July 05, 2016. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/8/3/29.

Heaberlin:2016:EWN

[HD16] Bradi Heaberlin and Simon DeDeo. The evolution of Wikipedia’s norm network. Future Internet, 8 (2):14, April 20, 2016. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/8/
2/14.

Hackl:2019:ESU

[HD19] Jürgen Hackl and Thibaut Dubernet. Epidemic spreading in urban areas using agent-based transportation models. Future Internet, 11(4):92, April 08, 2019. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/11/4/92.

Hamarsheh:2023:NIT

[HDBD23] Qadri Hamarsheh, Omar Daoud, Mohammed Baniyounis, and Ahlam Damati. Narrowband Internet-of-Things to enhance the vehicular communications performance. Future Internet, 15(1):16, December 28, 2023. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/15/1/16.

Hanzl:2012:HGU

[HDK^{+}12] Małgorzata Hanzl, Karol Dzik, Paulina Kowalczyk, Krystian Kwieciński, Ewa Stankiewicz, and Agata L. Wierzbicka. Human geomatics in urban design-two case studies. Future Internet, 4(1):347–361, March 22, 2012. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/4/1/347.

Hare:2014:EMC

[HDL^{+}14] Jonathon S. Hare, David P. Dupplaw, Paul H. Lewis, Wendy Hall, and Kirk Martinez. Exploiting multimedia in creating and analysing multimedia Web archives. Future Internet, 6(2):242–260, April 24, 2014. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/6/2/242.

Honecker:2022:CDT

[HDT22] Fabian Honecker, Julian Dreyer, and Ralf Tönjes. Comparison of distributed tamper-proof storage methods for public key infrastructures. Future Internet, 14(11):336, November 18, 2022. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/14/
                                        11/336.

Hewage:2020:MQE

[HE20] Chaminda Hewage and Erhan Ekmekcioglu. Multimedia quality of experience (QoE): Current status and future direction. Future Internet, 12(7): 121, July 20, 2020. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/12/
                                        7/121.

Heller:2015:SID

[Hel15] Alfred Heller. The sensing Internet— a discussion on its impact on rural areas. Future Internet, 7(4):363–371, September 28, 2015. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/7/4/363.

Hanka:2011:DPK

[HEP^{+}11] Oliver Hanka, Michael Eichhorn, Martin Pfannenstein, Jörg Eberspächer, and Eckehard Steinbach. A distributed public key infrastructure based on threshold cryptography for the HiiMap next generation Internet architecture. Future Internet, 3(1):14– 30, February 01, 2011. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/3/
1/14.

Hermann:2014:CIP

[HF14] Steffen Hermann and Benjamin Fabian. A comparison of Internet protocol (IPv6) security guidelines. Future Internet, 6(1):1–60, January 10, 2014. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/6/1/1.

Hassan:2020:AAA

[HF20] Walaa Hassan and Tamer Farag. Adaptive allocation algorithm for multiradio multi-channel wireless mesh networks. Future Internet, 12(8):127, July 29, 2020. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/12/8/127.

Huang:2022:OPC

[HF22] Zhaohui Huang and Vasilis Friderikos. Optimal proactive caching for multiview streaming mobile augmented reality. Future Internet, 14(6):166, May 30, 2022. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/14/6/166.

Hunter:2012:TAE

[HG12] Jane Hunter and Anna Gerber. Towards annotopiaenabling the semantic interoperability of Webbased annotations. Future Internet, 4(3):788– 806, August 30, 2012. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/4/
                                        3/788.

Herrera:2021:ERL

[HGF21] Mauricio Herrera and Alex Godoy-Faúndez. Exploring the roles of local mobility patterns, socioeconomic conditions, and lockdown policies in shaping the patterns of COVID-19 spread. Future Internet, 13(5):112, April 28, 2021. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/13/5/112.

Hu:2023:CMN

[HGL^{+}23] Yang Hu, Liangliang Gong, Xinyang Li, Hui Li, Ruoxin Zhang, and Rentao Gu. A carrying method for 5G network slicing in smart grid communication services based on neural network. Future Internet, 15(7):247–??, July 2023. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
7/247.

Hao:2020:MAS

[HH20] Jingjing Hao and Guangsheng Han. On the modeling of automotive security: a survey of methods and perspectives. Future Internet, 12(11):198, November 16, 2020. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/12/
11/198.

Hung:2024:EHT

[HHL24] Jeih-Weih Hung, Pin-Chen Huang, and Li-Yin Li. Employing Huber and TAP losses to improve inter-SubNet in speech enhancement. Future Internet, 16(10), October 4, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/10/360.

Han:2023:RRA

[HHLZ23] Lu Han, Xiaohong Huang, Dandan Li, and Yong Zhang. RingFFL: a ringarchitecture-based fair federated learning framework. Future Internet, 15(2):68, February 09, 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
2/68.

Hasan:2023:RDL

[HHM^{+}23] Md. Tarek Hasan, Md. Al Emran Hossain, Md. Saddam Hossain Mukta, Arifa Akter, Mohiuddin Ahmed, and Salekul Islam. A review on deep-learning-based cyberbullying detection. Future Internet, 15(5):179, May 11, 2023. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
5/179.

Hussein:2018:MBA

[HJA18] Isam A. Hussein, Basil H. Jasim, and Ramzy S. Ali. A modified BA anticollision protocol for coping with capture effect and interference in RFID systems. Future Internet, 10 (10):96, October 01, 2018. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/10/
10/96.

Hameed:2018:SBC

[HK18] Sufian Hameed and Hassan Ahmed Khan. SDN based collaborative scheme for mitigation of DDoS attacks. Future Internet, 10 (3):23, February 27, 2018. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/10/
3/23.

Hossain:2024:HAI

[HKH^{+}24] Mahmud Hossain, Golam Kayas, Ragib Hasan, Anthony Skjellum, Shahid Noor, and S. M. Riazul Islam. A holistic analysis of Internet of Things (IoT) security: Principles, practices, and new perspectives. Future Internet, 16 (2):40, January 24, 2024. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
2/40.

Herouala:2022:CTB

[HKZ^{+}22] Abdelkader Tayeb Herouala, Chaker Abdelaziz Kerrache, Benameur Ziani, Carlos T. Calafate, Nasreddine Lagraa, and Abdou el Karim Tahari. Controlling the trade-off between resource efficiency and user satisfaction in NDNs based on naïve Bayes data classification and Lagrange method. Future Internet, 14(2):48, January 31, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
2/48.

Hsu:2019:BUP

[HL19] Chien-Lung Hsu and Yi-Chuan Liao. Bridging user perception and stickiness in business microblog contexts: a moderated mediation model. Future Internet, 11(6):134, June 19, 2019. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/11/6/134.

Haralambous:2023:BSW

[HL23] Yannis Haralambous and Philippe Lenca. Beyond the semantic Web: Towards an implicit pragmatic Web and a Web of social representations. Future Internet, 15(7): 239–??, July 2023. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/15/
                                        7/239.

Henriques:2024:CAF

[HLC24] Lucas Henriques, Felipe Prata Lima, and Cecilia Castro. Combining advanced feature-selection methods to uncover atypical energy-consumption patterns. Future Internet, 16(7):??, June 28, 2024. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/16/
                                        7/229.

Hamada:2021:SCL

[HLG21] Louiza Hamada, Pascal Lorenz, and Marc Gilg. Security challenges for light emitting systems. Future Internet, 13(11):276, October 28, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/11/276.

Ho:2010:IVA

[HLLT10] Yeh-Chin Ho, Yi-Bing Lin, Ren-Huang Liou, and Yuan-Kuang Tu. Implementing value added applications in next generation networks. Future Internet, 2(3):282–294, August 06, 2010. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/2/3/282.

Han:2018:MAP

[HLZ^{+}18] Chong Han, Zilong Li, Jian Zhou, Lijuan Sun, and Siyu Chen. A multiple access protocol based on gray forecast for satellite network. Future Internet, 10(6):45, June 01, 2018. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/10/
6/45.

He:2023:GTB

[HLZ23] Jingsha He, Yue Li, and Nafei Zhu. A game theorybased model for the dissemination of privacy information in online social networks. Future Internet, 15(3):92, February 27, 2023. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/15/3/92.

Hazari:2020:ITS

[HM20] Shihab Shahriar Hazari and Qusay H. Mahmoud. Improving transaction speed and scalability of blockchain systems via parallel proof of work. Future Internet, 12(8):125, July 27, 2020. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/12/
                                        8/125.

Harnes:2024:SAT

[HM24] Håkon Harnes and Donn Morrison. SoK: Analysis techniques for WebAssembly. Future Internet, 16(3):84, February 29, 2024. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/16/3/84.

Hu:2021:UBM

[HMR^{+}21] Rui Hu, Bruno Michel, Dario Russo, Niccolò Mora, Guido Matrella, Paolo Ciampolini, Francesca Cocchi, Enrico Montanari, Stefano Nunziata, and Thomas Brunschwiler. An unsupervised behavioral modeling and alerting system based on passive sensing for elderly care. Future Internet, 13(1):6, December 30, 2021. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
1/6.

Hoang:2018:BDB

[HN18] Xuan Dau Hoang and Quynh Chi Nguyen. Botnet detection based on machine learning techniques using DNS query data. Future Internet, 10(5):43, May 18, 2018. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/10/5/43.

Hasslinger:2023:SAA

[HNHH23] Gerhard Hasslinger, Konstantinos Ntougias, Frank Hasslinger, and Oliver Hohlfeld. Scope and accuracy of analytic and approximate results for FIFO, clock-based and LRU caching performance. Future Internet, 15(3):91, February 24, 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
3/91.

Hota:2021:ACM

[HNK^{+}21] Lopamudra Hota, Biraja Prasad Nayak, Arun Kumar, G. G. Md. Nawaz Ali, and Peter Han Joo Chong. An analysis on contemporary MAC layer protocols in vehicular networks: State-of-the-art and future directions. Future Internet, 13(11):287, November 17, 2021. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
11/287.

Hernandez-Orellana:2021:CDI

[HOPGRV21] Marisol Hernández-Orellana, Adolfina Pérez-Garcias, and Ángel Roco-Videla. Characterization of the digital identity of Chilean university students considering their personal learning environments. Future Internet, 13(3):74, March 16, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/3/74.

Hortelano:2019:AVG

[Hor19] Lorenzo J. Torres Hortelano. Audio-visual genres and polymediation in successful Spanish YouTubers. Future Internet, 11 (2):40, February 11, 2019. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
2/40.

Hernandez:2023:ADI

[HOV23] Luis Alberto Martínez Hernández, Ana Lucila Sandoval Orozco, and Luis Javier García Villalba. Analysis of digital information in storage devices using supervised and unsupervised natural language processing techniques. Future Internet, 15(5):155, April 23, 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
5/155.

Hoang:2016:RHI

[HP16] Xuan Dau Hoang and Hong Ky Pham. A review on hot-IP finding methods and its application in early DDoS target detection. Future Internet, 8 (4):52, October 25, 2016. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/8/
4/52.

Hyla:2019:EIM

[HP19] Tomasz Hyla and Jerzy Pejaś. eHealth integrity model based on permissioned blockchain. Future Internet, 11(3):76, March 24, 2019. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/11/3/76.

Harborth:2020:EIE

[HP20] David Harborth and Sebastian Pape. Empirically investigating extraneous influences on the “APCO” model — childhood brand nostalgia and the positivity bias. Future Internet, 12(12):220, December 02, 2020. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/12/
12/220.

Helmstetter:2021:CLS

[HP21] Stefan Helmstetter and Heiko Paulheim. Collecting a large scale dataset for classifying fake news tweets using weak supervision. Future Internet, 13(5):114, April 29, 2021. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/13/
                                        5/114.

Harutyunyan:2024:DCA

[HPB^{+}24] Ashot Harutyunyan, Arnak Poghosyan, Tigran Bunarjyan, Andranik Haroyan, Marine Harutyunyan, Lilit Harutyunyan, and Nelson Baloian. Discovery of cloud applications from logs. Future Internet, 16 (6):216, June 18, 2024. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/16/
                                        6/216.

Harrison:2012:COG

[HPC12] Teresa M. Harrison, Theresa A. Pardo, and Meghan Cook. Creating open government ecosystems: a research and development agenda. Future Internet, 4(4):900– 928, October 23, 2012. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/4/
                                        4/900.

Haz:2024:FOS

[HPF^{+}24] Amma Liesvarastranta Haz, Yohanes Yohanie Fridelin Panduman, Nobuo Funabiki, Evianita Dewi Fajrianti, and Sritrusta Sukaridhoto. Fully opensource meeting minutes generation tool. Future Internet, 16(11):429, November 20, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/11/429.

Hatami:2024:SRT

[HQC^{+}24] Mohsen Hatami, Qian Qu, Yu Chen, Hisham Kholidy, Erik Blasch, and Erika Ardiles-Cruz. A survey of the real-time metaverse: Challenges and opportunities. Future Internet, 16 (10), October 18, 2024. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
10/379.

Hammood:2019:BBC

[HRAA19] Dalal Abdulmohsin Hammood, Hasliza A. Rahim, Ahmed Alkhayyat, and R. Badlishah Ahmad. Body-to-body cooperation in Internet of medical things: Toward energy efficiency improvement. Future Internet, 11(11):239, November 14, 2019. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
11/239.

Hueso-Romero:2021:STM

[HRGQHOA21] José Javier Hueso-Romero, Javier Gil-Quintana, Helen Hasbun, and Sara Osuna-Acedo. The social and transfer massive open online course: Post-digital learning. Future Internet, 13(5):119, April 30, 2021. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/13/
                                        5/119.

Hu:2021:SCA

[HRKL21] Minda Hu, Ashwin Rao, Mayank Kejriwal, and Kristina Lerman. Socioeconomic correlates of antiscience attitudes in the US. Future Internet, 13(6): 160, June 19, 2021. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/13/
                                        6/160.

Hudson-Smith:2009:FI

[HS09] Andrew Hudson-Smith. The future Internet. Future Internet, 1(1):1–2, July 17, 2009. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/1/1/1.

Hardaha:2018:SDR

[HS18a] Prakash Narayan Hardaha and Shailendra Singh. Structured data REST protocol for end to end data mashup. Future Internet, 10(10):98, October 04, 2018. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/10/10/98.

Huang:2018:MIU

[HS18b] Yi-Ting Huang and Sheng-Fang Su. Motives for Instagram use and topics of interest among young adults. Future Internet, 10(8):77, August 09, 2018. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/10/
8/77.

Huang:2020:ANS

[HS20] Hailong Huang and Andrey V. Savkin. Autonomous navigation of a solar-powered UAV for secure communication in urban environments with eavesdropping avoidance. Future Internet, 12(10): 170, October 10, 2020. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/12/
10/170.

Hall:2019:BIA

[HSC19] Calum C. Hall, Lynsay A.

Shepherd, and Natalie Coull. BlackWatch: Increasing attack awareness within Web applications. Future Internet, 11(2):44, February 15, 2019. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
2/44.

Han:2022:CFM

[HSC^{+}22] Junyan Han, Huili Shi, Longfei Chen, Hao Li, and Xiaoyuan Wang. The carfollowing model and its applications in the V2X environment: a historical review. Future Internet, 14(1):14, December 27, 2022. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/14/1/14.

Sun:2017:SQE

[hSgLH17] Mao hua Sun, Yuan gang Li, and Bing He. Study on a quality evaluation method for college English classroom teaching. Future Internet, 9(3):41, July 30, 2017. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/9/3/41.

Habib:2022:BTB

[HSI^{+}22] Gousia Habib, Sparsh Sharma, Sara Ibrahim, Imtiaz Ahmad, Shaima Qureshi, and Malik Ishfaq. Blockchain technology: Benefits, challenges, applications, and integration of blockchain technology with cloud computing. Future Internet, 14(11): 341, November 21, 2022. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/14/
                                        11/341.

Huang:2023:NNA

[HSL^{+}23] Songtao Huang, Jun Shen, Qingquan Lv, Qingguo Zhou, and Binbin Yong. A novel NODE approach combined with LSTM for short-term electricity load forecasting. Future Internet, 15(1):22, December 30, 2023. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/15/1/22.

Hao:2019:IVD

[HSM19] Ruohong Hao, Bingjia Shao, and Rong Ma. Impacts of video display on purchase intention for digital and home appliance products-empirical study from China. Future Internet, 11(11):224, October 24, 2019. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/11/11/224.

Hernandez-Suarez:2023:MAI

[HSSPTM^{+}23] Aldo Hernandez-Suarez, Gabriel Sanchez-Perez, Linda Karina Toscano-Medina, Hector Manuel Perez-Meana, Jose Portillo-Portillo, and Jesus Olivares-Mercado. Methodological approach for identifying Websites with infringing content via text transformers and dense neural networks. Future Internet, 15(12):397, December 09, 2023. CODEN ????

ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/15/12/397.

Hurst:2023:DAM

[HSTK23] William Hurst, Orestis Spyrou, Bedir Tekinerdogan, and Caspar Krampe. Digital art and the metaverse: Benefits and challenges. Future Internet, 15(6):188–??, June 2023. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/15/
                                        6/188.

Hatlevik:2014:UMA

[HT14] Ove Edvard Hatlevik and Karoline Tømte. Using multilevel analysis to examine the relationship between upper secondary students Internet safety awareness, social background and academic aspirations. Future Internet, 6(4):717–734, November 14, 2014. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/6/4/717.

Huang:2015:PDE

[Hua15] Yi-Ting Huang. Participatory design to enhance ICT learning and community attachment: a case study in rural Taiwan. Future Internet, 7(1):50–66, March 02, 2015. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/7/1/50.

Hoppari:2021:PGI

[HUM^{+}21] Mika Hoppari, Mikko Uitto, Jukka Mäkelä, Ilkka Harjula, and Seppo Rantala. Performance of the 5th generation indoor wireless technologies — empirical study. Future Internet, 13(7):180, July 09, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/7/180.

Hanna:2017:SSD

[HVS17] Dalal Hanna, Prakash Veeraraghavan, and Ben Soh. SDMw: Secure dynamic middleware for defeating port and OS scanning. Future Internet, 9 (4):67, October 21, 2017. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/9/
4/67.

Huckle:2016:SB

[HW16] Steve Huckle and Martin White. Socialism and the blockchain. Future Internet, 8(4):49, October 18, 2016. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/8/4/49.

Hu:2012:SLP

[HWCH12] Yuh-Jong Hu, Win-Nan Wu, Kua-Ping Cheng, and Ya-Ling Huang. Semantic legal policies for data exchange and protection across super-peer domains in the cloud. Future Internet, 4(4):929–954, October 25, 2012. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/4/4/929.

Hong:2017:BEF

[HWCL17] Zhen Hong, Zehua Wang, Wei Cai, and Victor C. M. Leung. Blockchainempowered fair computational resource sharing system in the D2D network. Future Internet, 9(4):85, November 17, 2017. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/9/
4/85.

He:2018:MBF

[HWZP18] Jingsha He, Jianan Wu, Nafei Zhu, and Muhammad Salman Pathan. MinHash-based fuzzy keyword search of encrypted data across multiple cloud servers. Future Internet, 10(5):38, May 01, 2018. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/10/
5/38.

He:2017:APP

[HXHP17] Jingsha He, Qi Xiao, Peng He, and Muhammad Salman Pathan. An adaptive privacy protection method for smart home environments using supervised learning.

Future Internet, 9(1):7, March 05, 2017. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/9/1/7.

Hu:2022:SSG

[HY22] Jingyuan Hu and Zhouwang Yang. Single-shot global and local context refinement neural network for head detection. Future Internet, 14(12):384, December 19, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
12/384.

Huang:2013:SSV

[HYL13] Jun Steed Huang, Oliver Yang, and Funmilyo Lawal. Sending safety video over WiMAX in vehicle communications. Future Internet, 5(4):535–567, October 31, 2013. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/5/4/535.

Huang:2019:CFE

[HYLP19] Yongrui Huang, Jianhao Yang, Siyu Liu, and Jiahui Pan. Combining facial expressions and electroencephalography to enhance emotion recognition. Future Internet, 11 (5):105, May 02, 2019. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
5/105.

Hang:2024:LLM

[HYMT24] Ching-Nam Hang, Pei-Duo Yu, Roberto Morabito, and Chee-Wei Tan. Large language models meet next-generation networking technologies: a review. Future Internet, 16 (10), October 7, 2024. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/16/
                                        10/365.

Hu:2022:SVD

[HZCZ22] Zhixi Hu, Yi Zhu, Xiaoying Chen, and Yu Zhao. Safety verification of driving resource occupancy rules based on functional language. Future Internet, 14(2):60, February 17, 2022. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/14/2/60.

He:2024:SEA

[HZQ24] Peixiong He, Yi Zhou, and Xiao Qin. A survey on energy-aware security mechanisms for the Internet of Things. Future Internet, 16(4):128, April 08, 2024. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/16/4/128.

Han:2020:ECF

[HZW^{+}20] Junyan Han, Jinglei Zhang, Xiaoyuan Wang, Yaqi Liu, Quanzheng Wang, and Fusheng Zhong. An extended car-following model considering generalized preceding vehicles in V2X environment. Future Internet, 12(12):216, November 28, 2020. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/12/
12/216.

Hao:2021:EIS

[HZWJ21] Chuanyan Hao, Anqi Zheng, Yuqi Wang, and Bo Jiang. Experiment information system based on an online virtual laboratory. Future Internet, 13 (2):27, January 24, 2021. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
2/27.

Iqbal:2021:IPC

[IAG^{+}21] Umar Iqbal, Ashraf Abosekeen, Jacques Georgy, Areejah Umar, Aboelmagd Noureldin, and Michael J. Korenberg. Implementation of parallel cascade identification at various phases for integrated navigation system. Future Internet, 13(8): 191, July 26, 2021. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
8/191.

Imran:2024:ERA

[IAM24] Muhammad Imran, Annalisa Appice, and Donato Malerba. Evaluating realistic adversarial attacks against machine learning models for Windows PE malware detection. Future Internet, 16(5):168, May 12, 2024. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/16/5/168.

Ijaz:2020:RHP

[IBBA20] Qaiser Ijaz, El-Bay Bourennane, Ali Kashif Bashir, and Hira Asghar. Revisiting the high-performance reconfigurable computing for future datacenters. Future Internet, 12(4):64, April 06, 2020. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/12/4/64.

Imhof:2011:NRK

[ICF^{+}11] Mark Imhof, Matthew Cox, Angela Fadersen, Wayne Harvey, Sonia Thompson, David Rees, and Christopher Pettit. Natural resource knowledge and information management via the Victorian resources online Website. Future Internet, 3(4):248– 280, November 09, 2011. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/3/
                                        4/248.

Ismail:2023:SWS

[IDR23] Shereen Ismail, Diana W.

Dawoud, and Hassan Reza. Securing wireless sensor networks using machine learning and blockchain: a review. Future Internet, 15(6):200– ??, June 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
6/200.

Ibrahim:2021:FBC

[IFF21] Ahmed H. Ibrahim, Zaki T.

Fayed, and Hossam M. Faheem. Fog-based CDN framework for minimizing latency of Web services using fog-based HTTP browser. Future Internet, 13(12):320, December 17, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/12/320.

Islam:2010:NEI

[IG10] Salekul Islam and Jean-Charles Grégoire. Network edge intelligence for the emerging next-generation Internet. Future Internet, 2(4):603–623, November 05, 2010. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/2/4/603.

Ilie:2023:ADH

[IGL^{+}23] Dragos Ilie, Håkan Grahn, Lars Lundberg, Alexander Westerhagen, Bo Granbom, and Anders Höök. Avoiding detection by hostile nodes in airborne tactical networks. Future Internet, 15(6):204–??, June 2023. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
6/204.

Ibarra-Hernandez:2024:MLS

[IHCSG^{+}24] Roilhi F. Ibarra-Hernández, Francisco R. Castillo-Soria, Carlos A. Gutiérrez, Abel García-Barrientos, Luis Alberto Vásquez-Toledo, and J. Alberto Del-Puerto-Flores. Machine learning strategies for reconfigurable intelligent surface-assisted communication systems — a review. Future Internet, 16(5):173, May 17, 2024. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
5/173.

Islam:2022:NML

[IHMG22] Mohammad S. Islam, Shahid Husain, Jawed Mustafa, and Yuantong Gu. A novel machine learning prediction model for aerosol transport in upper 17-generations of the human respiratory tract. Future Internet, 14(9):247, August 24, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
9/247.

Ismail:2016:FST

[IIOW16] Umar Mukhtar Ismail, Shareeful Islam, Moussa Ouedraogo, and Edgar Weippl. A framework for security transparency in cloud computing. Future Internet, 8(1):5, February 17, 2016. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/8/1/5.

Imadali:2013:ESS

[IKP^{+}13] Sofiane Imadali, Athanasia Karanasiou, Alexandru Petrescu, Ioannis Sifniadis, Eleftheria Velidou, Véronique Vèque, and Pantelis Angelidis. eHealth service support in future IPv6 vehicular networks. Future Internet, 5(3):317–335, June 27, 2013. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/5/3/317.

Ivanov:2021:IUM

[IKS^{+}21] Vladimir M. Ivanov, Anton M. Krivtsov, Sergey V. Strelkov, Nikolay V. Kalakutskiy, Andrey I. Yaremenko, Marina Yu. Petropavlovskaya, Maria N. Portnova, Olga V. Lukina, and Andrey P. Litvinov. Intraoperative use of mixed reality technology in median neck and branchial cyst excision. Future Internet, 13 (8):214, August 18, 2021.

CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/13/
                                        8/214.

Irons:2014:DFI

[IL14] Alastair Irons and Harjinder Singh Lallie. Digital forensics to intelligent forensics. Future Internet, 6(3):584–596, September 12, 2014. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/6/3/584.

Islam:2024:DFV

[IMLL24] Tasin Islam, Alina Miron, Xiaohui Liu, and Yongmin Li. Dynamic fashion video synthesis from static imagery. Future Internet, 16(8):??, August 8, 2024. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/16/
                                        8/287.

Ibadah:2018:SCR

[IMR^{+}18] Nisrine Ibadah, Khalid Minaoui, Mohammed Rziza, Mohammed Oumsis, and César Benavente-Peces. Smart collection of real-time vehicular mobility traces. Future Internet, 10(8):78, August 09, 2018. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/10/
                                        8/78.

Ibba:2016:DLC

[IP16] Simona Ibba and Filippo Eros Pani. Digital libraries: The challenge of integrating Instagram with a taxonomy for content management. Future Internet, 8(2):16, May 10, 2016. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/8/2/16.

Ibba:2018:ICO

[IPL^{+}18] Simona Ibba, Andrea Pinna, Maria Ilaria Lunesu, Michele Marchesi, and Roberto Tonelli. Initial coin offerings and agile practices. Future Internet, 10(11):103, October 23, 2018. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/10/11/103.

Iordache:2024:EAV

[IPP24] Stefan Iordache, Catalina Camelia Patilea, and Ciprian Paduraru. Enhancing autonomous vehicle safety with blockchain technology: Securing vehicle communication and AI systems. Future Internet, 16(12):471, December 18, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/12/471.

Imbimbo:2020:ACI

[ISG20] Enrico Imbimbo, Federica Stefanelli, and Andrea Guazzini. Adolescent’s collective intelligence: Empirical evidence in real and online classmates groups. Future Internet, 12(5):81, April 29, 2020. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/12/5/81.

Iglesias-Sanuy:2022:ELB

[ISLARP^{+}22] Pablo Iglesias-Sanuy, José Carlos López-Ardao, Miguel Rodríguez-Pérez, Sergio Herrería-Alonso, Andrés Suárez-González, and Raúl F. Rodríguez-Rubio. An efficient location-based forwarding strategy for named data networking and LEO satellite communications. Future Internet, 14(10):285, September 29, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/10/285.

Ivanov:2024:CLM

[IT24] Valeriy Ivanov and Maxim Tereshonok. Cross-layer methods for ad hoc networks — review and classification. Future Internet, 16(1):29, January 16, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/1/29.

Isobe:2021:MAL

[ITH^{+}21] Shinnosuke Isobe, Satoshi Tamura, Satoru Hayamizu, Yuuto Gotoh, and Masaki Nose. Multi-angle lipreading with angle classificationbased feature extraction and its application to audio-visual speech recognition. Future Internet, 13(7):182, July 15, 2021. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
7/182.

Ismailov:2020:VOS

[ITZ20] Max Ismailov, Michail Tsikerdekis, and Sherali Zeadally. Vulnerabilities to online social network identity deception detection research and recommendations for mitigation. Future Internet, 12(9):148, August 31, 2020. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/12/
9/148.

Imoize:2023:OQS

[IUIL23] Agbotiname Lucky Imoize, Friday Udeji, Joseph Isabona, and Cheng-Chi Lee. Optimizing the quality of service of mobile broadband networks for a dense urban environment. Future Internet, 15(5):181, May 12, 2023. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
5/181.

Iqbal:2023:OTE

[IUK^{+}23] Mansoor Iqbal, Zahid Ullah, Izaz Ahmad Khan, Sheraz Aslam, Haris Shaheer, Mujtaba Humayon, Muhammad Asjad Salahuddin, and Adeel Mehmood. Optimizing task execution: The impact of dynamic time quantum and priorities on round Robin scheduling. Future Internet, 15(3):104, March 08, 2023. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/15/3/104.

Isyaku:2020:SDN

[IZK^{+}20] Babangida Isyaku, Mohd Soperi Mohd Zahid, Maznah Bte Kamat, Kamalrulnizam Abu Bakar, and Fuad A. Ghaleb. Software defined networking flow table management of Open-Flow switches performance and security challenges: A survey. Future Internet, 12 (9):147, August 31, 2020. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/12/
                                        9/147.

Iqbal:2023:ESC

[IZK^{+}23] Asad Iqbal, Muhammad Zubair, Muhammad Asghar Khan, Insaf Ullah, Ghani Ur-Rehman, Alexey V. Shvetsov, and Fazal Noor. An efficient and secure certificateless aggregate signature scheme for vehicular ad hoc networks. Future Internet, 15(8):266, August 10, 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
8/266.

Jaloudi:2019:CPI

[Jal19] Samer Jaloudi. Communication protocols of an industrial Internet of Things environment: a comparative study. Future Internet, 11(3):66, March 07, 2019. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/11/3/66.

Javed:2021:PPA

[JAS21] Yousra Javed, Elham Al Qahtani, and Mohamed Shehab. Privacy policy analysis of banks and mobile money services in the Middle East. Future Internet, 13(1):10, January 03, 2021. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/13/1/10.

Jain:2023:VGB

[JBS^{+}23] Shushant Kumar Jain, Rinkoo Bhatia, Neeraj Shrivastava, Sharad Salunke, Mohammad Farukh Hashmi, and Neeraj Dhanraj Bokde. Virtual grid-based routing for query-driven wireless sensor networks. Future Internet, 15(8):259, July 30, 2023. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/15/8/259.

Jesus:2022:SMQ

[JCPV22] Thiago C. Jesus, Daniel G. Costa, Paulo Portugal, and Francisco Vasques. A survey on monitoring quality assessment for wireless visual sensor networks. Future Internet, 14(7): 213, July 19, 2022. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/14/
                                        7/213.

Jahromi:2021:STC

[JDH21] Hamed Z. Jahromi, Declan Delaney, and Andrew Hines. A sign of things to come: Predicting the perception of above-the-fold time in Web browsing. Future Internet, 13 (2):50, February 17, 2021. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/13/
                                        2/50.

Javed:2024:ILM

[JEJ^{+}24] Abbas Javed, Amna Ehtsham, Muhammad Jawad, Muhammad Naeem Awais, Ayyaz ul Haq Qureshi, and Hadi Larijani. Implementation of lightweight machine learning-based intrusion detection system on IoT devices of smart homes. Future Internet, 16 (6):200, June 5, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
6/200.

Jeon:2020:DEC

[Jeo20] Soobin Jeon. Data exchange in cluster structure for longevity of IoT. Future Internet, 12(2):32, February 12, 2020. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/12/
2/32.

Jimenez:2021:CUE

[JGV^{+}21] Ivonne Angelica Castiblanco Jimenez, Laura Cristina Cepeda García, Maria Grazia Violante, Federica Marcolin, and Enrico Vezzetti. Commonly used external TAM variables in e-learning, agriculture and virtual reality applications. Future Internet, 13(1):7, December 31, 2021. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/13/1/7.

Jiang:2024:UCM

[JGY^{+}24] Shui Jiang, Yanning Ge, Xu Yang, Wencheng Yang, and Hui Cui. UAV control method combining reptile meta-reinforcement learning and generative adversarial imitation learning. Future Internet, 16(3):105, March 20, 2024. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/16/
                                        3/105.

Jiang:2024:TOR

[JHBW24] Guiwen Jiang, Rongxi Huang, Zhiming Bao, and Gaocai Wang. A task offloading and resource allocation strategy based on multi-agent reinforcement learning in mobile edge computing. Future Internet, 16(9), September 11, 2024. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/16/9/333.

Jiang:2020:PTI

[JHC^{+}20] Bo Jiang, Yanbai He, Rui Chen, Chuanyan Hao, Sijiang Liu, and Gangyao Zhang. Progressive teaching improvement for small scale learning: A case study in China. Future Internet, 12(8):137, August 17, 2020. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/12/8/137.

Johnson:2018:HAI

[JHL^{+}18] Shardrom Johnson, Jinwu Han, Yuanchen Liu, Li Chen, and Xinlin Wu. Hybrid approach with improved genetic algorithm and simulated annealing for thesis sampling. Future Internet, 10(8):71, July 30, 2018. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/10/
8/71.

Jiang:2019:MII

[Jia19] Shengming Jiang. Marine Internet for Internetworking in oceans: a tutorial. Future Internet, 11(7):146, July 05, 2019. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
7/146.

Joshi:2017:FCB

[JII17] Pusp Raj Joshi, Shareeful Islam, and Syed Islam. A framework for cloud based e-government from the perspective of developing countries. Future Internet, 9(4):80, November 09, 2017. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/9/4/80.

Jung:2021:BIM

[JJ21] Hyunjun Jung and Dongwon Jeong. Blockchain implementation method for interoperability between CBDCs. Future Internet, 13(5):133, May 18, 2021. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
5/133.

Jatoth:2022:ICB

[JJFC22] Chandrashekar Jatoth, Rishabh Jain, Ugo Fiore, and Subrahmanyam Chatharasupalli. Improved classification of blockchain transactions using feature engineering and ensemble learning. Future Internet, 14(1):16, December 28, 2022. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/14/1/16.

Jahedi:2020:VSH

[JK20] Zahra Jahedi and Thomas Kunz. The value of simple heuristics for virtualized network function placement. Future Internet, 12(10):161, September 25, 2020. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/12/10/161.

Jin:2024:WFI

[JK24] Shanyu Jin and Dongwoo Kim. WiFi fingerprint indoor localization employing Adaboost and probability-one access point selection for multifloor campus buildings. Future Internet, 16(12): 466, December 13, 2024. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
12/466.

Jadhav:2023:ITA

[JKM^{+}23] Hindavi Kishor Jadhav, Vinoth Babu Kumaravelu, Arthi Murugadass, Agbotiname Lucky Imoize, Poongundran Selvaprabhu, and Arunkumar Chandrasekhar. Intelligent transmit antenna selection schemes for high-rate fully generalized spatial modulation. Future Internet, 15 (8):281, August 21, 2023. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
8/281.

Jin:2019:JLD

[JLGF19] Mingshuang Jin, Hongbin Luo, Shuai Gao, and Bohao Feng. Joint locationdependent pricing and request mapping in ICNbased telco CDNs for 5G. Future Internet, 11(6):125, June 03, 2019. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
6/125.

Jeong:2013:FAA

[JLH^{+}13] Sooman Jeong, Kisung Lee, Jungwoo Hwang, Seongjin Lee, and Youjip Won. Framework for analyzing Android I/O stack behavior: From generating the workload to analyzing the trace. Future Internet, 5(4):591–610, December 13, 2013. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/5/4/591.

Jia:2019:IBC

[JLW19] Yun-Fei Jia, Shan Li, and Renbiao Wu. Incorporating background checks with sentiment analysis to identify violence risky Chinese microblogs. Future Internet, 11(9):200, September 19, 2019. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/11/
                                        9/200.

Ji:2018:DIR

[JLZ18] Wusheng Ji, Li Li, and Weiwei Zhou. Design and implementation of a RFID reader/router in RFID-WSN hybrid system. Future Internet, 10(11):106, November 03, 2018. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/10/
                                        11/106.

Jamshidi:2023:FDD

[JMJ^{+}23] Mohammad (Behdad) Jamshidi, Omid Moztarzadeh, Alireza Jamshidi, Ahmed Abdelgawad, Ayman S. El-Baz, and Lukas Hauer. Future of drug discovery: The synergy of edge computing, Internet of Medical Things, and deep learning. Future Internet, 15(4): 142, April 07, 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
4/142.

Jian:2022:STW

[JMZ^{+}22] Bijian Jian, Chunbo Ma, Dejian Zhu, Yixiao Sun, and Jun Ao. Seeing through wavy water–air interface: a restoration model for instantaneous images distorted by surface waves. Future Internet, 14(8):236, July 29, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/8/236.

John:2024:IBR

[JNARV^{+}24] Jobish John, Md. Noor-A-Rahim, Aswathi Vijayan, H. Vincent Poor, and Dirk Pesch. Industry 4.0 and beyond: The role of 5G, WiFi 7, and time-sensitive networking (TSN) in enabling smart manufacturing. Future Internet, 16 (9), September 21, 2024. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
9/345.

James:2022:EDD

[JOL22] Nicholas James, Lee-Yeng Ong, and Meng-Chew Leow. Exploring distributed deep learning inference using Raspberry Pi Spark cluster. Future Internet, 14(8):220, July 25, 2022. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/14/8/220.

Jang:2013:CIS

[JP13] S. Mo Jang and Yong Jin Park. The citizen as issue specialists in a changing media environment. Future Internet, 5(4):568– 579, November 13, 2013. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/5/
                                        4/568.

Jiang:2020:HTD

[JP20] Fenyu Jiang and Chris Phillips. High throughput data relay in UAV wireless networks. Future Internet, 12(11):193, November 09, 2020. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/12/
                                        11/193.

Jimenez:2021:DTB

[JP21] María Artemisa Sangermán Jiménez and Pedro Ponce. Differentiated teaching based on standardized metrics integrating fuzzy logic type 2 detection theory: High school case-PrepaTec, Mexico. Future Internet, 13(4):98, April 13, 2021. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/13/4/98.

Jimenez:2021:YVV

[JPVC21] María Artemisa Sangermán Jiménez, Pedro Ponce, and Esteban Vázquez-Cano. YouTube videos in the virtual flipped classroom model using brain signals and facial expressions. Future Internet, 13(9):224, August 30, 2021. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
9/224.

Jaramillo-Ramirez:2021:SDF

[JRP21] Daniel Jaramillo-Ramirez and Manuel Perez. Spectrum demand forecasting for IoT services. Future Internet, 13(9):232, September 08, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/9/232.

Johnphill:2023:SHC

[JSAO^{+}23] Obinna Johnphill, Ali Safaa Sadiq, Feras Al-Obeidat, Haider Al-Khateeb, Mohammed Adam Taheir, Omprakash Kaiwartya, and Mohammed Ali. Selfhealing in cyber-physical systems using machine learning: a critical analysis of theories and tools. Future Internet, 15(7): 244–??, July 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
7/244.

Jamshidi:2023:MMI

[JSJM23] Mohammad (Behdad) Jamshidi, Arash Dehghaniyan Serej, Alireza Jamshidi, and Omid Moztarzadeh. The meta-metaverse: Ideation and future directions. Future Internet, 15(8):252, July 27, 2023. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/15/
                                        8/252.

Jussila:2021:TAM

[JSPH21] Jari Jussila, Anu Helena Suominen, Atte Partanen, and Tapani Honkanen. Text analysis methods for misinformation-related research on Finnish language Twitter. Future Internet, 13(6):157, June 17, 2021. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/13/
                                        6/157.

Jiang:2023:DPM

[JT23] Lili Jiang and Vicenç Torra. Data protection and multi-database datadriven models. Future Internet, 15(3):93, February 27, 2023. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/15/3/93.

Jalata:2021:MAN

[JTA^{+}21] Ibsa K. Jalata, Thanh-Dat Truong, Jessica L. Allen, Han-Seok Seo, and Khoa Luu. Movement analysis for neurological and musculoskeletal disorders using graph convolutional neural network. Future Internet, 13(8):194, July 28, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/8/194.

Jamroz:2023:OAS

[JUH^{+}23] Zainab Jamroz, Insaf Ullah, Bilal Hassan, Noor Ul Amin, Muhammad Asghar Khan, Pascal Lorenz, and Nisreen Innab. An optimal authentication scheme through dual signature for the Internet of Medical Things. Future Internet, 15(8):258, July 30, 2023. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
8/258.

Jamal:2024:HMA

[JUN^{+}24] Mansoor Jamal, Zaib Ullah, Muddasar Naeem, Musarat Abbas, and Antonio Coronato. A hybrid multi-agent reinforcement learning approach for spectrum sharing in vehicular networks. Future Internet, 16(5):152, April 28, 2024. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
5/152.

Jurgenson:2012:WAM

[Jur12] Nathan Jurgenson. When atoms meet bits: Social media, the mobile Web and augmented revolution. Future Internet, 4(1):83– 91, January 23, 2012. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/4/
                                        1/83.

Jiang:2024:LDT

[JWD^{+}24] Yuning Jiang, Wei Wang, Jianguo Ding, Xin Lu, and Yanguo Jing. Leveraging digital twin technology for enhanced cybersecurity in cyber–physical production systems. Future Internet, 16(4):134, April 17, 2024. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/16/
                                        4/134.

Jiang:2017:EER

[JWRX17] Fan Jiang, Honglin Wang, Hao Ren, and Shuai Xu. Energy-efficient resource and power allocation for underlay multicast device-to-device transmission. Future Internet, 9 (4):84, November 14, 2017. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/9/
                                        4/84.

Jamal:2019:DLB

[JXA19] Nasir Jamal, Chen Xianqiao, and Hamza Aldabbas. Deep learningbased sentimental analysis for large-scale imbalanced Twitter data. Future Internet, 11(9):190, September 02, 2019. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
9/190.

Khrais:2021:RMA

[KA21] Laith T. Khrais and Abdullah M. Alghamdi. The role of mobile application acceptance in shaping ecustomer service. Future Internet, 13(3):77, March 19, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/3/77.

Kulakli:2023:RTI

[KA23] Atik Kulakli and Cenk Lacin Arikan. Research trends of the Internet of Things in relation to business model innovation: Results from co-word and content analyses. Future Internet, 15 (2):81, February 17, 2023. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
2/81.

Khamaiseh:2022:RDK

[KAAAA22] Samer Khamaiseh, Abdullah Al-Alaj, Mohammad Adnan, and Hakam W. Alomari. The robustness of detecting known and unknown DDoS saturation attacks in SDN via the integration of supervised and semi-supervised classifiers. Future Internet, 14(6):164, May 27, 2022. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
6/164.

Kabashkin:2023:EES

[Kab23] Igor Kabashkin. End-to-end service availability in heterogeneous multi-tier cloud–fog–edge networks. Future Internet, 15(10): 329, October 06, 2023. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
10/329.

Kaleem:2024:ORP

[KAEAB24] Sarah Kaleem, Muhammad Asim, Mohammed El-Affendi, and Muhammad Babar. Optimizing requirements prioritization for IoT applications using extended analytical hierarchical process and an advanced grouping framework. Future Internet, 16 (5):160, May 6, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
5/160.

Kiptoo:2019:ATE

[KAL^{+}19] Mark Kipngetich Kiptoo, Oludamilare Bode Adewuyi, Mohammed Elsayed Lotfy, Theophilus Amara, Keifa Vamba Konneh, and Tomonobu Senjyu. Assessing the technoeconomic benefits of flexible demand resources scheduling for renewable energy-based smart microgrid planning. Future Internet, 11(10):219, October 22, 2019. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/11/10/219.

Khan:2017:TAC

[KALY17] Muhidul Islam Khan, Muhammad Mahtab Alam, Yannick Le Moullec, and Elias Yaacoub. Throughputaware cooperative reinforcement learning for adaptive resource allocation in device-to-device communication. Future Internet, 9(4):72, November 01, 2017. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/9/4/72.

Kamal:2019:IBS

[KAM^{+}19] Miraal Kamal, Manal Atif, Hafsa Mujahid, Tamer Shanableh, A. R. Al-Ali, and Ahmad Al Nabulsi. IoT based smart city bus stops. Future Internet, 11(11):227, October 26, 2019. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/11/11/227.

Kumar:2020:PPA

[KAS^{+}20] Ajit Kumar, Vinti Agarwal, Shishir Kumar Shandilya, Andrii Shalaginov, Saket Upadhyay, and Bhawna Yadav. PACER: Platform for Android malware classification, performance evaluation and threat reporting. Future Internet, 12(4):66, April 12, 2020. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/12/
4/66.

Khanal:2022:UBI

[KAS^{+}22] Yurika Pant Khanal, Abeer Alsadoon, Khurram Shahzad, Ahmad B. Al-Khalil, Penatiyana W. C. Prasad, Sabih Ur Rehman, and Rafiqul Islam. Utilizing blockchain for IoT privacy through enhanced ECIES with secure hash function. Future Internet, 14(3):77, February 28, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
3/77.

Kouhoue:2021:EVV

[KBBH21] Austin Waffo Kouhoué, Yoann Bonavero, Thomas Bouétou Bouétou, and Marianne Huchard. Exploring variability of visual accessibility options in operating systems. Future Internet, 13(9):230, September 04, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/9/230.

Kouahla:2022:SBI

[KBF^{+}22] Zineddine Kouahla, Ala-Eddine Benrazek, Mohamed Amine Ferrag, Brahim Farou, Hamid Seridi, Muhammet Kurulay, Adeel Anjum, and Alia Asheralieva. A survey on big IoT data indexing: Potential solutions, recent advancements, and open issues. Future Internet, 14 (1):19, December 31, 2022. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
1/19.

Koirala:2023:EID

[KBF23] Ashish Koirala, Rabindra Bista, and Joao C. Ferreira. Enhancing IoT device security through network attack data analysis using machine learning algorithms. Future Internet, 15(6):210–??, June 2023. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
6/210.

Kenyeres:2024:MTM

[KBHP24] Martin Kenyeres, Ivana Budinská, Ladislav Hluchý, and Agostino Poggi. Modern trends in multi-agent systems. Future Internet, 16(2):54, February 08, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/2/54.

Kirigin:2022:GBT

[KBP22] Tajana Ban Kirigin, Sanda Bujacić Babić, and Benedikt Perak. Graph-based taxonomic semantic class labeling. Future Internet, 14(12):383, December 19, 2022. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/14/12/383.

Korla:2019:MLW

[KC19] Swaroopa Korla and Shanti Chilukuri. T-Move: a light-weight protocol for improved QoS in contentcentric networks with producer mobility. Future Internet, 11(2):28, January 27, 2019. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/11/2/28.

Kavallieratos:2019:TAS

[KCK^{+}19] Georgios Kavallieratos, Nabin Chowdhury, Sokratis Katsikas, Vasileios Gkioulos, and Stephen Wolthusen. Threat analysis for smart homes. Future Internet, 11(10):207, September 25, 2019. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/11/10/207.

Khan:2024:EED

[KCT^{+}24] Muhammad Nafees Ulfat Khan, Weiping Cao, Zhiling Tang, Ata Ullah, and Wanghua Pan.

Energy-efficient De-duplication mechanism for healthcare data aggregation in IoT. Future Internet, 16(2):66, February 19, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
2/66.

Khan:2022:RSR

[KCY22] Hassan Mahmood Khan, Fang-Fang Chua, and Timothy Tzen Vun Yap. ReSQoV: a scalable resource allocation model for QoS-satisfied cloud services. Future Internet, 14 (5):131, April 26, 2022. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
5/131.

Karajeh:2021:CDI

[KDD^{+}21] Ola Karajeh, Dirar Darweesh, Omar Darwish, Noor Abu-El-Rub, Belal Alsinglawi, and Nasser Alsaedi. A classifier to detect informational vs. noninformational heart attack tweets. Future Internet, 13 (1):19, January 16, 2021. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
1/19.

Kenyon:2024:CPE

[KDE24] Anthony Kenyon, Lipika Deka, and David Elizondo. Characterising payload entropy in packet flows — baseline entropy analysis for network anomaly detection. Future Internet, 16(12):470, December 16, 2024. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/16/12/470.

Karabila:2023:ECF

[KDEA^{+}23] Ikram Karabila, Nossayba Darraz, Anas El-Ansari, Nabil Alami, and Mostafa El Mallahi. Enhancing collaborative filteringbased recommender system using sentiment analysis. Future Internet, 15 (7):235–??, July 2023. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/15/
                                        7/235.

Kareem:2024:SEB

[KDG24] Yunus Kareem, Djamel Djenouri, and Essam Ghadafi. A survey on emerging blockchain technology platforms for securing the Internet of Things. Future Internet, 16(8):??, August 8, 2024. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/16/
                                        8/285.

Khatiwoda:2024:CCD

[KDJ24] Naba Raj Khatiwoda, Babu Ram Dawadi, and Sashidhar Ram Joshi. Capacity and coverage dimensioning for 5G standalone mixed-cell architecture: an impact of using existing 4G infrastructure. Future Internet, 16(11): 423, November 14, 2024. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
11/423.

Kitsiou:2022:RUD

[KDKG22] Angeliki Kitsiou, Charikleia Despotidi, Christos Kalloniatis, and Stefanos Gritzalis. The role of users’ demographic and social attributes for accepting biometric systems: a Greek case study. Future Internet, 14(11):328, November 13, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/11/328.

Khalil:2022:LTC

[KE22] Haytham Khalil and Khalid Elgazzar. Location transparency call (LTC) system: an intelligent phone dialing system based on the Phone of Things (PoT) architecture. Future Internet, 14(4):111–??, March 31, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/4/111.

Kalgaonkar:2022:NES

[KES22] Priyank Kalgaonkar and Mohamed El-Sharkawy. NextDet: Efficient sparseto-dense object detection with attentive feature aggregation. Future Internet, 14(12):355, November 28, 2022. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/14/12/355.

Kalgaonkar:2024:NAB

[KES24] Priyank Kalgaonkar and Mohamed El-Sharkawy. NeXtFusion: Attentionbased camera-radar fusion network for improved three-dimensional object detection and tracking. Future Internet, 16(4):114, March 28, 2024. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/16/
                                        4/114.

Kuznetsov:2024:NTE

[KFK^{+}24] Oleksandr Kuznetsov, Emanuele Frontoni, Kateryna Kuznetsova, Ruslan Shevchuk, and Mikolaj Karpinski. NFT technology for enhanced global digital registers: a novel approach to tokenization. Future Internet, 16(7):??, July 17, 2024. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/16/
                                        7/252.

Kolios:2010:LAS

[KFP10] Panayiotis Kolios, Vasilis Friderikos, and Katerina Papadaki. Look-ahead strategies based on storecarry and forward relaying for energy efficient cellular communications. Future Internet, 2(4):587– 602, November 04, 2010. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/2/
4/587.

Khan:2018:RHD

[KG18] Sara Khan and Claudio Germak. Reframing HRI design opportunities for social robots: Lessons learnt from a service robotics case study approach using UX for HRI. Future Internet, 10(10): 101, October 10, 2018. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/10/
10/101.

Karras:2024:TAB

[KGK^{+}24] Aristeidis Karras, Anastasios Giannaros, Christos Karras, Leonidas Theodorakopoulos, Constantinos S. Mammassis, George A. Krimpas, and Spyros Sioutas. TinyML algorithms for big data management in large-scale IoT systems. Future Internet, 16(2):42, January 25, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/2/42.

Kakkavas:2020:RAA

[KGKP20] Grigorios Kakkavas, Despoina Gkatzioura, Vasileios Karyotis, and Symeon Papavassiliou. A review of advanced algebraic approaches enabling network tomography for future network infrastructures. Future Internet, 12(2):20, January 22, 2020. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/12/
                                        2/20.

Kammerer:2021:ASD

[KGRP21] Klaus Kammerer, Manuel Göster, Manfred Reichert, and Rüdiger Pryss. Ambalytics: a scalable and distributed system architecture concept for bibliometric network analyses. Future Internet, 13 (8):203, August 04, 2021. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/13/
                                        8/203.

Kirikkayis:2023:BFM

[KGWR23] Yusuf Kirikkayis, Florian Gallik, Michael Winter, and Manfred Reichert. BPMNE4IoT: a framework for modeling, executing and monitoring IoTdriven processes. Future Internet, 15(3):90, February 22, 2023. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/15/3/90.

Kos:2010:CGS

[KH10] Alexander Kos and Hans-Jürgen Himmler. CWM Global Search — the Internet search engine for chemists and biologists. Future Internet, 2(4):635– 644, December 03, 2010. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/2/
4/635.

Kuai:2019:LBD

[KH19] Meng Kuai and Xiaoyan Hong. Location-based deferred broadcast for adhoc named data networking. Future Internet, 11(6): 139, June 24, 2019. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
6/139.

Khanfor:2024:DAI

[Kha24] Abdullah Khanfor. Determinants to adopt industrial Internet of Things in small and medium-sized enterprises. Future Internet, 16(9), September 20, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/9/340.

Kaddoura:2021:PDD

[KHAA21] Sanaa Kaddoura, Ramzi A.

Haraty, Karam Al Kontar, and Omar Alfandi. A parallelized database damage assessment approach after cyberattack for healthcare systems. Future Internet, 13(4):90, March 31, 2021.

CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/13/
                                        4/90.

Kufakunesu:2024:CAA

[KHAM24] Rachel Kufakunesu, Gerhard P. Hancke, and Adnan M. Abu-Mahfouz. Collision avoidance adaptive data rate algorithm for LoRaWAN. Future Internet, 16(10), October 19, 2024. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/16/10/380.

Kobza:2023:EBE

[KHC^{+}23] Ondrej Kobza, David Herel, Jan Cuhel, Tommaso Gargiani, Jan Pichl, Petr Marek, Jakub Konrad, and Jan Sedivy. Enhancements in BlenderBot 3: Expanding beyond a singular model governance and boosting generational performance. Future Internet, 15(12):384, November 28, 2023. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/15/12/384.

Kobza:2024:ADT

[KHC^{+}24] Ondrej Kobza, David Herel, Jan Cuhel, Tommaso Gargiani, Petr Marek, and Jan Sedivy. Alquist 5.0: Dialogue trees meet generative models, a novel approach for enhancing SocialBot conversations. Future Internet, 16(9), September 21, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
9/344.

Ko:2020:PBN

[KHM20] Kyi Thar Ko, Htet Htet Hlaing, and Masahiro Mambo. A PEKS-based NDN strategy for name privacy. Future Internet, 12(8):130, July 31, 2020. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/12/
8/130.

Kim:2024:VVD

[KHM24] Young Sook Kim, Seng-Phil Hong, and Marko Majer. Validation of valuedriven token economy: Focus on blockchain content platform. Future Internet, 16(5):178, May 20, 2024. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
5/178.

Kinkelin:2011:UTS

[KHN^{+}11] Holger Kinkelin, Ralph Holz, Heiko Niedermayer, Simon Mittelberger, and Georg Carle. On using TPM for secure identities in future home networks. Future Internet, 3(1):1–13, January 07, 2011. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/3/1/1.

Kavrestad:2022:ECG

[KHN^{+}22] Joakim Kävrestad, Allex Hagberg, Marcus Nohlberg, Jana Rambusch, Robert Roos, and Steven Furnell. Evaluation of contextual and game-based training for phishing detection. Future Internet, 14(4):104– ??, March 25, 2022. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/14/
                                        4/104.

Kallempudi:2022:TSS

[KHP^{+}22]  Goutham  Kallempudi,

Khurram Azeem Hashmi, Alain Pagani, Marcus Liwicki, Didier Stricker, and Muhammad Zeshan Afzal. Toward semi-supervised graphical object detection in document images. Future Internet, 14(6):176, June 08, 2022. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/14/
                                        6/176.

Khrais:2020:RAI

[Khr20] Laith T. Khrais. Role of artificial intelligence in shaping consumer demand in e-commerce. Future Internet, 12(12):226, December 08, 2020. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.

mdpi.com/1999-5903/12/
12/226.

Kolsch:2019:SBP

[KHRG19] Johannes Kölsch, Christopher Heinz, Axel Ratzke, and Christoph Grimm. Simulation-based performance validation of homomorphic encryption algorithms in the Internet of Things. Future Internet, 11(10):218, October 22, 2019. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/11/10/218.

Kan:2020:LCP

[KI20] MingSung Kan and Atsushi Ito. Language cognition and pronunciation training using applications. Future Internet, 12 (3):42, February 25, 2020. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/12/
3/42.

Kondratyeva:2022:CDB

[KIB^{+}22] Anastasia Kondratyeva, Daria Ivanova, Vyacheslav Begishev, Ekaterina Markova, Evgeni Mokrov, Yuliya Gaidamaka, and Konstantin Samouylov. Characterization of dynamic blockage probability in industrial millimeter wave 5G deployments. Future Internet, 14(7):193, June 27, 2022. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/14/
                                        7/193.

Kruger:2015:ITE

[KID^{+}15] Dov Kruger, Sarah Inman, Zhiyu Ding, Yijin Kang, Poornima Kuna, Yujie Liu, Xiakun Lu, Stephen Oro, and Yingzhu Wang. Improving teacher effectiveness: Designing better assessment tools in learning management systems. Future Internet, 7(4):484– 499, December 18, 2015. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/7/
                                        4/484.

Khando:2023:ETD

[KIG23] Khando Khando, M. Sirajul Islam, and Shang Gao. The emerging technologies of digital payments and associated challenges: a systematic literature review. Future Internet, 15(1):21, December 30, 2023. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/15/1/21.

Kim:2020:MBS

[KK20] Mihui Kim and Youngmin Kim. Multi-blockchain structure for a crowdsensingbased smart parking system. Future Internet, 12 (5):90, May 16, 2020. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.

mdpi.com/1999-5903/12/
5/90.

Kenyeres:2021:CSD

[KK21] Martin Kenyeres and Jozef Kenyeres. Comparative study of distributed consensus gossip algorithms for network size estimation in multi-agent systems. Future Internet, 13(5):134, May 18, 2021. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
5/134.

Kenyeres:2023:DAC

[KK23a] Martin Kenyeres and Jozef Kenyeres. Distributed average consensus algorithms in d-regular bipartite graphs: Comparative study. Future Internet, 15(5):183, May 16, 2023. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
5/183.

Khoei:2023:MLM

[KK23b] Tala Talaei Khoei and Naima Kaabouch. Machine learning: Models, challenges, and research directions. Future Internet, 15(10):332, October 09, 2023. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/15/10/332.

Kulatunga:2010:ENC

[KKBD10] Chamil Kulatunga, Jesse Kielthy, Dmitri Botvich, and William Donnelly. Exploiting the in-network capabilities of multicast to discover proximate IPTV channels. Future Internet, 2(4):431–445, September 29, 2010. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/2/4/431.

Kochetkova:2023:STM

[KKBG23] Irina Kochetkova, Anna Kushchazli, Sofia Burtseva, and Andrey Gorshenin. Short-term mobile network traffic forecasting using seasonal ARIMA and Holt–Winters models. Future Internet, 15(9):290, August 28, 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
9/290.

Kabamba:2024:VLO

[KKD24] Herve M. Kabamba, Matthew Khouzam, and Michel R. Dagenais. Vnode: Lowoverhead transparent tracing of Node.js-based microservice architectures. Future Internet, 16(1):13, December 29, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
1/13.

Kontogiannis:2017:PFN

[KKEV17] Sotirios Kontogiannis, George Kokkonis, Soultana Ellinidou, and Stavros Valsamidis. Proposed fuzzy-NN algorithm with LoRaCommunication protocol for clustered irrigation systems. Future Internet, 9(4):78, November 07, 2017. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/9/
4/78.

Kavallieratos:2020:CSC

[KKG20] Georgios Kavallieratos, Sokratis Katsikas, and Vasileios Gkioulos. Cybersecurity and safety coengineering of cyberphysical systems — a comprehensive survey. Future Internet, 12(4):65, April 11, 2020. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/12/4/65.

Karpinski:2020:GAT

[KKK^{+}20] Mikolaj Karpinski, Svitlana Kuznichenko, Nadiia Kazakova, Oleksii Fraze-Frazenko, and Daniel Jancarczyk. Geospatial assessment of the territorial road network by fractal method. Future Internet, 12(11):201, November 17, 2020. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/12/11/201.

Kashevnik:2020:HPA

[KKL^{+}20] Alexey Kashevnik, Mikhail Kruglov, Igor Lashkov, Nikolay Teslya, Polina Mikhailova, Evgeny Ripachev, Vladislav Malutin, Nikita Saveliev, and Igor Ryabchikov. Human psychophysiological activity estimation based on smartphone camera and wearable electronics. Future Internet, 12(7):111, July 01, 2020. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/12/
                                        7/111.

Krammer:2022:USI

[KKM^{+}22] Peter Krammer, Marcel Kvassay, Ján Mojzis, Martin Kenyeres, Milos Ockay, Ladislav Hluchý, Lubos Pavlov, and Lubos Skurcák. Using satellite imagery to improve local pollution models for high-voltage transmission lines and insulators. Future Internet, 14(4):99–??, March 23, 2022. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/14/4/99.

Kakkavas:2022:TIL

[KKP22a] Grigorios Kakkavas, Vasileios Karyotis, and Symeon Papavassiliou. Topology inference and link parameter estimation based on end-to-end measurements.

Future Internet, 14(2):45, January 28, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
2/45.

Kallitsis:2022:PED

[KKP22b] Georgios Kallitsis, Vasileios Karyotis, and Symeon Papavassiliou. On the potential of enhancing delaytolerant routing protocols via age of information. Future Internet, 14(8):242, August 17, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
8/242.

Kirova:2019:IMV

[KKS^{+}19] Veronika Kirova, Kirill Karpov, Eduard Siemens, Irina Zander, Oksana Vasylenko, Dmitry Kachan, and Sergii Maksymov. Impact of modern virtualization methods on timing precision and performance of high-speed applications. Future Internet, 11(8):179, August 16, 2019. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
8/179.

Kungne:2020:RBL

[KKT20] Willy Kengne Kungne, Georges-Edouard Kouamou, and Claude Tangha. A rule-based language and verification framework of dynamic service composition. Future Internet, 12 (2):23, January 26, 2020. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/12/
                                        2/23.

Konig:2020:CBS

[KKTK20] Lukas König, Yuliia Korobeinikova, Simon Tjoa, and Peter Kieseberg. Comparing blockchain standards and recommendations. Future Internet, 12(12):222, December 07, 2020. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/12/12/222.

Krouska:2022:PEA

[KKTS22] Akrivi Krouska, Katerina Kabassi, Christos Troussas, and Cleo Sgouropoulou. Personalizing environmental awareness through smartphones using AHP and PROMETHEE II. Future Internet, 14(2):66, February 21, 2022. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/14/
                                        2/66.

Kwon:2021:IRM

[KL21] Junhyung Kwon and Sangkyun Lee. Improving the robustness of model compression by onmanifold adversarial training. Future Internet, 13 (12):300, November 25, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/12/300.

Koulianos:2023:BTS

[KL23] Athanasios Koulianos and Antonios Litke. Blockchain technology for secure communication and formation control in smart drone swarms. Future Internet, 15(10):344, October 19, 2023. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/15/10/344.

Kochetkova:2024:CQS

[KLG^{+}24] Irina Kochetkova, Kseniia Leonteva, Ibram Ghebrial, Anastasiya Vlaskina, Sofia Burtseva, Anna Kushchazli, and Konstantin Samouylov. Controllable queuing system with elastic traffic and signals for resource capacity planning in 5G network slicing. Future Internet, 16(1):18, December 31, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
1/18.

Karyotakis:2019:SPS

[KLKA19] Minos-Athanasios Karyotakis, Evangelos Lamprou, Matina Kiourexidou, and Nikos Antonopoulos. SEO practices: a study about the way news Websites allow the users to comment on their news articles. Future Internet, 11 (9):188, August 30, 2019. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/11/
                                        9/188.

Kapsis:2019:TSI

[KLKP19] Theodore T. Kapsis, Nikolaos K. Lyras, Charilaos I. Kourogiorgas, and Athanasios D. Panagopoulos. Time series irradiance synthesizer for optical GEO satellite downlinks in 5G networks. Future Internet, 11(6):131, June 13, 2019. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/11/6/131.

Kuang:2017:HTA

[KLTC17] Zhufang Kuang, Gongqiang Li, Junshan Tan, and Zhigang Chen. High throughput and acceptance ratio multipath routing algorithm in cognitive wireless mesh network. Future Internet, 9(4):91, November 25, 2017. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/9/4/91.

Kou:2016:TVR

[KLZ16] Weili Kou, Hui Li, and Kailai Zhou. Turning video resource management into cloud computing. Future Internet, 8(3): 35, July 21, 2016. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/8/3/35.

Kamal:2024:AHT

[KM24] Hesham Kamal and Maggie Mashaly. Advanced hybrid transformer-CNN deep learning model for effective intrusion detection systems with class imbalance mitigation using resampling techniques. Future Internet, 16(12):481, December 23, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
12/481.

Kampourakis:2024:SDS

[KMK24] Vyron Kampourakis, Georgios Michail Makrakis, and Constantinos Kolias. From seek-and-destroy to splitand-destroy: Connection partitioning as an effective tool against low-rate DoS attacks. Future Internet, 16(4):137, April 19, 2024. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
4/137.

Kennedy:2012:SAP

[KMS^{+}12] Sean Kennedy, Owen Molloy, Robert Stewart, Paul Jacob, Maria Maleshkova, and Frank Doheny. A semantically automated protocol adapter for mapping SOAP Web services to RESTful HTTP format to enable the Web infrastructure, enhance Web service interoperability and ease Web service migration. Future Internet, 4(2): 372–395, April 11, 2012. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/4/
                                        2/372.

Koryzis:2023:DTP

[KMV^{+}23] Dimitris Koryzis, Dionisis Margaris, Costas Vassilakis, Konstantinos Kotis, and Dimitris Spiliotopoulos. Disruptive technologies for parliaments: a literature review. Future Internet, 15(2):66, February 05, 2023. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/15/2/66.

Kambourakis:2018:SPW

[KMW18] Georgios Kambourakis, Felix Gomez Marmol, and Guojun Wang. Security and privacy in wireless and mobile networks. Future Internet, 10(2):18, February 09, 2018. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/10/2/18.

Kuaban:2024:ITA

[KNN^{+}24] Godlove Suila Kuaban, Valery Nkemeni, Onyeka J. Nwobodo, Piotr Czekalski, and Fabien Mieyeville. Internet of Things adoption in technology ecosystems within the central African region: The case of silicon mountain. Future Internet, 16(10), October 16, 2024. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
10/376.

Kolding:2013:QSP

[KOJP13] Troels Kolding, Pawel Ochal, Niels Terp Kjeldgaard Jørgensen, and Klaus Pedersen. QoS selfprovisioning and interference management for cochannel deployed 3G femtocells. Future Internet, 5 (2):168–189, May 02, 2013. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/5/
2/168.

Kosenkov:2016:CCN

[Kos16] Alexander Kosenkov. Cyber conflicts as a new global threat. Future Internet, 8(3):45, September 09, 2016. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/8/3/45.

Knapp:2015:UEM

[KP15] Herschel Knapp and Sanjog Pangarkar. Utilizing the ECHO model in the veterans health affairs system: Guidelines for setup, operations and preliminary findings. Future Internet, 7(2):184–195, June 08, 2015. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/7/2/184.

Karagiannis:2019:DLS

[KP19] Georgios A. Karagiannis and Athanasios D. Panagopoulos. Dynamic lognormal shadowing framework for the performance evaluation of next generation cellular systems. Future Internet, 11(5): 106, May 02, 2019. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/11/
                                        5/106.

Kabassi:2023:ARE

[KPB23] Katerina Kabassi, Anastasia Papadaki, and Athanasios Botonis. Adapting recommendations on environmental education programs. Future Internet, 15 (1):28, January 04, 2023. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/15/
                                        1/28.

Kim:2019:NIS

[KPC19] Sanghyun Kim, Hyunsun Park, and Moon Jong Choi. Negative impact of social network services based on stressor–stressoutcome: The role of experience of privacy violations. Future Internet, 11(6):137, June 20, 2019.

CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/

6/137. See comment [FC19].

Kljaic:2023:OCC

[KPC^{+}23] Zdenko Kljaić, Danijel Pavković, Mihael Cipek, Maja Trstenjak, Tomislav Josip Mlinarić, and Mladen Niksić. An overview of current challenges and emerging technologies to facilitate increased energy efficiency, safety, and sustainability of railway transport. Future Internet, 15(11):347, October 25, 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
11/347.

Kotenko:2013:DIH

[KPCS13] Igor Kotenko, Olga Polubelova, Andrey Chechulin, and Igor Saenko. Design and implementation of a hybrid ontologicalrelational data repository for SIEM systems. Future Internet, 5(3):355– 375, July 09, 2013. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/5/
3/355.

Kim:2024:TDL

[KPH^{+}24] Haedam Kim, Suhyun Park, Hyemin Hong, Jieun Park, and Seongmin Kim. A transferable deep learning framework for improving the accuracy of Internet of Things intrusion detection. Future Internet, 16(3):80, February 28, 2024. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/16/3/80.

Kontopoulou:2023:RAV

[KPKM23] Vaia I. Kontopoulou, Athanasios D. Panagopoulos, Ioannis Kakkos, and George K. Matsopoulos. A review of ARIMA vs. machine learning approaches for time series forecasting in data driven networks. Future Internet, 15(8):255, July 30, 2023. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
8/255.

Kapsoulis:2020:KYC

[KPP^{+}20a] Nikolaos Kapsoulis, Alexandros Psychas, Georgios Palaiokrassas, Achilleas Marinakis, Antonios Litke, and Theodora Varvarigou. Know your customer (KYC) implementation with smart contracts on a privacy-oriented decentralized architecture. Future Internet, 12(2):41, February 24, 2020. CO-DEN ???? ISSN 1999-

5903. URL https://www.

mdpi.com/1999-5903/12/
2/41.

Kapsoulis:2020:CBS

[KPP^{+}20b] Nikolaos Kapsoulis, Alexandros Psychas, Georgios Palaiokrassas, Achilleas Marinakis, Antonios Litke, Theodora Varvarigou, Charalampos Bouchlis, Amaryllis Raouzaiou, Gonçal Calvo, and Jordi Escudero Subirana. Consortium blockchain smart contracts for musical rights governance in a collective management organizations (CMOs) use case. Future Internet, 12(8):134, August 11, 2020. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/12/
8/134.

Kempelis:2024:CVM

[KPRP24] Arturs Kempelis, Inese Polaka, Andrejs Romanovs, and Antons Patlins. Computer vision and machine learning-based predictive analysis for urban agricultural systems. Future Internet, 16(2):44, January 28, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/2/44.

Kaliyaperumal:2024:NHU

[KPT^{+}24] Prabu Kaliyaperumal, Sudhakar Periyasamy, Manikandan Thirumalaisamy, Balamurugan Balusamy, and Francesco Benedetto. A novel hybrid unsupervised learning approach for enhanced cybersecurity in the IoT. Future Internet, 16(7):??, July 18, 2024. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/16/
                                        7/253.

Kratzke:2020:VHC

[Kra20] Nane Kratzke. Volunteer down: How COVID-19 created the largest idling supercomputer on Earth. Future Internet, 12(6):98, June 06, 2020. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/12/6/98.

Kratzke:2022:CNO

[Kra22] Nane Kratzke. Cloudnative observability: The many-faceted benefits of structured and unified logging—a multi-case study. Future Internet, 14(10): 274, September 26, 2022. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/14/
                                        10/274.

Kurtukova:2022:CCS

[KRSF22] Anna Kurtukova, Aleksandr Romanov, Alexander Shelupanov, and Anastasia Fedotova. Complex cases of source code authorship identification using a hybrid deep neural network. Future Internet, 14(10):287, September 30, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/10/287.

Khazane:2024:HRM

[KRSK24] Hassan Khazane, Mohammed Ridouani, Fatima Salahdine, and Naima Kaabouch. A holistic review of machine learning adversarial attacks in IoT networks. Future Internet, 16(1):32, January 19, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/1/32.

Kaschesky:2014:DVF

[KS14] Michael Kaschesky and Luigi Selmi. 7R data value framework for open data in practice: Fusepool. Future Internet, 6(3):556– 583, September 08, 2014. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/6/
3/556.

Kejriwal:2019:MPI

[KS19] Mayank Kejriwal and Pedro Szekely. myDIG: Personalized illicit domainspecific knowledge discovery with no programming. Future Internet, 11(3):59, March 04, 2019. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/11/3/59.

Kabashkin:2024:AIT

[KS24] Igor Kabashkin and Leonid Shoshin. Artificial intelligence of things as new paradigm in aviation health monitoring systems. Future Internet, 16 (8):??, August 2, 2024. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/16/
                                        8/276.

Khan:2021:SCL

[KSA^{+}21] Shawal Khan, Ishita Sharma, Mazzamal Aslam, Muhammad Zahid Khan, and Shahzad Khan. Security challenges of location privacy in VANETs and state-of-the-art solutions: a survey. Future Internet, 13(4):96, April 10, 2021. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/13/
                                        4/96.

Kalafatidis:2023:LCS

[KSD^{+}23] Sarantis Kalafatidis, Sotiris Skaperas, Vassilis Demiroglou, Lefteris Mamatas, and Vassilis Tsaoussidis. Logicallycentralized SDN-based NDN strategies for wireless mesh smart-city networks. Future Internet, 15 (1):19, December 29, 2023. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/15/
                                        1/19.

Kousaridas:2020:CBO

[KSE^{+}20] Apostolos Kousaridas, Andreas Schimpe, Sebastian Euler, Xavier Vilajosana, Mikael Fallgren, Giada Landi, Francesca Moscatelli, Sokratis Barmpounakis, Francisco Vázquez-Gallego, Roshan Sedar, Rodrigo Silva, Laurent Dizambourg, Stefan Wendt, Maciej Muehleisen, Kurt Eckert, Jérôme Härri, and Jesus Alonso-Zarate. 5G cross-border operation for connected and automated mobility: Challenges and solutions. Future Internet, 12(1):5, December 24, 2020. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/12/1/5.

Kamruzzaman:2024:MLB

[KSG24a] Md Kamruzzaman, Nurul I. Sarkar, and Jairo Gutierrez. Machine learning-based resource allocation algorithm to mitigate interference in D2D-Enabled cellular networks. Future Internet, 16(11): 408, November 6, 2024. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
11/408.

Kumar:2024:ESC

[KSG^{+}24b] Pradeep Kumar, Guo-Liang Shih, Bo-Lin Guo, Siva Kumar Nagi, Yibeltal Chanie Manie, Cheng-Kai Yao, Michael Augustine Arockiyadoss, and Peng-Chun Peng. Enhancing smart city safety and utilizing AI expert systems for violence detection. Future Internet, 16 (2):50, January 31, 2024. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/16/
                                        2/50.

Kuantama:2023:FWB

[KSJZ23] Endrowednes Kuantama, Avishkar Seth, Alice James, and Yihao Zhang. Flying watchdog-based guard patrol with check point data verification. Future Internet, 15(10):340, October 16, 2023. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/15/
                                        10/340.

Kohana:2019:WBN

[KSO19] Masaki Kohana, Shinji Sakamoto, and Shusuke Okamoto. Web browser network based on a BA model for a Web-based virtual world. Future Internet, 11(7):147, July 05, 2019. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/11/7/147.

Konsgen:2019:MME

[KSSF19] Andreas Könsgen, Md. Shahabuddin, Amanpreet Singh, and Anna Förster. A mathematical model for efficient and fair resource assignment in multipath transport. Future Internet, 11(2):39, February 10, 2019. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/11/2/39.

Kapetanidou:2019:RBT

[KST19] Ioanna Angeliki Kapetanidou, Christos-Alexandros Sarros, and Vassilis Tsaoussidis. Reputation-based trust approaches in named data networking. Future Internet, 11(11):241, November 18, 2019. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
11/241.

Koukis:2024:AAD

[KST24] Georgios Koukis, Konstantina Safouri, and Vassilis Tsaoussidis. All about delay-tolerant networking (DTN) contributions to future Internet. Future Internet, 16(4):129, April 09, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/4/129.

Kieseberg:2021:SID

[KSW21] Peter Kieseberg, Sebastian Schrittwieser, and Edgar Weippl. Secure internal data markets. Future Internet, 13(8):208, August 12, 2021. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/13/8/208.

Kapassa:2022:BTA

[KT22] Evgenia Kapassa and Marinos Themistocleous. Blockchain technology applied in IoV demand response management: a systematic literature review. Future Internet, 14 (5):136, April 29, 2022. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/14/
                                        5/136.

Khan:2023:RDP

[KTAH^{+}23] Imran Moez Khan, Andrew Thompson, Akram Al-Hourani, Kandeepan Sithamparanathan, and Wayne S. T. Rowe. RSSI and device pose fusion for fingerprinting-based indoor smartphone localization systems. Future Internet, 15(6):220–??, June 2023. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/15/6/220.

Kapassa:2021:BAI

[KTCI21] Evgenia Kapassa, Marinos Themistocleous, Klitos Christodoulou, and Elias Iosif. Blockchain application in Internet of Vehicles: Challenges, contributions and current limitations. Future Internet, 13(12):313, December 10, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/12/313.

Kubba:2024:ELR

[KTD24] Abbas Kubba, Hafedh Trabelsi, and Faouzi Derbel. Enhanced long-range network performance of an oil pipeline monitoring system using a hybrid deep extreme learning machine model. Future Internet, 16(11):425, November 17, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
11/425.

Katsinis:2017:MIM

[KTP17] Georgios Katsinis, Eirini Eleni Tsiropoulou, and Symeon Papavassiliou. Multicell interference management in device to device underlay cellular networks. Future Internet, 9(3):44, August 07, 2017. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/9/3/44.

Karembai:2018:TPI

[KTS18] Anil Kumar Karembai, Jeffrey Thompson, and Patrick Seeling. Towards prediction of immersive virtual reality image quality of experience and quality of service. Future Internet, 10(7):63, July 07, 2018. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/10/7/63.

Kuppusamy:2019:EFP

[KTUF19] Vishnupriya Kuppusamy, Udaya Miriya Thanthrige, Asanga Udugama, and Anna Förster. Evaluating forwarding protocols in opportunistic networks: Trends, advances, challenges and best practices. Future Internet, 11 (5):113, May 11, 2019. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
5/113.

Kianersi:2022:ERM

[KUBS22] Darius Kianersi, Suraj Uppalapati, Anirudh Bansal, and Jeremy Straub. Evaluation of a reputation management technique for autonomous vehicles. Future Internet, 14(2):31, February 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/2/31.

Kyriakos:2023:BIT

[KV23] Christos Kyriakos and Manolis Vavalis. Business intelligence through machine learning from satellite remote sensing data. Future Internet, 15(11): 355, October 27, 2023. CODEN ???? ISSN 1999-

5903. URL https://www.

mdpi.com/1999-5903/15/
11/355.

Kouretsis:2022:MAK

[KVL^{+}22] Alexandros Kouretsis, Iraklis Varlamis, Laida Limniati, Minas Pergantis, and Andreas Giannakoulopoulos. Mapping art to a knowledge graph: Using data for exploring the relations among visual objects in Renaissance art. Future Internet, 14(7):206, July 03, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/7/206.

Kapadais:2019:FDS

[KVST19] Konstantinos Kapadais, Iraklis Varlamis, Christos Sardianos, and Konstantinos Tserpes. A framework for the detection of search and rescue patterns using shapelet classification. Future Internet, 11(9):192, September 04, 2019. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
9/192.

Kotenidis:2022:ICM

[KVVD22] Efthimis Kotenidis, Nikolaos Vryzas, Andreas Veglis, and Charalampos Dimoulas. Integrating chatbot media automations in professional journalism: an evaluation framework. Future Internet, 14(11): 343, November 21, 2022.

CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/14/
                                        11/343.

Kalyani:2024:ASD

[KVWC24] Yogeswaranathan Kalyani, Liam Vorster, Rebecca Whetton, and Rem Collier. Application scenarios of digital twins for smart crop farming through cloud–fog–edge infrastructure. Future Internet, 16(3):100, March 16, 2024. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/16/3/100.

Kato:2019:STE

[KWI^{+}19] Shigeru Kato, Naoki Wada, Ryuji Ito, Takaya Shiozaki, Yudai Nishiyama, and Tomomichi Kagawa. Snack texture estimation system using a simple equipment and neural network model. Future Internet, 11(3):68, March 08, 2019. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/11/3/68.

Kim:2024:RBT

[KXZ^{+}24] Heejoo Kim, Zhe Xiao, Xiaocai Zhang, Xiuju Fu, and Zheng Qin. Rethinking blockchain technologies for the maritime industry: an overview of the current landscape. Future Internet, 16(12):454, December 3, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
12/454.

Kankanamge:2022:GCE

[KYG22] Nayomi Kankanamge, Tan Yigitcanlar, and Ashantha Goonetilleke. Gamifying community education for enhanced disaster resilience: an effectiveness testing study from Australia. Future Internet, 14(6):179, June 09, 2022. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
6/179.

Khalil:2021:IDI

[KYS21] Samar Samir Khalil, Sherin M.

Youssef, and Sherine Nagy Saleh. iCaps-Dfake: an integrated capsule-based model for deepfake image and video detection. Future Internet, 13(4):93, April 05, 2021. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/13/4/93.

Keller:2023:BIU

[KYT^{+}23] Damián Keller, Azeema Yaseen, Joseph Timoney, Sutirtha Chakraborty, and Victor Lazzarini. Banging interaction: a ubimusdesign strategy for the musical Internet. Future Internet, 15(4):125, March 27, 2023. CODEN ????

ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/15/4/125.

Krol:2020:AIW

[KZ20] Karol Król and Dariusz Zdonek. Aggregated indices in Website quality assessment. Future Internet, 12(4):72, April 17, 2020. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/12/
                                        4/72.

Khalil:2024:SON

[KZ24] Ayman Khalil and Besma Zeddini. A secure opportunistic network with efficient routing for enhanced efficiency and sustainability. Future Internet, 16(2):56, February 08, 2024. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/16/2/56.

Kadusic:2022:SPS

[KZRH22] Esad Kadusic, Natasa Zivic, Christoph Ruland, and Narcisa Hadzajlic. A smart parking solution by integrating NB-IoT radio communication technology into the core IoT platform. Future Internet, 14(8):219, July 25, 2022. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/14/
                                        8/219.

Lousado:2020:MSE

[LA20] José Paulo Lousado and Sandra Antunes. Monitoring and support for elderly people using LoRa communication technologies: IoT concepts and applications. Future Internet, 12(11):206, November 20, 2020. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/12/11/206.

Latifinavid:2023:DVB

[LA23] Masoud Latifinavid and Aydin Azizi. Development of a vision-based unmanned ground vehicle for mapping and tennis ball collection: a fuzzy logic approach. Future Internet, 15(2):84, February 19, 2023. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/15/2/84.

Lettieri:2018:EMA

[LAG^{+}18] Nicola Lettieri, Antonio Altamura, Rosalba Giugno, Alfonso Guarino, Delfina Malandrino, Alfredo Pulvirenti, Francesco Vicidomini, and Rocco Zaccagnino. Ex machina: Analytical platforms, law and the challenges of computational legal science. Future Internet, 10(5):37, April 26, 2018. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/10/5/37.

Lai:2024:ASR

[Lai24] Joel Weijia Lai. Adapting self-regulated learning in an age of generative artificial intelligence chatbots. Future Internet, 16 (6):218, June 20, 2024. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/16/
                                        6/218.

Lallie:2014:PCM

[Lal14] Harjinder Singh Lallie. The problems and challenges of managing crowd sourced audio-visual evidence. Future Internet, 6(2):190–202, April 01, 2014. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/6/2/190.

Lopez-Ardao:2023:RAI

[LARPHA23] José Carlos López-Ardao, Miguel Rodríguez-Pérez, and Sergio Herrería-Alonso. Recent advances in informationcentric networks (ICNs). Future Internet, 15(12): 392, December 01, 2023. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/15/
                                        12/392.

Laurini:2015:GOG

[Lau15]  Robert Laurini.  Geo-

graphic ontologies, gazetteers and multilingualism. Future Internet, 7(1):1–23, January 05, 2015. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/7/1/1.

Liccardo:2023:VAD

[LB23] Annalisa Liccardo and Francesco Bonavolontà. VR, AR, and 3-D user interfaces for measurement and control. Future Internet, 15(1):18, December 29, 2023. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/15/1/18.

Lazzaro:2024:AAD

[LB24] Sara Lazzaro and Francesco Buccafurri. Achieving accountability and data integrity in message queuing telemetry transport using blockchain and interplanetary file system. Future Internet, 16(7):??, July 13, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/7/246.

Lawo:2024:WFB

[LBA^{+}24] Daniel Christian Lawo, Rana Abu Bakar, Abraham Cano Aguilera, Filippo Cugini, José Luis Imaña, Idelfonso Tafur Monroy, and Juan Jose Vegas Olmos. Wireless and fiber-based post-quantum-cryptography-secured IPsec tunnel. Future Internet, 16(8):??, August 21, 2024. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/16/
                                        8/300.

Liaqat:2023:DCP

[LBB23] Rao Muzamal Liaqat, Philip Branch, and Jason But. Design considerations and performance evaluation of gossip routing in LoRa-based linear networks. Future Internet, 15(11):366, November 11, 2023. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/15/11/366.

Lopez-Bernal:2021:ETB

[LBBPM21] Diego Lopez-Bernal, David Balderas, Pedro Ponce, and Arturo Molina. Education 4.0: Teaching the basics of KNN, LDA and simple perceptron algorithms for binary classification problems. Future Internet, 13(8):193, July 27, 2021. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/13/8/193.

Loreti:2018:PAB

[LBC18] Pierpaolo Loreti, Lorenzo Bracciale, and Alberto Caponi. Push attack: Binding virtual and real identities using mobile push notifications. Future Internet, 10(2):13, January 31, 2018. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/10/2/13.

Latvakoski:2014:THA

[LBG^{+}14] Juhani Latvakoski, Mahdi Ben Alaya, Herve Ganem, Bashar Jubeh, Antti Iivari, Jeremie Leguay, Jaume Martin Bosch, and Niclas Granqvist. Towards horizontal architecture for autonomic M2M service networks. Future Internet, 6(2):261–301, May 06, 2014. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/6/2/261.

Lucena:2022:CUU

[LBK22] Felipe Lucena, Fabio Marcelo Breunig, and Hermann Kux. The combined use of UAV-based RGB and DEM images for the detection and delineation of orange tree crowns with Mask R-CNN: an approach of labeling and unified framework. Future Internet, 14(10):275, September 27, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
10/275.

LaQuatra:2023:BIE

[LC23] Moreno La Quatra and Luca Cagliero. BART-IT: an efficient sequenceto-sequence model for Italian text summarization.

Future Internet, 15(1):15, December 27, 2023. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/15/
                                        1/15.

Lai:2024:PFL

[LCC^{+}24] Ying-Hsun Lai, Shin-Yeh Chen, Wen-Chi Chou, Hua-Yang Hsu, and Han-Chieh Chao. Personalized federated learning with adaptive feature extraction and category prediction in non-IID datasets. Future Internet, 16(3):95, March 11, 2024. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/16/3/95.

Liu:2024:IIH

[LCCC24] Jui-Chuan Liu, Heng-Xiao Chi, Ching-Chun Chang, and Chin-Chen Chang. An innovative information hiding scheme based on block-wise pixel reordering. Future Internet, 16 (1):34, January 22, 2024. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/16/
                                        1/34.

Lopez:2017:KTC

[LCMV17] Lorena Isabel Barona López, Ángel Leonardo Valdivieso Caraguay, Marco Antonio Sotelo Monge, and Luis Javier García Villalba. Key technologies in the context of future networks: Operational and management requirements. Future Internet, 9(1):1, December 22, 2017. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/9/1/1.

Lord:2012:TSH

[LCS12] Phillip Lord, Simon Cockell, and Robert Stevens. Three steps to heaven: Semantic publishing in a real world workflow. Future Internet, 4(4):1004–1015, November 08, 2012. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/4/
4/1004.

Lopez:2017:TIM

[LCV^{+}17] Lorena Isabel Barona López, Ángel Leonardo Valdivieso Caraguay, Jorge Maestre Vidal, Marco Antonio Sotelo Monge, and Luis Javier García Villalba. Towards incidence management in 5G based on situational awareness. Future Internet, 9(1):3, January 17, 2017. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/9/1/3.

Liu:2018:PFE

[LCW^{+}18] Kanghuai Liu, Zhigang Chen, Jia Wu, Yutong Xiao, and Heng Zhang. Predict and forward: An efficient routing-delivery scheme based on node profile in opportunistic networks. Future Internet, 10 (8):74, August 06, 2018. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/10/
8/74.

Lockett:2023:IIM

[LCY^{+}23] Adam Lockett, Ioannis Chalkias, Cagatay Yucel, Jane Henriksen-Bulmer, and Vasilis Katos. Investigating IPTV malware in the wild. Future Internet, 15(10):325, September 28, 2023. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/15/10/325.

LeuchAlencar:2024:DFI

[LDd^{+}24] Ancilon Leuch Alencar, Marcelo Dornbusch Lopes, Anita Maria da Rocha Fernandes, Julio Cesar Santos dos Anjos, Juan Francisco De Paz Santana, and Valderi Reis Quietinho Leithardt. Detection of forged images using a combination of passive methods based on neural networks. Future Internet, 16(3):97, March 14, 2024. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
3/97.

Leech:2022:FMB

[LDM22] Sonya Leech, Jonathan Dunne, and David Malone. A framework to model bursty electronic data interchange messages for queueing systems. Future Internet, 14(5):149, May 12, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/5/149.

Liang:2023:RAB

[LDZ^{+}23] Haoming Liang, Jinze Du, Hongchen Zhang, Bing Han, and Yan Ma. Relational action bank with semantic–visual attention for few-shot action recognition. Future Internet, 15 (3):101, March 03, 2023. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
3/101.

Li:2024:PPA

[LDZP24] Lihui Li, Hanwen Deng, Zhongyi Zhai, and Sheng-Lung Peng. Privacypreserving authentication based on PUF for VANETs. Future Internet, 16(9), September 8, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
9/326.

Lee:2017:DPD

[Lee17a] Il-Gu Lee. Digital predistortion of carrier frequency offset for reliable wi-fi enabled IoTs. Future Internet, 9(3):46, August 09, 2017. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/9/3/46.

Lee:2017:IAO

[Lee17b] Il-Gu Lee. Interferenceaware opportunistic dynamic energy saving mechanism for wi-fi enabled IoTs. Future Internet, 9 (3):38, July 18, 2017. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/9/
                                        3/38.

Lee:2018:SIF

[Lee18] Il-Gu Lee. Secure interframe space communications for wireless LANs. Future Internet, 10(6):47, June 04, 2018. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/10/6/47.

Lee:2020:ITI

[Lee20] In Lee. Internet of Things (IoT) cybersecurity: Literature review and IoT cyber risk management. Future Internet, 12(9):157, September 18, 2020. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/12/
                                        9/157.

Lelli:2019:ITI

[Lel19] Francesco Lelli. Interoperability of the time of Industry 4.0 and the Internet of Things. Future Internet, 11(2):36, February 03, 2019. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/11/2/36.

Lerman:2016:IVO

[Ler16] Kristina Lerman. Information is not a virus, and other consequences of human cognitive limits. Future Internet, 8(2): 21, May 13, 2016. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/8/2/21.

Lettieri:2016:CSS

[Let16] Nicola Lettieri. Computational social science, the evolution of policy design and rule making in smart societies. Future Internet, 8(2):19, May 12, 2016. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/8/
2/19.

Lischke:2016:ABN

[LF16] Matthias Lischke and Benjamin Fabian. Analyzing the Bitcoin network: The first four years. Future Internet, 8(1):7, March 07, 2016. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/8/1/7.

Liu:2023:MSA

[LF23] Fan Liu and Jiandong Fang. Multi-scale audio spectrogram transformer for classroom teaching interaction recognition. Future Internet, 15(2):65, February 02, 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
2/65.

Li:2024:SEE

[LF24] Yushan Li and Satoshi Fujita. A synergistic Elixir-EDA-MQTT framework for advanced smart transportation systems. Future Internet, 16(3):81, February 28, 2024. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/16/3/81.

Lynn:2020:UDF

[LFGR20] Theo Lynn, Grace Fox, Anna Gourinovitch, and Pierangelo Rosati. Understanding the determinants and future challenges of cloud computing adoption for high performance computing. Future Internet, 12 (8):135, August 11, 2020. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/12/
8/135.

Lombardo:2019:MAA

[LFM^{+}19]  Gianfranco  Lombardo,

Paolo Fornacciari, Monica Mordonini, Michele Tomaiuolo, and Agostino Poggi. A multi-agent architecture for data analysis. Future Internet, 11 (2):49, February 18, 2019. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
2/49.

Li:2020:CCM

[LG20] KeDi Li and Ning Gui.

CMS: A continuous machinelearning and serving platform for industrial big data. Future Internet, 12 (6):102, June 10, 2020. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/12/
6/102.

Lymperis:2023:SPS

[LG23] Dimitrios Lymperis and Christos Goumopoulos. SEDIA: A platform for semantically enriched IoT data integration and development of smart city applications. Future Internet, 15(8):276, August 18, 2023. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/15/8/276.

Loor:2024:MSI

[LGCM24] Fernando Loor, Veronica Gil-Costa, and Mauricio Marin. Metric space indices for dynamic optimization in a peer to peer-based image classification crowdsourcing platform. Future Internet, 16 (6):202, June 6, 2024. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/16/
                                        6/202.

Lettieri:2019:PET

[LGMZ19] Nicola Lettieri, Alfonso Guarino, Delfina Malandrino, and Rocco Zaccagnino. Platform economy and techno-regulation-experimenting with reputation and nudge. Future Internet, 11(7):163, July 23, 2019. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/11/7/163.

Loukas:2013:RCT

[LGV13] George Loukas, Diane Gan, and Tuan Vuong. A review of cyber threats and defence approaches in emergency management. Future Internet, 5(2):205– 236, May 07, 2013. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/5/
                                        2/205.

Liu:2023:DSP

[LGY23] Yunchuan Liu, Amir Ghasemkhani, and Lei Yang. Drifting streaming peaks-over-threshold-enhanced self-evolving neural networks for short-term wind farm generation forecast. Future Internet, 15 (1):17, December 28, 2023. CODEN ???? ISSN 1999-

                                        5903. URL https://www.

mdpi.com/1999-5903/15/
1/17.

Latvakoski:2019:TCH

[LH19] Juhani Latvakoski and Jouni Heikkinen. A trustworthy communication hub for cyber-physical systems. Future Internet, 11(10):211, October 08, 2019. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/11/10/211.

Lei:2021:VCB

[LH21a] Zhou Lei and Yiyong Huang. Video captioning based on channel soft attention and semantic reconstructor. Future Internet, 13(2):55, February 23, 2021. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/13/2/55.

Lepasepp:2021:SLR

[LH21b] Tuuli Katarina Lepasepp and William Hurst. A systematic literature review of Industry 4.0 technologies within medical device manufacturing. Future Internet, 13(10):264, October 13, 2021. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
10/264.

Liu:2021:HCS

[LH21c] Hongyu Liu and Rui Han.

A hierarchical cache size allocation scheme based on content dissemination in information-centric networks. Future Internet, 13(5):131, May 15, 2021. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
5/131.

Li:2024:EEP

[LHD^{+}24] Xiu Li, Aron Henriksson, Martin Duneld, Jalal Nouri, and Yongchao Wu. Evaluating embeddings from pre-trained language models and knowledge graphs for educational content recommendation. Future Internet, 16(1):12, December 29, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
1/12.

Liu:2024:TTD

[LHL24a] Jing Liu, Xuesong Hai, and Keqin Li. TDLearning: Trusted distributed collaborative learning based on blockchain smart contracts. Future Internet, 16 (1):6, December 25, 2024. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
1/6.

Liu:2024:UUB

[LHL24b] Qiang Liu, Rui Han, and Yang Li. Utilizing user bandwidth resources in information-centric networking through blockchainbased incentive mechanism. Future Internet, 16 (1):11, December 28, 2024. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
1/11.

Loskyll:2012:CBO

[LHSS12] Matthias Loskyll, Ines Heck, Jochen Schlick, and Michael Schwarz. Contextbased orchestration for control of resource-efficient manufacturing processes. Future Internet, 4(3):737– 761, August 14, 2012. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/4/
3/737.

Liu:2018:HMS

[LHW18] Jianqiang Liu, Shuai Huo, and Yi Wang. A hierarchical mapping system for flat identifier to locator resolution based on active degree. Future Internet, 10(8):75, August 08, 2018. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/10/8/75.

Liu:2024:RBM

[LHW24] Yuqi Liu, Rui Han, and Xu Wang. A reordering buffer management method at edge gateway in hybrid IP-ICN multipath transmission system. Future Internet, 16(12):464, December 11, 2024. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/16/
                                        12/464.

Li:2018:IPI

[Li18] YangQun Li. An integrated platform for the Internet of Things based on an open source ecosystem. Future Internet, 10(11): 105, October 31, 2018. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/10/
                                        11/105.

Li:2023:VIA

[Li23] Qiang Li. A V2V identity authentication and key agreement scheme based on identity-based cryptograph. Future Internet, 15 (1):25, January 03, 2023. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/15/
                                        1/25.

Luczak:2023:EMC

[LIM23] Lukasz Piotr Luczak, Przemysław Ignaciuk, and Michał Morawski. Evaluating MPTCP congestion control algorithms: Implications for streaming in open Internet. Future Internet, 15(10):328, October 04, 2023. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/15/10/328.

Li:2023:LGB

[LJR23] Zuopeng Li, Hengshuai Ju, and Zepeng Ren. A learning game-based approach to task-dependent edge resource allocation. Future Internet, 15(12):395, December 07, 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
12/395.

Li:2024:CLA

[LJR24] Zuopeng Li, Hengshuai Ju, and Zepeng Ren. Correction: Li et al. A Learning Game-Based Approach to Task-Dependent Edge Resource Allocation. Future Internet 2023, 15,

395. Future Internet, 16 (4):141, April 22, 2024. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
4/141.

Lee:2023:VSS

[LJWX23] Uddom Lee, Peng Jiang, Hongyi Wu, and Chunsheng Xin. View synthesis with scene recognition for cross-view image localization. Future Internet, 15(4):126, March 28, 2023. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
4/126.

Lohr:2020:IXH

[LK20] Christophe Lohr and Jérôme Kerdreux. Improvements of the xAAL home automation system. Future Internet, 12(6):104, June 12, 2020. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/12/
                                        6/104.

Leivadeas:2019:VPO

[LKIL19] Aris Leivadeas, George Kesidis, Mohamed Ibnkahla, and Ioannis Lambadaris. VNF placement optimization at the edge and cloud. Future Internet, 11(3):69, March 09, 2019. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/11/3/69.

Li:2020:TAT

[LL20a] Haiyan Li and Hongtao Lu. AT-Text: Assembling text components for efficient dense scene text detection. Future Internet, 12(11):200, November 17, 2020. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/12/11/200.

Lo:2020:IEL

[LL20b] Chih-Fong Lo and Chin-Huang Lin. The impact of English learning motivation and attitude on well-being: Cram school students in Taiwan. Future Internet, 12(8):131, August 06, 2020. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/12/
8/131.

Lorenzo-Lledo:2020:DVQ

[LL20c] Alejandro Lorenzo-Lledó.

Design and validation of a questionnaire to measure future Spanish teachers’ perceptions of cinema in pre-school and primary education: Towards active and technological learning. Future Internet, 12(9):149, September 03, 2020. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/12/
9/149.

Lehniger:2023:TWE

[LL23a] Kai Lehniger and Peter Langendörfer. Through the window: Exploitation and countermeasures of the ESP32 register window overflow. Future Internet, 15(6):217–??, June 2023. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
6/217.

Li:2023:CSR

[LL23b] Yanxia Li and Yang Li.

Collaborative storage and resolution method between layers in hierarchical ICN name resolution systems. Future Internet, 15 (2):74, February 13, 2023. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/15/
                                        2/74.

Loh:2024:THS

[LLB24] Peter K. K. Loh, Aloysius Z. Y. Lee, and Vivek Balachandran. Towards a hybrid security framework for phishing awareness education and defense. Future Internet, 16 (3):86, March 01, 2024. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/16/
                                        3/86.

Lodovisi:2018:PAH

[LLBB18] Chiara Lodovisi, Pierpaolo Loreti, Lorenzo Bracciale, and Silvello Betti. Performance analysis of hybrid optical-acoustic AUV swarms for marine monitoring. Future Internet, 10 (7):65, July 10, 2018. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/10/
                                        7/65.

Lin:2024:IRS

[LLCC24] Yijie Lin, Jui-Chuan Liu, Ching-Chun Chang, and Chin-Chen Chang. An innovative recompression scheme for VQ index tables. Future Internet, 16 (8):??, August 19, 2024. CODEN ???? ISSN 1999-

                                        5903. URL https://www.

mdpi.com/1999-5903/16/
8/297.

Li:2023:CTA

[LLCL23] Yufeng Li, Mengxiao Liu, Chenhong Cao, and Jiangtao Li. Communicationtraffic-assisted mining and exploitation of buffer overflow vulnerabilities in ADASs. Future Internet, 15(5):185, May 18, 2023. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
5/185.

Lee:2020:TBS

[LLHW20] Yu-Tse Lee, Jhan-Jia Lin, Jane Yung-Jen Hsu, and Ja-Ling Wu. A time bank system design on the basis of hyperledger fabric blockchain. Future Internet, 12(5):84, May 08, 2020. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/12/5/84.

Liu:2016:OBR

[LLJ16] Xin Liu, Zhongfu Li, and Shaohua Jiang. Ontologybased representation and reasoning in building construction cost estimation in China. Future Internet, 8(3):39, August 03, 2016. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/8/
3/39.

Li:2022:BBB

[LLJ22] Xinlu Li, Yuanyuan Lei, and Shengwei Ji. BERTand BiLSTM-based sentiment analysis of online Chinese buzzwords. Future Internet, 14(11):332, November 14, 2022. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/14/
                                        11/332.

Liu:2021:RTO

[LLL21] Jun Liu, Xiaohui Lian, and Chang Liu. Research on task-oriented computation offloading decision in space–air–ground integrated network. Future Internet, 13(5):128, May 13, 2021. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/13/5/128.

Li:2021:MAS

[LLLD21] Shaoyong Li, Liang Lv, Xiaoya Li, and Zhaoyun Ding. Mobile app startup prediction based on federated learning and attributed heterogeneous network embedding. Future Internet, 13(10):256, October 07, 2021. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/13/
                                        10/256.

Lee:2019:VPD

[LLR19] Jae-Gil Lee, Kwan Min Lee, and Seoung-Ho Ryu. Vehicle politeness in driving situations. Future Internet, 11(2):48, February 16, 2019. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/11/2/48.

Liu:2024:TAH

[LLSS24] Haibo Liu, Yang Liao, Changting Shi, and Jing Shen. Task allocation of heterogeneous multiunmanned systems based on improved sheep flock optimization algorithm. Future Internet, 16(4):124, April 07, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
4/124.

Li:2017:SCT

[LLW^{+}17] Dongyuan Li, Chengshuai Li, Zidong Wang, Deqiang Wang, Jianping Xing, and Bo Zhang. Signal consensus in TSP of the same grid in road network. Future Internet, 9(4):69, October 24, 2017. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/9/4/69.

Lin:2021:ETB

[LLW21] Jhan-Jia Lin, Yu-Tse Lee, and Ja-Ling Wu. The effect of thickness-based dynamic matching mechanism on a hyperledger fabric-based TimeBank system. Future Internet, 13(3):65, March 06, 2021. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/13/
                                        3/65.

Liu:2024:JBF

[LLW^{+}24] Lidong Liu, Shidang Li, Mingsheng Wei, Jinsong Xu, and Bencheng Yu. Joint beam-forming optimization for active-RIS-assisted Internet-of-Things networks with SWIPT. Future Internet, 16(1):20, January 06, 2024. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/16/
                                        1/20.

Li:2023:FDQ

[LLY23] Meng Li, Jiqiang Liu, and Yeping Yang. Financial data quality evaluation method based on multiple linear regression. Future Internet, 15(10):338, October 14, 2023. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/15/
                                        10/338.

Li:2024:AIS

[LLY24] Meng Li, Jiqiang Liu, and Yeping Yang. Automated identification of sensitive financial data based on the topic analysis. Future Internet, 16(2):55, February 08, 2024. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/16/2/55.

Li:2022:CFM

[LLYL22] Yaning Li, Hongsheng Li, Baoguo Yu, and Jun Li. A CSI fingerprint method for indoor pseudolite positioning based on RT-ANN. Future Internet, 14(8):235, July 29, 2022. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
8/235.

Liang:2021:END

[LLZ21] Kun Liang, Jingjing Liu, and Yiying Zhang. The effects of non-directional online behavior on students’ learning performance: a user profile based analysis method. Future Internet, 13(8):199, July 31, 2021. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
8/199.

Li:2019:IAT

[LLZL19] Wenkuan Li, Peiyu Liu, Qiuyue Zhang, and Wenfeng Liu. An improved approach for text sentiment classification based on a deep neural network via a sentiment attention mechanism. Future Internet, 11 (4):96, April 11, 2019. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
4/96.

Li:2023:FSR

[LMF23] Jinying Li, Ananda Maiti, and Jiangang Fei. Features and scope of regulatory technologies: Challenges and opportunities with Industrial Internet of Things. Future Internet, 15(8):256, July 30, 2023. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/15/
                                        8/256.

Li:2019:SPB

[LMH19] Yanli Li, Wendan Ma, and Yue Han. A spatial prediction-based motioncompensated frame rate up-conversion. Future Internet, 11(2):26, January 23, 2019. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/11/2/26.

Li:2018:IIS

[LMK18] Wenjuan Li, Weizhi Meng, and Lam For Kwok. Investigating the influence of special on-off attacks on challenge-based collaborative intrusion detection networks. Future Internet, 10(1):6, January 08, 2018. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/10/
                                        1/6.

Li:2018:UNL

[LMLW18] Yanli Li, Lala Mei, Ran Li, and Changan Wu. Using noise level to detect frame repetition forgery in video frame rate up-conversion. Future Internet, 10(9):84, August 24, 2018. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/10/9/84.

Liao:2019:CSA

[LMPT19] Fei Liao, Liangli Ma, Jingjing Pei, and Linshan Tan. Combined self-attention mechanism for Chinese named entity recognition in military. Future Internet, 11(8):180, August 18, 2019. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
8/180.

Leotta:2019:SHH

[LMSC19] Francesco Leotta, Massimo Mecella, Daniele Sora, and Tiziana Catarci. Surveying human habit modeling and mining techniques in smart spaces. Future Internet, 11(1):23, January 19, 2019. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
1/23.

Lim:2021:PMV

[LMY21] Seng Boon Lim, Jalaluddin Abdul Malek, and Tan Yigitcanlar. Postmaterialist values of smart city societies: International comparison of public values for good enough governance. Future Internet, 13(8):201, August 03, 2021. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/13/8/201.

Lu:2019:SDC

[LN19] Haoye Lu and Amiya Nayak. A scheme to design community detection algorithms in various networks. Future Internet, 11 (2):41, February 12, 2019. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/11/
                                        2/41.

Li:2017:REI

[LNA17] Haipeng Li, Hidenori Nakazato, and Syed Hassan Ahmed. Request expectation index based cache replacement algorithm for streaming content delivery over ICN. Future Internet, 9(4):83, November 14, 2017. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/9/
                                        4/83.

Le:2012:SHO

[LNB12] Quynh Lê, Hoang Boi Nguyen, and Tony Barnett. Smart homes for older people: Positive aging in a digital world. Future Internet, 4(2):607– 617, June 19, 2012. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/4/
2/607.

LoIacono:2019:NGR

[LNG19] Luigi Lo Iacono, Hoai Viet Nguyen, and Peter Leo Gorski. On the need for a general REST-security framework. Future Internet, 11(3):56, February 27, 2019. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/11/3/56.

Liu:2023:TRC

[LNH23] Hongyu Liu, Hong Ni, and Rui Han. A transmission rate control method for active congestion reduction based on network node bandwidth allocation. Future Internet, 15(12):385, November 29, 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
12/385.

Lightbody:2023:AIS

[LNT^{+}23] Dominic Lightbody, Duc-Minh Ngo, Andriy Temko, Colin C. Murphy, and Emanuel Popovici. Attacks on IoT: Side-channel power acquisition framework for intrusion detection. Future Internet, 15 (5):187, May 21, 2023. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
5/187.

Lightbody:2024:DIS

[LNT^{+}24] Dominic Lightbody, Duc-Minh Ngo, Andriy Temko, Colin C. Murphy, and Emanuel Popovici. Dragon Pi: IoT side-channel power data intrusion detection dataset and unsupervised convolutional autoencoder for intrusion detection. Future Internet, 16(3):88, March 05, 2024. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/16/3/88.

Lim:2021:RCT

[LOL21] Zhou-Yi Lim, Lee-Yeng Ong, and Meng-Chew Leow. A review on clustering techniques: Creating better user experience for online roadshow. Future Internet, 13(9):233, September 13, 2021. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/13/
                                        9/233.

Lim:2023:NRM

[LOL23] Shueh-Ting Lim, Lee-Yeng Ong, and Meng-Chew Leow. New RFI model for behavioral audience segmentation in Wi-Fi advertising system. Future Internet, 15(11):351, October 26, 2023. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/15/11/351.

Lombardo:2022:MLB

[LPA^{+}22]  Gianfranco  Lombardo,

Mattia Pellegrino, George Adosoglou, Stefano Cagnoni, Panos M. Pardalos, and Agostino Poggi. Machine learning for bankruptcy prediction in the American stock market: Dataset and benchmarks. Future Internet, 14(8):244, August 22, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/8/244.

Lux:2010:CGV

[LPM10] Mathias Lux, Arthur Pitman, and Oge Marques. Can global visual features improve tag recommendation for image annotation? Future Internet, 2(3):341–362, August 27, 2010. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/2/3/341.

Lamprogeorgos:2022:ATS

[LPPG22] Aristeidis Lamprogeorgos, Minas Pergantis, Michail Panagopoulos, and Andreas Giannakoulopoulos. Aesthetic trends and semantic Web adoption of media outlets identified through automated archival data extraction. Future Internet, 14(7):204, June 30, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.

                                        mdpi.com/1999-5903/14/
                                        7/204.

Lemmi:2024:SBE

[LPVM24] Laura Lemmi, Carlo Puliafito, Antonio Virdis, and Enzo Mingozzi. SRv6based edge service continuity in 5G mobile networks. Future Internet, 16(4):138, April 19, 2024. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/16/
                                        4/138.

Latif:2017:HWC

[LQFI17] Siddique Latif, Junaid Qadir, Shahzad Farooq, and Muhammad Ali Imran. How 5G wireless (and concomitant technologies) will revolutionize healthcare? Future Internet, 9 (4):93, December 11, 2017. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/9/
                                        4/93.

Liu:2019:CTS

[LQY^{+}19] Weihuang Liu, Jinhao Qian, Zengwei Yao, Xintao Jiao, and Jiahui Pan. Convolutional two-stream network using multi-facial feature fusion for driver fatigue detection. Future Internet, 11(5):115, May 14, 2019. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/11/5/115.

Lancelot:2023:PEL

[LRD23] Jonathan Lancelot, Bhaskar P.

Rimal, and Edward M. Dennis. Performance evaluation of a lane correction module stress test: a field test of Tesla Model 3. Future Internet, 15(4): 138, March 31, 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
4/138.

Lopez:2019:TFD

[LRdLS19] César Pérez López, María Jesús Delgado Rodríguez, and Sonia de Lucas Santos. Tax fraud detection through neural networks: An application using a sample of personal income taxpayers. Future Internet, 11(4):86, March 30, 2019. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/11/4/86.

Li:2024:CCM

[LRL^{+}24] Wei Li, Mengzhen Ren, Yazhi Liu, Chenyu Li, Hui Qian, and Zhenyou Zhang. Congestion control mechanism based on backpressure feedback in data center networks. Future Internet, 16(4):131, April 15, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/4/131.

Laneuville:2015:QAU

[LS15] Odette Laneuville and Dorota Sikora. Quantitative analysis of the usage of a pedagogical tool combining questions listed as learning objectives and answers provided as online videos. Future Internet, 7 (2):140–151, May 15, 2015. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/7/
                                        2/140.

Li:2021:NUA

[LS21] Xiaohui Li and Andrey V. Savkin. Networked unmanned aerial vehicles for surveillance and monitoring: a survey. Future Internet, 13(7):174, July 02, 2021. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/13/7/174.

Lozic:2023:FFC

[LS23] Edisa Lozić and Benjamin Stular. Fluent but not factual: a comparative analysis of ChatGPT and other AI chatbots’ proficiency and originality in scientific writing for humanities. Future Internet, 15(10):336, October 13, 2023. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/15/
                                        10/336.

Lei:2017:NHC

[LSC^{+}17] Zhou Lei, Exiong Sun, Shengbo Chen, Jiang Wu, and Wenfeng Shen. A novel hybrid-copy algorithm for live migration of virtual machine. Future Internet, 9(3):37, July 18, 2017. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/9/3/37.

Li:2019:GTA

[LSF19] Yang Li, Leyi Shi, and Haijie Feng. A gametheoretic analysis for distributed honeypots. Future Internet, 11(3):65, March 05, 2019. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/11/3/65.

Li:2021:PSB

[LSH21] Wei Li, Peng Sun, and Rui Han. Path segmentationbased hybrid caching in information-centric networks. Future Internet, 13 (1):16, January 12, 2021. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
1/16.

Li:2019:BTM

[LSJ19] Xiaoyu Li, Changyin Sun, and Fan Jiang. Beam training for millimeterwave communication based on tabu table enhanced Rosenbrock algorithm. Future Internet, 11(10):214, October 12, 2019. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
10/214.

Lowell:2012:EAS

[LSM^{+}12] Kim Lowell, Lindsay Smith, Ian Miller, Christopher Pettit, and Eloise Seymour. Extension Activity Support System (EASY): a Web-based prototype for facilitating farm management. Future Internet, 4(1):42–64, January 04, 2012. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/4/1/42.

Leiding:2021:MEM

[LSN21] Benjamin Leiding, Priyanka Sharma, and Alexander Norta. The machine-to-everything (M2X) economy: Business enactments, collaborations, and e-governance. Future Internet, 13(12):319, December 19, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/12/319.

Lembo:2022:GGL

[LSSD22] Domenico Lembo, Valerio Santarelli, Domenico Fabio Savo, and Giuseppe De Giacomo. Graphol: A graphical language for ontology modeling equivalent to OWL 2. Future Internet, 14(3):78, February 28, 2022. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/14/3/78.

Li:2021:ATF

[LSZ21] Chuanhong Li, Lei Song, and Xuewen Zeng. An adaptive throughput-first packet scheduling algorithm for DPDK-based packet processing systems. Future Internet, 13(3):78, March 19, 2021. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/13/3/78.

Liu:2024:SDS

[LT24a] Bowen Liu and Qiang Tang. Secure data sharing in federated learning through blockchain-based aggregation. Future Internet, 16(4):133, April 15, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/4/133.

Liu:2024:SBS

[LT24b] Na Liu and Wei-Tek Tsai.

SmartDED: a blockchainand smart contract-based digital electronic detonator safety supervision system. Future Internet, 16 (5):171, May 16, 2024. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
5/171.

Lombardo:2022:MUW

[LTM^{+}22]  Gianfranco  Lombardo,

Michele Tomaiuolo, Monica Mordonini, Gaia Codeluppi, and Agostino Poggi. Mobility in unsupervised word embeddings for knowledge extraction — the scholars’ trajectories across research topics. Future Internet, 14(1):25, January 09, 2022. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/14/
                                        1/25.

Lopes:2024:VTP

[LTP24] Ricardo Lopes, Marcello Trovati, and Ella Pereira. Volumetric techniques for product routing and loading optimisation in Industry 4.0: a review. Future Internet, 16(2):39, January 24, 2024. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/16/2/39.

Libbi:2021:GST

[LTTS21] Claudia Alessandra Libbi, Jan Trienes, Dolf Trieschnigg, and Christin Seifert. Generating synthetic training data for supervised de-identification of electronic health records. Future Internet, 13(5): 136, May 20, 2021. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.

mdpi.com/1999-5903/13/
5/136.

Lin:2021:ETN

[LTW21] Yi-Bing Lin, Chien-Chao Tseng, and Ming-Hung Wang. Effects of transport network slicing on 5G applications. Future Internet, 13(3):69, March 11, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/3/69.

Lei:2018:JAA

[LTY18] Jianjun Lei, Jiarui Tao, and Shanshan Yang. Joint AP association and bandwidth allocation optimization algorithm in highdense WLANs. Future Internet, 10(8):73, August 06, 2018. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/10/8/73.

Li:2011:SOA

[LW11] Haifeng Li and Bo Wu. A service-oriented architecture for proactive geospatial information services. Future Internet, 3(4):298– 318, December 19, 2011. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/3/
4/298.

Lu:2015:LJW

[LW15] Cunbo Lu and Liangtian Wan. A low-jitter wireless transmission based on buffer management in coding-aware routing. Future Internet, 7(3):307– 328, August 31, 2015. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/7/
3/307.

Li:2019:ACO

[LW19] Gang Li and Zhijun Wu. Ant colony optimization task scheduling algorithm for SWIM based on load balancing. Future Internet, 11(4):90, April 02, 2019. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/11/4/90.

Liu:2021:PPE

[LW21] Jianhua Liu and Zibo Wu. PECSA: Practical edge computing service architecture applicable to adaptive IoT-based applications. Future Internet, 13(11):294, November 19, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/11/294.

Lin:2024:EFO

[LW24] Hsin-Tung Lin and Chih-Yu Wen. Edge federated optimization for heterogeneous data. Future Internet, 16(4):142, April 22, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/4/142.

Li:2022:ISR

[LWBM22] Kangying Li, Jiayun Wang, Biligsaikhan Batjargal, and Akira Maeda. Intuitively searching for the rare colors from digital artwork collections by text description: A case demonstration of Japanese ukiyo-e print retrieval. Future Internet, 14(7):212, July 18, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
7/212.

Lin:2024:OSA

[LWH24] Ming-Yen Lin, Ping-Chun Wu, and Sue-Chen Hsueh. Optimizing session-aware recommenders: a deep dive into GRU-based latent interaction integration. Future Internet, 16 (2):51, February 01, 2024. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
2/51.

Lan:2014:NDE

[LWHR14] Xiaoyu Lan, Liangtian Wan, Guangjie Han, and Joel J. P. C. Rodrigues. A novel DOA estimation algorithm using array rotation technique. Future Internet, 6(1):155– 170, March 17, 2014. CO-DEN ???? ISSN 1999-

5903. URL https://www.

                                        mdpi.com/1999-5903/6/
                                        1/155.

Li:2022:UPC

[LWW22] Gengxian Li, Chundong Wang, and Huaibin Wang. Unreachable peers communication scheme in decentralized networks based on peer-to-peer overlay approaches. Future Internet, 14(10):290, October 12, 2022. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/14/10/290.

Lan:2023:ACB

[LWW^{+}23] Chong Lan, Yongsheng Wang, Chengze Wang, Shirong Song, and Zheng Gong. Application of ChatGPT-based digital human in animation creation. Future Internet, 15(9):300, September 02, 2023. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/15/9/300.

Liu:2024:SSU

[LWW^{+}24] Yuze Liu, Weihong Wu, Ying Wang, Jiang Liu, and Fan Yang. SUDC: Synchronous update with the division and combination of SRv6 policy. Future Internet, 16(4):140, April 22, 2024. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/16/4/140.

Li:2018:CTC

[LWX18] Yue Li, Xutao Wang, and Pengjian Xu. Chinese text classification model based on deep learning. Future Internet, 10(11):113, November 20, 2018. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/10/
11/113.

Li:2015:DRM

[LWY^{+}15] Hua Li, Huan Wang, Wenqing Yin, Yongwei Li, Yan Qian, and Fei Hu. Development of a remote monitoring system for henhouse environment based on IoT technology. Future Internet, 7(3):329–341, September 10, 2015. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/7/3/329.

Lei:2020:DBC

[LWY20] Jianjun Lei, Ying Wang, and Hong Yun. Decouplingbased channel access mechanism for improving throughput and fairness in dense multi-rate WLANs. Future Internet, 12(1):3, December 23, 2020. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/12/1/3.

Liu:2021:FSC

[LWZ^{+}21] Kunlin Liu, Ping Wang, Wenbo Zhou, Zhenyu Zhang, Yanhao Ge, Honggu Liu, Weiming Zhang, and Nenghai Yu. Face swapping consistency transfer with neural identity carrier. Future Internet, 13(11):298, November 22, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/11/298.

Liu:2022:DBA

[LWZL22] Yazhi Liu, Dongyu Wei, Chunyang Zhang, and Wei Li. Distributed bandwidth allocation strategy for QoE fairness of multiple video streams in bottleneck links. Future Internet, 14(5):152, May 18, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/5/152.

Liotou:2023:CEA

[LXG^{+}23] Eirini Liotou, Dionysis Xenakis, Vasiliki Georgara, Georgios Kourouniotis, and Lazaros Merakos. Cache-enabled adaptive video streaming: a QoEbased evaluation study. Future Internet, 15(7): 221–??, July 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
7/221.

Liu:2019:EEM

[LXL^{+}19] Ziqi Liu, Gaochao Xu, Peng Liu, Xiaodong Fu, and Yang Liu. Energyefficient multi-user routing in a software-defined multi-hop wireless network. Future Internet, 11 (6):133, June 17, 2019. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
6/133.

Li:2020:JOR

[LXL^{+}20] Long Li, Gaochao Xu, Peng Liu, Yang Li, and Jiaqi Ge. Jointly optimize the residual energy of multiple mobile devices in the MEC-WPT system. Future Internet, 12(12):233, December 20, 2020. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/12/
12/233.

Liu:2024:DNO

[LXQ24] Lisi Liu, Zijie Xu, and Xiaobin Qu. Development of a novel open control system implementation method under industrial IoT. Future Internet, 16(8):??, August 14, 2024. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
8/293.

Liu:2017:NIT

[LXY17] Lingjun Liu, Zhonghua Xie, and Cui Yang. A novel iterative thresholding algorithm based on plug-and-play priors for compressive sampling. Future Internet, 9(3):24, June 24, 2017. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/9/3/24.

Luo:2019:TSE

[LY19] Xiangfeng Luo and Yawen Yi. Topic-specific emotion mining model for online comments. Future Internet, 11(3):79, March 24, 2019. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/11/3/79.

Li:2022:AIE

[LYD22] Jiaqi Li, Jiali You, and Haojiang Deng. Adjacencyinformation-entropy-based cooperative name resolution approach in ICN. Future Internet, 14(3):68–??, February 23, 2022. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/14/
                                        3/68.

Lupetti:2017:DCP

[LYMG17] Maria Luce Lupetti, Yuan Yao, Haipeng Mi, and Claudio Germak. Design for children’s playful learning with robots. Future Internet, 9(3):52, September 18, 2017. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/9/3/52.

Lin:2019:OES

[LYP^{+}19] Guoying Lin, Yuyao Yang, Feng Pan, Sijian Zhang, Fen Wang, and Shuai Fan. An optimal energy-saving strategy for home energy management systems with bounded customer rationality. Future Internet, 11 (4):88, April 02, 2019. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
4/88.

Li:2024:MSM

[LYY^{+}24] Di Li, Zhibang Yang, Siyang Yu, Mingxing Duan, and Shenghong Yang. A micro-segmentation method based on VLAN-VxLAN mapping technology. Future Internet, 16 (9), September 4, 2024. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
9/320.

Lai:2011:OSE

[LZ11] Sabrina Lai and Corrado Zoppi. An ontology of the strategic environmental assessment of city masterplans. Future Internet, 3(4):362–378, December 20, 2011. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/3/4/362.

Lei:2023:SPA

[LZ23] Kuncheng Lei and Zhenrong Zhang. System performance analysis of sensor networks for RF energy harvesting and information transmission. Future Internet, 15(5):172, April 30, 2023. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/15/5/172.

Lotito:2021:RAS

[LZC21] Quintino Francesco Lotito, Davide Zanella, and Paolo Casari. Realistic aspects of simulation models for fake news epidemics over social networks. Future Internet, 13(3):76, March 17, 2021. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
3/76.

Liu:2017:SFI

[LZL^{+}17] Xiruo Liu, Meiyuan Zhao, Sugang Li, Feixiong Zhang, and Wade Trappe. A security framework for the Internet of Things in the future Internet architecture. Future Internet, 9(3):27, June 28, 2017. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/9/3/27.

Liu:2021:LBO

[LZL^{+}21] Yazhi Liu, Jiye Zhang, Wei Li, Qianqian Wu, and Pengmiao Li. Load balancing oriented predictive routing algorithm for data center networks. Future Internet, 13(2):54, February 22, 2021. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/13/2/54.

Liao:2023:IBF

[LZL^{+}23] Wei-Shun Liao, Ou Zhao, Keren Li, Hikaru Kawasaki, and Takeshi Matsumura. Implementation of in-band full-duplex using software defined radio with adaptive filter-based selfinterference cancellation. Future Internet, 15(11): 360, November 03, 2023. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
11/360.

Li:2022:IRG

[LZLW22] Shuailong Li, Wei Zhang, Yuquan Leng, and Xiaohui Wang. The important role of global state for multi-agent reinforcement learning. Future Internet, 14(1):17, December 30, 2022. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/14/1/17.

Li:2018:RSC

[LZS18] Rongheng Li, Jian Zhang, and Wenfeng Shen. Replicas strategy and cache optimization of video surveillance systems based on cloud storage. Future Internet, 10(4):34, April 10, 2018. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/10/4/34.

Li:2022:GCD

[LZY22] Junda Li, Chunxu Zhang, and Bo Yang. Global contextual dependency network for object detection. Future Internet, 14(1):27, January 13, 2022. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/14/1/27.

Liu:2024:COB

[LZY^{+}24] Yazhi Liu, Pengfei Zhong, Zhigang Yang, Wei Li, and Siwei Li. Computation offloading based on a distributed overlay network cache-sharing mechanism in multi-access edge computing. Future Internet, 16(4):136, April 19, 2024. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/16/
                                        4/136.

Mousa:2023:IDL

[MA23] Amthal K. Mousa and Mohammed Najm Abdullah. An improved deep learning model for DDoS detection based on hybrid stacked autoencoder and checkpoint network. Future Internet, 15(8):278, August 19, 2023. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.

mdpi.com/1999-5903/15/
8/278.

Mackay:2022:ESI

[Mac22] Michael Mackay. Editorial for the special issue on 5G enabling technologies and wireless networking. Future Internet, 14(11):342, November 21, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
11/342.

Miraz:2018:INT

[MAEP18] Mahdi H. Miraz, Maaruf Ali, Peter S. Excell, and Richard Picking. Internet of Nano-Things, things and everything: Future growth trends. Future Internet, 10(8):68, July 28, 2018. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/10/8/68.

Mahlous:2017:SES

[Mah17] Ahmed Redha Mahlous.

SCMC: An efficient scheme for minimizing energy in WSNs using a set cover approach. Future Internet, 9 (4):95, December 13, 2017. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/9/
4/95.

Malik:2020:GTA

[Mal20] Om P. Malik. Global trends and advances towards a smarter grid and smart cities. Future Internet, 12(2):37, February 14, 2020. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/12/2/37.

Md:2021:TOU

[MAM^{+}21] Abdul Quadir Md, Divyank Agrawal, Monark Mehta, Arun Kumar Sivaraman, and Kong Fah Tee. Time optimization of unmanned aerial vehicles using an augmented path. Future Internet, 13(12): 308, November 29, 2021. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
12/308.

Mourtzis:2023:FHM

[MAP23] Dimitris Mourtzis, John Angelopoulos, and Nikos Panopoulos. The future of the human–machine interface (HMI) in Society 5.0. Future Internet, 15(5): 162, April 27, 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
5/162.

Marganski:2013:VRV

[Mar13] Alison Marganski. Virtual relationship violence and perspectives on punishment: Do gender or nationality matter? Future Internet, 5(3):301– 316, June 26, 2013. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/5/
3/301.

Martelli:2017:PVN

[Mar17] Cristina Martelli. A point of view on new education for smart citizenship. Future Internet, 9(1):4, February 01, 2017. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/9/
1/4.

Matricciani:2020:GSS

[Mat20] Emilio Matricciani. Geocentric spherical surfaces emulating the geostationary orbit at any latitude with zenith links. Future Internet, 12(1):16, January 18, 2020. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/12/1/16.

Mitsis:2019:IDD

[MATP19] Giorgos Mitsis, Pavlos Athanasios Apostolopoulos, Eirini Eleni Tsiropoulou, and Symeon Papavassiliou. Intelligent dynamic data offloading in a competitive mobile edge computing market. Future Internet, 11(5):118, May 21, 2019. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/11/5/118.

Mahfouz:2020:ECN

[MAVS20] Ahmed Mahfouz, Abdullah Abuhussein, Deepak Venugopal, and Sajjan Shiva. Ensemble classifiers for network intrusion detection using a novel network attack dataset. Future Internet, 12(11):180, October 26, 2020. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/12/
11/180.

Manzoor:2019:CBR

[MAW^{+}19] Hamza Manzoor, Kamil Akhuseyinoglu, Jackson Wonderly, Peter Brusilovsky, and Clifford A. Shaffer. Crossing the borders: Reuse of smart learning objects in advanced content access systems. Future Internet, 11(7):160, July 19, 2019. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/11/7/160.

Majeed:2023:DLB

[MAW^{+}23] Abdul Majeed, Abdullah M. Alnajim, Athar Waseem, Aleem Khaliq, Aqdas Naveed, Shabana Habib, Muhammad Islam, and Sheroz Khan. Deep learning-based symptomizing cyber threats using adaptive 5G shared slice security approaches. Future Internet, 15(6):193– ??, June 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
6/193.

Mahapatra:2022:FBP

[MB22] Tanmaya Mahapatra and Syeeda Nilofer Banoo. Flow-based programming for machine learning. Future Internet, 14(2):58, February 15, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
2/58.

Martins:2020:MTR

[MBAS20] Nuno Martins, Daniel Brandão, Heitor Alvelos, and Sara Silva. Emarketplace as a tool for the revitalization of Portuguese craft industry: The design process in the development of an online platform. Future Internet, 12(11):195, November 12, 2020. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/12/11/195.

Muller:2012:DPM

[MBF^{+}12] Sebastian Müller, Franziska Brecht, Benjamin Fabian, Steffen Kunz, and Dominik Kunze. Distributed performance measurement and usability assessment of the Tor anonymization network. Future Internet, 4(2):488–513, May 15, 2012. CODEN ????

ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/4/2/488.

Marir:2022:SBF

[MBH22] Souad Marir, Faiza Belala, and Nabil Hameurlain. A strategy-based formal approach for fog systems analysis. Future Internet, 14(2):52, February 09, 2022. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/14/2/52.

Milicevic:2022:DRN

[MBLC22] Mario Milicevic, Vedran Batos, Adriana Lipovac, and Zeljka Car. Deep regression neural networks for proportion judgment. Future Internet, 14(4): 100–??, March 23, 2022. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/14/
                                        4/100.

Mooney:2012:CHE

[MC12] Peter Mooney and Padraig Corcoran. Characteristics of heavily edited objects in OpenStreetMap. Future Internet, 4(1):285– 305, March 20, 2012. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/4/
                                        1/285.

Misirli:2024:AMM

[MC24] Javid Misirli and Emiliano Casalicchio. An analysis of methods and metrics for task scheduling in fog computing. Future Internet, 16(1):16, December 30, 2024. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/16/1/16.

Marin:2023:BSB

[MCA23] Oana Marin, Tudor Cioara, and Ionut Anghel. Blockchain solution for buildings’ multi-energy flexibility trading using multi-token standards. Future Internet, 15(5):177, May 10, 2023. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/15/5/177.

Manzoni:2019:IVG

[MCCHO19] Pietro Manzoni, Carlos T. Calafate, Juan-Carlos Cano, and Enrique Hernández-Orallo. Indoor vehicles geolocalization using LoRaWAN. Future Internet, 11(6):124, May 31, 2019. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/11/6/124.

Moreno-Cabanillas:2024:DCS

[MCCOCE24] Andrea Moreno-Cabanillas, Elizabet Castillero-Ostio, and Antonio Castillo-Esparcia. Digital communication and social organizations: an evaluation of the communication strategies of the most-valued NGOs worldwide. Future Internet, 16(1):26, January 13, 2024. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/16/1/26.

Molinaro:2020:VCN

[MCH^{+}20] Antonella Molinaro, Claudia Campolo, Jérôme Härri, Christian Esteve Rothenberg, and Alexey Vinel. 5G-V2X communications and networking for connected and autonomous vehicles. Future Internet, 12(7):116, July 08, 2020. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/12/7/116.

McKenna:2020:HSE

[McK20] H. Patricia McKenna. Human-smart environment interactions in smart cities: Exploring dimensionalities of smartness. Future Internet, 12(5):79, April 27, 2020. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/12/5/79.

Mahfoudh:2010:NCC

[MCM^{+}10] Saoucene Mahfoudh, Gerard Chalhoub, Pascale Minet, Michel Misson, and Ichrak Amdouni. Node coloring and color conflict detection in wireless sensor networks. Future Internet, 2(4):469–504, October 13, 2010. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/2/4/469.

Morales-Caporal:2024:IBL

[MCPLBH^{+}24] Roberto Morales-Caporal, Rodolfo Eleazar Pérez-Loaiza, Edmundo Bonilla-Huerta, Julio Hernández-Pérez, and José de Jesús Rangel-Magdaleno. IoTbased LPG level sensor for domestic stationary tanks with data sharing to a filling plant to optimize distribution routes. Future Internet, 16(12):479, December 21, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
12/479.

Medeiros:2023:DAS

[MCR^{+}23] Eduardo Medeiros, Leonel Corado, Luís Rato, Paulo Quaresma, and Pedro Salgueiro. Domain adaptation speech-to-text for lowresource European Portuguese using deep learning. Future Internet, 15(5): 159, April 24, 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
5/159.

Ma:2024:BBZ

[MCS^{+}24] Zhe Ma, Xuhesheng Chen, Tiejiang Sun, Xukang Wang, Ying Cheng Wu, and Mengjie Zhou. Blockchainbased zero-trust supply chain security integrated with deep reinforcement learning for inventory optimization. Future Internet, 16(5):163, May 10, 2024. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/16/
                                        5/163.

Marin:2023:RBT

[MCT^{+}23] Oana Marin, Tudor Cioara, Liana Toderean, Dan Mitrea, and Ionut Anghel. Review of blockchain tokens creation and valuation. Future Internet, 15(12):382, November 27, 2023. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/15/12/382.

Malik:2022:RTN

[MDB22] Faheem Ahmed Malik, Laurent Dala, and Krishna Busawon. Real-time nanoscopic rider safety system for smart and green mobility based upon varied infrastructure parameters. Future Internet, 14 (1):9, December 25, 2022. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/14/
                                        1/9.

Matosevic:2021:UML

[MDM21] Goran Matosević, Jasminka Dobsa, and Dunja Mladenić. Using machine learning for Web page classification in search engine optimization. Future Internet, 13(1):9, January 02, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/1/9.

Mohammed:2023:EFA

[MDPHB23] Medina Ayta Mohammed, Carmen De-Pablos-Heredero, , and José Luis Montes Botella. Exploring the factors affecting countries’ adoption of blockchainenabled central bank digital currencies. Future Internet, 15(10):321, September 28, 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
10/321.

Mukhopadhyay:2024:QAI

[MDRR24] Adwitiya Mukhopadhyay, Aryadevi Remanidevi Devidas, Venkat P. Rangan, and Maneesha Vinodini Ramesh. A QoSaware IoT edge network for mobile telemedicine enabling in-transit monitoring of emergency patients. Future Internet, 16(2):52, February 06, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
2/52.

Maher:2024:HEN

[MDS^{+}24] Carol Maher, Kylie A.

Dankiw, Ben Singh, Svetlana Bogomolova, and Rachel G. Curtis. Inhome evaluation of the neo care artificial intelligence sound-based fall detection system. Future Internet, 16(6):197, June 2, 2024. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
6/197.

Margariti:2020:MST

[MDT20] Spiridoula V. Margariti, Vassilios V. Dimakopoulos, and Georgios Tsoumanis. Modeling and simulation tools for fog computing — a comprehensive survey from a cost perspective. Future Internet, 12(5):89, May 16, 2020. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/12/
5/89.

Marosi:2022:IDA

[MEH^{+}22] Attila Csaba Marosi, Márk Emodi, Ákos Hajnal, Róbert Lovas, Tamás Kiss, Valerie Poser, Jibinraj Antony, Simon Bergweiler, Hamed Hamzeh, James Deslauriers, and József Kovács. Interoperable data analytics reference architectures empowering digital-twin-aided manufacturing. Future Internet, 14(4):114–??, April 06, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/4/114.

Mesbah:2023:AIS

[MEJA23] Mohamed Mesbah, Mahmoud Said Elsayed, Anca Delia Jurcut, and Marianne Azer. Analysis of ICS and SCADA systems attacks using honeypots. Future Internet, 15(7):241– ??, July 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
7/241.

Mansour:2018:EMG

[MEO18a] Asmae Ait Mansour, Nourddine Enneya, and Mohamed Ouadou. Enhanced matching game for decoupled uplink downlink context-aware handover. Future Internet, 10(4):35, April 15, 2018. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/10/4/35.

Mansour:2018:VAH

[MEO18b] Asmae Ait Mansour, Nourddine Enneya, and Mohamed Ouadou. A velocity-aware handover trigger in two-tier heterogeneous networks. Future Internet, 10(1):9, January 15, 2018. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/10/1/9.

Muneer:2020:CAM

[MF20] Amgad Muneer and Suliman Mohamed Fati. A comparative analysis of machine learning techniques for cyberbullying detection on Twitter. Future Internet, 12(11):187, October 29, 2020. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/12/
11/187.

Matviichuk:2022:SOI

[MFH22] Liudmyla Matviichuk, Stefano Ferilli, and Nataliia Hnedko. Study of the organization and implementation of e-learning in wartime inside Ukraine. Future Internet, 14(10): 295, October 15, 2022. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
10/295.

Miranda:2020:PSM

[MFO^{+}20] Julián Miranda, Angélica Flórez, Gustavo Ospina, Ciro Gamboa, Carlos Flórez, and Miguel Altuve. Proposal for a system model for offline seismic event detection in Colombia. Future Internet, 12(12):231, December 18, 2020. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/12/12/231.

Maier:2010:QPT

[MG10] Martin Maier and Navid Ghazisaidi. QoS provisioning techniques for future fiber-wireless (FiWi) access networks. Future Internet, 2(2):126–155, April 29, 2010. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/2/2/126.

Martin:2012:RW

[MG12] Ludger Martin and Thomas Gottron. Readability and the Web. Future Internet, 4(1):238–252, March 12, 2012. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/4/1/238.

Massaro:2021:IAE

[MGBG21] Alessandro Massaro, Daniele Giannone, Vitangelo Birardi, and Angelo Maurizio Galiano. An innovative approach for the evaluation of the Web page impact combining user experience and neural network score. Future Internet, 13(6):145, May 31, 2021. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
6/145.

Maynard:2014:SCA

[MGFF14] Diana Maynard, Gerhard Gossen, Adam Funk, and Marco Fisichella. Should I care about your opinion? Detection of opinion interestingness and dynamics in social media. Future Internet, 6(3):457–481, August 13, 2014. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/6/3/457.

Meng:2022:IFI

[MGJ22] Weizhi Meng, Thanassis Giannetsos, and Christian D. Jensen. Information and future Internet security, trust and privacy. Future Internet, 14(12): 372, December 12, 2022. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/14/
                                        12/372.

Moreno-Guerrero:2020:SDE

[MGLBMMSC20] Antonio-José Moreno-Guerrero, Jesús López-Belmonte, José-Antonio Marín-Marín, and Rebeca Soler-Costa. Scientific development of educational artificial intelligence in Web of Science. Future Internet, 12 (8):124, July 24, 2020. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/12/
                                        8/124.

Mangler:2023:DXE

[MGM^{+}23] Juergen Mangler, Joscha Grüger, Lukas Malburg, Matthias Ehrendorfer, Yannis Bertrand, Janik-Vasily Benzin, Stefanie Rinderle-Ma, Estefania Serral Asensio, and Ralph Bergmann. DataStream XES extension: Embedding IoT sensor data into extensible event stream logs. Future Internet, 15 (3):109, March 14, 2023. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
3/109.

Montenegro:2012:LUP

[MGUD12] Nuno Montenegro, Jorge C.

Gomes, Paulo Urbano, and José P. Duarte. A land use planning ontology: LBCS. Future Internet, 4(1):65–82, January 06, 2012. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/4/1/65.

Mazzenga:2017:FBF

[MGV17] Franco Mazzenga, Romeo Giuliano, and Francesco Vatalaro. FttC-based fronthaul for 5G dense/ultradense access network: Performance and costs in realistic scenarios. Future Internet, 9(4):71, October 27, 2017. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/9/4/71.

Menth:2013:GLL

[MHK13] Michael Menth, Matthias Hartmann, and Dominik Klein. Global locator, local locator, and identifier split (GLI-split). Future Internet, 5(1):67–94, March 11, 2013. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/5/1/67.

Musa:2023:DNN

[MHL^{+}23] Abubakar Ahmad Musa, Adamu Hussaini, Weixian Liao, Fan Liang, and Wei Yu. Deep neural networks for spatial-temporal cyber-physical systems: a survey. Future Internet, 15(6):199–??, June 2023. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/15/
                                        6/199.

Musa:2023:ORA

[MHQ^{+}23] Abubakar Ahmad Musa, Adamu Hussaini, Cheng Qian, Yifan Guo, and Wei Yu. Open radio access networks for smart IoT systems: State of art and future directions. Future Internet, 15(12):380, November 27, 2023. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/15/12/380.

Maaradji:2023:SCM

[MHS23] Abderrahmane Maaradji, Hakim Hacid, and Assia Soukane. From service composition to Mashup Editor: a multiperspective taxonomy. Future Internet, 15(2):59, January 31, 2023. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/15/2/59.

Merrick:2016:SGT

[MHSH16] Kathryn Merrick, Medria Hardhienata, Kamran Shafi, and Jiankun Hu. A survey of game theoretic approaches to modelling decision-making in information warfare scenarios. Future Internet, 8(3):34, July 22, 2016. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/8/3/34.

Mazhar:2023:ACS

[MIK^{+}23] Tehseen Mazhar, Hafiz Muhammad Irfan, Sunawar Khan, Inayatul Haq, Inam Ullah, Muhammad Iqbal, and Habib Hamam. Analysis of cyber security attacks and its solutions for the smart grid using machine learning and blockchain methods. Future Internet, 15 (2):83, February 19, 2023. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
2/83.

Minati:2020:CCS

[Min20] Gianfranco Minati. Complex cognitive systems and their unconscious. related inspired conjectures for artificial intelligence. Future Internet, 12(12):213, November 27, 2020. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/12/
12/213.

Mana:2024:SAR

[MJG24] Antonio Maña, Francisco J. Jaime, and Lucía Gutiérrez. A secure auditable remote registry pattern for IoT systems. Future Internet, 16(11): 405, November 4, 2024. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/16/
                                        11/405.

Mikhailov:2020:TBA

[MK20] Sergei Mikhailov and Alexey Kashevnik. Tourist behaviour analysis based on digital pattern of life — an approach and case study. Future Internet, 12(10):165, September 28, 2020. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/12/10/165.

Masinde:2020:SEL

[MKDG20] Newton Masinde, Liat Khitman, Iakov Dlikman, and Kalman Graffi. Systematic evaluation of LibreSocial — a peer-to-peer framework for online social networks. Future Internet, 12(9):140, August 20, 2020. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/12/9/140.

Muller:2021:DCH

[MKHR21] Romy Müller, Franziska Kessler, David W. Humphrey, and Julian Rahm. Data in context: How digital transformation can support human reasoning in cyber-physical production systems. Future Internet, 13(6):156, June 17, 2021. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
6/156.

Markus:2017:CAI

[MKK17a] Andras Markus, Attila Kertesz, and Gabor Kecskemeti. Cost-aware IoT extension of DISSECT-CF. Future Internet, 9(3):47, August 14, 2017. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/9/3/47.

Menesidou:2017:CKM

[MKK17b] Sofia Anna Menesidou, Vasilios Katos, and Georgios Kambourakis. Cryptographic key management in delay tolerant networks: a survey. Future Internet, 9(3):26, June 27, 2017. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/9/
3/26.

Mojzis:2022:TRB

[MKK^{+}22] Ján Mojzis, Peter Krammer, Marcel Kvassay, Lenka Skovajsová, and Ladislav Hluchý. Towards reliable baselines for document-level sentiment analysis in the Czech and Slovak languages. Future Internet, 14(10):300, October 19, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/10/300.

Mavroeidi:2019:GVP

[MKKG19] Aikaterini-Georgia Mavroeidi, Angeliki Kitsiou, Christos Kalloniatis, and Stefanos Gritzalis. Gamification vs. privacy: Identifying and analysing the major concerns. Future Internet, 11 (3):67, March 07, 2019. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
3/67.

Mambou:2018:NCV

[MKKS18] Sebastien Mambou, Ondrej Krejcar, Kamil Kuca, and Ali Selamat. Novel cross-view human action model recognition based on the powerful viewinvariant features technique. Future Internet, 10(9):89, September 13, 2018. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/10/9/89.

Mtonga:2019:MLB

[MKM^{+}19] Kambombo Mtonga, Santhi Kumaran, Chomora Mikeka, Kayalvizhi Jayavel, and Jimmy Nsenga. Machine learning-based patient load prediction and IoT integrated intelligent patient transfer systems.

Future Internet, 11(11): 236, November 12, 2019. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
11/236.

Mpimis:2022:CDG

[MKPG22] Thanassis Mpimis, Theodore T.

Kapsis, Athanasios D. Panagopoulos, and Vassilis Gikas. Cooperative D-GNSS aided with multi attribute decision making module: a rigorous comparative analysis. Future Internet, 14(7):195, June 27, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/7/195.

Meqdad:2022:IDO

[MKR22] Maytham N. Meqdad, Seifedine Kadry, and Hafiz Tayyab Rauf. Improved dragonfly optimization algorithm for detecting IoT outlier sensors. Future Internet, 14(10):297, October 17, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
10/297.

Maslej-Kresnakova:2022:UDA

[MKSJ22] Viera Maslej-Kresnáková, Martin Sarnovský, and Júlia Jacková. Use of data augmentation techniques in detection of antisocial behavior using deep learning methods. Future Internet, 14(9):260, August 31, 2022. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/14/9/260.

Malarczyk:2023:IRT

[MKSK23] Mateusz Malarczyk, Grzegorz Kaczmarczyk, Jaroslaw Szrek, and Marcin Kaminski. Internet of Robotic Things (IoRT) and metaheuristic optimization techniques applied for wheellegged robot. Future Internet, 15(9):303, September 06, 2023. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/15/9/303.

Mimidis-Kentis:2019:NGP

[MKSV^{+}19] Angelos Mimidis-Kentis, Jose Soler, Paul Veitch, Adam Broadbent, Marco Mobilio, Oliviero Riganelli, Steven Van Rossem, Wouter Tavernier, and Bessem Sayadi. The next generation platform as a service: Composition and deployment of platforms and services. Future Internet, 11(5):119, May 21, 2019. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/11/5/119.

Mikoczy:2011:ECN

[MKvD11] Eugen Mikóczy, Ivan Kotuliak, and Oskar van Deventer. Evolution of the converged NGN service platforms towards future networks. Future Internet, 3(1):67–86, March 04, 2011. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/3/1/67.

Matsuoka:2024:OOP

[MKY24] Ryo Matsuoka, Koichi Kobayashi, and Yuh Yamashita. Online optimization of pickup and delivery problem considering feasibility. Future Internet, 16 (2):64, February 17, 2024. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
2/64.

Markopoulos:2020:ISO

[ML20] Evangelos Markopoulos and Mika Luimula. Immersive safe oceans technology: Developing virtual onboard training episodes for maritime safety. Future Internet, 12(5):80, April 28, 2020. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/12/5/80.

Manolopoulos:2022:VME

[MLK22]  Ioannis  Manolopoulos,

Dimitrios Loukatos, and Kimon Kontovasilis. A versatile MANET experimentation platform and its evaluation through experiments on the performance of routing protocols under diverse conditions. Future Internet, 14(5):154, May 19, 2022. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/14/
                                        5/154.

Munjal:2022:MAD

[MLL^{+}22] Rashmi Munjal, William Liu, Xuejun Li, Jairo Gutierrez, and Peter Han Joo Chong. Multiattribute decision making for energy-efficient public transport network selection in smart cities. Future Internet, 14(2):42, January 26, 2022. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/14/2/42.

Motoyoshi:2014:MTF

[MLM14] Gen Motoyoshi, Kenji Leibnitz, and Masayuki Murata. Mobility tolerant firework routing for improving reachability in MANETs. Future Internet, 6(1):171–189, March 24, 2014. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/6/1/171.

Ma:2023:DSB

[MLPH23] Khai-Minh Ma, Duc-Hung Le, Cong-Kha Pham, and Trong-Thuc Hoang. Design of an SoC based on 32-bit RISC-V processor with low-latency lightweight cryptographic cores in FPGA. Future Internet, 15(5):186, May 19, 2023. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/15/5/186.

Ma:2024:RSS

[MLWS24] Min Ma, Shanrong Liu, Shufei Wang, and Shengnan Shi. Refined semisupervised modulation classification: Integrating consistency regularization and pseudo-labeling techniques. Future Internet, 16 (2):38, January 23, 2024. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
2/38.

Mountaser:2021:SBE

[MM21] Ghizlane Mountaser and Toktam Mahmoodi. An SDR-based experimental study of reliable and lowlatency Ethernet-based fronthaul with MAC-PHY split. Future Internet, 13 (7):170, June 30, 2021. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
7/170.

Mrvelj:2023:FVF

[MM23] Stefica Mrvelj and Marko Matulin. FLAME-VQA: a fuzzy logic-based model for high frame rate video quality assessment. Future Internet, 15(9):295, September 01, 2023. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/15/
                                        9/295.

Mahfoudh:2010:EER

[MMA10] Saoucene Mahfoudh, Pascale Minet, and Ichrak Amdouni. Energy efficient routing and node activity scheduling in the OCARI wireless sensor network. Future Internet, 2(3):308– 340, August 17, 2010. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/2/
                                        3/308.

Manugunta:2022:DLB

[MMD22] Ramya Krishna Manugunta, Rytis Maskeliunas, and Robertas Damasevicius. Deep learning based semantic image segmentation methods for classification of Web page imagery. Future Internet, 14(10): 277, September 27, 2022. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/14/
                                        10/277.

Marabissi:2019:RCI

[MMF^{+}19] Dania Marabissi, Lorenzo Mucchi, Romano Fantacci, Maria Rita Spada, Fabio Massimiani, Andrea Fratini, Giorgio Cau, Jia Yunpeng, and Lucio Fedele. A real case of implementation of the future 5G city. Future Internet, 11(1):4, December 22, 2019. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/11/1/4.

Marti:2016:DED

[MMH16] Patrizia Marti, Carl Megens, and Caroline Hummels. Data-enabled design for social change: Two case studies. Future Internet, 8(4):46, September 23, 2016. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/8/4/46.

Marques:2021:DFB

[MMM21] Claudio Marques, Silvestre Malta, and João Magalhães. DNS firewall based on machine learning. Future Internet, 13(12):309, November 30, 2021. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
12/309.

Menth:2019:IEA

[MMMH19] Michael Menth, Habib Mostafaei, Daniel Merling, and Marco Häberle. Implementation and evaluation of activity-based congestion management using P4 (P4-ABC). Future Internet, 11(7):159, July 19, 2019. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/11/7/159.

Mastroeni:2019:SLA

[MMN19] Loretta Mastroeni, Alessandro Mazzoccoli, and Maurizio Naldi. Service level agreement violations in cloud storage: Insurance and compensation sustainability. Future Internet, 11(7):142, June 30, 2019. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/11/
                                        7/142.

Miguel:2024:UCV

[MMQ^{+}24] Joaquim Miguel, Pedro Mendonça, Agnelo Quelhas, João M. L. P. Caldeira, and Vasco N. G. J. Soares. Using computer vision to collect information on cycling and hiking trails users. Future Internet, 16(3):104, March 20, 2024. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/16/
                                        3/104.

Mavropoulos:2019:CAC

[MMS^{+}19] Thanassis Mavropoulos, Georgios Meditskos, Spyridon Symeonidis, Eleni Kamateri, Maria Rousi, Dimitris Tzimikas, Lefteris Papageorgiou, Christos Eleftheriadis, George Adamopoulos, Stefanos Vrochidis, and Ioannis Kompatsiaris. A contextaware conversational agent in the rehabilitation domain. Future Internet, 11(11):231, November 01, 2019. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/11/11/231.

Marabissi:2022:INA

[MMS22] Dania Marabissi, Lorenzo Mucchi, and Andrea Stomaci. IoT nodes authentication and ID spoofing detection based on joint use of physical layer security and machine learning. Future Internet, 14(2):61, February 17, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/2/61.

Marletta:2024:AIR

[MMT24] Daniele Marletta, Alessandro Midolo, and Emiliano Tramontana. Automatically injecting robustness statements into distributed applications. Future Internet, 16(11):416, November 10, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
11/416.

Morshedi:2021:EPV

[MN21] Maghsoud Morshedi and Josef Noll. Estimating PQoS of video conferencing on Wi-Fi networks using machine learning. Future Internet, 13(3):63, March 03, 2021. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/13/3/63.

Mohideen:2024:BCI

[MNH^{+}24] Mohamed Azarudheen Mohamed Mohideen, Muhammad Shahroz Nadeem, James Hardy, Haider Ali, Umair Ullah Tariq, Fariza Sabrina, Muhammad Waqar, and Salman Ahmed. Behind the code: Identifying zero-day exploits in WordPress. Future Internet, 16(7):??, July 19, 2024. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/16/
                                        7/256.

Mahbub:2024:FFV

[MNP^{+}24] Khaled Mahbub, Antonio Nehme, Mohammad Patwary, Marc Lacoste, and Sylvain Allio. FI-VADMI: a framework for in-vehicle anomaly detection by monitoring and isolation. Future Internet, 16(8):??, August 8, 2024. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/16/
                                        8/288.

Ming:2021:DCV

[MNWK21] Teo Rhun Ming, Noris Mohd Norowi, Rahmita Wirza, and Azrina Kamaruddin. Designing a collaborative virtual conference application: Challenges, requirements and guidelines. Future Internet, 13(10):253, September 29, 2021. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
10/253.

Militano:2017:NID

[MOAI17] Leonardo Militano, Antonino Orsino, Giuseppe Araniti, and Antonio Iera. NB-IoT for D2D-enhanced content uploading with social trustworthiness in 5G systems. Future Internet, 9 (3):31, July 08, 2017. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/9/
3/31.

Mottura:2024:DAC

[Mot24] Stefano Mottura. Does anyone care about the opinion of people on participating in a “Social” metaverse? A review and a draft proposal for a surveying tool. Future Internet, 16(7):??, July 2, 2024. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
7/236.

Moubayed:2024:CED

[Mou24] Abdallah Moubayed. A complete EDA and DL pipeline for softwarized 5G network intrusion detection. Future Internet, 16 (9), September 10, 2024. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/16/
                                        9/331.

Multisilta:2013:SLW

[MP13] Jari Multisilta and Arttu Perttula. Supporting learning with wireless sensor data. Future Internet, 5(1):95–112, March 19, 2013. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/5/1/95.

Martini:2016:SOA

[MP16] Barbara Martini and Federica Paganelli. A serviceoriented approach for dynamic chaining of virtual network functions over multi-provider softwaredefined networks. Future Internet, 8(2):24, June 01, 2016. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/8/2/24.

Marchiori:2020:SSS

[MP20] Massimo Marchiori and Lino Possamai. Strategies of success for social networks: Mermaids and temporal evolution. Future Internet, 12(2):25, February 04, 2020. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/12/2/25.

Munoz:2023:PAD

[MPCS^{+}23] Ernesto Cadena Muñoz, Gustavo Chica Pedraza, Rafael Cubillos-Sánchez, Alexander Aponte-Moreno, and Mónica Espinosa Buitrago. PUE attack detection by using DNN and entropy in cooperative mobile cognitive radio networks. Future Internet, 15(6):202–??, June 2023. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
6/202.

Marotta:2020:BAQ

[MPG20] Lorena Marotta, Andrea Pesce, and Andrea Guazzini. Before and after the quarantine: an approximate study on the psychological impact of COVID-19 on the Italian population during the lockdown period. Future Internet, 12(12):229, December 15, 2020. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/12/12/229.

Moustakidis:2023:PFT

[MPK^{+}23] Serafeim Moustakidis, Spyridon Plakias, Christos Kokkotis, Themistoklis Tsatalas, and Dimitrios Tsaopoulos. Predicting football team performance with explainable AI: Leveraging SHAP to identify key team-level performance metrics. Future Internet, 15(5):174, May 05, 2023. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/15/
                                        5/174.

Miranda:2021:CHC

[MPL^{+}21] Pedro R. Miranda, Daniel Pestana, João D. Lopes, Rui Policarpo Duarte, Mário P. Véstias, Horácio C. Neto, and José T. de Sousa. Configurable hardware core for IoT object detection. Future Internet, 13(11):280, October 30, 2021. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/13/11/280.

Munoz:2024:MPA

[MPM24] Ernesto Cadena Muñoz, Gustavo Chica Pedraza, and Alexander Aponte Moreno. Multiple PUE attack detection in cooperative mobile cognitive radio networks. Future Internet, 16(12):456, December 4, 2024. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/16/12/456.

Mandarino:2023:PFD

[MPT23] Valerio Mandarino, Giuseppe Pappalardo, and Emiliano Tramontana. Proof of flow: a design pattern for the green energy market. Future Internet, 15(9):313, September 17, 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
9/313.

Matricciani:2023:TFL

[MR23a] Emilio Matricciani and Carlo Riva. Transfer functions and linear distortions in ultra-wideband channels faded by rain in GeoSurf satellite constellations. Future Internet, 15 (1):27, January 03, 2023. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
1/27.

Mohammed:2023:FIA

[MR23b] Sarfaraz Ahmed Mohammed and Anca L. Ralescu. Future Internet architectures on an emerging scale — a systematic review. Future Internet, 15(5):166, April 29, 2023. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
5/166.

Miao:2024:MBS

[MRM^{+}24] Jinpeng Miao, Dasari Rajasekhar, Shivakant Mishra, Sanjeet Kumar Nayak, and Ramanarayan Yadav. A microservicebased smart agriculture system to detect animal intrusion at the edge. Future Internet, 16(8):??, August 16, 2024. CODEN ????

ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/16/8/296.

Matsumoto-Royo:2021:ODL

[MRRMC21] Kiomi Matsumoto-Royo, Maria Soledad Ramírez-Montoya, and Paulette Conget. Opportunities to develop lifelong learning tendencies in practicebased teacher education: Getting ready for Education 4.0. Future Internet, 13(11):292, November 19, 2021. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/13/11/292.

Massa:2016:FEU

[MS16] Daniele Massa and Lucio Davide Spano. Face-Mashup: An end-user development tool for social network data. Future Internet, 8(2):10, March 29, 2016. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/8/2/10.

Martinez:2020:IBV

[MS20] Juan Miguel Martínez Martínez and Antonio Tudela Sancho. Interactions between virtual spaces and schools: a collective case study. Future Internet, 12(12):217, December 02, 2020. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/12/12/217.

Muthavhine:2022:SID

[MS22] Khumbelo Muthavhine and Mbuyu Sumbwanyambe. Securing IoT devices against differential-linear (DL) attack used on Serpent algorithm. Future Internet, 14 (2):55, February 13, 2022. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
2/55.

Mokrov:2023:PAC

[MS23] Evgeni Mokrov and Konstantin Samouylov. Performance assessment and comparison of deployment options for 5G millimeter wave systems. Future Internet, 15(2):60, January 31, 2023. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/15/2/60.

Moedjahedy:2022:CCC

[MSAA22] Jimmy Moedjahedy, Arief Setyanto, Fawaz Khaled Alarfaj, and Mohammed Alreshoodi. CCrFS: Combine correlation features selection for detecting phishing Websites using machine learning. Future Internet, 14(8):229, July 27, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/8/229.

Moubayed:2022:EEI

[MSAD22] Abdallah Moubayed, Abdallah Shami, and Anwer Al-Dulaimi. On end-to-end intelligent automation of 6G networks. Future Internet, 14(6):165, May 29, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/6/165.

Mackita:2019:ERM

[MSC19] Masky Mackita, Soo-Young Shin, and Tae-Young Choe. ERMOC-TAVE: a risk management framework for IT systems which adopt cloud computing. Future Internet, 11(9):195, September 10, 2019. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/11/9/195.

Moreno:2021:OSF

[MSC^{+}21] Jesús Fernando Cevallos Moreno, Rebecca Sattler, Raúl P. Caulier Cisterna, Lorenzo Ricciardi Celsi, Aminael Sánchez Rodríguez, and Massimo Mecella. Online service function chain deployment for live-streaming in virtualized content delivery networks: a deep reinforcement learning approach. Future Internet, 13(11): 278, October 29, 2021. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
11/278.

Manzoor:2024:SSS

[MSC^{+}24] Habib Ullah Manzoor, Attia Shabbir, Ao Chen, David Flynn, and Ahmed Zoha. A survey of security strategies in federated learning: Defending models, data, and privacy. Future Internet, 16(10), October 15, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
10/374.

Moreno:2016:MIB

[MSFM16] Julio Moreno, Manuel A.

Serrano, and Eduardo Fernández-Medina. Main issues in big data security. Future Internet, 8(3):44, September 01, 2016. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/8/
3/44.

Mescheryakov:2020:DCE

[MSIP20] Serg Mescheryakov, Dmitry Shchemelinin, Konstantin Izrailov, and Victor Pokussov. Digital cloud environment: Present challenges and future forecast. Future Internet, 12(5):82, April 29, 2020. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/12/5/82.

Mukasine:2023:CAM

[MSJ^{+}23] Angelique Mukasine, Louis Sibomana, Kayalvizhi Jayavel, Kizito Nkurikiyeyezu, and Eric Hitimana. Correlation analysis model of environment parameters using IoT framework in a biogas energy generation context. Future Internet, 15(8):265, August 09, 2023. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/15/8/265.

Mousavi:2018:FCM

[MSK^{+}18] Seyed Davoud Mousavi, Rasool Sadeghi, Mohamadreza Karimi, Erfan Karimian, and Mohammad Reza Soltan Aghaei. A fair cooperative MAC protocol in IEEE 802.11 WLAN. Future Internet, 10(5):39, May 03, 2018. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/10/
5/39.

Mehmood:2023:UKT

[MSL^{+}23] Tahir Mehmood, Ivan Serina, Alberto Lavelli, Luca Putelli, and Alfonso Gerevini. On the use of knowledge transfer techniques for biomedical named entity recognition. Future Internet, 15(2):79, February 17, 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
2/79.

Mantere:2013:NTF

[MSN13] Matti Mantere, Mirko Sailio, and Sami Noponen. Network traffic features for anomaly detection in specific industrial control system network. Future Internet, 5(4):460–473, September 25, 2013. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/5/4/460.

Moumgiakmas:2021:CVF

[MSP21] Seraphim S. Moumgiakmas, Gerasimos G. Samatas, and George A. Papakostas. Computer vision for fire detection on UAVs — from software to hardware. Future Internet, 13(8):200, July 31, 2021. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
8/200.

Mearns:2014:TMS

[MSR^{+}14] Graeme Mearns, Rebecca Simmonds, Ranald Richardson, Mark Turner, Paul Watson, and Paolo Missier. Tweet my street: a cross-disciplinary collaboration for the analysis of local Twitter data. Future Internet, 6(2):378– 396, May 27, 2014. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/6/
2/378.

Meng:2021:PNF

[MSX^{+}21] Fanhui Meng, Haoming Sun, Jiarong Xie, Chengjun Wang, Jiajing Wu, and Yanqing Hu. Preference for number of friends in online social networks. Future Internet, 13(9):236, September 16, 2021. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/13/9/236.

Mangi:2023:SMC

[MSZ23] Fawad Ali Mangi, Guoxin Su, and Minjie Zhang. Statistical model checking in process mining: a comprehensive approach to analyse stochastic processes. Future Internet, 15(12): 378, November 26, 2023. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/15/
                                        12/378.

Mamais:2017:BVP

[MT17a] Stylianos S. Mamais and George Theodorakopoulos. Behavioural verification: Preventing report fraud in decentralized advert distribution systems. Future Internet, 9(4):88, November 20, 2017. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/9/
                                        4/88.

Mamais:2017:PSD

[MT17b] Stylianos S. Mamais and George Theodorakopoulos. Private and secure distribution of targeted advertisements to mobile phones. Future Internet, 9 (2):16, May 01, 2017. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/9/
2/16.

Mazumdar:2020:CST

[MT20] Suvodeep Mazumdar and Dhavalkumar Thakker. Citizen science on Twitter: Using data analytics to understand conversations and networks. Future Internet, 12(12):210, November 26, 2020. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/12/
12/210.

Midolo:2023:ATS

[MT23] Alessandro Midolo and Emiliano Tramontana. An automatic transformer from sequential to parallel Java code. Future Internet, 15(9):306, September 08, 2023. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/15/9/306.

Mallikarachchi:2020:DCR

[MTA^{+}20] Thanuja Mallikarachchi, Dumidu Talagala, Hemantha Kodikara Arachchi, Chaminda Hewage, and Anil Fernando. A decodingcomplexity and ratecontrolled video-coding algorithm for HEVC. Future Internet, 12(7):120, July 16, 2020. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/12/7/120.

Murugappan:2021:CIP

[MTF^{+}21] Manickam Murugappan, John Victor Joshua Thomas, Ugo Fiore, Yesudas Bevish Jinila, and Subhashini Radhakrishnan. COVID-Net: Implementing parallel architecture on sound and image for high efficacy. Future Internet, 13(11): 269, October 26, 2021. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
11/269.

Murorunkwere:2022:FDU

[MTHN22] Belle Fille Murorunkwere, Origene Tuyishimire, Dominique Haughton, and Joseph Nzabanita. Fraud detection using neural networks: a case study of income tax. Future Internet, 14(6):168, May 31, 2022. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
6/168.

Mudassar:2024:PPD

[MTK^{+}24] Bakhtawar Mudassar, Shahzaib Tahir, Fawad Khan, Syed Aziz Shah, Syed Ikram Shah, and Qammer Hussain Abbasi. Privacy-preserving data analytics in Internet of Medical Things. Future Internet, 16(11):407, November 5, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
11/407.

Margariti:2024:TCS

[MTKS24] Spiridoula V. Margariti, Ioannis G. Tsoulos, Evangelia Kiousi, and Eleftherios Stergiou. Traffic classification in softwaredefined networking using genetic programming tools. Future Internet, 16 (9), September 19, 2024. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
9/338.

Mozaffariahrar:2022:SWF

[MTM22] Erfan Mozaffariahrar, Fabrice Theoleyre, and Michael Menth. A survey of Wi-Fi 6: Technologies, advances, and challenges. Future Internet, 14(10):293, October 14, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
10/293.

Mugridge:2011:EWB

[MUS11] Rick Mugridge, Mark Utting, and David Streader. Evolving Web-based test automation into agile business specifications. Future Internet, 3(2):159– 174, June 03, 2011. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/3/
                                        2/159.

Maniou:2020:ECN

[MV20] Theodora A. Maniou and Andreas Veglis. Employing a chatbot for news dissemination during crisis: Design, implementation and evaluation. Future Internet, 12(7):109, June 30, 2020. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/12/7/109.

Marin-Vega:2023:ERX

[MVAHBL^{+}23] Humberto Marín-Vega, Giner Alor-Hernández, Maritza Bustos-López, Ignacio López-Martínez, and Norma Leticia Hernández-Chaparro. Extended reality (XR) engines for developing gamified apps and serious games: a scoping review. Future Internet, 15(12):379, November 27, 2023. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/15/12/379.

Moreno-Vozmediano:2024:IRO

[MVMHL24] Rafael Moreno-Vozmediano, Rubén S. Montero, Eduardo Huedo, and Ignacio M. Llorente. Intelligent resource orchestration for 5G edge infrastructures. Future Internet, 16(3):103, March 19, 2024. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
3/103.

Maier:2024:MMS

[MW24] Florian Maier and Markus Weinberger. Metaverse meets smart cities — applications, benefits, and challenges. Future Internet, 16(4):126, April 08, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/4/126.

Milicic:2020:GTA

[MWL20] Gregor Milicic, Sina Wetzel, and Matthias Ludwig. Generic tasks for algorithms. Future Internet, 12(9):152, September 03, 2020. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/12/9/152.

Meng:2019:MTR

[MWY19] Xiangli Meng, Lingda Wu, and Shaobo Yu. Multitopology routing algorithms in SDN-based space information networks. Future Internet, 11(1):15, January 12, 2019. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
1/15.

Mei:2021:MMF

[MXL^{+}21] Jing Mei, Huahu Xu, Yang Li, Minjie Bian, and Yuzhe Huang. MFCNet: Mining features context network for RGB-IR person re-identification. Future Internet, 13(11):290, November 18, 2021. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/13/
                                        11/290.

Ma:2016:MDR

[MXT^{+}16] Tinghuai Ma, Xichao Xu, Meili Tang, Yuanfeng Jin, and Wenhai Shen. MH-Base: a distributed realtime query scheme for meteorological data based on HBase. Future Internet, 8 (1):6, March 01, 2016. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/8/
                                        1/6.

Morrow:2019:BTI

[MZ19] Monique J. Morrow and Mehran Zarrebini. Blockchain and the tokenization of the individual: Societal implications. Future Internet, 11(10):220, October 22, 2019. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/11/10/220.

Mir:2022:EGR

[MZH22] Mohd-Yaseen Mir, Hengbing Zhu, and Chih-Lin Hu. Enhanced geographic routing with one- and twohop movement information in opportunistic ad hoc networks. Future Internet, 14(7):214, July 20, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/7/214.

Musa:2022:CIC

[MZLP22] Salahadin Seid Musa, Marco Zennaro, Mulugeta Libsie, and Ermanno Pietrosemoli. Convergence of information-centric networks and edge intelligence for IoV: Challenges and future directions. Future Internet, 14(7):192, June 25, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
7/192.

Mahmood:2019:SDH

[MZS19] Adnan Mahmood, Wei Emma Zhang, and Quan Z. Sheng. Software-defined heterogeneous vehicular networking: The architectural design and open challenges. Future Internet, 11(3):70, March 11, 2019. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
3/70.

Nekovee:2023:VET

[NA23] Maziar Nekovee and Ferheen Ayaz. Vision, enabling technologies, and scenarios for a 6G-enabled Internet of Verticals (6G-IoV). Future Internet, 15 (2):57, January 30, 2023. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
2/57.

Nunez-Agurto:2024:NTC

[NAFM^{+}24] Daniel Nuñez-Agurto, Walter Fuertes, Luis Marrone, Eduardo Benavides-Astudillo, Christian Coronel-Guerrero, and Franklin Perez. A novel traffic classification approach by employing deep learning on software-defined networking. Future Internet, 16(5): 153, April 29, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
5/153.

Nadal:2023:MSD

[NAV^{+}23] Laia Nadal, Mumtaz Ali, Francisco Javier Vílchez, Josep Maria Fàbrega, and Michela Svaluto Moreolo. The multiband over spatial division multiplexing sliceable transceiver for future optical networks. Future Internet, 15(12):381, November 27, 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
12/381.

Nordfeldt:2012:YPV

[NB12] Sam Nordfeldt and Carina Berterö. Young patients’ views on the open Web 2.0 childhood diabetes patient portal: a qualitative study. Future Internet, 4(2):514– 527, May 18, 2012. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/4/
2/514.

Nayak:2024:IDS

[NB24] Nanavath Kiran Singh Nayak and Budhaditya Bhattacharyya. An intrusion detection system for 5G SDN network utilizing binarized deep spiking capsule fire hawk neural networks and blockchain technology. Future Internet, 16(10), October 3, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/10/359.

Nikic:2024:LDR

[NBL^{+}24] Vladimir Nikić, Dusan Bortnik, Milan Lukić, Dejan Vukobratović, and Ivan Mezei. Lightweight digit recognition in smart metering system using narrowband Internet of Things and federated learning. Future Internet, 16(11):402, October 31, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/11/402.

Nafeh:2022:JSV

[NBT22] Majd Nafeh, Arash Bozorgchenani, and Daniele Tarchi. Joint scalable video coding and transcoding solutions for fogcomputing-assisted DASH video applications. Future Internet, 14(9):268, September 17, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
9/268.

Nitu:2024:BLB

[ND24] Melania Nitu and Mihai Dascalu. Beyond lexical boundaries: LLMgenerated text detection for Romanian digital libraries. Future Internet, 16(2):41, January 25, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/2/41.

Nguyen-Duc:2019:MVP

[NDKBL19] Anh Nguyen-Duc, Khan Khalid, Sohaib Shahid Bajwa, and Tor Lønnestad. Minimum viable products for Internet of Things applications: Common pitfalls and practices. Future Internet, 11(2):50, February 18, 2019. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/11/2/50.

Nguyen:2023:HHE

[NDS^{+}23] Duy Tung Khanh Nguyen, Dung Hoang Duong, Willy Susilo, Yang-Wai Chow, and The Anh Ta. HeFUN: Homomorphic encryption for unconstrained secure neural network inference. Future Internet, 15(12): 407, December 18, 2023. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
12/407.

Nikou:2021:FMA

[NE21] Stavros A. Nikou and Anastasios A. Economides. A framework for mobile-assisted formative assessment to promote students’ self-determination. Future Internet, 13(5):116, April 30, 2021. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
5/116.

Nguyen:2019:RSS

[Ngu19] Van Suong Nguyen. Research on a support system for automatic ship navigation in fairway. Future Internet, 11(2):38, February 03, 2019. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/11/2/38.

Nikulchev:2021:ISE

[NIK^{+}21] Evgeny Nikulchev, Dmitry Ilin, Pavel Kolyasnikov, Shamil Magomedov, Anna Alexeenko, Alexander N. Kosenkov, Andrey Sokolov, Artem Malykh, Victoria Ismatullina, and Sergey Malykh. Isolated sandbox environment architecture for running cognitive psychological experiments in Web platforms. Future Internet, 13(10):245, September 24, 2021. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
10/245.

Nilsen:2012:CBP

[Nil12] Line Lundvoll Nilsen. Collaboration between professionals: The use of videoconferencing for delivering e-health. Future Internet, 4(2):362–371, April 02, 2012. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/4/2/362.

Nguyen:2024:LIN

[NJH^{+}24] Kieu Anh Nguyen, Yi-Jia Jiang, Chiao-Shin Huang, Meng-Hsun Kuo, and Walter Chen. Leveraging Internet news-based data for rockfall hazard susceptibility assessment on highways. Future Internet, 16 (8):??, August 21, 2024. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
8/299.

Nawaz:2024:ETH

[NKSF24] Omer Nawaz, Siamak Khatibi, Muhammad Nauman Sheikh, and Markus Fiedler. Eye tracking and human influence factors’ impact on quality of experience of mobile gaming. Future Internet, 16(11): 420, November 13, 2024. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
11/420.

Nguyen:2024:TSB

[NLD^{+}24] Thi-Hao Nguyen, Van-Hung Le, Huu-Son Do, Trung-Hieu Te, and Van-Nam Phan. TQU-SLAM benchmark dataset for comparative study to build visual odometry based on extracted features from feature descriptors and deep learning. Future Internet, 16(5):174, May 17, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/5/174.

Nguyen:2012:BAS

[NLPV12] Dinh Khoa Nguyen, Francesco Lelli, Mike P. Papazoglou, and Willem-Jan Van den Heuvel. Blueprinting approach in support of cloud computing. Future Internet, 4(1):322–346, March 21, 2012. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/4/1/322.

Ngo:2023:HNH

[NLT^{+}23] Duc-Minh Ngo, Dominic Lightbody, Andriy Temko, Cuong Pham-Quoc, Ngoc-Thinh Tran, Colin C. Murphy, and Emanuel Popovici. HH-NIDS: Heterogeneous hardwarebased network intrusion detection framework for IoT security. Future Internet, 15(1):9, December 26, 2023. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/15/1/9.

Naudts:2021:VCM

[NMH^{+}21] Dries Naudts, Vasilis Maglogiannis, Seilendria Hadiwardoyo, Daniel van den Akker, Simon Vanneste, Siegfried Mercelis, Peter Hellinckx, Bart Lannoo, Johann Marquez-Barja, and Ingrid Moerman. Vehicular communication management framework: a flexible hybrid connectivity platform for CCAM services. Future Internet, 13(3):81, March 22, 2021. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/13/3/81.

Nkemeni:2023:ECR

[NMT23] Valery Nkemeni, Fabien Mieyeville, and Pierre Tsafack. Energy consumption reduction in wireless sensor network-based water pipeline monitoring systems via energy conservation techniques. Future Internet, 15(12):402, December 14, 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
12/402.

Noguez:2021:VHM

[NNRR^{+}21] Julieta Noguez, Luis Neri, Víctor Robledo-Rella, Rosa María Guadalupe García-Castelán, Andres Gonzalez-Nucamendi, David Escobar-Castillejos, and Arturo Molina. VIS-HAPT: a methodology proposal to develop visuohaptic environments in Education 4.0. Future Internet, 13(10):255, October 05, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/10/255.

Nguyen:2022:QMA

[NNT22] Duc Nguyen, Nam Pham Ngoc, and Truong Cong Thang. QoE models for adaptive streaming: a comprehensive evaluation. Future Internet, 14(5): 151, May 13, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
5/151.

Newell:2012:CSC

[NPB12] David A. Newell, Margaret M. Pembroke, and William E. Boyd. Crowd sourcing for conservation: Web 2.0 a powerful tool for biologists. Future Internet, 4(2):551–562, May 24, 2012. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/4/2/551.

Nundloll:2019:DDE

[NPB^{+}19] Vatsala Nundloll, Barry Porter, Gordon S. Blair, Bridget Emmett, Jack Cosby, Davey L. Jones, Dave Chadwick, Ben Winterbourn, Philip Beattie, Graham Dean, Rory Shaw, Wayne Shelley, Mike Brown, and Izhar Ullah. The design and deployment of an endto-end IoT infrastructure for the natural environment. Future Internet, 11(6):129, June 07, 2019. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/11/
                                        6/129.

Nikoloudakis:2018:CSR

[NPM^{+}18] Yannis Nikoloudakis, Spyridon Panagiotakis, Thrasivoulos Manios, Evangelos Markakis, and Evangelos Pallis. Composting as a service: a real-world IoT implementation. Future Internet, 10(11):107, November 05, 2018. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/10/
11/107.

Ntafloukas:2023:VAA

[NPMPM23] Konstantinos Ntafloukas, Liliana Pasquale, Beatriz Martinez-Pastor, and Daniel P. McCrum. A vulnerability assessment approach for transportation networks subjected to cyber–physical attacks. Future Internet, 15(3):100, February 28, 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
3/100.

Norbu:2024:FAT

[NPWC24] Tenzin Norbu, Joo Yeon Park, Kok Wai Wong, and Hui Cui. Factors affecting trust and acceptance for blockchain adoption in digital payment systems: a systematic review. Future Internet, 16(3):106, March 21, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
3/106.

Nunes:2024:AFM

[NRS24] Pedro Nunes, Eugénio Rocha, and José Santos. Adaptive framework for maintenance scheduling based on dynamic preventive intervals and remaining useful life estimation. Future Internet, 16(6):214, June 17, 2024. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/16/
                                        6/214.

Nilsen:2012:TPV

[NS12] Line Lundvoll Nilsen and Terje Solvoll. Traditional practice vs. new tools and routines in stroke treatment. Future Internet, 4(3):688–699, August 06, 2012. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/4/3/688.

Niarchos:2022:SPF

[NSD^{+}22] Michail Niarchos, Marina Eirini Stamatiadou, Charalampos Dimoulas, Andreas Veglis, and Andreas Symeonidis. A semantic preprocessing framework for breaking news detection to support future drone journalism services. Future Internet, 14(1):26, January 10, 2022. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/14/1/26.

Nozari:2021:ISG

[NSJGN21] Hamed Nozari, Agnieszka Szmelter-Jarosz, and Javid Ghahremani-Nahr. The ideas of sustainable and green marketing based on the Internet of Everything — the case of the dairy industry. Future Internet, 13(10):266, October 19, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/10/266.

Nikolaidis:2023:TER

[NST23] Fotis Nikolaidis, Moysis Symeonides, and Demetris Trihinas. Towards efficient resource allocation for federated learning in virtualized managed environments. Future Internet, 15(8):261, July 31, 2023. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
8/261.

Nardini:2017:FRB

[NSV17] Giovanni Nardini, Giovanni Stea, and Antonio Virdis. A fast and reliable broadcast service for LTE-advanced exploiting multihop device-to-device transmissions. Future Internet, 9(4):89, November 25, 2017. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/9/4/89.

Nyblom:2020:RPS

[NWG20] Philip Nyblom, Gaute Wangen, and Vasileios Gkioulos. Risk perceptions on social media use in Norway. Future Internet, 12(12):211, November 26, 2020. CODEN ????

ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/12/12/211.

Neis:2014:RDF

[NZ14] Pascal Neis and Dennis Zielstra. Recent developments and future trends in volunteered geographic information research: The case of OpenStreetMap. Future Internet, 6(1):76– 106, January 27, 2014. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/6/
                                        1/76.

Niyonambaza:2020:PMP

[NZU20] Irene Niyonambaza, Marco Zennaro, and Alfred Uwitonze. Predictive maintenance (PdM) structure using Internet of Things (IoT) for mechanical equipment used into hospitals in rwanda. Future Internet, 12(12):224, December 07, 2020. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/12/12/224.

Neis:2012:SNE

[NZZ12] Pascal Neis, Dennis Zielstra, and Alexander Zipf. The street network evolution of crowdsourced maps: OpenStreetMap in Germany 2007–2011. Future Internet, 4(1):1–21, December 29, 2012. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.

mdpi.com/1999-5903/4/
1/1.

Neis:2013:CVG

[NZZ13] Pascal Neis, Dennis Zielstra, and Alexander Zipf. Comparison of volunteered geographic information data contributions and community development for selected world regions. Future Internet, 5(2):282– 300, June 03, 2013. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/5/
2/282.

Ojo:2021:EEL

[OAG21] Mike Oluwatayo Ojo, Davide Adami, and Stefano Giordano. Experimental evaluation of a LoRa wildlife monitoring network in a forest vegetation area. Future Internet, 13(5):115, April 29, 2021. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
5/115.

Owoh:2024:MDB

[OAH^{+}24] Nsikak Owoh, John Adejoh, Salaheddin Hosseinzadeh, Moses Ashawa, Jude Osamor, and Ayyaz Qureshi. Malware detection based on API call sequence analysis: a gated recurrent unit–generative adversarial network model approach. Future Internet, 16(10), October 13, 2024.

CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/16/
                                        10/369.

Omasheye:2022:JRF

[OAI^{+}22] Okiemute Roberts Omasheye, Samuel Azi, Joseph Isabona, Agbotiname Lucky Imoize, Chun-Ta Li, and Cheng-Chi Lee. Joint random forest and particle swarm optimization for predictive pathloss modeling of wireless signals from cellular networks. Future Internet, 14(12):373, December 12, 2022. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/14/
                                        12/373.

Ogul:2013:PAM

[OB13] Murat Ogul and Selçuk Baktir. Practical attacks on mobile cellular networks and possible countermeasures. Future Internet, 5(4):474–489, September 30, 2013. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/5/4/474.

Orehovacki:2023:EPQ

[OBK23] Tihomir Orehovacki, Luka Blasković, and Matej Kurevija. Evaluating the perceived quality of mobile banking applications in Croatia: an empirical study. Future Internet, 15 (1):8, December 26, 2023.

CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
1/8.

Olszowski:2020:CIP

[OC20] Rafał Olszowski and Marcin Chmielowski. Collective intelligence in Polish– Ukrainian Internet projects. Debate models and research methods. Future Internet, 12(6):106, June 20, 2020. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/12/6/106.

Oliveira:2020:FME

[OCCB20] João Oliveira, Gonçalo Carvalho, Bruno Cabral, and Jorge Bernardino. Failure mode and effect analysis for cyber-physical systems. Future Internet, 12(11):205, November 20, 2020. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/12/11/205.

Ogata:2023:DSA

[OF23] Keigo Ogata and Satoshi Fujita. Decentralized storage with access control and data persistence for ebook stores. Future Internet, 15(12):406, December 18, 2023. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/15/12/406.

Office:2014:ARF

[Off14] Future Internet Editorial Office. Acknowledgement to reviewers of future Internet in 2013. Future Internet, 6(1):107– 108, February 24, 2014. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/6/
                                        1/107.

Office:2015:ARF

[Off15] Future Internet Editorial Office. Acknowledgement to reviewers of Future Internet in 2014. Future Internet, 7(1):24–25, January 09, 2015. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/7/1/24.

Office:2016:ARF

[Off16] Future Internet Editorial Office. Acknowledgement to reviewers of Future Internet in 2015. Future Internet, 8(1):3, January 22, 2016. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/8/1/3.

Office:2017:ARF

[Off17] Future Internet Editorial Office. Acknowledgement to reviewers of Future Internet in 2016. Future Internet, 9(1):2, January 12, 2017. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/9/1/2.

Office:2018:ARF

[Off18] Future Internet Editorial Office. Acknowledgement to reviewers of Future Internet in 2017. Future Internet, 10(1):7, January 12, 2018. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/10/1/7.

Office:2019:ARF

[Off19] Future Internet Editorial Office. Acknowledgement to reviewers of Future Internet in 2018. Future Internet, 11(1):14, January 10, 2019. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/11/1/14.

Office:2022:ARF

[Off22] Future Internet Editorial Office. Acknowledgment to reviewers of Future Internet in 2021. Future Internet, 14(2):53, February 10, 2022. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/14/2/53.

Orogun:2019:SCS

[OFK19] Adebola Orogun, Oluwaseun Fadeyi, and Ondrej Krejcar. Sustainable communication systems: a graphlabeling approach for cellular frequency allocation in densely-populated areas. Future Internet, 11 (9):186, August 26, 2019. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
9/186.

Oikonomidis:2021:MAR

[OFV21] Theodoros Oikonomidis, Konstantinos Fouskas, and Maro Vlachopoulou. A multidimensional analysis of released COVID-19 location-based mobile applications. Future Internet, 13(11):268, October 25, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/11/268.

Ostrikova:2024:DSS

[OGB^{+}24] Darya Ostrikova, Elizaveta Golos, Vitalii Beschastnyi, Egor Machnev, Yuliya Gaidamaka, and Konstantin Samouylov. Dynamic SNR, spectral efficiency, and rate characterization in 5G/6G mmWave/sub-THz systems with macro- and micro-mobilities. Future Internet, 16(7):??, July 6, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/7/240.

Oliveira:2020:MCI

[OGR^{+}20] Thays A. Oliveira, Yuri B. Gabrich, Helena Ramalhinho, Miquel Oliver, Miri W. Cohen, Luiz S. Ochi, Serigne Gueye, Fábio Protti, Alysson A. Pinto, Diógenes V. M. Ferreira, Igor M. Coelho, and Vitor N. Coelho. Mobility, citizens, innovation and technology in digital and smart cities. Future Internet, 12(2):22, January 26, 2020. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/12/2/22.

Otoum:2024:EHD

[OHSN24] Yazan Otoum, Chaosheng Hu, Eyad Haj Said, and Amiya Nayak. Enhancing heart disease prediction with federated learning and blockchain integration. Future Internet, 16(10), October 14, 2024. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
10/372.

Oie:2012:SNU

[Øie12] Kjetil Vaage Øie. Sensing the news: User experiences when reading locative news. Future Internet, 4(1):161–178, February 21, 2012. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/4/1/161.

Ober:2013:SAB

[OKH13] Micha Ober, Stefan Katzenbeisser, and Kay Hamacher. Structure and anonymity of the Bitcoin transaction graph. Future Internet, 5 (2):237–250, May 07, 2013. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/5/
                                        2/237.

Olmedo:2020:PAN

[OLCMdA20] Gonzalo Olmedo, Román Lara-Cueva, Diego Martínez, and Celso de Almeida. Performance analysis of a novel TCP protocol algorithm adapted to wireless networks. Future Internet, 12(6):101, June 09, 2020. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/12/
                                        6/101.

Ocampo-Lopez:2023:PDL

[OLMBRCRC23] Carlos Ocampo-López, Oscar Muñoz-Blandón, Leidy Rendón-Castrillón, and Margarita Ramírez-Carmona. Post-digital learning for rural development: a case study of Open Biotec MOOCs in Colombia. Future Internet, 15(4):141, April 06, 2023. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/15/
                                        4/141.

Ogundokun:2022:MIT

[OMD^{+}22] Roseline Oluwaseun Ogundokun, Sanjay Misra, Mychal Douglas, Robertas Damasevicius, and Rytis Maskeliunas. Medical Internet-of-Things based breast cancer diagnosis using hyperparameteroptimized neural networks. Future Internet, 14(5):153, May 18, 2022. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
5/153.

Ortiz-Marcos:2021:CAI

[OMTFFL21] José Manuel Ortiz-Marcos, María Tomé-Fernández, and Christian Fernández-Leyva. Cyberbullying analysis in intercultural educational environments using binary logistic regressions. Future Internet, 13(1):15, January 09, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/1/15.

Ou:2023:DCF

[ON23] Yuanyou Ou and Baoning Niu. Dual-channel feature enhanced collaborative filtering recommendation algorithm. Future Internet, 15(6):215–??, June 2023. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
6/215.

ODonovan:2019:SAR

[OO19] Peter O’Donovan and Dominic T. J. O’Sullivan. A systematic analysis of realworld energy blockchain initiatives. Future Internet, 11(8):174, August 10, 2019. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/11/8/174.

Oliveira:2018:LMP

[OOM^{+}18] Ewerton L. S. Oliveira, Davide Orrù, Luca Morreale, Tiago P. Nascimento, and Andrea Bonarini. Learning and mining player motion profiles in physically interactive robogames. Future Internet, 10(3):22, February 26, 2018. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/10/3/22.

Obonna:2023:DMM

[OOM^{+}23]  Ugochukwu  Onyekachi

Obonna, Felix Kelechi Opara, Christian Chidiebere Mbaocha, Jude-Kennedy Chibuzo Obichere, Isdore Onyema Akwukwaegbu, Miriam Mmesoma Amaefule, and Cosmas Ifeanyi Nwakanma. Detection of man-in-the-middle (MitM) cyberattacks in oil and gas process control networks using machine learning algorithms. Future Internet, 15 (8):280, August 21, 2023. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
8/280.

Osial:2022:AAI

[OP22] Magdalena Osial and Agnieszka Pregowska. The application of artificial intelligence in magnetic hyperthermia based research. Future Internet, 14(12): 356, November 28, 2022. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
12/356.

Otta:2023:SSM

[OPGH23] Soumya Prakash Otta, Subhrakanta Panda, Maanak Gupta, and Chittaranjan Hota. A systematic survey of multi-factor authentication for cloud infrastructure. Future Internet, 15(4):146, April 10, 2023. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
4/146.

Owen:2024:VDI

[OR24] Aron E. Owen and Jonathan C.

Roberts. Visualisation design ideation with AI: a new framework, vocabulary, and tool. Future Internet, 16(11):406, November 5, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/11/406.

Oliveira:2019:MLP

[ORK^{+}19] Luiz Oliveira, Joel J. P. C.

Rodrigues, Sergei A. Kozlov, Ricardo A. L. Rabêlo, and Victor Hugo C. de Albuquerque. MAC layer protocols for Internet of Things: a survey. Future Internet, 11(1):16, January 14, 2019. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/11/1/16.

Raviv:2019:SMU

[oRL19] Li on Raviv and Amir Leshem. Scheduling for multi-user multi-input multi-output wireless networks with priorities and deadlines. Future Internet, 11(8):172, August 05, 2019. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/11/8/172.

Orlova:2022:DTA

[Orl22] Ekaterina V. Orlova. Design technology and AIbased decision making model for digital twin engineering. Future Internet, 14(9):248, August 24, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/9/248.

ONeill:2023:SMC

[OS23] Vyas O’Neill and Ben Soh. Spot market cloud orchestration using taskbased redundancy and dynamic costing. Future Internet, 15(9):288, August 27, 2023. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/15/9/288.

Orrillo:2024:ERA

[OSdS24] Héctor Orrillo, André Sabino, and Mário Marques da Silva. Evaluation of radio access protocols for V2X in 6G scenariobased models. Future Internet, 16(6):203, June 6, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/6/203.

Okamoto:2016:RBI

[OT16] Kazuaki Okamoto and Kazumasa Takami. Routing based on information about the routes of fixedroute traveling nodes and on destination areas aimed at reducing the load on the DTN. Future Internet, 8 (2):15, April 27, 2016. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/8/
2/15.

Onorati:2020:PLD

[OTR^{+}20] Dario Onorati, Pierfrancesco Tommasino, Leonardo Ranaldi, Francesca Fallucchi, and Fabio Massimo Zanzotto. Pat-in-the-loop: Declarative knowledge for controlling neural networks. Future Internet, 12(12):218, December 02, 2020. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/12/12/218.

Orkphol:2019:WSD

[OY19] Korawit Orkphol and Wu Yang. Word sense disambiguation using cosine similarity collaborates with Word2vec and Word-Net. Future Internet, 11 (5):114, May 12, 2019. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/11/
                                        5/114.

Ozadowicz:2023:TQE

[Oza23] Andrzej Ozadowicz. Technical, qualitative and energy analysis of wireless control modules for distributed smart home systems. Future Internet, 15(9):316, September 20, 2023. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/15/9/316.

Pileggi:2013:ASG

[PA13] Salvatore F. Pileggi and Robert Amor. Addressing semantic geographic information systems. Future Internet, 5(4):585– 590, November 26, 2013. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/5/
                                        4/585.

Patsias:2023:TAM

[PAK^{+}23] Vasilios Patsias, Petros Amanatidis, Dimitris Karampatzakis, Thomas Lagkas, Kalliopi Michalakopoulou, and Alexandros Nikitas. Task allocation methods and optimization techniques in edge computing: A systematic review of the literature. Future Internet, 15(8):254, July 28, 2023. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
8/254.

Peivandizadeh:2024:SKE

[PAM^{+}24] Ali Peivandizadeh, Haitham Y.

Adarbah, Behzad Molavi, Amirhossein Mohajerzadeh, and Ali H. Al-Badi. A secure key exchange and authentication scheme for securing communications in the Internet of Things environment. Future Internet, 16(10), September 30, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/10/357.

Papavassiliou:2020:SDN

[Pap20]  Symeon  Papavassiliou.

Software defined networking (SDN) and network function virtualization (NFV). Future Internet, 12(1):7, January 02, 2020. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/12/1/7.

Pethers:2023:RAD

[PB23] Brent Pethers and Abubakar Bello. Role of attention and design cues for influencing cyber-sextortion using social engineering and phishing attacks. Future Internet, 15(1):29, January 07, 2023. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/15/
                                        1/29.

Parchin:2019:DBM

[PBAAN19] Naser Ojaroudi Parchin, Haleh Jahanbakhsh Basherlou, Raed A. Abd-Alhameed, and James M. Noras. Dual-band monopole antenna for RFID applications. Future Internet, 11 (2):31, January 30, 2019. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/11/
                                        2/31.

Pacini:2020:SCS

[PBB20] Giovanna Pacini, Cinzia Belmonte, and Franco Bagnoli. Science cafés, science shops and the lockdown experience in Florence and Rome. Future Internet, 12(7):115, July 08, 2020. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/12/7/115.

Pflanzner:2022:LAB

[PBK22] Tamas Pflanzner, Hamza Baniata, and Attila Kertesz. Latency analysis of blockchainbased SSI applications. Future Internet, 14(10): 282, September 29, 2022. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
10/282.

Paraschiv:2024:ADN

[PBP^{+}24]  Elena-Anca  Paraschiv,

Lidia Bajenaru, Cristian Petrache, Ovidiu Bica, and Dragos-Nicolae Nicolau. AI-driven neuromonitoring: Advancing schizophrenia detection and management through deep learning and EEG analysis. Future Internet, 16(11):424, November 16, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/11/424.

Peng:2023:HNO

[PC23] Chao-Chung Peng and Yi-Ho Chen. A hybrid neural ordinary differential equation based digital twin modeling and online diagnosis for an industrial cooling fan. Future Internet, 15(9):302, September 04, 2023. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/15/9/302.

Papazis:2024:HFM

[PC24] Kon Papazis and Naveen Chilamkurti. Hydrakon, a framework for measuring indicators of deception in emulated monitoring systems. Future Internet, 16(12):455, December 4, 2024. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/16/12/455.

Pau:2018:SPC

[PCC^{+}18] Giovanni Pau, Tiziana Campisi, Antonino Canale, Alessandro Severino, Mario Collotta, and Giovanni Tesoriere. Smart pedestrian crossing management at traffic light junctions through a fuzzy-based approach. Future Internet, 10(2):15, February 01, 2018. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/10/2/15.

Plachouras:2014:ACA

[PCF^{+}14] Vassilis Plachouras, Florent Carpentier, Muhammad Faheem, Julien Masanès, Thomas Risse, Pierre Senellart, Patrick Siehndel, and Yannis Stavrakas. ARCOMEM crawling architecture. Future Internet, 6(3):518–541, August 19, 2014. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/6/3/518.

Petrolini:2022:ADS

[PCM22] Michael Petrolini, Stefano Cagnoni, and Monica Mordonini. Automatic detection of sensitive data using transformer-based classifiers. Future Internet, 14 (8):228, July 27, 2022. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
8/228.

Piccardo:2024:BTP

[PCM24] Giorgio Piccardo, Lorenzo Conti, and Alessio Martino. Blockchain technology and its potential to benefit public services provision: a short survey. Future Internet, 16(8):??, August 9, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
8/290.

Padilla-Cuevas:2021:OBC

[PCROB21] Josué Padilla-Cuevas, José A.

Reyes-Ortiz, and Maricela Bravo. Ontology-based context event representation, reasoning, and enhancing in academic environments. Future Internet, 13(6):151, June 10, 2021. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
6/151.

Papageorgiou:2024:SUL

[PCVH24] Eleftheria Papageorgiou, Christos Chronis, Iraklis Varlamis, and Yassine Himeur. A survey on the use of large language models (LLMs) in fake news. Future Internet, 16(8):??, August 19, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
8/298.

Pticek:2023:MAI

[PD23] Martina Pticek and Jasminka Dobsa. Methods of annotating and identifying metaphors in the field of natural language processing. Future Internet, 15(6): 201–??, June 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
6/201.

Peker:2021:ECE

[PDI21] Serhat Peker, Gonca Gokce Menekse Dalveren, and Yavuz Inal. The effects of the content elements of online banner ads on visual attention: Evidence from aneye-tracking study. Future Internet, 13(1):18, January 15, 2021. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/13/1/18.

Pibernik:2019:EFA

[PDMK19] Jesenka Pibernik, Jurica Dolic, Hrvoje Abraham Milicevic, and Bojan Kanizaj. The effects of the floating action button on quality of experience. Future Internet, 11(7):148, July 06, 2019. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
7/148.

Pisani:2019:FVC

[PdRB19] Flávia Pisani, Vanderson Martins do Rosario, and Edson Borin. Fog vs. cloud computing: Should I stay or should I go? Future Internet, 11(2):34, February 02, 2019. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/11/2/34.

Papathanasiou:2023:MFP

[PDT23] Dimitris Papathanasiou, Konstantinos Demertzis, and Nikos Tziritas. Machine failure prediction using survival analysis. Future Internet, 15(5):153, April 22, 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
5/153.

Pahlevan:2015:DEC

[PDTM15] Atousa Pahlevan, Jean-Luc Duprat, Alex Thomo, and Hausi Müller. Dynamis: Effective contextaware Web service selection using dynamic attributes. Future Internet, 7 (2):110–139, May 12, 2015. CODEN ???? ISSN 1999-

5903. URL https://www.

                                        mdpi.com/1999-5903/7/
                                        2/110.

Paltrinieri:2013:PIE

[PE13] Roberta Paltrinieri and Piergiorgio Degli Esposti. Processes of inclusion and exclusion in the sphere of prosumerism. Future Internet, 5(1):21–33, January 10, 2013. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/5/1/21.

Pecori:2018:VLA

[Pec18] Riccardo Pecori. A virtual learning architecture enhanced by fog computing and big data streams. Future Internet, 10(1):4, January 03, 2018. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/10/1/4.

Perifanou:2023:TVI

[PEN23] Maria Perifanou, Anastasios A. Economides, and Stavros A. Nikou. Teachers’ views on integrating augmented reality in education: Needs, opportunities, challenges and recommendations. Future Internet, 15(1):20, December 29, 2023. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/15/1/20.

Petrie:2011:ECI

[Pet11] Charles Petrie. Enterprise coordination on the Internet. Future Internet, 3(1):49–66, February 17, 2011. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/3/1/49.

Petrov:2020:ACC

[Pet20] Mikhail Petrov. An approach to changing competence assessment for human resources in expert networks. Future Internet, 12(10):169, October 03, 2020. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/12/10/169.

Pena-Fernandez:2021:REE

[PFdRGG21] Simón Peña-Fernández, Miguel Ángel Casado del Río, and Daniel García-González. From rigidity to exuberance: Evolution of news on online newspaper homepages. Future Internet, 13(6):150, June 09, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/6/150.

Pollino:2012:COS

[PFL^{+}12] Maurizio Pollino, Grazia Fattoruso, Luigi La Porta, Antonio Bruno Della Rocca, and Valentina James. Collaborative open source geospatial tools and maps supporting the response planning to disastrous earthquake events.

Future Internet, 4(2):451– 468, May 07, 2012. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/4/
2/451.

Pileggi:2012:WSM

[PFLT12] Salvatore F. Pileggi, Carlos Fernandez-Llatas, and Vicente Traver. When the social meets the semantic: Social Semantic Web or Web 2.5. Future Internet, 4(3):852–864, September 21, 2012. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/4/3/852.

Paxinou:2024:TSA

[PFT^{+}24] Evgenia Paxinou, Georgios Feretzakis, Rozita Tsoni, Dimitrios Karapiperis, Dimitrios Kalles, and Vassilios S. Verykios. Tracing student activity patterns in e-learning environments: Insights into academic performance. Future Internet, 16(6): 190, May 29, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
6/190.

Papadopoulou:2014:CTK

[PG14] Chrysaida-Aliki Papadopoulou and Maria Giaoutzi. Crowdsourcing as a tool for knowledge acquisition in spatial planning. Future Internet, 6(1):109– 125, March 05, 2014. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/6/
1/109.

Pelle:2019:EAF

[PG19] István Pelle and András Gulyás. An extensible automated failure localization framework using NetKAT, Felix, and SDN traceroute. Future Internet, 11(5):107, May 04, 2019. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/11/5/107.

Paasche:2023:FSA

[PG23] Simon Paasche and Sven Groppe. A finite state automaton for green data validation in a real-world smart manufacturing environment with special regard to time-outs and overtaking. Future Internet, 15(11):349, October 26, 2023. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/15/11/349.

Papenfuss:2024:EEE

[PGFH24] Dennis Papenfuß, Bennet Gerlach, Stefan Fischer, and Mohamed Ahmed Hail. Enhancing energy efficiency in IoT-NDN via parameter optimization. Future Internet, 16(2):61, February 16, 2024. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/16/
                                        2/61.

Pombo:2020:ELC

[PGSL20] Iñigo Pombo, Leire Godino, Jose Antonio Sánchez, and Rafael Lizarralde. Expectations and limitations of cyber-physical systems (CPS) for advanced manufacturing: A view from the grinding industry. Future Internet, 12(9):159, September 22, 2020. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/12/
                                        9/159.

Premnath:2016:SPC

[PH16] Sriram Nandha Premnath and Zygmunt J. Haas. Supporting privacy of computations in mobile big data systems. Future Internet, 8(2):17, May 10, 2016. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/8/2/17.

Papa:2020:SDM

[PH20] Louis Edward Papa and Thaier Hayajneh. A survey of defensive measures for digital persecution in the global south. Future Internet, 12(10):166, September 29, 2020. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/12/
                                        10/166.

Pires:2020:HDN

[PHG^{+}20] Ivan Miguel Pires, Faisal Hussain, Nuno M. Garcia, Petre Lameski, and Eftim Zdravevski. Homogeneous data normalization and deep learning: a case study in human activity classification. Future Internet, 12(11):194, November 10, 2020. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/12/
11/194.

Pires:2020:IHA

[PHGZ20] Ivan Miguel Pires, Faisal Hussain, Nuno M. Garcia, and Eftim Zdravevski. Improving human activity monitoring by imputation of missing sensory data: Experimental study. Future Internet, 12(9):155, September 17, 2020. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/12/
9/155.

Pickering:2021:TVI

[Pic21] Brian Pickering. Trust, but verify: Informed consent, AI technologies, and public health emergencies. Future Internet, 13 (5):132, May 18, 2021. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
5/132.

Pilloni:2018:HDW

[Pil18] Virginia Pilloni. How data will transform industrial processes: Crowdsensing, crowdsourcing and big data as pillars of Industry 4.0. Future Internet, 10(3):24, March 01, 2018. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/10/
                                        3/24.

Pinart:2010:ATD

[Pin10] Carolina Pinart. Anticipation of traffic demands to guarantee QoS in IP/ optical networks. Future Internet, 2(3):417– 430, September 21, 2010. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/2/
                                        3/417.

Pires:2022:SOT

[Pir22] Ivan Miguel Pires. Smart objects and technologies for social good. Future Internet, 14(12):370, December 09, 2022. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/14/
                                        12/370.

Perez-Jorge:2022:GSA

[PJMM22] David Pérez-Jorge and María Carmen Martínez-Murciano. Gamification with scratch or app inventor in higher education: a systematic review. Future Internet, 14(12):374, December 13, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
12/374.

Pieska:2024:PIN

[PKB^{+}24] Marcus Pieska, Andreas Kassler, Anna Brunstrom, Veselin Rakocevic, and Markus Amend. Performance impact of nested congestion control on transport-layer multipath tunneling. Future Internet, 16(7):??, June 28, 2024. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
7/233.

Pokrovskaia:2021:DRI

[PKCF21] Nadezhda N. Pokrovskaia, Olga N. Korableva, Lucio Cappelli, and Denis A. Fedorov. Digital regulation of intellectual capital for open innovation: Industries’ expert assessments of tacit knowledge for controlling and networking outcome. Future Internet, 13(2):44, February 10, 2021. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/13/2/44.

Papadopoulos:2024:RAF

[PKF24] Christos Papadopoulos, Konstantinos-Filippos Kollias, and George F. Fragulis. Recent advancements in federated learning: State of the art, fundamentals, principles, IoT applications and future trends. Future Internet, 16(11):415, November 9, 2024. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/16/11/415.

Panwar:2023:BAE

[PKMS23] Arvind Panwar, Manju Khari, Sanjay Misra, and Urvashi Sugandh. Blockchain in agriculture to ensure trust, effectiveness, and traceability from farm fields to groceries. Future Internet, 15(12): 404, December 16, 2023. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/15/
                                        12/404.

Pokhrel:2024:DME

[PKS^{+}24] Shiva Raj Pokhrel, Jonathan Kua, Deol Satish, Sebnem Ozer, Jeff Howe, and Anwar Walid. DDPG-MPCC: an experience driven multipath performance oriented congestion control. Future Internet, 16(2):37, January 23, 2024. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/16/
                                        2/37.

Park:2022:CBH

[PL22a] Minwoo Park and Euichul Lee. Correlation between human emotion and temporal–spatial contexts by analyzing environmental factors. Future Internet, 14(7):203, June 30, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/7/203.

Pedone:2022:QKD

[PL22b] Ignazio Pedone and Antonio Lioy. Quantum key distribution in Kubernetes clusters. Future Internet, 14(6):160, May 25, 2022. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
6/160.

Park:2017:CMR

[PLA17] Kyungmo Park, Sanghyo Lee, and Yonghan Ahn. Construction Management Risk System (CMRS) for construction management (CM) firms. Future Internet, 9(1):5, February 10, 2017. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/9/1/5.

Pokrovskaia:2021:DCT

[PLA^{+}21] Nadezhda N. Pokrovskaia, Veronika L. Leontyeva, Marianna Yu. Ababkova, Lucio Cappelli, and Fabrizio D’Ascenzo. Digital communication tools and knowledge creation processes for enriched intellectual outcome-experience of short-term e-learning courses during pandemic. Future Internet, 13(2):43, February 05, 2021. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/13/
                                        2/43.

Pellegrino:2024:MHL

[PLA^{+}24] Mattia Pellegrino, Gianfranco Lombardo, George Adosoglou, Stefano Cagnoni, Panos M. Pardalos, and Agostino Poggi. A multihead LSTM architecture for bankruptcy prediction with time series accounting data. Future Internet, 16(3):79, February 27, 2024. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/16/3/79.

Pliatsios:2023:SSS

[PLG23] Antonios Pliatsios, Dimitrios Lymperis, and Christos Goumopoulos. S2NetM: a semantic social network of things middleware for developing smart and collaborative IoT-based solutions. Future Internet, 15(6):207–??, June 2023. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/15/
                                        6/207.

Parsons:2011:TDD

[PLL11] David Parsons, Ramesh Lal, and Manfred Lange.

Test driven development: Advancing knowledge by conjecture and confirmation. Future Internet, 3(4):281–297, December 14, 2011. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/3/4/281.

Pan:2020:PLB

[PLL20] Wei Pan, Jide Li, and Xiaoqiang Li. Portfolio learning based on deep learning. Future Internet, 12(11):202, November 18, 2020. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/12/11/202.

Pinna:2023:ISA

[PLOT23] Andrea Pinna, Maria Ilaria Lunesu, Stefano Orrù, and Roberto Tonelli. Investigation on self-admitted technical debt in opensource blockchain projects. Future Internet, 15(7): 232–??, July 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
7/232.

Perger:2014:GGW

[PLSF14] Christoph Perger, Ellsworth LeDrew, Linda See, and Steffen Fritz. Geography Geo-Wiki in the classroom: Using crowdsourcing to enhance geographical teaching. Future Internet, 6(4):597–611, September 29, 2014. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/6/4/597.

Pan:2020:MSN

[PLYD20] Yirong Pan, Xiao Li, Yating Yang, and Rui Dong. Multi-source neural model for machine translation of agglutinative language. Future Internet, 12(6):96, June 03, 2020. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/12/6/96.

Pruvost:2017:EDM

[PM17] Hippolyte Pruvost and Peter Mooney. Exploring data model relations in OpenStreetMap. Future Internet, 9(4):70, October 24, 2017. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/9/4/70.

Patwe:2024:BES

[PM24] Sonali Patwe and Sunil B. Mane. Blockchain-enabled secure and interoperable authentication scheme for metaverse environments. Future Internet, 16(5): 166, May 11, 2024. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/16/
                                        5/166.

Papadopoulou:2022:MNN

[PMA^{+}22] Olga Papadopoulou, Themistoklis Makedas, Lazaros Apostolidis, Francesco Poldi, Symeon Papadopoulos, and Ioannis Kompatsiaris. MeVer NetworkX: Network analysis and visualization for tracing disinformation. Future Internet, 14(5):147, May 10, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/5/147.

Papamartzivanos:2021:PMC

[PMGG21]  Dimitrios  Papamartzi-

vanos, Sofia Anna Menesidou, Panagiotis Gouvas, and Thanassis Giannetsos. A perfect match: Converging and automating privacy and security impact assessment on-the-fly. Future Internet, 13(2): 30, January 27, 2021. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
2/30.

Panda:2022:NLI

[PMP22a] Padmalochan Panda, Alekha Kumar Mishra, and Deepak Puthal. A novel logo identification technique for logo-based phishing detection in cyber-physical systems. Future Internet, 14 (8):241, August 15, 2022. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
8/241.

Ponce:2022:SFA

[PMP^{+}22b] Pedro Ponce, Omar Mata, Esteban Perez, Juan Roberto Lopez, Arturo Molina, and Troy McDaniel. S4 features and artificial intelligence for designing a robot against COVID-19-robocov. Future Internet, 14(1):22, January 06, 2022. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/14/1/22.

Patrickson:2023:DCD

[PMT^{+}23] Bronwin Patrickson, Mike Musker, Dan Thorpe, Yasmin van Kasteren, and Niranjan Bidargaddi. Indepth co-design of mental health monitoring technologies by people with lived experience. Future Internet, 15(6):191– ??, June 2023. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/15/
                                        6/191.

Poleto:2023:ISA

[PNdC^{+}23] Thiago Poleto, Thyago Celso Cavalcante Nepomuceno, Victor Diogho Heuer de Carvalho, Ligiane Cristina Braga de Oliveira Friaes, Rodrigo Cleiton Paiva de Oliveira, and Ciro José Jardim Figueiredo. Information security applications in smart cities: a bibliometric analysis of emerging research. Future Internet, 15(12):393, December 01, 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
12/393.

Psomakelis:2020:SSD

[PNM^{+}20]  Evangelos  Psomakelis,

Anastasios Nikolakopoulos, Achilleas Marinakis, Alexandros Psychas, Vrettos Moulos, Theodora Varvarigou, and Andreas Christou. A scalable and semantic data as a service marketplace for enhancing cloud-based applications. Future Internet, 12(5):77, April 25, 2020. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/12/5/77.

Paajala:2022:UPK

[PNMK22] Iikka Paajala, Jesse Nyyssölä, Juho Mattila, and Pasi Karppinen. Users’ perceptions of key blockchain features in games. Future Internet, 14(11):321, November 04, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
11/321.

Pashazadeh:2023:CSE

[PNS23] Ali Pashazadeh, Giovanni Nardini, and Giovanni Stea. A comprehensive survey exploring the multifaceted interplay between mobile edge computing and vehicular networks. Future Internet, 15(12): 391, November 30, 2023. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/15/
                                        12/391.

Poulter:2020:ESU

[POC20] Andrew John Poulter, Steven J. Ossont, and Simon J. Cox. Enabling the secure use of dynamic identity for the Internet of Things — using the Secure Remote Update Protocol (SRUP). Future Internet, 12(8):138, August 18, 2020. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/12/8/138.

Portela:2021:DSK

[Por21] Filipe Portela. Data science and knowledge discovery. Future Internet, 13(7):178, July 07, 2021. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/13/
                                        7/178.

Pelekoudas-Oikonomou:2023:PHF

[PORM^{+}23] Filippos Pelekoudas-Oikonomou, José C. Ribeiro, Georgios Mantas, Georgia Sakellari, and Jonathan Gonzalez. Prototyping a hyperledger fabric-based security architecture for IoMTbased health monitoring systems. Future Internet, 15(9):308, September 11, 2023. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/15/9/308.

Pozzebon:2024:EFC

[Poz24] Alessandro Pozzebon. Edge and fog computing for the Internet of Things. Future Internet, 16(3):101, March 16, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
3/101.

Patil:2022:DBD

[PP22a] Pranita Patil and Kevin Purcell. Decorrelationbased deep learning for bias mitigation. Future Internet, 14(4):110– ??, March 29, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
4/110.

Polyakov:2022:ALP

[PP22b] Nikita Polyakov and Anna Platonova. Assessing latency of packet delivery in the 5G 3GPP integrated access and backhaul architecture with half-duplex constraints. Future Internet, 14(11):345, November 21, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/11/345.

Pirotti:2022:ITR

[PPDC22] Francesco Pirotti, Marco Piragnolo, Marika D’Agostini, and Raffaele Cavalli. Information technologies for real-time mapping of human well-being indicators in an urban historical garden. Future Internet, 14(10):280, September 29, 2022. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/14/10/280.

Paradiso:2016:CBE

[PPGC16] Francesca Paradiso, Federica Paganelli, Dino Giuli, and Samuele Capobianco. Context-based energy disaggregation in smart homes. Future Internet, 8(1):4, January 27, 2016. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/8/1/4.

Paakkonen:2021:AEE

[PPKS21] Pekka Pääkkönen, Daniel Pakkala, Jussi Kiljander, and Roope Sarala. Architecture for enabling edge inference via model transfer from cloud domain in a Kubernetes environment. Future Internet, 13(1):5, December 29, 2021. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/13/
                                        1/5.

Polasko:2021:IMU

[PPM21] Ken Polasko, Pedro Ponce, and Arturo Molina. An income model using historical data, power-law distributions and Monte Carlo method for university technology transfer offices. Future Internet, 13 (5):122, May 06, 2021. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
5/122.

Parada:2019:RBE

[PPMMS19] Raúl Parada, Alfonso Palazón, Carlos Monzo, and Joan Melià-Seguí. RFID based embedded system for sustainable food management in an IoT network paradigm. Future Internet, 11(9):189, September 01, 2019. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
9/189.

Pecorella:2018:NSF

[PPN18] Tommaso Pecorella, Laura Pierucci, and Francesca Nizzi. ”Network sentiment” framework to improve security and privacy for smart home. Future Internet, 10(12):125, December 19, 2018. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/10/
12/125.

Popescu:2021:ISR

[PPP21] Traian Mihai Popescu, Alina Madalina Popescu, and Gabriela Prostean. IoT security risk management strategy reference model (IoTSRM2). Future Internet, 13(6):148, June 04, 2021. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/13/6/148.

Phan:2023:ISM

[PPP^{+}23] Vu Hien Phan, Danh Phan Hong Pham, Tran Vu Pham, Kashif Naseer Qureshi, and Cuong Pham-Quoc. An IoT system and MODIS images enable smart environmental management for Mekong Delta. Future Internet, 15 (7):245–??, July 2023. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/15/
                                        7/245.

Paez:2020:ABE

[PPR^{+}20] Rafael Páez, Manuel Pérez, Gustavo Ramírez, Juan Montes, and Lucas Bouvarel. An architecture for biometric electronic identification document system based on blockchain. Future Internet, 12(1):10, January 11, 2020. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/12/1/10.

Paar:2011:LCM

[PR11] Philip Paar and Jörg Rekittke. Low-cost mapping and publishing methods for landscape architectural analysis and design in slum-upgrading projects. Future Internet, 3(4):228–247, October 20, 2011. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/3/4/228.

Prandini:2012:RRA

[PR12] Marco Prandini and Marco Ramilli. Raising risk awareness on the adoption of Web 2.0 technologies in decision making processes. Future Internet, 4(3):700– 718, August 09, 2012. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/4/
3/700.

Ponce:2020:ASO

[PR20] Alan Ponce and Raul Alberto Ponce Rodriguez. An analysis of the supply of open government data. Future Internet, 12(11): 186, October 29, 2020. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/12/
11/186.

Pathmaperuma:2022:CUA

[PRDK22] Madushi H. Pathmaperuma, Yogachandran Rahulamathavan, Safak Dogan, and Ahmet Kondoz. CNN for user activity detection using encrypted inapp mobile data. Future Internet, 14(2):67, February 21, 2022. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/14/2/67.

Priscoli:2010:FCA

[Pri10] Francesco Delli Priscoli. A fully cognitive approach for future Internet. Future Internet, 2(1):16–29, January 22, 2010. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/2/1/16.

Priyadarshini:2023:AST

[Pri23] Ishaani Priyadarshini. Autism screening in toddlers and adults using deep learning and fair AI techniques. Future Internet, 15(9):292, August 28, 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
9/292.

Pinto:2022:BDC

[PRS22] Flavio Pinto, Yogachandran Rahulamathavan, and James Skinner. Blockchain for doping control applications in sports: A conceptual approach. Future Internet, 14(7):210, July 14, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/7/210.

Pandey:2016:UFI

[PS16] Pankaj Pandey and Einar Snekkenes. Using financial instruments to transfer the information security risks. Future Internet, 8(2):20, May 17, 2016. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/8/2/20.

Piezzo:2017:FSS

[PS17] Chiara Piezzo and Kenji Suzuki. Feasibility study of a socially assistive humanoid robot for guiding elderly individuals during walking. Future Internet, 9 (3):30, July 08, 2017. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/9/
3/30.

Pau:2019:SIN

[PSC19] Giovanni Pau, Alessandro Severino, and Antonino Canale. Special issue “New perspectives in intelligent transportation systems and mobile communications towards a smart cities context”. Future Internet, 11(11):228, October 28, 2019. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
11/228.

Przytarski:2022:QPB

[PSGM22] Dennis Przytarski, Christoph Stach, Clémentine Gritti, and Bernhard Mitschang. Query processing in blockchain systems: Current state and future challenges. Future Internet, 14(1):1, December 21, 2022. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/14/1/1.

Pizzi:2018:ETM

[PSM^{+}18] Sara Pizzi, Chiara Suraci, Leonardo Militano, Antonino Orsino, Antonella Molinaro, Antonio Iera, and Giuseppe Araniti. Enabling trustworthy multicast wireless services through D2D communications in 5G networks. Future Internet, 10(7):66, July 11, 2018. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/10/7/66.

Phengsuwan:2021:USM

[PST^{+}21] Jedsada Phengsuwan, Tejal Shah, Nipun Balan Thekkummal, Zhenyu Wen, Rui Sun, Divya Pullarkatt, Hemalatha Thirugnanam, Maneesha Vinodini Ramesh, Graham Morgan, Philip James, and Rajiv Ranjan. Use of social media data in disaster management: a survey. Future Internet, 13 (2):46, February 12, 2021. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
2/46.

Poltronieri:2023:RLV

[PSTZ23] Filippo Poltronieri, Cesare Stefanelli, Mauro Tortonesi, and Mattia Zaccarini. Reinforcement learning vs. computational intelligence: Comparing service management approaches for the cloud continuum. Future Internet, 15(11):359, October 31, 2023. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/15/11/359.

Paglialonga:2023:APS

[PTK^{+}23] Alessia Paglialonga, Rebecca Theal, Bruce Knox, Robert Kyba, David Barber, Aziz Guergachi, and Karim Keshavjee. Applying patient segmentation using primary care electronic medical records to develop a virtual peer-to-peer intervention for patients with type 2 diabetes. Future Internet, 15 (4):149, April 14, 2023. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
4/149.

Pierro:2020:ORE

[PTM20] Giuseppe Antonio Pierro, Roberto Tonelli, and Michele Marchesi. An organized repository of Ethereum smart contracts’ source codes and metrics. Future Internet, 12(11):

197, November 15, 2020. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/12/
                                        11/197.

Pilania:2022:FVS

[PTZM22] Urmila Pilania, Rohit Tanwar, Mazdak Zamani, and Azizah Abdul Manaf. Framework for video steganography using integer wavelet transform and JPEG compression. Future Internet, 14 (9):254, August 25, 2022. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/14/
                                        9/254.

Paligu:2020:BFI

[PV20] Furkan Paligu and Cihan Varol. Browser forensic investigations of WhatsApp Web utilizing IndexedDB persistent storage. Future Internet, 12(11):184, October 28, 2020. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/12/
                                        11/184.

Paligu:2022:BFI

[PV22] Furkan Paligu and Cihan Varol. Browser forensic investigations of Instagram utilizing IndexedDB persistent storage. Future Internet, 14(6):188, June 17, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/6/188.

Pergantis:2023:SOA

[PVKG23] Minas Pergantis, Iraklis Varlamis, Nikolaos Grigorios Kanellopoulos, and Andreas Giannakoulopoulos. Searching online for art and culture: User behavior analysis. Future Internet, 15(6):211– ??, June 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
6/211.

Pappalardo:2022:EOC

[PVM22] Martina Pappalardo, Antonio Virdis, and Enzo Mingozzi. Energy-optimized content refreshing of ageof-information-aware edge caches in IoT systems. Future Internet, 14(7):197, June 28, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
7/197.

Peng:2021:RDB

[PW21] Yun Peng and Jianmei Wang. Rumor detection based on attention CNN and time series of context information. Future Internet, 13(11):267, October 25, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/11/267.

Pokhrel:2022:GTR

[PW22] Shiva Raj Pokhrel and Carey Williamson. A game-theoretic rent-seeking framework for improving multipath TCP performance. Future Internet, 14 (9):257, August 29, 2022. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/14/
                                        9/257.

Peng:2018:NMP

[PXC18] Guang-Qian Peng, Guangtao Xue, and Yi-Chao Chen. Network measurement and performance analysis at server side. Future Internet, 10(7):67, July 16, 2018. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/10/7/67.

Piao:2021:DSS

[PYC21] Yangheran Piao, Kai Ye, and Xiaohui Cui. A data sharing scheme for GDPRcompliance based on consortium blockchain. Future Internet, 13(8):217, August 21, 2021. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/13/
                                        8/217.

Peng:2021:OCD

[PZC21] Yubo Peng, Bofeng Zhang, and Furong Chang. Overlapping community detection of bipartite networks based on a novel community density. Future Internet, 13(4):89, March 31, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/4/89.

Pathan:2018:ETB

[PZH^{+}18] Muhammad Salman Pathan, Nafei Zhu, Jingsha He, Zulfiqar Ali Zardari, Muhammad Qasim Memon, and Muhammad Iftikhar Hussain. An efficient trustbased scheme for secure and quality of service routing in MANETs. Future Internet, 10(2):16, February 05, 2018. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/10/2/16.

Quarati:2024:LOG

[QA24] Alfonso Quarati and Riccardo Albertoni. Linked open government data: Still a viable option for sharing and integrating public data? Future Internet, 16(3):99, March 15, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/3/99.

Quttoum:2017:ARA

[QAM17] Ahmad Nahar Quttoum, Ayoub Alsarhan, and Abidalrahman Moh’d. ARAAC: a rational allocation approach in cloud data center networks. Future Internet, 9(3):50, September 06, 2017. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/9/3/50.

Qasaimeh:2021:SDE

[QAQA21] Malik Qasaimeh, Raad S. Al-Qassas, and Mohammad Ababneh. Software design and experimental evaluation of a reduced AES for IoT applications. Future Internet, 13(11): 273, October 27, 2021. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/13/
                                        11/273.

Qiu:2022:FCD

[QDL22] Haoxuan Qiu, Yanhui Du, and Tianliang Lu. The framework of cross-domain and model adversarial attack against deepfake. Future Internet, 14(2):46, January 29, 2022. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/14/
                                        2/46.

Qu:2024:MTO

[QHX^{+}24] Qian Qu, Mohsen Hatami, Ronghua Xu, Deeraj Nagothu, Yu Chen, Xiaohua Li, Erik Blasch, Erika Ardiles-Cruz, and Genshe Chen. The microverse: a task-oriented edge-scale metaverse. Future Internet, 16(2):60, February 13, 2024. CODEN ????

ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/2/60.

Qian:2024:IPF

[QLK^{+}24] Jingshi Qian, Jiahe Li, Nobuyoshi Komuro, Won-Suk Kim, and Younghwan Yoo. Improved particle filter in machine learningbased BLE fingerprinting method to reduce indoor location estimation errors. Future Internet, 16(6):211, June 15, 2024. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
6/211.

Qiao:2021:DSR

[QLL^{+}21] Wenxin Qiao, Hao Lu, Yu Lu, Lijie Meng, and Yicen Liu. A dynamic service reconfiguration method for satellite-terrestrial integrated networks. Future Internet, 13(10):260, October 09, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/10/260.

Qian:2022:DTC

[QLR^{+}22] Cheng Qian, Xing Liu, Colin Ripley, Mian Qian, Fan Liang, and Wei Yu. Digital twin — cyber replica of physical things: Architecture, applications and future research directions. Future Internet, 14 (2):64, February 21, 2022.

CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
2/64.

Qin:2019:TCB

[QMN19] Jiwei Qin, Liangli Ma, and Jinghua Niu. THBase: a coprocessor-based scheme for big trajectory data management. Future Internet, 11(1):10, January 03, 2019. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/11/1/10.

Qasim:2024:IET

[QNJ24] Kian Raheem Qasim, Noor M. Naser, and Ahmed J. Jabur. An IoTenhanced traffic light control system with Arduino and IR sensors for optimized traffic patterns. Future Internet, 16(10), October 18, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
10/377.

Qu:2024:ISF

[QQM24] Jiantao Qu, Chunyu Qi, and He Meng. An imbalanced sequence feature extraction approach for the detection of LTE-R cells with degraded communication performance. Future Internet, 16(1):30, January 16, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.

mdpi.com/1999-5903/16/
1/30.

Qian:2024:SIS

[QQX^{+}24] Mian Qian, Cheng Qian, Guobin Xu, Pu Tian, and Wei Yu. Smart irrigation systems from cyber–physical perspective: State of art and future directions. Future Internet, 16(7):??, June 29, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/7/234.

Qadir:2017:ANU

[QSF^{+}17] Junaid Qadir, Arjuna Sathiaseelan, Umar Bin Farooq, Muhammad Usama, Muhammad Ali Imran, and Muhammad Shafique. Approximate networking for universal Internet access. Future Internet, 9(4):94, December 11, 2017. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/9/4/94.

Qiu:2024:ODE

[QSML24] Wen Qiu, Xun Shao, Hiroshi Masui, and William Liu. Optimizing drone energy use for emergency communications in disasters via deep reinforcement learning. Future Internet, 16(7):??, July 11, 2024. CODEN ???? ISSN 1999-

5903. URL https://www.

mdpi.com/1999-5903/16/
7/245.

Quan:2018:NTL

[QWR18] Li Quan, Zhiliang Wang, and Fuji Ren. A novel two-layered reinforcement learning for task offloading with tradeoff between physical machine utilization rate and delay. Future Internet, 10(7):60, July 01, 2018. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/10/7/60.

Qu:2021:EFP

[QXC^{+}21] Qian Qu, Ronghua Xu, Yu Chen, Erik Blasch, and Alexander Aved. Enable fair proof-of-work (PoW) consensus for blockchains in IoT by Miner Twins (MinT). Future Internet, 13(11):291, November 19, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/11/291.

Roffia:2018:DLD

[RAA^{+}18] Luca Roffia, Paolo Azzoni, Cristiano Aguzzi, Fabio Viola, Francesco Antoniazzi, and Tullio Salmon Cinotti. Dynamic linked data: a SPARQL event processing architecture. Future Internet, 10(4):36, April 20, 2018. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/10/4/36.

Roy:2022:MPU

[RAA^{+}22] Sanjiban Sekhar Roy, Ali Ismail Awad, Lamesgen Adugnaw Amare, Mabrie Tesfaye Erkihun, and Mohd Anas. Multimodel phishing URL detection using LSTM, bidirectional LSTM, and GRU models. Future Internet, 14(11):340, November 21, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/11/340.

Rodriguez-Abitia:2021:ADT

[RABC21] Guillermo Rodríguez-Abitia and Graciela Bribiesca-Correa. Assessing digital transformation in universities. Future Internet, 13 (2):52, February 20, 2021. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
2/52.

Rancea:2024:ECH

[RAC24] Alexandru Rancea, Ionut Anghel, and Tudor Cioara. Edge computing in healthcare: Innovations, opportunities, and challenges. Future Internet, 16(9), September 10, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
9/329.

Raheman:2022:FCA

[Rah22] Fazal Raheman. The future of cybersecurity in the age of quantum computers. Future Internet, 14(11):335, November 16, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/11/335.

Radam:2023:UMS

[RAJJ23] Neamah S. Radam, Sufyan T. Faraj Al-Janabi, and Khalid Sh. Jasim. Using metaheuristics (SA-MCSDN) optimized for multi-controller placement in software-defined networking. Future Internet, 15(1):39, January 16, 2023. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/15/1/39.

Roy:2020:ERF

[RAK^{+}20] Animesh Chandra Roy, Mohammad Shamsul Arefin, A. S. M. Kayes, Mohammad Hammoudeh, and Khandakar Ahmed. An empirical recommendation framework to support location-based services. Future Internet, 12(9):154, September 17, 2020. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/12/9/154.

Rasool:2023:QCH

[RAR^{+}23] Raihan Ur Rasool, Hafiz Farooq Ahmad, Wajid Rafique, Adnan Qayyum, Junaid Qadir, and Zahid Anwar. Quantum computing for healthcare: a review. Future Internet, 15(3):94, February 27, 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
3/94.

Ramzan:2024:BCB

[RAU^{+}24] Muhammad Sher Ramzan, Anees Asghar, Ata Ullah, Fawaz Alsolami, and Iftikhar Ahmad. A bee colony-based optimized searching mechanism in the Internet of Things. Future Internet, 16(1):35, January 22, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
1/35.

Ray:2023:OWI

[Ray23] Partha Pratim Ray. An overview of WebAssembly for IoT: Background, tools, state-of-the-art, challenges, and future directions. Future Internet, 15 (8):275, August 18, 2023. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
8/275.

Ren:2022:AVN

[RBG^{+}22] Wenbo Ren, Xinran Bian, Jiayuan Gong, Anqing Chen, Ming Li, Zhuofei Xia, and Jingnan Wang. Analysis and visualization of new energy vehicle battery data. Future Internet, 14(8):225, July 26, 2022. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
8/225.

Rodriguez-Breijo:2021:CPO

[RBSMRN21] Vanessa Rodríguez-Breijo, Núria Simelio, and Pedro Molina-Rodríguez-Navas. Council Press Offices as sources of political information: Between journalism for accountability and propaganda. Future Internet, 13(2):34, January 29, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/2/34.

Raheman:2022:WZV

[RBVV22] Fazal Raheman, Tejas Bhagat, Brecht Vermeulen, and Peter Van Daele. Will zero vulnerability computing (ZVC) ever be possible? Testing the hypothesis. Future Internet, 14(8): 238, July 30, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
8/238.

R:2018:SDM

[RC18] Jithin R and Priya Chandran. Secure and dynamic memory management architecture for virtualization technologies in IoT devices. Future Internet, 10(12):119, November 30, 2018. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/10/12/119.

Rovira:2019:RRC

[RCGSL19] Cristòfol Rovira, Lluís Codina, Frederic Guerrero-Solé, and Carlos Lopezosa. Ranking by relevance and citation counts, a comparative study: Google Scholar, Microsoft Academic, WoS and Scopus. Future Internet, 11(9):202, September 19, 2019. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
9/202.

Rovira:2021:LBG

[RCL21] Cristòfol Rovira, Lluís Codina, and Carlos Lopezosa. Language bias in the Google Scholar ranking algorithm. Future Internet, 13(2):31, January 27, 2021. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/13/2/31.

Ruan:2018:TDS

[RD18] Yibin Ruan and Jiazhu Dai. TwinNet: a double sub-network framework for detecting universal adversarial perturbations. Future Internet, 10(3):26, March 06, 2018. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/10/3/26.

Rodriguez-Diaz:2022:MLB

[RDASB22]  Nuria  Rodriguez-Diaz,

Decky Aspandi, Federico M. Sukno, and Xavier Binefa. Machine learning-based Lie detector applied to a novel annotated game dataset. Future Internet, 14(1):2, December 21, 2022. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/14/1/2.

Risse:2014:AAS

[RDD^{+}14] Thomas Risse, Elena Demidova, Stefan Dietze, Wim Peters, Nikolaos Papailiou, Katerina Doka, Yannis Stavrakas, Vassilis Plachouras, Pierre Senellart, Florent Carpentier, Amin Mantrach, Bogdan Cautis, Patrick Siehndel, and Dimitris Spiliotopoulos. The ARCOMEM architecture for socialand semantic-driven Web archiving. Future Internet, 6(4):688–716, November 04, 2014. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/6/4/688.

Recupero:2021:TEL

[Rec21] Diego Reforgiato Recupero. Technology enhanced learning using humanoid robots. Future Internet, 13(2):32, January 27, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/2/32.

Reis:2023:DCV

[Rei23] Manuel J. C. S. Reis. Developments of computer vision and image processing: Methodologies and applications. Future Internet, 15(7):233–??, July 2023. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/15/7/233.

Roumeliotis:2022:MTO

[REP22] Anargyros J. Roumeliotis, Christos N. Efrem, and Athanasios D. Panagopoulos. Minimization of nth order rate matching in satellite networks with one to many pairings. Future Internet, 14(10):286, September 30, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
10/286.

Ravi:2023:AGI

[RES23] Niranjan Ravi and Mohamed El-Sharkawy. Addressing the gaps of IoU loss in 3D object detection with IIoU. Future Internet, 15(12):399, December 11, 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.

                                        mdpi.com/1999-5903/15/
                                        12/399.

Reali:2021:TLN

[RF21] Gianluca Reali and Mauro Femminella. Two-layer network caching for different service requirements. Future Internet, 13(4):85, March 27, 2021. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/13/4/85.

Romero:2023:TRA

[RF23] Virginia M. Romero and Eduardo B. Fernandez. Towards a reference architecture for cargo ports. Future Internet, 15(4):139, April 04, 2023. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/15/
                                        4/139.

Reali:2024:AIR

[RF24] Gianluca Reali and Mauro Femminella. Artificial intelligence to reshape the healthcare ecosystem. Future Internet, 16(9), September 20, 2024. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/16/
                                        9/343.

Rudwan:2023:HFS

[RFD23] Mohammed Suleiman Mohammed Rudwan and Jean Vincent Fonou-Dombeu. Hybridizing fuzzy string matching and machine learning for improved ontology alignment. Future Internet, 15 (7):229–??, July 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
7/229.

Rodrigues:2022:QPP

[RFS22] Pedro Rodrigues, Filipe Freitas, and José Simão. QuickFaaS: Providing portability and interoperability between FaaS platforms. Future Internet, 14(12):360, November 30, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/12/360.

Rodriguez-Ferrandiz:2021:MDS

[RFSOHMSB21] Raúl Rodríguez-Ferrándiz, Cande Sánchez-Olmos, Tatiana Hidalgo-Marí, and Estela Saquete-Boro. Memetics of deception: Spreading local meme hoaxes during COVID-19 1st year. Future Internet, 13(6):152, June 10, 2021. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
6/152.

Ranaldi:2022:DCA

[RFZ22] Leonardo Ranaldi, Francesca Fallucchi, and Fabio Massimo Zanzotto. Dis-cover AI minds to preserve human knowledge. Future Internet, 14(1):10, December 24, 2022. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/14/1/10.

Rajawat:2023:QML

[RGB^{+}23] Anand Singh Rajawat, S. B. Goyal, Pradeep Bedi, Tony Jan, Md Whaiduzzaman, and Mukesh Prasad. Quantum machine learning for security assessment in the Internet of Medical Things (IoMT). Future Internet, 15(8):271, August 15, 2023. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/15/
                                        8/271.

Rahman:2024:CRM

[RGB^{+}24] Md Motiur Rahman, Deepti Gupta, Smriti Bhatt, Shiva Shokouhmand, and Miad Faezipour. A comprehensive review of machine learning approaches for anomaly detection in smart homes: Experimental analysis and future directions. Future Internet, 16(4):139, April 19, 2024. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/16/4/139.

Rojas:2021:DBR

[RGCM21] Oscar Rojas, Veronica Gil-Costa, and Mauricio Marin. A DFT-based running time prediction algorithm for Web queries. Future Internet, 13(8):204, August 04, 2021. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
8/204.

RuizVinals:2024:MFA

[RGD24] Carmen Ruiz Viñals, Marta Gil Ibáñez, and José Luis Del Olmo Arriaga. Metaverse and fashion: an analysis of consumer online interest. Future Internet, 16(6): 199, June 4, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
6/199.

Reis:2022:BMI

[RGdSK22] Jacqueline Zonichenn Reis, Rodrigo Franco Gonçalves, Marcia Terra da Silva, and Nikolai Kazantsev. Business models for the Internet of Services: State of the art and research agenda. Future Internet, 14(3):74, February 25, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/3/74.

Rajpal:2021:FBH

[RGM^{+}21] Danveer Rajpal, Akhil Ranjan Garg, Om Prakash Mahela, Hassan Haes Alhelou, and Pierluigi Siano.

A fusion-based hybridfeature approach for recognition of unconstrained offline handwritten Hindi characters. Future Internet, 13(9):239, September 18, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/9/239.

Rosado:2012:SAM

[RGMFM12] David G. Rosado, Rafael Gómez, Daniel Mellado, and Eduardo Fernández-Medina. Security analysis in the migration to cloud environments. Future Internet, 4(2):469–487, May 08, 2012. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/4/2/469.

Rubart:2022:AIC

[RGP22] Jessica Rubart, Valentin Grimm, and Jonas Potthast. Augmenting industrial control rooms with multimodal collaborative interaction techniques. Future Internet, 14(8):224, July 26, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
8/224.

Romero:2024:BDG

[RH24] Abraham Romero and Roberto Hernandez. Blockchaindriven generalization of policy management for multiproduct insurance companies. Future Internet, 16(10), September 30, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/10/356.

Riana:2021:IFH

[RHHN21] Dwiza Riana, Achmad Nizar Hidayanto, Sri Hadianti, and Darmawan Napitupulu. Integrative factors of e-health laboratory adoption: a case of Indonesia. Future Internet, 13 (2):26, January 24, 2021. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
2/26.

Roychowdhury:2017:ABS

[RHV17] Sohini Roychowdhury, Paul Hage, and Joseph Vasquez. Azure-based smart monitoring system for anemialike pallor. Future Internet, 9(3):39, July 26, 2017. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/9/
3/39.

Rauniyar:2017:DDR

[RIS^{+}17] Ashish Rauniyar, Mohammad Irfan, Oka Danil Saputra, Jin Woo Kim, Ah Ra Lee, Jae Min Jang, and Soo Young Shin. Design and development of a real-time monitoring system for multiple lead-acid batteries based on Internet of Things. Future Internet, 9(3):28, June 29, 2017. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/9/
3/28.

Roohitavaf:2019:AAF

[RK19] Mohammad Roohitavaf and Sandeep Kulkarni. Automatic addition of fault-tolerance in presence of unchangeable environment actions. Future Internet, 11(7):144, July 04, 2019. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/11/7/144.

Rejeb:2022:CPP

[RKM^{+}22] Abderahman Rejeb, John G. Keogh, Wayne Martindale, Damion Dooley, Edward Smart, Steven Simske, Samuel Fosso Wamba, John G. Breslin, Kosala Yapa Bandara, Subhasis Thakur, Kelly Liu, Bridgette Crowley, Sowmya Desaraju, Angela Ospina, and Horia Bradau. Charting past, present, and future research in the Semantic Web and interoperability. Future Internet, 14(6): 161, May 25, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
6/161.

Romanov:2021:AIR

[RKS^{+}21] Aleksandr Romanov, Anna Kurtukova, Alexander Shelupanov, Anastasia Fedotova, and Valery Goncharov. Authorship identification of a Russianlanguage text using support vector machine and deep neural networks. Future Internet, 13(1):3, December 25, 2021. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/13/1/3.

Rejeb:2019:LIT

[RKT19] Abderahman Rejeb, John G.

Keogh, and Horst Treiblmaier. Leveraging the Internet of Things and blockchain technology in supply chain management. Future Internet, 11(7):161, July 20, 2019. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/11/7/161.

Rustamov:2020:DDA

[RKY20] Fayozbek Rustamov, Juhwan Kim, and JooBeom Yun. DeepDiver: Diving into abysmal depth of the binary for hunting deeply hidden software vulnerabilities. Future Internet, 12(4):74, April 18, 2020. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/12/
4/74.

Rasheed:2024:BBI

[RL24] Sana Rasheed and Soulla Louca. Blockchain-based implementation of national census as a supplementary instrument for enhanced transparency, accountability, privacy, and security. Future Internet, 16(1):24, January 11, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/1/24.

Rangelov:2023:TIM

[RLB^{+}23] Denis Rangelov, Philipp Lämmel, Lisa Brunzel, Stephan Borgert, Paul Darius, Nikolay Tcholtchev, and Michell Boerger. Towards an integrated methodology and toolchain for machine learning-based intrusion detection in urban IoT networks and platforms. Future Internet, 15(3):98, February 28, 2023. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/15/3/98.

Rajaei:2023:SPD

[RM23] Mohammad Javad Rajaei and Qusay H. Mahmoud. A survey on pump and dump detection in the cryptocurrency market using machine learning. Future Internet, 15(8):267, August 11, 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
8/267.

Roche:2012:WBC

[RMB^{+}12] Stéphane Roche, Boris Mericskay, Wided Batita, Matthieu Bach, and Mathieu Rondeau. WikiGIS basic concepts: Web 2.0 for geospatial collaboration. Future Internet, 4(1): 265–284, March 13, 2012. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/4/
1/265.

Ramirez-Montoya:2021:CTP

[RMLAZOPC21] María Soledad Ramírez-Montoya, María Isabel Loaiza-Aguirre, Alexandra Zúñiga-Ojeda, and May Portuguez-Castro. Characterization of the teaching profile within the framework of education 4.0. Future Internet, 13(4): 91, April 01, 2021. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
4/91.

Rahayu:2020:RTU

[RNFS20] Flourensia Sapty Rahayu, Lukito Edi Nugroho, Ridi Ferdiana, and Djoko Budiyanto Setyohadi. Research trend on the use of IT in digital addiction: an investigation using a systematic literature review. Future Internet, 12(10):174, October 18, 2020. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/12/
                                        10/174.

Ridolfi:2023:IEF

[RNST23] Lorenzo Ridolfi, David Naseh, Swapnil Sadashiv Shinde, and Daniele Tarchi. Implementation and evaluation of a federated learning framework on Raspberry PI platforms for IoT 6G applications. Future Internet, 15(11):358, October 31, 2023. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/15/11/358.

Roka:2023:ODC

[Rók23] Rastislav Róka. Optimization of the decision criterion for increasing the bandwidth utilization by means of the novel effective DBA algorithm in NG-PON2 networks. Future Internet, 15(7):242– ??, July 2023. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/15/
                                        7/242.

Rotondo:2012:UCP

[Rot12] Francesco Rotondo. The U-City paradigm: Opportunities and risks for edemocracy in collaborative planning. Future Internet, 4(2):563–574, June 05, 2012. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/4/2/563.

Roy:2014:ODO

[Roy14] Jeffrey Roy. Open data and open governance in Canada: a critical examination of new opportunities and old tensions. Future Internet, 6(3):414– 432, June 27, 2014. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/6/
3/414.

Rizvi:2017:DCB

[RPB^{+}17] Syed Tahir Hussain Rizvi, Denis Patti, Tomas Björklund, Gianpiero Cabodi, and Gianluca Francini. Deep classifiers-based license plate detection, localization and recognition on GPU-powered mobile platform. Future Internet, 9 (4):66, October 21, 2017. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/9/
4/66.

Rifa-Pous:2011:CEC

[RPHJ11] Helena Rifà-Pous and Jordi Herrera-Joancomartí. Computational and energy costs of cryptographic algorithms on handheld devices. Future Internet, 3(1):31–48, February 14, 2011. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/3/1/31.

Rajendran:2024:RTE

[RR24a] Venushini Rajendran and R Kanesaraj Ramasamy. Real-time evaluation of the improved eagle strategy model in the Internet of Things. Future Internet, 16(11):409, November 6, 2024. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/16/11/409.

Rosario:2024:ITD

[RR24b] Albérico Travassos Rosário and Ricardo Raimundo. Internet of Things and distributed computing systems in business models. Future Internet, 16(10), October 21, 2024. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/16/
                                        10/384.

Romero-Rodriguez:2020:CII

[RRAGMMGG20] José-María Romero-Rodríguez, Santiago Alonso-García, José-Antonio Marín-Marín, and Gerardo Gómez-García. Considerations on the implications of the Internet of Things in Spanish universities: The usefulness perceived by professors. Future Internet, 12(8):123, July 24, 2020. CODEN ???? ISSN 1999-

                                        5903. URL https://www.

mdpi.com/1999-5903/12/
8/123.

Rodrigues:2022:QBM

[RRdBSS22] Laécio Rodrigues, Joel J.

P. C. Rodrigues, Antonio de Barros Serra, and Francisco Airton Silva. A queueing-based model performance evaluation for Internet of people supported by fog computing. Future Internet, 14(1):23, January 08, 2022. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/14/1/23.

Ramakrishnan:2020:STC

[RRLT20] Abinaya Megan Ramakrishnan, Aparna Nicole Ramakrishnan, Sarah Lagan, and John Torous. From symptom tracking to contact tracing: A framework to explore and assess COVID-19 apps. Future Internet, 12(9):153, September 08, 2020. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/12/
9/153.

Rajendran:2022:IES

[RRMI22] Venushini Rajendran, R Kanesaraj Ramasamy, and Wan-Noorshahida Mohd-Isa. Improved eagle strategy algorithm for dynamic Web service composition in the IoT: a conceptual approach. Future Internet, 14(2):56, February 15, 2022. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/14/2/56.

Reddy:2017:ETE

[RRP17] Nalavala Ramanjaneya Reddy, Pakanati Chenna Reddy, and Mokkala Padmavathamma. Efficient traffic engineering strategies for improving the performance of TCP friendly rate control protocol. Future Internet, 9(4):74, November 01, 2017. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/9/
                                        4/74.

Restrepo:2020:CTT

[RRS20] Sandra Restrepo, Dubis Rincón, and Edwin Sepulveda. Cognitive training for the treatment of addictions mediated by information and communication technologies (ICT). Future Internet, 12(2):38, February 14, 2020. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/12/2/38.

Rinaldi:2020:KDM

[RRT20] Antonio Maria Rinaldi, Cristiano Russo, and Cristian Tommasino. A knowledge-driven multimedia retrieval system based on semantics and deep features. Future Internet, 12(11):183, October 28, 2020. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/12/11/183.

Ren:2023:KDB

[RRT^{+}23] Yingying Ren, Ryan D.

Restivo, Wenkai Tan, Jian Wang, Yongxin Liu, Bin Jiang, Huihui Wang, and Houbing Song. Knowledge distillation-based GPS spoofing detection for small UAV. Future Internet, 15(12):389, November 30, 2023. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/15/12/389.

Rametta:2017:DSN

[RS17] Corrado Rametta and Giovanni Schembra. Designing a softwarized network deployed on a fleet of drones for rural zone monitoring. Future Internet, 9 (1):8, March 20, 2017. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/9/
1/8.

Ruffo:2022:FES

[RS22] Giancarlo Ruffo and Alfonso Semeraro. Fake-NewsLab: Experimental study on biases and pitfalls preventing us from distinguishing true from false news. Future Internet, 14(10):283, September 29, 2022. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/14/10/283.

Ricci:2023:DMM

[RSF23] Marina Ricci, Alessandra Scarcelli, and Michele Fiorentino. Designing for the metaverse: a multidisciplinary laboratory in the industrial design program. Future Internet, 15(2):69, February 10, 2023. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/15/
                                        2/69.

Rossler:2021:RVF

[RSG21] Jannik Rößler, Jiachen Sun, and Peter Gloor. Reducing videoconferencing fatigue through facial emotion recognition. Future Internet, 13(5):126, May 12, 2021. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/13/5/126.

Ruiu:2024:DIE

[RSG24] Pietro Ruiu, Salvatore Saiu, and Enrico Grosso. Digital identity in the EU: Promoting eIDAS solutions based on biometrics. Future Internet, 16 (7):??, June 28, 2024. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/16/
                                        7/228.

Ruta:2024:ICT

[RSL^{+}24] Michele Ruta, Floriano Scioscia, Giuseppe Loseto, Agnese Pinto, Corrado Fasciano, Giovanna Capurso, and Eugenio Di Sciascio. Internet of Conscious Things: Ontologybased social capabilities for smart objects. Future Internet, 16(9), September 8, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/9/327.

Ramadan:2021:UML

[RSMC21] Lara Ramadan, Isam Shahrour, Hussein Mroueh, and Fadi Hage Chehade. Use of machine learning methods for indoor temperature forecasting. Future Internet, 13(10):242, September 23, 2021. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
10/242.

Rodrigues:2024:ICC

[RSV^{+}24] Gabriel Arquelau Pimenta Rodrigues, André Luiz Marques Serrano, Guilherme Fay Vergara, Robson de Oliveira Albuquerque, and Georges Daniel Amvame Nze. Impact, compliance, and countermeasures in relation to data breaches in publicly traded U.S. companies. Future Internet, 16 (6):201, June 5, 2024. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/16/
                                        6/201.

Rexha:2018:ITF

[RSX18] Blerim Rexha, Gresa Shala, and Valon Xhafa. Increasing trustworthiness of face authentication in mobile devices by modeling gesture behavior and location using neural networks. Future Internet, 10 (2):17, February 05, 2018. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/10/
                                        2/17.

Roumeliotis:2023:COA

[RT23] Konstantinos I. Roumeliotis and Nikolaos D. Tselikas. ChatGPT and open-AI models: a preliminary review. Future Internet, 15(6):192–??, June 2023. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/15/
                                        6/192.

Rozanec:2022:EAI

[RTN^{+}22] Joze Rozanec, Elena Trajkova, Inna Novalija, Patrik Zajec, Klemen Kenda, Blaz Fortuna, and Dunja Mladenić. Enriching artificial intelligence explanations with knowledge fragments. Future Internet, 14(5):134, April 29, 2022.

CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
5/134.

Razak:2022:MUA

[RYAA22] Siti Fatimah Abdul Razak, Sumendra Yogarayan, Mohd Fikri Azli Abdullah, and Afizan Azman. Modeling user acceptance of in-vehicle applications for safer road environment. Future Internet, 14(5): 148, May 11, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
5/148.

Ramstetter:2010:ASN

[RYY10] Jerry Rick Ramstetter, Yaling Yang, and Danfeng Yao. Applications and security of next-generation, user-centric wireless systems. Future Internet, 2 (3):190–211, July 28, 2010. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/2/
3/190.

Rao:2019:SSN

[RZCL19] Xuli Rao, Jiaxu Zhao, Zhide Chen, and Feng Lin. Substitute seed nodes mining algorithms for influence maximization in multi-social networks. Future Internet, 11(5):112, May 10, 2019. CO-DEN ???? ISSN 1999-

5903. URL https://www.

                                        mdpi.com/1999-5903/11/
                                        5/112.

Ren:2018:QSM

[RZQS18] Chuanxiang Ren, Wenbo Zhang, Lingqiao Qin, and Bo Sun. Queue spillover management in a connected vehicle environment. Future Internet, 10 (8):79, August 10, 2018. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/10/
                                        8/79.

Shalaginov:2021:SRC

[SA21] Andrii Shalaginov and Muhammad Ajmal Azad. Securing resource-constrained IoT nodes: Towards intelligent microcontrollerbased attack detection in distributed smart applications. Future Internet, 13(11):272, October 27, 2021. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/13/11/272.

Soares:2024:TNB

[SA24] Rendrikson Soares and André Araújo. Towards a new business model for streaming platforms using blockchain technology. Future Internet, 16 (6):207, June 13, 2024. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/16/
                                        6/207.

Sacco:2019:SIS

[Sac19] Giovanni Maria Sacco.

SAES: An introduction to self-adapting exploratory structures. Future Internet, 11(3):54, February 26, 2019. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/11/3/54.

Sabatucci:2023:EDP

[SACG23] Luca Sabatucci, Agnese Augello, Giuseppe Caggianese, and Luigi Gallo. Envisioning digital practices in the metaverse: a methodological perspective. Future Internet, 15(12):394, December 06, 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
12/394.

Santonicola:2024:ACA

[SACP24] Emanuele Santonicola, Ennio Andrea Adinolfi, Simone Coppola, and Francesco Pascale. Automotive cybersecurity application based on CAR-DIAN. Future Internet, 16 (1):10, December 28, 2024. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
1/10.

Scazzariello:2021:MSA

[SADP21] Mariano Scazzariello, Lorenzo Ariemma, Giuseppe Di Battista, and Maurizio Patrignani. Megalos: a scalable architecture for the virtualization of large network scenarios. Future Internet, 13(9):227, August 30, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/9/227.

Shepherd:2014:RRS

[SAF14] Lynsay A. Shepherd, Jacqueline Archibald, and Robert Ian Ferguson. Reducing risky security behaviours: Utilising affective feedback to educate users. Future Internet, 6(4):760–772, November 27, 2014. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/6/4/760.

Saadatfar:2024:NDG

[SAJ24] Hamid Saadatfar, Hamid Gholampour Ahangar, and Javad Hassannataj Joloudari. A new dynamic gamebased pricing model for cloud environment. Future Internet, 16(2):49, January 31, 2024. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/16/2/49.

Scata:2021:CPN

[SAL21] Marialisa Scatá, Barbara Attanasio, and Aurelio La Corte. Cognitive profiling of nodes in 6G through multiplex social network and evolutionary collective dynamics. Future Internet, 13(5):135, May 20, 2021. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
5/135.

Suomalainen:2018:SAS

[SAM^{+}18] Jani Suomalainen, Kimmo Ahola, Mikko Majanen, Olli Mämmelä, and Pekka Ruuska. Security awareness in software-defined multi-domain 5G networks. Future Internet, 10(3):27, March 08, 2018. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/10/
3/27.

Shahid:2024:ONP

[SAR24] Kamal Shahid, Saleem Naseer Ahmad, and Syed Tahir Hussain Rizvi. Optimizing network performance: a comparative analysis of EIGRP, OSPF, and BGP in IPv6-based load-sharing and link-failover systems. Future Internet, 16(9), September 20, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
9/339.

Sultan:2018:BDP

[SAZ18] Kashif Sultan, Hazrat Ali, and Zhongshan Zhang. Big data perspective and challenges in next generation networks. Future Internet, 10(7):56, June 21, 2018. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/10/
                                        7/56.

S:2022:QSL

[SB22] Shrisha H. S. and Uma Boregowda. Quality-of-service-linked privileged content-caching mechanism for named data networks. Future Internet, 14(5):157, May 20, 2022. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/14/
                                        5/157.

Szabo:2023:NAW

[SB23] Zoltán Szabó and Vilmos Bilicki. A new approach to Web application security: Utilizing GPT language models for source code inspection. Future Internet, 15(10):326, September 28, 2023. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/15/10/326.

Safarov:2023:ELB

[SBN^{+}23] Furkat Safarov, Mainak Basak, Rashid Nasimov, Akmalbek Abdusalomov, and Young Im Cho. Explainable lightweight block attention module framework for network-based IoT attack detection. Future Internet, 15(9):297, September 01, 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
9/297.

Saleem:2024:EET

[SBRZ24] Gulshan Saleem, Usama Ijaz Bajwa, Rana Hammad Raza, and Fan Zhang. Edge-enhanced Tempo-FuseNet: a two-stream framework for intelligent multiclass video anomaly recognition in 5G and IoT environments. Future Internet, 16(3):83, February 29, 2024. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/16/3/83.

Safkhani:2018:SRO

[SBS18] Masoumeh Safkhani, Nasour Bagheri, and Mahyar Shariat. On the security of rotation operation based ultra-lightweight authentication protocols for RFID systems. Future Internet, 10(9):82, August 21, 2018. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/10/9/82.

Sarasa-Cabezuelo:2020:MCI

[SC20] Antonio Sarasa-Cabezuelo.

A model for creating interactive eBooks for eLearning. Future Internet, 12 (12):223, December 07, 2020. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/12/12/223.

Shen:2023:ECB

[SC23] Tsu-Chuan Shen and Edward T.-H. Chu. Edgecomputing-based peoplecounting system for elevators using MobileNet– single-stage object detection. Future Internet, 15(10):337, October 14, 2023. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/15/10/337.

Scassa:2014:POG

[Sca14] Teresa Scassa. Privacy and open government. Future Internet, 6(2):397– 413, June 18, 2014. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/6/
                                        2/397.

Sengupta:2019:HAT

[SCA^{+}19] Saumendra Sengupta, Chen-Fu Chiang, Bruno Andriamanalimanana, Jorge Novillo, and Ali Tekeoglu. A hybrid adaptive transaction injection protocol and its optimization for verification-based decentralized system. Future Internet, 11(8):167, July 27, 2019. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/11/8/167.

Saia:2021:WIM

[SCB21] Roberto Saia, Salvatore Carta, and Olaf Bergmann. Wireless Internet, multimedia, and artificial intelligence: New applications and infrastructures. Future Internet, 13(9):240, September 21, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/9/240.

Serodio:2023:ESI

[SCC^{+}23] Carlos Serôdio, José Cunha, Guillermo Candela, Santiago Rodriguez, Xosé Ramón Sousa, and Frederico Branco. The 6G ecosystem as support for IoE and private networks: Vision, requirements, and challenges. Future Internet, 15(11):348, October 25, 2023. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/15/11/348.

Subirats:2012:KRP

[SCM12] Laia Subirats, Luigi Ceccaroni, and Felip Miralles. Knowledge representation for prognosis of health status in rehabilitation. Future Internet, 4(3):762– 775, August 20, 2012. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/4/
3/762.

Salariseddigh:2024:DIF

[SDDB24] Mohammad Javad Salariseddigh, Ons Dabbabi, Christian Deppe, and Holger Boche. Deterministic K-identification for future communication networks: The binary symmetric channel results. Future Internet, 16(3):78, February 26, 2024. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/16/
                                        3/78.

Sun:2019:TMT

[SDL^{+}19] Yue Sun, Songmin Dai, Jide Li, Yin Zhang, and Xiaoqiang Li. Toothmarked tongue recognition using gradient-weighted class activation maps. Future Internet, 11(2):45, February 15, 2019. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/11/
                                        2/45.

Simoes:2021:DAS

[SDM21] Rhodney Simões, Kelvin Dias, and Ricardo Martins. Dynamic allocation of SDN controllers in NFVbased MEC for the Internet of Vehicles. Future Internet, 13(11):270, October 26, 2021. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/13/
                                        11/270.

Spyridou:2022:MVN

[SDM22] Paschalia (Lia) Spyridou, Constantinos Djouvas, and Dimitra Milioni. Modeling and validating a news recommender algorithm in a mainstream medium-sized news organization: an experimental approach. Future Internet, 14(10):284, September 29, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
10/284.

Serrano:2022:UPC

[SDS^{+}22] Ronaldo Serrano, Ckristian Duran, Marco Sarmiento, Tuan-Kiet Dang, Trong-Thuc Hoang, and Cong-Kha Pham. A unified PUF and crypto core exploiting the metastability in latches. Future Internet, 14(10):298, October 17, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/10/298.

Schweighofer:2015:ACW

[SE15] Patrick Schweighofer and Martin Ebner. Aspects to be considered when implementing technologyenhanced learning approaches: a literature review. Future Internet, 7(1):26–49, February 03, 2015. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/7/1/26.

Salah:2023:ENS

[SE23] Zaher Salah and Esraa Abu Elsoud. Enhancing network security: A machine learning-based approach for detecting and mitigating Krack and Kr00k attacks in IEEE 802.11. Future Internet, 15 (8):269, August 14, 2023. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/15/
                                        8/269.

Svorobej:2019:SFE

[SEB^{+}19] Sergej Svorobej, Patricia Takako Endo, Malika Bendechache, Christos Filelis-Papadopoulos, Konstantinos M. Giannoutakis, George A. Gravvanis, Dimitrios Tzovaras, James Byrne, and Theo Lynn. Simulating fog and edge computing scenarios: An overview and research challenges. Future Internet, 11(3):55, February 26, 2019. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/11/3/55.

Sempere:2011:ADV

[Sem11] Andrew Sempere. Architecture and design for virtual conferences: a case study. Future Internet, 3 (3):175–184, July 06, 2011. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/3/
                                        3/175.

Saleh:2024:EDL

[SESO24] Sherine Nagy Saleh, Mazen Nabil Elagamy, Yasmine N. M. Saleh, and Radwa Ahmed Osman. An explainable deep learning-enhanced IoMT model for effective monitoring and reduction of maternal mortality risks. Future Internet, 16(11):411, November 8, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/11/411.

Sultana:2022:IRS

[SF22] Ajmery Sultana and Xavier Fernando. Intelligent reflecting surface-aided device-to-device communication: a deep reinforcement learning approach. Future Internet, 14(9):256, August 29, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
9/256.

Scheel:2020:VIA

[SFD20] Christian Scheel, Francesca Fallucchi, and Ernesto William De Luca. Visualization, interaction and analysis of heterogeneous textbook resources. Future Internet, 12(10):176, October 21, 2020. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/12/10/176.

Shehni:2018:NLW

[SFEK18] Rezvan Almas Shehni, Karim Faez, Farshad Eshghi, and Manoochehr Kelarestaghi. A new lightweight watchdog-based algorithm for detecting Sybil nodes in mobile WSNs. Future Internet, 10 (1):1, December 21, 2018. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/10/
1/1.

Sturley:2024:VVC

[SFSN^{+}24] Hamish Sturley, Augustin Fournier, Andoni Salcedo-Navarro, Miguel Garcia-Pineda, and Jaume Segura-Garcia. Virtualization vs. containerization, a comparative approach for application deployment in the computing continuum focused on the edge. Future Internet, 16(11):427, November 19, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
11/427.

Seiger:2023:IMD

[SFW23] Ronny Seiger, Marco Franceschetti, and Barbara Weber. An interactive method for detection of process activity executions from IoT data. Future Internet, 15(2):77, February 16, 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
2/77.

Sarma:2012:STP

[SG12] Amardeo Sarma and Joao Girao. Supporting trust and privacy with an identity-enabled architecture. Future Internet, 4 (4):1016–1025, November 19, 2012. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/4/4/1016.

Sgantzos:2019:AII

[SG19] Konstantinos Sgantzos and Ian Grigg. Artificial intelligence implementations on the blockchain. Use cases and future applications. Future Internet, 11(8):170, August 02, 2019. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/11/8/170.

Sun:2021:MNP

[SG21a] Jiachen Sun and Peter Gloor. E-mail network patterns and body language predict risk-taking attitude. Future Internet, 13(1):17, January 14, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/1/17.

Sun:2021:APP

[SG21b] Jiachen Sun and Peter A.

Gloor. Assessing the predictive power of online social media to analyze COVID-19 outbreaks in the 50 U.S. states. Future Internet, 13(7):184, July 20, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/7/184.

Stach:2023:SIS

[SG23] Christoph Stach and Clémentine Gritti. Special issue on security and privacy in blockchains and the IoT volume II. Future Internet, 15(8):272, August 16, 2023. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/15/8/272.

Spathoulas:2019:CBB

[SGDT19] Georgios Spathoulas, Nikolaos Giachoudis, Georgios-Paraskevas Damiris, and Georgios Theodoridis. Collaborative blockchain-based detection of distributed denial of service attacks based on Internet of Things botnets. Future Internet, 11(11):226, October 25, 2019. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/11/11/226.

Sagu:2022:HDL

[SGG^{+}22] Amit Sagu, Nasib Singh Gill, Preeti Gulia, Jyotir Moy Chatterjee, and Ishaani Priyadarshini. A hybrid deep learning model with self-improved optimization algorithm for detection of security attacks in IoT environment. Future Internet, 14(10):301, October 19, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
10/301.

Shi:2020:ERS

[SGLZ20] Yanjun Shi, Yijia Guo, Lingling Lv, and Keshuai Zhang. An efficient resource scheduling strategy for V2X microservice deployment in edge servers. Future Internet, 12(10): 172, October 15, 2020. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/12/
10/172.

Seyghaly:2024:SSS

[SGMBK24] Rasool Seyghaly, Jordi Garcia, Xavi Masip-Bruin, and Jovana Kuljanin. SBNNR: Small-size batoptimized KNN regression. Future Internet, 16 (11):422, November 14, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/11/422.

Santamaria-Granados:2021:TRS

[SGMMRG21] Luz Santamaria-Granados, Juan Francisco Mendoza-Moreno, and Gustavo Ramirez-Gonzalez. Tourist recommender systems based on emotion recognition — a scientometric review. Future Internet, 13(1):2, December 24, 2021. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/13/1/2.

Scholz:2013:CSF

[SGR^{+}13] Markus Scholz, Dawud Gordon, Leonardo Ramirez, Stephan Sigg, Tobias Dyrks, and Michael Beigl. A concept for support of firefighter frontline communication. Future Internet, 5(2):113–127, April 16, 2013. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/5/2/113.

Shi:2022:HRL

[SGW22] Jialin Shi, Chenyi Guo, and Ji Wu. A hybrid robust-learning architecture for medical image segmentation with noisy labels. Future Internet, 14 (2):41, January 26, 2022. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/14/
                                        2/41.

Samarnggoon:2023:DVW

[SGWK23] Keattikorn Samarnggoon, Supara Grudpan, Noppon Wongta, and Konlawat Klaynak. Developing a virtual world for an openhouse event: a metaverse approach. Future Internet, 15(4):124, March 27, 2023. CODEN ???? ISSN 1999-

                                        5903. URL https://www.

mdpi.com/1999-5903/15/
4/124.

Song:2020:DPD

[SH20] Ha Yoon Song and Hyochang Han. A design of a parcel delivery system for point to point delivery with IoT technology. Future Internet, 12(4):70, April 17, 2020. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/12/4/70.

Song:2021:SAC

[SH21] Guizhe Song and Degen Huang. A sentiment-aware contextual model for realtime disaster prediction using Twitter data. Future Internet, 13(7):163, June 25, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/7/163.

Suttle:2021:UCM

[SHA^{+}21] Ryan Suttle, Scott Hogan, Rachel Aumaugher, Matthew Spradling, Zak Merrigan, and Jeremy Straub. University community members’ perceptions of labels for online media. Future Internet, 13(11):281, October 31, 2021. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
11/281.

Saffre:2023:FBS

[SHA23] Fabrice Saffre, Hanno Hildmann, and Antti Anttonen. Force-based self-organizing MANET/ FANET with a UAV swarm. Future Internet, 15(9):315, September 19, 2023. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/15/9/315.

Simon:2023:CAH

[SHB23] Marek Simon, Ladislav Huraj, and Nicolas Búcik. A comparative analysis of high availability for Linux container infrastructures. Future Internet, 15(8): 253, July 28, 2023. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/15/
                                        8/253.

Sufiev:2019:DSC

[SHBS19] Hadar Sufiev, Yoram Haddad, Leonid Barenboim, and José Soler. Dynamic SDN controller load balancing. Future Internet, 11(3):75, March 21, 2019. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/11/
                                        3/75.

Seet:2018:RAC

[SHC18] Boon-Chong Seet, Syed Faraz Hasan, and Peter Han-Joo Chong. Recent advances on cellular D2D communications. Future Internet, 10(1):10, January 17, 2018. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/10/1/10.

Sarma:2022:PAD

[SHC22] Subhra Sankha Sarma, Ranjay Hazra, and Peter Han Joo Chong. Performance analysis of DF relay-assisted D2D communication in a 5G mmWave network. Future Internet, 14(4):101– ??, March 24, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
4/101.

Sim:2024:CWT

[SHC24] Kyungmin Sim, Honyeong Heo, and Haehyun Cho. Combating Web tracking: Analyzing Web tracking technologies for user privacy. Future Internet, 16 (10), October 5, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
10/363.

Satar:2021:TVC

[SHHM21] Siti Dhalila Mohd Satar, Masnida Hussin, Zurina Mohd Hanapi, and Mohamad Afendee Mohamed. Towards virtuous cloud data storage using access policy hiding in ciphertext policy attributebased encryption. Future Internet, 13(11):279, October 30, 2021. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/13/
                                        11/279.

Sun:2024:PCO

[SHLH24] Enbo Sun, Jiaxuan Han, Yiquan Li, and Cheng Huang. A packet contentoriented remote code execution attack payload detection model. Future Internet, 16(7):??, July 2, 2024. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/16/7/235.

Song:2019:JUD

[SHN^{+}19] Xin Song, Xiuwei Han, Yue Ni, Li Dong, and Lei Qin. Joint uplink and downlink resource allocation for D2D communications system. Future Internet, 11(1):12, January 06, 2019. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/11/1/12.

So-In:2010:DRR

[SIJA10] Chakchai So-In, Raj Jain, and Abdel-Karim Al Tamimi. Deficit round Robin with fragmentation scheduling to achieve generalized weighted fairness for resource allocation in IEEE 802.16e mobile WiMAX networks. Future Internet, 2(4):446–468, October 12, 2010. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/2/4/446.

Stojmenova:2012:AUC

[SIZD12] Emilija Stojmenova, Bojan Imperl, Tomaz Zohar, and Dejan Dinevski. Adapted user-centered design: a strategy for the higher user acceptance of innovative e-health services. Future Internet, 4 (3):776–787, August 27, 2012. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/4/3/776.

Stenliden:2012:HCW

[SJ12] Linnea Stenliden and Mikael Jern. How can we study learning with geovisual analytics applied to statistics? Future Internet, 4(1):22–41, December 30, 2012. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/4/1/22.

Su:2020:HGR

[SJ20] Zichun Su and Jialin Jiang. Hierarchical gated recurrent unit with semantic attention for event prediction. Future Internet, 12(2):39, February 14, 2020. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/12/2/39.

Safavi:2021:RPR

[SJ21] Sadaf Safavi and Mehrdad Jalali. RecPOID: POI recommendation with friendship aware and deep CNN. Future Internet, 13(3):79, March 22, 2021. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/13/3/79.

Salam:2023:CBD

[SJA^{+}23] Abdu Salam, Qaisar Javaid, Masood Ahmad, Ishtiaq Wahid, and Muhammad Yeasir Arafat. Clusterbased data aggregation in flying sensor networks enabled Internet of Things. Future Internet, 15(8):279, August 20, 2023. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/15/
                                        8/279.

Sahajwani:2018:LLR

[SJG18] Manish Sahajwani, Alok Jain, and Radheyshyam Gamad. Log likelihood ratio based relay selection scheme for amplify and forward relaying with three state Markov channel. Future Internet, 10(9): 87, September 06, 2018. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/10/
                                        9/87.

Salahdine:2019:SEA

[SK19] Fatima Salahdine and Naima Kaabouch. Social engineering attacks: a survey. Future Internet, 11 (4):89, April 02, 2019. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
4/89.

Spiekermann:2023:CNF

[SK23] Daniel Spiekermann and Jörg Keller. Challenges of network forensic investigation in fog and edge computing. Future Internet, 15(10):342, October 18, 2023. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/15/10/342.

Siachos:2024:EAI

[SK24] Ilias Siachos and Nikos Karacapilidis. Explainable artificial intelligence methods to enhance transparency and trust in digital deliberation settings. Future Internet, 16(7):??, July 6, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/7/241.

Sone:2023:OST

[SKA^{+}23] Takuro Sone, Shin Kato, Ray Atarashi, Jin Nakazato, Manabu Tsukada, and Hiroshi Esaki. An ontology for spatio-temporal media management and an interactive application. Future Internet, 15(7): 225–??, July 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
7/225.

Stamatopoulos:2024:EAC

[SKFM24] Dionysios Stamatopoulos, Menelaos Katsantonis, Panagiotis Fouliras, and Ioannis Mavridis. Exploring the architectural composition of cyber ranges: a systematic review. Future Internet, 16(7):??, June 28, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/7/231.

Sultana:2023:NHE

[SKHK23] Habiba Sultana, A. H. M. Kamal, Gahangir Hossain, and Muhammad Ashad Kabir. A novel hybrid edge detection and LBP code-based robust image steganography method. Future Internet, 15(3):108, March 10, 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
3/108.

Serketzis:2019:IFT

[SKI^{+}19] Nikolaos Serketzis, Vasilios Katos, Christos Ilioudis, Dimitrios Baltatzis, and Georgios Pangalos. Improving forensic triage efficiency through cyber threat intelligence. Future Internet, 11(7): 162, July 23, 2019. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
7/162.

Schizas:2022:TUL

[SKKS22] Nikolaos Schizas, Aristeidis Karras, Christos Karras, and Spyros Sioutas. TinyML for ultra-low power AI and large scale IoT deployments: a systematic review. Future Internet, 14(12):363, December 06, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
12/363.

Spruit:2021:ABG

[SKM21] Marco Spruit, Marcin Kais, and Vincent Menger. Automated business goal extraction from e-mail repositories to bootstrap business understanding. Future Internet, 13(10): 243, September 23, 2021. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
10/243.

Stamou:2019:ANM

[SKT^{+}19] Adamantia Stamou, Grigorios Kakkavas, Konstantinos Tsitseklis, Vasileios Karyotis, and Symeon Papavassiliou. Autonomic network management and cross-layer optimization in software defined radio environments. Future Internet, 11(2):37, February 03, 2019. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/11/2/37.

Su:2024:RSC

[SL24a] Bing Su and Jiwu Liang. Research on secure community opportunity network based on trust model. Future Internet, 16(4):121, April 01, 2024. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/16/
                                        4/121.

Su:2024:WTP

[SL24b] Guoxin Su and Li Liu. Workflow trace profiling and execution time analysis in quantitative verification. Future Internet, 16(9), September 3, 2024. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/16/
                                        9/319.

Sprunger:2024:MRP

[SLB24] John David Sprunger, Alvin Lim, and David M. Bevly. A method for the rapid propagation of emergency event notifications in a long vehicle convoy. Future Internet, 16 (5):154, April 29, 2024. CODEN ???? ISSN 1999-

                                        5903. URL https://www.

mdpi.com/1999-5903/16/
5/154.

Salazar:2022:AML

[SLP^{+}22] Luiz Henrique A. Salazar, Valderi R. Q. Leithardt, Wemerson Delcio Parreira, Anita M. da Rocha Fernandes, Jorge Luis Victória Barbosa, and Sérgio Duarte Correia. Application of machine learning techniques to predict a Patient’s no-show in the healthcare sector. Future Internet, 14(1):3, December 22, 2022. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/14/1/3.

Sandner:2020:RCI

[SLS20] Philipp Sandner, Anna Lange, and Philipp Schulden. The role of the CFO of an industrial company: An analysis of the impact of blockchain technology. Future Internet, 12(8):128, July 30, 2020. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/12/
8/128.

Son:2021:DKG

[SLSK21] Jiseong Son, Chul-Su Lim, Hyoung-Seop Shim, and Ji-Sun Kang. Development of knowledge graph for data management related to flooding disasters using open data. Future Internet, 13(5):124, May 11, 2021. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/13/5/124.

Sun:2017:ABM

[SLW17] Peng Sun, Guangming Li, and Fuqiang Wang. An adaptive back-off mechanism for wireless sensor networks. Future Internet, 9(2):19, June 01, 2017. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/9/
                                        2/19.

Sun:2012:WBG

[SLY^{+}12] Min Sun, Jing Li, Chaowei Yang, Gavin A. Schmidt, Myra Bambacus, Robert Cahalan, Qunying Huang, Chen Xu, Erik U. Noble, and Zhenlong Li. A Webbased geovisual analytical system for climate studies. Future Internet, 4 (4):1069–1085, December 14, 2012. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/4/4/1069.

Sakurai:2023:BCD

[SLZJ23] Guilherme Yukio Sakurai, Jessica Fernandes Lopes, Bruno Bogaz Zarpelão, and Sylvio Barbon Junior. Benchmarking change detector algorithms from different concept drift perspectives. Future Internet, 15(5):169, April 29, 2023.

CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
5/169.

Sun:2021:IVG

[SLZY21] Lili Sun, Xueyan Liu, Min Zhao, and Bo Yang. Interpretable variational graph autoencoder with noninformative prior. Future Internet, 13(2):51, February 18, 2021. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/13/2/51.

Sofia:2019:OPB

[SM19] Rute C. Sofia and Paulo M.

Mendes. An overview on push-based communication models for informationcentric networking. Future Internet, 11(3):74, March 21, 2019. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/11/3/74.

Santos:2023:BBL

[SMB23] André F. Santos, José Marinho, and Jorge Bernardino. Blockchain-based loyalty management system. Future Internet, 15(5):161, April 27, 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
5/161.

Sakellari:2013:ITB

[SMG13] Georgia Sakellari, Christina Morfopoulou, and Erol Gelenbe. Investigating the tradeoffs between power consumption and quality of service in a backbone network. Future Internet, 5(2):268–281, May 24, 2013. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/5/2/268.

Smith:2013:AAN

[Smi13] Danielle Taana Smith. African Americans and network disadvantage: Enhancing social capital through participation on social networking sites. Future Internet, 5(1):56– 66, March 06, 2013. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/5/
1/56.

Smith:2015:OSP

[Smi15] Marlene A. Smith. Output from statistical predictive models as input to eLearning dashboards. Future Internet, 7(2):170– 183, June 02, 2015. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/7/
2/170.

Symvoulidis:2022:HIB

[SMK^{+}22] Chrysostomos Symvoulidis, George Marinos, Athanasios Kiourtis, Argyro Mavrogiorgou, and Dimosthenis Kyriazis. Health-Fetch: an influence-based, context-aware prefetch scheme in citizen-centered health storage clouds. Future Internet, 14(4):112– ??, April 01, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
4/112.

Singh:2022:LLS

[SMN22] Surjit Singh, Vivek Mehla, and Srete Nikolovski. LSS-DNF: a lightweight secure software defined network framework for future Internet in 5G–6G. Future Internet, 14(12):369, December 08, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
12/369.

Silva:2024:EDE

[SMR24] Lucas Silva, José Metrôlho, and Fernando Ribeiro. Efficient data exchange between WebAssembly modules. Future Internet, 16 (9), September 20, 2024. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
9/341.

Singh:2024:SBS

[SMS^{+}24] Radheshyam Singh, Leo Mendiboure, José Soler, Michael Stübert Berger, Tidiane Sylla, Marion Berbineau, and Lars Dittmann. SDN-based secure common emergency service for railway and road coexistence scenarios. Future Internet, 16(4):122, April 02, 2024. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/16/4/122.

Samala:2021:IDU

[SMVP21] Thirupathi Samala, Vijaya Kumar Manupati, Maria Leonilde R. Varela, and Goran Putnik. Investigation of degradation and upgradation models for flexible unit systems: a systematic literature review. Future Internet, 13 (3):57, February 25, 2021. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/13/
                                        3/57.

Sharma:2023:CSD

[SN23a] Sachin Sharma and Avishek Nag. Cognitive software defined networking and network function virtualization and applications. Future Internet, 15(2):78, February 17, 2023. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/15/
                                        2/78.

Si:2023:RBD

[SN23b] Honghao Si and Baoning Niu. Research on blockchain data availability and storage scalability. Future Internet, 15(6):

212–??, June 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
6/212.

Sun:2023:MTA

[SN23c] Yiming Sun and Tatsuo Nakajima. Mitigating technological anxiety through the application of natural interaction in mixed reality systems. Future Internet, 15(6):216– ??, June 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
6/216.

Song:2019:TPA

[SNH^{+}19] Xin Song, Yue Ni, Xiuwei Han, Lei Qin, and Li Dong. Time and power allocation for energy efficiency maximization in wireless-powered full-duplex relay systems. Future Internet, 11(10): 205, September 20, 2019. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
10/205.

Singh:2020:BFB

[SNKG20] Parminder Singh, Anand Nayyar, Avinash Kaur, and Uttam Ghosh. Blockchain and fog based architecture for Internet of Everything in smart cities. Future Internet, 12(4):61, March 26, 2020. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/12/4/61.

Su:2022:CBR

[SNS22] Yue Su, Kien Nguyen, and Hiroo Sekiya. A comparison of blockchain recovery time in static and mobile IoT-blockchain networks. Future Internet, 14(11): 330, November 14, 2022. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/14/
                                        11/330.

Song:2021:TLC

[SNZ21] Yaqin Song, Hong Ni, and Xiaoyong Zhu. Twolevel congestion control mechanism (2LCCM) for information-centric networking. Future Internet, 13(6):149, June 07, 2021. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/13/
                                        6/149.

Salunke:2024:PBP

[SO24] Sanket Vilas Salunke and Abdelkader Ouda. A performance benchmark for the PostgreSQL and MySQL databases. Future Internet, 16(10), October 19, 2024. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/16/10/382.

Shuaieb:2020:RRF

[SOA^{+}20] Wafa Shuaieb, George Oguntala, Ali AlAbdullah, Huthaifa Obeidat, Rameez Asif, Raed A. Abd-Alhameed, Mohammed S. Bin-Melha, and Chakib Kara-Zaïtri. RFID RSS fingerprinting system for wearable human activity recognition. Future Internet, 12(2):33, February 12, 2020. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/12/2/33.

Sheridan:2019:OBR

[SOB^{+}19] Paul Sheridan, Mikael Onsjö, Claudia Becerra, Sergio Jimenez, and George Dueñas. An ontologybased recommender system with an application to the Star Trek television franchise. Future Internet, 11(9):182, August 22, 2019. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/11/9/182.

Soderstrom:2013:DDY

[Söd13] Sylvia Söderström. Digital differentiation in young people’s Internet use — eliminating or reproducing disability stereotypes. Future Internet, 5(2):190– 204, May 07, 2013. CO-DEN ???? ISSN 1999-

5903. URL https://www.

                                        mdpi.com/1999-5903/5/
                                        2/190.

Salah:2024:EEA

[SOE^{+}24] Zaher Salah, Hamza Abu Owida, Esraa Abu Elsoud, Esraa Alhenawi, Suhaila Abuowaida, and Nawaf Alshdaifat. An effective ensemble approach for preventing and detecting phishing attacks in textual form. Future Internet, 16(11):414, November 8, 2024. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/16/11/414.

Sofia:2019:GTI

[Sof19] Rute C. Sofia. Guidelines towards informationdriven mobility management. Future Internet, 11(5):111, May 10, 2019. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/11/
                                        5/111.

Shoeibi:2024:NSD

[SOK24] Mehrdad Shoeibi, Anita Ershadi Oskouei, and Masoud Kaveh. A novel six-dimensional chimp optimization algorithm — deep reinforcement learningbased optimization scheme for reconfigurable intelligent surface-assisted energy harvesting in batteryless IoT networks. Future Internet, 16(12):460, December 6, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
12/460.

Sorschag:2012:FOI

[Sor12] Robert Sorschag. A flexible object-of-interest annotation framework for online video portals. Future Internet, 4(1):179– 215, February 22, 2012. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/4/
1/179.

Strizhov:2016:SPS

[SOR16] Mikhail Strizhov, Zachary Osman, and Indrajit Ray. Substring position search over encrypted cloud data supporting efficient multiuser setup. Future Internet, 8(3):28, July 04, 2016. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/8/
3/28.

Shaibu:2023:PPL

[SOS^{+}23] Farouq E. Shaibu, Elizabeth N. Onwuka, Nathaniel Salawu, Stephen S. Oyewobi, Karim Djouani, and Adnan M. Abu-Mahfouz. Performance of path loss models over mid-band and high-band channels for 5G communication networks: a review. Future Internet, 15(11):362, November 07, 2023. CODEN ????

ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/15/11/362.

Santos-Olmo:2016:ISC

[SOSC^{+}16] Antonio Santos-Olmo, Luis Enrique Sánchez, Ismael Caballero, Sara Camacho, and Eduardo Fernandez-Medina. The importance of the security culture in SMEs as regards the correct management of the security of their assets. Future Internet, 8(3):30, July 07, 2016. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/8/3/30.

Schade:2012:SOI

[SOSK12] Sven Schade, Frank Ostermann, Laura Spinsanti, and Werner Kuhn. Semantic observation integration. Future Internet, 4(3):807–829, September 03, 2012. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/4/3/807.

Santos-Olmo:2016:AAR

[SOSR^{+}16] Antonio Santos-Olmo, Luis Enrique Sánchez, David G. Rosado, Eduardo Fernández-Medina, and Mario Piattini. Applying the actionresearch method to develop a methodology to reduce the installation and maintenance times of information security management systems. Future Internet, 8(3):36, July 22, 2016. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/8/3/36.

Surantha:2022:ISN

[SP22] Nico Surantha and Noffal A. Putra. Integrated SDN-NFV 5G network performance and management-complexity evaluation. Future Internet, 14(12):378, December 14, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/12/378.

Sharma:2011:OLO

[SPB^{+}11] Subhash Sharma, Christopher Pettit, Ian Bishop, Pang Chan, and Falak Sheth. An online landscape object library to support interactive landscape planning. Future Internet, 3(4):319–343, December 20, 2011. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/3/4/319.

Schroth:2011:TTV

[SPC^{+}11] Olaf Schroth, Ellen Pond, Cam Campbell, Petr Cizek, Stephen Bohus, and Stephen R. J. Sheppard. Tool or toy? Virtual globes in landscape planning. Future Internet, 3(4):204–227, October 20, 2011. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/3/4/204.

Satapathy:2022:PSD

[SPC22] Ranjan Satapathy, Shweta Rajesh Pardeshi, and Erik Cambria. Polarity and subjectivity detection with multitask learning and BERT embedding. Future Internet, 14(7):191, June 22, 2022. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/14/7/191.

Serina:2023:SAF

[SPGS23] Lorenzo Serina, Luca Putelli, Alfonso Emilio Gerevini, and Ivan Serina. Synonyms, antonyms and factual knowledge in BERT heads. Future Internet, 15(7):230–??, July 2023. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/15/7/230.

Shrestha:2024:FIC

[SPK24] Shyam Kumar Shrestha, Shiva Raj Pokhrel, and Jonathan Kua. On the fairness of Internet congestion control over WiFi with deep reinforcement learning. Future Internet, 16(9), September 10, 2024. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/16/
                                        9/330.

Suraci:2022:TDD

[SPM^{+}22] Chiara Suraci, Sara Pizzi, Federico Montori, Marco Di Felice, and Giuseppe Araniti. 6G to take the digital divide by storm: Key technologies and trends to bridge the gap. Future Internet, 14 (6):189, June 19, 2022. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
6/189.

Sun:2021:SSA

[SPS^{+}21] Haoli Sun, Bingfeng Pi, Jun Sun, Takeshi Miyamae, and Masanobu Morinaga. SASLedger: a secured, accelerated scalable storage solution for distributed ledger systems. Future Internet, 13(12): 310, November 30, 2021. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
12/310.

Saura:2017:UDM

[SPSS17] José Ramón Saura, Pedro Palos-Sánchez, and Luis Manuel Cerdá Suárez. Understanding the digital marketing environment with KPIs and Web analytics. Future Internet, 9 (4):76, November 04, 2017. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/9/
4/76.

Singh:2023:BEN

[SPT23] Raman Singh, Zeeshan Pervez, and Hitesh Tewari. Blockchain-enabled NextGen service architecture for mobile Internet offload. Future Internet, 15(5): 173, May 05, 2023. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/15/
                                        5/173.

Shah:2017:NID

[SQK^{+}17] Reehan Ali Shah, Yuntao Qian, Dileep Kumar, Munwar Ali, and Muhammad Bux Alvi. Network intrusion detection through discriminative feature selection by using sparse logistic regression. Future Internet, 9(4):81, November 10, 2017. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/9/4/81.

Sindermann:2020:IRB

[SRM20] Cornelia Sindermann, René Riedl, and Christian Montag. Investigating the relationship between personality and technology acceptance with a focus on the smartphone from a gender perspective: Results of an exploratory survey study. Future Internet, 12(7):110, June 30, 2020. CODEN ???? ISSN 1999-

                                        5903. URL https://www.

mdpi.com/1999-5903/12/
7/110.

Sanchez-Romero:2021:SEC

[SRMJ21] Cristina Sánchez-Romero and Eva María Muñoz-Jiménez. Social and educational coexistence in adolescents’ perception in current social problems through networks. Future Internet, 13(6):141, May 27, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/6/141.

Schnell:2023:DSW

[SRS23] Karen Schnell, Kaushik Roy, and Madhuri Siddula. A descriptive study of webpage designs for posting privacy policies for different-sized US hospitals to create an assessment framework. Future Internet, 15(3):112, March 17, 2023. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/15/3/112.

Shahjehan:2024:RMW

[SRS^{+}24] Waleed Shahjehan, Rajkumar Singh Rathore, Syed Waqar Shah, Mohammad Aljaidi, Ali Safaa Sadiq, and Omprakash Kaiwartya. A review on millimeter-wave hybrid beamforming for wireless intelligent transport systems. Future Internet, 16 (9), September 14, 2024.

CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/16/
                                        9/337.

Singh:2023:OMT

[SRSDF23] Satyanand Singh, Joanna Rosak-Szyrocka, István Drotár, and Xavier Fernando. Oceania’s 5G multi-tier fixed wireless access Link’s long-term resilience and feasibility analysis. Future Internet, 15(10):334, October 10, 2023. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/15/10/334.

Somarakis:2014:PIT

[SS14] Giorgos Somarakis and Anastasia Stratigea. Public involvement in taking legislative action as to the spatial development of the tourist sector in Greece — the “OpenGov” platform experience. Future Internet, 6(4):735–759, November 25, 2014. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/6/4/735.

Singh:2015:EST

[SS15] Dilraj Singh and Amardeep Singh. Enhanced secure trusted AODV (ESTA) protocol to mitigate blackhole attack in mobile ad hoc networks. Future Internet, 7(3):342– 362, September 23, 2015.

CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/7/
3/342.

Shivani:2017:RIE

[SS17] J. L. Divya Shivani and Ranjan K. Senapati. Robust image embedded watermarking using DCT and listless SPIHT. Future Internet, 9(3):33, July 12, 2017. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/9/3/33.

Spradling:2022:EFI

[SS22] Matthew Spradling and Jeremy Straub. Evaluation of the factors that impact the perception of online content trustworthiness by income, political affiliation and online usage time. Future Internet, 14(11):320, November 03, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
11/320.

Serrano:2023:PNM

[SS23] Salvatore Serrano and Marco Scarpa. A Petri net model for cognitive radio Internet of Things networks exploiting GSM bands. Future Internet, 15 (3):115, March 21, 2023. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
3/115.

Standl:2021:PMM

[SSB21] Bernhard Standl and Nadine Schlomske-Bodenstein. A pattern mining method for teaching practices. Future Internet, 13(5):106, April 23, 2021. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/13/
                                        5/106.

Shi:2024:SIB

[SSD24] Haohan Shi, Xiyu Shi, and Safak Dogan. Speech inpainting based on multilayer long short-term memory networks. Future Internet, 16(2):63, February 17, 2024. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/16/2/63.

Severino:2023:PHN

[SSDS23] Ricardo Severino, José Simão, Nuno Datia, and António Serrador. Protecting hybrid ITS networks: a comprehensive security approach. Future Internet, 15(12):388, November 30, 2023. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/15/
                                        12/388.

Said:2018:PCE

[SSF^{+}18] Anwar Said, Syed Waqas Haider Shah, Hasan Farooq, Adnan Noor Mian, Ali Imran, and Jon Crowcroft.

Proactive caching at the edge leveraging influential user detection in cellular D2D networks. Future Internet, 10(10):93, September 21, 2018. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/10/10/93.

Saripalle:2020:IUT

[SSH20] Rishi Saripalle, Mehdi Sookhak, and Mahboobeh Haghparast. An interoperable UMLS terminology service using FHIR. Future Internet, 12(11):199, November 16, 2020. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/12/
11/199.

Serghini:2024:DCN

[SSM^{+}24] Omar Serghini, Hayat Semlali, Asmaa Maali, Abdelilah Ghammaz, and Salvatore Serrano. 1-D convolutional neural network-based models for cooperative spectrum sensing. Future Internet, 16(1):14, December 29, 2024. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/16/1/14.

Silva:2019:CPM

[SSO^{+}19] Marianne Silva, Gabriel Signoretti, Julio Oliveira, Ivanovitch Silva, and Daniel G. Costa. A crowdsensing platform for monitoring of vehicular emissions: a smart city perspective. Future Internet, 11(1):13, January 08, 2019. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/11/1/13.

Sanchez:2016:DES

[SSOÁ^{+}16] Luis Enrique Sánchez, Antonio Santos-Olmo, Esther Álvarez, Monica Huerta, Sara Camacho, and Eduardo Fernández-Medina. Development of an expert system for the evaluation of students’ curricula on the basis of competencies. Future Internet, 8(2):22, May 18, 2016. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/8/2/22.

Sehra:2017:AOD

[SSR17] Sukhjit Singh Sehra, Jaiteg Singh, and Hardeep Singh Rai. Assessing Open-StreetMap data using intrinsic quality indicators: An extension to the QGIS processing toolbox. Future Internet, 9(2):15, April 21, 2017. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/9/2/15.

Shilov:2021:RHF

[SSR^{+}21] Lev Shilov, Semen Shanshin, Aleksandr Romanov, Anastasia Fedotova, Anna Kurtukova, Evgeny Kostyuchenko, and Ivan Sidorov. Reconstruction of a 3D human foot shape model based on a video stream using photogrammetry and deep neural networks. Future Internet, 13(12):315, December 14, 2021. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
12/315.

Silva:2019:BUM

[SSS^{+}19] Cristiano M. Silva, Lucas D. Silva, Leonardo A. L. Santos, João F. M. Sarubbi, and Andreas Pitsillides. Broadening understanding on managing the communication infrastructure in vehicular networks: Customizing the coverage using the delta network. Future Internet, 11(1):1, December 20, 2019. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
1/1.

Spradling:2021:PFN

[SSS21] Matthew Spradling, Jeremy Straub, and Jay Strong. Protection from ‘fake news’: The need for descriptive factual labeling for online content. Future Internet, 13(6):142, May 28, 2021. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/13/6/142.

Singh:2023:BEC

[SST23] Raman Singh, Sean Sturley, and Hitesh Tewari. Blockchain-enabled Chebyshev polynomial-based group authentication for secure communication in an Internet of Things network. Future Internet, 15 (3):96, February 28, 2023. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/15/
                                        3/96.

Sun:2020:PEO

[SSX20] Wei Sun, Hui Su, and Huacheng Xie. Policyengineering optimization with visual representation and separation-of-duty constraints in attributebased access control. Future Internet, 12(10):164, September 27, 2020. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/12/
                                        10/164.

Studen:2020:SMQ

[ST20] Laura Studen and Victor Tiberius. Social media, quo vadis? Prospective development and implications. Future Internet, 12 (9):146, August 28, 2020. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/12/
                                        9/146.

Salameh:2022:CTA

[ST22] Ahmed I. Salameh and Mohamed El Tarhuni. From 5G to 6G—challenges, technologies, and applications. Future Internet, 14 (4):117–??, April 12, 2022. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
4/117.

Stach:2022:SIS

[Sta22] Christoph Stach. Special issue on security and privacy in blockchains and the IoT. Future Internet, 14(11):317, November 01, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/11/317.

Stach:2023:DNO

[Sta23] Christoph Stach. Data is the new oil — sort of: a view on why this comparison is misleading and its implications for modern data administration. Future Internet, 15(2):71, February 12, 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
2/71.

Sharma:2021:SMF

[STCP21] Gourav Prateek Sharma, Wouter Tavernier, Didier Colle, and Mario Pickavet. Scheduling for media function virtualization. Future Internet, 13(7):167, June 28, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/7/167.

Shen:2020:SDR

[STG^{+}20] Bingqing Shen, Weiming Tan, Jingzhi Guo, Hongming Cai, Bin Wang, and Shuaihe Zhuo. A study on design requirement development and satisfaction for future virtual world systems. Future Internet, 12(7):112, July 06, 2020. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/12/
7/112.

Sikelis:2021:OBF

[STK21] Konstantinos Sikelis, George E. Tsekouras, and Konstantinos Kotis. Ontology-based feature selection: a survey. Future Internet, 13 (6):158, June 18, 2021. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
6/158.

Stranieri:2022:ISP

[Str22] Silvia Stranieri. An indoor smart parking algorithm based on fingerprinting. Future Internet, 14(6): 185, June 14, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
6/185.

Spyrou:2023:DDS

[STS23] Evangelos D. Spyrou, Ioannis Tsoulos, and Chrysostomos Stylios. Distributed denial of service classification for softwaredefined networking using grammatical evolution. Future Internet, 15(12): 401, December 13, 2023. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
12/401.

Sharma:2022:TSI

[STV^{+}22] Purushottam Sharma, Devesh Tulsian, Chaman Verma, Pratibha Sharma, and Nancy Nancy. Translating speech to Indian sign language using natural language processing. Future Internet, 14(9):253, August 25, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
9/253.

Sadik:2017:CAH

[SU17a] Ahmed R. Sadik and Bodo Urban. Combining adaptive holonic control and ISA-95 architectures to self-organize the interaction in a worker-industrial robot cooperative workcell. Future Internet, 9(3): 35, July 14, 2017. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/9/3/35.

Sadik:2017:FSS

[SU17b] Ahmed R. Sadik and Bodo Urban. Flow shop scheduling problem and solution in cooperative roboticscase-study: One cobot in cooperation with one worker. Future Internet, 9(3):48, August 16, 2017. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/9/
                                        3/48.

Sadik:2017:OBA

[SU17c] Ahmed R. Sadik and Bodo Urban. An ontologybased approach to enable knowledge representation and reasoning in worker-cobot agile manufacturing. Future Internet, 9(4):90, November 24, 2017. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/9/4/90.

Sharma:2022:FWN

[SUF^{+}22] Sachin Sharma, Saish Urumkar, Gianluca Fontanesi, Byrav Ramamurthy, and Avishek Nag. Future wireless networking experiments escaping simulations. Future Internet, 14 (4):120–??, April 14, 2022. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/14/
                                        4/120.

Sufi:2023:NAB

[Suf23] Fahim Sufi. A new AIbased semantic cyber intelligence agent. Future Internet, 15(7):231–??, July 2023. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/15/7/231.

Suleiman:2022:CAF

[Sul22] Husam Suleiman. A costaware framework for QoSbased and energy-efficient scheduling in cloud–fog computing. Future Internet, 14(11):333, November 14, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/11/333.

Satkauskas:2024:MAD

[SV24] Nerijus Satkauskas and Algimantas Venckauskas. Multi-agent dynamic fog service placement approach. Future Internet, 16(7):??, July 13, 2024. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
7/248.

Sousa-Vieira:2023:SCA

[SVFV23] María E. Sousa-Vieira and Manuel Fernández-Veiga. Study of coded ALOHA with multi-user detection under heavy-tailed and correlated arrivals. Future Internet, 15(4):132, March 30, 2023. CODEN ????

ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/15/4/132.

Samuel:2024:FCS

[SVGGO24] Cyril Naves Samuel, François Verdier, Severine Glock, and Patricia Guitton-Ouhamou. A fair crowdsourced automotive data monetization approach using substrate hybrid consensus blockchain. Future Internet, 16(5):156, April 30, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/5/156.

Sayeed:2022:ACI

[SVK^{+}22] Aqsa Sayeed, Chaman Verma, Neerendra Kumar, Neha Koul, and Zoltán Illés. Approaches and challenges in Internet of Robotic Things. Future Internet, 14(9):265, September 14, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
9/265.

S:2022:AEV

[SVV^{+}22] Sofana Reka S, Prakash Venugopal, Ravi V, Hassan Haes Alhelou, Amer Al-Hinai, and Pierluigi Siano. Analysis of electric vehicles with an economic perspective for the future electric market. Future Internet, 14(6):172, May 31, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/6/172.

Sansevere:2021:LPB

[SW21a] Kayla S. Sansevere and Nathan Ward. Linking phubbing behavior to selfreported attentional failures and media multitasking. Future Internet, 13(4): 100, April 14, 2021. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
4/100.

Shekhtman:2021:EBB

[SW21b] Louis Shekhtman and Erez Waisbard. EngraveChain: a blockchainbased tamper-proof distributed log system. Future Internet, 13(6):143, May 29, 2021. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
6/143.

Shi:2023:RST

[SWC^{+}23] Qi Shi, Daheng Wang, Wen Chen, Jinpei Yu, Weiting Zhou, Jun Zou, and Guangzu Liu. Research on spaceborne target detection based on Yolov5 and image compression. Future Internet, 15(3):114, March 19, 2023. CODEN ???? ISSN 1999-

5903. URL https://www.

mdpi.com/1999-5903/15/
3/114.

Song:2018:NSC

[SWD^{+}18] Anping Song, Zuoyu Wu, Xuehai Ding, Qian Hu, and Xinyi Di. Neurologist standard classification of facial nerve paralysis with deep neural networks. Future Internet, 10(11):111, November 16, 2018. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/10/
11/111.

Sun:2019:RMO

[SWGL19] Wei Sun, Shiwei Wei, Huaping Guo, and Hongbing Liu. Role-mining optimization with separationof-duty constraints and security detections for authorizations. Future Internet, 11(9):201, September 19, 2019. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/11/9/201.

Song:2020:BAB

[SWJ20] Yang Song, Yawen Wang, and Dahai Jin. A Bayesian approach based on Bayes minimum risk decision for reliability assessment of Web service composition. Future Internet, 12(12): 221, December 04, 2020. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/12/
12/221.

Scanlan:2022:CHP

[SWP^{+}22] Joel Scanlan, Paul A.

Watters, Jeremy Prichard, Charlotte Hunn, Caroline Spiranovic, and Richard Wortley. Creating honeypots to prevent online child exploitation. Future Internet, 14(4):121– ??, April 14, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
4/121.

Sun:2018:PCD

[SWY^{+}18] Bo Sun, Ming Wei, Chunfeng Yang, Zhihuo Xu, and Han Wang. Personalised and coordinated demandresponsive feeder transit service design: a genetic algorithms approach. Future Internet, 10(7):61, July 01, 2018. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/10/7/61.

Sun:2018:ODD

[SWZ18] Bo Sun, Ming Wei, and Senlai Zhu. Optimal design of demand-responsive feeder transit services with passengers’ multiple time windows and satisfaction. Future Internet, 10(3):30, March 12, 2018. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/10/3/30.

Song:2019:JOP

[SXXH19] Xin Song, Siyang Xu, Zhigang Xie, and Xiuwei Han. Joint optimal power allocation and relay selection scheme in energy harvesting two-way relaying network. Future Internet, 11(2):47, February 15, 2019. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/11/2/47.

Song:2020:SCU

[SXXW20] Xin Song, Lin Xia, Siyang Xu, and Yue Wang. Secure communication for uplink cellular networks assisted with full-duplex device-to-device user. Future Internet, 12(10):175, October 18, 2020. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/12/10/175.

Safitri:2018:ICP

[SYB^{+}18] Cutifa Safitri, Yoshihide Yamada, Sabariah Baharun, Shidrokh Goudarzi, Quang Ngoc Nguyen, Keping Yu, and Takuro Sato. An intelligent content prefix classification approach for quality of service optimization in information-centric networking. Future Internet, 10(4):33, April 09, 2018. CODEN ???? ISSN 1999-

5903. URL https://www.

mdpi.com/1999-5903/10/
4/33.

Shi:2021:CMS

[SYGY21] Yanjun Shi, Hao Yu, Yijia Guo, and Zhiheng Yuan. A collaborative merging strategy with lane changing in multilane freeway on-ramp area with V2X network. Future Internet, 13(5):123, May 10, 2021. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
5/123.

Sun:2021:AGA

[SYLL21] Tianfang Sun, Pin Yang, Mengming Li, and Shan Liao. An automatic generation approach of the cyber threat intelligence records based on multisource information fusion. Future Internet, 13(2):40, February 02, 2021. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
2/40.

Song:2019:PAR

[SYZ19] Xiangpeng Song, Hongbin Yang, and Congcong Zhou. Pedestrian attribute recognition with graph convolutional network in surveillance scenarios. Future Internet, 11(11):245, November 19, 2019. CO-DEN ???? ISSN 1999-

5903. URL https://www.

                                        mdpi.com/1999-5903/11/
                                        11/245.

Sun:2019:HLS

[SZC19] Xiaolei Sun, Yu Zhang, and Jing Chen. High-level smart decision making of a robot based on ontology in a search and rescue scenario. Future Internet, 11(11):230, October 31, 2019. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/11/11/230.

Sun:2018:DTS

[SZFC18] Dong Sun, Kaixin Zhao, Yaming Fang, and Jie Cui. Dynamic traffic scheduling and congestion control across data centers based on SDN. Future Internet, 10(7):64, July 09, 2018. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/10/
                                        7/64.

Shi:2021:RAC

[SZL21] Jia Shi, Xuewen Zeng, and Yang Li. A register access control scheme for SNR system to counter CPA attack based on malicious user blacklist. Future Internet, 13(10):262, October 13, 2021. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/13/
                                        10/262.

Skondras:2023:GSR

[SZT23] Panagiotis Skondras, Panagiotis Zervas, and Giannis Tzimas. Generating synthetic resume data with large language models for enhanced job description classification. Future Internet, 15(11):363, November 09, 2023. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/15/11/363.

Song:2023:PFP

[SZW^{+}23] Beibei Song, Dehua Zhou, Jiahe Wu, Xiaowei Yuan, Yiming Zhu, and Chuansheng Wang. Protecting function privacy and input privacy in the publicly verifiable outsourcing computation of polynomial functions. Future Internet, 15 (4):152, April 21, 2023. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
4/152.

Sun:2023:IPP

[SZZL23] Nigang Sun, Chenyang Zhu, Yuanyi Zhang, and Yining Liu. An identity privacy-preserving scheme against insider logistics data leakage based on onetime-use accounts. Future Internet, 15(11):361, November 05, 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.

                                        mdpi.com/1999-5903/15/
                                        11/361.

Tuffley:2015:EEO

[TA15] David Tuffley and Amy Antonio. Enhancing educational opportunities with computer-mediated assessment feedback. Future Internet, 7(3):294– 306, August 11, 2015. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/7/
                                        3/294.

Trihinas:2021:FTS

[TAAK21] Demetris Trihinas, Michalis Agathocleous, Karlen Avogian, and Ioannis Katakis. FlockAI: a testing suite for ML-driven drone applications. Future Internet, 13(12):317, December 16, 2021. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/13/12/317.

Tadj:2023:EID

[TAD23] Timothy Tadj, Reza Arablouei, and Volkan Dedeoglu. On evaluating IoT data trust via machine learning. Future Internet, 15(9):309, September 12, 2023. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/15/
                                        9/309.

Talamo:2020:BBP

[TADS20] Maurizio Talamo, Franco Arcieri, Andrea Dimitri, and Christian H. Schunck. A blockchain based PKI validation system based on rare events management. Future Internet, 12(2):40, February 14, 2020. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/12/
2/40.

Tagarev:2020:TDC

[Tag20]  Todor Tagarev.  To-

wards the design of a collaborative cybersecurity networked organisation: Identification and prioritisation of governance needs and objectives. Future Internet, 12(4):62, March 28, 2020. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/12/4/62.

Taleb:2023:EMS

[TAHO23] Anas Abu Taleb, Qasem Abu Al-Haija, and Ammar Odeh. Efficient mobile sink routing in wireless sensor networks using bipartite graphs. Future Internet, 15(5):182, May 14, 2023. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
5/182.

Taha:2022:ELE

[TAMA22] Ashraf A. Taha, Hagar O.

Abouroumia, Shimaa A. Mohamed, and Lamiaa A. Amar. Enhancing the lifetime and energy efficiency of wireless sensor networks using Aquila Optimizer algorithm. Future Internet, 14(12):365, December 07, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/12/365.

Taswell:2010:DIM

[Tas10]  Carl Taswell.  A dis-

tributed infrastructure for metadata about metadata: The HDMM architectural style and PORTAL-DOORS system. Future Internet, 2(2):156– 189, June 01, 2010. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/2/
2/156.

Taha:2023:CSS

[TBT^{+}23] Ahmad Taha, Basel Barakat, Mohammad M. A. Taha, Mahmoud A. Shawky, Chun Sing Lai, Sajjad Hussain, Muhammad Zainul Abideen, and Qammer H. Abbasi. A comparative study of single and multi-stage forecasting algorithms for the prediction of electricity consumption using a UK-National Health Service (NHS) hospital dataset. Future Internet, 15(4):134, March 31, 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
4/134.

Tosi:2021:HSA

[TC21] Davide Tosi and Alessandro Siro Campi. How schools affected the COVID-19 pandemic in Italy: Data analysis for Lombardy Region, Campania Region, and Emilia Region. Future Internet, 13(5):109, April 27, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/5/109.

Tatulea:2019:IFB

[TCB^{+}19] Paula Tatulea, Florina Calin, Remus Brad, Lucian Brâncovean, and Mircea Greavu. An image feature-based method for parking lot occupancy. Future Internet, 11(8):169, August 01, 2019. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
8/169.

Tejedor:2019:ASU

[TCG19] Santiago Tejedor, Laura Cervi, and Gerard Gordon. Analysis of the structure and use of digital resources on the Websites of the main football clubs in Europe. Future Internet, 11(5):104, April 29, 2019. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
5/104.

Teng:2018:OAU

[TCH18] Qianru Teng, Yimin Chen, and Chen Huang. Occlusionaware unsupervised learning of monocular depth, optical flow and camera pose with geometric constraints. Future Internet, 10(10):92, September 21, 2018. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/10/10/92.

Tang:2023:IVS

[TCH23] Xiangdong Tang, Fei Chen, and Yunlong He. Intelligent video streaming at network edge: an attention-based multiagent reinforcement learning solution. Future Internet, 15(7):234–??, July 2023. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/15/7/234.

Tu:2022:TOB

[TCYZ22] Youpeng Tu, Haiming Chen, Linjie Yan, and Xinyan Zhou. Task offloading based on LSTM prediction and deep reinforcement learning for efficient edge computing in IoT. Future Internet, 14(2):30, February 2022. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
2/30.

Tefera:2024:ADM

[TDM^{+}24] Minyechil Alehegn Tefera, Amare Mulatie Dehnaw, Yibeltal Chanie Manie, Cheng-Kai Yao, Shegaw Demessie Bogale, and Peng-Chun Peng. Advanced denoising and meta-learning techniques for enhancing smart health monitoring using wearable sensors. Future Internet, 16(8):??, August 5, 2024. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
8/280.

Truong:2021:FFR

[TDT^{+}21] Thanh-Dat Truong, Chi Nhan Duong, Minh-Triet Tran, Ngan Le, and Khoa Luu. Fast flow reconstruction via robust invertible n × n convolution. Future Internet, 13(7):179, July 08, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/7/179.

Tan:2024:EES

[TDW24] Wuzheng Tan, Shenglong Du, and Jian Weng. Enhancing efficiency and security in unbalanced PSI-CA protocols through cloud computing and homomorphic encryption in mobile networks. Future Internet, 16(6):205, June 7, 2024. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/16/6/205.

Topal:2010:TRA

[TEE10] Sebahattin Topal, Ismet Erkmen, and Aydan M. Erkmen. Towards the robotic “Avatar”: An extensive survey of the cooperation between and within networked mobile sensors. Future Internet, 2(3):363–387, September 14, 2010. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/2/3/363.

Temoa:2019:RLB

[TFKY19] Djorwé Témoa, Anna Förster, Kolyang, and Serge Doka Yamigno. A reinforcement learning based intercell interference coordination in LTE networks. Future Internet, 11(1):19, January 17, 2019. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/11/
                                        1/19.

Todoli-Ferrandis:2024:PMI

[TFSBSPSC24] David Todoli-Ferrandis, Javier Silvestre-Blanes, Víctor Sempere-Payá, and Salvador Santonja-Climent. Polling mechanisms for industrial IoT applications in long-range wide-area networks. Future Internet, 16(4):130, April 12, 2024. CODEN ???? ISSN 1999-

                                        5903. URL https://www.

mdpi.com/1999-5903/16/
4/130.

Teets:2013:LRC

[TG13] Michael Teets and Matthew Goldner. Libraries’ role in curating and exposing big data. Future Internet, 5(3):429–438, August 20, 2013. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/5/3/429.

Tomas-Gabarron:2013:OVT

[TGELGH13] Juan-Bautista Tomas-Gabarron, Esteban Egea-Lopez, and Joan Garcia-Haro. Optimization of vehicular trajectories under Gaussian noise disturbances. Future Internet, 5(1):1–20, December 27, 2013. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/5/1/1.

Takano:2019:OSO

[TGS^{+}19] Hirotaka Takano, Ryota Goto, Thin Zar Soe, Nguyen Duc Tuyen, and Hiroshi Asano. Operation scheduling optimization for microgrids considering coordination of their components. Future Internet, 11(11):223, October 24, 2019. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/11/11/223.

Tipantuna:2020:NEE

[TH20] Christian Tipantuña and Xavier Hesselbach. NFVenabled efficient renewable and non-renewable energy management: Requirements and algorithms. Future Internet, 12(10):171, October 14, 2020. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/12/
10/171.

Tashman:2021:CFC

[TH21] Deemah Tashman and Walaa Hamouda. Cascaded κ–μ fading channels with colluding and non-colluding eavesdroppers: Physical-layer security analysis. Future Internet, 13(8):205, August 04, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/8/205.

Tu:2021:IRS

[THL21] Shu-Fen Tu, Ching-Sheng Hsu, and Yu-Tzu Lu. Improving RE-SWOT analysis with sentiment classification: a case study of travel agencies. Future Internet, 13(9):226, August 30, 2021. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
9/226.

Takasuka:2018:DSD

[THT18] Hidenori Takasuka, Koichi Hirai, and Kazumasa Takami. Development of a social DTN for message communication between SNS group members. Future Internet, 10 (4):32, April 04, 2018. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/10/
4/32.

Tran:2022:DRS

[THV^{+}22] Thanh-Nam Tran, Van-Cuu Ho, Thoai Phu Vo, Khanh Ngo Nhu Tran, and Miroslav Voznak. Design of relay switching to combat an eavesdropper in IoT-NOMA wireless networks. Future Internet, 14(3):71, February 24, 2022. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/14/3/71.

Tonnies:2011:SOA

[TKH^{+}11] Sascha Tönnies, Benjamin Köhncke, Patrick Hennig, Ingo Brunkhorst, and Wolf-Tilo Balke. A service oriented architecture for personalized universal media access. Future Internet, 3(2):87–116, April 01, 2011. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/3/2/87.

Tsiropoulou:2018:QEC

[TKT^{+}18] Eirini Eleni Tsiropoulou, George Kousis, Athina Thanou, Ioanna Lykourentzou, and Symeon Papavassiliou. Quality of experience in cyberphysical social systems based on reinforcement learning and game theory. Future Internet, 10(11): 108, November 07, 2018. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/10/
                                        11/108.

Tian:2017:IRB

[TL17] Hao Tian and Peifeng Liang. Improved recommendations based on trust relationships in social networks. Future Internet, 9 (1):9, March 21, 2017. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/9/
                                        1/9.

Touseau:2019:CWT

[TL19] Lionel Touseau and Nicolas Le Sommer. Contribution of the Web of things and of the opportunistic computing to the smart agriculture: a practical experiment. Future Internet, 11(2):33, February 01, 2019. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/11/2/33.

Tang:2024:CSW

[TL24a] Maolin Tang and Wei Li.

Continuous space wireless communication tower placement by hybrid simulated annealing. Future Internet, 16(4):117, March 29, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/4/117.

Toivonen:2024:VAH

[TL24b] Heidi Toivonen and Francesco Lelli. The varieties of agency in human–smart device relationships: The four agency profiles. Future Internet, 16(3):90, March 07, 2024. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/16/3/90.

Tian:2015:DTR

[TLC15] Bo Tian, Kecheng Liu, and Yuanzhong Chen. Dynamical trust and reputation computation model for B2C e-commerce. Future Internet, 7(4):405– 428, October 27, 2015. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/7/
4/405.

Thapa:2020:CML

[TLK^{+}20] Niraj Thapa, Zhipeng Liu, Dukka B. KC, Balakrishna Gokaraju, and Kaushik Roy. Comparison of machine learning and deep learning models for network intrusion detection systems. Future Internet, 12(10):167, September 30, 2020. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/12/10/167.

Tomaiuolo:2020:STD

[TLM^{+}20] Michele Tomaiuolo, Gianfranco Lombardo, Monica Mordonini, Stefano Cagnoni, and Agostino Poggi. A survey on troll detection. Future Internet, 12(2):31, February 10, 2020. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/12/2/31.

Trim:2023:IHV

[TLV23] Peter R. J. Trim, Yang-Im Lee, and An Vu. Insights into how Vietnamese retailers utilize social media to facilitate knowledge creation through the process of value co-creation. Future Internet, 15(4):123, March 26, 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
4/123.

Teixeira:2022:VBA

[TMP22] Diogo Teixeira, Silvestre Malta, and Pedro Pinto. A vote-based architecture to generate classified datasets and improve performance of intrusion detection systems based on supervised learning. Future Internet, 14(3):72, February 25, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/3/72.

Trihinas:2024:EDM

[TMS24] Demetris Trihinas, Panagiotis Michael, and Moysis Symeonides. Evaluating DL model scaling trade-offs during inference via an empirical benchmark analysis. Future Internet, 16(12):468, December 13, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/12/468.

Torres-Madronero:2020:CPI

[TMTMB20] Esperanza Milena Torres-Madroñero, Maria C. Torres-Madroñero, and Luz Dary Ruiz Botero. Challenges and possibilities of ICT-mediated assessment in virtual teaching and learning processes. Future Internet, 12(12): 232, December 18, 2020. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/12/
12/232.

Tsourela:2020:ITI

[TN20] Maria Tsourela and Dafni-Maria Nerantzaki. An Internet of Things (IoT) acceptance model. Assessing consumer’s behavior toward IoT products and applications. Future Internet, 12(11):191, November 03, 2020. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/12/11/191.

Tan:2021:ETS

[TOLG21] Yee-Fan Tan, Lee-Yeng Ong, Meng-Chew Leow, and Yee-Xian Goh. Exploring time-series forecasting models for dynamic pricing in digital signage advertising. Future Internet, 13(10):241, September 22, 2021. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/13/
                                        10/241.

Tomczyk:2021:EDP

[Tom21] Lukasz Tomczyk. Evaluation of digital piracy by youths. Future Internet, 13 (1):11, January 04, 2021. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/13/
                                        1/11.

Tosi:2023:ESI

[Tos23] Davide Tosi. Editorial for the special issue on “Software Engineering and Data Science”, volume II. Future Internet, 15(9):312, September 16, 2023. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.

mdpi.com/1999-5903/15/
9/312.

Tsoumanis:2020:ITI

[TPD^{+}20] Georgios Tsoumanis, Asterios Papamichail, Vasileios Dragonas, George Koufoudakis, Constantinos T. Angelis, and Konstantinos Oikonomou. Implementation of a Topology Independent MAC (TiMAC) policy on a low-cost IoT system. Future Internet, 12(5):86, May 11, 2020. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/12/
5/86.

Tihinen:2021:DMC

[TPJ21] Maarit Tihinen, Ari Pikkarainen, and Jukka Joutsenvaara. Digital manufacturing challenges education — Smart-Lab concept as a concrete example in tackling these challenges. Future Internet, 13(8):192, July 26, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/8/192.

Tchakounte:2020:CRF

[TPKA20] Franklin Tchakounté, Athanase Esdras Yera Pagor, Jean Claude Kamgang, and Marcellin Atemkeng. CIAA-RepDroid: A fine-grained and probabilistic reputation scheme for Android apps based on sentiment analysis of reviews. Future Internet, 12 (9):145, August 27, 2020. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/12/
                                        9/145.

Thakur:2023:CAI

[TPP^{+}23] Nirmalya Thakur, Kesha A. Patel, Audrey Poon, Rishika Shah, Nazif Azizi, and Changhee Han. A comprehensive analysis and investigation of the public discourse on Twitter about exoskeletons from 2017 to 2023. Future Internet, 15(10): 346, October 22, 2023. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/15/
                                        10/346.

Tran:2018:OMD

[Tra18] Tien Anh Tran. The optimization of marine diesel engine rotational speed control process by fuzzy logic control based on particle swarm optimization algorithm. Future Internet, 10(10):99, October 04, 2018. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/10/10/99.

Tomas:2022:DAE

[TRB22] Joana Tomás, Deolinda Rasteiro, and Jorge Bernardino. Data anonymization: an experimental evaluation using open-source tools. Future Internet, 14(6):

167, May 30, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
6/167.

Tsai:2021:SBE

[TRC^{+}21] Pang-Wei Tsai, Aris Cahyadi Risdianto, Meng Hui Choi, Satis Kumar Permal, and Teck Chaw Ling. SD-BROV: an enhanced BGP hijacking protection with route validation in software-defined eXchange. Future Internet, 13(7):171, June 30, 2021. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
7/171.

Treiblmaier:2021:ENW

[Tre21] Horst Treiblmaier. Exploring the next wave of blockchain and distributed ledger technology: The overlooked potential of scenario analysis. Future Internet, 13(7):183, July 19, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/7/183.

Treiblmaier:2022:WCA

[Tre22] Horst Treiblmaier. What is coming across the horizon and how can we handle it? Bitcoin scenarios as a starting point for rigorous and relevant research. Future Internet, 14(6):162, May 26, 2022.

CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/14/
                                        6/162.

Tavakolian:2023:HRL

[TRHU23] Alireza Tavakolian, Alireza Rezaee, Farshid Hajati, and Shahadat Uddin. Hospital readmission and length-of-stay prediction using an optimized hybrid deep model. Future Internet, 15(9):304, September 06, 2023. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/15/9/304.

Trinugroho:2014:IIP

[Tri14] Yohanes Baptista Dafferianto Trinugroho. Information integration platform for patient-centric healthcare services: Design, prototype and dependability aspects. Future Internet, 6(1):126– 154, March 06, 2014. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/6/
                                        1/126.

Toman:2024:LAS

[TRLR24] Patrick Toman, Nalini Ravishanker, Nathan Lally, and Sanguthevar Rajasekaran. Latent autoregressive student- t prior process models to assess impact of interventions in time series. Future Internet, 16(1):8, December 28, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/1/8.

Tsikteris:2024:TMT

[TRST24] Sean Tsikteris, Aisha B.

Rahman, Md. Sadman Siraj, and Eirini Eleni Tsiropoulou. TRUST-ME: Trust-based resource allocation and server selection in multi-access edge computing. Future Internet, 16(8):??, August 4, 2024. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
8/278.

Tomer:2022:DIA

[TS22] Vikas Tomer and Sachin Sharma. Detecting IoT attacks using an ensemble machine learning model. Future Internet, 14(4): 102–??, March 24, 2022. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
4/102.

Tomer:2024:RIM

[TSD24] Vikas Tomer, Sachin Sharma, and Mark Davis. Resilience in the Internet of Medical Things: a review and case study. Future Internet, 16(11):430, November 20, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
11/430.

Tropea:2022:SWS

[TSDG22] Mauro Tropea, Mattia Giovanni Spina, Floriano De Rango, and Antonio Francesco Gentile. Security in wireless sensor networks: a cryptography performance analysis at MAC layer. Future Internet, 14(5):145, May 10, 2022. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/14/5/145.

Takabatake:2024:TEN

[TSHO24] Wataru Takabatake, Yohei Shibata, Kenji Hoshino, and Tomoaki Ohtsuki. Time-efficient neural-network-based dynamic area optimization algorithm for high-altitude platform station mobile communications. Future Internet, 16 (9), September 11, 2024. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/16/
                                        9/332.

Tsimpourla:2014:RDL

[Tsi14] Clare Tsimpourla. Rikki don’t lose that number: Enumerated human rights in a society of infinite connections. Future Internet, 6(3):482–497, August 19, 2014. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/6/3/482.

Teixeira:2018:SST

[TSZ^{+}18] Marcio Andrey Teixeira, Tara Salman, Maede Zolanvari, Raj Jain, Nader Meskin, and Mohammed Samaka. SCADA system testbed for cybersecurity research using machine learning approach. Future Internet, 10(8):76, August 09, 2018. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/10/8/76.

Trofimova:2022:ERA

[TT22] Yelena Trofimova and Pavel Tvrdík. Enhancing reactive ad hoc routing protocols with trust. Future Internet, 14(1):28, January 15, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
1/28.

Tsoulos:2024:UOT

[TTK24] Ioannis G. Tsoulos, Alexandros Tzallas, and Evangelos Karvounis. Using optimization techniques in grammatical evolution. Future Internet, 16(5): 172, May 16, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
5/172.

Tang:2019:NTC

[TTKZ19] Lujie Tang, Bing Tang, Linyao Kang, and Li Zhang.

A novel task caching and migration strategy in multi-access edge computing based on the genetic algorithm. Future Internet, 11(8):181, August 20, 2019. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/11/8/181.

Tara:2024:DSO

[TTN24] Andrei Tara, Hjalmar K. Turesson, and Nicolae Natea. Dynamic storage optimization for communication between AI agents. Future Internet, 16(8):??, August 1, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
8/274.

Tanaka:2019:MSA

[TTO^{+}19] Kazuaki Tanaka, Kota Takenouchi, Kohei Ogawa, Yuichiro Yoshikawa, Shuichi Nishio, and Hiroshi Ishiguro. Maintaining the sense of agency in semiautonomous robot conferencing. Future Internet, 11(7):143, July 03, 2019. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
7/143.

Tseng:2024:CTD

[TTS24] Chinyang Henry Tseng, Woei-Jiunn Tsaur, and Yueh-Mao Shen. Classification tendency difference index model for feature selection and extraction in wireless intrusion detection. Future Internet, 16 (1):25, January 12, 2024. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
1/25.

Tirziu:2024:EFD

[TVAT24] Eugenia Tîrziu, Ana-Mihaela Vasilevschi, Adriana Alexandru, and Eleonora Tudora. Enhanced fall detection using YOLOv7-W6-pose for real-time elderly monitoring. Future Internet, 16(12):472, December 19, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
12/472.

Tariq:2020:VAU

[TWM20] Muhammad Atiq Ur Rehman Tariq, Cheuk Yin Wai, and Nitin Muttil. Vulnerability assessment of ubiquitous cities using the analytic hierarchy process. Future Internet, 12(12):235, December 21, 2020. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/12/
12/235.

Tseng:2024:MCI

[TWW24] Shu-Ming Tseng, Yan-Qi Wang, and Yung-Chung Wang. Multi-class intrusion detection based on transformer for IoT networks using CIC-IoT-2023 Dataset. Future Internet, 16(8):??, August 8, 2024. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/16/
                                        8/284.

Truyen:2023:VAR

[TXJ23] Eddy Truyen, Hongjie Xie, and Wouter Joosen. Vendor-agnostic reconfiguration of Kubernetes clusters in cloud federations. Future Internet, 15(2):63, February 01, 2023. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/15/
                                        2/63.

Tian:2023:RSP

[TYH23] Zhenyu Tian, Jiali You, and Linlin Hu. A reverse shortest path tree-based multicast joining node selection method. Future Internet, 15(5):156, April 23, 2023. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/15/5/156.

Trajanovski:2022:ABB

[TZ22] Tolijan Trajanovski and Ning Zhang. An automated behaviour-based clustering of IoT botnets. Future Internet, 14(1):6, December 23, 2022. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.

mdpi.com/1999-5903/14/
1/6.

Tsourma:2021:AEF

[TZE^{+}21] Maria Tsourma, Alexandros Zamichos, Efthymios Efthymiadis, Anastasios Drosou, and Dimitrios Tzovaras. An AI-enabled framework for real-time generation of news articles based on big EO data for disaster reporting. Future Internet, 13(6):161, June 19, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/6/161.

Theodorou:2021:AES

[TZS^{+}21] Traianos-Ioannis Theodorou, Alexandros Zamichos, Michalis Skoumperdis, Anna Kougioumtzidou, Kalliopi Tsolaki, Dimitris Papadopoulos, Thanasis Patsios, George Papanikolaou, Athanasios Konstantinidis, Anastasios Drosou, and Dimitrios Tzovaras. An AIenabled stock prediction platform combining news and social sensing with financial statements. Future Internet, 13(6):138, May 21, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/6/138.

Utsu:2020:TBS

[UANU20] Keisuke Utsu, Mariko Abe, Shuji Nishikawa, and Osamu Uchida. Twitterbased safety confirmation system for disaster situations. Future Internet, 12 (1):14, January 16, 2020. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/12/
1/14.

Uslander:2018:ASE

[UB18] Thomas Usländer and Thomas Batz. Agile service engineering in the industrial Internet of Things. Future Internet, 10(10):100, October 09, 2018. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/10/10/100.

Udugama:2019:MST

[UDF^{+}19] Asanga Udugama, Jens Dede, Anna Förster, Vishnupriya Kuppusamy, Koojana Kuladinithi, Andreas Timm-Giel, and Zeynep Vatandas. My Smartphone tattles: Considering popularity of messages in opportunistic data dissemination. Future Internet, 11 (2):29, January 29, 2019. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
2/29.

Urushidani:2010:DRA

[UFK^{+}10] Shigeo Urushidani, Kensuke Fukuda, Michihiro Koibuchi, Motonori Nakamura, Shunji Abe, Yusheng Ji, Michihiro Aoki, and Shigeki Yamada. Dynamic resource allocation and QoS control capabilities of the Japanese academic backbone network. Future Internet, 2(3):295– 307, August 09, 2010. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/2/
3/295.

Hassan:2018:IMF

[UGJA18] Tehseen Ul Hassan, Fei Gao, Babur Jalal, and Sheeraz Arif. Interference management in femtocells by the adaptive network sensing power control technique. Future Internet, 10 (3):25, March 01, 2018. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/10/
3/25.

Uhm:2022:ZIP

[UJ22] Daiho Uhm and Sunghae Jun. Zero-inflated patent data analysis using generating synthetic samples. Future Internet, 14(7):211, July 16, 2022. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
7/211.

Islam:2024:HSS

[UMS24] Rafiq Ul Islam, Pasquale Mazzei, and Claudio Savaglio. Healthiness and safety of smart environments through edge intelligence and Internet of Things technologies. Future Internet, 16(10), October 14, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
10/373.

Utomo:2023:FAT

[URH^{+}23] Sapdo Utomo, Adarsh Rouniyar, Hsiu-Chun Hsu, , and Pao-Ann Hsiung. Federated adversarial training strategies for achieving privacy and security in sustainable smart city applications. Future Internet, 15(11):371, November 20, 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
11/371.

Uchida:2019:VDT

[USS19] Noriki Uchida, Goshi Sato, and Yoshitaka Shibata. Vehicular delaytolerant networks with image recognition-based adaptive array antenna for winter road surveillance in local areas. Future Internet, 11(9):203, September 19, 2019. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/11/9/203.

Uto:2013:MRI

[Uto13] Nelson Uto. A methodology for retrieving information from malware encrypted output files: Brazilian case studies. Future Internet, 5(2):140– 167, April 25, 2013. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/5/
2/140.

Utsu:2019:TWW

[UUT^{+}19] Keisuke Utsu, Shun Ueta, Sachi Tajima, Yoshitaka Kajita, Yuji Murakami, and Osamu Uchida. Townwatching workshop using disaster information tweeting and mapping system. Future Internet, 11(7): 150, July 07, 2019. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
7/150.

Ulven:2021:SRC

[UW21] Joachim Bjørge Ulven and Gaute Wangen. A systematic review of cybersecurity risks in higher education. Future Internet, 13(2):39, February 02, 2021. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/13/2/39.

Vasilas:2024:BHS

[VBB24a] Teodora Vasilas, Claudiu Bacila, and Remus Brad. Beat the heat: Syscall attack detection via thermal side channel. Future Internet, 16(8):??, August 21, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/8/301.

Vasilas:2024:CVA

[VBB24b] Teodora Vasilas, Claudiu Bacila, and Remus Brad. Correction: Vasilas et al. Beat the Heat: Syscall Attack Detection via Thermal Side Channel. Future Internet 2024, 16, 301. Future Internet, 16(10), October 21, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
10/383.

Volosciuc:2024:GIB

[VBB^{+}24c] Cristian Volosciuc, Razvan Bogdan, Bianca Blajovan, Cristina Stângaciu, and Marius Marcu. GreenLab, an IoT-based small-scale smart greenhouse. Future Internet, 16(6):195, May 31, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/6/195.

Venkatraman:2022:SCT

[VBWW22] Sitalakshmi Venkatraman, Fahri Benli, Ye Wei, and Fiona Wahr. Smart classroom teaching strategy to enhance higher order thinking skills (HOTS) — an agile approach for Education 4.0. Future Internet, 14(9):255, August 28, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/9/255.

Vangelista:2019:WCI

[VC19] Lorenzo Vangelista and Marco Centenaro. Worldwide connectivity for the Internet of Things through LoRaWAN. Future Internet, 11(3):57, March 02, 2019. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/11/3/57.

Vizzari:2023:PSR

[VC23] Giuseppe Vizzari and Thomas Cecconello. Pedestrian simulation with reinforcement learning: a curriculum-based approach. Future Internet, 15(1):12, December 27, 2023. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/15/1/12.

Viscusi:2014:ASV

[VCB14] Gianluigi Viscusi, Marco Castelli, and Carlo Batini. Assessing social value in open data initiatives: a framework. Future Internet, 6(3):498–517, August 19, 2014. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/6/3/498.

Vaiani:2024:ERV

[VCG24] Lorenzo Vaiani, Luca Cagliero, and Paolo Garza. Emotion recognition from videos using multimodal large language models. Future Internet, 16(7): ??, July 13, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
7/247.

Villegas-Ch:2021:IVA

[VCGOMC^{+}21] William Villegas-Ch, Joselin García-Ortiz, Karen Mullo-Ca, Santiago Sánchez-Viteri, and Milton Roman-Cañizares. Implementation of a virtual assistant for the academic management of a university with the use of artificial intelligence. Future Internet, 13(4):97, April 13, 2021. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
4/97.

Villegas-Ch:2022:AST

[VCJAMN22] William Villegas-Ch., Angel Jaramillo-Alcázar, and Aracely Mera-Navarrete. Assistance system for the teaching of natural numbers to preschool children with the use of artificial intelligence algorithms. Future Internet, 14(9):266, September 15, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.

mdpi.com/1999-5903/14/
9/266.

Villegas-Ch:2020:ITM

[VCPPRC20] William Villegas-Ch, Xavier Palacios-Pacheco, and Milton Román-Cañizares. An Internet of Things model for improving process management on university campus. Future Internet, 12(10):162, September 25, 2020. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/12/10/162.

Villarejo-Carballido:2022:KCL

[VCPT22] Beatriz Villarejo-Carballido, Cristina M. Pulido, and Santiago Tejedor. Key competences for lifelong learning through the “Animal crossing: New horizons” video game. Future Internet, 14(11):329, November 13, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
11/329.

Villegas-Ch:2021:ASL

[VCRCSV^{+}21] William Villegas-Ch., Milton Roman-Cañizares, Santiago Sánchez-Viteri, Joselin García-Ortiz, and Walter Gaibor-Naranjo. Analysis of the state of learning in University students with the use of a Hadoop framework. Future Internet, 13(6):140, May 24, 2021. CODEN ????

ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/13/6/140.

Vogel:2020:WOI

[VDE^{+}20] Bahtijar Vogel, Yuji Dong, Blerim Emruli, Paul Davidsson, and Romina Spalazzese. What is an open IoT platform? Insights from a systematic mapping study. Future Internet, 12(4):73, April 18, 2020. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/12/4/73.

Vanneste:2019:DUS

[VdHH^{+}19] Simon Vanneste, Jens de Hoog, Thomas Huybrechts, Stig Bosmans, Reinout Eyckerman, Muddsair Sharif, Siegfried Mercelis, and Peter Hellinckx. Distributed uniform streaming framework: An elastic fog computing platform for event stream processing and platform transparency. Future Internet, 11(7):158, July 19, 2019. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/11/
                                        7/158.

Vogel:2021:RCD

[VDKL21] Elisabeth Vogel, Zoya Dyka, Dan Klann, and Peter Langendörfer. Resilience in the cyberworld: Definitions, features and models. Future Internet, 13(11):293, November 19, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/11/293.

vanderSchyff:2023:OPF

[vdSFRF23] Karl van der Schyff, Greg Foster, Karen Renaud, and Stephen Flowerday. Online privacy fatigue: a scoping review and research agenda. Future Internet, 15(5):164, April 28, 2023. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/15/5/164.

Vinceslas:2023:ADD

[VDSK23] Leny Vinceslas, Safak Dogan, Srikumar Sundareshwar, and Ahmet M. Kondoz. Abstracting data in distributed ledger systems for higher level analytics and visualizations. Future Internet, 15(1):33, January 11, 2023. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/15/1/33.

Vemuru:2022:ICE

[Vem22] Krishnamurthy V. Vemuru. Implementation of the canny edge detector using a spiking neural network. Future Internet, 14(12):371, December 11, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/12/371.

Vestoso:2018:GBP

[Ves18] Margherita Vestoso. The GDPR beyond privacy: Data-driven challenges for social scientists, legislators and policy-makers. Future Internet, 10(7):62, July 06, 2018. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/10/7/62.

Vo:2019:ISA

[VFFHF19] Tri Hoang Vo, Woldemar Fuhrmann, Klaus-Peter Fischer-Hellmann, and Steven Furnell. Identityas-a-service: An adaptive security infrastructure and privacy-preserving user identity for the cloud environment. Future Internet, 11(5):116, May 15, 2019. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/11/
                                        5/116.

Voulgaris:2019:BTI

[VFS^{+}19] Spyros Voulgaris, Nikos Fotiou, Vasilios A. Siris, George C. Polyzos, Mikael Jaatinen, and Yannis Oikonomidis. Blockchain technology for intelligent environments. Future Internet, 11(10):213, October 11, 2019. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/11/10/213.

Veglis:2020:SEO

[VG20] Andreas Veglis and Dimitrios Giomelakis. Search engine optimization. Future Internet, 12(1):6, December 31, 2020. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/12/1/6.

Varadarajan:2021:EDC

[VGK21] Vijayakumar Varadarajan, Dweepna Garg, and Ketan Kotecha. An efficient deep convolutional neural network approach for object detection and recognition using a multi-scale anchor box in real-time. Future Internet, 13(12):307, November 29, 2021. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
12/307.

Vila:2019:RTM

[VGS^{+}19] Gaël Vila, Christelle Godin, Oumayma Sakri, Etienne Labyt, Audrey Vidal, Sylvie Charbonnier, Simon Ollander, and Aurélie Campagne. Realtime monitoring of passenger’s psychological stress. Future Internet, 11(5):102, April 26, 2019. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
5/102.

Vovides:2016:ELU

[VI16] Yianna Vovides and Sarah Inman. Elusive learningusing learning analytics to support reflective sensemaking of illstructured ethical problems: a learner-managed dashboard solution. Future Internet, 8(2):26, June 11, 2016. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/8/2/26.

Vgena:2019:TAL

[VKK^{+}19] Katerina Vgena, Angeliki Kitsiou, Christos Kalloniatis, Dimitris Kavroudakis, and Stefanos Gritzalis. Toward addressing location privacy issues: New affiliations with social and location attributes. Future Internet, 11(11): 234, November 01, 2019. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
11/234.

Vidros:2017:ADO

[VKKA17] Sokratis Vidros, Constantinos Kolias, Georgios Kambourakis, and Leman Akoglu. Automatic detection of online recruitment frauds: Characteristics, methods, and a public dataset. Future Internet, 9 (1):6, March 03, 2017. CO-DEN ???? ISSN 1999-

5903. URL https://www.

mdpi.com/1999-5903/9/
1/6.

Vgena:2022:DRS

[VKKG22] Katerina Vgena, Angeliki Kitsiou, Christos Kalloniatis, and Stefanos Gritzalis. Determining the role of social identity attributes to the protection of users’ privacy in social media. Future Internet, 14(9):249, August 24, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
9/249.

Vladyko:2019:DEC

[VKM^{+}19] Andrei Vladyko, Abdukodir Khakimov, Ammar Muthanna, Abdelhamied A. Ateya, and Andrey Koucheryavy. Distributed edge computing to assist ultra-low-latency VANET applications. Future Internet, 11(6):128, June 04, 2019. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
6/128.

Vryzas:2022:PWA

[VKV^{+}22] Nikolaos Vryzas, Anastasia Katsaounidou, Lazaros Vrysis, Rigas Kotsakis, and Charalampos Dimoulas. A prototype Web application to support human-centered audiovisual content authentication and crowdsourcing. Future Internet, 14 (3):75, February 27, 2022. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/14/
                                        3/75.

Velusamy:2018:DCA

[VL18a] Gandhimathi Velusamy and Ricardo Lent. Dynamic cost-aware routing of Web requests. Future Internet, 10(7):57, June 21, 2018. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/10/7/57.

Vishwakarma:2018:EJM

[VL18b] Gopal Vishwakarma and Wonjun Lee. Exploiting JTAG and its mitigation in IOT: a survey. Future Internet, 10(12):121, December 03, 2018. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/10/
                                        12/121.

Velusamy:2020:SSD

[VL20] Gandhimathi Velusamy and Ricardo Lent. Smart site diversity for a high throughput satellite system with software-defined networking and a virtual network function. Future Internet, 12(12):225, December 07, 2020. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.

mdpi.com/1999-5903/12/
12/225.

Vegas:2024:OCA

[VL24] Jesús Vegas and César Llamas. Opportunities and challenges of artificial intelligence applied to identity and access management in industrial environments. Future Internet, 16(12):469, December 16, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/12/469.

Vazquez-Lopez:2021:IRA

[VLBRMP21] Alba Vázquez-López, Martín Barrasa-Rioja, and Manuel Marey-Perez. ICT in rural areas from the perspective of dairy farming: a systematic review. Future Internet, 13(4):99, April 13, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/4/99.

Volk:2019:SIA

[VLC^{+}19] Florian Völk, Konstantinos Liolis, Marius Corici, Joe Cahill, Thomas Schlichter Robert T.Schwarz , Eric Troudt, and Andreas Knopp. Satellite integration into 5G: Accent on first over-the-air tests of an edge node concept with integrated satellite backhaul. Future Internet, 11(9):193, September 05, 2019. CODEN ????

ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/11/9/193.

Vrahatis:2024:GAN

[VLK24] Aristidis G. Vrahatis, Konstantinos Lazaros, and Sotiris Kotsiantis. Graph attention networks: a comprehensive review of methods and applications. Future Internet, 16(9), September 3, 2024. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/16/
                                        9/318.

Vazquez-Lopez:2021:FAG

[VLMP21] Alba Vázquez-López and Manuel Marey-Perez. Factors affecting e-government adoption by dairy farmers: a case study in the North-West of Spain. Future Internet, 13(8):206, August 05, 2021. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/13/
                                        8/206.

Vilella:2020:IDT

[VLPR20] Salvatore Vilella, Mirko Lai, Daniela Paolotti, and Giancarlo Ruffo. Immigration as a divisive topic: Clusters and content diffusion in the Italian Twitter debate. Future Internet, 12(10):173, October 15, 2020. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/12/10/173.

Vergallo:2024:MEC

[VM24a] Roberto Vergallo and Luca Mainetti. Measuring the effectiveness of carbonaware AI training strategies in cloud instances: a confirmation study. Future Internet, 16(9), September 13, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/9/334.

Vizzarri:2024:UCS

[VM24b] Alessandro Vizzarri and Franco Mazzenga. 6G use cases and scenarios: a comparison analysis between ITU and other initiatives. Future Internet, 16(11):404, November 1, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/11/404.

Villafuerte:2023:AIV

[VMAG23] Naythan Villafuerte, Santiago Manzano, Paulina Ayala, and Marcelo V. García. Artificial intelligence in virtual telemedicine triage: a respiratory infection diagnosis tool with electronic measuring device. Future Internet, 15 (7):227–??, July 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
7/227.

Volkau:2022:INS

[VMD^{+}22] Ihar Volkau, Abdul Mujeeb, Wenting Dai, Marius Erdt, and Alexei Sourin. The impact of a number of samples on unsupervised feature extraction, based on deep learning for detection defects in printed circuit boards. Future Internet, 14(1):8, December 23, 2022. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/14/1/8.

Varlamis:2022:SUG

[VMGT22] Iraklis Varlamis, Dimitrios Michail, Foteini Glykou, and Panagiotis Tsantilas. A survey on the use of graph convolutional networks for combating fake news. Future Internet, 14 (3):70, February 24, 2022. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/14/
                                        3/70.

Vik:2018:VCF

[VNJ18] Ådne Stenberg Vik, Bjørn Christian Nørbech, and Debora Jeske. Virtual career fairs: Perspectives from Norwegian recruiters and exhibitors. Future Internet, 10(2):19, February 12, 2018. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/10/2/19.

Vitorino:2022:APP

[VOP22] João Vitorino, Nuno Oliveira, and Isabel Praça. Adaptative perturbation patterns: Realistic adversarial learning for robust intrusion detection. Future Internet, 14(4):108– ??, March 29, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
4/108.

Varela:2022:CIN

[VP22] Leonilde Varela and Goran D.

Putnik. Collaborative and intelligent networks and decision systems and services for supporting engineering and production management. Future Internet, 14(11):318, November 02, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/11/318.

Ventruto:2020:FED

[VPCE20] Federica Ventruto, Marco Pulimeno, Massimo Cafaro, and Italo Epicoco. On frequency estimation and detection of heavy hitters in data streams. Future Internet, 12(9):158, September 18, 2020. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/12/
9/158.

Venu:2021:EDC

[VR21] Sagar Kora Venu and Sridhar Ravula. Evaluation of deep convolutional generative adversarial networks for data augmentation of chest X-ray images. Future Internet, 13(1):8, December 31, 2021. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/13/1/8.

Viola:2018:IER

[VRA^{+}18] Fabio Viola, Luca Roffia, Francesco Antoniazzi, Alfredo D’Elia, Cristiano Aguzzi, and Tullio Salmon Cinotti. Interactive 3D exploration of RDF graphs through semantic planes. Future Internet, 10(8):81, August 17, 2018. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/10/8/81.

Vladyko:2021:TPA

[VSE21] Andrei Vladyko, Anastasia Spirkina, and Vasiliy Elagin. Towards practical applications in modeling blockchain system. Future Internet, 13(5):125, May 12, 2021. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/13/5/125.

Vu:2021:SBI

[VSEH^{+}21] Simon Nam Thanh Vu, Mads Stege, Peter Issam El-Habr, Jesper Bang, and Nicola Dragoni. A survey on botnets: Incentives, evolution, detection and current trends. Future Internet, 13(8):198, July 31, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/8/198.

vonSchwerin:2024:SCB

[vSR24] Magdalena von Schwerin and Manfred Reichert. A systematic comparison between open- and closed-source large language models in the context of generating GDPRcompliant data categories for processing activity records. Future Internet, 16(12):459, December 5, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/12/459.

Vasilakopoulos:2024:UAI

[VTPT24] Zisis Vasilakopoulos, Theocharis Tavantzis, Rafail Promikyridis, and Efthimios Tambouris. The use of artificial intelligence in eParticipation: Mapping current research. Future Internet, 16 (6):198, June 3, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
6/198.

Vrysis:2021:WIA

[VVK^{+}21] Lazaros Vrysis, Nikolaos Vryzas, Rigas Kotsakis, Theodora Saridou, Maria Matsiola, Andreas Veglis, Carlos Arcila-Calderón, and Charalampos Dimoulas. A Web interface for analyzing hate speech. Future Internet, 13(3):80, March 22, 2021. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/13/3/80.

Villora:2019:CDA

[VYN19] Beatriz Villora, Santiago Yubero, and Raúl Navarro. Cyber dating abuse and masculine gender norms in a sample of male adults. Future Internet, 11(4):84, March 28, 2019. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/11/4/84.

Wilson:2024:BEP

[WADL^{+}24] Stanly Wilson, Kwabena Adu-Duodu, Yinhao Li, Ringo Sham, Mohammed Almubarak, Yingli Wang, Ellis Solaiman, Charith Perera, Rajiv Ranjan, and Omer Rana. Blockchainenabled provenance tracking for sustainable material reuse in construction supply chains. Future Internet, 16(4):135, April 17, 2024. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/16/4/135.

Wang:2017:SEE

[Wan17] Xinhua Wang. Spectrum and energy efficiency of uplink massive MIMO system with D2D underlay. Future Internet, 9(2):12, April 13, 2017. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/9/2/12.

Wang:2023:ATC

[Wan23] Yang Wang. Advances techniques in computer vision and multimedia. Future Internet, 15(9):294, September 01, 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
9/294.

Wang:2012:VAS

[WBG12] Jue Wang, Keith J. Bennett, and Edward A. Guinness. Virtual astronaut for scientific visualization— a prototype for Santa Maria Crater on Mars. Future Internet, 4(4):1049–1068, December 13, 2012. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/4/
4/1049.

Wu:2023:CAE

[WBK^{+}23] Zixiu Wu, Simone Balloccu, Vivek Kumar, Rim Helaoui, Diego Reforgiato Recupero, and Daniele Riboni. Creation, analysis and evaluation of AnnoMI, a dataset of expertannotated counselling dialogues. Future Internet, 15(3):110, March 14, 2023. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/15/
                                        3/110.

Wiessner:2024:UAT

[WBS^{+}24] Paul Wiessner, Grigor Bezirganyan, Sana Sellami, Richard Chbeir, and Hans-Joachim Bungartz. Uncertainty-aware time series anomaly detection. Future Internet, 16(11): 403, October 31, 2024. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/16/
                                        11/403.

Wang:2017:SME

[WC17] Jinpeng Wang and Gérard Chalhoub. Study of mobility enhancements for RPL in convergecast scenarios. Future Internet, 9(4):86, November 17, 2017. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/9/
                                        4/86.

Wu:2020:HSD

[WC20] Zhijun Wu and Bohua Cui. A hybrid SWIM data naming scheme based on TLC structure. Future Internet, 12(9):142, August 25, 2020. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/12/9/142.

Wang:2023:FWP

[WC23] Sheng-Ming Wang and Wei-Min Cheng. Fast way to predict parking lots availability: For shared parking lots based on dynamic parking fee system. Future Internet, 15(3):89, February 22, 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
3/89.

Wang:2016:DII

[WCF^{+}16] Yun Wang, William Chu, Sarah Fields, Colleen Heinemann, and Zach Reiter. Detection of intelligent intruders in wireless sensor networks. Future Internet, 8(1):2, January 20, 2016. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/8/1/2.

Wang:2019:ADU

[WCM19] Jinpeng Wang, Gérard Chalhoub, and Michel Misson. Adaptive downward/upward routing protocol for mobile-sensor networks. Future Internet, 11(1):18, January 15, 2019. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/11/1/18.

Wang:2018:RBM

[WCSZ18] Chuang Wang, Xu’nan Chen, Abdel-Hamid Ali Soliman, and Zhixiang Zhu. RFID based manufacturing process of cloud MES. Future Internet, 10(11):104, October 30, 2018. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/10/11/104.

Wu:2020:BLS

[WCYL20] Pin Wu, Xuting Chang, Yang Yang, and Xiaoqiang Li. BASN-learning steganography with a binary attention mechanism. Future Internet, 12(3):43, February 27, 2020. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/12/
3/43.

Wang:2023:CST

[WCZ23] Jinlong Wang, Dong Cui, and Qiang Zhang. Chinese short-text sentiment prediction: a study of progressive prediction techniques and attentional fine-tuning. Future Internet, 15(5):158, April 23, 2023. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/15/5/158.

Wei:2024:MWM

[WD24] Zhongcheng Wei and Yanhu Dong. Multi-WiIR: Multi-user identity legitimacy authentication based on WiFi device. Future Internet, 16(4):127, April 08, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/4/127.

Wadatkar:2023:MTV

[WGN23] Prachi V. Wadatkar, Rosario G.

Garroppo, and Gianfranco Nencioni. 5G-MEC testbeds for V2X applications. Future Internet, 15(5):175, May 09, 2023. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
5/175.

Wang:2021:RIG

[WHB^{+}21] Xiaoyuan Wang, Junyan Han, Chenglin Bai, Huili Shi, Jinglei Zhang, and Gang Wang. Research on the impacts of generalized preceding vehicle information on traffic flow in V2X environment. Future Internet, 13(4):88, March 30, 2021. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/13/4/88.

Wheeler:2009:LSM

[Whe09] Steve Wheeler. Learning space mashups: Combining Web 2.0 tools to create collaborative and reflective learning spaces. Future Internet, 1(1):3–13, July 13, 2009. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/1/1/3.

Wang:2022:LCG

[WHW22] Zhihao Wang, Ru Huo, and Shuo Wang. A lightweight certificateless group key agreement method without pairing based on blockchain for smart grid. Future Internet, 14(4):119–??, April 14, 2022. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/14/4/119.

Wang:2018:ECS

[WHXL18] Baocheng Wang, Yafei Hu, Yu Xiao, and Yi Li. An EV charging scheduling mechanism based on price negotiation. Future Internet, 10 (5):40, May 03, 2018. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/10/
                                        5/40.

Wang:2021:AMG

[WJ21] Xiaoliang Wang and Peiquan Jin. Adaptive multigrained buffer management for database systems. Future Internet, 13(12): 303, November 26, 2021. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/13/
                                        12/303.

Wu:2022:CFE

[WJBZ22] Meng Wu, Tingting Jiang, Chenyang Bu, and Bin Zhu. Coarse-to-fine entity alignment for Chinese heterogeneous encyclopedia knowledge base. Future Internet, 14(2):39, January 25, 2022. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/14/2/39.

Walter:2024:DES

[WK24] Markus Walter and Jörg Keller. Design and evaluation of steganographic channels in fifthgeneration new radio. Future Internet, 16(11):410, November 6, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
11/410.

Watorek:2022:MCC

[WKD22] Marcin Watorek, Jarosław Kwapień, and Stanisław Drozdz. Multifractal crosscorrelations of bitcoin and ether trading characteristics in the post-COVID-19 time. Future Internet, 14(7):215, July 21, 2022. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
7/215.

Weichselbraun:2021:ADD

[WKF^{+}21]  Albert  Weichselbraun,

Philipp Kuntschik, Vincenzo Francolino, Mirco Saner, Urs Dahinden, and Vinzenz Wyss. Adapting data-driven research to the fields of social sciences and the humanities. Future Internet, 13(3):59, February 26, 2021. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/13/3/59.

Wu:2018:UHG

[WKPC18] Yu Wu, Jessica Kropczynski, Raquel Prates, and John M. Carroll. Understanding how GitHub supports curation repositories. Future Internet, 10(3):29, March 10, 2018. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/10/
                                        3/29.

Wang:2018:HRA

[WKW18] Wentao Wang, Xuan Ke, and Lingxia Wang. A HMM-R approach to detect L-DDoS attack adaptively on SDN controller. Future Internet, 10(9):83, August 23, 2018. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/10/9/83.

Wang:2020:SBD

[WL20] Jiajia Wang and Guangming Li. Study on bridge displacement monitoring algorithms based on multitargets tracking. Future Internet, 12(1):9, January 08, 2020. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/12/1/9.

Wang:2021:HPP

[WL21] Baocheng Wang and Zetao Li. Healthchain: a privacy protection system for medical data based on blockchain. Future Internet, 13(10):247, September 24, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/10/247.

Wang:2022:NPO

[WL22] Jin Wang and Jun Luo.

No perfect outdoors: Towards a deep profiling of GNSS-based location contexts. Future Internet, 14 (1):7, December 23, 2022. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
1/7.

Wang:2024:RDH

[WLCC24] Xu Wang, Jui-Chuan Liu, Ching-Chun Chang, and Chin-Chen Chang. Reversible data hiding in encrypted 3D mesh models based on multi-group partition and closest pair prediction. Future Internet, 16(6):210, June 15, 2024. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
6/210.

Wu:2017:ACM

[WLCS17] Jiang Wu, Zhou Lei, Shengbo Chen, and Wenfeng Shen. An access control model for preventing virtual machine escape attack. Future Internet, 9(2): 20, June 02, 2017. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/9/
2/20.

Wang:2022:TSP

[WLG^{+}22] Tingting Wang, Zhuolin Li, Xiulin Geng, Baogang Jin, and Lingyu Xu. Time series prediction of sea surface temperature based on an adaptive graph learning neural model. Future Internet, 14(6):171, May 31, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/6/171.

Wang:2017:MIM

[WLH^{+}17] Weijun Wang, Ying Li, Yinghui Huang, Hui Liu, and Tingting Zhang. A method for identifying the mood states of social network users based on cyber psychometrics. Future Internet, 9(2):22, June 16, 2017. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/9/2/22.

Wan:2013:LEC

[WLHR13] Liangtian Wan, Lutao Liu, Guangjie Han, and Joel J. P. C. Rodrigues. A low energy consumption DOA estimation approach for conformal array in ultrawideband. Future Internet, 5(4):611–630, December 16, 2013. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/5/4/611.

Wang:2021:FSP

[WLJM21] Shunli Wang, Rui Li, Jie Jiang, and Yao Meng. Fine-scale population estimation based on building classifications: a case study in Wuhan. Future Internet, 13(10):251, September 28, 2021. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
10/251.

Wang:2019:SBE

[WLL^{+}19] Yufeng Wang, Shuangrong Liu, Songqian Li, Jidong Duan, Zhihao Hou, Jia Yu, and Kun Ma. Stackingbased ensemble learning of self-media data for marketing intention detection. Future Internet, 11(7): 155, July 10, 2019. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
7/155.

Wang:2020:HCA

[WLL20] Baocheng Wang, Zetao Li, and Haibin Li. Hybrid consensus algorithm based on modified proofof-probability and DPoS. Future Internet, 12(8):

122, July 24, 2020. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/12/
                                        8/122.

Wang:2024:MAD

[WLLY24] Huiting Wang, Yazhi Liu, Wei Li, and Zhigang Yang. Multi-agent deep reinforcement learningbased fine-grained traffic scheduling in data center networks. Future Internet, 16(4):119, March 31, 2024. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/16/
                                        4/119.

Wu:2023:MUC

[WLWZ23] Qianqian Wu, Qiang Liu, Zefan Wu, and Jiye Zhang. Maximizing UAV coverage in maritime wireless networks: a multiagent reinforcement learning approach. Future Internet, 15(11):369, November 16, 2023. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/15/11/369.

Wu:2020:SQI

[WLYL20] Zhijun Wu, Rong Li, Panpan Yin, and Changliang Li. Steganalysis of quantization index modulation steganography in G.723.1 codec. Future Internet, 12 (1):17, January 19, 2020. CODEN ???? ISSN 1999-

                                        5903. URL https://www.

mdpi.com/1999-5903/12/
1/17.

Wu:2018:PPM

[WLZ18] Yin Wu, Bowen Li, and Fuquan Zhang. Predictive power management for wind powered wireless sensor node. Future Internet, 10(9):85, September 06, 2018. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/10/9/85.

Wang:2024:SSG

[WML24] Ziyu Wang, Jing Ma, and Edmund M-K Lai. A survey of scenario generation for automated vehicle testing and validation. Future Internet, 16(12):480, December 23, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
12/480.

Wali:2023:TAI

[WMS23] Abdul Wali, Saipunidzam Mahamad, and Suziah Sulaiman. Task automation intelligent agents: a review. Future Internet, 15(6):196–??, June 2023. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
6/196.

Wu:2023:NID

[WQH23] Kongpei Wu, Huiqin Qu, and Conggui Huang. A network intrusion detection method incorporating Bayesian attack graph and incremental learning part. Future Internet, 15 (4):128, March 28, 2023. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
4/128.

Wu:2023:SBB

[WRHH23] Lang Wu, Weijian Ruan, Jinhui Hu, and Yaobin He. A survey on blockchainbased federated learning. Future Internet, 15(12): 400, December 12, 2023. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
12/400.

Wehbe:2021:BBS

[WS21] Rania Wehbe and Isam Shahrour. A BIM-based smart system for fire evacuation. Future Internet, 13 (9):221, August 25, 2021. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
9/221.

Wei:2022:SSL

[WSD22] Dawei Wei, Feifei Shi, and Sahraoui Dhelim. A selfsupervised learning model for unknown Internet traffic identification based on surge period. Future Internet, 14(10):289, October 10, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/10/289.

Wang:2023:DSE

[WSF23] Xu Wang, Bin Shi, and Yili Fang. Distributed systems for emerging computing: Platform and application. Future Internet, 15(4):151, April 20, 2023. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
4/151.

Wang:2020:NRA

[WSM20] Jingpu Wang, Xin Song, and Yatao Ma. A novel resource allocation scheme in NOMA-based cellular network with D2D communications. Future Internet, 12(1):8, January 06, 2020. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/12/1/8.

Warch:2024:ATG

[WSN24] Dominik Warch, Patrick Stellbauer, and Pascal Neis. Advanced techniques for geospatial referencing in online media repositories. Future Internet, 16 (3):87, March 01, 2024. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
3/87.

West:2013:REO

[WT13] Rebecca J. West and Bhoomi K. Thakore. Racial exclusion in the online world. Future Internet, 5 (2):251–267, May 24, 2013. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/5/
2/251.

Weerasinghe:2023:TCB

[WUH^{+}23] Nuwan Weerasinghe, Muhammad Arslan Usman, Chaminda Hewage, Eckhard Pfluegel, and Christos Politis. Threshold cryptography-based secure vehicle-to-everything (V2X) communication in 5G-enabled intelligent transportation systems. Future Internet, 15(5):157, April 23, 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
5/157.

Wang:2020:PMV

[WWB^{+}20] Jiao Wang, Jay Weitzen, Oguz Bayat, Volkan Sevindik, and Mingzhe Li. Performance model for video service in 5G networks. Future Internet, 12(6):99, June 08, 2020. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/12/6/99.

Wang:2021:ITC

[WWD21] Yinfeng Wang, Longxiang Wang, and Xiaoshe Dong. An intelligent TCP congestion control method based on deep Q network.

Future Internet, 13(10): 261, October 09, 2021. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
10/261.

Wang:2021:DAA

[WWLZ21] Jing Wang, ZhongCheng Wu, Fang Li, and Jun Zhang. A data augmentation approach to distracted driving detection. Future Internet, 13(1):1, December 22, 2021. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
1/1.

Wang:2019:WSS

[WWYQ19] Huan Wang, Bin Wu, Yuancheng Yao, and Mingwei Qin. Wideband spectrum sensing based on reconfigurable filter bank in cognitive radio. Future Internet, 11(11):244, November 18, 2019. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
11/244.

Wang:2020:UME

[WX20] Xiaofan Wang and Lingyu Xu. Unsteady multielement time series analysis and prediction based on spatial-temporal attention and error forecast fusion. Future Internet, 12 (2):34, February 13, 2020.

CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/12/
                                        2/34.

Wang:2022:EER

[WX22] Yinglin Wang and Xinyu Xu. ERGCN: Enhanced relational graph convolution network, an optimization for entity prediction tasks on temporal knowledge graphs. Future Internet, 14(12):376, December 13, 2022. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/14/12/376.

Wang:2024:MOD

[WX24] Peng Wang and Yonghua Xiong. A method to optimize deployment of directional sensors for coverage enhancement in the sensing layer of IoT. Future Internet, 16(8):??, August 22, 2024. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/16/8/302.

Wang:2022:MLF

[WXL22] Zhao Wang, Qingguo Xu, and Weimin Li. Multilayer feature fusion-based community evolution prediction. Future Internet, 14(4):113–??, April 06, 2022. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/14/4/113.

Wang:2023:DSS

[WXL^{+}23] Mian Wang, Cong’an Xu, Yun Lin, Zhiyi Lu, Jinlong Sun, and Guan Gui. A distributed sensor system based on cloud-edge-end network for industrial Internet of Things. Future Internet, 15(5):171, April 30, 2023. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/15/5/171.

Wu:2019:MIM

[WY19] Cong Wu and Jianhui Yang. Multimedia independent multipath routing algorithms for Internet of Things based on a node hidden communication model. Future Internet, 11(11):240, November 15, 2019. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/11/11/240.

Wu:2018:SMI

[WYL18] Pin Wu, Yang Yang, and Xiaoqiang Li. StegNet: Mega image steganography capacity with deep convolutional network. Future Internet, 10(6):54, June 15, 2018. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/10/6/54.

Wang:2021:DIV

[WYL^{+}21] Juan Wang, Yang Yu, Yi Li, Chengyang Fan, and Shirong Hao. Design and implementation of virtual security function based on multiple enclaves. Future Internet, 13(1):12, January 06, 2021. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/13/1/12.

Wang:2017:EEP

[WYS17] Xinhua Wang, Yan Yang, and Jinlu Sheng. Energy efficient power allocation for the uplink of distributed massive MIMO systems. Future Internet, 9 (2):21, June 09, 2017. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/9/
2/21.

Wei:2022:BTM

[WYW^{+}22] Lijun Wei, Yuhan Yang, Jing Wu, Chengnian Long, and Yi-Bing Lin. A bidirectional trust model for service delegation in Social Internet of Things. Future Internet, 14(5):135, April 29, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/5/135.

Wang:2018:PRA

[WYX18] Bo Wang, Feiyue Ye, and Jialu Xu. A personalized recommendation algorithm based on the user’s implicit feedback in ecommerce. Future Internet, 10(12):117, November 29, 2018. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/10/12/117.

Wang:2024:DST

[WYZ24] Dong Wang, Hongji Yang, and Hua Zhou. Dynamic spatial–temporal self-attention network for traffic flow prediction. Future Internet, 16(6):189, May 25, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
6/189.

Wu:2018:CEE

[WZ18] Yue Wu and Junyi Zhang.

Chinese event extraction based on attention and semantic features: a bidirectional circular neural network. Future Internet, 10(10):95, September 26, 2018. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/10/10/95.

Wang:2019:HTC

[WZC19] Xuan Wang, Bofeng Zhang, and Furong Chang. Hot topic community discovery on cross social networks. Future Internet, 11(3):60, March 04, 2019. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/11/3/60.

Wei:2018:DBG

[WZG^{+}18] Jinlong Wei, Ji Zhou, Elias Giacoumidis, Paul A. Haigh, and Jianming Tang. DSP-based 40 GB/s lane rate next-generation access networks. Future Internet, 10(12):118, November 30, 2018. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/10/
                                        12/118.

Wang:2020:NLC

[WZL^{+}20] Jiong Wang, Hua Zhang, Dongliang Lin, Huibin Feng, Tao Wang, Hongyan Zhang, and Xiaoding Wang. A novel lowcomplexity fault diagnosis algorithm for energy Internet in smart cities. Future Internet, 12(2):26, February 06, 2020. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/12/2/26.

Wu:2021:TLM

[WZL21] Pin Wu, Rukang Zhu, and Zhidan Lei. Transfer learning for multi-premise entailment with relationship processing module. Future Internet, 13(3):71, March 13, 2021. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/13/3/71.

Wang:2023:IPA

[WZL23] Juan Wang, Jing Zhong, and Jiangqi Li. IoTportrait: Automatically identifying IoT devices via transformer with incremental learning. Future Internet, 15(3):102, March 07, 2023. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/15/3/102.

Wu:2019:RSS

[WZLL19] Zhijun Wu, Shengyan Zhou, Liang Liu, and Jin Lei. Research on SWIM services dynamic migration method. Future Internet, 11(9):187, August 27, 2019. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/11/9/187.

Wang:2019:IMS

[WZLW19] Fei Wang, Zhenfang Zhu, Peiyu Liu, and Peipei Wang. Influence maximization in social network considering memory effect and social reinforcement effect. Future Internet, 11 (4):95, April 11, 2019. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
4/95.

Wang:2022:TSR

[WZS^{+}22] Yingxun Wang, Hushairi Zen, Mohamad Faizrizwan Mohd Sabri, Xiang Wang, and Lee Chin Kho. Towards strengthening the resilience of IoV networks — a trust management perspective. Future Internet, 14(7):202, June 30, 2022. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/14/7/202.

Wu:2020:MAR

[WZX20] Zhijun Wu, Yun Zhang, and Enzhong Xu. Multiauthority revocable access control method based on CP-ABE in NDN. Future Internet, 12(1):15, January 16, 2020. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/12/1/15.

Xie:2018:NSA

[XCSM18] Lei Xie, Shengbo Chen, Wenfeng Shen, and Huaikou Miao. A novel selfadaptive VM consolidation strategy using dynamic multi-thresholds in IaaS clouds. Future Internet, 10 (6):52, June 13, 2018. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/10/
                                        6/52.

Xiong:2022:RPB

[XCW^{+}22] Huanliang Xiong, Muxi Chen, Canghai Wu, Yingding Zhao, and Wenlong Yi. Research on progress of blockchain consensus algorithm: a review on recent progress of blockchain consensus algorithms. Future Internet, 14(2):47, January 30, 2022. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/14/2/47.

Xing:2024:MIS

[XDH24] Mengchi Xing, Haojiang Deng, and Rui Han. A method for 5G–ICN seamless mobility support based on router buffered data. Future Internet, 16 (3):96, March 13, 2024. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
3/96.

Xu:2019:DGR

[XDS^{+}19] Haiyan Xu, Yanhui Ding, Jing Sun, Kun Zhao, and Yuanjian Chen. Dynamic group recommendation based on the attention mechanism. Future Internet, 11(9):198, September 17, 2019. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
9/198.

Xu:2019:FLE

[XGN19] Dong Xu, Ruping Ge, and Zhihua Niu. Forwardlooking element recognition based on the LSTM-CRF model with the integrity algorithm. Future Internet, 11(1):17, January 14, 2019. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/11/1/17.

Xia:2022:RUT

[XGY^{+}22] Zhuofei Xia, Jiayuan Gong, Hailong Yu, Wenbo Ren, and Jingnan Wang. Research on urban traffic incident detection based on vehicle cameras. Future Internet, 14(8):227, July 26, 2022. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/14/8/227.

Xie:2010:TFS

[XJ10] Iris Xie and Soohyung Joo. Tales from the field: Search strategies applied in Web searching. Future Internet, 2(3):259– 281, August 06, 2010. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/2/
                                        3/259.

Xevgenis:2023:AZS

[XKKL23] Michael Xevgenis, Dimitrios G. Kogias, Panagiotis A. Karkazis, and Helen C. Leligou. Addressing ZSM security issues with blockchain technology. Future Internet, 15 (4):129, March 28, 2023. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/15/
                                        4/129.

Xu:2022:MLP

[XLJ^{+}22] Xiaoting Xu, Tin Lai, Sayka Jahan, Farnaz Farid, and Abubakar Bello. A machine learning predictive model to detect water quality and pollution. Future Internet, 14(11):324, November 08, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/11/324.

Xu:2023:DBM

[XLM^{+}23] Tongyang Xu, Yuan Liu, Zhaotai Ma, Yiqiang Huang, and Peng Liu. A DQN-based multiobjective participant selection for efficient federated learning. Future Internet, 15(6):209–??, June 2023. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
6/209.

Xiao:2024:AAM

[XLW^{+}24] Shaoping Xiao, Junchao Li, Zhaoan Wang, Yingbin Chen, and Soheyla Tofighi. Advancing additive manufacturing through machine learning techniques: a state-of-the-art review. Future Internet, 16(11): 419, November 13, 2024. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
11/419.

Xu:2021:EPE

[XNC21] Ronghua Xu, Deeraj Nagothu, and Yu Chen. EconLedger: a proofof-ENF consensus based lightweight distributed ledger for IoVT networks. Future Internet, 13(10): 248, September 24, 2021. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/13/
                                        10/248.

Xu:2023:MDS

[XNZ23a] Yong Xu, Hong Ni, and Xiaoyong Zhu. A multipath data-scheduling strategy based on path correlation for informationcentric networking. Future Internet, 15(4):148, April 11, 2023. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/15/4/148.

Xu:2023:NMT

[XNZ23b] Yong Xu, Hong Ni, and Xiaoyong Zhu. A novel multipath transmission scheme for informationcentric networking. Future Internet, 15(2):80, February 17, 2023. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/15/2/80.

Xu:2019:DTS

[XQLZ19] Ling Xu, Jianzhong Qiao, Shukuan Lin, and Wanting Zhang. Dynamic task scheduling algorithm with deadline constraint in heterogeneous volunteer computing platforms. Future Internet, 11(6):121, May 28, 2019. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/11/6/121.

Xu:2022:DBD

[XSM22] Xingjian Xu, Lijun Sun, and Fanjun Meng. Distributed big data storage infrastructure for biomedical research featuring high-performance and rich-features. Future Internet, 14(10):273, September 24, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
10/273.

Xu:2024:MLS

[XSSA24] Hanyue Xu, Kah Phooi Seng, Jeremy Smith, and Li Minn Ang. Multi-level split federated learning for large-scale AIoT system based on smart cities. Future Internet, 16(3):82, February 28, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
3/82.

Xu:2018:EAS

[XSX18] Yi-Han Xu, Qiu-Ya Sun, and Yu-Tong Xiao. An environmentally aware scheme of wireless sensor networks for forest fire monitoring and detection. Future Internet, 10(10): 102, October 19, 2018. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/10/
10/102.

Xu:2018:QRU

[XY18] Jialu Xu and Feiyue Ye. Query recommendation using hybrid query relevance. Future Internet, 10(11):112, November 19, 2018. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/10/11/112.

Xu:2021:AWM

[XY21] Yao Xu and Qin Yu. Adaptive weighted multi-level fusion of multi-scale features: a new approach to pedestrian detection. Future Internet, 13(2):38, February 02, 2021. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
2/38.

Xu:2019:ETS

[XZ19] Da Xu and Tao Zhang. Efficient tensor sensing for RF tomographic imaging on GPUs. Future Internet, 11(2):46, February 15, 2019. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/11/2/46.

Xing:2024:SME

[XZWL24] Weicheng Xing, Tianqing Zhu, Jenny Wang, and Bo Liu. A survey on MLLMs in education: Application and future directions. Future Internet, 16(12):467, December 13, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/12/467.

Xie:2024:DGR

[XZY^{+}24] Mingxuan Xie, Tao Zou, Junchen Ye, Bowen Du, and Runhe Huang. Dynamic graph representation learning for passenger behavior prediction. Future Internet, 16(8):??, August 15, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
8/295.

Yang:2024:AFU

[YA24] Jeong Yang and Anoop Abraham. Analyzing the features, usability, and performance of deploying a containerized mobile Web application on serverless cloud platforms. Future Internet, 16(12):475, December 19, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
12/475.

Yang:2017:TND

[Yan17] Yao Yang. Towards a new digital era: Observing local e-government services adoption in a Chinese municipality. Future Internet, 9(3):53, September 20, 2017. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/9/3/53.

Yang:2019:RFA

[Yan19] Keng Yang. Research on factors affecting solvers’ participation time in online crowdsourcing contests. Future Internet, 11 (8):176, August 12, 2019. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/11/
                                        8/176.

Yimam:2021:IVS

[YAV^{+}21] Seid Muhie Yimam, Abinew Ali Ayele, Gopalakrishnan Venkatesh, Ibrahim Gashaw, and Chris Biemann. Introducing various semantic models for Amharic: Experimentation and evaluation with multiple tasks and datasets. Future Internet, 13(11):275, October 27, 2021. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/13/11/275.

Yakunin:2022:CIT

[YB22] Alexander V. Yakunin and Svetlana S. Bodrunova.

Cumulative impact of testing factors in usability tests for human-centered Web design. Future Internet, 14(12):359, November 30, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/12/359.

Yanagida:2024:MMD

[YB24] Ryo Yanagida and Saleem Noel Bhatti. Mobility–multihoming duality. Future Internet, 16(10), October 1, 2024. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
10/358.

Yuloskov:2021:SCR

[YBMK21] Artem Yuloskov, Mohammad Reza Bahrami, Manuel Mazzara, and Iouri Kotorov. Smart cities in Russia: Current situation and insights for future development. Future Internet, 13(10):252, September 28, 2021. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
10/252.

Yang:2023:DRL

[YBO^{+}23] Siyuan Yang, Mondher Bouazizi, Tomoaki Ohtsuki, Yohei Shibata, Wataru Takabatake, Kenji Hoshino, and Atsushi Nagate. Deep reinforcement learning evolution algorithm for dynamic antenna control in multi-cell configuration HAPS system. Future Internet, 15(1):34, January 12, 2023. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/15/
                                        1/34.

Yang:2016:NQP

[YC16] Weijun Yang and Yuanfeng Chen. A novel QoS provisioning algorithm for optimal multicast routing in WMNs. Future Internet, 8(3):38, August 01, 2016. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/8/
                                        3/38.

Yan:2022:CDC

[YC22] Hui Yan and Chaoyuan Cui. CacheHawkeye: Detecting cache side channel attacks based on memory events. Future Internet, 14 (1):24, January 08, 2022. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/14/
                                        1/24.

Yu:2023:MMS

[YCD^{+}23] Yong Yu, Shudong Chen, Rong Du, Da Tong, Hao Xu, and Shuai Chen. MSEN: a multi-scale evolutionary network for modeling the evolution of temporal knowledge graphs. Future Internet, 15(10):327, September 30, 2023. CODEN ????

ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/15/10/327.

Yang:2018:VBH

[YCHG18] Wanli Yang, Yimin Chen, Chen Huang, and Mingke Gao. Video-based human action recognition using spatial pyramid pooling and 3D densely convolutional networks. Future Internet, 10(12):115, November 22, 2018. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/10/
12/115.

Yan:2017:HPE

[YCS17] Chengxin Yan, Ningjiang Chen, and Zhang Shuo. High-performance elastic management for cloud containers based on predictive message scheduling. Future Internet, 9(4):87, November 28, 2017. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/9/
4/87.

Yamamoto:2024:IWW

[YFY24] Yudai Yamamoto, Aoi Fukushima, and Shingo Yamaguchi. Implementation of white-hat worms using Mirai source code and its optimization through parameter tuning. Future Internet, 16(9), September 13, 2024. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/16/9/336.

Yankson:2020:TET

[YG20] Samuel Yankson and Mahdi Ghamkhari. Transactive energy to thwart load altering attacks on power distribution systems. Future Internet, 12 (1):4, December 24, 2020. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/12/
                                        1/4.

Yang:2022:CTC

[YGQ^{+}22] Hua Yang, Teresa Gonçalves, Paulo Quaresma, Renata Vieira, Rute Veladas, Cátia Sousa Pinto, João Oliveira, Maria Cortes Ferreira, Jéssica Morais, Ana Raquel Pereira, Nuno Fernandes, and Carolina Gonçalves. Clinical trial classification of SNS24 calls with neural networks. Future Internet, 14(5):130, April 26, 2022. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/14/
                                        5/130.

Yang:2021:AAB

[YHG^{+}21] Zhi-Peng Yang, Lu Hua, Ning-Jie Gao, Ru Huo, Jiang Liu, and Tao Huang. An accelerating approach for blockchain information transmission based on NDN. Future Internet, 13 (2):47, February 14, 2021. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
2/47.

Yousif:2021:ITD

[YHN21] Mohamed Yousif, Chaminda Hewage, and Liqaa Nawaf. IoT technologies during and beyond COVID-19: a comprehensive review. Future Internet, 13(5):105, April 23, 2021. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
5/105.

Yao:2019:RSA

[YHSW19] Jiaying Yao, Zhigeng Han, Muhammad Sohail, and Liangmin Wang. A robust security architecture for SDN-based 5G networks. Future Internet, 11(4):85, March 28, 2019. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/11/4/85.

Yiu:2021:DSC

[Yiu21a] Neo C. K. Yiu. Decentralizing supply chain anti-counterfeiting and traceability systems using blockchain technology. Future Internet, 13(4):84, March 25, 2021. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/13/4/84.

Yiu:2021:TBE

[Yiu21b] Neo C. K. Yiu. Toward blockchain-enabled supply chain anti-counterfeiting and traceability. Future Internet, 13(4):86, March 29, 2021. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/13/4/86.

Yang:2018:CPG

[YJSL18] Hongbin Yang, Shuxiong Jiang, Wenfeng Shen, and Zhou Lei. Certificateless provable group shared data possession with comprehensive privacy preservation for cloud storage. Future Internet, 10(6):49, June 07, 2018. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/10/6/49.

Yoon:2021:RPP

[YK21] Youngjoon Yoon and Hyogon Kim. Resolving persistent packet collisions through broadcast feedback in cellular V2X communication. Future Internet, 13(8):211, August 16, 2021. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/13/8/211.

Yang:2019:SEO

[YKC19] Jin Sol Yang, Myung-Sook Ko, and Kwang Sik Chung. Social emotional opinion decision with newly coined words and emoticon polarity of social networks services. Future Internet, 11(8): 165, July 25, 2019. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
8/165.

Yu:2018:DBP

[YKY18] Lingli Yu, Decheng Kong, and Xiaoxin Yan. A driving behavior planning and trajectory generation method for autonomous electric bus. Future Internet, 10(6):51, June 10, 2018. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/10/6/51.

Yolchuyev:2021:DCP

[YL21] Agil Yolchuyev and Janos Levendovszky. Data chunks placement optimization for hybrid storage systems. Future Internet, 13(7):181, July 11, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/7/181.

Yu:2017:EEM

[YLH^{+}17] Qingyao Yu, Guangming Li, Xiaojie Hang, Kun Fu, and Tianqi Li. An energy efficient MAC protocol for wireless passive sensor networks. Future Internet, 9 (2):14, April 19, 2017. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/9/
                                        2/14.

Yuan:2019:MRC

[YLJ^{+}19] Baoxi Yuan, Yang Li, Fan Jiang, Xiaojie Xu, Yingxia Guo, Jianhua Zhao, Deyue Zhang, Jianxin Guo, and Xiaoli Shen. MU R-CNN: a two-dimensional code instance segmentation network based on deep learning. Future Internet, 11(9): 197, September 13, 2019. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/11/
                                        9/197.

Yuan:2019:LBD

[YLK^{+}19] Jie Yuan, Erxia Li, Chaoqun Kang, Fangyuan Chang, Tingting Yuan, and Xiaoyong Li. Latencybased dynamic controller assignment in hybrid SDNs: Considering the impact of legacy routers. Future Internet, 11(8): 168, July 28, 2019. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/11/
                                        8/168.

Yang:2021:CFB

[YLL21] FengLei Yang, Fei Liu, and ShanShan Liu. Collaborative filtering based on a variational Gaussian mixture model. Future Internet, 13(2):37, February 01, 2021. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/13/2/37.

Yu:2021:CMA

[YLPW21] Jie Yu, Yaliu Li, Chenle Pan, and Junwei Wang. A classification method for academic resources based on a graph attention network. Future Internet, 13 (3):64, March 04, 2021. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
3/64.

Yu:2024:HAC

[YLX^{+}24] Rui Yu, Sooyeon Lee, Jingyi Xie, Syed Masum Billah, and John M. Carroll. Human–AI collaboration for remote sighted assistance: Perspectives from the LLM era. Future Internet, 16(7):??, July 18, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/7/254.

Yang:2022:TPD

[YLY^{+}22] Cheng Yang, Tianliang Lu, Shangyi Yan, Jianling Zhang, and Xingzhan Yu. N-trans: Parallel detection algorithm for DGA domain names. Future Internet, 14(7):209, July 13, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/7/209.

Yang:2022:MAN

[YLYL22] Sa Yang, Suoping Li, Nana Yang, and Ying Lin. Modelling analysis of a novel frameless slotted-ALOHA protocol based on the number of detectable conflicting users. Future Internet, 14(10):279, September 28, 2022. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/14/10/279.

Yaser:2022:IDD

[YMH22] Ahmed Latif Yaser, Hamdy M. Mousa, and Mahmoud Hussein. Improved DDoS detection utilizing deep neural networks and feedforward neural networks as autoencoder. Future Internet, 14(8):240, August 12, 2022. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/14/8/240.

Yu:2020:BOT

[YMJ20] Jingyuan Yu and Juan Muñoz-Justicia. A bibliometric overview of Twitter-related studies indexed in Web of Science. Future Internet, 12(5):91, May 20, 2020. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/12/5/91.

Ye:2024:HSA

[YMSP24] Anjia Ye, Ananda Maiti, Matthew Schmidt, and Scott J. Pedersen. A hybrid semi-automated workflow for systematic and literature review processes with large language model analysis. Future Internet, 16(5):167, May 12, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/5/167.

Yeboah-Ofori:2019:CST

[YOI19] Abel Yeboah-Ofori and Shareeful Islam. Cyber security threat modeling for supply chain organizational environments. Future Internet, 11(3):63, March 05, 2019. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/11/3/63.

Younes:2023:ADS

[You23] Maram Bani Younes. Assisting drivers at stop signs in a connected vehicle environment. Future Internet, 15(7):238– ??, July 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
7/238.

Yang:2020:CEE

[YPG20] Jie Yang, Ziyu Pan, and Lihong Guo. Coverage and energy efficiency analysis for two-tier heterogeneous cellular networks based on Matérn hardcore process. Future Internet, 12(1):1, December 19, 2020. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/12/1/1.

Yuan:2019:FOS

[YPLZ19] Peiyan Yuan, Xiaoxiao Pang, Ping Liu, and En Zhang. FollowMe: One social importance-based collaborative scheme in MONs. Future Internet, 11(4):98, April 17, 2019. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/11/
                                        4/98.

Yerima:2012:AMB

[YPMM12] Suleiman Y. Yerima, Gerard P. Parr, Sally I. Mc-Clean, and Philip J. Morrow. Adaptive measurementbased policy-driven QoS management with fuzzyrule-based resource allocation. Future Internet, 4 (3):646–671, July 04, 2012. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/4/
                                        3/646.

Yang:2019:JOP

[YPXH19] Jie Yang, Ziyu Pan, Hengfei Xu, and Han Hu. Joint optimization of picobase-station density and transmit power for an energy-efficient heterogeneous cellular network. Future Internet, 11(10):208, September 27, 2019. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
10/208.

Yuan:2018:ICP

[YPZ18] Peiyan Yuan, Xiaoxiao Pang, and Xiaoyan Zhao. Influence of crowd participation features on mobile edge computing. Future Internet, 10(10):94, September 25, 2018. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/10/
10/94.

Yao:2024:DIO

[YQ24] Yu Yao and Quan Qian.

Dynamic industrial optimization: a framework integrates online machine learning for processing parameters design. Future Internet, 16(3):94, March 10, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/3/94.

Yu:2023:LAS

[YSAY23] Liangkun Yu, Xiang Sun, Rana Albelaihi, and Chen Yi. Latency-aware semisynchronous client selection and model aggregation for wireless federated learning. Future Internet, 15(11):352, October 26, 2023. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/15/11/352.

Yovita:2020:PAC

[YSEK20] Leanna Vidya Yovita, Nana Rachmana Syambas, Ian Joseph Matheus Edward, and Noriaki Kamiyama. Performance analysis of cache based on popularity and class in named data network. Future Internet, 12(12):227, December 09, 2020. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/12/
12/227.

Younis:2024:ECN

[YST24] Samir A. Younis, Dalia Sobhy, and Noha S. Tawfik. Evaluating convolutional neural networks and vision transformers for baby cry sound analysis. Future Internet, 16 (7):??, July 7, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
7/242.

Yamashita:2015:RTH

[YT15] Ryoma Yamashita and Kazumasa Takami. Receivertriggered handshake protocol for DTN in disaster area. Future Internet, 7 (2):152–169, May 27, 2015. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/7/
2/152.

Yashima:2018:RAC

[YT18] Tamotsu Yashima and Kazumasa Takami. Route availability as a communication quality metric of a mobile ad hoc network. Future Internet, 10(5):41, May 04, 2018. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/10/5/41.

Yajima:2019:YPU

[YT19] Hayato Yajima and Kazumasa Takami. A yielding protocol that uses intervehicle communication to improve the traffic of vehicles on a low-priority road at an unsignalized intersection. Future Internet, 11(5):110, May 09, 2019. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
5/110.

Ye:2016:DSA

[YW16] Ting Ye and Baowei Wang.

Density self-adaptive hybrid clustering routing protocol for wireless sensor networks. Future Internet, 8(3):27, June 29, 2016. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/8/
3/27.

Yang:2022:IBC

[YWLY22] Cunwei Yang, Weiqing Wang, Fengying Li, and Degang Yang. An IoTbased COVID-19 prevention and control system for enclosed spaces. Future Internet, 14(2):40, January 26, 2022. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/14/2/40.

Yan:2022:MSA

[YWS22] Shangyi Yan, Jingya Wang, and Zhiqiang Song. Microblog sentiment analysis based on dynamic character-level and wordlevel features and multihead self-attention pooling. Future Internet, 14(8): 234, July 29, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
8/234.

Yuan:2019:FLD

[YWW^{+}19] Quanbo Yuan, Huijuan Wang, Botao Wu, Yaodong Song, and Hejia Wang. A fusion load disaggregation method based on clustering algorithm and support vector regression optimization for low sampling data. Future Internet, 11(2):51, February 19, 2019. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
2/51.

Yuan:2023:SAC

[YWW^{+}23] Ye Yuan, Wang Wang, Guangze Wen, Zikun Zheng, and Zhemin Zhuang. Sentiment analysis of Chinese product reviews based on fusion of DUALchannel BiLSTM and selfattention. Future Internet, 15(11):364, November 10, 2023. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/15/11/364.

Yuan:2024:ELT

[YWZ^{+}24] Xinyang Yuan, Jiarong Wang, Haozhi Zhao, Tian Yan, and Fazhi Qi. Empowering LLMs with toolkits: an open-source intelligence acquisition method. Future Internet, 16(12):461, December 7, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/12/461.

Yu:2021:JOE

[YXL^{+}21] Zhiyan Yu, Gaochao Xu, Yang Li, Peng Liu, and Long Li. Joint offloading and energy harvesting design in multiple time blocks for FDMA based wireless powered MEC. Future Internet, 13(3):70, March 12, 2021. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/13/3/70.

Yuan:2017:CBM

[YY17] Peiyan Yuan and Hai Yu.

A combinational buffer management scheme in mobile opportunistic network. Future Internet, 9 (4):82, November 14, 2017. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/9/
                                        4/82.

Yang:2020:MBS

[YYW20] Wenzhong Yang, Tingting Yuan, and Liejun Wang. Micro-blog sentiment classification method based on the personality and bagging algorithm. Future Internet, 12(4):75, April 20, 2020. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/12/4/75.

Yang:2023:HIH

[YYY^{+}23] Xingyuan Yang, Jie Yuan, Hao Yang, Ya Kong, Hao Zhang, and Jinyu Zhao. A highly interactive honeypot-based approach to network threat management. Future Internet, 15 (4):127, March 28, 2023. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/15/
                                        4/127.

Yin:2023:BTT

[YZ23] Zhenzhong Yin and Bin Zhang. Bus travel time prediction based on the similarity in drivers’ driving styles. Future Internet, 15(7):222–??, July 2023.

CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
7/222.

Yu:2020:SSV

[YZC^{+}20] Miao Yu, Jianwei Zhuge, Ming Cao, Zhiwei Shi, and Lin Jiang. A survey of security vulnerability analysis, discovery, detection, and mitigation on IoT devices. Future Internet, 12(2):27, February 06, 2020. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/12/2/27.

Yang:2024:PCA

[YZCR24] Jindong Yang, Xiran Zhang, Wenhao Chen, and Fei Rong. Prophet– CEEMDAN–ARBiLSTMbased model for shortterm load forecasting. Future Internet, 16(6):192, May 31, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
6/192.

Yang:2022:FIB

[YZP22] Jie Yang, Jiajia Zhu, and Ziyu Pan. A fairness index based on rate variance for downlink non-orthogonal multiple access system. Future Internet, 14(9):261, August 31, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.

                                        mdpi.com/1999-5903/14/
                                        9/261.

Yang:2020:MCB

[YZZ^{+}20] Pin Yang, Huiyu Zhou, Yue Zhu, Liang Liu, and Lei Zhang. Malware classification based on shallow neural network. Future Internet, 12(12):219, December 02, 2020. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/12/
                                        12/219.

Zafiropoulos:2012:CPA

[Zaf12] Kostas Zafiropoulos. Connectivity practices and activity of Greek political blogs. Future Internet, 4(3):719–736, August 14, 2012. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/4/3/719.

Zargayouna:2022:UMA

[Zar22] Mahdi Zargayouna. On the use of the multi-agent environment for mobility applications. Future Internet, 14(5):132, April 27, 2022. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/14/5/132.

Zarie:2024:MBV

[ZAS^{+}24] Mira M. Zarie, Abdelhamied A. Ateya, Mohammed S. Sayed, Mohammed ElAffendi, and Mohammad Mahmoud Abdellatif. Microservicebased vehicular network for seamless and ultrareliable communications of connected vehicles. Future Internet, 16(7):??, July 19, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/7/257.

Zafiropoulos:2014:SFG

[ZAV14] Kostas Zafiropoulos, Konstantinos Antoniadis, and Vasiliki Vrana. Sharing followers in e-government Twitter accounts: The case of Greece. Future Internet, 6(2):337–358, May 14, 2014. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/6/2/337.

Zaway:2024:OWC

[ZBK^{+}24] Lassaad Zaway, Nader Ben Amor, Jalel Ktari, Mohamed Jallouli, Larbi Chrifi Alaoui, and Laurent Delahoche. Optimization of wheelchair control via multi-modal integration: Combining webcam and EEG. Future Internet, 16 (5):158, May 3, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
5/158.

Zannou:2021:SNS

[ZBN21] Abderrahim Zannou, Abdelhak Boulaalam, and El Habib Nfaoui. SIoT: a new strategy to improve the network lifetime with an efficient search process. Future Internet, 13(1):4, December 29, 2021. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/13/
                                        1/4.

Zhang:2021:DRP

[ZC21] Xiaolin Zhang and Chao Che. Drug repurposing for Parkinson’s disease by integrating knowledge graph completion model and knowledge fusion of medical literature. Future Internet, 13(1):14, January 08, 2021. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/13/1/14.

Zilli:2023:INW

[ZCdRR^{+}23] Vitória Francesca Biasibetti Zilli, Cesar David Paredes Crovato, Rodrigo da Rosa Righi, Rodrigo Ivan Goytia Mejia, Giovani Pesenti, and Dhananjay Singh. I4.0I: a new way to rank how involved a company is in the Industry 4.0 era. Future Internet, 15 (2):73, February 13, 2023. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/15/
                                        2/73.

Zhang:2024:EDI

[ZCG^{+}24] Jingzhe Zhang, Tiange Chen, Wenjie Gong, Jiayue Liu, and Jiangjie Chen. Exploring data input problems in mixed reality environments: Proposal and evaluation of natural interaction techniques. Future Internet, 16(5):150, April 27, 2024. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
5/150.

Zhang:2019:ODN

[ZCHG19] Ying Zhang, Yimin Chen, Chen Huang, and Mingke Gao. Object detection network based on feature fusion and attention mechanism. Future Internet, 11 (1):9, January 02, 2019. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
1/9.

Zhang:2017:MCU

[ZCZ17] Jingbo Zhang, Lili Cai, and Shufang Zhang. Malicious cognitive user identification algorithm in centralized spectrum sensing system. Future Internet, 9 (4):79, November 08, 2017. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/9/
4/79.

Zheng:2022:DRB

[ZDP^{+}22] Jason Zheng, Chidinma Dike, Stefan Pancari, Yi Wang, George C.

Giakos, Wafa Elmannai, and Bingyang Wei. An in-depth review on blockchain simulators for IoT environments. Future Internet, 14(6):182, June 10, 2022. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/14/6/182.

Ziaimatin:2013:STD

[ZGH13] Hasti Ziaimatin, Tudor Groza, and Jane Hunter. Semantic and time-dependent expertise profiling models in community-driven knowledge curation platforms. Future Internet, 5(4):490– 514, October 11, 2013. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/5/
                                        4/490.

Zhao:2020:ITC

[ZGL20] Yafei Zhao, Paolo Vincenzo Genovese, and Zhixing Li. Intelligent thermal comfort controlling system for buildings based on IoT and AI. Future Internet, 12(2):30, February 10, 2020. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/12/2/30.

Zhu:2018:BDL

[ZGZ^{+}18] Yonghua Zhu, Xun Gao, Weilin Zhang, Shenkai Liu, and Yuanyuan Zhang.

A bi-directional LSTM-CNN model with attention for aspect-level text classification. Future Internet, 10(12):116, November 24, 2018. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/10/12/116.

Zhao:2021:DPP

[ZH21] Yuexuan Zhao and Jing Huang. Dirichlet process prior for Student’s t graph variational autoencoders. Future Internet, 13(3):75, March 16, 2021. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/13/3/75.

Zhao:2019:IQE

[Zha19] Haichuan Zhao. Information quality or entities’ interactivity? Understanding the determinants of social network-based brand community participation. Future Internet, 11(4):87, April 01, 2019. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/11/4/87.

Zhao:2024:RMM

[ZHH24] Rui Zhao, Jutao Hao, and Huan Huo. Research on multi-modal pedestrian detection and tracking algorithm based on deep learning. Future Internet, 16(6):194, May 31, 2024. CODEN ???? ISSN 1999-

5903. URL https://www.

                                        mdpi.com/1999-5903/16/
                                        6/194.

Zhao:2023:TSB

[ZHW23] Ke Zhao, Rui Han, and Xu Wang. Time segmentation-based hybrid caching in 5G-ICN bearer network. Future Internet, 15(1):30, January 07, 2023. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/15/1/30.

Zardari:2019:DAD

[ZHZ^{+}19] Zulfiqar Ali Zardari, Jingsha He, Nafei Zhu, Khalid Hussain Mohammadani, Muhammad Salman Pathan, Muhammad Iftikhar Hussain, and Muhammad Qasim Memon. A dual attack detection technique to identify black and Gray hole attacks using an intrusion detection system and a connected dominating set in MANETs. Future Internet, 11(3):61, March 05, 2019. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/11/3/61.

Zhai:2018:IEM

[ZJ18] Linbo Zhai and Wenwen Jiang. Intelligent environment monitoring system for university laboratories. Future Internet, 10(11): 110, November 16, 2018. CODEN ???? ISSN 1999-

                                        5903. URL https://www.

mdpi.com/1999-5903/10/
11/110.

Zhu:2023:TAM

[ZJJS^{+}23] Jinting Zhu, Julian Jang-Jaccard, Amardeep Singh, Paul A. Watters, and Seyit Camtepe. Taskaware meta learning-based Siamese neural network for classifying control flow obfuscated malware. Future Internet, 15(6):214– ??, June 2023. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
6/214.

Zhou:2018:SST

[ZJX18] Liang Zhou, Sheng-Ming Jiang, and Chen-Lin Xiong. Studying semi-TCP and its application in marine Internet. Future Internet, 10(6):44, May 25, 2018. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/10/6/44.

Zantalis:2019:RML

[ZKKK19] Fotios Zantalis, Grigorios Koulouras, Sotiris Karabetsos, and Dionisis Kandris. A review of machine learning and IoT in smart transportation. Future Internet, 11(4):94, April 10, 2019. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/11/4/94.

Zafiropoulos:2012:AAG

[ZKV12] Kostas Zafiropoulos, Ioannis Karavasilis, and Vasiliki Vrana. Assessing the adoption of e-government services by teachers in Greece. Future Internet, 4 (2):528–544, May 21, 2012. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/4/
                                        2/528.

Zoppi:2014:OAA

[ZL14] Corrado Zoppi and Sabrina Lai. An ontology of the appropriate assessment of municipal master plans related to Sardinia (Italy). Future Internet, 6(2):223–241, April 23, 2014. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/6/2/223.

Zhang:2019:MAN

[ZL19a] Qiuyue Zhang and Ran Lu. A multi-attention network for aspect-level sentiment analysis. Future Internet, 11(7):157, July 16, 2019. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/11/7/157.

Zhang:2019:DGR

[ZL19b] Xinyu Zhang and Xiaoqiang Li. Dynamic gesture recognition based on MEMP network. Future Internet, 11(4):91, April 03, 2019. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/11/4/91.

Zhang:2020:IDB

[ZL20] Yue Zhang and Fangai Liu. An improved deep belief network prediction model based on knowledge transfer. Future Internet, 12(11):188, October 29, 2020. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/12/11/188.

Zhao:2021:TPE

[ZL21] Xujian Zhao and Wei Li. Trend prediction of event popularity from microblogs. Future Internet, 13(9):220, August 24, 2021. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/13/9/220.

Zhang:2022:TVD

[ZL22] Yujian Zhang and Daifu Liu. Toward vulnerability detection for Ethereum smart contracts using graph-matching network. Future Internet, 14(11): 326, November 11, 2022. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
11/326.

Zhang:2018:MIP

[ZLL18] Lijun Zhang, Kai Liu, and Jian Liu. Multidiscipline integrated platform based on probabilistic analysis for manufacturing engineering processes. Future Internet, 10(8):70, July 30, 2018. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/10/8/70.

Zhu:2021:IPR

[ZLL^{+}21] Qigang Zhu, Yifan Liu, Ming Liu, Shuaishuai Zhang, Guangyang Chen, and Hao Meng. Intelligent planning and research on urban traffic congestion. Future Internet, 13(11): 284, November 08, 2021. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
11/284.

Zhou:2018:AOR

[ZLLW18] Zhibin Zhou, Pin Liu, Qin Liu, and Guojun Wang. An anonymous offline RFID grouping-proof protocol. Future Internet, 10(1):2, January 01, 2018. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/10/
1/2.

Zhou:2024:CSU

[ZLOW24] Mengcai Zhou, Minglun Li, Kenta Ono, and Makoto Watanabe. A comparative study of the user interaction behavior and experience in a homeoriented multi-user interface (MUI) during family collaborative cooking. Future Internet, 16(12):478, December 20, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
12/478.

Zhao:2019:NAB

[ZLS19] Ning Zhao, Yuhe Liu, and Junjie Shen. Nonlinear analysis of built-in sensor in smart device under the condition of voice actuating. Future Internet, 11(3): 81, March 26, 2019. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
3/81.

Zhang:2022:LCG

[ZLT22] Xinhe Zhang, Wenbo Lv, and Haoran Tan. Lowcomplexity GSM detection based on maximum ratio combining. Future Internet, 14(5):159, May 23, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/5/159.

Zhao:2021:APA

[ZLW21] Yulin Zhao, Junke Li, and Jiang-E Wang. Analysis and prediction of “AI + Education” attention based on Baidu index — taking Guizhou Province as an example. Future Internet, 13(5):120, April 30, 2021. CODEN ????

ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/13/5/120.

Zhang:2019:FFT

[ZLXY19] Jingren Zhang, Fang’ai Liu, Weizhi Xu, and Hui Yu. Feature fusion text classification model combining CNN and Bi-GRU with multi-attention mechanism. Future Internet, 11(11):237, November 12, 2019. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/11/11/237.

Zhang:2020:KMC

[ZLY^{+}20] Wenbo Zhang, Xiao Li, Yating Yang, Rui Dong, and Gongxu Luo. Keeping models consistent between pretraining and translation for low-resource neural machine translation. Future Internet, 12(12): 215, November 27, 2020. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/12/
                                        12/215.

Zhang:2024:MCM

[ZLZ^{+}24] Bing Zhang, Hui Li, Shuai Zhang, Jing Sun, Ning Wei, Wenhong Xu, and Huan Wang. Multiconstraint and multipolicy path hopping active defense method based on SDN. Future Internet, 16(4):143, April 22, 2024.

CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
4/143.

Zhou:2023:PEP

[ZLZZ23] Qingyan Zhou, Hao Li, Youhua Zhang, and Junhong Zheng. Product evaluation prediction model based on multi-level deep feature fusion. Future Internet, 15(1):31, January 09, 2023. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/15/1/31.

Zali:2022:PPS

[ZMDCEMI22] Siti-Aisyah Zali, Shahbe Mat-Desa, Zarina Che-Embi, and Wan-Noorshahida Mohd-Isa. Post-processing for shadow detection in drone-acquired images using U-NET. Future Internet, 14(8):231, July 28, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/8/231.

ZareRavasan:2022:ESI

[ZMKR22] Ahad ZareRavasan, Taha Mansouri, Michal Krcál, and Saeed Rouhani. Editorial for the special issue on blockchain: Applications, challenges, and solutions. Future Internet, 14(5):155, May 19, 2022. CODEN ???? ISSN 1999-

5903. URL https://www.

mdpi.com/1999-5903/14/
5/155.

Zhang:2022:SLP

[ZMZ^{+}22] Weiwei Zhang, Xin Ma, Yuzhao Zhang, Ming Ji, and Chenghui Zhen. SMY-OLO: Lightweight pedestrian target detection algorithm in low-altitude scenarios. Future Internet, 14 (1):21, January 04, 2022. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
1/21.

Zhang:2022:EDM

[ZN22] Yiyi Zhang and Tatsuo Nakajima. Exploring the design of a mixed-reality 3D minimap to enhance pedestrian satisfaction in urban exploratory navigation. Future Internet, 14(11):325, November 10, 2022. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/14/11/325.

Zhang:2024:AKA

[ZOAR24] Jiayi Zhang, Abdelkader Ouda, and Raafat Abu-Rukba. Authentication and key agreement protocol in hybrid edge–fog– cloud computing enhanced by 5G networks. Future Internet, 16(6):209, June 14, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/6/209.

Zhang:2024:PPF

[ZPX24] Yu Zhang, Xiaowei Peng, and Hequn Xian. pFed-BASC: Personalized federated learning with blockchainassisted semi-centralized framework. Future Internet, 16(5):164, May 11, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/5/164.

Zhu:2016:NMF

[ZQCC16] Zhiqin Zhu, Guanqiu Qi, Yi Chai, and Yinong Chen. A novel multifocus image fusion method based on stochastic coordinate coding and local density peaks clustering. Future Internet, 8(4):53, November 11, 2016. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/8/
4/53.

Zu:2024:SMS

[ZQL^{+}24] Lijun Zu, Wenyu Qi, Hongyi Li, Xiaohua Men, Zhihui Lu, Jiawei Ye, and Liang Zhang. UP-SDCG: a method of sensitive data classification for collaborative edge computing in financial cloud environment. Future Internet, 16 (3):102, March 18, 2024. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
3/102.

Zaidi:2015:VIS

[ZR15] Kamran Zaidi and Muttukrishnan Rajarajan. Vehicular Internet: Security & privacy challenges and opportunities. Future Internet, 7(3):257–275, July 24, 2015. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/7/3/257.

Zhang:2018:ECP

[ZS18] Yu Zhang and Bingjia Shao. The effect of customer participation types on online recovery satisfaction: a mental accounting perspective. Future Internet, 10(10):97, October 03, 2018. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/10/10/97.

Zeadally:2019:YBT

[ZSB19] Sherali Zeadally, Farhan Siddiqui, and Zubair Baig. 25 years of Bluetooth technology. Future Internet, 11(9):194, September 09, 2019. CODEN ???? ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/11/9/194.

Zhang:2024:GAM

[ZSB24] Peng Zhang, Jiayu Shi, and Maged N. Kamel Boulos. Generative AI in medicine and healthcare: Moving beyond the ‘peak of inflated expectations’.

Future Internet, 16(12): 462, December 9, 2024. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
12/462.

Zia:2020:IRA

[ZSF20] Kashif Zia, Muhammad Shafi, and Umar Farooq. Improving recommendation accuracy using social network of owners in social Internet of Vehicles. Future Internet, 12(4):69, April 16, 2020. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/12/4/69.

Zhao:2023:CPP

[ZSRC23] Jiaxu Zhao, Binting Su, Xuli Rao, and Zhide Chen. A cross-platform personalized recommender system for connecting e-commerce and social network. Future Internet, 15(1):13, December 27, 2023. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/15/1/13.

Zubani:2022:PCD

[ZSS^{+}22] Matteo Zubani, Luca Sigalini, Ivan Serina, Luca Putelli, Alfonso E. Gerevini, and Mattia Chiari. A performance comparison of different cloud-based natural language understanding services for an Italian elearning platform. Future Internet, 14(2):62, February 18, 2022. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/14/2/62.

Zhang:2018:IAS

[ZSZ18] Jingbo Zhang, Zhenyang Sun, and Shufang Zhang. The improved adaptive silence period algorithm over time-variant channels in the cognitive radio system. Future Internet, 10 (2):12, January 29, 2018. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/10/
                                        2/12.

Ziouzios:2020:DAS

[ZTBD20] Dimitris Ziouzios, Dimitris Tsiktsiris, Nikolaos Baras, and Minas Dasygenis. A distributed architecture for smart recycling using machine learning. Future Internet, 12 (9):141, August 24, 2020. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/12/
                                        9/141.

Ziakis:2019:IFI

[ZVKK19] Christos Ziakis, Maro Vlachopoulou, Theodosios Kyrkoudis, and Makrina Karagkiozidou. Important factors for improving Google search rank. Future Internet, 11(2):32, January 30, 2019. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/11/
2/32.

Zafiropoulos:2012:BCC

[ZVV12] Kostas Zafiropoulos, Vasiliki Vrana, and Dimitrios Vagianos. Bloggers’ community characteristics and influence within Greek political blogosphere. Future Internet, 4(2):396– 412, April 19, 2012. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/4/
2/396.

Zhang:2024:PPO

[ZWL^{+}24] Chen Zhang, Celimuge Wu, Min Lin, Yangfei Lin, and William Liu. Proximal policy optimization for efficient D2D-Assisted computation offloading and resource allocation in multiaccess edge computing. Future Internet, 16(1):19, January 02, 2024. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/16/
1/19.

Zhang:2019:RMP

[ZWPL19] Wenjie Zhang, Pin Wu, Yan Peng, and Dongke Liu. Roll motion prediction of unmanned surface vehicle based on coupled CNN and LSTM. Future Internet, 11(11):243, November 18, 2019. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/11/
                                        11/243.

Zhang:2020:INP

[ZWS20] Ping Zhang, Rongqin Wang, and Nianfeng Shi. IgA nephropathy prediction in children with machine learning algorithms. Future Internet, 12(12): 230, December 17, 2020. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/12/
                                        12/230.

Zhang:2017:FBI

[ZWWW17] Yunpeng Zhang, Chengyou Wang, Xiaoli Wang, and Min Wang. Feature-based image watermarking algorithm using SVD and APBT for copyright protection. Future Internet, 9 (2):13, April 19, 2017. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/9/
                                        2/13.

Zhang:2017:RIW

[ZWZ17] Heng Zhang, Chengyou Wang, and Xiao Zhou. A robust image watermarking scheme based on SVD in the spatial domain. Future Internet, 9(3):45, August 07, 2017. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/9/3/45.

Zheng:2019:CPS

[ZWZ19] Xin Zheng, Gaocai Wang, and Qifei Zhao. A cache placement strategy with energy consumption optimization in informationcentric networking. Future Internet, 11(3):64, March 05, 2019. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/11/3/64.

Zhang:2024:RIS

[ZX24] Yuzhu Zhang and Hao Xu.

Reconfigurable-intelligent-surface-enhanced dynamic resource allocation for the social Internet of Electric Vehicle Charging Networks with causalstructure-based reinforcement learning. Future Internet, 16(5):165, May 11, 2024. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/16/5/165.

Zhao:2022:EST

[ZXJ22] Jie Zhao, Fangwei Xiong, and Peiquan Jin. Enhancing short-term sales prediction with microblogs: a case study of the movie box office. Future Internet, 14(5):141, May 04, 2022. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
5/141.

Zhu:2022:FER

[ZXMB22] Hongtao Zhu, Huahu Xu, Xiaojin Ma, and Minjie Bian. Facial expression recognition using dual path feature fusion and stacked attention. Future Internet, 14(9):258, August 30, 2022. CO-DEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/14/
                                        9/258.

Zhou:2021:DMP

[ZXWZ21] Xingchen Zhou, Ming Xu, Yiming Wu, and Ning Zheng. Deep model poisoning attack on federated learning. Future Internet, 13(3):73, March 14, 2021. CODEN ???? ISSN 1999-

                                        5903. URL https://www.
                                        mdpi.com/1999-5903/13/
                                        3/73.

Zhang:2022:JCN

[ZXXB22] Yaojie Zhang, Huahu Xu, Junsheng Xiao, and Minjie Bian. JoSDW: Combating noisy labels by dynamic weight. Future Internet, 14(2):50, February 02, 2022. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/14/2/50.

Zhou:2015:ORR

[ZY15] Dan Zhou and Zhong Yao. Optimal referral reward considering customer’s budget constraint.

Future Internet, 7(4):516– 529, December 21, 2015. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/7/
4/516.

Zhao:2020:UAI

[ZY20a] Jie Zhao and Can Yan.

User acceptance of information feed advertising: a hybrid method based on SEM and QCA. Future Internet, 12(12):209, November 26, 2020. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/12/
12/209.

Zhou:2020:LSM

[ZY20b] Yi Zhou and Fenglei Yang.

Latent structure matching for knowledge transfer in reinforcement learning. Future Internet, 12(2):36, February 13, 2020. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/12/
2/36.

Zhou:2017:LPP

[ZYLM17] Kaijun Zhou, Lingli Yu, Ziwei Long, and Siyao Mo. Local path planning of driverless car navigation based on jump point search method under urban environment. Future Internet, 9(3):51, September 12, 2017. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/9/3/51.

Zhao:2024:DID

[ZYWS24] Yadi Zhao, Lei Yan, Jian Wu, and Ximing Song. Design and implementation of a digital twin system for log rotary cutting optimization. Future Internet, 16(1):7, December 25, 2024. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/16/1/7.

Zhang:2018:FLM

[ZZ18] Lei Zhang and Xiaoli Zhi. A fast and lightweight method with feature fusion and multi-context for face detection. Future Internet, 10(8):80, August 17, 2018. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/10/8/80.

Zhao:2019:SAB

[ZZ19] Lingling Zhao and Anping Zhao. Sentiment analysis based requirement evolution prediction. Future Internet, 11(2):52, February 21, 2019. CODEN ???? ISSN 1999-5903.

                                        URL https://www.mdpi.
                                        com/1999-5903/11/2/52.

Zhou:2021:HPG

[ZZ21] Chao Zhou and Tao Zhang. High performance graph data imputation on multiple GPUs. Future Internet, 13(2):36, January 31, 2021. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/13/2/36.

Zhu:2023:HCW

[ZZL^{+}23] Yiming Zhu, Dehua Zhou, Yuan Li, Beibei Song, and Chuansheng Wang. How can we achieve query keyword frequency analysis in privacy-preserving situations? Future Internet, 15(6):197–??, June 2023. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/15/
6/197.

Zhao:2019:EDL

[ZZSX19] Hongwei Zhao, Weishan Zhang, Haoyun Sun, and Bing Xue. Embedded deep learning for ship detection and recognition. Future Internet, 11(2):53, February 21, 2019. CODEN ???? ISSN 1999-5903.

URL https://www.mdpi.
com/1999-5903/11/2/53.

Zhang:2019:MEE

[ZZWP19] Wenming Zhang, Yiwen Zhang, Qilin Wu, and Kai Peng. Mobility-enabled edge server selection for multi-user composite services. Future Internet, 11 (9):184, August 25, 2019. CODEN ???? ISSN 1999-

5903. URL https://www.

mdpi.com/1999-5903/11/
9/184.

Zou:2019:RCC

[ZZY19] Yongyan Zou, Yanzhi Zhang, and Xin Yi. Research on cooperative communication strategy and intelligent agent directional source grouping algorithms for Internet of Things. Future Internet, 11(11):233, November 01, 2019. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/11/11/233.

Zhao:2021:CPD

[ZZYC21] Yu Zhao, Yi Zhu, Qiao Yu, and Xiaoying Chen. Crossproject defect prediction method based on manifold feature transformation. Future Internet, 13 (8):216, August 20, 2021. CODEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/13/
8/216.

Zhu:2022:DIC

[ZZZ^{+}22] Yancong Zhu, Juan Zhang, Zhaoxi Zhang, Gina Clepper, Jingpeng Jia, and Wei Liu. Designing an interactive communication assistance system for hearingimpaired College students based on gesture recognition and representation. Future Internet, 14(7):198, June 29, 2022. CO-DEN ???? ISSN 1999-

5903. URL https://www.
mdpi.com/1999-5903/14/
7/198.

Zhang:2023:IUN

[ZZZ^{+}23] Weijie Zhang, Lanping Zhang, Xixi Zhang, Yu Wang, Pengfei Liu, and Guan Gui. Intelligent unsupervised network traffic classification method using adversarial training and deep clustering for secure Internet of Things. Future Internet, 15(9):298, September 01, 2023. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/15/9/298.

Zhai:2023:CRD

[ZZZL23a] Qiuhong Zhai, Wenhao Zhu, Xiaoyu Zhang, and Chenyun Liu. Contrastive refinement for dense retrieval inference in the open-domain question answering task. Future Internet, 15(4):137, March 31, 2023. CODEN ???? ISSN 1999-5903. URL

https://www.mdpi.com/
1999-5903/15/4/137.

Zhu:2023:HTG

[ZZZL23b] Wenhao Zhu, Xiaoyu Zhang, Qiuhong Zhai, and Chenyun Liu. A hybrid text generation-based query expansion method for open-domain question answering. Future Internet, 15(5):180, May 12, 2023. CODEN ????

ISSN 1999-5903. URL

                                        https://www.mdpi.com/
                                        1999-5903/15/5/180.